### v35

# Good News for Polynomial Root-finding

Abstract
Univariate polynomial root-finding has been studied for four millennia and very intensively in the last decades. Our novel near optimal root-finders approximate all zeros of a polynomial $p$ almost as fast as one accesses its coefficients with the precision required for the solution within a prescribed error bound.  [^1] Under Las Vegas randomization we allow output errors but detect them at a dominated computational cost. Furthermore, our root-finders can be applied to a black box polynomial, defined by an oracle (that is, black box subroutine) for its evaluation rather than by its coefficients. [^2] The pioneering root-finder by Louis and Vempala at FOCS 2016 is the only other known black box root-finder, but it only approximates the absolutely largest root of a real-rooted polynomial. Such root-finders are particularly fast for polynomials that can be evaluated fast such as the sum of a few shifted monomials of the form $(x-c)^{d}$. Our divide and conquer algorithm of ACM STOC 1995 is the only other known near optimal polynomial root-finder, but it extensively uses the coefficients, is quite involved, and has never been implemented, while according to extensive numerical experiments with standard test polynomials, already initial implementations of our new root-finders compete with user’s choice package of root-finding subroutine MPSolve and supersede it more and more significantly where the degree of a polynomial grows large. [^3] The root-finder of 1995 is deterministic, while our current ones apply Las Vegas randomization, that is, allow output errors but detect them at a dominated computational cost. Randomization enables us to avoid (with a high probability) division by near vanishing values in our computations with bounded precision. Our root-finders are readily extended to support approximation of the eigenvalues of a matrix within a record Las Vegas expected bit operation time bound. Our auxiliary algorithms and techniques for computations with black box polynomials can be of independent interest.
Keywords:
polynomial computations, matrix eigenvalues, complexity, polynomial zeros, black box polynomials, root-squaring, symbolic-numeric computing.
2020 Math. Subject Classification:
65H05, 26C10, 30C15
1 Introduction
1.1 Polynomial Root-finding: Problems, Impact, and Earlier Studies
According to [7, 95, 96, 80], univariate polynomial root-finding, that is, the solution of a polynomial equation $p(x)=0$, had been the central problem of Mathematics and Computational Mathematics for about four millennia since Sumerian times; it was still a central problem well into the 19th century. In its long way from its rudimentary level to its modern version, this study has been responsible for the introduction of mathematical rigor and notation as well as the concepts of irrational, negative, and complex numbers, algebraic groups, rings, and fields etc. Polynomial root-finding is still an important subject of modern computations, in particular for computer algebra, signal processing, geometric modeling, control, and financial mathematics [77, 79, 108].
The existence of complex roots of the equation $p(x)=0$ is called the Fundamental Theorem of Algebra. The attempts of its proof by d’Alembert (1746), Euler (1749), de Foncenex (1759), Lagrange (1772), Laplace (1795), and Gauss (1799) had gaps (cf. [129]), and the first rigorous proof is due to an amateurish mathematician Argand (1806, revised in 1813).
Already about 500 BC it was proved in Pythagorean school that the roots of the equation $x^{2}=2$ were irrational radicals, while even with the help of radicals one cannot express the roots rationally if the degree $d$ of $p(x)$ exceeds 4, according to Ruffini (1799,1813) and Abel (1824); moreover, one cannot do this already for specific equations such as $x^{5}-x-1=0$ or $x^{5}-4x-2$, according to Galois theory of 1830, 1832. Weyl in his constructive proof of Fundamental Theorem of Algebra in 1924 in [144], however, rationally approximated all $d$ zeros of $p$ within any fixed positive error tolerance. This progress, the advance of computer technology, and various applications to modern computations have led to the study of the following
Root-finding problems: Given the $d+1$ complex coefficients of a polynomial $p(x)$ of a degree $d$, approximate within a fixed error tolerance $\epsilon$ either (i) its every zero (aka every root of the equation $p(x)=0$) or (ii) its every zero lying in a fixed Region of Interest $\mathbb{S}$ [^4] We count the zeros of $p$ with multiplicities and do not distinguish multiple roots from root clusters.
Hundreds of efficient root-finders have been proposed [77, 79] and keep appearing (see [118, 21, 38, 67, 73, 123, 104, 22, 23, 61, 106, 107, 72, 57, 59, 78, 60, 56, 111], and the bibliography therein). Smale in [129] proposed to study Problem 1 in terms of arithmetic complexity and computational precision and later (jointly with Blum, Cucker, and Shub) (cf. [13, 28]) linked this study to rounding errors and conditioning of the roots. In [130] Schönhage proposed to estimate the complexity of polynomial root-finders in terms of bit operations involved and presented an algorithm that approximated all zeros of $p$ within $\epsilon=1/2^{b}$ by using $\tilde{O}(bd^{3})$ bit operations, that is, $bd^{3}$ bit operations up to a polylogarithmic factor in $bd$. Intensive study in the 1980s and 1990s has culminated in [92] (cf. also [93, 99] and surveys in [95, 96]) with our divide and conquer root-finder that decreased Schönhage’s record to $\tilde{O}((b+d)d^{2})$. This near optimal bound, that is, optimal up to polylogarithmic factors, implies that the root-finder runs nearly as fast as one accesses the coefficients with the precision required for the task.
If we only consider polynomials whose zeros and tiny clusters of zeros are pairwise sufficiently isolated (as we specify in Def. B.1 of Sec. B), then the known lower bounds on the computational precision and the bit operation complexity of root-finding decrease by a factor of $d$, and so do the upper bounds supported by the algorithms of [92, 98] as well; hence the algorithms remain near optimal.
1.2 Reduction of root-finding problems to the case of the unit disc
Hereafter ${}^{T}$ shows transposition of a vector or a matrix, $\mathbb{R}^{k}$ and $\mathbb{C}^{k}$ denote the spaces of the column vectors $(v_{1},\dots,v_{k})^{T}$ of length $k$ with real and complex coordinates, respectively, and $D(c,\rho):=\{x:~{}|x-c|\leq\rho\}$ and $C(c,\rho):=\{x:~{}|x-c|=\rho\}$ denote the disc of radius $\rho>0$ centered at $c\in\mathbb{C}$ and its boundary circle, respectively.
Next we link the above Root-finding problem (i) to its case (ii) where the Region of Interest $\mathbb{S}$ is the unit disc $D(0,1)$.
Problem 0: Root approximation in the unit disc.
INPUT: positive $\epsilon\in\mathbb{R}$, $d\in\mathbb{N}$, and $(p_{0},p_{1},\dots,p_{d})^{T}\in\mathbb{C}^{d}$, defining the polynomial
$p=p(x):=\sum_{i=0}^{d}p_{i}x^{i}:=p_{d}\prod_{j=1}^{d}(x-z_{j}),$ (1.1)
APPROXIMATE within $\epsilon$ all its zeros lying in the unit disc $D(0,1)$. [^5] We will also refer to $z_{1},\dots,z_{d}$ as “roots of equation $p=0$” or just “roots”.
(i) Two ways from the zeros of $p$ lying in the unit disc to all its zeros.
(a) Given an algorithm $\mathcal{A}$ that solves Problem 0, we can compute approximations $\tilde{y}_{j}$ within $\epsilon$ to the reciprocals $y_{j}=1/x_{j}$ of all zeros $x_{j}$ of $p$ lying outside the disc $D(0,1)$ [130] – simply by applying this algorithm to the reverse polynomial
$p_{\rm rev}(x):=x^{d}p\Big{(}\frac{1}{x}\Big{)}=\sum_{i=0}^{d}p_{i}x^{d-i},~{}p_{\rm rev}(x)=p_{0}\prod_{j=1}^{d}\Big{(}x-\frac{1}{z_{j}}\Big{)}~{}{\rm if}~{}p_{0}\neq 0.$ (1.2)
Write $\Delta_{j}:=\tilde{y}_{j}-y_{j}$ and deduce that $\frac{1}{y_{j}}-\frac{1}{\tilde{y}_{j}}=\frac{\Delta_{j}}{\tilde{y}_{j}~{}y_{j}}$, and so $\frac{\tilde{y}_{j}}{y_{j}}-1=\frac{\Delta_{j}}{y_{j}}=\Delta_{j}x_{j}$. Recall that $|\Delta_{j}|\leq\epsilon$ by assumption and obtain that $|\frac{\tilde{y}_{j}}{y_{j}}-1|\leq\epsilon|x_{j}|$, thus arriving at the following
Proposition 1.1.
The reciprocal of an approximation within $\epsilon$ to a zero of $p_{\rm rev}(x)$ in the disc $D(0,1)$ approximates the associated zero of $p(x)$ lying outside that disc within a relative error at most $\epsilon$.
(b) If in addition to an algorithm $\mathcal{A}$ we are given a value
$R\geq\max_{j=1}^{d}\{|x_{j}|\},$ (1.3)
then we can scale the variable $x\mapsto y=\frac{x}{R}$, $p(x)\mapsto p(y)=p(\frac{x}{R})$ and then approximate every root $x_{y}$, for $j=1,\dots,d$, within $\epsilon R$ by solving Problem 0.
(ii) From the zeros of $p$ in the unit disc to all zeros in a Region of Interest.
For any pair of complex $c$ and positive $\rho$ we can reduce root-finding in the disc $D(c,\rho)$ to Problem 0 by means of applying the linear map
$x\mapsto\frac{x-c}{\rho},~{}p(x)\mapsto t(x)=p\Big{(}\frac{x-c}{\rho}\Big{)},~{}D(c,\rho)\mapsto D(0,1),$ (1.4)
which shifts the variable $x$ by $c$ and then scales $x-c$ by $1/\rho$.
Now, given any bounded region of interest $\mathbb{S}$ on the complex plane, we can successively (a) cover it with a minimal disc $D(c,\rho)$ (see Remark 1.4 of Sec. 1.6.2 for closer covers), (b) map this disc into the disc $D(0,1)$, (c) solve Problem 0 for the resulting polynomial $t(x)$ of (1.4), and finally, (d) apply map (1.4) in the reverse direction to obtain approximations to all zeros of $p$ lying in the intersection of the $\epsilon$-neighborhood $\mathbb{S}_{\epsilon}$ of the region $\mathbb{S}$ with the covering disc $D(c,\rho)$.
1.3 Black Box Polynomial Root-finding and Eigenvalue Approximation: Problem Statements
A nontrivial variant of Newton’s iterations in the FOCS 2016 paper [73] by Louis and Vempala approximates within a fixed positive $\epsilon\leq 1/2$ an absolutely largest zero $z$ of a real-rooted polynomial $p$ of a degree $d$ at the cost of its evaluation with a precision in $O((b+\log(R))\log(R~{}d)))$ at $O((b+\log(R))\log(d))$ points for $b:=\log_{2}(R/\epsilon)$ and a fixed $R$ satisfying (1.3). While the previous polynomial root-finders only worked for polynomials given in the monomial basis, that is, with their coefficients, the algorithm of [73] never involves them and works for a black box polynomial $p$ – given with an oracle (black box subroutine) for its evaluation. Next we modify Problem 0 accordingly.
Problem 1: Black Box Polynomial Root Approximation.
INPUT: a positive $\epsilon\in\mathbb{R}$, $d\in\mathbb{N}$, and a $d$th degree black box polynomial $p$.
APPROXIMATE within $\epsilon$ the zeros of $p$ lying in the unit disc $D(0,1)$.
By applying their root-finder to the characteristic polynomial $p(x)=\det(xI-M)$ of a $d\times d$ symmetric matrix $M$, Louis and Vempala approximated its absolutely largest eigenvalue by using a record Las Vegas expected number of bit operations. [^6] Under Las Vegas randomization one allows output errors but detects them at a dominated computational cost. Accordingly, we extend Problem 0 as follows:
Problem 2: Approximation of the Eigenvalues of a Matrix. [^7] As in [73], the problem covers or can be extended to the case where an input is a polynomial matrix, aka matrix polynomial.
INPUT: positive $\epsilon\in\mathbb{R}$, $d\in\mathbb{N}$, and a $d\times d$ matrix $M$.
APPROXIMATE within $\epsilon$ the eigenvalues of $M$ lying in the unit disc $D(0,1)$.
The algorithms for Problems 0, 1, and 2 may output multiple approximations within $\epsilon$ to the same roots and eigenvalues, but in the next modification we avoid such a confusion.
Problems $0^{*}$, $1^{*}$, $2^{*}$: Isolation of Clusters of the Zeros of $p$ and the Eigenvalues of $M$. Under the assumptions of Problems 0, 1, and 2 compute at most $d$ disjoint discs of radii at most $\epsilon$ (hereafter $\epsilon$-discs) covering all zeros or eigenvalues that lie in the disc $D(0,1)$ and also compute the number of the zeros or eigenvalues lying in each $\epsilon$-disc.
Remark 1.1.
If we collapse all zeros of $p$ lying in every $\epsilon$-disc into a single multiple zero placed in the center of the disc, then these zeros will lie at the distances greater than $2\epsilon$ from each other. If polynomial $p$ is square-free (having only simple zeros) and has integer coefficients, then we would arrive at the solution of the well-known Problem of Root-isolation (see Problem ISOL and Remark C.4 in Sec. C.5). These comments can be immediately extended to Problem $*{}^{*}$.
1.4 Some Benefits of Black Box Polynomial Root-finders
(i) Black box polynomial root-finders are particularly efficient for $d$th degree polynomials $p$ that can be evaluated fast, say, in $O(\log(d))$ arithmetic operations (ops), such as Mandelbrot-like polynomials $p:=M_{k}(x)$, where $M_{0}:=x,~{}M_{i+1}(x):=xM_{i}^{2}(x)+1$, $i=0,1,\dots,k$, $d=2^{k}-1$, and the sums of a bounded number of shifted monomials, such as $p:=\alpha(x-a)^{d}+\beta(x-b)^{d}+\gamma(x-c)^{d}$ for six constants $a$, $b$, $c$, $\alpha$, $\beta$, and $\gamma$. For comparison, Horner’s algorithm uses $2d$ ops to evaluate at a point $x$ a general polynomial $p=p(x)$ of (1.1) given with its $d+1$ coefficients, and this bound is optimal (see [85, 66] or [135, Section “Pan’s method"]). One can evaluate $p$ at $d_{+}>d$ points and then interpolate to it, but this would destroy sparsity, blow up the precision and bit operation cost of the computations, and dramatically slow down root-finding in a region containing a small number of roots.
(ii) The known efficient polynomial root-finders extensively shift (translate) and scale variable $x$ and in particular routinely apply map (1.4) to reduce root-finding in any disc $D(c,\rho)$ to root-finding in the unit disc $D(0,1)$. All this, however, is done at the price of losing sparsity and blowing up computational precision and bit operation complexity of root-finding.
(iii) The same problems severely limit application of the DLG classical root-squaring iterations of Dandelin 1826, Lobachevsky 1834, and Gräffe 1837 (see [49, 82, 109] for the history), which compute polynomials $p_{2^{i}}(x):=\prod_{j=1}^{d}(x-z_{j}^{2^{i}})$, for $i=0,1,\dots$, as follows:
$p_{0}(x):=\frac{p(x)}{p_{d}},~{}p_{2^{i+1}}(x^{2}):=(-1)^{d}p_{2^{i}}(x)p_{2^{i}}(-x),~{}i=0,1,\dots$ (1.5)
Among various helpful properties of iterations (1.5), root-squaring also squares the distances from the origin to the roots, thus moving the external roots away from the unit circle $C(0,1)$, that is, strengthens its isolation from the roots.
By avoiding computation of the coefficients of $t(x)$ and $p_{2^{i}}(x)$ of maps (1.4) and (1.5), respectively, black box polynomial root-finders keep all benefits of using these maps but avoid the cited notorious problems.
For yet another benefit we can apply black box root-finders to polynomials in Bernstein form [71], whose evaluation is known to be numerically stable.
Remark 1.2.
Polynomial root-finding in the 19-th and well into the 20-th century relied on these iterations combined with Vieta’s formulas, which link the polynomial coefficients $p_{0},p_{1},\dots,p_{d}$ to the elementary symmetric polynomials of the roots. The combination enabled efficient approximation of the minimal and maximal distance of the roots from the origin and due to the map (1.4) from any complex point (see [82, 109, 111] for details). The approach became obsolete because of severe problems of numerical stability of the DLG iterations, but they disappeared in the case of root-finders that work for black box polynomials. Extensive tests in [111] with standard test polynomials have consistently demonstrated efficiency of a single particular implementation of this approach, but even the straightforward implementation of iterations (1.5) promises to be highly efficient as well.
1.5 Our complexity estimates
We adopt the model of [73] but otherwise rely on distinct approach and techniques (with few exceptions listed below). Indeed, the root-finder of [73] cannot help solve even Problem 1: even in the explicit setting, where the coefficients of $p$ are given, repeatedly deflating $p$, by peeling off large roots, would accumulate errors that grow too quickly, and it is not entirely clear how to even do this in the black box setting. Even if this recipe worked it would only solve the well-known and much easier problem of root-finding for a real-rooted polynomial (cf. [89, 24, 31]). Moreover, the resulting algorithm for that much easier problem would involve more than $b^{2}d^{2}$ bit operations, which would by a factor of $b$ exceed the upper bound of [92, 99] for general polynomial root-finding. [^8] For root-finding problems one should assume that $b$ exceeds $d$ because the precision of representation of $p$ (input length) increases by a factor of $d$ in routine shifts and scaling of the variable $x$.
Recursive combination of [73] with the Inverse Power Iterations [46] would enable approximation of all eigenvalues of a symmetric (although not unsymmetric) matrix, together with its eigenvectors. One can be challenged to estimate the increased computational complexity of the solution of this extended problem of matrix computations, but we study other approaches.
In particular we accelerate approximation of $m$ zeros of $p$ for $m\leq d$ by a factor of $m$ by means of incorporation of fast multipoint polynomial evaluation of [40, 81, 65, 78], which we could not use if we successively approximated zeros of $p$ or matrix eigenvalues by using deflation.
As in [73] we reduce root-finding to the evaluation of $p(x)$ at sufficiently many points. The ops involved in this evaluation are more numerous than the other ops involved in our algorithms, even in the case where $p(x)=x^{d}$.
Unlike [73] we use Las Vegas randomization – to control with a high probability (whp) precision and bit-operation complexity of our computation. In particular we readily prove that whp $p(x)$ and $p^{\prime}(x)$ are not too close to 0 for a random variable $x$, and so we perform division safely whp when we approximate Newton’s ratio NR$(x):=\frac{p(x)}{p^{\prime}(x)}$ or its reciprocal – Newtons Inverse Ratio NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ at random points $x$.
For these basic steps of our root-finders we fix a complex $\delta$ whose absolute value is sufficiently small in context and approximate $p^{\prime}(x)$ with the ratio $\frac{p(x)-p(x-\delta)}{\delta}$. This ratio is the first order divided difference, equal to $p^{\prime}(y)$ for some $y\in[x-\delta,x]$ by virtue of Taylor-Lagrange’s formula, [^9] Unlike [73] we do not need to approximate higher order derivatives and thus involve no higher order divided differences; by using Las Vegas randomization, however, we still ensure whp the same bound $O(b\log(d))$ of [73] on the computational precision. and so $y\mapsto x$ and $p^{\prime}(y)\mapsto p^{\prime}(x)$ as $\delta\mapsto 0$. Substitute the ratio for $p^{\prime}(x)$ and obtain that
${\rm NIR}(x)\approx\frac{p(x)-p(x-\delta)}{p(x)\delta}=\frac{p^{\prime}(y)}{p(x)}~{}{\rm for~{}some}~{}y\in[x-\delta,x].$ (1.6)
Our root-finders approximate the roots lying in the unit disc $D(0,1)$ together with some roots lying nearby. At the end we only keep the roots lying in the disc $D(0,1+\epsilon)$ but at the intermediate steps may approximate some extraneous roots in a larger domain, but we modify subdivision in Sec. 1.6.2 so that all extraneous roots lie in the disc $D(0,1+\beta)$ for any fixed positive $\beta$.
Definition 1.1.
Hereafter $m=m(\beta)$ denotes the number of the zeros of $p$ lying in the disc $D(0,1+\beta)$ for a fixed positive constant $\beta$ of our choice. [^10] See our Cor. 1.2 and compare more restrictive choice of $\beta=2$ in [22, 23].
Our Las Vegas randomized root-finders support the Main Theorems 1-4 below.
Theorem 1.1.
(Main Theorem 1.) One can solve Problems 1 and $*{}^{*}$ by applying Las Vegas randomized queries of the polynomial evaluation oracle with a computational precision $O(b\log(d))$ at $\tilde{O}(dm)$ locations for $b:=\log_{2}(1/\epsilon)$ and $m$ of Def. 1.1.
By following [73] one can readily extend Main Theorem 1 to matrix eigenvalue approximation:
Theorem 1.2.
(Main Theorem 2.) One can solve Problems 2 and $*{}^{*}$ at Las Vegas randomized bit operation cost
$\mathbb{B}_{\rm eigen}=\tilde{O}(\tilde{b}d^{\omega}m)$ (1.7)
for $m$ of Def. 1.1, $\omega$ denoting an exponent of feasible or unfeasible matrix multiplication (MM),  [^11] The current records are about 2.7734 for exponents of feasible MM [86, 103, 105, 69, 54], unbeaten since 1982, and about 2.372 for exponents of unfeasible MM [6, 35]. The exponents below 2.77 are only reached for unfeasible MM – supported by algorithms that are slower than the classical $d\times d$ MM unless $\log_{2}(\log_{2}(\log_{2}(d))))>100$, say. and $\tilde{b}:=\log_{2}(||M||_{F}/\epsilon)$ for the Frobenius norm $||M||_{F}$ of a matrix $M$.
Thm. 1.1 implies the bound $\tilde{O}(bd^{2}m)$ on Las Vegas randomized bit operation complexity of Problem 1 because one can evaluate any polynomial $p$ of degree $d$ in $2d$ ops and perform an op with the precision of $s$ bits by using $O(s\log(s)\log(\log(s)))=\tilde{O}(s)$ bit operations, which extends the celebrated bound
$\mu(s):=O(s\log(s)\log(\log(s)))=\tilde{O}(s)$ (1.8)
on the number of bit operations for multiplication of two integers modulo $2^{s}$ [137] (see [66, 130, 131] for this extension).
In the case of the Mandelbrot polynomials $p:=M_{k}(x)$ and the sums of a constant number of shifted monomials one can immediately decrease by a factor of $d/\log(d)$ both of the above upper bounds of $2d$ ops and $\tilde{O}(bd^{2}m)$ bit-operations.
Even in the worst case of general polynomial $p$ given with its coefficients, however, we decrease the upper bound on the bit operation cost of our root-finders by roughly a factor of $d$ – to match the near optimal bound of [92, 99] in our next theorem. We achieve this by combining our root-finders supporting Main Theorem 1 with the known algorithms for fast multipoint polynomial evaluation [40, 81, 65, 78].
Theorem 1.3.
(Main Theorem 3.) One can solve Problems 0 and $*{}^{*}$ at Las Vegas randomized bit operation cost $\tilde{O}((b+d)dm)$ for $b:=\log_{2}(1/\epsilon)$, $m$ of Def. 1.1, and general polynomial $p$ of (1.1) normalized such that $||p(x)||_{1}:=\sum_{i=1}^{d}|p_{i}|=1$.
Finally, we decrease our upper bounds of Thms. 1.1-1.3 by a factor of $d$ under an additional assumption on the disposition of the zeros of $p$ and eigenvalues of $M$, namely, under the following model: [^12] Its assumption is a little stronger than that of our Def. B.1, supporting such a decrease in [92, 99], but it is in good accordance with the results of extensive tests in [57, 60] for standard test polynomials (see Sec. 13.4.1 and the beginning of Sec. 4.1).
Random Root Model: The zeros of $p$ (in particular for $p$ denoting the characteristic polynomial of a matrix $M$) are independent identically distributed (iid) random variables sampled under the uniform probability distribution on the disc $D(0,R)$ for a large $R\in\mathbb{R}$ such that the value $\frac{R^{2}}{d}-1$ exceeds a positive constant.
Theorem 1.4.
(Main Theorem 4) The upper bounds of Thms. 1.1-1.3 on computational precision and bit operation complexity can be decreased by a factor of $d$ under the Random Root Model.
In view of Remarks 1.1 and C.4 we can readily extend our Main Theorems to Polynomial Root Isolation.
1.6 Our algorithms: design and analysis (outline)
For a polynomial $p$ given in monomial basis (with its coefficients) our deterministic algorithms of [92, 99] support the Main Theorems 3 and 4, but our present Las Vegas randomized algorithms does this in the case of a black box polynomial as well.
1.6.1 Subdivision root-finders and reduction to e/i tests
We depart from the popular subdivision root-finders, traced back to [144, 50, 124, 97], which extend root-finding in a line segment by means of bisection to polynomial root-finding in a square on the complex plane and have been extensively used in Computational Geometry under the name of Quad-tree Construction. Let a square contain precisely $m>0$ zeros of $p$ (with multiplicities). Call it suspect, recursively divide every suspect square into four congruent sub-squares, and to each of them apply exclusion/inclusion (e/i) test, to decide whether the square contains any root or does not. If it does, call the square suspect and process it in the next iteration; otherwise discard it (see Fig. 1a).
(a) Four roots marked by asterisks lie in suspect sub-squares; the other (empty) sub-squares are discarded.
(b) Both exclusion and $\sqrt{2}$-soft inclusion criteria hold for e/i test applied to the disc bounded by internal circle.
Figure 1:
Observation 1.1.
A subdivision iteration applied to a square with $m$ zeros of $p$ (i) decreases by twice the diameter of a suspect square and (ii) processes at most $4m$ suspect squares, (iii) whose centers approximate every zero of $p$ within the distance of one half of their diameter.
Proof.
Readily verify claim (i). Claim (ii) holds because a zero of $p$ can lie in at most four suspect squares (and even in at most two unless it is the shared vertex of the four). Claim (iii) holds because a subdivision iteration never discards a square containing a root. ∎
Corollary 1.1.
Let subdivision iterations be applied to a square $S$ with side length $\Delta$ (and diameter $\Delta\sqrt{2}$) containing precisely $m$ zeros of $p$. Then (i) the $k$-th iteration outputs at most $4m$ suspect squares whose centers approximate within $\epsilon:=\Delta/2^{k-0.5}$ all the $m$ zeros of $p$ lying in $S$, so that $k\leq\lceil\log_{2}(\frac{\Delta}{2\epsilon}\sqrt{2})\rceil$, and (ii) the first $k$ iterations together involve at most $4mk$ e/i tests.
1.6.2 Soft e/i tests and compression of a Region of Interest
Instead of testing whether a suspect square $S$ contains some zeros of $p$, Henrici in [50] tested whether the minimal disc $D(c,\rho)$ covering $S$ does this. We adopt his approach, which implies that the zeros of $p$ lying outside the square but in the disc can affect the result of our e/i test. Furthermore, in the presence of rounding errors no e/i test can decide whether a zero of $p$ lies in or outside a disc if it actually lies on or very close to its boundary circle. To address both problems we fix $\sigma>1$ and apply soft e/i tests as follows.
Definition 1.2.
Given a real $\sigma>1$, a polynomial $p$, a square $S$, and its minimal covering disc $D(c,\rho)$, the $\sigma$-soft e/i test reports either exclusion or $\sigma$-soft inclusion. Exclusion certifies that the square $S$ contains no zeros of $p$ and therefore should be discarded from subdivision process. $\sigma$-soft inclusion certifies that the disc $D(c,\sigma\rho)$, said to be the $\sigma$-cover of $S$, contains a root; then the square $S$ is said to be suspect and is subdivided into four congruent sub-squares in the next subdivision step. Criteria for both exclusion and $\sigma$-soft inclusion can hold simultaneously (see Fig. 1b), but the test ends as soon as it verifies any of them.
Let us estimate how transition to $\sigma$-soft e/i tests affects Cor. 1.1.131313We counter the adverse impact of rounding errors by increasing $\sigma$ but only need tiny increase for small $\epsilon$ and will ignore it in our estimates.
Observation 1.2.
(i) The number $n(\sigma)$ of the $\sigma$-covers of suspect squares that contains a fixed zero of $p$ is $O(1)$, that is, bounded by a constant, if so is $\sigma$. (ii) In particular $n(\sigma)\leq 4$ if $\sigma<\sqrt{2}$.
Remark 1.3.
We only use $\sigma$-soft e/i tests for $\sigma<\sqrt{2}$ in our root-finders.
Seeking roots in the unit disc $D(0,1)$ we apply subdivision iterations to its minimal covering square $S(0,1)=\{x:~{}|\Re(x)|\leq 1,~{}|\Im(x)|\leq 1\}$, having side length 2 (see Fig. 1b) and $\sigma$-cover $D(0,\sigma\sqrt{2})$.
Observation 1.3.
Let $U_{i}=U_{i}(\sigma)$ denote the union of the $\sigma$-covers of all suspect squares processed at the $i$-th subdivision step with $\sigma$-soft e/i tests. Then (i) $U_{i+1}(\sigma)\subset U_{i}(\sigma)$, that is, domain $U_{i}(\sigma)$ shrinks as $i$ increases, (ii) $U_{i}(\sigma)\subset D(0,2)$ for all $i\geq 0$ if $\sigma<\sqrt{2}$, and (iii) $U_{h}(\sigma)\subset D(1+\beta)$ for any fixed positive $\beta$ and some $h=h(\sigma,\beta)$, which is a constant, $O(1)$, if so is $\sigma$.
Under the Random Root Model the numbers of roots lying in the discs $D(0,1)$ and $D(0,2)$ differ by a constant factor $f=O(1)$ whp, and so for $\sigma<\sqrt{2}$ the estimates of Cor. 1.1 hold up to that factor whp, due to claim (ii) of the latter observation.
Based on its claims (i) and (iii) we next extend Cor. 1.1 deterministically, by incorporating the following two recipes of Extension and Delay for discarding squares in subdivision process.
(a) Extend discarding: In addition to discarding squares according to $\sigma$-soft e/i tests, discard all squares that lie outside the disc $D(0,1)$.
(b) Delay discarding: For a fixed positive $\beta$ begin applying $\sigma$-soft e/i tests and discarding squares only at the $h$th subdivision step for $h=h(\sigma,\beta)$ of claim (iii) of Observation 1.3.
Clearly, the $h$th subdivision iteration may process up to $4^{h}=O(1)$ suspect squares, although a little fewer for larger $h$ if the extension of discarding is applied.
Now we can extend Cor. 1.1 as follows.
Corollary 1.2.
Fix $\beta>0$ and $\sigma>1$, let the disc $D(0,1+\beta)$ contain precisely $m=m(\beta)$ zeros of $p$ (cf. Def. 1.1), and apply subdivision iterations (with $\sigma$-soft e/i tests under the recipes of extension and delay of discarding) to a minimal square $S(0,1)$ covering the unit disc $D(0,1)$. Then (i) the $k$-th iteration involves at most $n(\sigma)m$ suspect squares for $n(\sigma)$ of Observation 1.2, $n(\sigma)\leq 4$ for $\sigma<\sqrt{2}$; the centers of those squares approximate within $2^{1.5-k}$ all the zeros of $p$ lying in the disc $D(0,1)$ and (ii) the first $k$ subdivision iterations involve no suspect squares and $\sigma$-soft e/i tests for $k<h=h(\sigma,\beta)$ and at most $(k-h-1)n(\sigma)m+4^{h}$ suspect squares and $\sigma$-soft e/i tests for $k\geq h$.
Remark 1.4.
We extend Cor. 1.1 for subdivision process applied to the unit disc $D(0,1)$ but can similarly do this for any fixed Region of Interest.
1.6.3 Cauchy-based root-counting and e/i tests
We have highly efficient heuristic algorithms based on the DLG iterations (1.5) that closely approximate the smallest distance to the roots from a complex point (see Remark 1.2); this can be immediately extended to soft e/i tests.
Next we cover efficient e/i tests that also enable formal support of our Main Theorems 1– 4.
Let #$\#(\mathcal{D})$ denote the number of zeros of $p$ lying in a complex domain $\mathcal{D}$ with the boundary $\Gamma$ and recall from [1] that
[…]
that is, #$\#(\mathcal{D})$ is the average of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ on the boundary $\Gamma$.
Having an approximation of the integral of (1.9) within less than 1/2, we can round it to the closest integer and obtain $s_{0}=\#(D(0,1))$. This is more than enough for our e/i tests, for which we only need to know whether $s_{0}=0$ or not.
In Secs. 1.6.4 and 1.6.5 we outline our two approaches to devising e/i tests for a black box polynomial based on (1.9) for $\mathcal{D}=D(0,1)$ We cover these two approaches in some detail in Part II, made up of Secs. 29, and Part III, made up of Secs. 10-15, respectively.
The subdivision root-finders that incorporate these e/i tests are quite fast but only support root approximation (rather than isolation) and the complexity estimates exceeding those of our Main Theorems by a factor of $b$. In Part IV, made up of Secs. 16-20 and outlined in Sec. 1.6.7, we fix both deficiencies, thus proving our Main Theorems.
1.6.4 Our soft e/i tests in Part II: outline
In Part II we approximate the integral of (1.9) with finite sums made up of the scaled values of NIR$(x)$ at the scaled $q$th roots of unity. Namely, for a fixed $q>1$ let $\zeta:=\zeta_{q}$ denote a primitive $q$th root of unity:
$\zeta:=\zeta_{q}:=\exp\Big{(}\frac{2\pi{\bf i}}{q}\Big{)},~{}{\bf i}:=\sqrt{-1},$ (1.10)
and devise randomized $\sigma$-soft e/i tests as follows:
(i) choose a real $\rho$ at random in a fixed range covered by $[1,\sigma]$,
(ii) fix a sufficiently large integer $q$ of order $m\log(d)$, and
(iii) compute approximation $s_{0,q}$ to the integral (1.9) for the disc $D(0,\rho)$ and the polynomial $t(x)=p(\rho x)$ replacing $p(x)$.
[…]
(iv) Obtain #$\#(D(0,\rho))$ as the closest integer approximation of $s_{0,q}$, prove that the recipe is correct whp, and extend the root-counter to e/i test.
(v) Repeat the test $v$ times to decrease the probability of its output error exponentially in $v$.
The $v$ tests are reduced to the evaluation of $p(x)$ at $O(mv\log(d))$ points; in Part II we also devise a deterministic soft e/i test at the price of increasing the above bound on the number of points to order of $m^{2}\log(d)$.
1.6.5 Our generic soft e/i test in Part III
(i) In Part III we decrease our estimates of Part II for $q$ in e/i tests by a factor of $m$ by avoiding approximation of the integer #$\#(D(0,1))$ and only deciding whether this integer vanishes or not, which is sufficient for e/i tests. In these e/i tests we elaborate upon the following generic one, where we use $h$-lifting of roots generalizing DLG root-squaring (1.5):
$p_{0}(x):=\frac{p(x)}{p_{d}},~{}p_{h}(x):=\prod_{g=0}^{h-1}p_{0}(\zeta_{h}^{g}x)=(-1)^{d}\prod_{j=1}^{d}(x-z_{j}^{h}),~{}\zeta_{h}:=\exp\Big{(}\frac{2\pi{\bf i}}{h}\Big{)},$ (1.11)
for $h=1,2,\dots$ and $t_{2^{k}}(x)=p_{k}(x)$ for all $k$.
Algorithm 1.1.
A generic soft e/i test.
INPUT: a black box polynomial $p(x)$ of a degree $d$, given by an oracle for its evaluation, a real $\sigma>1$ and a positive integer $h$ (both specified in Sec. 11).
INITIALIZATION: Fix a positive $q\in\mathbb{N}$.
COMPUTATIONS:
(1) Select $q$ points $x_{1},\dots,x_{q}$ on the circle $C(0,1)$.
(2) By extending (1.6) to $p_{h}(x)$ of (1.5) (we specify this extension in (10.5)) closely approximate
$v_{i}:=|{\rm NIR}_{p_{h}}(x_{i})|=\Big{|}\frac{p^{\prime}_{h}(x_{i})}{p_{h}(x_{i})}\Big{|},~{}i=1,\dots,q.$ (1.12)
(3) Certify that #$(D(0,\sigma))>0$, that is, claim $\sigma$-soft inclusion, if $1+d/v_{j}\leq\sigma^{h}$ for some $i\leq q$. Otherwise, claim that #$(D(0,1))=0$, that is, claim exclusion.
(ii) To prove correctness of soft inclusion claim of Alg. 1.1, first differentiate $p(x)=p_{d}\prod_{j=1}^{d}(x-z_{j})$ and obtain the following well-known equation:
${\rm NIR}(x):=\frac{p^{\prime}(x)}{p(x)}=\sum_{j=1}^{d}\frac{1}{x-z_{j}}.$ (1.13)
Eqn. (1.13) implies two useful corollaries:
Corollary 1.3.
Map (1.4) scales NIR by $\rho$.
Proof.
Eqn. (1.13) implies that NIR$(x-c)=$NIR$(x)$ and that NIR$(x/a)=a$NIR$(x)$. ∎
Corollary 1.4.
#$\#(D(c,\rho))>0$ if $\rho\leq d~{}|p(c)/p^{\prime}(c)|$.
Proof.
Let $x=c$ in Eqn. (1.13). If $|c-z_{j}|<\rho$ for all $j$, then
$\Big{|}\frac{p^{\prime}(c)}{p(c)}\Big{|}=\Big{|}\sum_{j=1}^{d}\frac{1}{c-z_{j}}\Big{|}\leq\sum_{j=1}^{d}\frac{1}{|c-z_{j}|}<\frac{d}{\rho},~{}{\rm and~{}so}~{}\Big{|}\frac{p(c)}{p^{\prime}(c)}\Big{|}>\frac{\rho}{d}.$
We use Cor. 1.3 later but next apply Cor. 1.4 to prove the following lemma, which implies correctness of the soft inclusion claim of Alg. 1.1.
Lemma 1.1.
Let $\rho:=d\Big{|}\frac{p_{h}(x)}{p^{\prime}_{h}(x)}\Big{|}$ and $\sigma:=(|x|+\rho)^{1/h}$ for a positive integer $h$. Then #$(D(0,\sigma))>0$.
Proof.
Apply Cor. 1.4 to the polynomial $p_{h}(x)$ and obtain that the equation $\rho=d\Big{|}\frac{p_{h}(x_{)}}{p^{\prime}_{h}(x)}\Big{|}$ implies that the disc $D(x,\rho)$ contains a zero of $p_{h}(x)$. Hence the disc $D(0,|x|+\rho)$ contains a zero of $p_{h}(x)$ by virtue of triangle inequality. Therefore, the disc $D(0,\sigma)$ contains a zero of $p(x)$ for $\sigma=(|x|+\rho)^{1/h}$. ∎
Apply this lemma for $x=x_{i}$, $|x_{i}|=1$, $\rho=d/v_{i}$, and $v_{i}=|{\rm NIR}_{p_{h}}(x_{i})|$ (cf. Alg. 1.1). Deduce that #$(D(0,\sigma))>0$ for $\sigma^{h}=1+d/v_{i}$, that is, Alg. 1.1 is correct when it claims $\sigma$-soft inclusion.
Observation 1.4.
$\sigma$-soft inclusion test by means of Alg. 1.1 is correct where
$d\geq 2,~{}x_{1}=1,~{}\frac{1}{|{\rm NIR}(1)|}=4\sqrt{d}-\frac{1}{d},~{}h=40\log_{2}(d),~{}{\rm and}~{}\sigma<1.1.$
Proof.
Indeed, in this case
$1+\rho=1+\frac{d}{|{\rm NIR}(1)|}=4d\sqrt{d}=2^{2+1.5\log_{2}(d)}.$
Apply Lemma 1.1 and obtain that #$(D(0,\sigma))>0$ for $\sigma=(1+\rho)^{1/h}=2^{\nu}~{}{\rm for}~{}\nu=\frac{2+1.5\log_{2}(d)}{40\log_{2}(d)}$.
Deduce that $\nu\leq\frac{7}{80}$ because $\frac{2}{40\log_{2}(d)}\leq 0.05$ for $d\geq 2$ and because $\frac{2+1.5\log_{2}(d)}{40\log_{2}(d)}<\frac{3}{80}$.
Hence $\sigma\leq 2^{\frac{7}{80}}=1.062527\dots<1.1$. ∎
(b) We prove correctness of the exclusion claim of Alg. 1.1 separately for its two specifications, both of which strengthen isolation of the unit circle $C(0,1)$ by means of root-squaring (1.5) or root-lifting (1.11). In Sec. 11.3 we choose any integer $q$ exceeding $d$ and then prove deterministic correctness of exclusion claim, provided that $|{{\rm NIR}_{p_{h}}(x_{i})}|<1/(2\sqrt{q})$ for all $x_{i}$ being $q$th roots of unity. This bound is compatible with the bound of Example 1.4 because $4\sqrt{d}-\frac{1}{d}>3\sqrt{d}$ for $d\geq 2$. We also extend this deterministic e/i test by allowing any $q>m$ (rather than $q>d$) provided that $|{{\rm NIR}_{p_{h}}(x_{i})}|<1/(3\sqrt{q})$ for all $x_{i}$ being $q$th roots of unity.
In Sec. 11.4 we sample $q=O(\log(d))$ iid random values of $x_{i}$ under the uniform probability distribution on the circle $C(0,1)$ and then prove that the exclusion claim is correct whp under the Random Root Model if $|{{\rm NIR}_{p_{h}}(x_{i})}|\leq 1/d^{a}$ for a sufficiently large constant $a$.
(iv) Counting query locations. Approximation of $p_{h}(x)$ is reduced to approximation of $p(x)$ at all the $h$th roots of $x$ for $h$ being a power of 2 based on (1.5) or for any integer $h>1$ if we apply (1.11), and so Alg. 1.1 evaluates polynomial $p(x)$ either at $2jh$ points for $j\leq q$, when it claims inclusion, or at $2qh$ points, when it claims exclusion. In all these estimates we choose factor $h$ in $O(\log(d))$, and so we can ignore this factor in our $\tilde{O}$ estimates and say that our soft e/i test by means of Alg. 1.1 evaluates polynomial $p(x)$ at $\tilde{O}(m)$ points. Combine these estimates with Cor. 1.2 to bound by $\tilde{O}(bm^{2})$ the overall number of points $x$ at which we evaluate $p(x)$ in subdivision root-finders incorporating our e/i tests.
Under the Random Root Model of Sec. 1.5 the zeros of $p$ are isolated whp from the unit circle $C(0,1)$, and we decrease the bound $\tilde{O}(bm^{2})$ to $\tilde{O}(bm)$ whp by exploit their even stronger isolation ensured by means of $O(\log(\log(d)))$ root-squaring steps (1.5) or alternatively with $O(\log(d))$ steps of root-lifting of (1.11).
1.6.6 Computational precision and bit-operation complexity
Louis and Vempala in [73] bound by $O(b\log(d))$ the computational precision of their novel high order Newton’s iterations. We also prove this bound for our root-finders by using a small part of their analysis together with Las Vegas randomization and some rather customary techniques of numerical analysis.
We extend this bound combined with the bound on the number of queries in our root-finders to the bit operation complexity estimates: we use the known fast algorithms for multipoint polynomial evaluation. We satisfy the assumptions of the latter algorithms by using map (1.4).
We devise two groups of supporting algorithms in Parts II and III of the paper, and in both cases arrive at the estimates for the numbers of evaluation points and bit-operations involved in our algorithms that exceed those of our Main Theorems by factors of $b$ for the roots and $\tilde{b}$ for the eigenvalues.
1.6.7 Acceleration of root-finding in Part IV: outline
In Part IV we accelerate our algorithms to yield the Main Theorems. Here is an outline of that acceleration.
(i) Component tree and compact components. Apply subdivision process for a fixed $\sigma$, $1<\sigma<\sqrt{2}$, to Problem 1 and partition the union of all suspect squares tested at every subdivision step into connected components.
Represent all partitions of components with a component tree as follows. At most $m$ components, said to be the leaves, are covered by isolated $\epsilon$-discs for a fixed error tolerance $\epsilon\geq 0$. The other (at most $m-1$) vertices of the tree are components-parents, each broken by a subdivision step into at least two components-children. The initial suspect square is the root of the tree. The tree has at most $22m-$ 2 edges, connecting parents with their children and representing partition of the components but not showing modification of components between their partitions.
Recall that every component contains $O(m)$ suspect squares whose diameter is halved at every subdivision step. Hence in $O(\log_{2}(mb))$ steps, for $b:=\log_{2}(1/\epsilon)$, every component has diameter below $\epsilon/2$, and then its $\sigma$-cover lies in an $\epsilon$-disc.
Call a component compact if it consists of at most four squares sharing a vertex (see Fig. 2). Recall that a subdivision step never includes new roots into a component and deduce that unless a component stays compact it is soon broken into more than one component or shrinks enough to be covered by an $\epsilon$-disc.
[…]
Figure 2: Roots (asterisks) define compact components. A subdivision step halves their diameters and accordingly strengthens isolation of their minimal covering squares and discs.
(ii) Acceleration of subdivision process with compression of components. The bound $O(m)$ does not generally hold for the overall number of suspect squares at all subdivision steps. This is because subdivision steps become inefficient where a component stays compact and not broken in many consecutive subdivision steps (see Fig. 3), but we replace these steps with compression of such a component $C$ into a near minimal disc $D$ that shares with $C$ all the zeros of $p$ lying in it.
Figure 3: Five blue marked centers of suspect squares converge to a green-marked root cluster with linear rate; two red marked Schröder’s iterates of (20.1) converge with quadratic rate.
If the disc $D$ is an $\epsilon$-disc we just output it. Otherwise we apply subdivision iterations to its minimal superscribing square $S$. Since we compress the component into a near minimal disc $D$, we obtain readily that these iterations very soon output more than one component or a sub-domain of an $\epsilon$-disc. From this we deduce that our subdivision process with compression reduces the solution of Problem 1 to at most $22m-$ 2 compression steps and $O(m)$ soft e/i steps overall.
(iii) Compression of components. We reduce a compression of a compact component to the following main stages (see more details in Sec. 20.5):
(i) compute a point $c$ lying in or near the minimal disc that covers all zeros of $p$ lying in the component $C$ and let $c$ denote the center of the compressed disc that we seek;
(ii) compute a close upper bound $\rho$ on the distance from $c$ to the farthest zero of $p$ among its at most $m$ zeros lying in the compact component $C$.
We prove that Schröder’s single celebrated iteration (20.1) following $O(\log(d))$ auxiliary subdivision steps performs step (i) above, that is, computes the desired center $c$.
Beginning stage (ii) above we have a range $[\rho_{-},\rho_{+}]$ where the bound $\rho$ lies, for $10<\rho_{-}<\rho_{+}<$ 1. Hence we can approximate it in $\lceil\log_{2}(1/\epsilon)\rceil$ bisection steps, but we greatly accelerate convergence to $\rho$ by applying bisection to the range of the exponents, $[\log_{2}(\rho_{-}),\log_{2}(\rho_{+})]$.
We extend our soft e/i tests to define the pivot step of such a modified bisection. Summarizing we reduce the solution of Problem 1 to $\tilde{O}(m)$ soft e/i tests and $\tilde{O}(m)$ Schröder’s iterations.
In this way we accelerate the root-finders of Parts II and III by a factor of $b$. Together with our recipes for detection and correction of output errors of root-finders in the next subsection, this completes the proof of our Main Theorems.
Remark 1.5.
Many “natural" directions to proving our Main Theorems have led us into dead end. E.g., we defined a tree whose vertices represented non-adjacent suspect squares; then we readily bounded by $12m-$ 1 their overall number at all partition stages but could not count them so readily between these stages until we decided to count all suspect squares. See more examples of this kind in the next remark and Remark 11.1.
Remark 1.6.
For bounding precision of computing, Louis and Vempala in [73] presented a major technical innovation, which enabled us to deduce a priori estimates for the bit operation cost of our subdivision root-finders versus routine a posteriori estimates in the previous papers on subdivision root-finding (see, e.g., [118, 23]). In particular, the bit operation cost estimates of the Main Theorem of [23] contain the term
$\log(1/{\rm Discr}(p(x))),$
with “Discr" standing for “Discriminant", such that
$\log(1/{\rm Discr}(p(x/2^{b})))=(d-1)db\log(1/{\rm Discr}(p(x))).$
Subdivision iterations of the known root-finders, in particular, of those of [23], scale the variable $x$ order of $m$ times by factors of $2^{b}$. Hence the bit-complexity specified in the Main Theorem of [23] has order of $mbd^{2}$, which exceeds our bound and that of [92, 99] by a factor of $d$.
1.7 Error detection and correction
A randomized root-finder can lose some roots, although with a low probability, but we can detect such a loss at the end of root-finding process, simply by observing that among the $m\leq d$ roots lying in an input disc only $m-w$ tame roots, say, $z_{w+1},z_{w+1},\dots,z_{m}$ have been closely approximated, [^14] We call a complex $c$ a tame root for an error tolerance $\epsilon$ if $c$ lies in an isolated disc $D(c,\epsilon)$, whose index #$\#(D(c,\epsilon))$ we can readily compute (cf. Sec. 20.4). while $w>0$ wild roots, say, $z_{1},z_{2},\dots,z_{w}$ remain at large. We can compute the coefficients of the wild factor $q(x):=p(x)/\prod_{j=w+1}^{m}(x-z_{j})$, say, by applying evaluation-interpolation techniques, then apply to $q(x)$ the same or another root-finder, and repeat this computation recursively until we approximate all $m$ roots in the disc. [^15] By this recipe we detect output errors at the very end of computations in any root-finder that can only lose roots but cannot wrongly generate them. Our subdivision root-finders belong to this class, but with them we can detect the loss of a root even earlier – whenever in a subdivision process applied to a square with $m$ roots we detect that $m$ exceeds the sum of the indices of all suspect squares, that is, the overall number of roots contained in them.
Readily bounds the complexity and error probability of this recursive process:
Proposition 1.2.
Let the above recipe for error detection and correction be applied to the output of a randomized root-finder performing at the cost $\alpha$ (under a fixed model), let it be correct with a probability $1/\beta$ for $\beta>1$ and let the algorithm be applied recursively, each time with error detection at the cost $\delta$. Then in $v$ recursive applications, the output is certified to be correct at the cost of less than $(\alpha+\delta)v$ with a probability at least $1-1/\beta^{v}$.
Explicit deflation can blow up the coefficient length and destroy sparsity unless the degree $w$ is sufficiently small, but we can avoid these mishaps by combining a fixed black box polynomial root-finder with implicit deflation, that is, by applying that root-finder to the quotient polynomial $q(x)$ without computing its coefficients – based on the following expression, implied by (1.13):
$\frac{q^{\prime}(x)}{q(x)}=\frac{p^{\prime}(x)}{p(x)}-\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}.$ (1.14)
In many cases one can accelerate simultaneous numerical approximation of the sum $\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}$ at many points $x$ by applying Fast Multipole Method (FMM) of [48, 27, 30].
1.8 Our novelties and related works
1.8.1 Our novelties (a short list)
Root-finding for polynomial given in the monomial basis, with their coefficients, has been studied for four millennia and very intensively in the last decades, but we had to introduce a number of nontrivial novel techniques (i) to handle efficiently black box polynomials and (ii) to accelerate the known subdivision algorithms to a near optimal level.
Some of our techniques for black box polynomial root-finding, e.g., randomization for root-finding, numerical stabilization of root-squaring, acceleration of black box subdivision root-finding by means of compression of a disc without losing any zero of $p$, acceleration of bisection by means of applying it to the exponent range, approximation of root radii and near real roots, and even our simple but novel recipes for detection of output errors of polynomial root-finders can be of independent interest.
1.8.2 Related works
Unlike deterministic solution in [92, 99] we use Las Vegas randomization but propose algorithms that work for a black box polynomial, are simpler to implement, and in the case of a polynomial $p$ given with its coefficients support the same near optimal complexity bound of [92, 99].
The pioneering paper [73] proposed a novel algorithm for approximation of an absolutely largest zero of a real-rooted black box polynomial, estimated the computational precision of that algorithm, and extended it to approximation of an absolutely largest eigenvalue of a symmetric matrix. We adopt their model of computation, their extension to approximation of matrix eigenvalues, and some of their patterns for estimating the computational precision of root-finders, but otherwise use distinct approach and techniques.
Namely, we accelerate the classical subdivision root-finders, based on e/i tests. All known e/i test operate with the coefficients of $p$: they involve higher order derivatives $p^{(i)}(x)$, $i=1,2,\dots,d$ [50, 124], Newton’s identities and fast root-squaring algorithms based on convolution [97], or Pellet’s theorem [22, 23]. Neither of these approaches works for black box polynomials.
Our soft e/i tests involve the classical root-squaring techniques of Dandelin 1826, Lobachevsky 1834, and Gräffe 1837, but we stabilize them numerically avoiding notorious problems that severely limit all previous applications of maps (1.4) of shift and scaling and (1.5) of DLG iterations.
Renegar in [124] accelerated subdivision process by a factor of $b/\log(b)$ by incorporating Newton’s root-finding iterations; the papers [97, 22, 23] combined his approach with alternative functional root-finding iterations. We apply a distinct technique of compression of a disc instead.
Estimating the bit operation complexity of our root-finders we use [130, Eqn. (9.4)] and [78, Thm. 2] for multipoint polynomial evaluation. [^16] The ingenious algorithm of Moroz supporting his [78, Thm. 2] departs from the simple observation that a polynomial $p(x-c)\mod(x-c)^{\ell-1}$ approximates $p(x)$ in the disc $D(c,\rho)$ within $O(\rho^{\ell+1})$ for small $\rho$. To approximate $p(c)$ for any $c\in D(0,1)$ Moroz used $\tilde{O}((\ell+\tau)d)$ bit operations to compute a set of complex $c_{i}$ and positive $\rho_{i}$ for $i=1,\dots,K$ and $K=O(d/\ell)$ such that $D(0,1)\subseteq\cup_{i=1}^{K}D(c_{i},\rho_{i})$; then he ensured approximation to $p(c)$ for every point $c\in D(0,1)$ within $\epsilon:=3\sum_{{}_{i}=0}^{d}|p_{i}|/2^{\ell}$ by means of approximation within $\epsilon/2$ of the lower degree polynomial $p(x-c_{i})\mod(x-c)^{\ell+1}$ for an appropriate $i$. At this point, application of [65, Alg. 5.3 and Thm. 3.9] yielded new record fast algorithm for multipoint polynomial approximation.
In Sec. 11.4 we accelerate e/i tests by a factor of $m$ under a random root model and accordingly accelerate the solution of Problems 1-3 and $*{}^{*}$$*{}^{*}$ by such a factor. Moroz in [78] and Imbach and Moroz in [56] accelerated their solution under an upper bound on condition numbers of all zeros of $p(x)$ instead. In spite of the appeal to the concept of condition number, used for the design and analysis of polynomial root-finding in [13, 28, 21], this is a little stronger restriction than bounding cluster sizes of Def. B.1, under which [92, 99] proved such a complexity bound.
Without using the algorithm of [78, Thm. 2], our current algorithms still support such a near optimal bit operation cost bound for Problem 1 with no restriction on an input polynomial $p$ in the case where $m^{2}=\tilde{O}(d)$, $1\leq m\leq d$ (see [108, Remark 20.1]) as well as under the Random Root Model of Sec. 1.5.
1.9 Implementation: State of the Art
The algorithm of [130] was implemented by Xavier Gourdon in [44] for the Magma and PARI/GP computer algebra systems, while the near optimal but more involved algorithm of [92, 99] has never been implemented.
Our current algorithms are much simpler for implementation than that of [92, 99] and, if properly implemented, can become user’s choice root-finders. Already incorporation in 2020 of some of our techniques into the previous best implementation of subdivision root-finder of [61] has enabled its noticeable acceleration (see [57]). The more advanced publicly available implementation in [60] “relies on novel ideas and techniques" from [108] (see [60, Sec. 1.2]). According to the results of extensive tests with standard test polynomials, the algorithm has performed dramatically faster than [61] or any other known variant of complex subdivision root-finder, has at least competed with user’s choice package MPSolve even for approximation of all $d$ zeros of a polynomial $p$ given with its coefficients, [^17] The complexity of a subdivision root-finder in a region is proportional to the number of roots in it, while MPSolve is about as fast and slow for all roots as for their fixed subset. MPSolve implements Ehrlich-Aberth’s root-finding iterations, which empirically converge to all $d$ roots very fast right from the start but only under empirical support and only under initialization that operates with the coefficients of $p$ [125]. and tended to supersede MPSolve significantly as the degree $d$ of $p$ grew large. Such a performance could only be a dream for the designers of subdivision root-finders before 2022, but there is a large room for further improvement because [60] still inherits some earlier algorithms from [22, 61], which are slower by order of magnitude than our current ones (see some specific directions next).
1.10 Tentative directions for further progress
Next we comment on some specific directions for significant (and perhaps dramatic) acceleration of the root-finders of [57, 60], based partly on our present work and partly on its expected further advance. Since our algorithms are near optimal, we only seek relaxation of the assumptions of the Random Root Model and empirical acceleration.
(A) Acceleration of e/i tests. They are the most used ingredient of our subdivision root-finders as well as those of [57, 60].
(i) In Part III we by-pass approximation of the Cauchy integral of (1.9) and only evaluate the polynomial $p$ at or near the unit circle $C(0,1)$ (see Sec. 11.4).
(ii) We achieve additional acceleration under the Random Root Model because we only need to perform root-lifting and then to evaluate $p$ at a small number of random points $x$. The cost of root-lifting is dominated in our formal study but can substantially slow down implementation. We conjecture that empirically the algorithm should work without lifting or with a small number of lifting steps.
(iii) E/i tests by means of approximation of the distance from a complex point to the roots seems to be competitive or maybe even superior empirically (see Remark 1.2).
(iv) We can try to exploit potential benefits of fast numerical approximation of Cauchy integrals with the Exponentially Convergent Adaptive Trapezoidal Quadrature (see [141]), efficiently implemented in Boost C++ libraries. This promises no significant acceleration for the worst case inputs but may be useful empirically.
(B) Acceleration of compression of a disc.
We accelerate the compression in [57, 60] at two stages:
(i) The algorithms of Secs. 20.2 and 20.3 are much faster than those of [57, 60], based on Cauchy sum, for computation of a point that lie in or near the convex hull of the root set in a fixed disc.
(ii) The algorithm of Sec. 12.1 runs significantly faster than those in [57, 60] for approximation of the smallest root radius from a given point on the complex plane, and those based on the DLG iterations (see Remark 1.2) are potentially competitive. Notice that we specified and analyzed extension from approximation of the distance to roots to approximation of the distance to the $m$ closest root in a disc containing precisely $m$ roots (see Sec. 12).
(C) Acceleration of real and near real root-finding. Our novel algorithm of Sec. 9 can be readily implemented and seems to be the first that efficiently solves this important problem.
(D) Acceleration of approximation of matrix eigenvalues. We adopt extension of root-finding to proposed in [73] because it implies record low upper estimates in the case of matrices and matrix polynomials of very large size, but this extension relies on the computation of the coefficients of the characteristic polynomial of an input matrix or matrix polynomial, which is prone to numerical stability problems and avoided in practice of numerical matrix computations. In Remark 7.1 of Sec. 7 we point out an alternative extension, which avoids such a notorious stage and is dramatically faster where the input matrix or matrix polynomial is data sparse.
1.11 Organization of the rest of the paper
In Parts II and III we devise and analyze our two alternative e/i tests. Deterministic e/i test of Part II is slower than that of Part III, but the root-finders of both parts use Las Vegas randomization to bound computational precision and bit-operation cost. These bounds are roughly the same in both parts, except thatin Part III we yield acceleration by a factor of $d$ under the Random Root Model. All the bit-operation complexity bounds of Parts II and III are is inferior by a factor of $b$ to that of Thm. 1.3.
We complete the proof of Thm. 1.3 in Part IV by means of our novel acceleration of subdivision iterations based on compression of a disc without losing its roots.
In Sec. A of the Appendix we comment on applications and implementation of root-finding. In its Secs. B and C we outline the fundamental results of Schönhage from [130] on polynomial factorization, its links to root-finding, and root-isolation and some relevant bit-complexity estimates. In Sec. D we discuss root-finding by means of functional iterations, namely, Newton’s, Weierstrass’s (aka Durand-Kerner’s), and Ehrlich’s (aka Aberth’s), and strengthen them with incorporation of Fast Multipole Algorithm (FMM).
1.12 Independence of Part III from Part II
In Parts II an III we solve Problems 0, 1, $*{}^{*}$, and $*{}^{*}$ in two different ways. In Part II we rely on approximation of a Cauchy integral associated with root-counting in a disc on the complex plane. In Part III we rely on lifting the zeros of $p$ to a fixed power $h=O(\log(d))$. In addition, in Part III, unlike Part II, we devise a faster root-finder under a random root model. We make our presentation in Part III independent of Part II; in particular we include again some definitions, figures, and theorems.
PART II: Classical subdivision root-finder with novel exclusion/inclusion (e/i) tests based on approximation of Cauchy integrals
[…]
2 Background: basic definitions
  • $S(c,\rho):=\{x:~{}|\Re(c-x)|\leq\rho$, $|\Im(c-x)|\leq\rho\},~{}D(c,\rho):=\{x:~{}|x-c|\leq\rho\}$,
    $C(c,\rho):=\{x:~{}|x-c|=\rho\}$, and $A(c,\rho,\rho^{\prime}):=\{x:~{}\rho\leq|x-c|\leq\rho^{\prime}\}$
    denote a square, a disc, a circle (circumference), and an annulus (ring) on the complex plane, respectively.
  • $\Delta(\mathbb{R})$, $X(\mathbb{R})$, $CH(\mathbb{R})$, and $\#(\mathbb{R})$ denote the diameter, the root set, the convex hull, and the index (root set’s cardinality) of a region $\mathbb{R}$ on the complex plane, respectively.
  • A disc $D(c,\rho)$, a circle $C(c,\rho)$, or a square $S(c,\rho)$ is said to be $\theta$-isolated for $\theta>1$ if $X(D(c,\rho))=X(D(c,\theta\rho))$, $X(C(c,\rho))=X(A(c,\rho/\theta,\rho/\theta))$, or $X(S(c,\rho))=X(S(c,\theta\rho))$, respectively.
  • The largest upper bound on such a value $\theta$ is said to be the isolation of the disc $D(c,\rho)$, the circle $C(c,\rho)$, or the square $S(c,\rho)$, respectively (see Fig. 3), and is denoted $i(D(c,\rho))$, $i(C(c,\rho))$, and $i(S(c,\rho))$, respectively.
  • $r_{1}(c,t)=|y_{1}-c|,\dots,r_{d}(c,t)=|y_{d}-c|$ in non-increasing order are the $d$ root radii, that is, the distances from a complex center $c$ to the zeros $y_{1}\dots,y_{d}$ of a $d$th degree polynomial $t(x)$. $r_{j}(c):=r_{j}(c,p)$, $r_{j}:=r_{j}(0)$ for $j=1,\dots,d$.
[…]
Observation 2.1.
Eqn. (1.4) maps the zeros $z_{j}$ of $p(x)$ into the zeros $y_{j}=\frac{z_{j}-c}{\rho}$ of $t(y)$, for $j=1,\dots,d$, and preserves the index #$\#(D(c,\rho))$ and the isolation $i(D(c,\rho))$.
$\rho$$\rho\theta$$\rho\theta^{2}$c********
Figure 4: The roots are marked by asterisks. The red circle $C(c,\rho\theta)$ and the disc $D(c,\rho\theta)$ have isolation $\theta$. The disc $D(c,\rho)$ has isolation $\theta^{2}$.
3 Cauchy root-counting in a disc
3.1 The power sums of the roots in a disc
Recall the following equations generalizing (1.9) (cf. [1]):
[…]
For $\mathcal{D}=D(0,1)$ and $\zeta$ of (1.10) we will approximate $s_{h}$ with finite sums
$s_{h,q}:=\frac{1}{q}\sum_{g=0}^{q-1}\zeta^{(h+1)g}~{}\frac{p^{\prime}(\zeta^{g})}{p(\zeta^{g})}~{}{\rm for}~{}h=0,1,\dots,q-1$ (3.2)
and recall that $s_{0}(D(0,1))=\#(D(0,1))$ (cf. (1.9)), while an e/i test is just the decision whether #$\#(D(0,1))>0$.
3.2 Cauchy sums in the unit disc as weighted power sums of the roots
Cor. 3.1 of the following lemma expresses the $h$th Cauchy sum $s_{h,q}$ as the sum of the $h$th powers of the roots $z_{j}$ with the weights $\frac{1}{1-x^{q}}$.
Lemma 3.1.
For a complex $z$, two integers $h\geq 0$ and $q>1$, and $\zeta$ of (1.10), it holds that
$\frac{1}{q}\sum_{g=0}^{q-1}\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=\frac{z^{h}}{1-z^{q}}.$ (3.3)
Proof.
First let $|z|<1$ and obtain
$\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=\frac{\zeta^{hg}}{1-\frac{z}{\zeta^{g}}}=\zeta^{hg}\sum_{u=0}^{\infty}\Big{(}\frac{z}{\zeta^{g}}\Big{)}^{u}=\sum_{u=0}^{\infty}\frac{z^{u}}{\zeta^{(u-h)g}}.$
The equation in the middle follows from Newman’s expansion
$\frac{1}{1-y}=\sum_{u=0}^{\infty}y^{u}~{}{\rm for}~{}y=\frac{z}{\zeta^{g}}.$
We can apply it because $|y|=|z|$ while $|z|<1$ by assumption.
Sum the above expressions in $g$, recall that
$\zeta^{u}\neq 1~{}{\rm and}~{}\sum_{s=0}^{q-1}\zeta^{u}=0~{}{\rm for}~{}0<u<q,~{}\zeta^{q}=1,$ (3.4)
and deduce that
$\frac{1}{q}\sum_{g=0}^{q-1}\frac{z^{u}}{\zeta^{(u-h)g}}=z^{u}$
for $u=h+ql$ and an integer $l$, while $\sum_{g=0}^{q-1}\frac{z^{u}}{\zeta^{(u-h)g}}=0$ for all other values $u$. Therefore
$\frac{1}{q}\sum_{g=0}^{q-1}\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=z^{h}\sum_{l=0}^{\infty}z^{ql}.$
Apply Newman’s expansion for $y=z^{q}$ and deduce (3.3) for $|z|<1$. By continuity extend (3.3) to the case of $|z|=1.$
Now let $|z|>1$. Then
$\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=-\frac{\zeta^{(h+1)g}}{z}~{}~{}\frac{1}{1-\frac{\zeta^{g}}{z}}=-\frac{\zeta^{(h+1)g}}{z}\sum_{u=0}^{\infty}\Big{(}\frac{\zeta^{g}}{z}\Big{)}^{u}=-\sum_{u=0}^{\infty}\frac{\zeta^{(u+h+1)g}}{z^{u+1}}.$
Sum these expressions in $g$, write $u:=ql-h-1$, apply (3.4), and obtain
$\frac{1}{q}\sum_{g=0}^{q-1}\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=-\sum_{l=1}^{\infty}\frac{1}{z^{ql-h}}=-z^{h-q}\sum_{l=0}^{\infty}\frac{1}{z^{ql}}.$
Apply Newman’s expansion for $y=z^{q}$ again and obtain
$\frac{1}{q}\sum_{g=0}^{q-1}\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=-\frac{z^{h-q}}{1-\frac{1}{z^{q}}}=\frac{z^{h}}{1-z^{q}}.$
Hence (3.3) holds in the case where $|z|>1$ as well. ∎
Sum equations (3.3) for $z=z_{j}$ and $j=1,\dots,d$, combine Eqns. 1.13 and (3.2), and obtain
Corollary 3.1.
For the zeros $z_{j}$ of $p(x)$ and all $h$, the Cauchy sums $s_{h,q}$ of (3.2) satisfy $s_{h,q}=\sum_{j=1}^{d}\frac{z_{j}^{h}}{1-z_{j}^{q}}~{}{\rm unless}~{}z_{j}^{q}=1~{}{\rm for~{}some}~{}j.$
3.3 Approximation of the power sums in the unit disc
Cor. 3.1 implies that the values $|s_{h,q-}s_{h}|$ decrease exponentially in $q-h$ if the unit disc $D(0,1)$ is isolated.
Theorem 3.1.
[130, Eqn. (12.10)]. [^20] Unlike the proof in [130], we rely on Cor. 3.1 of independent interest. Let the unit disc $D(0,1)$ be $\theta$-isolated and let $d_{\rm in}$ and $d_{\rm out}=d-d_{\rm in}$ denote the numbers of the roots lying in and outside the disc $D(0,1)$, respectively. Then
$|s_{h,q}-s_{h}|\leq\frac{d_{\rm in}\theta^{-h}+d_{\rm out}\theta^{h}}{\theta^{q}-1}\leq\frac{d\theta^{h}}{\theta^{q}-1}~{}{\rm for}~{}h=0,1,\dots,q-1.$ (3.5)
Proof.
Enumerate the roots $z_{j}$ so that $|z_{j}|<1$ if and only if $j\leq d_{\rm in}$; then Cor. 3.1 implies that
$s_{h,q}-s_{h}=s_{h,q,{\rm in}}+s_{h,q,{\rm out}}$ (3.6)
where
$s_{h,q,{\rm in}}=\sum_{j=1}^{d_{\rm in}}z_{j}^{h}\Big{(}\frac{1}{1-z_{j}^{q}}-1\Big{)}~{}{\rm and}~{}s_{h,q,{\rm out}}=\sum_{j=d_{\rm in}+1}^{d}\frac{z_{j}^{h}}{1-z_{j}^{q}}=\sum_{j=d-d_{\rm out}+1}^{d}\frac{z_{j}^{h}}{1-z_{j}^{q}}.$ (3.7)
Notice that $\frac{1}{1-z_{j}^{q}}-1=\frac{z^{q}}{1-z_{j}^{q}}$, and so $|z_{j}^{h}(\frac{1}{1-z_{j}^{q}}-1)|\leq\frac{\theta^{-h}}{\theta^{q}-1}$ for $z_{j}\in D(0,1)$, while $\Big{|}\frac{z_{j}^{h}}{1-z_{j}^{q}}\Big{|}\leq\frac{\theta^{h}}{\theta^{q}-1}$ for $z_{j}\notin D(0,1)$. Combine these bounds with Eqns. (3.6) and (3.7) and obtain (3.5). ∎
By choosing $q$ that exceeds $\log_{\theta}(2d+1)$ we ensure that $|s_{0}-s_{0,q}|<1/2$.
Example 3.1.
Let the unit circle $C(0,1)$ be 2-isolated. Then Thm. 3.1 implies that $|s_{0,q}-s_{0}|<1/2$ for $d\leq 1,000$ if $q=10$ and for $d\leq 1,000,000$ if $q=20$.
Algorithm 3.1.
INPUT: a black box polynomial $p$ of a degree $d$ and $\theta>1$.
INITIALIZATION: Compute the integer $q=\lfloor\log_{\theta}(4d+2)\rfloor>\log_{\theta}(2d+1)$.
COMPUTATIONS: Compute Cauchy sum $s_{0,q}$ and output an integer $\bar{s}_{0}$ closest to it.
Observation 3.1.
(i) Alg. 3.1 evaluates $p(x)$ at $q=\lfloor\log_{\theta}(4d+2)\rfloor$ points. (ii) It outputs $\bar{s}_{0}=\#(D(0,1))$ if the circle $C(0,1)$ is $\theta$-isolated for $\theta>1$. (iii) If the algorithm outputs $\bar{s}_{0}>0$, then #$\#(D(0,\theta))>0$.
Proof.
Thm. 3.1 immediately implies claim (ii) but also implies that #$\#(D(0,1))>0$ if $\bar{s}_{0}>0$ unless #$\#(A(0,1/\theta,\theta))>0$. In both cases #$\#(D(0,\theta))>0$. ∎
3.4 Extension to any disc
Map (1.4), preserving #$\#(D(0,1))$ (cf. Observation 2.1), enables us to extend definition (3.2) of Cauchy sum to any disc $D(c,\rho)$ as follows:
$s_{0,q}(p,c,\rho):=s_{0,q}(t,0,1):=\frac{\rho}{q}\sum_{g=0}^{q-1}\zeta^{g}~{}\frac{p^{\prime}(c+\rho\zeta^{g})}{p(c+\rho\zeta^{g})},$ (3.8)
for $\zeta$ of (1.10), that is, $s_{h,q}(p,c,\rho)$ is the Cauchy sums $s_{h,q}(t,0,1)$ of the zeros of the polynomial $t(y)=p(\frac{y-c}{\rho})$ in the unit disc $D(0,1)$.
Observation 3.2.
Given a polynomial $p(x)$, a complex $c$, a positive $\rho$, a positive integer $q$, and the $q$-th roots of unity, we can compute $s_{0,q}(p,c,\rho)$ by means of the evaluation of NIR$(\zeta^{g})$ for $g=0,1,\dots,q-1$, reduced to evaluation of $p$ at $2q$ points, and in addition performing $q+1$ divisions, $12q-$ 1 multiplications, and $12q-$ 1 additions.
[…]
Proposition 3.1.
[…]
By applying Thm. 3.1 we can estimate the errors $|s_{h,q}(t,0,1)-s_{h}(t,0,1)|$ of the approximation of the $h$th power sums $s_{h}(t,0,1)$ of the roots of $t(x)$ in the unit disc $D(0,1)$. Next we specify this for $h=1$.
Proposition 3.2.
(Cf. [60].) Let for $\theta>1$ a $\theta$-isolated disc $D(c,\rho)$ contain exactly $m$ roots. Fix a positive integer $q$ and define the Cauchy sum $s_{1,q}(p,c,\rho)$ of (3.8). Then
$|mc+\rho s_{1,q}(p,c,\rho)-s_{1}(p,c,\rho)|\leq\frac{\rho d\theta}{\theta^{q}-1}.$
3.5 The real part of the 0-th Cauchy sum
Clearly, $|\Re(s_{0,q})-s_{0}|\leq|s_{0,q}-s_{0}|$ because $s_{0}$ is an integer. Next we estimate the bound $|\Re(s_{0,q})-s_{0}|$ for the unit disc $D(0,1)$, point out some tentative directions to algorithmic gain for root-finding, but leave further elaboration as a challenge for formal and empirical investigation. [^21] We will not use the results of this subsection, but they can be of independent interest.
Recall that $s_{0,q}=\sum_{j=1}^{d}\frac{1}{1-z_{j}^{q}}$ unless $z_{j}^{q}=1$ for some $q$ (cf. Cor. 3.1), write $z:=z_{j}^{q}$ and estimate the contribution of a roots $z_{j}$ to $\Re(s_{0,q})$.
Proposition 3.3.
Let $z=\rho\cdot(\cos(\phi)+\sin(\phi)\bf i)$ for a non-negative $\rho$ and a real $\phi\in[0,2\pi)$. Then
$\Re\Big{(}\frac{1}{1-z}\Big{)}=\frac{1-\rho\cos(\phi)}{1+\rho^{2}-2\rho\cos(\phi)},$ (3.9)
$1-\Re\Big{(}\frac{1}{1-z}\Big{)}=\rho\frac{\rho-\cos(\phi)}{1+\rho^{2}-2\rho\cos(\phi)}.$ (3.10)
Proof.
Multiply both numerator and denominator of the fraction
$\frac{1}{1-z}=\frac{1}{1-\rho\cdot(\cos(\phi)+\sin(\phi)\bf i)}$
by $1-\rho\cdot(\cos(\phi)-\sin(\phi)\bf i)$, substitute $\cos^{2}(\phi)+\sin^{2}(\phi)=1$, and obtain (3.9). Then readily verify (3.10). ∎
Corollary 3.2.
Under the assumptions of Prop. 3.3 it holds that
$\Re\Big{(}\frac{1}{1-z}\Big{)}\geq 0~{}{\rm if}~{}\rho\leq 1,$ (3.11)
$\Re\Big{(}\frac{1}{1-z}\Big{)}=\frac{1}{2}~{}{\rm if}~{}\rho=1~{}{\rm and~{}if}~{}\phi\neq 0,$ (3.12)
$\Re\Big{(}\frac{1}{1-z}\Big{)}\geq\frac{1-0.8\rho}{1+\rho^{2}-1.6\rho}~{}{\rm if}~{}\rho\leq 5/4~{}{\rm and~{}if}~{}\cos(\phi)\leq 0.8,$ (3.13)
$\Re\Big{(}\frac{1}{1-z}\Big{)}=\frac{1}{1-\rho}~{}{\rm if}~{}\phi=0~{}{\rm and~{}if}~{}\rho\neq 1,$ (3.14)
$\frac{1}{1+\rho}\leq\Big{|}\Re\Big{(}\frac{1}{1-z}\Big{)}\Big{|}\leq\frac{1}{|1-\rho|},$ (3.15)
$\frac{\rho}{1+\rho}\leq\Big{|}1-\Re\Big{(}\frac{1}{1-z}\Big{)}\Big{|}\leq\frac{\rho}{|1-\rho|}.$ (3.16)
Proof.
Readily deduce (3.11), (3.12), and (3.14) from (3.9).
Write
$y:=\cos(\phi),~{}f(y):=\Re\Big{(}\frac{1}{1-z}\Big{)},~{}{\rm and}~{}g(y):=\frac{1}{f(y)}.$
Then deduce from (3.9) that $g(y)=2+\frac{\rho^{2}-1}{1-\rho y}$ Hence $g^{\prime}(y)=\frac{(\rho^{2}-1)\rho}{(1-\rho y)^{2}}$ is a monotone increasing or monotone decreasing function, singular if and only if $y=\frac{1}{\rho}$. Extend these properties to $f(y)$ and $1-f(y)$ if $\rho<1$. This implies (3.13), (3.15), and (3.16) provided that $\rho<1$ because
$f(1)=\frac{1}{1-\rho},~{}1-f(1)=\frac{\rho}{1-\rho},~{}f(-1)=\frac{1}{|1+\rho|},~{}1-f(-1)=\frac{\rho}{|1+\rho|},$
and $f(0.8)=\frac{1-0.8\rho}{1+\rho^{2}-1.6\rho}$.
The proof can be extended to the case where $\rho\geq 1$ except that the zero $y=\frac{1}{\rho}$ of $f(y)$ turns into a pole of $\frac{1}{f(y)}$. We complete the proof of the bounds (3.15) for all positive $\rho$ by observing that $f(\frac{1}{\rho})$ is undefined for $\rho=1$ and that otherwise $f(\frac{1}{\rho})=0$ and $11-f(\frac{1}{\rho})=$ 1. ∎
Next we outline but do not elaborate upon some potential links of Cor. 3.2 to root-counting. Let the unit disc $D(0,1)$ be $\theta$-isolated, combine bounds (3.15) and (3.16), and obtain
$|\Re(s_{0,q})-s_{0}|\leq\frac{d}{\theta^{q}-1}.$
The bound is similar to (3.5) for $h=0$ but is overly pessimistic: it is only reached where $\cos(\phi)=1$ for $z=z_{j}$ and all $j$.
Equation (3.14) implies numerically unstable dependence of the fraction $\Re(\frac{1}{1-z})$ on $\rho$ where $\cos(\phi)\approx 1\approx\rho$, but bound (3.13) shows that outside a rather small neighborhood of the extreme values of $\cos(\phi)$ and $\rho$ the fraction is smooth and approaches 1/2 as $\rho\rightarrow 1$ and/or $\cos(\phi)\rightarrow-1$.
Remark 3.1.
In view of bound (3.13) random rotation of the circle $C(0,1)$ is a heuristic recipe for countering instability in $z$ of the fraction $\Re(\frac{1}{1-z})$ because $\cos(\phi)>0.8$ for only one of four equally spaced points $\phi$ on the unit circle $C(0,1)$ and hence occurs with a probability less than 1/4 for a random choice of $\phi$ under the uniform probability distribution in the range $[0,2\pi)$. Moreover, random rotation of the circle can limit the adverse and numerically unstable impact of boundary roots, lying on or near the boundary circle $C(0,1)$, on an e/i test.
We conclude with a simple but surprising result on root-counting on an isolated circle. We do not use it in this paper.
Corollary 3.3.
Suppose that $p(1)\neq 0$ and that $p(x)$ has precisely $m$ roots in an annulus $A(0,\frac{1}{\theta},\theta)$ for a fixed $\theta>1$ and that all of them lie on the unit circle $C(0,1)$. Let $s^{\prime}$ and $s^{\prime\prime}$ denote the 00th Cauchy sums for the discs $D(0,\frac{1}{\theta})$ and $D(0,\theta)$, respectively, and write $s_{0,q}=s^{\prime\prime}-s^{\prime}$. Then #$(C(0,1))=2\Re(s_{0,q})$.
Proof.
Combine Cor. 3.1 for $h=0$ with equation (3.12). ∎
4 Root-counting and $\ell$-tests
4.1 Overview
We seek fast root-counters and $\ell$-tests based on computing Cauchy sum $s_{0,q}$ for a smaller integer $q$, but the output can deceive us if we decrease $q$ too much: the polynomial $p(x):=(1-wx^{d-m})(x^{m}-dx+w)$ has zero near the origin if $w\approx 0$, while NIR$(x)=p^{\prime}(x)/p(x)$ vanishes for $x$ equal to any $(d-1)$st root of unity.
Such deceptive e/i tests have been extremely rare guests in the extensive experiments in [57, 60] for $q$ of order $\log^{2}(d))$ and were completely excluded where the values $|s_{h,q}|$ were required to be small for $h=0,1,2$, but formal support for this empirical behavior is still a challenge.
In the next two subsections we formally support Las Vegas randomized root-counters and $\ell$-tests that evaluate polynomial $p$ at $O(m\log(d))$ points.
One can probably obtain similar results by computing the Cauchy sum $s_{0,q}$ for the polynomial $p(ax)$ at $O(m\log(d))$ random values $a\in C(0,1)$, but we leave formal support of this conjecture as a research challenge.
Recall that randomization in our root-finders is of Las Vegas type (see Sec. 1.7), while we apply randomization in Sec. 6 anyway – to bound computational precision in our root-finders. For the sake of completeness we also present a deterministic $\ell$-test in Sec. 4.4.
4.2 A basic randomized root-counter
Next we fix a parameter $\gamma>1$ and devise a root-counter for a disc $D(c,\rho)$ with a fixed center $c$ (wlog we let $c=0$) and a radius $\rho$ sampled at random in a fixed small range. Then we prove that under such a Las Vegas randomization the test fails with a probability at most $1/\gamma$.
Algorithm 4.1.
Basic randomized root-counter.
INPUT: $\gamma\geq 1$ and a $d$-th degree black box polynomial $p$ having at most $m$ zeros in the disc $D(0,\sqrt{2})$.
INITIALIZATION: Sample a random value $w$ in the range $[0.2,0.4]$ under the uniform probability distribution in that range and write $\rho:=2^{w}$ and $q=\lfloor 10m\gamma\log_{2}(4d+2)\rfloor$.
[…]
Proposition 4.1.
Alg. 4.1 queries evaluation oracle at $q=\lfloor 10m\gamma\log_{2}(4d+2)\rfloor$ locations and outputs the value $\bar{s}_{0}=\#(D(0,\rho))$ with a probability at least $1-1/\gamma$.
Proof.
Readily verify the bound on the number of query locations claimed in the proposition.
By assumption, the annulus $A(0,1,\sqrt{2})$ contains at most $m$ roots.
Hence at most $m$ root radii $r_{j}=2^{e_{j}}$, for $j=1,\dots,m^{\prime}\leq m$, lie in the range $[1,\sqrt{2}]$, that is, $0\leq e_{j}\leq 0.5$ for at most $m^{\prime}\leq m$ integers $j$.
Fix $m^{\prime}$ intervals, centered at $e_{j}$, for $j=1,\dots,m^{\prime}$, each of a length at most $1/(5m^{\prime}\gamma)$. Their overall length is at most $1/(5\gamma)$. Let $\mathbb{U}$ denote their union.
Sample a random $u$ in the range $[0.2,0.4]$ under the uniform probability distribution in that range and notice that Probability$(u\in\mathbb{U})\leq 1/\gamma$.
Hence with a probability at least $1-1/\gamma$ the circle $C(0,\rho)$ is $\theta$-isolated for $\theta=2^{\frac{1}{10m\gamma}}$, in which case Alg. 4.1 outputs $\bar{s}_{0}=s_{0}=\#(D(0,\rho))$ by virtue of claim (ii) of Prop. 3.1. ∎
4.3 Refining the bound on error probability
By reapplying Alg. 4.1 $v$ times we evaluate the polynomial $p$ at $v$ times more points, but we will prove that this decreases the Las Vegas error probability dramatically – below $1/\gamma^{v}$.
First let $\ell=1$ or $\ell=m$ and then again reduce the root radius approximation problem to the decision problem of $\ell$-test. Namely, narrow Alg. 4.1 to an $\ell$-test for a fixed $\ell$ in the range $1\leq\ell\leq m$, so that with a probability at least $1-1/\gamma$ this $\ell$-test outputs 0 if #$\#(D(0,1))<\ell$ and outputs 1 if #$\#(D(0,2^{0.5}))\geq\ell$. Moreover, output 1 is certified by virtue of claim (iii) of Prop. 3.1 if $\ell=1$; similarly the output 0 is certified if $\ell=m$.
Next extend the $\ell$-test by means of applying Alg. 4.1 for $v$ iid random variables $w$ in the range $0.2\leq w\leq 0.4$. Call this $\ell$-test Alg. $\ell$.
[…]
Recall Prop. 4.1 and then readily verify the following proposition.
Proposition 4.2.
(i) For two integers $1\leq\ell\leq m$ and $v\geq 1$ and a real $\gamma\geq 1$, Alg. $\ell$ evaluates $p(x)$ at $\lfloor 10m\gamma\log_{2}(4d+2)\rfloor v$ points; this evaluation involves more ops than the other stages of the algorithm, even in the cases of the Mandelbrot and sparse polynomials $p$ (cf. Observation 3.2). (ii) Output 1 of Alg. 4.1v,1 and the output 0 of Alg. 4.1v,$m$ are certified. (iii) Output 0 of Alg. 4.1v,1 and output 1 of Alg. 4.1v,$m$ are correct with a probability at least $1-1/\gamma^{v}$.
The proposition bounds the cost of $\ell$-tests for $\ell=1$ (e/i tests) and $\ell=m$. This is sufficient for our root-finding, but our next natural extension covers $\ell$-tests for any integer $\ell$ in the range $1<\ell<m$.
Assign the values 0 or 1 to the output integers $g$ as follows: compute the average of the $v$ outputs in $v$ independent applications of the test and output $g$ denoting the integer closest to that average; if the average is 1/2, write $g=0$ to break ties.
This extension of Alg. $\ell$ is only interesting for $1<\ell<m$; otherwise its cost/error estimates are inferior to those of Prop. 4.2.
Proposition 4.3.
For an integer $\ell$ such that $1<\ell<m\leq d$ and real $v\geq 1$ and $\gamma>1$, Algs. $\ell$ queries evaluation oracle at $\lfloor 10m\gamma\log_{2}(4d+2)\rfloor v$ locations, and its output is correct with a probability at least $1-(4/\gamma)^{v/2}$.
Proof.
Since $D(0,1)\subset D(0,\sqrt{2})$, we only consider the following three cases:
(i) #$\#(D(0,1))<\ell$ and #$\#(D(0,2^{0.5}))\geq\ell$,
(ii) #$\#(D(0,2^{0.5}))<\ell$, and then also #$\#(D(0,1))<\ell$,
(iii) #$\#(D(0,1))\geq\ell$, and then also #$\#(D(0,2^{0.5}))\geq\ell$.
In case (i) the output $g$ is always correct.
In case (ii) the output 0 is correct, while the output 1 is wrong and occurs if and only if more than $v/2$ tests go wrong, by outputting 1.
Likewise, in case (iii) the output 1 is correct, while the output 0 occurs if and only if more than $v/2$ tests go wrong, by outputting 0.
In both cases (ii) and (iii) we deal with the binomial distribution having failure probability $1/\gamma$, and so (cf. [25]) the probability $P$ of at least $v/2$ failures among $v$ tests is at most $\sum_{j=\bar{v}}^{v}\begin{pmatrix}v\\ j\end{pmatrix}T_{j}$ for $T_{j}=\gamma^{-j}(1-1/\gamma)^{v-j}\leq\gamma^{-j}$.
Therefore, $P\leq\sum_{j=0}^{v}\begin{pmatrix}v\\ j\end{pmatrix}=T_{\bar{v}}2^{v}\max_{j}T_{j}$. Substitute $\sum_{j=0}^{v}\begin{pmatrix}v\\ j\end{pmatrix}=2^{v}$ and $\bar{v}=\lceil v/2\rceil$ and obtain $P<2^{v}T_{\bar{v}}\leq 2^{v}/\gamma^{v/2}=(4/\gamma)^{v/2}$. ∎
4.4 Deterministic $\ell$-test
In view of our remark at the end of Sec. 4.1, randomization is only a nominal restriction for $\ell$-tests, but next we get rid of it, although at the price of increasing the numbers of points of evaluation of $p$ and bit operations involved in the tests by a factor of $m$.
[…]
Next we elaborate upon this test.
Without loss of generality (wlog) let $c=0$, $\theta\rho=1$, $C(c,\theta\rho)=C(0,1)$.
Algorithm 4.2.
Deterministic $\ell$-test.
INPUT: A black box polynomial $p$ of a degree $d$ and two positive integers $\ell$ and $m\geq\ell$ such that the unit disc $D(0,1)$ contains at most $m$ roots.
INITIALIZATION: Fix
$\sigma>1,~{}\theta=\sigma^{\frac{1}{4m+1}},~{}q=\lfloor\log_{\theta}(4d+2)\rfloor,~{}\rho_{i}=\frac{1}{\theta^{2i+1}},~{}i=0,1,\dots,2m.$ (4.1)
OUTPUT: 0 if #$\#(D(0,1/\sigma))<\ell$ or 1 if #$\#(D(0,1))\geq\ell$.
[…]
Output 1 if at least $m+1$ integers $\bar{s}_{0}$ output by the algorithm (among all $12m+$ 1 such integers) exceed $\ell-1$. Otherwise output 0. (For $\ell=1$ we can output 1 already on a single output $\bar{s}_{0}>0$; for $\ell=m$ we can output 0 already on a single output $\bar{s}_{0}<m$.)
Proposition 4.4.
Alg. 4.2 is a $\sigma$-soft $\ell$-test for $\sigma>1$ of our choice (see (4.1)).
Proof.
[…]
To complete the proof, combine this property with the relationships $\rho_{2m}=1/\theta^{4m+1}$ (cf. (4.1)), $\rho_{0}=1/\theta<1$, and #$\#(D(0,\rho_{i+1}))\leq\#(D(0,\rho_{i}))$, which hold because $D(0,\rho_{i+1})\subseteq(D(0,\rho_{i})$ for $i=0,1,\dots,2m-1$. ∎
Observation 4.1.
[Cf. Observation 3.2.] Alg. 4.2 evaluates polynomial $p(x)$ at $(2m+1)q$ points for $q=\lfloor\log_{\theta}(4d+2)\rfloor=\lfloor(4m+1)\log_{\sigma}(4d+2)\rfloor$, which dominates the cost of the remaining ops involved, even in the cases of the Mandelbrot and sparse input polynomials and which is in $O(m^{2}\log(d))$ if $\sigma-1$ exceeds a positive constant.
5 Root radius estimation based on $\ell$-test
In this section we assume that we are given an algorithm for $\ell$-test and go back from the decision problem of $\ell$-test to approximation of $r_{d-\ell+1}(c^{\prime})$, the $\ell$th smallest root radius from a fixed complex point $c^{\prime}$; we assume that $r_{d-\ell+1}(c^{\prime})$ lies in a fixed range $[a_{-},a_{+}]$. [^22] For our root-finders we only need to approximate $r_{d-\ell+1}(c^{\prime})$ for $\ell=1$ and $\ell=m$ and only with a constant bound on the relative error, but we will estimate $r_{d-\ell+1}(c^{\prime})$ within $1/2^{b}$ for any real $b$ and $\ell$, $1\leq\ell\leq m$, which is of independent interest for root-finding. Wlog let this be the range $[\epsilon,1]$ for a fixed small positive $\epsilon=1/2^{b}$. The bisection algorithm is too slow for us, and we accelerate it by applying it to approximate the exponent of a target value. Call that algorithm bisection of exponent or BoE and apply it to bracket the $\ell$th smallest root radius $r_{d-\ell+1}$ in a range $[\rho_{-},\rho_{+}]$ such that $0<\rho_{<}-\rho_{+}\leq\epsilon$ or $0<\rho_{+}<\beta\rho_{-}$ for two fixed tolerance values $\epsilon=1/2^{b}$ and $\beta=2^{\phi}>1$.
[…]
Since the test is soft, so is BoE as well: for a range $[b_{-},b_{+}]$ with midpoint $\widehat{b}=0.5(b_{+}+b_{-})$ regular rigid bisection continues the search in the range $[b_{-},\widehat{b}]$ or $[\widehat{b},b_{+}]$, but our soft BoE continues it in the range $[b_{-},(1+\alpha)\widehat{b}]$ or $[(1-\alpha)\widehat{b},b_{+}]$ for a fixed small positive $\alpha$. The approximations still converges linearly and just a little slower than the regular rigid bisection: the approximation error decreases by a factor of $\frac{2}{1+\alpha}$ rather than by twice. [^23] If we applied BoE to the range $[0,1]$, we would have to bisect the infinite range $[-\infty,0]$, but we deal with the narrower range $[-b,0]$ since we have narrowed the range $[0,1]$ to $[\epsilon,1]$.
Algorithm 5.1.
Bracketing the $\ell$-th smallest root radius based on $\ell$-test.
[…]
OUTPUT: A range $[\rho_{-},\rho_{+}]$ such that $[\rho_{-},\rho_{+}]=[0,\epsilon]$ or $\rho_{-}\leq r_{d-\ell+1}\leq\rho_{+}\leq\beta\rho_{-}$.
INITIALIZATION: Write $\bar{b}:=\log_{2}(\frac{\sigma}{\epsilon})$, $b_{-}:=-\bar{b}$, and $b_{+}:=0$.
[…]
$\rho_{-}=2^{b_{-}},~{}\rho_{+}=2^{b_{+}},~{}\rho:=2^{\widehat{b}},~{}{\rm and}~{}\widehat{b}:=\lceil(b_{-}+b_{+})/2\rceil.$ (5.1)
Write $b_{-}:=(1-\alpha)\widehat{b}$ if the $\ell$-test outputs 0; write $b_{+}:=(1+\alpha)\widehat{b}$ if the test outputs 1. Then update $\widehat{b}$ and $\rho$ according to (5.1).
Continue recursively: stop and output the current range $[\rho_{-},\rho_{+}]=[2^{b_{-}},2^{b_{+}}]$ if $b_{+}-b_{-}\leq\phi$, that is, if $\rho_{+}\leq\beta\rho_{-}$; otherwise update $\widehat{b}$ and $\rho$ and reapply the computations.
The disc $D(0,\epsilon)$ contains at least $\ell$ roots, that is, the range $[0,\epsilon]$ brackets the $\ell$th smallest root radius $r_{d-\ell+1}$ if the $\sigma$-soft $\ell$-test applied to the disc $D(0,\frac{\epsilon}{\sigma})$ outputs 1. Otherwise, the disc $D(0,\frac{\epsilon}{\sigma})$ contains less than $\ell$ roots, and so the root radius $r_{d-\ell+1}$ lies in the range $[\frac{\epsilon}{\sigma},1]$, in which we shall continue our search.
Readily verify that our policy of recursively updating the range $[b_{-},b_{+}]$ maintains this property throughout and that updating stops where $b_{+}-b_{-}\leq\phi$, that is, where $\rho_{+}\leq\beta\rho_{-}$. This implies correctness of Alg. 5.1.
Let us estimate the number $\mathbb{I}$ of its invocations of an $\ell$-test.
Proposition 5.1.
Alg. 5.1 invokes an $\ell$-test $\mathbb{I}$ times where $\mathbb{I}\leq\lceil\log_{\nu}(\frac{\bar{b}}{\phi})\rceil+1$ for $\nu=\frac{2}{1+\alpha}$, $\mathbb{I}=O(\log(b/\phi))$.
Proof.
The computations in Alg. 5.1 stop after the first application of an $\ell$-test if it outputs 1. It remains to count the subsequent $\ell$-tests if it outputs 0.
Each test decreases the length $\bar{b}=b_{+}-b_{-}$ of its input range $[b_{-},b_{+}]$ for the exponent to $0.5(1+\alpha)\bar{b}$. Hence $i$ applications change it into $\frac{(1+\alpha)^{i}}{2^{i}}\bar{b}$ for $i=1,2,\dots$. These values linearly converge to 0 as $i$ increases to the infinity, and the proposition follows. ∎
Remark 5.1.
In the case where we know that #$\#(D(0,\rho))\geq\ell$ for a fixed $\rho$ in the range $[\epsilon,1]$, we only need to refine the initial range $[\rho,1]$ for the root radius $r_{d-\ell+1}$.
Clearly, $\mathbb{I}$ applications of $\ell$-test increase the number of evaluation points for the ratio $\frac{p^{\prime}(x)}{p(x)}$ by at most a factor of $\mathbb{I}$, and if the $\ell$-test is randomized, then the error probability also grows by at most a factor of $\mathbb{I}$. Let us specify these estimates for Alg. 5.1.
Corollary 5.1.
Let precisely $m$ zeros of a black box polynomial $p(x)$ of a degree $d$ lie in the unit disc, that is, let #$\#(D(0,1))=m$, and let $i(D(0,1))\geq\theta$, for $\theta-1$ exceeding a positive constant. Given two positive tolerance values $\epsilon=1/2^{b}$ and $\phi$, an integer $\ell$, $1\leq\ell\leq m$, and a black box $\ell$-test, which evaluates polynomial $p(x)$ at $\mathcal{NR}$ points, Alg. 5.1 approximates the $\ell$-th smallest root radius $|z_{j}|$ within the relative error $2^{\phi}$ or determines that $|z_{j}|\leq\epsilon$, at the cost of evaluation of polynomial $p(x)$ at $\mathbb{I}\cdot\mathcal{NR}$ points for $\mathbb{I}=O(\log(\frac{b}{\phi}))$ estimated in Prop. 5.1.
6 Precision of computing in our root-finders
6.1 Doubling computational precision towards its a posteriori estimation
Given an algorithm $\mathbb{A}$ for black box polynomial computations and target tolerance TOL to output error, how should one choose computational precision? A simple practical recipe of [10, 14, 21] was the basis for highly successful performance root-finders of MPSolve: namely, first apply algorithm $\mathbb{A}$ with a lower precision and then recursively double it until an output error decreases below TOL, at which point it outputs precision within a factor of rwo from optimal, thus also supporting near optimal a posteriori bit operation complexity estimate for algorithm $\mathbb{A}$.
In the rest of this section we obtain a priori estimate by proving that precision of $O(b\log(d))$ bits is sufficient for correctness of our subdivision root-finders.
6.2 Reduction to approximation of $x$NIR$(x)$ within $\frac{3}{8}$
Next we estimate precision of computations in our deterministic and randomized e/i tests that supports approximation of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ required in our root-finders; we only need to consider these tests performed for $\theta$-isolated discs $D(c,\rho)$ with $\theta-1>0$ of order $\frac{1}{m}$ where we approximate NIR$(x)$ with the first order divided difference of (1.6).
Map (1.4) reduces such a test for $p(x)$ to the disc $D(0,1)$ for the polynomial $t(x)=p(\frac{x-c}{\rho})$ and decreases the error estimate by a factor of $\rho$, which is at most $2^{b}$ because we only apply our e/i tests to the disc of radius greater than $\epsilon=1/2^{b}$. By examining the impact of map (1.4) on the error and precision bounds we observe that the latter bound can increase by at most a factor of $b$. Thus we arrive at the claimed precision bound $O(b\log(d))$ in case of any disc $D(c,\rho)$ by proving bound $O(\log(d))$ in the case of the unit disc $D(0,1)$.
The output of an e/i test is an integer, and we only need to certify that the output errors of Algs. 4.2 and 4.1 are less than $\frac{1}{2}$.
Our e/i tests compute the $q$ values $\rho x\frac{p^{\prime}(c+\rho x)}{p(c+\rho x)}=x\frac{t^{\prime}(x)}{t(x)}$, for $x=\zeta^{g}$, $\zeta$ of (3.2), and $g=0,1,\dots,q-1$, sum these values, and divide the sum by $q$.
By slightly abusing notation write ${\rm NIR}(x):=\frac{t^{\prime}(x)}{t(x)}$ rather than ${\rm NIR}(x)=\frac{p^{\prime}(x)}{p(x)}$. According to straightforward error analysis (cf. [20, Appendix A of Ch.3] and the references therein) rounding errors introduced by these additions and division are dominated, and we can ignore them; then it is sufficient to compute $\zeta^{g}{\rm NIR}(\zeta^{g})$ within, say, $\frac{3}{8}$ for every $g$. Indeed, in that case the sum of $q$ error bounds is at most $\frac{3q}{8}$ and decreases below $\frac{1}{2}$ in division by $q$.
6.3 Precision of representation of $x$NIR$(x)$ within $\frac{1}{8}$
Recall Eqn. (1.13) and represent $x{\rm NIR}(x)$ for $|x|=1$ within $\frac{1}{8}$ by using a precision of $O(\log(d))$ bits.
Proposition 6.1.
Write $C:=C(c,\rho)$ and assume that $|x|=1$ and the circle $C$ is $\theta$-isolated for $\theta>1$. Then $|x{\rm NIR}(x)|\leq\frac{d}{1-1/\theta}=\frac{\theta d}{\theta-1}$.
Proof.
Eqn. (1.13) implies that $|{\rm NIR}(x)|=|\sum_{j=1}^{d}\frac{1}{x-z_{j}}|\leq\sum_{j=1}^{d}\frac{1}{|x-z_{j}|}$ where $|x-z_{j}|\geq(1-1/\theta)\rho$ for $|z_{j}|<1$, while $|x-z_{j}|\geq(\theta-1)\rho$ for $|z_{j}|>1$ since $i(C)\geq\theta$. Combine these bounds, write $m:=\#(D(c,\rho))$, and obtain $|x{\rm NIR}(x)|\leq\frac{m}{1-1/\theta}+\frac{d-m}{\theta-1}\leq\frac{d}{1-1/\theta}$ for $|x|=1$. ∎
Corollary 6.1.
One can represent $x{\rm NIR}(x)$ within $\frac{1}{8}$ for $|x|=1$ by using a precision of $3+\lceil\log_{2}(\frac{d}{1-1/\theta})\rceil$ bits, which is in $O(\log(d))$ provided that $\frac{1}{\theta-1}=d^{O(1)}$.
6.4 What is left to estimate?
Instead of $x{\rm NIR}(x)=\frac{xt^{\prime}(x)}{t(x)}$ we actually approximate $\frac{xt^{\prime}(y)}{t(x)}$ where $|y-x|\leq\delta$ and $t^{\prime}(y)$ is equal to the divided difference of (1.6). Thus, we shall increase the above error bound $\frac{1}{8}$ by adding to it the upper bounds $\frac{1}{8}$ on $\alpha:=|\frac{xt^{\prime}(x)}{t(x)}-\frac{xt^{\prime}(y)}{t(x)}|$ for $|x|=1$ and on the rounding error $\beta$ of computing $x\frac{t(x)-t(x-\delta)}{\delta t(x)}$. Next we estimate $\alpha$ in terms of $d$, $\theta$, and $\delta$ – by extending the proof of [73, Lemma 3.6], then readily ensure that $\alpha<\frac{1}{8}$ by choosing $\delta$ with a precision $\log_{2}(1/|\delta|)$ of order $\log(d)$, and finally estimate $\beta$. We simplify our estimates by dropping the factor $x$ where $|x|=1$.
6.5 Precision of approximation with divided difference
Recall the following lemma, which is [73, Fact 3.5].
Lemma 6.1.
For $t(x)=\prod_{j=1}^{d}(x-y_{j})$ and a non-negative integer $j\leq d$ it holds that
$t^{(j)}(x)=j!~{}t(x)\sum_{S_{j,d}}~{}\prod_{j\in S_{j,d}}\frac{1}{x-y_{j}}$
where the summation $\sum_{S_{j,d}}$ is over all subsets $S_{j,d}$ of the set $\{1,\dots,d\}$ having cardinality $j$.
Proposition 6.2.
Under the assumptions of Prop. 6.1 let $w=w(x):=\max_{j=1}\frac{1}{|x-y_{j}|}$. Then
$|t(x)|\alpha=|\frac{t(x)-t(x-\delta)}{\delta}-t^{\prime}(x)|\leq|t(x)\frac{(d~{}w)^{2}\delta}{1-\delta~{}d~{}w}|$.
Proof.
The claimed bound on $|\xi^{\prime}|$ follows from (1.13). It remains to apply the first five lines of the proof of [73, Lemma 3.6] with $f$, $\alpha$, $n$, and $\xi$ replaced by $t$, $-\delta$, $d$, and $w$, s respectively. Namely, first obtain from Taylor’s expansion that $t(x)-t(x-\delta)=\sum_{j=0}^{\infty}\frac{\delta^{j}}{j!}t^{(j)}(x)-t(x)$.
Substitute the expressions of Lemma 6.1 and obtain
$t(x)-t(x-\delta)=\delta t^{\prime}(x)+\sum_{j=2}^{\infty}\delta^{j}t(x)\sum_{S_{j,d}}~{}\prod_{j\in S_{j,d}}\frac{1}{x-y_{j}}.$
Combine this equation with the assumed bounds on $\frac{1}{|x-y_{j}|}$ and deduce that
$\Big{|}\frac{t(x)-t(x-\delta)}{\delta}-t^{\prime}(x)\Big{|}\leq\Big{|}\frac{f(x)}{\delta}\sum_{j=2}^{\infty}\delta^{j}w^{j}d^{j}\Big{|}\leq\Big{|}f(x)\frac{(d~{}w)^{2}~{}\delta}{1-\delta~{}d~{}w}\Big{|}.$
Recall that in our estimates for $\alpha$ and $\delta$ we can apply Prop. 6.1 for $w\leq\frac{\theta}{\theta-1}$ and obtain
Corollary 6.2.
Under the assumptions of Prop. 6.1, let $w\leq\frac{\theta}{\theta-1}$. Then
$\alpha<1/8~{}{\rm for}~{}|\delta|\leq\frac{(\theta-1)^{2}}{10d^{2}\theta^{2}}.$
6.6 The impact of rounding errors
To approximate $|x{\rm NIR}_{c,\rho}(x)|$ within $\frac{3}{8}$ it remains to approximate $\beta$ within $\frac{1}{8}$.
Proposition 6.3.
Suppose that a black box oracle computes for us the values $t(y)$ within a relative error bounds $\nabla(y)$ for $y=x$, $y=x-\delta$, and $\delta$ of Cor. 6.2. Then we can ensure that $|\beta|<\frac{1}{8}$ by choosing a proper $\nabla(y)$ of order $1/d^{O(1)}$, represented with a precision of $\log(\frac{1}{\nabla(y)})=O(\log(d))$ bits.
Proof.
Write $\beta\delta=\frac{t(x-\delta)(1+\nabla(x-\delta))}{t(x)(1+\nabla(x))}-\frac{t(x-\delta)}{t(x)}=\frac{t(x-\delta)}{t(x)}~{}\frac{\nabla(x-\delta)-\nabla(x)}{1+\nabla(x)}$.
Taylor-Lagrange’s formula implies that $\frac{t(x-\delta)}{t(x)}=1+\frac{t(x-\delta)-t(x)}{t(x)}=1+\delta\frac{t^{\prime}(u)}{t(x)}$, for $u\in[x-\delta,x]$. Hence $|\frac{t(x-\delta)}{t(x)}|\leq 1+|(1+\alpha^{\prime})\delta|$ for $\alpha^{\prime}=\frac{t^{\prime}(u)-t(x)}{t(x)}$.
By extending Cor. 6.2 obtain $|\alpha^{\prime}|\leq\frac{1}{8}$.
Hence $|\frac{t(x-\delta)}{t(x)}|\leq w:=1+\frac{9|\delta|}{8}$, while $|\beta\delta|\leq w\frac{|\nabla(x)|+|\nabla(x-\delta)|}{1-|\nabla(x)|}$.
Now choose $\nabla(y)$ such that $|\nabla(y)|\leq\frac{|\delta|}{17w}$ for $y=x$ and $y=x-\delta$, and so $w(|\nabla(x)|+|\nabla(x-\delta)|)\leq\frac{2}{17\delta}$. Then verify that $\log(\frac{1}{|\nabla(y)|})=O(\log(d))$, $1-|\nabla(x)|\geq 1-\frac{|\delta|}{17w}>\frac{16}{17}$, and $|\delta\beta|<\frac{|\delta|}{8}$. ∎
By combining Cors. 6.1 and 6.2 and Prop. 6.3 obtain
Corollary 6.3.
Suppose that a black box oracle can output approximations of a $d$-th degree polynomial $t(x)$ with any relative error in $1/d^{O(1)}$ for $|x|\leq 1$ and let the unit disc be $\theta$-isolated for $t(x)$ and for $\frac{1}{\theta-1}=d^{O(1)}$. Then we can perform e/i test for $t(x)$ on the unit disc by using a precision of $O(b\log(d))$ bits.
7 Approximation of matrix eigenvalues
Given a $d\times d$ matrix $M$ that has characteristic polynomial $p(x)=\det(xI-M)$, our root-finders can approximate all $m$ its zeros (that is, the eigenvalues of $M$) lying in a fixed disc $D$ on the complex plane isolated from external eigenvalues.
Recall Gershgorin’s bound for any eigenvalue $\lambda$ of a matrix $M=(a_{i,j})_{i,j=1}^{d}$ [134, Thm. 1.3.2]: $|\lambda-a_{i,i}|\leq\sum_{j\neq i}|a_{i,j}|$ for some $i$, $1\leq i\leq d$.
Hence, clearly, $|\lambda|\leq||M||_{F}$.
Combine this bound with the precision estimates of the previous section for our root-finders [^24] Practically, one can combine our root-finders with the recipe of doubling the precision of computing of [10, 14, 21]. and with the following result of Storjohann [133] (cf. Remark 7.1):
Theorem 7.1.
One can evaluate the determinant of a $d\times d$ integer matrix $A$ by using an expected number of $O(d^{\omega}\log^{2}(d)\log(|A||_{F}))$ bit operations.
This implies a record bit operation complexity bound $\tilde{O}(m^{2}\tilde{b}^{2}d^{\omega})$ (cf. (1.7)) for approximation within $\epsilon$ of all $m$ eigenvalues of $M$ lying in the disc $D$ provided that $\tilde{b}=\log_{2}(||M||_{F}/\epsilon)$.
Remark 7.1.
Our root-finders are essentially reduced to evaluation of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ at sufficiently many points $x$. Hence we can rely on the following expressions for approximation of the eigenvalues of a matrix or a matrix polynomial:
$\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}((xI-M)^{-1})}$
where $M$ is a matrix and $p(x)=\det(xI-M)$ (cf. (1.13)) and
$\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}(M^{-1}(x)M^{\prime}(x))}$
where $M(x)$ is a matrix polynomial and $p(x)=\det(M(x))$ [19, Eqn. (5)]. [^25] These expressions have been obtained based on Jacobi’s formula ${\rm d}(\ln(\det(A)))={\rm trace}(A^{-1}{\rm d}A),$ which links differentials of a nonsingular matrix $A$ and of its determinant. Based on the above expressions one can reduce approximation of eigenvalues of a matrix or a matrix polynomial to its inversion. This computation tends to be more numerically stable than computing determinants and is significantly accelerated in the highly important case where where an input matrix or matrix polynomial can be inverted fast, e.g., is data sparse.
8 Bit operation cost of the classical subdivision root-finding with novel e/i tests based on approximation of a Cauchy integral
8.1 Our goal and a basic theorem
Next we estimate the bit operation complexity of our root-finders by combining them with fast multipoint evaluation of a polynomial by means of the algorithm of Fiduccia [40] and Moenck and Borodin [81] (see [18, Sec. 4.5] for historical account), non-trivially extended in [78] to support the following theorem:
Theorem 8.1.
Given a positive $b$, the coefficients of a polynomial $t(x):=\sum_{i=0}^{d}t_{i}x^{i}$ such that $||t(x)||_{1}=\sum_{i=0}^{d}|t_{i}|\leq 2^{\tau}$, and $q$ complex points $x_{1},\dots,x_{q}$ in the unit disc $D(0,1)$, one can approximate the values $t(x_{1}),\dots,t(x_{q})$ within $1/2^{b}$ by using $\tilde{O}((d+q)(b+\tau))$ bit operations.
Proof.
For $q=d$ this is [78, Thm. 2], based on using novel data structure for hyperbolic approximation of $p$, proposed and analyzed by Moroz in [78]. Trivially extend this theorem to the cases where $q>d$ (by considering $t(x)$ a polynomial of degree $q$ with $q-d$ leading coefficients 0) and where $q<d$ (by removing $d-q$ points $x_{i}$, with no increase of the complexity bound). ∎
We only apply this theorem for $\tau=0$, in which case its bit operation bound turns into
$\mathbb{B}_{\rm Moroz}=\tilde{O}((d+q)b)$ (8.1)
For comparison, the previous State of the Art was represented by the bound
$\mathbb{B}_{\rm Kirr}=\tilde{O}((d+q)(b+q)),$ (8.2)
due to [65, Thm. 3.9, Alg. 5.3, and Appendix A.3] by Peter Kirrinnis, which relied on combining his and Schönhage’s study of the bit operation cost of basic polynomial computations (cf. [130]) with the algorithm of [40, 81]. [^26] Precision of computing required in the latter algorithm grows fast as $d$ increases; e.g., the IEEE standard double precision of the customary numerical computations is not enough to support that algorithm already for $d>50$; for $b=O(\log(d))$ one can fix this deficiency by applying the Fast Multipole Method of Greengard and Rokhlin [48, 27, 101, 102] to the equivalent task of multiplication of Vandermonde matrix by a vector. One can, however, support bounds of Eqns. (8.1) and (8.2) on the overall bit operation complexity by computing with multiple precision for larger values $b$. Moroz obtained [78, Thm. 2] by applying bound (8.2) to his lower degree polynomials approximating $p(x)$.
The algorithms of [40, 81, 65, 78] extensively involve the coefficients of $p(x)$.
8.2 Reduction of our root-finders to multipoint polynomial approximation
Given a complex $c$, a pair of positive $b$ and $\rho$, and $d+1$ coefficients of a polynomial $p=p(x)$ of (1.1) such that the disc $D(c,\rho)$ is isolated and contains precisely $m$ zeros of $p$, we approximate all these zeros within $R/2^{b}$ for $R=|c|+\rho$ by applying subdivision root-finding iterations with deterministic or Las Vegas randomized e/i tests of Sec. 4. Then every subdivision step is reduced to application of e/i tests to $\bar{m}=O(m)$ discs $D(c_{\lambda},\rho_{\lambda})$ for $\rho_{\lambda}\geq R/2^{b}$ and all $\lambda$.
Such a test amounts to approximation of $p(x)$ at $O(q\log(d))$ equally spaced points $x$ on each of $v$ circles sharing the center $c_{\lambda}$ and having radii in the range $[\rho_{\lambda}/\sigma,\rho_{\lambda}\sigma]$ for $1<\sigma<\sqrt{2}$. Here $v=2m+1$ and $q=m^{2}$ or $v=O(1)$ and $q=m$ in our e/i tests of Sec. 4, while $q=1$ under the random root model.
We only use Thm. 8.1 for $\tau=0$ to cover the case where a polynomial $t(x)$ with $||t(x)||_{1}\leq 1$ is evaluated at points $x_{j}\in D(0,1)$. To ensure these assumptions we first scale the variable $x\mapsto Rx$ to map the discs $D(c_{\lambda},\rho_{\lambda}\sigma)\mapsto D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})\subseteq D(0,1)$, for $\bar{c}_{\lambda}=c_{\lambda}/R$, $\bar{\rho}_{\lambda}=\rho_{\lambda}/R\geq 1/2^{b}$, and all $\lambda$.
Then we write $t(x):=p(Rx)/\psi$ for $\psi=O(R^{d}||p(x)||_{1})$ such that $||t(x)||_{1}=1$ and approximate $t(x)$ within $1/2^{\bar{b}}$ or equivalently within the relative error bound $1/(|t(x)|2^{\bar{b}})$ at $\bar{q}:=qv$ points $x$ in $\bar{m}$ e/i tests. [^27] Scaling $p$ and $x$ above is done once and for all e/i tests; to compensate for this we only need to change our bound on the output precision $b$ by $\lceil|\log_{2}R|\rceil$ bits.
Due to Cor. 6.3 it is sufficient to ensure the bound $1/(|t(x)|2^{\bar{b}})=1/d^{O(1)}$; we achieve this by choosing any $\bar{b}\geq\log(\frac{1}{|t(x)|})+O(\log(d))$.
In our e/i tests we only need to approximate $t(x)$ at the points $x$ lying on $\theta_{\lambda}$-isolated circles where $\frac{1}{\theta_{\lambda}-1}=O(m)$. Map the disc $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ into the unit disc $D(0,1)$. This does not change isolation of the boundary circle and may increase the norm $||t(x)||_{1}$ by at most factor of $2^{b+d}$ because $|\bar{c}_{\lambda}|\leq 1$ and $\bar{\rho}_{\lambda}\geq 1/2^{b}$. Now substitute $||t(x)||_{1}\leq 2^{b+d}$ and $\log(\frac{||t(x)||_{1}}{|t(x)|})=O(d\log(m))$ (cf. [130, Eqn.(9.4)]) [^28] This bound is readily deduced from the following non-trivial result [130, Thm. 4.2]: Let $t(x)=f\prod_{j=1}^{m}(x-z_{j})\prod_{j=m+1}^{d}(\frac{x}{z_{j}}-1)$, $||t(x)||_{1}=1$, $|z_{j}|<1$ for $j\leq m$, and $|z_{j}|>1$ for $j>m$. Then $|f|\geq 1/2^{d}$. and obtain that $\bar{b}=\log(\frac{1}{|t(x)|})+O(\log(d))=O(d\log(m)+b)$.
Remark 8.1.
For $m=d$ the above bound on $\log(\frac{||t(x)||_{1}}{|t(x)|})$ of [130, Eqn.(9.4)]) is sharp up to a constant factor and is reached for the polynomial $t(x)=(x+1-\frac{1}{m})^{d}/(2-\frac{1}{m})^{d}$, $\rho_{\lambda}=1$, and $x=-1$. For $m<d$ [130, Thm. 4.5] implies a little stronger bounds, namely, $\log(\frac{||t(x)||_{1}}{|t(x)|})=O(d+m\log(m))$ and hence $\bar{b}=O(m\log(m)+d+b)$ because all discs $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ lie in $\theta$-isolated unit disc $D(0,1)$ for $\theta-1$ exceeding a positive constant.
8.3 The bit operation complexity of our e/i tests and root-finders
Apply Thm. 8.1 with $q$ and $b$ replaced by $\bar{q}=O(m^{2})$ and $\bar{b}=O(b+d+m\log(m))$, respectively, [^29] Under the Random Root Model of Sec. 1.5 we can whp decrease this bound by a factor of $m$ and then extend this decrease to the related subsequent estimates. and obtain that the bit operation cost of our $O(m)$ e/i tests of Sec. 4 at any fixed subdivision step is in $\tilde{O}((b+d+m)m^{2})$ for $1\leq m\leq d$. This implies the upper estimate
$\mathbb{B}_{\rm roots}=\tilde{O}((b+d)m^{2}b)$ (8.3)
for the overall bit operation cost of our $O(b)$ subdivision steps, which includes the time-complexity of the evaluation oracle. Based on [65, Alg. 5.3] and Eqn. (8.2) obtain the following (a little larger but still reasonably small) upper bound on the bit operation cost of our subdivision root-finders,
$\mathbb{B}_{\rm roots}^{\prime}=\tilde{O}((m^{2}+d+b)(m^{2}+d)b).$ (8.4)
8.4 Higher bit operation complexity of general polynomial root-finding based on map (1.4) and root-squaring
Our bit operation complexity estimates for root-finding rely on restriction of computations to approximation of NIR without involving coefficients of $p$. Knowing the coefficients of the polynomial $t(x)=p(\frac{x-c}{\rho})$ one can devise very fast e/i test for the disc $D(c,\rho)$ by applying Pellet’s theorem to $t_{h}(x)$ for $h$ of order $\log(d)$, whose zero are the $h$th powers of the zeros of $t(x)$ [22, 23].
The paper [23] applies subdivision iterations of [144, 50, 124, 97] with e/i tests performed by means of Pellet’s theorem, which involves all coefficients of an input polynomial, thus losing the benefits where, say, $p(x)$ is the Mandelbrot polynomial. The tests must be applied to the discs covering all suspect squares, that is, up to $4m$ squares per subdivision iteration.
A subdivision algorithm of [23] first maps the covering disc into the unit disc $D(0,1)$ (cf. (1.4)), thus losing the benefits where $p(x)$ is the sum of a small number of shifted monomials. Then the algorithm performs order of $\log(\log(d))$ root-squaring steps (one can apply order of $\log(d)$ root-lifting steps instead but within the same bit operation complexity estimates), and finally applies Pellet’s theorem to the resulting polynomial $t_{c,\rho}(x)=\sum_{i=0}^{d}t_{i}x^{i}$.
Transition from $p(x)$ to $t_{c,\rho}(x)$, however, involves scaling of the variable $x$ by factors ranging from $R/2^{b}$ to $R=\max_{j-1}^{d}|z_{j}|$ (besides root-squaring and shift of the variable). Hence for every $i$, $i=0,1,\dots,d$, the coefficient $t_{i}$ must be computed within $1/2^{b-i}$ to approximate the zeros of $p$ within $1/2^{b}$ (see Observation B.2).
This involves at least $bg$ bits for $g=1,\dots,d$, that is, at least $0.5(d-1)db$ bits and $0.25(d-1)db$ bit-operations overall for a single e/i test.
[…]
The bit-count above and the bit operation cost bounds for the root-finders of [92, 99, 23] decrease by a factor of $d$ where the minimal pairwise distance between the zeros of $p$ exceeds a positive constant, but the ratio of the bounds of [23] and [92, 99] remains at least $m$, where $m$ can be as large as $d$.
Similar argument shows that even representation of the leading coefficients $t_{d}$ of $t(x)$ alone after scaling for a singe e/i test involves odder of $bd$ bits, which is amplified to $d^{2}b$ for $t_{d}^{d}$ in every e/i test (see the end of Sec. B.4) and up to $d^{2}bm$ per a subdivision iteration with $m$ e/i tests.
It may be interesting that a shift by $c$ such that $|c|\leq 2$ is much less costly than scaling:
Theorem 8.2.
Bit operation cost of the shift of the variable, [132, Lemma 2.3] (cf. [65, Lemma 3.6]). Given a positive $b$, the coefficients of a polynomial $p(x):=\sum_{i=0}^{d}p_{i}x^{i}$ such that $||p(x)||_{1}\leq 1$, and a complex $c$ such that $|c|\leq 2$, one can approximate within $1/2^{b}$ the coefficients $t_{0},t_{1},\dots,t_{d}$ of the polynomial $t(x):=\sum_{i=0}^{d}t_{i}x^{i}:=p(x-c)$ by performing $O(\mu((b+d)d))$ bit operations for $\mu(s)$ of (1.8).
We can relax the restriction $|c|\leq 2$ above by means of scaling the variable $x\mapsto xc$, but as we showed above, this would greatly increase the bound on the overall number of bit operations involved.
9 Real and near real root-finding
9.1 Background, outline, and organization of our presentation
The algorithm of [67] specializes subdivision of [23] to real root-finding for univariate polynomial represented in monomial basis. Currently this is user’s choice algorithm, although we have accelerated it in [59] by narrowing the region for the search of the real roots, based on application of Newton’s polygon – by following [130] and [88, Sec.4] (cf. [89, Sec. 5]).
Unlike the customary real root-finders, next we approximate both real and near real roots, lying on and near a fixed real line segment. This covers approximation of real zeros of a polynomial given by its values or coefficients that are only available numerically with rounding errors.
Taylor’s shifts with scaling (1.4) readily reduce root-finding and root-counting for $p(x)$ in an isolated line segment to the case of the real line segment $[-1,1]$. Zhukovsky’s map of the next subsection further reduces the task to root-finding and root-counting for an auxiliary polynomial $s(z)$ on isolated unit circle $C(0,1)$ (see Remark 9.1).
If the polynomial $s(z)$ has the same root set on the unit circle $C(0,1)$ and in the annulus $A(0,\theta_{-},\theta_{+})$, for $0<\theta_{-}<1<\theta_{+}$, and if its the boundary circles are isolated, then we obtain an efficient root-counter, by applying our Alg. 3.1 to the isolated discs $D(0,\theta_{-})$ and $D(0,\theta_{+})$.
We can readily extend our subdivision algorithms to the the case of an input annulus rather than a disc. For a fixed error bound $\epsilon>0$, its complexity is dominated by the cost of the evaluation of NIR$(x)$ at $O(w\log(1/\epsilon))$ points for $w$ denoting the number of real zeros of $p(x)$ in a fixed range; realistically this number tends to be much smaller than $d$ (cf. [63, 37]).
We organize our presentation in this section as follows. In Sec. 9.2 we specify the two-way map, based on Zhukovsky’s function, from the real segment $[-1,1]$ to the unit circle $C(0,1)$ and backward. The map $[-1,1]\Longleftrightarrow C(0,1)$ can be extended to the neighborhoods of the segment and the circle (cf. Remark 9.1). We can readily compute the number $w$ of roots on the isolated unit circle $C(0,1)$ and can deflate a factor $f$ of $p$ with root set made up of the $w$ zeros of $p$ in $C(0,1)$ by extending deflation algorithms of [130] from a disc to an annulus. In Secs. 9.3 and 9.4 we extend these computations to the segment $[-1,1]$ and its neighborhood and consequently to approximation of both real and near real roots.
9.2 Zhukovsky’s function
For back and forth transition between the segment and the circle we can apply the two-to-one Zhukovsky’s function $z=J(x)$ and its one-to-two inverse, for complex variables $x$ and $z$. It maps the circle $C(0,1)$ to the segment $[-1,1]$ and vice versa:
$x=J(z):=\frac{1}{2}\Big{(}z+\frac{1}{z}\Big{)};~{}z=J^{-1}(x):=x\pm\sqrt{x^{2}-1}.$ (9.1)
9.3 A real root-finder
Algorithm 9.1.
Root-finding on a line segment.
Input:
A polynomial $p=p(x)$ of (1.1).
Output:
The number $w$ of its zeros on the unit segment $[-1,1]$ and approximations to all these zeros.
Computations:
  1. 1.
    Compute the values $v_{h}=p(\Re(\zeta_{2d}^{h}))$ of the polynomial $p$ at the Chebyshev points $\Re(\zeta_{2d}^{h})=\cos(\frac{\pi h}{2d})$ for $h=0,1,\dots,2d-1$ and $\zeta_{2d}$ of (1.10).
  2. 2.
    Interpolate to the polynomial $s(z)$ of degree $2d$ such that
    $s(z)=z^{d}p(x)~{}{\rm for}~{}x=\frac{1}{2}(z+z^{-1})$ (9.2)
    from its values
    $s(\zeta_{2d}^{h})=(-1)^{h}v_{h}~{}{\rm for}~{}h=0,1,\dots,2d-1$
    by means of Inverse DFT. [Recall that $\zeta_{2d}^{dh}=(-1)^{h}$.]
  3. 3.
    Approximate a factor (e.g., the monic factor) $g=g(z)$ of $p$ whose root set is made up of all roots of the polynomial $s(z)$ that lie on the unit circle $C(0,1)$. [The zeros of $s(z)$ appear in pairs: $s(z)=0$ if $s(z)=0$ by virtue of (9.2). Furthermore, if $p(1)=0$ and/or $p(-1)=0$, then 1 and/or $-1$ are also the zeros of $g(z)$ with double multiplicity.] Approximate a factor (e.g., the monic factor) $g=g(z)$ of $p$ whose root set is made up of all zeros of the polynomial $s(z)$ that lie on the unit circle $C(0,1)$. At this stage extend deflation algorithms of [130] to the the case of an input annulus about the unit circle $C(0,1)$. Output $w=0.5\deg(g)$.
  4. 4.
    By applying MPSolve, subdivision, or another root-finder approximate all $2w$ zeros of the polynomial $g(z)$. Let $y_{1},\dots,y_{w}$ denote the first $w$ of them in the order of increasing their arguments from 0.
  5. 5.
    Compute and output the values $z_{j}=\frac{1}{2}(y_{j}+\frac{1}{y_{j}})$ for $j=1,\dots,w$; these are the $w$ zeros of $p$ lying in the segment $[-1,1]$.
Correctness of this algorithm follows from (9.1) and (9.2).
Its Stage 2 is DFT. Its Stage 1 of the evaluation at Chebyshev’s points can be performed by means of the algorithm of [43] or Discrete Cosine Transform (cf. [98, Sec. 3.11 and the notes to it in Sec. 3.12]). In both cases the computation is simplified if $d$ is a power of 2; we ensure this by replacing $p$ with $x^{u}p$ for the minimal non-negative integer $u$ such that $d+u$ is a power of 2.
The complexity at stage 3 is dominated by the cost of evaluation of $p(x)$ at $O(w\log(1/\epsilon))$ points for $\epsilon$ bounding the output errors.
The converse transition from the unit circle $C(0,1)$ to the unit segment $[-1,1]$ enables us to simplify stages 4 and 5 of Alg. 9.1 by moving from $g(z)$ to a polynomial of degree $w$ whose all zeros lie in the segment $[-1,1]$. Then again we achieve this by means of evaluation and interpolation:
Algorithm 9.2.
Transition to the unit segment.
  1. 1.
    Compute the values $u_{h}=g(\zeta_{2K}^{h})$ of the polynomial $g(z)$ at the $2K$-th roots of unity $\zeta_{2K}^{h}$ for $h=0,1,\dots,K-1$ and $K>w$.
  2. 2.
    Interpolate to the polynomial $f(x)$ of degree at most $w$ from its values $f(\Re(\zeta_{2K}^{h}))=(-1)^{h}u_{h}$ at the Chebyshev points $\Re(\zeta_{2K}^{h})=\cos(\frac{\pi h}{2K})$, for $h=0,1,\dots,K-1$. [Recall that $\zeta_{2K}^{Kh}=(-1)^{h}$.]
  3. 3.
    Approximate the $w$ zeros of the polynomial $f=f(x)$ by applying MPSolve, real subdivision, e.g., those of [67], or another real polynomial root-finder.
We propose to perform steps 1 and 2 above by means of forward DFT and inverse Cosine Transforms, which we would apply to the polynomial $z^{v}g(z)$, replacing $g(x)$, where $v$ is the minimal non-negative integer such that $w+v$ is a power 2.
Remark 9.1.
Represent complex numbers as $z:=u+v\bf i$. Then Zhukovsky’s map transforms a circle $C(0,\rho)$ for $\rho\neq 1$ into the ellipse $E(0,\rho)$ whose points $(u,v)$ satisfy the following equation,
$\frac{u^{2}}{s^{2}}+\frac{v^{2}}{t^{2}}=1~{}{\rm for}~{}s=\frac{1}{2}\Big{(}\rho+\frac{1}{\rho}\Big{)},~{}~{}t=\frac{1}{2}\Big{(}\rho-\frac{1}{\rho}\Big{)}.$
Consequently it transforms the annulus $A(0,\frac{1}{\theta},\theta)$ into the domain bounded by the ellipses $E(0,\frac{1}{\theta})$ and $E(0,\theta)$, and so the circle $C(0,1)$ is $\theta$-isolated if and only if no roots of $p$ lie in the latter domain.
9.4 Real root-counting
We can simplify Alg. 9.1 and use only evaluations if we restrict our task to counting the roots that lie on or near a line segment.
Here is a high level description of this algorithm, where we do not specify the parameters $\theta$ and $q$ and assume that the input includes the polynomial $s(z)$.
Algorithm 9.3.
Root-counting on a line segment.
A real $\theta>1$ and the polynomial $s(z)$ of equation (9.2) such that the unit circle $C(0,1)$ is t $\eta$-isolated for $\eta\geq\theta^{2}$, that is, the annulus $A(0,\frac{1}{\theta^{2}},\theta^{2})$ contains no zeros of $s(z)$ except possibly some on the unit circle $C(0,1)$.
The number $w$ of the roots in the unit segment $[-1,1]$ or FAILURE.
  1. 1.
    Compute the polynomial $s_{-}(z)=s(\frac{z}{\sqrt{\theta}})$.
  2. 2.
    Choose a sufficiently large $q$ and compute Cauchy sums $\tilde{s}_{0,q}$ of the roots of $s_{-}(z)$ in the unit disc $D(0,1)$.
  3. 3.
    If the value $s_{0,q}$ is sufficiently close to an integer $j$, then output $w:=d-j$ and stop. Otherwise output FAILURE.
Let us prove correctness of the algorithm. By assumption the annulus $A(0,\frac{1}{\theta},1)$ contains no roots, except possibly some on the circle $C(0,1)$, and so $j$ is the number of the zeros of $s(z)$ in the disc $D(0,\frac{1}{\theta})$ by virtue of Observation 3.1. Clearly, the polynomial $s(z)$ of degree $2d$ has as many zeros lying in and outside the unit disc, and so precisely $2d-2j$ zeros of $s(z)$ lie on the unit circle $C(0,1)$. This is twice as many zeros as $p(x)$ has on the unit segment $[-1,1]$.
PART III: Classical subdivision root-finder with novel e/i tests based on root-lifting
Organization of Part III: We devote the next section to background. In Secs. 11 and 12 we cover our e/i tests and their extensions to $m$-tests, respectively. In Secs. 13 and 15 we estimate computational precision of our root-finders and their bit operation cost, respectively. In short Sec. 14 we extend our study of root-finding to approximation of matrix eigenvalues.
10 Background
10.1 Definitions
We recall or redefine some concepts from the Introduction and Part II.
  • ${\rm NR}_{t}(x):=\frac{t(x)}{t^{\prime}(x)}$ is Newton’s ratio, ${\rm NIR}_{t}(x):=\frac{t^{\prime}(x)}{t(x)}$ is Newton’s inverse ratio for a polynomial $t(x)$, ${\rm NIR}_{t}(x){\rm NR}_{t}(x)=1$.
  • We write NR$(x):={\rm NR}_{p}(x)$ and NIR$(x):={\rm NIR}_{p}(x)$.
  • We approximate NIR$(x)$ at the cost of evaluation of $p(x)$ at two points by using the first order divided differences (cf. (1.6)).
  • Define squares, discs, circles (circumferences), and annuli on the complex plane: $S(c,\rho):=\{x:~{}|\Re(c-x)|\leq\rho$, $|\Im(c-x)|\leq\rho\}$$D(c,\rho):=\{x:~{}|x-c|\leq\rho\},$ $C(c,\rho):=\{x:~{}|x-c|=\rho\}$,  $A(c,\rho,\rho^{\prime}):=\{x:~{}\rho\leq|x-c|\leq\rho^{\prime}\}.$
  • $\Delta(\mathbb{S})$, $CH(\mathbb{S})$, $X(\mathbb{S})$ and #$\#(\mathbb{S})$ denote the diameter, root set, convex hull, and index (root set’s cardinality) of a set, or region $\mathbb{S}$ on the complex plane, respectively.
  • A disc $D(c,\rho)$, a circle $C(c,\rho)$, or a square $S(c,\rho)$ is $\theta$-isolated for $\theta\geq 1$ if $X(D(c,\rho))=X(D(c,\theta\rho))$, $X(C(c,\rho))=X(A(c,\rho/\theta,\rho\theta)$ or $X(S(c,\rho))=X(S(c,\theta\rho)),$ respectively.
    The isolation of a disc $D(c,\rho)$, a circle $C(c,\rho)$, or a square $S(c,\rho)$ is the largest upper bound on such a value $\theta$ (see Fig. 4) and is denoted $i(D(c,\rho))$, $i(C(c,\rho))$, or $i(S(c,\rho))$, respectively.
    We say isolated for “$\theta$-isolated" if $\theta-1$ exceeds a positive constant.
    $\rho$$\rho\theta$$\rho\theta^{2}$c********
    Figure 5: The roots are marked by asterisks. The red circle $C(c,\rho\theta)$ and the disc $D(c,\rho\theta)$ have isolation $\theta$. The disc $D(c,\rho)$ has isolation $\theta^{2}$.
  • A disc $D$ has rigidity $\eta(\mathbb{S})=\Delta(X(D))/\Delta(D)\leq 1$. A disc is $\eta$-rigid for any $\eta\leq\eta(D)$.
    We say rigid for “$\eta$-rigid" if $\eta$ exceeds a positive constant.
  • $|u|:=||{\bf u}||_{1}:=\sum_{i=0}^{q-1}|u_{i}|$, $||{\bf u}||:=||{\bf u}||_{2}:=(\sum_{i=0}^{q-1}|u_{i}|^{2})^{\frac{1}{2}}$, and $|u|_{\rm max}:=||{\bf u}||_{\infty}:=\max_{i=0}^{q-1}|u_{i}|$ denote the 1-, 2-, and $\infty$-norm, respectively, of a polynomial $u(x)=\sum_{i=0}^{q-1}u_{i}x^{i}$ and its coefficient vector [^30] Here and hereafter ${\bf v}$ denotes a column vector, while ${\bf v}^{T}$ denotes its transpose. ${\bf u}=(u_{i})_{i=0}^{d-1}$, and then (see [46, Eqs. (2.2.5)–(2.2.7)])
    $||{\bf u}||_{\infty}\leq||{\bf u}||_{2}\leq||{\bf u}||_{1}\leq{\sqrt{q}}~{}||{\bf u}||_{2}\leq q||{\bf u}||_{\infty}.$ (10.1)
  • $r_{1}(c,t)=|y_{1}-c|,\dots,r_{d}(c,t)=|y_{d}-c|$ in non-increasing order are the $d$ root radii, that is, the distances from a complex center $c$ to the roots $y_{1},\dots,y_{d}$ of a $d$th degree polynomial $t(x)$. Write $r^{h}_{j}(c,t):=(r_{j}(c,t))^{h}$, $r_{j}(c):=r_{j}(c,p)$.
  • Generalize e/i tests as follows:
    Definition 10.1.
[…]
Observation 10.1.
Taylor’s shift, or translation, of the variable $x\mapsto y:=x-c$ and its scaling combined (see (1.4)) map the zeros $z_{j}$ of $p(x)$ into the zeros $y_{j}=\frac{z_{j}-c}{\rho}$ of $t(y)$, for $j=1,\dots,d$, and preserve the index and the isolation of the disc $D(c,\rho)$.
10.2 Reverse polynomial
Immediately observe the following properties of the reverse polynomial of (1.2).
Proposition 10.1.
(i) The roots of $p_{\rm rev}(x)$ are the reciprocals of the roots of $p(x)$, and hence
$r_{j}(0,p)r_{d+1-j}(0,p_{\rm rev})=1~{}{\rm for}~{}j=1,\dots,d.$ (10.2)
(ii) The unit circle $C(0,1)$ has the same isolation for $p(x)$ and $p_{\rm rev}(x).~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
${\rm(iii)}~{}p^{\prime}_{\rm rev}(x)=dx^{d-1}p\Big{(}\frac{1}{x}\Big{)}-x^{d-2}p^{\prime}\Big{(}\frac{1}{x}\Big{)}~{}{\rm for}~{}x\neq 0,~{}p_{d}=p_{\rm rev}(0),~{}p_{d-1}=p^{\prime}_{\rm rev}(0).$
${\rm(iv)}~{}~{}~{}~{}\frac{p_{\rm rev}(x)}{p^{\prime}_{\rm rev}(x)}=\frac{1}{x}-\frac{{\rm NIR}(x)}{x^{2}},~{}x\neq 0,~{}{\rm NIR}(0)=\frac{p_{d-1}}{p_{d}}=-\sum_{j=1}^{d}z_{j}.~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
[…]
We will not use the following theorem but include it for the sake of completeness.
Theorem 10.1.
For $p(x)$ of (1.1) and $R>1$ it holds that
$\Big{|}\frac{p(R)}{R^{d}}-p_{d}\Big{|}\leq\frac{1}{R-1}\max_{i<d}|p_{i}|,$ (10.3)
$\Big{|}\frac{Rp^{\prime}(R)-dp(R)}{R^{d-1}}-p_{d-1}\Big{|}\leq\frac{1}{R-1}\max_{0<i<d}|p_{i}|.$ (10.4)
Proof.
To prove bound (10.3) notice that
$\frac{p(R)}{R^{d}}-p_{d}=\sum_{i=0}^{d-1}p_{i}R^{i-d}=\frac{1}{R}\sum_{i=0}^{d-1}p_{i}R^{i-d+1},$
and so
$\Big{|}\frac{p(R)}{R^{d}}-p_{d}\Big{|}\leq\frac{1}{R}\sum_{i=0}^{d-1}|p_{i}|R^{i-d+1}\leq\max_{i<d}|p_{i}|\sum_{i=0}^{d-1}|p_{i}|R^{i-d}\leq\frac{1}{R-1}\max_{i<d}|p_{i}|.$
Likewise, to prove bound (10.4) notice that
$Rp^{\prime}(R)-dp(R)-dR^{d-1}p_{d-1}=\sum_{i=1}^{d-1}ip_{i}R^{i-1}.$
Hence
$\frac{Rp^{\prime}(R)-dp(R)}{R^{d-1}}-p_{d-1}=\frac{1}{R}\sum_{i=1}^{d-1}\frac{i}{d}p_{i}R^{i-d+1},$
and so
$\Big{|}\frac{Rp^{\prime}(R)-dp(R)}{R^{d-1}}-p_{d-1}\Big{|}\leq\max_{0<i<d}|p_{i}|\sum_{i=1}^{d-1}R^{i-d}\leq\frac{1}{R-1}\max_{0<i<d}|p_{i}|.$
10.3 Root-squaring and root-lifting
The DLG root-squaring iterations (1.5) by Dandelin 1826, Lobachevsky 1834, and Gräffe 1837 compute the polynomials $p_{2^{i}}(x):=\prod_{j=1}^{d}(x-z_{j}^{2^{i}})$, for $i=0,1,\dots$. By applying $h$-lifting of (1.11) we extend the definition of $p_{h}(x)$ to all $h\in\mathbb{N}$.
As we noticed in the introduction, the known studies of DLG iterations always focus and stumble on numerically unstable computation of the coefficients of the polynomials $p_{i}(x)$ [82, 109], but we only compute their values bypassing those problems.
[…]
${\rm NIR}_{p_{h}}(x)=\frac{1}{{\rm NR}_{p_{h}}(x)}\approx\frac{p_{h}(x)-p_{h}(x-\delta)}{p_{h}(x)\delta}=\frac{p_{h}^{\prime}(y)}{p_{h}(x)}.$ (10.5)
Remark 10.1.
[…]
Observation 10.2.
(i) Root-lifting preserves the index #$\#(D(0,1))$ of the unit disc. (ii) $h$-lifting (achieved with root-squaring for $h=2^{k}$ and a positive integer $k$) maps the values $i(D(0,1))$, $\sigma(D(0,1))$, $\eta(D(0,1))$, and $r_{j}(0,p)$ for all $j$, that is, the isolation of the unit disc $D(0,1)$, its softness, its rigidity and root radii, to their $h$-th powers.
10.4 The distance to a closest zero of $p$ and fast inclusion tests
Next we estimate the distance from the origin to the closest zero $z_{d}$ of $p$. The distance to its farthest zero $z_{1}$ is given by the same estimates applied to the reverse polynomial $p_{\rm rev}(x)$.
For boundons $r_{d}$ and $r_{1}$ in terms of the coefficients of $p$, see [77, Sec. 1.10]; in particular see Fujihara’a bounds of [39]:
$r_{1}\leq 2\max_{i=1}^{d}|p_{d-j}/p_{d}|^{1/j}$ unless $p_{d}=0$ and $r_{d}\geq 2\min_{i=1}^{d}|p_{j}/p_{0}|^{1/j}$ unless $p_{0}=0$.
For a complex $c$ and positive integers $j$ and $h$, Eqn. (1.13) and Observation 10.2 combined imply the following relationships independent of $p_{d}$:
$r_{d}(c)\leq d~{}|{\rm NR}(c)|,$ (10.6)
$r^{h}_{j}(0,p)=r_{j}(0,p_{h}).$ (10.7)
The roots lying far from a point $c$ little affect NIR$(c)$, and we extend bounds (10.6) as follows.
Proposition 10.2.
If #$\#(D(c,1))=\#(D(c,\theta))=m$ for $\theta>1$, then $|{\rm NIR}(c)|-\frac{d-m}{\theta-1}<\frac{m}{r_{d}(c)}.$
Proof.
We can assume that $|z_{j}|>\theta$ if and only if $j>m$ and then deduce from Eqn. (1.13) that $|{\rm NIR}(c)|\leq\sum_{j=1}^{m}\frac{1}{|c-z_{j}|}+\sum_{j=m+1}^{d}\frac{1}{|c-z_{j}|}<\frac{m}{r_{d}(c)}+\frac{d-m}{\theta-1}$. ∎
Certify inclusion into the disc $D(0,\sigma)$ if $\rho_{d}(0,p_{h})\leq\sigma^{h}$, $\sigma>1$ and $h>0$, although this bound is very poor for worst case polynomials $p$:
Proposition 10.3.
The ratio $|{\rm NR}(0)|$ is infinite for $p(x)=x^{d}-u^{d}$ and $u\neq 0$, while
$r_{d}(c,p)=r_{1}(c,p)=|u|.$
Proof.
Notice that $p^{\prime}(0)=0\neq p(0)=-u^{d}$, $p(x)=\prod_{j=0}^{d-1}(x-u\zeta_{d}^{j})$ for $\zeta_{d}$ of (1.10). ∎
Next we argue that such deficiency only occurs for a narrow class of polynomials $p$ and comment on some heuristic recipes to counter it.
Notice that $\frac{1}{r_{d}(c)}\leq|{\rm NIR}(c)|=\frac{1}{d}\Big{|}\sum_{j=1}^{d}\frac{1}{c-z_{j}}\Big{|}$ by virtue of (1.13), and so approximation to the root radius $r_{d}(c,p)$ is poor if and only if severe cancellation occurs in the summation of the $d$ roots.
Formal probabilistic support for this observation is hard, but intuitively such a cancellation only occurs for a very narrow class of polynomials $p$. If so, the above non-costly test certifies inclusion for a very large class of polynomials $p$ [^34] One can try some empirical recipes. Clearly, the problem persists for the root radius $r_{d}(w)$ where $p^{\prime}(w)$ vanishes; rotation of the variable $p(x)\leftarrow t(x)=p(ax)$ for $|a|=1$ does not fix it, but the bound $r_{d}\leq|p_{0}/p_{d}|^{1/d}$ and shifts $p(x)\leftarrow t(x)=p(x-c)$ for $c\neq 0$ combined should fix it. We must also study the case where e/i tests output exclusion, however, because this occurs in more than 25% of all e/i tests in a subdivision process:
Proposition 10.4.
Let $\sigma_{i}$ suspect squares enter the $i$-th subdivision step for $i=1,\dots,t+1$. Let it stop at $f_{i}$ of them, discard $d_{i}$ and further subdivide $\sigma_{i}-d_{i}-f_{i}$ of them. Write $\sigma_{1}:=4$, $\sigma_{t+1}:=0$, $\Sigma:=\sum_{i=1}^{t}\sigma_{i}$, $D:=\sum_{i=1}^{t}d_{i}$, and $F:=\sum_{i=1}^{t}f_{i}$. Then $3\Sigma=4(D+F-1)$ and $D>\Sigma/4$.
Proof.
Notice that $\sigma_{i+1}=4(\sigma_{i}-d_{i}-f_{i})~{}{\rm for}~{}i=1,2,\dots,t+1$, sum these equations for all $i$, substitute $\sigma_{1}=4$ and $\sigma_{t+1}=0$, and obtain $3\Sigma+4\geq 4D+4F$. Represent subdivision process by a tree with $F\leq m$ leaves whose other nodes are components made up of suspect squares, to obtain pessimistic lower bounds $\Sigma+1\geq 2m\geq 2F$ and hence $D\geq\Sigma/4+1/2$. ∎
11 Our e/i tests
11.1 Overview
Map (1.4) and Cor. 1.3 together reduce an e/i test for any disc to the test for the unit disc $D=D(0,1)$, and we devise soft e/i tests for that disc assuming that #$\#(D(0,2))\leq m$ for a fixed $m$, $1\leq m\leq d$ (cf. claim (iv) of Cor. 1.2). Our Alg. 11.1 of Sec. 11.2 and Alg. 11.2 of Sec. 11.3 are deterministic specifications of Alg. 1.1, which is always correct when it claims inclusion, by virtue of Lemma 1.1. We prove that both of these specifications are also correct when they output exclusion. Namely, Algs. 11.1 and 11.2 evaluate ${\rm NIR}(x)$ at $hq$ equally spaced points of the unit circle $C(0,1)$ for $h=O(\log(d))$ and correctly claim exclusion if at all these points $|{\rm NIR}(x)|\leq\frac{1}{2\sqrt{q}}$ for $q>d$ or if $|{\rm NIR}(x)|\leq\frac{1}{3\sqrt{q}}$ for $q>m$, respectively. Our Example 11.1 shows that we cannot devise such deterministic algorithms for $q$ decreased below $m$, but in Sec. 11.4 we devise Las Vegas randomized e/i tests that evaluate polynomial $p(x)$ at $O(\log(d))$ points $x$ overall under the Random Root Model.
11.2 Basic deterministic e/i test
[…]
Algorithm 11.1.
Basic deterministic e/i test.
INPUT: a polynomial $p(x)$ of a degree $d$ and $\sigma$, $1<\sigma<\sqrt{2}$.
INITIALIZATION: Fix two integers $q>d$ and $h:=\lceil\log_{\sigma}(1+2d\sqrt{q})\rceil$, such that $\sigma^{h}\leq 1+2d\sqrt{q}$.
COMPUTATIONS: For $g=0,1,\dots,q-1$ recursively compute
$v_{g}:=\Big{|}\frac{p^{\prime}_{h}(\zeta^{g})}{p_{h}(\zeta^{g})}\Big{|}.$ (11.1)
Stop and certify $\sigma$-soft inclusion #$(D(0,\sigma))>0$ if $1+d/v_{g}\leq\sigma^{h}$ for an integer $g<q$. Otherwise stop and certify exclusion, that is, #$(D(0,1))=0$.
Alg. 11.1 outputs $\sigma$-soft inclusion correctly by virtue of Lemma 1.1. Next we prove that it also correctly claims exclusion.
Lemma 11.1.
A polynomial $p$ has no roots in the unit disc $D(0,1)$ if $|p(\gamma)|>2\max_{w\in D(0,1)}|p^{\prime}(w)|$ for a complex $\gamma\in C(0,1)$.
Proof.
By virtue of Taylor-Lagrange’s theorem $p(x)=p(\gamma)+(x-\gamma)p^{\prime}(w)$, for any pair of $x\in D(0,1)$ and $\gamma\in C(0,1)$ and for some $w\in D(0,1)$. Hence $|x-\gamma|\leq 2$, and so $|p(x)\geq|p(\gamma)|-2~{}|p^{\prime}(w)|.$ Therefore, $|p(x)|>0$ for $\gamma$ satisfying the assumptions of the lemma and for any $x\in D(0,1)$. ∎
Proposition 11.1.
#$\#(D(0,1))=0$ for a black box polynomial $p(x)$ of a degree $d$ if
[…]
Proof.
Bounds (11.1) and (11.2) together imply that $|p^{\prime}(\zeta^{g})|/v\leq|p(\zeta^{g})|$ for $g=0,\dots,q-1$ and hence $2\sqrt{q}||(p^{\prime}(\zeta^{g}))_{g=0}^{q-1}||_{\infty}<\widehat{P}~{}{\rm for}~{}\widehat{P}:=||(p(\zeta^{g}))_{g=0}^{q-1}||_{\infty}.$ Combine this bound with (10.1) to obtain
$2||(p^{\prime}(\zeta^{g}))_{g=0}^{q-1}||_{2}<\widehat{P}.$ (11.3)
Pad the coefficient vector of the polynomial $p^{\prime}(x)$ with $q-d-1$ initial coordinates 0 to arrive at $q$-dimensional vector ${\bf p^{\prime}}$. Then $(p^{\prime}(\zeta^{g}))_{g=0}^{q-1}=F{\bf p^{\prime}}$ for the $q\times q$ matrix $F:=(\zeta^{ig})_{i,g=0}^{q-1}$ of discrete Fourier transform. Substitute $F{\bf p^{\prime}}$ for $(p^{\prime}(\zeta^{g}))_{g=0}^{q-1}$ in (11.3) and obtain $2||F{\bf p^{\prime}}||_{2}<\widehat{P}$. Hence $2\sqrt{q}||{\bf p^{\prime}}||_{2}<\widehat{P}$ because the matrix $\frac{1}{\sqrt{q}}F$ is unitary (cf. [46, 98]). It follows (cf. (10.1)) that $2||{\bf p^{\prime}}||_{1}<\widehat{P}.$ Since $\max_{x\in D(0,1)}|p^{\prime}(x)|\leq||{\bf p^{\prime}}||_{1}$, deduce that $2\max_{x\in D(0,1)}|p^{\prime}(x)|<\widehat{P}=||(p(\zeta^{g}))_{g=0}^{q-1}||_{\infty}$. Complete the proof by combining this bound with Lemma 11.1. ∎
Now, if Alg. 11.1 outputs exclusion, then $1+\frac{d}{v_{g}}>\sigma^{h}$ for $g=0,1,\dots,q-1$, where $\sigma^{h}>1+2d\sqrt{q}$ according to initialization of Alg. 11.1. Hence $\frac{1}{v_{g}}>2\sqrt{q}$, and so the assumptions of Prop. 11.1 hold for $p_{h}(x)$ replacing $p(x)$. Therefore, the theorem implies that the disc $D(0,1)$ contains no zeros of $p_{h}(x)$, but then it contains no zeros of $p(x)$ as well (see Observation 10.2 (iii)), and so exclusion is correct.
Readily verify the following estimates.
Proposition 11.2.
Alg. 11.1 is a $\sigma$-soft e/i test, certifying inclusion at the cost of evaluation of polynomial $p(x)$ at $2hj$ points $x$ or exclusion at the cost of evaluation at $2hq$ points $x$ for three integers $h$, $j$, and $q$ such that $1\leq j\leq q$, $q>m$, and $h:=\lceil\log_{\sigma}(1+2d\sqrt{q})\rceil$, $h=O(\log(d))$ for $\sigma-1$ exceeding a positive constant.
11.3 A faster deterministic e/i test
[…]
Next we supply some relevant estimates.
Algorithm 11.2.
A faster deterministic e/i test.
INPUT: a black box polynomial $p(x)$ of a degree $d$, two integers $m$ and $q>m$, and real $\sigma$ and $\theta>1$ such that $1<\sigma<\sqrt{2}$,
$|z_{j}|\leq 1~{}{\rm for}~{}j\leq m,~{}|z_{j}|>\theta~{}{\rm for}~{}j>m,$
$\frac{d-m}{\theta^{h}-1}\leq\frac{1}{6\sqrt{q}}~{}{\rm for}~{}h:=\lceil\log_{\sigma}(1+6m\sqrt{q})\rceil~{}{\rm such~{}that}~{}\sigma^{h}\geq 1+6m\sqrt{q}.$
COMPUTATIONS: Recursively compute the values $v_{g}$ of (11.1) for $g=0,1,\dots,q-1$. Certify $\sigma$-soft inclusion, that is, #$(D(0,\sigma))>0$, if $v_{g}\geq\frac{1}{3\sqrt{q}}$ for some integer $g<q$. Otherwise certify exclusion, #$(D(0,1))=0$.
Prop. 10.2, for $p_{h}$ replacing $p$, and the assumed bound $\frac{d-m}{\theta^{h}-1}\leq\frac{1}{6\sqrt{q}}$ together imply
$\frac{m}{r_{d}(\zeta^{g},p_{h})}\geq v_{g}-\frac{d-m}{\theta^{h}-1}\geq v_{g}-\frac{1}{6\sqrt{q}}~{}{\rm for~{}all}~{}g.$
Alg. 11.2 claims inclusion if $v_{g}\geq\frac{1}{3\sqrt{q}}$ for some $g$, but then
$\frac{m}{r_{d}(\zeta^{g},p_{h})}\geq\frac{1}{6\sqrt{q}},~{}r_{d}(\zeta^{g},p_{h})\leq 6m\sqrt{q},$
and hence
$r_{d}(0,p_{h})\leq 1+6m\sqrt{q}~{}\leq\sigma^{h}~{}{\rm for}~{}|\zeta|=1.$
Now Observation 10.2 (iii) implies that Alg. 11.2 claims inclusion correctly.
Next prove that it also claims exclusion correctly.
Proposition 11.3.
Let a polynomial $p(x)$ of a degree $d$ have exactly $m$ roots in the disc $D(0,\theta)$ for
$\theta^{h}\geq 1+6(d-m)\sqrt{q},~{}h:=\lceil\log_{\sigma}(1+6m\sqrt{q})\rceil,~{}{\rm and~{}an~{}integer}~{}q>m$
and let $v<\frac{1}{3\sqrt{q}}$ for $v$ of (11.2). Then #$\#(D(0,1))=0$.
Proof.
Let $p(x)$ and its factor $f(x)$ of degree $m$ share the sets of their zeros in $D(0,\theta)$. Then (1.13) implies that $\Big{|}\frac{f^{\prime}(x)}{f(x)}\Big{|}<\Big{|}\frac{p^{\prime}(x)}{p(x)}\Big{|}+\frac{d-m}{\theta-1}~{}{\rm for}~{}|x|=1,$ and so $\max_{g=0}^{q-1}\Big{|}\frac{f^{\prime}(\zeta_{q}^{g})}{f(\zeta_{q}^{g})}\Big{|}<v+\frac{d-m}{\theta-1}$. Recall that $\frac{d-m}{\theta-1}\leq\frac{1}{6\sqrt{q}}$ and $v<\frac{1}{3\sqrt{q}}$ by assumptions.
Combine these bounds to obtain $\max_{g=0}^{q-1}\Big{|}\frac{f^{\prime}(\zeta_{q}^{g})}{f(\zeta_{q}^{g})}\Big{|}<\frac{1}{2\sqrt{q}}$. To complete the proof, apply Prop. 11.1 for $f(x)$ replacing $p(x)$ and for $m$ replacing $d$. ∎
Readily extend Prop. 11.2 to Alg. 11.2 by decreasing its bounds on $q$ and $h$ provided that #$\#(D(0,\theta))=m$ for $\theta\geq 1+6(d-m)\sqrt{q}$.
Proposition 11.4.
Alg. 11.2 is a $\sigma$-soft e/i test: it certifies either inclusion or exclusion at the cost of evaluation of $p(x)$ at $2hj$ or $2hq$ points $x$, respectively, for three integers $h$, $j$, and $q$ such that $1\leq j\leq q$, $q>m:=\#(D(0,\theta)),~{}\sigma^{h}>1+6m\sqrt{q}$, $\theta\geq 1+6(d-m)\sqrt{q}$, and $h:=\lceil\log_{\sigma}(1+6m\sqrt{q})\rceil$, so that $h=O(\log(m))$ if $\sigma-1$ exceeds a positive constant.
11.4 A randomized e/i test
13.4.1. Introductory comments. Would Algs. 11.1 and 11.2 stay correct if we apply them at fewer evaluation points? They have consistently stayed correct for $q$ of order $\log^{2}(d)$ in extensive numerical tests of [57, 60] applied to $p(x)$ rather than to $p_{i}(x)$ for $i>0$ (that is, used no root-lifting). The tests reported exclusion where the values $|s_{h,q}|$ were reasonably small for $s_{h,q}:=\frac{1}{q}\sum_{g=0}^{q-1}\zeta^{(h+1)g}~{}\frac{p^{\prime}(\zeta^{g})}{p(\zeta^{g})},~{}{\rm for}~{}h=0,1,2$ (cf. Eqn. 3.2 in Part II), and reported inclusion otherwise. Clearly, $|s_{h,q}|\leq\delta$ for all $h$ if $|p(\zeta^{g})|\leq\delta/q$ for $g=0,1,\dots,q-1$, while inclusion is certified unless $|p(\zeta^{g})|$ is small for all $g$.
The test results of [57, 60], however, cannot be extended to worst case input polynomials $p(x)$ already for $q=m-1$.
Example 11.1.
Let $q=m-1$, $p(x)=(1-ux^{d-m})(x^{m}-mx+u)$, $u\approx 0$ but $u\neq 0$. (i) First let $m=d$. Then $p(x)=(1-u)(x^{d}-dx+u)$, $p^{\prime}(x)=(1-u)d(x^{d-1}-1)$ and NIR$(x)=0$ for $x$ being any $(d-1)$-st roots of unity. Extension of our e/i test to $q=d-1$ would imply that #$\#(D(0,1))=0$, while $p(y)=0$ for $u\approx 0$ and some $y\approx 0$. (ii) For $0<m<d$ notice that $p^{\prime}(x)=uf(x)+m(1-ux^{d-m})(x^{m-1}-1)$ where $f(x)=(d-m)x^{d-m-1}(x^{m}-mx+u)$ and so $uf(x)\approx 0$ while $1-ux^{d-m}\approx 1$ for $|x|\leq 1$ and $u\approx 0$. Therefore, $p^{\prime}(x)\approx 0$ where $x$ is any $(m-1)$-st root of unity, and then again extension of our e/i test to $q=m-1$ would imply that #$\#(D(0,1))=0$, while $p(y)=0$ for $u\approx 0$ and some $y\approx 0$.
Clearly, [57, 60] have never encountered worst case inputs in their extensive tests. Next we formally support this observation by introducing probability structure in the space of polynomials $p(x)$ and $t(x)=p(\frac{x-c}{\rho})$. Shifts destroy all customary probability distributions in the space of the coefficients of $p(x)$, but we prove correctness whp of e/i tests that evaluate $p(x)$ at $O(\log(d))$ under the Random Root Model.
[…]
Lemma 11.2.
Fix any constant $\beta>1$ and a sufficiently large integer $h$ of order $\log(d)$. Then under the Random Root Model, Alg. 1.1(rrm) is correct with a probability at least $1-P$, where
$P\leq\frac{\gamma d}{\pi R^{2}}~{}{\rm for}~{}\gamma:=\pi\max\Big{\{}\frac{\beta^{2}}{\sigma^{2}},O\Big{(}\frac{1}{\beta^{2h}}\Big{)}\Big{\}}.$ (11.4)
Proof.
By virtue of Lemma 1.1 Alg. 1.1(rrm) outputs $\sigma$-soft inclusion correctly and therefore can be incorrect only where it outputs exclusion for a disc $D(0,1)$ containing a zero of $p$.
Let this zero be $z_{1}$, say. Then fix the values of the random variables $z_{2},\dots,z_{d}$ and $x_{1}\in C(0,1)$ and let $\gamma$ denote the area (Lebesgue’s measure) of the set $\mathbb{S}$ of the values of the variable $z_{1}$ for which
$v_{1}=|{\rm NIR}_{t_{h}}(x_{1})|<\frac{d}{\sigma^{h}-1}.$ (11.5)
Then Alg. 1.1(rrm) is incorrect with a probability at most
$P=\frac{\gamma d}{\pi R^{2}}$ (11.6)
where the factor of $d$ enables us to cover the cases where the value of any of the $d$ iid random variables $z_{1},\dots,z_{d}$, and not necessarily $z_{1}$, lies in the disc $D(0,1)$.
Fix two values $\bar{z},z\in\mathbb{S}$ such that
$\bar{z}={\rm argmax}_{y\in\mathbb{S}}~{}|y|~{}{\rm and}~{}z={\rm argmax}{{}_{y}\in\mathbb{S}}~{}|y-\bar{z}|$
and notice that
$\gamma\leq\pi\min\{|\bar{z}|^{2}|,|\bar{z}-z|^{2}\}.$ (11.7)
To estimate $|\bar{z}-z|$, first write $y:=z^{h}\in D(0,1),~{}\bar{y}:=\bar{z}^{h}\in D(0,1)$ and prove that
$|\Delta|=|\bar{y}-y|<\nu_{h}~{}{\rm for}~{}\nu_{h}:=\frac{8d}{\sigma^{h}-1}.$ (11.8)
[…]
$|\Delta_{t_{h}}|<\frac{2d}{\sigma^{h}-1}~{}{\rm for}~{}\Delta_{t_{h}}:={\rm NIR}^{\prime}-{\rm NIR}.$ (11.9)
Now let $y_{1}:=z_{1}^{h},\dots,y_{d}=z_{d}^{h}$ denote the $d$ zeros of $t_{h}(x)$ and deduce from (1.13) that
$|{\rm NIR}|=\Big{|}\frac{1}{x_{1}-y}+S\Big{|}~{}{\rm and}~{}|{\rm NIR}^{\prime}|=\Big{|}\frac{1}{x_{1}-\bar{y}}+S\Big{|}~{}{\rm for}~{}S:=\Big{|}\sum_{j=2}^{d}\frac{1}{x_{1}-y_{j}}\Big{|}.$
Hence
$|\Delta_{t_{h}}|=|{\rm NIR}-{\rm NIR}^{\prime}|=\Big{|}\frac{1}{x_{1}-y}-\frac{1}{x_{1}-\bar{y}}\Big{|}=\frac{|\Delta|}{|(x_{1}-y)(x_{1}-\bar{y})|}.$ (11.10)
Recall that $x_{1}\in C(0,1)$, while $y,\bar{y}\in D(0,1)$, and so $\max\{|x_{1}-y|,|x_{1}-\bar{y}|\}\leq 2.$
Combine this inequality with (11.10) and deduce that $|\Delta_{t_{h}}|\geq|\Delta|/4.$
Combine the latter bound with (11.10) and obtain (11.8).
Now fix a large integer $g$, substitute $\bar{y}=\bar{z}^{h}$ and $y=z^{h}$ for $h=g$ and $h=g+1$, and obtain
$\lambda_{g+1}:=z^{g+1}-\bar{z}^{g+1},~{}\lambda_{g}:=z^{g}-\bar{z}^{g},~{}{\rm and~{}so}~{}z^{g+1}=z\bar{z}^{g}+z\lambda_{g},~{}|\lambda_{h}|\leq\nu_{h}~{}{\rm for}~{}h=g,g+1.$
Therefore $\lambda_{g+1}+\bar{z}^{g+1}=\lambda_{g}z+\bar{z}^{g}z$, and so
$z-\bar{z}=\frac{\lambda_{g+1}-z\lambda_{g}}{\bar{z}^{g}},~{}|\bar{z}-z|\leq\frac{|\lambda_{g+1}|+|z|\cdot|\lambda_{g}|}{|\bar{z}|^{g}}<\frac{\nu_{g+1}+|z|\nu_{g}}{|\bar{z}|^{g}}=O\Big{(}\frac{d}{(\sigma|\bar{z}|)^{g}}\Big{)}.$
Hence $|z-\bar{z}|=O(1/\beta^{g})$ for a fixed $\beta>1$. Combine this bound with (11.6) and (11.7), choosing sufficiently large $g$ of order $\log d$ unless $\sigma|\bar{z}|<\beta$.
Combine this bound with (11.6) and (11.7), recall that $h=g$ or $h=g+1$, and obtain the lemma. ∎
For a fixed $\sigma>1$, e.g., for $\sigma=1.2$, choose a constant $\beta>1$ and sufficiently large $g=O(\log(d))$ such that $\gamma<\pi$ in (11.4). Then under the Random Root Model the probability $P$ of the output error for Alg. 1.1(rrm) is less than a constant $\nabla<1$. We can decrease the error probability below $\nabla^{v}$ by applying the algorithm $v$ times, in particular below $d^{a}$ for any fixed positive constant $a$ and for $v=O(\log(d))$.
Cor. 1.2 reduces classical subdivision root-finder to performing $O(mb)$ soft e/i tests, but in Part IV we decrease that bound to $w=O(m\log(b))$ tests. These tests magnify the error bounds by at most a factor of $w$. Since $m\leq d$, we can ensure correctness of the resulting e/i test whp provided that $\log(b)=o(d^{a-1})$. We arrive at the following
Corollary 11.1.
Under Random Root Model one can perform all $O(m\log(b)))$ soft Las Vegas randomized e/i tests involved in a subdivision root-finder which ensure output error bound $\epsilon=1/2^{b}$ whp provided that $\log(b)=d^{O(1)}$ and that every test is reduced to $v=O(\log(d))$ applications of Alg. 1.1(rrm).
Remark 11.1.
It is tempting to combine root-lifting or root-squaring with algorithms of [111], which for approximate Cauchy integral (1.9), to obtain alternative randomized e/i test supporting Cor. 11.1, but for that we would have to raise disc isolation from order of $1+1/m$ to a constant, and this would require lifting of order $m$.
11.5 Counting the evaluation queries of our root-finders
Combine Cors. 1.2 and 11.1 and Prop. 11.4 and obtain
Corollary 11.2.
We can reduce root-finding to deterministic evaluation of $p(x)$ at $O(m^{2}b\log(d))$ points as well as evaluation at $O(mb\log^{2}(d))$ points under the Random Root Model.
12 Reduction of $m$-test to e/i tests
Our algorithm in Sec. 12.1 reduces $m$-test to e/i tests motivated by application in Part IV. For these application our slower alternative reduction in Sec. 12.2 is inferior, but we include it because it relaxes isolation assumption and works for any disc, which can be of independent interest.
12.1 Simple reduction to a single e/i test in the case of an isolated disc
Clearly, we can immediately reduce the $d$-test in the unit disc $D(0,1)$ for $p(x)$ to performing 1-test (that is, e/i test) in that disc for the reverse polynomial $p_{\rm rev}(x)$ of (1.2).
Let us next reduce an $m$-test, for $0<m<d$, to a single application of 1-test provided that an input disc $D$ with $m$ roots is $\theta$-isolated for $\theta$ of order $\log(d)$; we can ensure that provision by applying root-lifting or root-squaring, in both cases at the cost of evaluation of $p(x)$ at $O(\log(d))$ points if an initial disc is just isolated.
Proposition 12.1.
Let the root sets of two polynomials $f(y)$ and $g(y)$ of degrees $m$ and $d-m$, respectively, lie outside the disc $D(0,1)$ and in the disc $D(0,1/\theta)$, respectively, for a sufficiently large real $\theta$. Then
${\rm NIR}_{f}(z)={\rm NIR}_{fg}(z)-\frac{d-m}{z}+\Delta~{}{\rm for}~{}|z|=1~{}{\rm and}~{}|\Delta|\leq\frac{d-m}{\theta-1}.$
Proof.
Notice that $\frac{1}{z-y_{j}}-\frac{1}{z}=\frac{y_{j}}{(z-y_{j})z}$, recall that $|y_{j}|\leq\frac{1}{\theta}$ for $\theta>1$ and all $d-m$ roots $y_{1},\dots,y_{d-m}$ of $g(y)$ by assumption, and deduce that
$\Big{|}\frac{1}{z-y_{j}}-\frac{1}{z}\Big{|}\leq\frac{1}{\theta-1}~{}{\rm for}~{}|z|=1~{}{\rm and}~{}j=1,\dots,d-m.$ (12.1)
[…]
Next fix an integer $q$ such that $m<q\leq 2m$, e.g., $q:=m+1$ or $q:=2^{h}$ for $h=\lceil\log_{2}(m+1)$. Assume that $\theta-1$ exceeds a positive constant and ensure that
$\theta>1+4(d-m)\sqrt{q}$ (12.2)
by applying root-lifting or root-squaring at the cost of evaluation of $p(x)$ at $O(\log(d))$ points.
[…]
$V_{f}:=\max_{g=0}^{q-1}|{\rm NIR}_{fg}(\zeta^{g})-(d-m)\zeta^{-g}|.$ (12.3)
Now ensure that $V_{f}<\frac{1}{2\sqrt{q}}$, say, by requiring that $\max_{g=0}^{q-1}|{\rm NIR}_{fg}(\zeta^{g})|<\frac{1}{4\sqrt{q}}$ because bound (12.2) implies that $\frac{d-m}{\theta-1}<\frac{1}{4\sqrt{q}}$.
12.2 $m$-test by means of e/i tests for any disc
Our alternative reduction of $m$-test for $1\leq m\leq d$ to a bounded number of e/i tests can be applied in the case of any disc, not necessarily isolated. If at least one of the e/i tests fails, then definitely there is a root in a narrow annulus about the boundary circle of an input disc, and hence the internal disc of the annulus would not contain $m$ roots. Otherwise we would obtain a root-free concentric annulus about the boundary circle of an input disc and then reduce $m$-test to root-counting in an isolated disc, which we can perform at a low cost (see, e.g., claim (ii) of Prop. 3.1).
Algorithm 12.1.
Reduction of $m$-test to e/i tests.
INPUT: a polynomial $p(x)$ of a degree $d$, two values $\bar{\theta}>1$ and $\bar{\sigma}>1$, an integer $m$ such that $1\leq m\leq d$ and #$\#(D(0,\bar{\sigma}))\leq m$, and two algorithms $\mathbb{ALG}~{}0$ and $\mathbb{ALG}~{}1$, which a black box polynomial perform $\bar{\sigma}$-soft e/i test for in any disc and root-counting in an isolated disc, respectively.
INITIALIZATION (see Fig. 6): Fix a positive $\rho$ and compute the integer
$v=\lceil 2\pi/\phi\rceil$ (12.4)
where $\phi$ is the angle defined by an arc of $C(0,1)$ with end points at the distance $\rho$. Then compute the values
$c_{j}=\exp(\phi_{j}{\bf i}),~{}j=0,1,\dots,v,$ (12.5)
such that $|c_{j}-c_{j-1}|=\rho$ for $j=1,2,\dots,v$ and $\phi_{j}-\phi_{j-1}=\phi$.
$c_{1}$$c_{2}$$c_{3}$$c_{4}$$c_{5}$$c_{6}$$c_{7}$$c_{8}$$c_{9}$$c_{0}$
Figure 6: $v=9$, $|c_{j}-c_{j-1}|=\rho>|c_{9}-c_{0}|$ for $j=1,2,\dots,9$.
OUTPUT: Either an upper bound $r_{d-m+1}<1-\rho\frac{\sqrt{3}}{2}$ or one of the two lower bounds $r_{d-m+1}\geq 1-\bar{\sigma}\rho$ or $r_{d-m+1}>1+(\sqrt{3}-1)\rho$.
COMPUTATIONS: Apply algorithm $\mathbb{ALG}~{}0$ (soft e/i test) to the discs $D(c_{j},\rho)$, $j=0,1,\dots,v$. (a) Stop and certify that $r_{d-m+1}\geq 1-\bar{\sigma}\rho$, that is, #$\#(D(0,1-\bar{\sigma}\rho))<m$ for the open disc $(D(0,1-\bar{\sigma}\rho)$, unless algorithm $\mathbb{ALG}~{}0$ outputs exclusion in all its $v+1$ applications.
(b) If it outputs exclusion $v+1$ times, then apply algorithm $\mathbb{ALG}~{}1$ (root-counter) to the $\theta$-isolated disc $D(0,1)$ for
${}\theta:=\min\Big{\{}1+(\sqrt{3}-1)\rho,\frac{1}{1-0.5\rho\sqrt{3}}\Big{\}}.$ (12.6)
Let $\bar{s}_{0}=\#(D(0,1))$ denotes its output integer. Then conclude that
(i) $r_{d-m+1}\leq 1-\rho\frac{\sqrt{3}}{2}$, and so #$\#(D(0,1-\rho\frac{\sqrt{3}}{2}))=m$ if $\bar{s}_{0}=m$, while
(ii) $r_{d-m+1}\geq 1+(\sqrt{3}-1)\rho$, and so #$\#(D(0,1+(\sqrt{3}-1)\rho))<m$ for the open disc $D(0,1+(\sqrt{3}-1)\rho)$ if $\bar{s}_{0}<m$.
To prove correctness of Alg. 12.1 we need the following result.
Theorem 12.1.
For the above values $c_{0},\dots,c_{v}$, and $\rho$, the domain $\cup_{j=0}^{v-1}D(c_{j},\rho)$ covers the annulus $A(0,1-0.5\rho\sqrt{3},1+(\sqrt{3}-1)\rho)$.
Proof.
Let $c_{j-1}$ and $c_{j}$ be the centers of two neighboring circles $C(c_{j-1},\rho)$ and $C(c_{j},\rho)$, let $z_{j}$ and $z^{\prime}_{j}$ be their intersection points such that $|z_{j}|=|z_{j}^{\prime}|-|z_{j}-z^{\prime}_{j}|$ for $j=1,2,\dots,v$, observe that the values $|z_{j}-z_{j}^{\prime}|$, $|z_{j}|$, $|z^{\prime}_{j}|$, and $|c_{j-1}-c_{j}|$ are invariant in $j$, and hence so is the annulus $A(0,|z_{j}|,|z_{j}^{\prime}|)$ of width $|z_{j}-z_{j}^{\prime}|$ as well; notice that it lies in the domain $\cup_{j=0}^{v-1}D(c_{j},\rho)$.
$c_{0}$$c_{1}$$z$$z^{\prime}$
Figure 7: $|c_{1}-c_{0}|=|c_{0}-z|=|c_{1}-z|=|c_{0}-z^{\prime}|=|c_{1}-z^{\prime}|=\rho$.
Let us prove that $|z_{j}-z_{j}^{\prime}|=\rho\sqrt{3}$ for $j=1$, and hence also for $j=2,\dots,v$.
Write $z:=z_{1}$ and $z^{\prime}:=z_{1}^{\prime}$ and observe (see Fig. 7) that $|c_{1}-c_{0}|=|c_{0}-z|=|c_{1}-z|=|c_{0}-z^{\prime}|=|c_{1}-z^{\prime}|=\rho$, and so the points $c_{0}$, $c_{1}$, and $z$ are the vertices of an equilateral triangle with a side length $\rho$. Hence $|z-z^{\prime}|=\rho\sqrt{3}$. The circle $C(c_{j},\rho)$ intersects the line segment $[z_{j},z_{j}^{\prime}]$ at a point $y_{j}$ such that $|y_{j}-z_{j}|>\rho\sqrt{3}/2$ and $|y_{j}-z_{j}^{\prime}|>(\sqrt{3}-1)\rho$ for $\rho<1$, and this implies the theorem if $|c_{v}-c_{0}|=\rho$. If, however, $|c_{v}-c_{0}|<\rho$ and if $C(c_{0},\rho)\cap C(c_{v},\rho)=\{\bar{z},\bar{z}\}$, then $|\bar{z}-\bar{z}^{\prime}|>|z-z^{\prime}|$, and the theorem follows in this case as well. ∎
Corollary 12.1.
The disc $D(0,1)$ is $\theta$-isolated for $\theta$ of (12.6) provided that the discs $D(c_{j},\rho)$ contain no roots for $j=0,1,\dots,v$.
Correctness of Alg. 12.1. (i) If algorithm $\mathbb{ALG}~{}1$ outputs inclusion in at least one of its applications, then one of the discs $D(c_{j},\bar{\sigma}\rho)$ contains a root, and then Alg. 12.1 correctly reports that $r_{d-m+1}\geq 1-\sigma\bar{\rho}$.
(ii) Otherwise Cor. 12.1 holds, and then correctness of Alg. 12.1 follows from correctness of $\mathbb{ALG}~{}1$.
Softness is at most $\frac{1}{1-\bar{\sigma}\rho}$, $\frac{1}{1-0.5\rho\sqrt{3}}$, and $1+(\sqrt{3}-1)\rho$ in cases (a), (b,i), and (b,ii), respectively.
Computational complexity. Alg. 12.1 invokes algorithm $\mathbb{ALG}~{}0$ at most $v+1$ times for $v$ of (12.4). If $\mathbb{ALG}~{}0$ outputs exclusion $v+1$ times, then algorithm $\mathbb{ALG}~{}1$ is invoked.
Remark 12.1.
One can modify Alg. 12.1 as follows (see [60]): instead of fixing a positive $\rho$ and then computing $v$ of (12.4) fix an integer $v>0$, apply algorithm $\mathbb{ALG}~{}0$ at $v$ equally spaced points $c_{j}=\exp(\frac{2j\pi{\bf i}}{v}),~{}j=0,1,\dots,v-1$, let $\phi=\frac{2\pi}{v}$, and then define $\rho$. We also arrive at this algorithm if we compute or guess this value $\rho$ and then apply Alg. 12.1, but notice that $c_{v}=c_{0}$ for $c_{j}$ of (12.5) and only invoke algorithm $\mathbb{ALG}~{}0$ for $c_{0},\dots,c_{v-1}$.
13 Precision of computing
13.1 Doubling computational precision towards its a posteriori estimation
Given an algorithm $\mathbb{A}$ for black box polynomial computations and target tolerance TOL to output error, how should one choose computational precision? A simple practical recipe of [10, 14, 21] was the basis for highly successful empirical performance of MPSolve: namely, first apply the algorithm $\mathbb{A}$ with a lower precision and then recursively double it until an output error decreases below TOL; at that point it outputs precision within a factor of two from optimal, thus also supporting near optimal a posteriori bit operation complexity estimate for algorithm $\mathbb{A}$.
13.2 A priori estimates for computational precision
In the rest of this section we deduce a priori estimate for computational precision by proving that a precision of $O(b\log(d))$ bits is sufficient for supporting correctness of our subdivision root-finders. Such an estimate is deduced in [73] based on advanced non-trivial techniques, but we only use a very small part of them because the paper [73] works with higher order divided differences, while we only use the first order ones combined with Las Vegas randomization and some rather routine techniques for numerical error estimation.
Estimating computational precision of our soft e/i tests we assume that they are applied to the unit disc $D(0,1)$ because we can extend the results to any disc $D(c,\rho)$ by applying map (1.4) and increasing the upper bounds on the precision by a factor of $\log_{2}(1/\rho)$, which is at most $b$ because we only need to handle discs of radii $\rho>\epsilon=1/2^{b}$ and not $\epsilon$-discs. Therefore, our bound $O(\log(d))$ on the computational precision of our e/i tests for the disc $D(0,1)$ will imply the target bound $O(b\log(d))$ for testing any disc.
We also make Assumption 0 (Relative error bound) that we can request and obtain from evaluation oracle (black box subroutine) the values of $p$ with a relative error bound in $1/d^{\nu}$ for any fixed constant $\nu>0$ of our choice; we can represent these value with the bit-precision $\log_{2}(d)+\log_{2}(\nu)$.
In our e/i tests we only need to compare the values $|{\rm NIR}_{t_{h}}(x)|$ at some complex points $x$ with some fixed values $v_{d}\geq 1/d^{\beta}$ for a constant $\beta>0$, and next we estimate that under Assumption 0 we can indeed do this by performing our computations with a precision of $O(\log(d))$ bits.
Multiplication with a precision $s$ contributes a relative error at most $2^{s}$ (cf. [20, Eqn. A.3 of Ch. 3]) and otherwise just sums the relative error bounds of the factors because we can ignore the dominated impact of higher order errors. Hence we deduce that $h=O(\log(d))$ multiplications for root-lifting to their $h$th powers increase these bounds at most by a factor of $h$, which is immaterial since we allow relative errors of order $1/d^{\nu}$ for a constant $\nu>0$ of our choice.
We only need to apply this analysis to approximation of NIR$(x)$ for $p_{h}(x)$ based on expression (1.6).
In our analysis we also use the following
Assumption 1 (Isolation Assumption): Whp all $q$ evaluation points $x\in C(0,1)$ of an e/i test for the unit disc $D(0,1)$ lie at least at the distance $\phi:=\frac{2\pi}{mqw}$ from all zeros of $p$ for any fixed $w>1$ of our choice.
We justify this assumption by applying our e/i tests under random rotation $x\mapsto vx$ for the random variable $v$ specified in our next proposition.
Proposition 13.1.
Let $\phi_{g}:=\min_{j=1}^{d}|v\zeta^{g}-z_{j}|$ and let $\phi:=\min_{g=0}^{q-1}\phi_{g}$ for the zeros $z_{j}$ of $p$, $\zeta$ of (1.10) and random variable $v$ sampled from a fixed arc of the circle $C(0,1)$ of length $2\pi/q$ under the uniform probability distribution on that arc. Let $i(D(0,1))=m$. Then $\phi\leq\frac{2\pi}{mqw}$ for $w>1$ with a probability at most $\frac{\pi}{w}$.
Proof.
Let $\psi_{j}$ denote the length of the arc $C(0,1)\cap D(z_{j},\frac{2\pi}{mqw})$, $j=1,2,\dots,d$. Then $\psi_{j}=0$ for at least $d-m$ integers $j$, and $\max_{j=1}^{d}\psi_{j}\leq\frac{\pi^{2}}{mqw}$ for sure. Hence $\sum_{j=1}^{d}\psi_{j}<\frac{\pi^{2}}{qw}$, and so $\phi_{g}\leq\frac{2\pi}{mqw}$ for a fixed $g$ with a probability at most $\frac{\pi}{qw}$. Therefore, $\phi\leq\frac{2\pi}{mqw}$ with a probability at most $\frac{\pi}{w}$. ∎
Next we prove that precision of $O(\log(d))$ bits supports our e/i tests under Assumptions 0 and 1 by applying Prop. 13.1 for sufficiently large $w=d^{O(1)}$. We will (i) prove that $|{\rm NIR}(z)|=d^{O(1)}$, (ii) approximate this value within $d^{\nu}\delta$ for any fixed constant $\nu$ and for $\delta$ of (1.6) chosen such that $\log(\frac{1}{\delta})=O(\log(d))$, and (iii) estimate that such a bound withstands the impact of rounding errors of computations for (1.6).
Eqn. (1.13) immediately implies
Proposition 13.2.
Under Assumption 1, $\max_{g=1}^{q}|{\rm NIR}(\zeta^{g})|\leq{d^{\eta}}$ for $\eta=O(1)$.
Corollary 13.1.
A precision of $\lceil(\eta+\beta)\log_{2}(d)\rceil$ bits is sufficient to represent ${\rm NIR}(x)$ within $\frac{1}{d^{\beta}}$ for $g=1,2,\dots,q$.
Instead of ${\rm NIR}(x)=\frac{t^{\prime}(x)}{t(x)}$ for $x\in C(0,1)$ we actually approximate $\frac{t^{\prime}(y)}{t(x)}$ where $|y-x|\leq\delta$ and $t^{\prime}(y)$ is equal to the divided difference of (1.6). Thus we shall increase the above error bound by adding upper bounds $\alpha$ on $|\frac{t^{\prime}(x)}{t(x)}-\frac{t^{\prime}(y)}{t(x)}|$ and $\beta$ on the rounding error of computing $\frac{t(x)-t(x-\delta)}{\delta t(x)}$, for $x=\zeta^{g}$ and $g=1,2,\dots,q$.
In the proof of the next theorem we use the following lemma, which is [73, Fact 3.5].
Lemma 13.1.
For $t(x)=\prod_{j=1}^{d}(x-y_{j})$ and a non-negative integer $j\leq d$ it holds that
$t^{(j)}(x)=j!~{}t(x)\sum_{S_{j,d}}~{}\prod_{j\in S_{j,d}}\frac{1}{x-y_{j}}$
where the summation $\sum_{S_{j,d}}$ is over all subsets $S_{j,d}$ of the set $\{1,\dots,d\}$ having cardinality $j$.
Proposition 13.3.
Under the assumptions of Prop. 13.2 it holds that
$|t(x)\alpha|=|\frac{t(x)-t(x-\delta)}{\delta}-t^{\prime}(x)|\leq|t(x)\frac{(d\xi^{\prime})^{2}\delta}{1-d\xi^{\prime}\delta}|$ for $\xi^{\prime}=\frac{1}{\rho}$.
Proof.
The claimed bound on $|\xi^{\prime}|$ follows from (1.13). It remains to apply the first five lines of the proof of [73, Lemma 3.6] with $f$, $\alpha$, $n$, and $\xi$ replaced by $t$, $-\delta$, $d$, and $w$, respectively. Namely, first obtain from Taylor’s expansion that $t(x)-t(x-\delta)=\sum_{j=0}^{\infty}\frac{\delta^{j}}{j!}t^{(j)}(x)-t(x)$.
Substitute the expressions of Lemma 13.1 and obtain
$t(x)-t(x-\delta)=\delta t^{\prime}(x)+\sum_{j=2}^{\infty}\delta^{j}t(x)\sum_{S_{j,d}}~{}\prod_{j\in S_{j,d}}\frac{1}{x-y_{j}}.$
Combine this equation with the assumed bounds on $\frac{1}{|x-y_{j}|}$ and deduce that
$\Big{|}\frac{t(x)-t(x-\delta)}{\delta}-t^{\prime}(x)\Big{|}\leq\Big{|}\frac{f(x)}{\delta}\sum_{j=2}^{\infty}\delta^{j}w^{j}d^{j}\Big{|}\leq\Big{|}f(x)\frac{(d~{}w)^{2}~{}\delta}{1-\delta~{}d~{}w}\Big{|}.$
Corollary 13.2.
We can ensure that $\alpha<1/d^{\bar{\eta}}$ for any fixed constant $\bar{\eta}$ by choosing $\delta$ of order $d^{O(1)}$, represented with a precision of $O(\log(d))$ bits.
Proposition 13.4.
Suppose that the values $t(y)$ have been computed by a black box oracle within a relative error bounds $\nabla(y)$ for $y=x$, $y=x-\delta$, and $\delta$ of Cor. 13.2. Then one can ensure that $\beta<\frac{1}{d^{\bar{\eta}}}$ for any fixed constant $\bar{\eta}$ by choosing a proper $\nabla(y)$ of order $1/d^{O(1)}$, represented with a precision of $O(\log(d))$ bits, which we can do under Assumption 0.
Proof.
Write
$\beta\delta=\frac{t(x-\delta)(1+\nabla(x-\delta))}{t(x)(1+\nabla(x))}-\frac{t(x-\delta)}{t(x)}=\frac{t(x-\delta)}{t(x)}~{}\frac{\nabla(x)-\nabla(x-\delta)}{1+\nabla(x)}.$
Taylor-Lagrange’s formula implies that
$\frac{t(x-\delta)}{t(x)}=1+\frac{t(x-\delta)-t(x)}{t(x)}=1+\delta\frac{t^{\prime}(u)}{t(x)}$, for $u\in[x-\delta,x]$.
Hence $|\frac{t(x-\delta)}{t(x)}|\leq 1+|\delta\alpha^{\prime}|+|\frac{t^{\prime}(x)}{t(x)}|$ for $\alpha^{\prime}=\frac{t^{\prime}(u)-t(x)}{t(x)}$.
By extending Cor. 13.2 obtain $|\alpha^{\prime}|\leq\frac{1}{d^{O(1)}}$ and deduce from (1.13) and Assumptions 0 and 1 that $|\frac{t^{\prime}(x)}{t(x)}|\leq\frac{1}{d^{O(1)}}$ (cf. the proof of Prop. 6.1). Hence $w:=|\frac{t(x-\delta)}{t(x)}|\leq d^{O(1)}$ and $|\beta\delta|\leq|w|\frac{|\nabla(x)|+|\nabla(x-\delta)|}{1-|\nabla(x)|}$.
Now choose $\nabla(y)$ such that $|\nabla(y)|\leq\frac{|\delta|}{20|w|}$ for $y=x$ and $y=x-\delta$. Then verify that $\log(\frac{1}{|\nabla(y)|})=O(\log(d))$, $1-|\nabla(x)|\geq 1-\frac{|\delta|}{20|w|}>\frac{19}{20}$, and $|\delta\beta|<\frac{|\delta|}{8}$. ∎
Combine Cors. 13.1 and 13.2 and Prop. 13.4 to obtain
Corollary 13.3.
Given a black box (oracle) for the evaluation of a polynomial $p(x)$ of a degree $d$ with a relative error of any order in $1/d^{O(1)}$ (this holds under Assumption 0), we can perform our black box subdivision root-finders with a precision of $O(\log(d))$ bits.
Remark 13.1.
We derived our precision bound $O(\log(d))$ (in the case of root-finding in the unit disc $D(0,1)$) based on operating entirely with NIR$(x)$. For comparison, if we relied on computing $p(x)$ and on the bound $\min_{x:~{}|x|=1}(|p(x)|/\sum_{i=0}^{d}|p_{i}|)\geq(\frac{\theta-1}{2\theta})^{d}$ of [130, Eqn. (9.4)], we would have only obtained a bound of order $m^{2}d$ on the computational precision of our root-finders.
14 Approximation of matrix eigenvalues
Given a $d\times d$ matrix $M$ that has characteristic polynomial $p(x)=\det(xI-M)$, our root-finders can approximate all $m$ its zeros (that is, the eigenvalues of $M$) lying in a fixed disc $D$ on the complex plane isolated from external eigenvalues.
Recall Gershgorin’s bound for any eigenvalue $\lambda$ of a matrix $M=(a_{i,j})_{i,j=1}^{d}$ [134, Thm. 1.3.2]: $|\lambda-a_{i,i}|\leq\sum_{j\neq i}|a_{i,j}|$ for some $i$, $1\leq i\leq d$.
Hence, clearly, $|\lambda|\leq||M||_{F}$.
Combine this bound with the precision estimates of the previous section for our root-finders [^35] For a posteriori estimates, one can combine our root-finders with the recipe of doubling the precision of computing of (see Sec. 13.1). and with the following result of Storjohann [133] (cf. Remark 14.1):
Theorem 14.1.
One can evaluate the determinant of a $d\times d$ integer matrix $A$ by using an expected number of $O(d^{\omega}\log^{2}(d)\log(|A||_{F}))$ bit operations.
This implies a record bit operation complexity bound $\tilde{O}(m^{2}\tilde{b}^{2}d^{\omega})$ for any matrix $M$ which decreases whp by a factor of $m$ for matrices $M$ whose eigenvalues are iid random variables sampled under the uniform probability distribution from a large disc on the complex plane (cf. (1.7)) for approximation within $\epsilon$ of all $m$ eigenvalues of $M$ lying in the disc $D$ provided that $\tilde{b}=\log_{2}(||M||_{F}/\epsilon)$. This supports a bound exceeding (1.7) by a factor of $b$, to be removed with acceleration of our root-finding in Part III.
Remark 14.1.
Our root-finders are essentially reduced to evaluation of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ at sufficiently many points $x$. Hence we can rely on the following expressions for approximation of the eigenvalues of a matrix or a matrix polynomial:
$\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}((xI-M)^{-1})}$
where $M$ is a matrix and $p(x)=\det(xI-M)$ (cf. (1.13)) and
$\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}(M^{-1}(x)M^{\prime}(x))}$
where $M(x)$ is a matrix polynomial and $p(x)=\det(M(x))$ [19, Eqn. (5)]. [^36] These expressions have been obtained based on Jacobi’s formula ${\rm d}(\ln(\det(A)))={\rm trace}(A^{-1}{\rm d}A),$ which links differentials of a nonsingular matrix $A$ and of its determinant. Based on the above expressions one can reduce approximation of eigenvalues of a matrix or a matrix polynomial to its inversion. This computation tends to be more numerically stable than computing determinants and is significantly accelerated the highly important case where an input matrix or matrix polynomial can be inverted fast, e.g., is data sparse.
15 Bit operation complexity of general polynomial root-finding
Next we estimate the bit operation complexity of our root-finders, which accelerate classical subdivision root-finding by incorporating our e/i tests and are applied to a general polynomial $p$ of (1.1) whose coefficients are given with a precision sufficient to support our root-finding.
We could simply combine the precision bound $O(b\log(d))$ of Sec. 13 for our e/i tests with the estimate $O(bm)$ of Cor. 11.2 for the number of evaluation points for our root-finders, the bound of $2d$ ops for the evaluation of $p(x)$, and (1.8) for bit-operation cost of an op and then would arrive at the bit-operation complexity bound $\tilde{O}(b^{2}dm)$ for our root-finders. [^37] Recall that under the Random Root Model of Sec. 1.5 we can whp decrease this bound by a factor of $m$.
Next, however, we decrease this bound roughly by a factor of $d$ by using benefits of fast multipoint evaluation of a polynomial by means of the algorithm of Fiduccia [40] and Moenck and Borodin [81] (see [18, Sec. 4.5] for historical account), non-trivially extended in [78] to support the following theorem:
Theorem 15.1.
Given a positive $b$, the coefficients of a polynomial $t(x):=\sum_{i=0}^{d}t_{i}x^{i}$ such that $||t(x)||_{1}=\sum_{i=0}^{d}|t_{i}|\leq 2^{\tau}$, and $q$ complex points $x_{1},\dots,x_{q}$ in the unit disc $D(0,1)$, one can approximate the values $p(x_{1}),\dots,p(x_{q})$ within $1/2^{b}$ by using $\tilde{O}((d+q)(b+\tau))$ bit operations.
Proof.
For $q=d$ this is [78, Thm. 2], based on using novel data structure for hyperbolic approximation of $p$, proposed and analyzed by Guillaume Moroz in [78]. Trivially extend this theorem to the cases where $q>d$ (by considering $t(x)$ a polynomial of degree $q$ with $q-d$ leading coefficients 0) and where $q<d$ (by removing $d-q$ points $x_{i}$, with no increase of the complexity bound). ∎
We only apply this theorem for $\tau=0$, in which case its bit operation cost bound turns into
$\mathbb{B}_{\rm Moroz}=\tilde{O}((d+q)b)$ (15.1)
For comparison, the previous State of the Art was represented by the bound
$\mathbb{B}_{\rm Kirr}=\tilde{O}((d+q)(b+q)),$ (15.2)
due to [65, Thm. 3.9, Alg. 5.3, and Appendix A.3] by Peter Kirrinnis, which relied on combining his and Schönhage’s study of bit operation complexity of basic polynomial computations (cf. [130]) with the algorithm of [40, 81]. [^38] Precision of computing required in the latter algorithm grows fast as $d$ increases; e.g., the IEEE standard double precision of the customary numerical computations is not enough to support that algorithm already for $d>50$; for $b=O(\log(d))$ one can fix this deficiency by applying the Fast Multipole Method of Greengard and Rokhlin [48, 27, 101, 102] to the equivalent task of multiplication of Vandermonde matrix by a vector. One can, however, support bounds of Eqns. (15.1) and (15.2) on the overall bit operation complexity by computing with multiple precision for larger values $b$. Moroz obtained [78, Thm. 2] by applying bound (15.2) to his lower degree polynomials approximating $p(x)$.
The algorithms of [40, 81, 65, 78] extensively involve the coefficients of $p(x)$.
Bound (15.1) combined with the precision bound $O(\log(d))$ of Sec. 13 implies the bit operation cost bound $\tilde{O}((d+q)b)$ dominated by an estimated bit operation cost $\mathbb{B}$ for the evaluation of $p(x)$ with relative error $1/d^{O(1)}$ at $\tilde{O}(db)$ points. It remains to estimate $\mathbb{B}$, which is our next goal.
15.1 Reduction of our root-finders to multipoint polynomial approximation
Given a complex $c$, a pair of positive $b$ and $\rho$, and $d+1$ coefficients of a polynomial $p=p(x)$ of (1.1) such that the disc $D(c,\rho)$ is isolated and contains precisely $m$ zeros of $p$, we approximate all these zeros within $R/2^{b}$ for $R=|c|+\rho$ by applying subdivision root-finding iterations with e/i tests of Sec. 11. Then every subdivision step is reduced to application of our e/i tests to $\bar{m}=O(m)$ discs $D(c_{\lambda},\rho_{\lambda})$, where $\rho_{\lambda}\geq R/2^{b}$ for all $\lambda$. Such a test amounts to approximation of $p(x)$ at $q$ equally spaced points $x$ on each circle $C(c_{\lambda},\rho_{\lambda})$.
In Thm. 15.1 a polynomial $t(x)$ with $||t(x)||_{1}\leq 1$ is approximated at some points $x\in D(0,1)$. To ensure the assumption about $x$, we scale the variable $x\mapsto Rx$ to map the discs $D(c_{\lambda},\rho_{\lambda}\sigma)\mapsto D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})\subseteq D(0,1)$, for $\bar{c}_{\lambda}=c_{\lambda}/R$, $\bar{\rho}_{\lambda}=\rho_{\lambda}/R\geq 1/2^{b}$, and all $\lambda$.
Then we write $t(x):=p(Rx)/\psi$ for $\psi=O(R^{d})$ such that $||t(x)||_{1}=1$ and approximate $t(x)$ within $1/2^{\bar{b}}$ or equivalently within the relative error bound $1/(|t(x)|2^{\bar{b}})$ at $\bar{m}$ points $x$ in $\bar{m}$ e/i tests. [^39] Under the Random Root Model of Sec. 1.5 we can whp decrease this bound by a factor of $\bar{m}/\log(\bar{m})$ and then extend this decrease to the related subsequent estimates. Due to Cor. 13.3 it is sufficient to ensure the bound $1/(|t(x)|2^{\bar{b}})=1/d^{O(1)}$; we achieve this by choosing $\bar{b}\geq\log(\frac{1}{|t(x)|})+O(\log(d))$.
In our e/i tests we only need to approximate $t(x)$ at the points $x$ lying on $\theta_{\lambda}$-isolated circles where $\frac{1}{\theta_{\lambda}-1}=O(m)$ (cf. Assumption 1 of Sec. 13.2). Then $\log(\frac{1}{|t(x)|})=O(d\log(m))$ by virtue of the bound of [130, Eqn. (9.4)]). [^40] That bound follows from the estimate $|f|\geq 1/2^{d}$ of [130, Thm. 4.2], which holds provided that $t(x)=f\prod_{j=1}^{m}(x-x_{j})\prod_{j=m+1}^{d}(\frac{x}{x_{j}}-1)$, $||t(x)||_{1}=1$, $|x_{j}|<1$ for $j\leq m$, and $|x_{j}|>1$ for $j>m$.
Substitute $\bar{\rho}_{\lambda}\geq 1/2^{b}$ and obtain $\bar{b}=\log(\frac{1}{|t(x)|})+O(\log(d))=O(d\log(m)+b)$.
Remark 15.1.
For $m=d$, the above bound on $\log(\frac{1}{|t(x)|})$ is sharp up to a constant factor and is reached for the polynomial $t(x)=(x+1-\frac{1}{m})^{d}/(2-\frac{1}{m})^{d}$, $\rho_{\lambda}=1$, and $x=-1$. For $m<d$, [130, Thm. 4.5] implies a little stronger bounds, namely, $\log(\frac{1}{|t(x)|})=O(d+m\log(m))$ and hence $\bar{b}=O(m\log(m)+d+b)$ because all discs $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ lie in $\theta$-isolated unit disc $D(0,1)$ for $\theta-1$ exceeding a positive constant.
15.2 Bit operation cost of our e/i tests and root-finders
Apply Thm. 15.1 with $q$ and $b$ replaced by $\bar{q}=O(m^{2})$ and $\bar{b}=O(b+d+m\log(m))$, respectively, and obtain that the bit operation cost of our $O(m)$ e/i tests of Sec. 11 at any fixed subdivision step is in $\tilde{O}((b+d+m)m^{2})$ for $1\leq m\leq d$. This implies the upper estimate
$\mathbb{B}_{\rm roots}=\tilde{O}((b+d)(m^{2}+d)b)$ (15.3)
for the overall bit operation cost of our $O(b)$ subdivision steps. [^41] Under the Random Root Model of Sec. 1.5 we can decrease this bound by a factor of $m$ whp and can extend this decrease to the related subsequent estimates.
Based on [65, Alg. 5.3] and Eqn. (15.2) we obtain a little larger but still reasonably small upper bound
$\mathbb{B}_{\rm roots}^{\prime}=\tilde{O}(((m^{2}+d)(m^{2}+d+b)b).$ (15.4)
In Part IV we decrease the number of subdivision steps of our root-finders from order of $b$ to $O(\log(b))$, thus decreasing bounds (15.3) and (15.4) by a factor of $b$. As by-product we ensure $\epsilon$-isolation and arrive at our Main Theorems 2, 3, and 4.
PART IV: Compression of a Disc and Acceleration of Subdivision Root-Finders
16 Overview
Classical subdivision iterations converge to all zeros of a polynomial $p$ with linear convergences rate (see Sec. 1.6), but clearly this is very slow when they process a component $\mathbb{C}$ made up of a single suspect square or up to four of them sharing a vertex. In such cases the algorithms of [124, 97] dramatically accelerate convergence to superlinear rate by means of Newton’s or Schröder’s iterations (see Fig. 7), but the authors of [124, 97] have not controlled the bit operation cost of their algorithms.
Figure 8: Five blue marked centers of suspect squares converge to a green-marked root cluster with linear rate; two red marked Schröder’s iterates converge with quadratic rate.
Alternatively, one can apply algorithms of [130] to split $p(x)$ into the product of two polynomials $f(x)$ and $g(x)$ whose zero sets lie in and outside a disc covering the isolated component $\mathbb{C}$, respectively, and then apply a root-finder to the polynomials $f=f(x)$ and $g=g(x)$. Splitting of $p(x)$ into factors, however, does not preserve sparsity of $p(x)$, and the factors may have significantly larger coefficient length than $p(x)$. (E.g., splitting $p(x)=x^{1000}+1$ into two factors of degrees 300 and 700.) [^42] The near optimal root-finder of [92, 99] relies on by splitting $p(x)$ into factors of comparable degrees, but under the assumption of computing with a rather high precision of order $d\log(d)$. These problems disappear if one of the factors, say, $g$ has a small degree: one can deflate it, approximate its zeros, and then approximate the zeros of $f(x)=p(x)/g(x)$ by applying implicit deflation (cf. (1.14)).
We stay within subdivision framework and like [124, 97, 22, 23] partition the union of all suspect squares at each subdivision step into pairwise isolated connected components but propose a novel algorithm that compresses such a component $\mathbb{C}$ into a near minimal disc containing all roots that lie in the component.
This algorithm enables us to decrease the estimates of Secs. 8.3 and 15.2 for the overall complexity of root-finding by this factor, thus providing formal support for our Main Theorems 2-4.
We organize our presentation in this part as follows. In the next section we study a tree representing recursive partition of the union of suspect squares in a subdivision process into connected components. In Sec. 18 we devise and analyze our first compression algorithm. In particular we observe in Sec. 18.4 that the whole subdivision process involves at most $12m-$ 1 compression steps. In Sec. 19 we prove that in addition it only involves $O(m)$ e/i steps. Together with our estimates for the complexity of an e/i step this enables us to bound the overall complexity of our root-finders and reach near optimal bound on the bit operation complexity of our root-finders. We reach that bound in Sec. 20 by devising and analyzing our alternative compression algorithm. In that section we also present an alternative fast root-counter in an isolated disc.
We use some results of Part I, Alg. 5.1 of Part II, and Sec. 10 (Background) of Part III. Otherwise our presentation in this part of the paper is self-contained. Our fast root-counter of Sec. 20.4 can serve as alternative to root-counters of Part II.
17 Component tree and compact components
17.1 Component tree: definition
Partition the union of all suspect squares at a subdivision step into components; a step can break some of them into more than one component.
Represent all such partitions with a component tree $\mathbb{T}$.
At most $m$ components, said to be the leaves of the tree, can be covered by isolated discs of radius at most $\epsilon$ for a fixed tolerance $\epsilon\geq 0$. The other vertices are components-parents, each broken by a subdivision step into at least two components-children. The initial suspect square is the root of the tree $\mathbb{T}$. The edges of the tree connect parents and children. The tree does not shows modifications of components between their partitions represented by the edges.
Observation 17.1.
Having at most $m$ leaves, the tree $\mathbb{T}$ has at most $m-1$ other nodes. These upper bounds are sharp – attained where a subdivision process defines the complete binary tree $\mathbb{T}$ with $m$ leaves.
17.2 Edge length, non-adjacent suspect squares, and compact components
The edges of the tree $\mathbb{T}$ represent subdivision steps between partitions of components. Cor. 17.1 will bound the number of such steps; one can view this as the edge length.
Call a pair of suspect squares non-adjacent if they share no vertices.
Call a component compact if it contains just a single non-adjacent suspect square, that is, consists of at most four suspect squares sharing a vertex (see Fig. 8).
Observation 17.2.
(See Fig. 2.) The $\sigma$-covers of two non-adjacent suspect squares are separated by at least the distance $a\Delta$ for the side length $\Delta$ and separation coefficient
$a:=2-\sigma\sqrt{2}>0~{}{\rm for}~{}1<\sigma<\sqrt{2}.$
Lemma 17.1.
Let $|z_{g}-z_{h}|\geq\sigma\Delta$, for a pair of roots $z_{g}$ and $z_{h}$, the softness $\sigma$ of e/i tests, and the side length $\Delta$ of a suspect square, and let $i>0.5+\log_{2}(m)$. Then in at most $i$ subdivision steps the roots $x$ and $y$ lie in distinct components.
Proof.
Two roots $z_{g}$ and $z_{h}$ lying in the same component at the $i$th subdivision step are connected by a chain of $\sigma$-covers of at most $m$ suspect squares, each having diameter $\sigma\Delta/2^{i-0.5}$. Hence
$|z_{g}-z_{h}|\leq m\sigma\Delta/2^{i-0.5},~{}i\leq\log_{2}\Big{(}\frac{m\sigma\Delta}{|z_{g}-z_{h}|}\Big{)}+0.5.$
Substitute the bound $|z_{g}-z_{h}|\geq\sigma\Delta$ and deduce that $m\leq 2^{i-0.5}$ or equivalently $i\leq 0.5+\log_{2}(m)$ if the roots $z_{g}$ and $z_{h}$ are connected by a chain of suspect squares, that is, lie in the same component. This implies the lemma. ∎
Corollary 17.1.
A component is broken in at most $1+\lfloor 0.5+\log_{2}(m)\rfloor$ successive subdivision steps unless it is compact.
17.3 Isolation of compact components
Observation 17.3.
(i) A compact component is surrounded by a frame of discarded squares (see Fig. 8) and hence is covered by a 2-isolated square. (ii) If it stays compact in $h$ successive subdivision steps, then its minimal covering square and disc become $2^{h+1}$-isolated and $2^{h}\sqrt{2}$-isolated, respectively.
Proof.
Readily verify claim (i). Claim (ii) follows because a subdivision step can only thicken the frame of discarded squares while decreasing at least by twice the diameter of any component that stays compact. ∎
18 $\epsilon$-compression: algorithm and analysis
18.1 Compression algorithm: high level description
Recall that the classical subdivision iterations converge slowly where they successively process only compact components (cf. Fig. 7), but then we accelerate this process by interrupting it with compression steps in at most three iterations.
Next we specify compression for a 3.5-isolated unit disc $D=D(0,1)$ covering a component; map (1.4) extends compression to any 3.5-isolated disc $D(c,\rho)$; the map $\rho\epsilon\mapsto\epsilon$ adjusts the error tolerance $\epsilon$ to scaling by $\rho$. Observation 17.3 implies that we can readily ensure $\theta$-isolation at the cost of performing $O(\log(\theta))$ additional subdivision steps.
Algorithm 18.1.
$\epsilon$-compression (high level description).
INPUT: a positive $\epsilon\leq 1$ and a black box polynomial $p(x)$ of a degree $d$ such that [^43] Newton’s and Schröder’s iterations of [124, 97] accelerate convergence of subdivision iterations to a root cluster in a disc if the disc is $5d^{2}$-isolated, while (18.1) only requires 3.5-isolation.
$i(D(0,1))\geq 3.5~{}{\rm and}~{}\#(D(0,1))\geq 1.$ (18.1)
OUTPUT: an integer $m$, a complex $c^{\prime}$, and a positive $\rho^{\prime}$ such that $m=\#(D(0,1))$, (i) $X(D(c^{\prime},\rho^{\prime}))=X(D(0,1))$, and (ii) $\rho^{\prime}\leq\epsilon$ or $8|z_{g}-z_{h}|>5\rho^{\prime}$ for a pair of roots $z_{g},z_{h}\in D(c^{\prime},\rho^{\prime})$.
INITIALIZATION: Fix $\delta\geq 0$ such that $8\delta\leq\epsilon\leq 1$.
COMPUTATIONS: Compute $m=\#(D(0,1))$, $c^{\prime}$, $\rho^{\prime}$ such that
$|c^{\prime}-c|\leq\delta~{}{\rm and}~{}\rho\leq\rho^{\prime}\leq\max\{\epsilon,\rho+2\delta\}$ (18.2)
for $\rho:=r_{d-m+1}(c^{\prime})$ and some $c\in CH(X(D(0,1)))$.
18.2 Correctness of compression algorithm
First prove correctness in a simple case where $\epsilon=\delta=0$, $c^{\prime}=c$, $\rho^{\prime}=\rho$.
In this case $c^{\prime}=c\in CH(X(D(0,1)))\subset D(0,1)$, and so $\rho^{\prime}=\rho=r_{d+1-m}(c^{\prime})\leq 2$. Hence $D(c^{\prime},\rho^{\prime})\subseteq D(0,3)$.
Furthermore, #$\#(D(c^{\prime},\rho^{\prime}))=m$ since $\rho^{\prime}=\rho:=r_{d+1-m}(c^{\prime})$, while #$\#(D(0,1))=m$ by assumption.
Now (18.1) implies that #$\#(D(0,3))=m$, which supports claim (i) for the output of Alg. 18.1.
Next let $|c^{\prime}-z_{g}|=\rho:=r_{d+1-m}(c^{\prime})$ for a zero $z_{g}$ of $p(x)$. Then all $m$ roots in $D(0,1)$ also lie in the disc $D(c^{\prime},\rho)=D(c^{\prime},\rho^{\prime})$ by virtue of claim (i), just proved.
Since $c^{\prime}=c\in CH(X(D(0,1))$, however, there is a root, $z_{h}$, lying in $D(c^{\prime},\rho^{\prime})$ but not in $D(z_{g},\rho)$, such that $|z_{g}-z_{h}|>\rho\geq\rho^{\prime}-2\delta$ (cf. (18.2)).
Hence $|z_{g}-z_{h}|>3\rho^{\prime}/4$ (implying claim (ii) for the output of Alg. 18.1) because $42\delta\leq\epsilon/4<\rho^{\prime}/$ 4.
Next extend this correctness proof to the general case based on bounds (18.2).
Figure 9: Roots (asterisks) define compact components. A subdivision step halves their diameters and at least doubles isolation of their minimal covering squares and discs.
Notice that $c^{\prime}\in D(0,1+\delta)$ because $c\in CH(X(D(0,1)))\subset D(0,1)$ and $|c^{\prime}-c|\leq\delta$.
Hence $\rho=r_{d-m+1}(c^{\prime})\leq 2+\delta$, and so $\rho^{\prime}\leq\rho+2\delta\leq 2+3\delta$ (cf. (18.2)) and $D(c^{\prime},\rho^{\prime})\subseteq D(0,3+4\delta)$.
Therefore $D(c^{\prime},\rho^{\prime})\subseteq D(0,3.5)$ because $8\delta\leq\epsilon\leq 1$ by assumption.
Now recall that #$\#(D(0,3.5))=\#(D(0,1))=m$, while #$\#(D(c^{\prime},\rho^{\prime}))\geq\#(D(c^{\prime},\rho))=m$ for $\rho^{\prime}\geq\rho$ and deduce claim (i) for the output of Alg. 18.1.
Next wlog assume that $c=0$, $c^{\prime}=\delta$, and $z_{g}=c^{\prime}+\rho$ for a root $z_{g}$.
Then $\Re(z_{g})\geq\rho-\delta\geq\rho^{\prime}-3\delta$ because $|c^{\prime}|=|c^{\prime}-c|\leq\delta$ and $\rho^{\prime}\leq\rho+2\delta$ by assumptions.
Furthermore, $\Re(z_{h})<0$ for some root $z_{h}$ because $c=0\in CH(X(D(0,1)))$. (Actually, $\Re(z_{h})$ reaches its maximum $-\frac{z_{h}}{m-1}$ where $z_{h}$ is a simple root and $-\frac{z_{h}}{m-1}$ is a root of multiplicity $m-1$.)
Therefore, $|z_{g}-z_{h}|\geq\Re(z_{g})-\Re(z_{h})>\rho^{\prime}-3\delta\geq\rho^{\prime}-\frac{3}{8}\epsilon$.
This implies claim (ii) for the output of Alg. 18.1 since $\rho^{\prime}>\epsilon$.
18.3 Elaboration upon the compression algorithm
We can readily compute $m=\#(D(0,1))$ for isolated input disc $D(0,1)$ of Alg. 18.1, e.g., by applying Alg. 3.1, and can extend this recipe to any disc $D(c,\rho)$ by applying map (1.4) and Prop. 3.1. Moreover, in Sec. 20.4 we describe a very fast alternative root-counter.
Furthermore, based on Prop. 3.2 we can compute close approximations $\tilde{s}_{1}$ to the sum $s_{1}$ of the $m$ zeros of $p$ lying in that disc and then closely approximate their average $c^{\prime}:=s_{1}/m$, lying in the convex hull $CH(D(c,\rho))$. Finally, we can approximate the $m$th smallest root radius $r_{d-m+1}(c^{\prime},p)$ by applying our root radius estimation of Sec. 5, based on $\ell$-tests for $\ell=m$. There we can apply $\ell$-tests of Part II or combine e/i tests of Sec. 11 with reduction of $m$-test to e/i tests in Sec. 12.
Computation of a sufficiently close approximation to the sum $s_{1}$ of the zeros of $p$ lying in the disc $D(0,1)$ and of the average $s_{1}/m$ is too costly to support Thms. 1.1 and 1.3 (we would need to evaluate $p(x)$ at $O(b\log(d)$ points and similarly for bit operation cost, but we fix this deficiency in Sec. 20 by presenting another variant of compression algorithm.
18.4 Counting $\epsilon$-compression steps
$c^{\prime}$ approximates within $\epsilon$ all $m$ roots lying in the disc $D(0,1)$ if $\rho^{\prime}\leq\epsilon$ for the output value $\rho^{\prime}$ of Alg. 18.1. Then call the compression final and stop computations.
Otherwise call the square $S=S(c^{\prime},\rho^{\prime})$ suspect and apply to it subdivision iterations.
Extend Lemma 17.1 to the case where $|z_{g}-z_{h}|\geq\frac{5}{8}\rho^{\prime}$ for a pair of roots $z_{g},z_{h}\in D(c^{\prime},\rho^{\prime})$, combine it with claim (ii) for the output of Alg. 18.1, and obtain that at most $2.5+\log_{2}(8\sigma m/5)$ subdivision steps break the square $S$ into more than one component.
Until that moment apply no new compression steps and hence apply less than $2m$ such steps overall – no more than there are nodes in the partition tree $\mathbb{T}$.
Now elaborate upon Observation 17.1 as follows.
Theorem 18.1.
For a fixed $\epsilon\geq 0$ let subdivision and compression steps be applied to a component $C$ of suspect squares containing at most $m$ roots. Then at most $m$ final compression steps among at most $12m-$ 1 compression steps overall output isolated discs of radii at most $\epsilon$ that cover all roots lying in $C$.
Corollary 18.1.
Under the assumptions of Thm. 18.1 all subdivision steps process at most $36m-$ 3 compact components, made up of at most $1224m-$ 12 suspect squares; this involves at most $1224m-$ 12 e/i tests.
Proof.
A compact component stays unbroken and uncompressed in at most $3$ consecutive subdivision steps, which end with a compression step or follow a non-final compression step, that is, such components are processed in at most $36m-$ 3 subdivision , which involve at most $1224m-$ 12 e/i tests overall. ∎
19 Counting e/i tests
19.1 An upper estimate
The upper estimate $12m-$ 1 of Thm. 18.1 is sharp – reached in the case of the complete binary tree $\mathbb{T}$ and implying that our accelerated subdivision process can involve $12m-$ 1 suspect squares (e/i steps). This reaches lower estimate up to a constant factor:
Theorem 19.1.
Let subdivision iterations with compression be applied to a suspect square containing at most $m$ roots. Then they involve $O(m)$ e/i tests overall.
19.2 The first lemma
Our proof of Thm. 19.1 begins with
Lemma 19.1.
In every component $C$ made up of $\sigma(C)$ suspect squares we can choose $n(C)\geq\sigma(C)/5$ non-adjacent suspect squares.
Proof.
[…]
19.3 The second lemma
Lemma 19.2.
Fix two integers $i\geq 0$ and $g\geq 3$, a value $\sigma$ of softness, $1<\sigma<\sqrt{2}$, define the separation fraction $a:=2-\sigma\sqrt{2}$ of Observation 17.2. For all $j$ let $j$-th subdivision step process $v_{j}$ components made up of $w_{j}$ suspect squares, $n_{j}$ of which are non-adjacent, let $\Delta_{j}$ denote their side length, and let the $j$-th steps perform no compression for $j=i,i+1,\dots,i+g$. Then
$w_{i+g}\geq(n_{i}-v_{i})a2^{g-0.5}~{}{\rm if}~{}\Delta_{i}\geq 2^{g}\Delta_{i+g}.$ (19.1)
Proof.
At the $i$th step a pair of non-adjacent suspect squares of the same component is separated by at least the gap $a\Delta_{i}$, and there are at least $n_{i}-v_{i}$ such gaps with overall length $(n_{i}-v_{i})a\Delta_{i}$. Neither of these gaps disappears at the $(i+g)$th step, and so they must be covered by $w_{i+g}$ suspect squares, each of diameter $\sqrt{2}~{}\Delta_{i+g}=\Delta_{i}/2^{g-0.5}$ under (19.1). Hence $(n_{i}-v_{i})a\Delta_{i}\leq w_{i+g}\Delta_{i}/2^{g-0.5}$. ∎
By virtue of Cor. 18.1 all compact components together involve $O(m)$ suspect squares and e/i tests; it remains to prove that we only involve $O(m)$ other e/i tests, for which $n_{j}\geq 2$ for all $j$.
Next we extend Lemma 19.2 by allowing compression steps. We only need to extend the equation $2\Delta_{j+1}=\Delta_{j}$ to the case where the $j$th subdivision step for $i\leq j\leq i+g$ may perform compression, after which we apply subdivision process to a single suspect square with side length at most $4\Delta_{j}$.
The next two subdivision steps perform only five e/i tests, that is, at most $510m-$ 5 e/i tests are performed at all pairs of subdivision steps following all (at most $12m-$ 1) compression steps.
Then again, we can exclude all these e/i tests and bound by $O(m)$ the number of e/i tests in the remaining subdivision process, which is not interrupted with compression steps anymore.
Furthermore, we resume subdivision process after a compression step with a suspect square whose side length decreased (in two discounted subdivision steps) from at most $4\Delta_{j}$ to at most $\Delta_{j}$.
Hence we can apply to it Lemma 19.2, because now it accesses the same number $n_{j}$ of non-adjacent suspect squares and an increased number $\sigma_{j}$ of all suspect squares (cf. (19.1)).
19.4 Counting e/i tests at partition steps
Let $h$ be the height of our partition tree $\mathbb{T}$ and count suspect squares processed in subdivision steps that partition components, temporarily ignoring the other steps.
Clearly, bounds (19.1) still hold and can only be strengthened. Notice that $v_{i}\geq 1$ for all $i$, sum $v_{i+g}$ as well as bounds (19.1) for a fixed $g>0$ and $i=0,1,\dots,h-g$, and obtain
$\sum_{i=0}^{h-g}n_{i}-\lambda\sum_{i=0}^{h-g}\sigma_{i+g}\leq\sum_{i=0}^{h-g}v_{i+g}~{}{\rm for}~{}\lambda:=\frac{\sqrt{2}}{a2^{g}}.$ (19.2)
Notice that
$\sum_{i=0}^{h-g}v_{i+g}\leq\sum_{i=0}^{h}v_{i}-\sum_{i=0}^{g-1}v_{i}<2m-g,$
Combine the above estimate and the bounds $n_{i}\leq m$ and (cf. Lemma 19.1) $w_{i}\leq 5n_{i},$ for all $i$, write $n^{\prime}:=\sum_{i=0}^{h}n_{i}$, obtain
$\sum_{i=0}^{h-g}n_{i}\leq n^{\prime}-\sum_{i=h-g+1}^{h}n_{i}\leq n^{\prime}-mg,~{}\sum_{i=0}^{h-g}w_{i+g}\leq\sum_{i=0}^{h}w_{i}\leq 5n^{\prime},$
and so
[…]
Obtain that
$0.5n^{\prime}\leq 2m-1-g+mg~{}{\rm for}~{}g=1+\Big{\lceil}\log_{2}\Big{(}\frac{5\sqrt{2}}{a}\Big{)}\Big{\rceil}.$
Hence
$n^{\prime}\leq 4m-2-2g+2mg,~{}{\rm and~{}so}~{}n^{\prime}=O(m)~{}{\rm for}~{}g=O(1).$ (19.3)
19.5 Extension to counting all e/i steps
Next consider a non-compact component, count all suspect squares processed in all subdivision steps applied until it is partitioned, and then prove in Lemma 19.3 that up to a constant factor the overall number of these suspect squares is given by their number at the partition step. Such a result will extend to all subdivision steps our count $n^{\prime}=O(m)$ of (19.3), so far restricted to partition steps. Together with Lemma 19.1 this will complete our proof of Thm. 19.1.
Lemma 19.3.
Let $k+1$ successive subdivision steps with $\sigma$-soft e/i tests be applied to a non-compact component for $\sigma$ of our choice in the range $1<\sigma<\sqrt{2}$. Let the $j$-th step output $n_{j}$ non-adjacent suspect squares for $j=0,1,\dots,k$. Then $\sum_{j=0}^{k-1}n_{j}\leq n_{k}\frac{10\sqrt{2}}{a}$ for the separation fraction $a:=2-\sigma\sqrt{2}$ (cf. Observation 17.2.)
Proof.
Substitute $g=k-i$ into (19.3) and obtain $w_{k}\geq(n_{i}-v_{i})a2^{k-i-0.5}$, for $i=0,1,\dots,k-1$.
Recall that $w_{j}\leq 5n_{j}$ (cf. Lemma 19.1) and that in our case $v_{i}=1$ and $n_{i}\geq 2$ and obtain that $n_{i}-v_{i}\geq n_{i}/2$ for all $i$.
Hence $5n_{k}\geq n_{i}a2^{k-i-1.5}$ for all $i$, and so
$\sum_{i=0}^{k-1}n_{i}\leq n_{k}\frac{10\sqrt{2}}{a}\sum_{j=1}^{k}2^{-j}<n_{k}\frac{10\sqrt{2}}{a}.$
Remark 19.1.
The overhead constants hidden in the "O" notation of our estimates are fairly large but overly pessimistic. E.g., (i) the roots $x$ and $y$ involved in the proof of Lemma 19.2 are linked by a path of suspect squares, but we deduce the estimates of the lemma by using only a part of this path, namely, the part that covers the gaps between non-adjacent suspect squares. (ii) Our proof of that lemma covers the worst case where a long chains between the roots $x$ and $y$ can consist of up to $4m$ suspect squares, but for typical inputs such chains are much shorter and can consist of just a few suspect squares. (iii) In the proof of Lemma 19.3 we assume that $n_{i}-1\geq n_{i}/2$; this holds for any $n_{i}\geq 2$ but the bounds become stronger as $n_{i}$ increases. (iv) In a component made up of five suspect squares at least two are non-adjacent, and so the bound of Lemma 19.1 can be strengthened at least in this case.
Remark 19.2.
Alg. 18.1 compresses an input disc $D(0,1)$ into $\eta$-rigid disc for $\eta=5/8$, but our proof of Thm. 19.1 only involves $\eta$ the argument in Sec. 18.4, and this argument and therefore the whole proof of the theorem can be immediately extended if it relies on compression into any rigid disc.
20 Computation of a point near root set and root-counting in a disc
20.1 Computation of a point near root set: introductory comments
Recall Schröder’s iterations [128], celebrated for fast convergence to a root of multiplicity $m$ [77, Sec. 5.4]:
$z=y-m~{}{\rm NR}(y).$ (20.1)
Due to map (1.4) we only need to consider the case where the input disc $D$ is the unit disc $D(0,1)$. We first assume that it contains all $d$ zeros of $p$ and readily verify
Observation 20.1.
Schröder’s single iteration (20.1) for $m=d$ and $p(x)=p_{d}(x-z_{1})^{d}$ outputs the zero $z_{1}$ of $p(x)$ unless $y=z_{1}$.
Next we extend this observation by proving that the output $z$ of Schröder’s iteration (20.1) lies near the convex hull $CH(X(D))$ provided that $m=d$, a disc $D$ contain all $d$ zeros of $p$, and point $y$ lies far enough from the disc $D$. Here are our specific estimates.
20.2 The case where a disc contains all $d$ roots
Theorem 20.1.
Write $D:=D(0,1)$, $CH:=CH(D)$, and $\Delta:=\Delta(D)$, fix $y$ such that $|y|>2$, let $z:=y-d~{}{\rm NR(y)}$, which is (20.1) for $m=d$. Let $\bar{x}\in CH$. Then $|z-\bar{x}|\leq\alpha~{}\Delta$ for $\alpha:=\frac{(|y|-1)^{2}}{(y|-2)^{2}}$.
Proof.
Recall that $|\bar{x},z_{1},\dots,z_{d}\in D(0,1)$, $|z_{j}|\leq 1$ and $|\bar{x}-z_{j}|\leq\Delta$ for $j=1,\dots,d$. Hence (cf. (1.13))
$\Big{|}{\rm NIR}(y)-\frac{d}{y-\bar{x}}\Big{|}\leq\sum_{j=1}^{d}\Big{|}\frac{1}{y-z_{j}}-\frac{1}{y-\bar{x}}\Big{|}\leq\sum_{j=1}^{d}\frac{|\bar{x}-z_{j}|}{|y-\bar{x}|\cdot|y-z_{j}|}\leq\frac{\Delta d}{(|y|-1)^{2}}.$
Therefore,
$|u-w|\leq\beta~{}{\rm for}~{}u:=\frac{{\rm NIR}(y)}{d},~{}w:=\frac{1}{y-\bar{x}},~{}{\rm and}~{}\beta:=\frac{\Delta}{(|y|-1)^{2}},$ (20.2)
and so
$|z-\bar{x}|=\Big{|}\bar{x}-y+\frac{d}{{\rm NIR}(y)}\Big{|}=\Big{|}\frac{1}{u}-\frac{1}{w}\Big{|}=\frac{|u-w|}{|uw|}\leq\frac{\beta d}{|{\rm NIR}(y)|}|y-\bar{x}|.$ (20.3)
Deduce from (1.13) and the bounds $\max_{j=1}^{d}|z_{j}|\leq 1<|y|$ that
$|{\rm NIR}(y)|=\Big{|}\sum_{j=1}^{d}\frac{1}{y-z_{j}}\Big{|}\geq d\max_{j=1}^{d}\frac{1}{|y-z_{j}|}\geq\frac{d}{|y|-1}.$
Substitute this bound into Eqn. (20.3) and obtain that
$|z-\bar{x}|\leq\beta~{}||y|-1|\cdot|y-\bar{x}|$
Substitute $\beta:=\frac{\Delta}{(|y|-\Delta)^{2}}$ of (20.2) and obtain
$|\bar{x}-z|\leq\alpha^{\prime}\Delta~{}{\rm for}~{}\alpha^{\prime}:=\frac{(|y|-1)|y-\bar{x}|}{(|y|-\Delta)^{2}}.$
Substitute the bounds $\Delta\leq 2$ and $|\bar{x}|\leq 1$ and obtain the theorem. ∎
Since $|z-\bar{x}|\geq r_{d}(z,p)$, Thm. 20.2 implies
Corollary 20.1.
Under the assumptions of Thm. 20.2 it holds that $r_{d}(z)\leq 2.25\Delta$ for $|y|\geq 4$, while $r_{d}(z)\leq 25\Delta/16$ for $|y|\geq 6$.
Readily verify
Observation 20.2.
Let #$\#(D)=m$ for a disc $D$ and let $\Delta:=\Delta(CH(X(D)))$ denote the diameter of the convex hull of the root set $X(D)$ in that disc. Write $r_{d-m}(z,p):=0$, $r:=r_{d-m+1}(z,p)$, and $\rho:=r_{d}(z,p)$ for a complex $z$ and let $r+\Delta\leq r_{d-m}(z,p)$, which implies that $r$ and $\rho$ do not depend on the $d-m$ zeros of $p$ lying outside the disc $D$. Then $r-\rho\leq\Delta\leq 2r$.
Corollary 20.2.
If $\rho\leq\nu\Delta$ under the assumptions of Observation 20.2, then the disc $D(z,r)$ is $1/(2+2\nu)$-rigid.
Now we devise our fast compression algorithm.
Algorithm 20.1.
$\epsilon$-compression 1.
INPUT: a positive tolerance $\epsilon=1/2^{b}\leq 1$ and a black box polynomial $p(x)$ of a degree $d$ such that
#$\#(D(0,1))=d.$ (20.4)
OUTPUT: a complex $z$, and a positive $r$ such that (i) $X(D(z,r))=X(D(0,1))$ r equivalently #$\#(D(z,r))=d$, and (ii) $r\leq\epsilon$ or the disc $D(z,r)$ is rigid.
INITIALIZATION: Fix $y\geq 4$.
COMPUTATIONS:
(i) Compute and output $z=y-d{\rm NR}~{}(y)$ (cf. (20.1)).
(ii) Compute an approximation $\tilde{r}\geq 0$ by applying Alg. 5.1 for $m=d$, a constant $\beta>1$, and $\epsilon$ replaced by $\epsilon/4$).
If $\tilde{r}\leq 3\epsilon/4$, output $z$ and $r=\epsilon$, defining $\epsilon$-compression into the disc $D(z,r)$ of radius $\epsilon$.
Otherwise output $r=\tilde{r}+0.25\epsilon$, defining compression into the rigid disc $D(z,r)$.
Correctness of the algorithm is readily verified if Alg. 5.1 outputs a range $[\rho_{-}\rho_{+}]$ of length at most $\epsilon/4$, for in this case $r_{1}\leq r=\tilde{r}+\epsilon/4\leq\epsilon$, and #$\#(D(z,\epsilon))=d$.
Otherwise $1/\beta\leq\tilde{r}/r_{1}\leq\beta$ for a constant $\beta>1$ of Alg. 5.1, while the disc $D(x,r_{1})$ is $\frac{1}{2\nu+2}$-rigid by virtue of Cor. 20.2 where $\nu\leq 2.25$ by virtue of Cor. 20.1 since $y>4$ in our case. It follows that the disc $\bar{D}:=D(z,r_{1}(z))$ is rigid.
This compression queries the evaluation oracle twice while performing Schröder’s single iteration for computing $z$ and otherwise runs at the cost of performing Alg. 5.1, estimated in Cor. 5.1. We can decrease the latter cost bound for a large class of polynomials $p$ by choosing larger tolerance $\epsilon$. Then we yield desired compression if the algorithm outputs $r$ exceeding 1/4 of that tolerance value. Otherwise we can recursively repeat the computations for tolerance recursively decreased to $\epsilon/4$.
20.3 Extension to any well-isolated disc
Now suppose that #$\#(D(0,1))=m<d$ and $i(D(0,1))\geq\theta$ for a sufficiently large $\theta$ of order $d^{O(1)}$. Then verify that the $d-m$ external zeros of $p$, lying outside the disc $D(0,1)$, little affect the proof of Thm. 20.1 and our whole study in the previous subsection.
Thus we first perform $O(\log(d))$ subdivision steps (see Observation 17.3). We do not need compression if these steps break the component into more than one component or compress it into an $\epsilon$-disc.
Otherwise the component remains compact and hence in $O(\log(d))$ subdivision steps shrinks enough to become $d^{\nu}$-isolated for any fixed constant $\nu$, in which case we apply our compression algorithm.
20.4 A fast root-counter based on Schröder’s iterations
Let the unit disc be sufficiently strongly isolated and let a single Schröder’s iteration (20.1) be applied at a point $y$ lying sufficiently far from the unit disc $D(0,1)$.
Then our study in this subsection implies that the iteration outputs a point $z=z(m)$ lying in or near the disc $D(0,1)$, say, in the disc $D(0,2)$, if and only if #$\#(D(0,1))=m$, that is, we arrive at a root-counter in $D(0,1)$ that evaluates $p(x)$ at two points.
Indeed, if #$\#(D(0,1))=m$, then, as we have showed in this section, the iteration outputs a point $z=z(m)$ lying near the convex hull of the root set in $D(0,1)$ and hence lying in or near the disc $D(0,1)$, say, in the disc $D(0,2)$ if the value $|y|$ is large enough.
In this case $|m{\rm NR}(y)-y|\leq 2$, and so $|{\rm NR}(y)-y|/m\leq 2/m$.
If #$\#(D(0,1))=\bar{m}\neq m$, however, then the iteration outputs a point $z=z(m)$ that lies at the distance at least $|\bar{m}-m|\cdot|{\rm NR}(y)|-2$ from the disc $D(0,2)$, which must be positive if $|y|$ is sufficiently (although reasonably) large.
20.5 Overall complexity of our root-finders
We have reduced our root-finding in an isolated disc to performing at most $12m-$ 1 compression steps and in addition $O(m)$ e/i tests. We have also reduced a compression step to (i) performing $O(\log(d))$ additional subdivision steps (each included to yield isolation of an input disc for compression and reduced to $O(m)$ e/i tests), (ii) application of Schröder’s single iteration, which evaluates polynomial $p(x)$ at two points, and (iii) performing $O(\log(d))$ e/i tests for approximation of the root radius $r_{d-m+1}(z,p)$.
Summarize these estimates, combine them with the estimates of the previous subsection, and arrive at the complexity bounds of our Main Theorems.
Remark 20.1.
If we incorporate Kirrinnis’s algorithm for multipoint polynomial evaluation, [65, Alg. 5.3], rather than Moroz’s algorithm of [78], then our root-finders would slow down but still would stay very fast and still would use near optimal number of bit operations in the cases where $m^{2}=\tilde{O}(d)$ and whp under the Random Root Model. Indeed, by using our compression steps we decrease the complexity estimate of (8.4) and (15.4) by a factor of $b/\log(b)$ to arrive at the bound $\mathbb{B}_{\rm roots}^{\prime}=\tilde{O}(((m^{2}+d)(m^{2}+d+b))$ on the number of bit operations. This bound turns into $\tilde{O}((d+b)d)$ if $m^{2}=\tilde{O}(d)$ and whp under the Random Root Model.
APPENDIX
Appendix A Applications and implementation of root-finders: brief overview
Numerous areas of mathematics and computational mathematics studied in the last 100 years had no direct link to polynomial root-finding. Furthermore, many computational problems arising in the sciences, engineering, business management, and statistics have been linearized and then solved by using tools from linear algebra, linear programming, fast Fourier transform (FFT), and Fast Multipole Method (FMM). [^44] Both FFT and FMM are among the Top Ten Algorithms of 20-th century (cf. [26]). For the history of FFT see D.E. Knuth [66]; on FMM see L.F. Greengard and V. Rokhlin Jr. [48] and [147]. Such tools sometimes involve root-finding for polynomials but usually of smaller degree $d$, where the available subroutines are sufficiently effective in most cases.
Root-finding for large degree polynomials, however, still has important applications to various areas of Scientific Computing, in particular to Computer Algebra, signal processing, geometric modeling, control, financial mathematics, and complex dynamics (see McNamee and Pan [79, Introduction]).
Unlike the near optimal polynomial root-finders of [92, 99], various non-optimal complex polynomial root-finders have been implemented (see e.g., MATLAB “roots”, Maple “solve”, MPSolve, and CCluster). In particular, Imbach, P., and Yap have implemented the root-finder of [22] in [61].
Since its creation in 2000 the package MPSolve by Bini and Fiorentino [14], revised in 2014 by Bini and Robol [21], has been remaining user’s choice for approximation of all $d$ complex roots of a polynomial $p$. It implements Ehrlich’s iterations of 1967 [36], aka Aberth’s of 1973 [3], which empirically converge very fast globally – right from the start, under proper initialization, involving the coefficients of $p$. Such convergence has been consistently observed in decades-long worldwide applications of these iterations as well as in the case of some other root-finders by means of functional iterations, notably Weierstrass’s iterations [143], aka Durand and Kerner’s [29, 64]).
Unsuccessful attempts of supporting this empirical behavior formally (e.g., see [4, 51]), however, have recently ended with surprising proofs that Weierstrass’s and Ehrlich’s iterations diverge in the case of inputs that lie on some complex manifolds, being an open domain in the case of Weierstrass’s iterations (see Reinke, Schleicher, and M. Stoll [127, 125]).
Serious competition to MPSolve came in 2001 from the package EigenSolve by Fortune [41], but the later version of MPSolve [21] combines the benefits of both packages. An implementations of Newton’s iterations for approximation of all roots by Randig, Schleicher, and R. Stoll in [126] and by Schleicher and R. Stoll in [138] competed with MPSolve in 2017, but then MPSolve has regained upper hand by means of incorporation of Fast Multipole Method (FMM), even though such implementation is simpler and can be more efficient for Newton’s than Ehrlich’s iterations (see our Sec. D.4).
The root-finder of [22], efficiently implemented in [61], slightly outperforms MPSolve for root-finding in a region of the complex plane containing a small number of roots, although it is by far inferior for the approximation of all $d$ roots of a polynomial $p(x)$. [^45] The computational cost of root-finders of [92, 99, 22, 23] decreases at least proportionally to the number of roots in a region of interest such as a disc or a rectangle, while Ehrlich’s, Weierstrass’s, and a number of other functional iterations approximate the roots in such a region almost as slow as all complex roots. [57, 60] have already superseded MPSolve according to extensive tests for standard test polynomials, and our present root-finders are significantly and probably even dramatically faster (see Sec. 1.9).
Real polynomial root-finding is highly important because in many applications, e.g., to optimization in algebraic geometry and geometric modeling, only real roots of a polynomial are of interest and typically are much less numerous than all $d$ complex roots (cf. Kac [63], Edelman and Kostlan [37]). Current user’s choice for this task is the root-finder in [67], although it has been already superseded in [59] according to extensive tests for standard test polynomials. Our novel algorithm of Sec. 9 can be readily implemented and seems to be the first that efficiently solves the important numerical problem of approximation of both real and near real zeros of a polynomial.
Appendix B Polynomial Factorization, its links to Root-finding, and some bit-complexity estimates
B.1 Polynomial Factorization and its links to Root-finding (the results)
Hereafter write $|u|=\sum_{i=0}^{d}|u_{i}|$ for a polynomial $u(x)=\sum_{i=0}^{d}u_{i}x^{i}$.
The root-finders of [130, 92, 99] solve Problems 0 and $*{}^{*}$ by extending the the solution of the following problem of independent importance:
Problem F. Approximate Polynomial Factorization: [^46] The solution of Problem F can be also extended to isolation of the zeros of a square-free polynomial with integer coefficients and approximation of matrix eigenvalues (see our Remark C.4 and Sec. 7) and has various other applications to modern computations, e.g., to time series analysis, Wiener filtering, noise variance estimation, co-variance matrix computation, and the study of multichannel systems [145, 17, 8, 33, 34, 142]. For a polynomial of degree $d$ the known sharp upper and lower bounds on the input precision and number of bit operations for root-finding exceed those for numerical factorization by a factor of $d$. Given a real $b$ and $d+1$ coefficients $p_{0},p_{1},\dots,p_{d}$ of (1.1), compute $2d$ complex numbers $u_{j},v_{j}$ for $j=1,\dots,d$ such that
$|p-\prod_{j=1}^{d}(u_{j}x-v_{j})|\leq\epsilon|p|~{}{\rm for}~{}\epsilon=1/2^{b}.$ (B.1)
Hereafter $\mathbb{B}_{0}(b,d)$ and $\mathbb{B}_{F}(b,d)$ denote the optimal number of bit operations required for the solution of Problems 0 and F, respectively, for an error bound $\epsilon=1/2^{b}$, and next we specify a class of polynomials whose zeros make up no large clusters (cf. [13, 28]).
Definition B.1.
Polynomials with bounded cluster sizes. The zero set of $p$ consists of clusters, made up of a constant number of zeros, each having a diameter at most $1/2^{3-b/d}$; furthermore, these clusters are pairwise separated by the distances at least $1/d^{\tilde{O}(1)}$.
[92, 93, 95, 96, 99] proved the following near optimal bound for Problem F and readily extended it to near optimal solution of Problem 0.
Theorem B.1.
$\mathbb{B}_{F}(b,d)=\tilde{O}((b+d)d)$.
In Sec. B.3 we reduce Problem 0 to Problem F and vice versa by applying Schönhage’s results of [130, Cor. 4.3 and Sec. 19]:
Theorem B.2.
(i) $\mathbb{B}_{0}(b,d)=\tilde{O}(\mathbb{B}_{F}(bd,d))$ for any polynomial $p(x)$ of degree $d$, (ii) $\mathbb{B}_{F}(b,d)=\tilde{O}(\mathbb{B}_{0}(b+d+\log(d),d))$ for any polynomial $p(x)$ of degree $d$, and (iii) $\mathbb{B}_{0}(b,d)=\tilde{O}(\mathbb{B}_{F}2(b,d))$ for polynomials of Def. B.1, with bounded cluster sizes.
By combining Thms. B.1 and B.2 obtain the following estimates.
Theorem B.3.
One can solve Problem 0 by using (i) $\tilde{O}((b+d)d^{2})$ bit operations for any polynomial $p$ of (1.1) and (ii) $\tilde{O}((b+d)d)$ bit operations for a polynomial $p$ with bounded cluster sizes.
Corollary B.1.
One can solve Problem 0 by using $\tilde{O}((b+d)d+(b+m)m^{2})$ bit operation for $m$ of Def. 1.1 of Sec. 1.5. The bound turns into $\tilde{O}((b+d)d$ if $m^{2}=\tilde{O}(d)$.
Proof.
Proceed in two steps:
(i) by using $\tilde{O}((b+d)d)$ bit operations solve Problem F and then
(ii) apply the algorithm that supports Thm. B.3 to the $m$th degree factor of the polynomial $\prod_{j=1}^{d}(u_{j}x-v_{j})$ of (B.1) that shares with this polynomial all its zeros lying in the disc $D(1+\beta)$ for positive $\beta$ of Def. 1.1.474747We can slightly speed up stage (i) because at that stage we actually only need to approximate factorization of the factor $f$ of degree $m$ of $p$ that shares with $p$ its zeros in $D$.
B.2 Lower bounds on the bit operation complexity
Readily verify the following information lower bounds.
Observation B.1.
To solve Problem 0 or F one must access $(d+1)b$ bits of input coefficients and perform at least $0.5(d+1)b$ bit operations – at most one operation for two input bits.
Next strengthen the lower bound for Problem 0.
Observation B.2.
Let $p(x)=(x-v_{1})^{j}f(x)$ for an integer $j>0$ and a monic polynomial $f(x)$ of degree $d-j$. Fix a real $b$. Then the polynomial $p_{j}(x)=p(x)-2^{-jb}f(x)$ has the $j$ zeros $v_{1}+\zeta^{g}2^{-b}$ for $g=0,1,\dots,j-1$.
Observation B.2 implies that one must access the coefficient $p_{d-j}$ within $2^{-jb}$ to approximate within $2^{-b}$ a $j$-multiple zero of $p$ or its zero in a cluster of $j$ zeros, and so a root-finder must process at least $jb$ bits of the coefficient $p_{d-j}$. Sum these bounds for $j=1,\dots,d$ and obtain
Corollary B.2.
One must access at least $0.5(d-1)db$ bits of the coefficients of $p$ and perform at least $0.25(d-1)db$ bit operations to approximate even a single zero of $p$ within $2^{-b}$.
Likewise, deduce lower bounds on the number of bit operations required for Problem 0.
Corollary B.3.
To solve Problem 0 and even to approximate within $\epsilon$ a single zero of $p$ in the disc $D$, one must access at least $0.5b\max\{d+1,(m-1)m\}$ bits of input coefficients and perform at least $0.25b\max\{d+1,(m-1)m\}$ bit operations for $b=\log_{2}(R/\epsilon)$ and $R=|c|+\rho$.
The lower bounds of this subsection imply that the algorithm of [92, 99] and its extension supporting Thm. B.1 solve Problems 0 and F of large size by using optimal number of bit operations up to polylog factors or, in simple terms.
B.3 Linking Factorization and Root-finding (with more details)
Next we prove Thm. B.2, by applying the results of [130, Cor. 4.3 and Sec. 19].
Recall that $\mathbb{B}_{0}(b,d)$ and $\mathbb{B}_{F}(b,d)$ denote the optimal number of bit operations for the solution of Problems 0 and F, respectively, for an error bound $\epsilon=1/2^{b}$, and furthermore, recall the lower bounds
$\mathbb{B}_{0}(b,d)\geq 0.5(d+1)b~{}{\rm and}~{}\mathbb{B}_{F}(b,d)\geq 0.25(d-1)db$ (B.2)
of [92, 99] and the following result (cf. [130, Thm. 19.1]).
Theorem B.4.
Let $r_{1}=\max_{j=1}^{d}|z_{j}|\leq 1$. Then a solution of Problem 0 within an error bound $1/2^{b}$ can be recovered at a dominated bit operation cost from the solution of Problem F within an error bound $1/2^{bd+2}$, and so $\mathbb{B}_{0}(b,d)\leq\mathbb{B}_{F}(bd+2,d)$.
Based on [130, Cor. 4.3] we can also readily relax the assumption that $p(x)$ is monic.
Furthermore, in view of (B.2), relaxing the bound $r_{1}\leq 1$ would little affect the latter estimate because we can apply map (1.4), which changes the error bound of Problem 0 by at most $\log_{2}(\max\{\rho,1/\rho\})$, for $\rho\geq 1/2^{b}$.
[132, Thm. 2.7], which we reproduce in Sec. C.4 as Thm. C.3, shows an alternative way to relaxing both assumptions – that $r_{1}\leq 1$ and $p(x)$ is monic.
Schönhage in [130, Sec. 19] has greatly strengthened the estimate of Thm. B.4 unless “the zeros of $p$ are clustered too much". Namely, he proved (see [130, Eqn. (19.8)]) the following result.
Lemma B.1.
Assume that the polynomial $p(x)$ is monic, the values $u_{j}=1$ for all $j$ and complex $v_{1},\dots,v_{d}$ satisfy bound (B.1), and $|z_{j}-v_{j}|\leq r:=2^{2-b/d}$ for all $j$. Furthermore, write $\Delta_{j,1}:=|v_{j}-v_{1}|$ for $j=2,3,\dots,d$ and suppose that $1.25\Delta_{j,1}\leq r$ for $j=2,3,\dots,d$ and $\Delta_{1}:=\min_{j}\Delta_{j,1}>r$. Then $|z_{1}-v_{1}|\leq 2^{2-b/m}/|\Delta_{1}-r|^{\frac{d-m}{m}}$.
If $1/|\Delta-r|^{\frac{d-m}{m}}=2^{\tilde{O}(d)}$, then Lemma B.1 implies that $|z_{1}-v_{1}|\leq 1/2^{b/m-\tilde{O}(d)}$, and we can similarly estimate $|v_{j}-z_{j}|$ for all $j=1,2,\dots,d$.
By extending these estimates we strengthen Thm. B.4 as follows.
Corollary B.4.
Under bounding cluster sizes of Def. B.1, it holds that $|z_{j}-v_{j}|\leq 1/2^{b/m-\tilde{O}(d)},~{}j=1,2,\dots,d$. Hence if $m=O(1)$ is a constant, then a solution of Problem 0 within an error bound $1/2^{b}$ can be recovered at a dominated bit operation cost from the solution of Problem F with a precision $1/2^{\bar{b}}$ for a sufficiently large $\bar{b}=\tilde{O}(b+d)$, and so $\mathbb{B}_{1}(b,d)\leq\mathbb{B}_{2}(\tilde{O}(b+d),d)$.
For converse reduction of Problem F to Problem 0 recall [130, Cor. 4.3] in the case of $k=2$, covering the product of two polynomials:
Theorem B.5.
Let $p=gh$ for three polynomials $p$ of (1.1), $g$, and $h$. Then $||g||_{1}||h||_{1}\leq 2^{d-1}||p||_{1}$.
Now consider the products $p=gh$ for $g=x-z_{j}$, for $j=1,\dots,d$. For a fixed $j$, an error of at most $\epsilon/2^{d}$ for the zero $z_{j}$ contributes at most $\epsilon||h||_{1}$ to overall factorization error of Problem F. Therefore, $||h||_{1}\leq 2^{d-1}||p||_{1}/||x-z_{j}||_{1}\leq 2^{d-1}||p||_{1}$ by virtue of Thm. B.5. Hence the approximation errors of all the $d$ zeros of $p$ together contribute at most $2^{d-1}d\epsilon$ to the relative error of the solution of Problem F (ignoring dominated impact of higher order terms), and we arrive at
Corollary B.5.
If $b=O(b-d-\log(d))$, then Problem F for an error bound $1/2^{b}$ can be reduced at a dominated bit operation cost to Problem 0 for an error bound $1/2^{\widehat{b}}$ and a sufficiently large $\widehat{b}=O(b+d+\log(d))$, and so $\mathbb{B}_{F}(b,d)=O(\mathbb{B}_{0}(b,d))$.
It follows that the complexity estimate of Thm. 1.3 can be also applied to the solution of Problem F of factorization provided that $b=O(b-d-\log(d))$. Then again, the bit operation complexity of this solution of Problem F is optimal (up to polylogarithmic factors) under the random root model provided that $d=\tilde{O}(b)$ (cf. Observation B.1).
B.4 Two root-finders of 2016 and 2018
Becker et al in [22, 23] claimed that their extension of subdivision root-finders of [124, 97] (i) runs within the same near optimal bit operation bounds of [92, 99] but (ii) unlike [92, 99] is easy to implement. This seemed to diminish incentive for further research in this area, but were the claims valid?
Implementation of the root-finder of [22] by Imbach et al in [61] provided no empirical support to optimality claimed by Becker et al because the root-finder of [60], using some novel techniques from [108] (see [60, Sec. 1.2]), runs much faster. Unfortunately, formal support for optimality claimed by Becker et al has been missing as well.
Indeed, for all proofs the paper [22] refers to a preprint of [23], with a distinct list of the authors and with the main algorithms and the Main Theorem significantly different from those of [22].
Furthermore, as we pointed out in the introduction, the the bit-complexity specified in the Main Theorem of [23] has order of $mbd^{2}$, that is, exceeds the near optimal bound of [92, 99] by a factor of $d$.
[…]
Theorem B.6.
Given a polynomial $(x)$ of (1.1) with complex coefficients and $p_{g}=1$. Let the polynomial $s(x)=x^{d}+p_{d-1}x^{d-1}+\cdots+|p_{i+1}|x^{i+1}-|p_{i}|x^{i}|+p_{i-1}|x^{i-1}+\cdots+|p_{0}|$ have two distinct positive roots $r$ and $R$, $r<R$. Then the polynomial $p$ has exactly $i$ zeros in or on the circle $C(0,r)=\{z:~{}|z|=r\}$ and no zeros in the annulus $A(0,r,R)=\{z:~{}r<|z|<R\}$.
[22, 23] apply Pellet’s theorem on each subdivision step to more than $4m$ polynomials $t(x)=\sum_{i=0}^{d}t_{i}x^{i}$ obtained from the polynomial $p$ by means of shifting the variable $x$, its scaling by up to $2^{b}$, and root-squaring.
To apply Pellet’s theorem to $t(x)$ one must compute the values $|\frac{t_{i}}{t_{d}}|$ for $i=0,1,\dots,d-1$. Hence, by virtue of Observation B.2, one must process at least $0.5(d-1)db$ bits of the coefficients of $t(x)$ to support approximation of its zeros within $1/2^{b}$.
The algorithms of [23] cannot avoid this because they apply e/i tests to the polynomials obtained in $\log(\log(d))$ steps of root-squaring (cf. Sec. 10.3) from polynomials $t_{c,\rho}(x)=p(\frac{x-c}{\rho})$ where $\rho>0$ can be as small as $1/2^{b}$.
[…]
Appendix C Recursive factorization of a polynomial and extensions to root-finding and root isolation
In this section we briefly recall some algorithms and complexity estimates from [130, 132]. We first recursively decompose a polynomial into the product of linear factors and then extend the factorization to approximation and isolation of the roots.
C.1 Auxiliary norm bounds
Recall two simple lemmas.
Lemma C.1.
It holds that $|u+v|\leq|u|+|v|$ and $|uv|\leq|u|~{}|v|$ for any pair of polynomials $u=u(x)$ and $v=v(x)$.
Lemma C.2.
Let $u=\sum_{i=0}^{d}u_{i}x^{i}$ and $|u|_{2}=(\sum_{i=0}^{d}|u_{i}|^{2})^{\frac{1}{2}}$. Then $|u|_{2}\leq|u|$.
The following lemma relates the norms of a polynomial and its factors.
Lemma C.3.
If $p=p(x)=\prod_{j=1}^{k}f_{j}$ for polynomials $f_{1},\dots,f_{k}$ and $\deg p\leq d$, then $\prod_{j=1}^{k}|f_{j}|\leq 2^{d}|p|_{2}\leq 2^{d}|p|.$
Proof.
The leftmost bound was proved by Mignotte in [75]. The rightmost bound follows from Lemma C.2. ∎
Remark C.1.
[132, Lemma 2.6] states with no proof a little stronger bound $\prod_{j=1}^{k}|f_{j}|\leq 2^{d-1}|p|$ under the assumptions of Lemma C.3. From various factors of the polynomial $p(x)=x^{d}-1$ such as $\prod_{j=1}^{\frac{d}{2}}(x-\zeta_{d}^{j})$ for even $d$ and $\zeta:=\zeta_{q}:=\exp(\frac{2\pi{\bf i}}{q})$ denoting a primitive $q$-th root of unity, one can see some limitations on strengthening this bound further.
C.2 Errors and complexity
Suppose that we split a polynomial $p$ into a pair of factors over an isolated circle and recursively apply this algorithm to the factors until they become linear of the form $ux+v$; some or all of them can be repeated. Finally, we arrive at complete approximate factorization
$p\approx p^{\ast}=p^{\ast}(x)=\prod_{j=1}^{d}(u_{j}x+v_{j}).$ (C.1)
Next, by following [130, Sec. 5], we estimate the norm of the residual polynomial
$\Delta^{\ast}=p^{\ast}-p.$ (C.2)
We begin with an auxiliary result.
Theorem C.1.
Let
$\Delta_{k}=|p-f_{1}\cdots f_{k}|\leq\frac{k\epsilon|p|}{d},$ (C.3)
$\Delta=|f_{1}-fg|\leq\epsilon_{k}|f_{1}|,$ (C.4)
for some non-constant polynomials $f_{1},\ldots,f_{k},f$ and $g$ such that
$\epsilon_{k}d2^{d}\prod_{j=1}^{k}|f_{j}|\leq\epsilon.$ (C.5)
Then
$|\Delta_{k+1}|=|p-fgf_{2}\cdots f_{k}|\leq\frac{(k+1)\epsilon|p|}{d}.$ (C.6)
Proof.
$\Delta_{k+1}=|p-f_{1}\cdots f_{k}+(f_{1}-fg)f_{2}\cdots f_{k}|.$ Apply Lemma C.1 and deduce that $\Delta_{k+1}\leq\Delta_{k}+|(f_{1}-fg)f_{2}\cdots f_{k}|$ and, furthermore, that
$|(f_{1}-fg)f_{2}\cdots f_{k}|\leq|f_{1}-fg|~{}|f_{2}\cdots f_{k}|=\Delta|f_{2}\cdots f_{k}|.$
Combine the latter inequalities and obtain $\Delta_{k+1}\leq\Delta_{k}+\Delta|f_{2}\cdots f_{k}|$. Combine this bound with (C.3)-(C.5) and Lemmas C.1 and C.3 and obtain (C.6). ∎
Write $f_{1}:=f$ and $f_{k+1}=g.$ Then (C.6) turns into (C.3) for $k$ replaced by $k+1.$ Now compute one of the factors $f_{j}$ as in (C.4), apply Thm. C.1, then recursively continue splitting the polynomial $p$ into factors of smaller degrees, and finally arrive at factorization (C.1) with
$|\Delta^{\ast}|\leq\epsilon|p|$
for $\Delta^{*}$ of (C.2). Let us call this computation Recursive Splitting Process provided that it begins with $k=1$ and $f_{1}=p$ and ends with $k=d$.
Theorem C.2.
In order to support (C.3) for $0<\epsilon\leq 1$ and $j=1,2,\ldots,d$ in the Recursive Splitting Process it is sufficient to choose $\epsilon_{k}$ in (C.4) satisfying
$\epsilon_{k}~{}d~{}2^{2d+1}\leq\epsilon~{}\textrm{for~{}all}~{}k.$ (C.7)
Proof.
Prove bound (C.3) by induction on $j$. Clearly, the bound holds for $k=1$. It remains to deduce (C.6) from (C.3) and (C.7) for any $k$. By first applying Lemma C.3 and then bound (C.3), obtain
$\prod_{i=1}^{k}|f_{i}|\leq 2^{d}|\prod_{i=1}^{k}f_{i}|\leq 2^{d}~{}\Big{(}1+\frac{k\epsilon}{d}\Big{)}~{}|p|.$
The latter bound cannot exceed $2^{d+1}|{p}|$ for $k\leq d$, $\epsilon\leq 1$. Consequently (C.7) ensures (C.5), and then (C.6) follows by virtue of Thm. C.1.
Remark C.2.
The theorem implies that we can ensure the output precision bound $b^{\prime}$ by working with the precision of
$\bar{b}\geq\bar{b}_{\rm inf}=2d+1+\log_{2}d+b^{\prime}$
bits throughout Recursive Splitting Process.
C.3 Overall complexity of recursive factorization and extension to Root Isolation
The overall complexity of recursive factorization is dominated by the sum of the bounds on the complexity of all splittings into pairs of factors. The algorithm of [99] adopts recursive factorization of [130] but decreases the overall complexity bound by roughly a factor of $d$. This is due to additional techniques, which ensure that in every recursive splitting of a polynomial into the product of two factors the ratio of their degrees is at most 11. Consequently, $O(\log(d)$ recursive splitting steps are sufficient in [99], and so the overall complexity of root-finding is proportional to that of the first splitting. In contrast, the degrees of a factor in any splittin in [130] can have low, e.g., 1, and so there can be allows up to $d-1$ splittings overall, with each of the first $0.5d$ of them almost as expensive as the very first one.
C.4 From factors to roots
Theorem C.3.
[132, Thm. 2.7]. Suppose that
$p=p_{d}\prod_{j=1}^{d}(x-z_{j}),~{}\tilde{p}=\tilde{p}_{d}\prod_{j=1}^{d}(x-y_{j}),~{}|\tilde{p}-p|\leq\epsilon|p|,~{}\epsilon\leq 2^{-7d}$
and
$|z_{j}|\leq 1~{}{\rm for}~{}j=1,\dots,d^{\prime}\leq d,~{}|z_{j}|\geq 1~{}{\rm for}~{}j=d^{\prime}+1,\dots,d.$
Then up to reordering the roots it holds that
$|y_{j}-z_{j}|<9\epsilon^{\frac{1}{d}}~{}{\rm for}~{}j=1,\dots,d^{\prime};~{}\Big{|}\frac{1}{y_{j}}-\frac{1}{z_{j}}\Big{|}<9\epsilon^{\frac{1}{d}}~{}{\rm for}~{}j=d^{\prime}+1,\dots,d.$
By virtue of Thm. C.3 for $b^{\prime}=O(bd)$ we can bound the bit operation complexity of the solution of Problem 0 by increasing the estimate for the complexity of factorization in Sec. C.3 by a factor of $d$.
Corollary C.1.
The bit operation complexity of the solution of Problem 0. Given a polynomial $p$ of degree $d$ and a real $b^{\prime}\geq 7d$, one can approximate all zeros of that polynomial within the error bound $2^{-b^{\prime}}$ by using $O(\mu(b^{\prime}+d)d^{2}\log(d)(\log^{2}d+\log(b^{\prime})))=\tilde{O}((b^{\prime}+d)d^{2})$ bit operations.
Remark C.3.
The study in this section covers the solution of Problem 1 for the worst case input polynomial $p$. If its roots are isolated from each other, then the upper estimates of Thm. C.3 and Cor. C.2 can be decreased by a factor of $d$ (cf. [99]).
C.5 Polynomial Root Isolation
Based on his results recalled in this appendix, Schönhage in [130, Sec. 20] estimates the bit operation complexity of the following well-known problem by reducing it to Problem F.
Problem ISOL. Polynomial Root Isolation. Given a polynomial $p$ of (1.1) that has integer coefficients and is square-free, that is, has only simple roots, compute $d$ disjoint discs on the complex plane, each containing exactly one root.
Remark C.4.
Our Problem $*{}^{*}$ can be viewed as the Problem of Polynomial Root Cluster Isolation. It turns into Problem ISOL if all the coefficients of $p$ are integers and the roots are simple, and so our algorithms for Problem $*{}^{*}$ as well as the Algorithms of [92, 99] solve Problem ISOL within a record complexity bounds.
In [130, Sec. 20] Schönhage reduces to factorization Problem F the root-isolation problem of computing $d$ disjoint discs each covering a single simple root of a square-free polynomial with integer coefficients. Problem ISOL is readily solved if the zeros are approximated within $\epsilon=1/2^{b}<0.5\delta$ for $\delta$ denoting the minimal pairwise distance between the zeros of $p$ (see the known lower bounds on $\delta$ in Kurt Mahler and Maurice Mignotte [74, 75, 76]), and Schönhage has estimated that solution of Problem F (e.g., by means of the algorithms of [130, 90, 99]) implies isolation provided that
$b=O((\tau+\log(d))+d)~{}{\rm and}~{}\tau=\log_{2}\max_{i=0}^{d}|p_{i}|.$ (C.8)
This implied the following result.
Corollary C.2.
The bit operation complexity of polynomial root isolation. Suppose that a polynomial $p$ of (1.1) has integer coefficients and only simple roots. Let $\sigma_{p}$ denotes its root separation, that is, the minimal distance between a pair of its roots. Write $\epsilon:=0.4\sigma_{p}$ and $b^{\prime}:=\log_{2}(\frac{1}{\epsilon})$. Let $\epsilon<2^{-d}$ and let $m=m_{p,\epsilon}$ denote the maximal number of the roots of the polynomial $p(x)$ in $\epsilon$-clusters of its roots. Then one can solve Problem ISOL by using $\tilde{O}(bdm)$ bit operations for $b=\frac{b^{\prime}}{m}$.
Appendix D Newton’s, Weierstrass’s, and Ehrlich’s iterations
D.1 Contents
In the next subsection we comment on convergence of the three listed iterative root-finders and on their stopping criterion. In Sec. D.3 we discuss the policy of doubling precision, the resulting dynamic partition of the roots into tame and wild, and the deflation of a wild factor. In Sec. D.4 we accelerate these iterations by incorporating FMM. In Sec. D.5 we compute a disc that contains all zeros of a black box polynomial $p$, which one needs in customary initialization of root-finding for a black box polynomial by means of Newton’s iterations (cf. [138]). Customary initialization is more demanding for Weierstrass’s and Ehrlich’s iterations (see Sec. D.6).
D.2 Definition
1. Definition. Consider Newton’s, Weierstrass’s (aka Durand-Kerner’s), and Ehrlich’s (aka Aberth’s) iterations [68, 55, 126, 138, 143, 29, 64, 36, 3, 21]):
$z\leftarrow z-N(z),~{}z_{j}\leftarrow z_{j}-W(z_{j}),~{}z_{j}\leftarrow z_{j}-E_{j}(z_{j}),~{}j=1,\dots,d,~{}{\rm where}$ (D.1)
$N(x)=W(x)=E_{j}(x)=0~{}{\rm for~{}all}~{}j~{}{\rm if}~{}p(x)=0$ (D.2)
and otherwise
$N(z):=\frac{p(z)}{p^{\prime}(z)},~{}W(x):=\frac{p(x)}{l^{\prime}(x)},~{}\frac{1}{E_{j}(x)}:=\frac{1}{N(x)}-\sum_{i=1,i\neq j}^{d}\frac{1}{x-z_{i}}$ (D.3)
for $l(x):=\prod_{j=1}^{d}(x-z_{j})$.
Newton’s iterations can be initiated at a single point $z$ or at many points and can approximate a single root, many or all $d$ roots at the cost decreasing as the number of roots decreases [68, 55, 126, 138].
In contrast, Weierstrass’s and Ehrlich’s iterations begin at $d$ points for a $d$th degree polynomial and approximate a subset of its $d$ roots about as fast and almost as slow as they approximate all its $d$ roots.
Near the roots Newton’s and Weierstrass’s iterations converge with quadratic rate, Ehrlich’s iterations with cubic rate [140], but empirically all of them consistently converge very fast globally, right from the start, assuming a proper initialization, which involves the coefficients of $p$ (see [14, 21]). [^48] Formal support for such global convergence is much weaker for Newton’s iterations (see [68, 55, 138], and the references therein) and is completely missing for Weierstrass’s and Ehrlich’s iterations.
D.3 Doubling precision and the deflation of a wild factor
Computational (working) precision sufficient in order to control the relative error of an output approximation to a root can be defined by root’s condition number (see the relevant discussion in [14, 21]). This number is not readily available, however, and instead one can recursively double precision of computing, until a fixed output accuracy is verified: first apply a fixed root-finding iterations with the IEEE double precision of 53 bits and then recursively double it as long as this is feasible and enables more accurate approximation of the roots, stopping where the required output accuracy is ensured.
This policy for root-finding can be traced back to [14], where it was proposed for MPSolve. It turned out to be highly efficient and was adopted in [41] for Eigensolve, in [21] for the next version of MPSolve, and in [126, 138] for approximation of all roots by means of Newton’s iterations.
According to recipe (D.2) the computed approximation $z$ is only updated until $p(z)=0$ or, in numerical implementation, until the roots become tame, for a fixed stopping criterion. In the latter case, one should restrict updating to the wild roots (cf. Sec. 1.7) and should deflate the wild factor, if it has small degree. When a root is tamed the minimal working precision is exceeded by less than a factor of 2. Recall from [130] and [132, Thm. 2.7] that working precision does not need to exceed the output precision $b^{\prime}$ by more than a factor of $d$, and so at most $O(\log(db^{\prime}))$ steps of doubling precision are sufficient.
For every working precision the $d$ roots of $p(x)$ are partitioned into $d-w$ tame and $w$ wild roots as above; then one can apply explicit or implicit deflation (cf. (1.14)).
[…]
D.4 Acceleration with Fast Multipole Method (FMM)
Given $3d$ complex values $w_{j}$, $z_{i}$, $v_{j}$ such that $z_{i}\neq v_{j}$ for all $i,j=1,\dots,d$, Fast Multipole Method (FMM) by Greengard and Rokhlin [48] and by Carrier, Greengard, and Rokhlin [27] computes the sums $\sum_{j=1}^{d}\frac{w_{j}}{x_{i}-z_{j}}$ as well as $\sum_{j=1,j\neq i}^{d}\frac{w_{j}}{z_{i}-z_{j}}$ for all $i$ by using $O(db)$ arithmetic operations, with the precision of order $b$, versus $2d^{2}$ arithmetic operations in the straightforward algorithm. The paper [117] proposes and extensively tests its acceleration to Superfast Multipole Method. The papers [101, 103] list the recent bibliography on FMM and related computations with Hierarchical Semiseparable (HSS) matrices. Relevant software libraries developed at the Max Planck Institute for Mathematics in the Sciences can be found in HLib,
The acceleration based on FMM is practically valuable in spite of the limitation of the working and output precision to the order of $\log(d)$ and the resulting limitation on the decrease of the bit operation complexity.
Next we incorporate FMM into Newton’s, Weierstrass’s, and Ehrlich’s iterations. First relate the roots of a polynomial $p(x)$ to the eigenvalues of a diagonal-plus-rank-one matrix
$A=D+{\bf v}{\bf e}^{T}~{}{\rm for}~{}D=\operatorname{diag}(z_{j})_{j=1}^{d},~{}{\bf e}=(1)_{j=1}^{d},~{}~{}{\bf w}=(w_{j})_{j=1}^{d},~{}w_{j}=W(z_{j}),$
$W(x)$ of (D.3) and all $j$ (cf. [15]). Let $I_{d}$ denote the $d\times d$ identity matrix, apply the Sherman-Morrison-Woodbury formula, and obtain
$xI_{d}-A=(xI_{d}-D)(I_{d}-(xI_{d}-D)^{-1}{\bf w}{\bf e}^{T}).$
Hence
$\frac{1}{p_{d}}p(x)=\det(xI_{d}-A)=-\sum_{i=1}^{d}w_{i}\prod_{j\neq i}(x-z_{j})+\prod_{j=1}^{d}(x-z_{j}).$ (D.4)
Indeed, the monic polynomials of degree $d$ on both sides have the same values for $x=z_{j}$, $j=1,\dots,d$ (cf. [15])). [^50] Gerschgorin’s theorem (cf. [46]) for the matrix $xI_{d}-A$ defines neighborhoods of its eigenvalues (the roots of $p$), which helps to analyze and to modify the iterations (cf. [21]). Equations (D.4) imply that
$p(z_{i})=-w_{i}\prod_{j\neq i}(z_{i}-z_{j}),~{}p^{\prime}(z_{i})=\prod_{j\neq i}(z_{i}-z_{j})\Big{(}1-\sum_{j\neq i}\frac{w_{i}+w_{j}}{z_{i}-z_{j}}\Big{)},$
and therefore
$\frac{w_{i}}{N(z_{i})}=\sum_{j\neq i}\frac{w_{i}+w_{j}}{z_{i}-z_{j}}-1.$ (D.5)
Hence, given the $2d$ values of $z_{j}$ and $w_{j}=W(z_{j})$ for all $j$, one can fast approximate $N(z_{j})$ and consequently $E_{j}(z_{j})$ for all $j$ (cf. (D.3)) by applying FMM.
Every iteration but the first one begins with some values $z_{j}$, $w_{j}$, and updated values $z_{j}^{\rm(new)}$ of (D.1), for $j=1,\dots,d$, and outputs updated values $w_{j}^{\rm(new)}=W(z_{j}^{\rm(new)})$ for all $j$. Then equations (D.3) and (D.5) define $N(z_{j}^{\rm(new)})$ and $E_{j}(z_{j}^{\rm(new)})$ for all $j$. To update the values $w_{j}$, apply [21, equation (17)]:
$w_{i}^{\rm new}=(z_{i}^{\rm new}-z_{i})\prod_{j\neq i}\frac{z_{i}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}\Big{(}\sum_{k=1}^{d}\frac{w_{k}}{z_{i}^{\rm new}-z_{k}}-1\Big{)},~{}i=1,\dots,d.$
In the case where $z_{i}^{\rm new}=z_{i}$ if and only if $i>k$ this updating is still used for $i=1,\dots,k$ but is simplified to
$w_{i}^{\rm new}=w_{i}\prod_{j=1}^{k}\frac{z_{i}-z_{j}}{z_{i}^{\rm new}-z_{j}}~{}{\rm for}~{}i=k+1,\dots,d.$
Now fast compute the values $\sum_{k=1}^{d}\frac{w_{k}}{z_{i}^{\rm new}-z_{k}}$ for all $i$ by applying FMM.
Also compute the values
$R_{i}=(z_{i}^{\rm new}-z_{i})\prod_{j\neq i}\frac{z_{i}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}$
fast in two ways.
(i) Use the following customary trick (see [30, Remark 3]): approximate the values $\ln(R_{i})$ for $i=1,\dots,d$ fast by applying FMM and then readily approximate $R_{i}=\exp(\ln(R_{i}))$ for all $i$. [30] proves numerical stability of the reduction to FMM of a similar albeit a little more involved rational expressions.
(ii) Observe that
$\ln\Big{(}\frac{R_{i}}{z_{i}^{\rm new}-z_{j}}\Big{)}=\sum_{j\neq i}\ln\Big{(}1-\frac{z_{j}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}\Big{)}~{}{\rm for}~{}i=1,\dots,d$
and that
$\ln\Big{(}1-\frac{z_{j}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}\Big{)}\approx\frac{z_{j}-z_{j}^{\rm new}}{z_{i}^{\rm new}-z_{j}^{\rm new}}~{}{\rm if}~{}|z_{j}^{\rm new}-z_{j}|\ll|z_{i}^{\rm new}-z_{j}^{\rm new}|.$
Let the latter bound hold for $j\in\mathbb{J}_{i}$ and let $\mathbb{J}^{\prime}_{i}$ denote the set of the remaining indices $j\neq i$. Then
$R_{i}\approx\exp(\sigma_{i})\prod_{j\in\mathbb{J}^{\prime}_{i}}\Big{(}1-\frac{z_{j}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}\Big{)},~{}\sigma_{i}=\sum_{j\in\mathbb{J}_{i}}\frac{z_{j}-z_{j}^{\rm new}}{z_{i}^{\rm new}-z_{j}^{\rm new}},~{}i=1,\dots,d.$
By applying FMM fast approximate $\sigma_{i}$ for all $i$ and compute the values $\prod_{j\in\mathbb{J}^{\prime}}\frac{z_{i}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}$ for all $i$ in at most $4\sum_{i=1}^{d}~{}|\mathbb{J}_{i}^{\prime}|-d$ arithmetic operations for $|\mathbb{J}^{\prime}_{i}|$ denoting the cardinality of the set $\mathbb{J}^{\prime}_{i}$. Compare the latter cost bound with order of $d^{2}$ arithmetic operations in the straightforward algorithm.
Clearly the above updating computation is simpler and more stable numerically for $W(z_{j})$ versus $N(z_{j})$ and even more so versus $E_{j}(z_{j})$, for $j=1,\dots,d$.
Let us comment on the changes where one incorporates FMM into Newton’s, Weierstrass’s, and Ehrlich’s iterations implicitly applied to the wild factor $f(x)$ of $p(x)$. The same expression (1.14) defines Ehrlich’s iterations for both $f(x)$ and $p(x)$ by virtue of Thm. 10.1. Newton’s iterations to a wild factor one additionally subtract the sum $\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}$ (cf. (1.14)) and can apply FMM to accelerate numerical computation of this sum. To apply Weierstrass’s iterations to a wild factor one performs additional division by $t(x)$ and decreases the degree of the divisor $l^{\prime}(x)$ from $d-1$ to $w-1$.
D.5 Computation of a disc covering all zeros of a black box polynomial
D.5.1 Introductory comments
Given coefficients of a polynomial $p$ of a degree $d$, we can readily compute its largest and smallest root radii $r_{1}$ and $r_{d}$, respectively (cf. [39, 50]). Next we compute a range $[\rho_{-},\rho_{+}]$ that brackets all root radii or, equivalently, compute an upper estimate for the largest root radius $r_{1}$ and a low estimate for the smallest root radius $r_{d}$ of $p(x)$.
We can apply the estimates of Sec. 10.4 to the reverse polynomial $p_{\rm rev}(x)$ if its coefficients are available, but next we study this problem for a black box polynomial. Its solution is required, e.g., in order to initialize algorithms for all $d$ roots by means of functional iterations.
Suppose that we have an algorithm that computes a disc $D(0,\rho_{-})$ containing no zeros of a black box polynomial $p(x)$; clearly, $\rho_{-}\leq r_{d}$. By applying this algorithm to the reverse polynomial $p_{\rm rev}(x)$ we will obtain the upper bound $1/\rho_{-,{\rm rev}}$ on $r_{1}$.
It remains to describe computation of the disc $D(0,\rho_{-})$. Empirically we can do this by applying the algorithms of [109, 47], but next we study formally some alternative algorithms.
To obtain a range $[\rho_{-},\rho_{+}]$ that brackets all root radii, we need upper estimate for the root radii $r_{1}$ and $r_{d}$, but next we only estimate $r_{d}$ because we can extend them to $r_{1}$ by estimating $r_{d}$ for the reverse polynomial $p_{\rm rev}(x)$. We obtain a low bound on $r_{d}$ by computing a disc that contains no zeros of a black box polynomial $p(x)$.
Hereafter let $p(x)$ denote a black box polynomial.
D.5.2 Probabilistic solution
Seek a disc $D(c,\rho)$ containing no zeros of $p(x)$ in its interior, where we can choose any positive $\rho$ Sample a random $c$ in the unit disc $D(0,1)$ under the uniform probability distribution in $D(0,1)$.
The disc has area $\pi$, while the union of the $d$ discs $D(x_{j},\rho)$, for $j=1,2,\dots,d$, has area at most $\pi d\rho^{2}$. Therefore, $c$ lies in that union with a probability at most $d\rho^{2}$, which we can decrease at will by decreasing $\rho$. Unless $c$ lies in such a union, we obtain a desired disc $D(c,\rho)$.
By applying our e/i tests of Secs. 11.411.3 to this disc, we can verify correctness of its choice.
D.5.3 Deterministic solution
We can deterministically compute a desired disc $D(0,\rho_{-})$ at a higher cost, based on a soft e/i test.
[…]
D.6 $d$ root radii and initialization
The algorithms of [68, 55, 138] approximate all $d$ roots of a polynomial $p$ by applying Newton’s iterations at $f(d)$ equally spaced points on a large circle containing inside all $d$ roots for $f(d)$ of order at least $d\log(d)$. Bounds (10.6) immediately define such a circle for a polynomial given by its coefficients.
MPSolve further elaborates upon initialization. It first applies Bini’s heuristic recipe of [10], reused in [14, 21], for the approximation of the $d$ root radii by involving all coefficients of $p$. Given such approximations MPSolve defines $k$ concentric annuli $A(c,\rho_{i},\rho_{i}^{\prime})$ and $k$ integers $m_{i}$ such that #$\#(A(c,\rho_{i},\rho_{i}^{\prime}))=m_{i}$ for $i=1,\dots,k$, $k\leq d$, and $\sum_{i=1}^{k}m_{i}=d$. MPSolve initializes Ehrlich’s iterations at $d$ points $z_{j}$, $j=1,2,\dots,d$, exactly $m_{i}$ of them being equally spaced on the circle $C(c,(\rho_{i}+\rho_{i}^{\prime})/2)$. This heuristic initialization recipe turns out to be highly efficient, and one is challenged to test whether it can also help accelerate convergence of Newton’s iterations.
The algorithm supporting the next theorem also approximates the $d$ root radii but is certified, deterministic, and faster than Bini’s. It was first outlined and briefly analyzed in [130] for fast approximation of a single root radius and then extended in [88, Sec. 4] (cf. also [89, Sec. 5]) to approximation of all root radii of a polynomial. As in [10, 14, 21] it uses all $d$ coefficients of $p$. A variation of this algorithm in [59] supports the following theorem.
Theorem D.1.
[59, Prop. 3]. Given the coefficients of a polynomial $p=p(x)$, one can approximate all the $d$ root radii $|x_{1}|,\dots,|x_{d}|$ within the relative error bound $4d$ at a Boolean cost in $O(d\log(||p||_{\infty}))$.
[59] extends the supporting algorithm to approximation of all root radii within a positive relative error bound $\Delta$ by performing $\lceil\log_{2}\log_{1+\Delta}(2d)\rceil$ DLG root-squaring iterations (1.5).
Combine Thm. D.1 with estimates for the Boolean cost of performing these iterations implicit in the proof of [130, Cor. 14.3] and obtain
Corollary D.1.
Given the coefficients of a polynomial $p=p(x)$ and a positive $\Delta$, one can approximate all the $d$ root radii $|x_{1}|,\dots,|x_{d}|$ within a relative error bound $\Delta$ in $\frac{1}{d^{O(1)}}$ at a Boolean cost in $O(d\log(||p||_{\infty})+d^{2}\log^{2}(d))$.
The algorithm supporting the corollary and said to be Alg. D.1a computes at a low cost $d$ suspect annuli $A(0,r_{j,-},r_{j,+})$ for $j=1,\dots,d$ (some of them may pairwise overlap), whose union $\mathbb{U}$ contains all $d$ roots. The algorithm was the point of departure for amazing progress in [56], which extends the novel ingenious algorithm of [78] for multipoint polynomial evaluation.
We conclude this subsection with a list of some other benefits of fast approximation of the $d$ root radii.
(i) In a subdivision step one can skip testing exclusion and discard a square unless it intersects $\mathbb{U}$.
(ii) The search for real roots can be narrowed to the intersection $\mathbb{I}$ of $\mathbb{U}$ with the real axis (see Fig. 10). [122] has tested a simple real root-finder that approximates the roots in $\mathbb{I}$. According to these tests the algorithm is highly efficient for the approximation of all isolated real roots. Clearly, the well-developed root-finding techniques, e.g., ones using derivatives and higher order derivatives (see [99, 62], and the references therein) can further strengthen these algorithms. By complementing the root radii approach with some advanced novel techniques, the paper [59] has noticeably accelerated the real root-finder [67], currently of user’s choice. One can extend the progress to fast approximation of near real roots by combining root radii computation with the techniques of Sec. 9.
(iii) Initialization based on Alg. D.1a should improve global convergence of Ehrlich’s, Weierstrass’s, and other functional iterations for root-finding as well as of Kirrinnis’s factorization algorithm of [65]. In particular Ehrlich’s iterations of MPSolve are celebrated for their very fast global empirical convergence (right from the start), but it has only been observed with such initialization; moreover, they diverge without it in an example of [125].
[…]
We can apply this recipe to every annulus containing a small number of zeros of $p$ and isolated from its external zeros and finally approximate the remaining wild zeros of $p$ by applying implicit deflation (cf. (1.14)).
$x$$y$may contain rootsmay contain roots
Figure 10: Two approximate root radii define two narrow suspect annuli and four suspect segments of the real axis.

### v36

@@
-Good News for Polynomial Root-finding
+Fast Polynomial Root-Finding and Approximation of the Eigenvalues of a Matrix
 Abstract
 Univariate polynomial root-finding has been studied for four millennia and very intensively in the last decades.
-Our novel near optimal root-finders approximate all zeros of a polynomial $p$ almost as fast as one accesses its coefficients with the precision required for the solution within a prescribed error bound.
-[^1] Under Las Vegas randomization we allow output errors but detect them at a dominated computational cost.
-Furthermore, our root-finders can be applied to a black box polynomial, defined by an oracle (that is, black box subroutine) for its evaluation rather than by its coefficients.
-[^2] The pioneering root-finder by Louis and Vempala at FOCS 2016 is the only other known black box root-finder, but it only approximates the absolutely largest root of a real-rooted polynomial.
-Such root-finders are particularly fast for polynomials that can be evaluated fast such as the sum of a few shifted monomials of the form $(x-c)^{d}$.
-Our divide and conquer algorithm of ACM STOC 1995 is the only other known near optimal polynomial root-finder, but it extensively uses the coefficients, is quite involved, and has never been implemented, while according to extensive numerical experiments with standard test polynomials, already initial implementations of our new root-finders compete with user’s choice package of root-finding subroutine MPSolve and supersede it more and more significantly where the degree of a polynomial grows large.
-[^3] The root-finder of 1995 is deterministic, while our current ones apply Las Vegas randomization, that is, allow output errors but detect them at a dominated computational cost.
-Randomization enables us to avoid (with a high probability) division by near vanishing values in our computations with bounded precision.
-Our root-finders are readily extended to support approximation of the eigenvalues of a matrix within a record Las Vegas expected bit operation time bound.
+Our new near-optimal root-finders approximate all zeros of a polynomial $p$ almost as fast as one accesses its coefficients with the precision required for the solution within a prescribed error bound.
+Furthermore, our root-finders can be applied to a black box polynomial, defined by an oracle (black box subroutine) for its evaluation rather than by its coefficients.
+Due to this feature our root-finders (i) support approximation of the eigenvalues of a matrix in a record Las Vegas expected bit operation time and (ii) are particularly fast for a polynomial that can be evaluated fast such as the sum of a few shifted monomials of the form $(x-c)^{d}$ or a Mandelbrot-like polynomial defined by a recurrence.
+Our divide and conquer algorithm of ACM STOC 1995 is the only other known near-optimal polynomial root-finder, but it extensively uses the coefficients, is quite involved, and has never been implemented, while according to extensive numerical experiments with standard test polynomials, already a slower initial implementation of our new root-finders competes with user’s choice package of root-finding subroutines MPSolve and supersedes it more and more significantly as the degree of a polynomial grows large.
 Our auxiliary algorithms and techniques for computations with black box polynomials can be of independent interest.
 Keywords:
-polynomial computations, matrix eigenvalues, complexity, polynomial zeros, black box polynomials, root-squaring, symbolic-numeric computing.
+polynomial root-finding, matrix eigenvalues, computational complexity, black box polynomials, symbolic-numeric computing.
 2020 Math. Subject Classification:
 65H05, 26C10, 30C15
-1 Introduction
+PART I: BACKGROUND AND OUR PROGRESS
+1 State of the Art and our Main Results
+In this part of the paper we state our main results and outline our algorithms. We present them in some detail in Parts II, III, and IV, which can be read independently of each other.
 1.1 Polynomial Root-finding: Problems, Impact, and Earlier Studies
 According to [7, 99, 100, 83], univariate polynomial root-finding, that is, the solution of a polynomial equation $p(x)=0$, had been the central problem of Mathematics and Computational Mathematics for about four millennia since Sumerian times; it was still a central problem well into the 19th century.
-In its long way from its rudimentary level to its modern version, this study has been responsible for the introduction of mathematical rigor and notation as well as the concepts of irrational, negative, and complex numbers, algebraic groups, rings, and fields etc.
-Polynomial root-finding is still an important subject of modern computations, in particular for computer algebra, signal processing, geometric modeling, control, and financial mathematics [77, 79, 108].
+In a long way from its rudimentary level to modern version, its study has been responsible for the introduction of mathematical rigor and notation as well as the concepts of irrational, negative, and complex numbers, algebraic groups, rings, and fields etc.
 The existence of complex roots of the equation $p(x)=0$ is called the Fundamental Theorem of Algebra.
@@
 Already about 500 BC it was proved in Pythagorean school that the roots of the equation $x^{2}=2$ were irrational radicals, while even with the help of radicals one cannot express the roots rationally if the degree $d$ of $p(x)$ exceeds 4, according to Ruffini (1799,1813) and Abel (1824); moreover, one cannot do this already for specific equations such as $x^{5}-x-1=0$ or $x^{5}-4x-2$, according to Galois theory of 1830, 1832.
-Weyl in his constructive proof of Fundamental Theorem of Algebra in 1924 in [144], however, rationally approximated all $d$ zeros of $p$ within any fixed positive error tolerance.
-This progress, the advance of computer technology, and various applications to modern computations have led to the study of the following
-Root-finding problems: Given the $d+1$ complex coefficients of a polynomial $p(x)$ of a degree $d$, approximate within a fixed error tolerance $\epsilon$ either (i) its every zero (aka every root of the equation $p(x)=0$) or (ii) its every zero lying in a fixed Region of Interest $\mathbb{S}$ [^4] We count the zeros of $p$ with multiplicities and do not distinguish multiple roots from root clusters.
+Nevertheless, in 1924 in [149], Herman Weyl published his constructive proof of the Fundamental Theorem of Algebra where he rationally approximated within any fixed positive error tolerance all roots of $p(x)=0$ lying in a fixed square in the complex plane, that is, all roots if the square is large enough.
+This was the solution of the following Problem 0 of modern polynomial root-finding:
+INPUT: $d+1$ complex coefficients $p_{0},p_{1},\dots,p_{d}$ of a polynomial
+$p=p(x):=\sum_{i=0}^{d}p_{i}x^{i}:=p_{d}\prod_{j=1}^{d}(x-z_{j}),$ (1.1)
+a positive error tolerance $\epsilon=1/2^{b}>0$, an integer $m$, $1\leq m\leq d$, and a bounded domain $\mathbb{D}$ in the complex plane, e.g., a disc or a rectangle, such that $z_{j}\in\mathbb{D}$ if and only if $j\leq m$.
+OUTPUT: approximations $\tilde{z}_{j}\approx z_{j}$, $j=1,\dots,m$, to the $m$ zeros of $p(x)$ (aka roots of $p(x)=0$) [^1] Hereafter we freely say just “root” or “zero” for short, always count them with their multiplicities, and do not distinguish multiple roots from pairwise isolated root clusters of diameters less than $\epsilon$.
+lying in the domain $\mathbb{D}$ such that
+$|\tilde{z}_{j}-z_{j}|\leq\epsilon~{}{\rm for}~{}j=1,\dots,m.$ (1.2)
+Example 1.1.
+[Approximation of all roots.] Fix $R\geq\max_{j}|z_{j}|$ (see [52] for simple expressions for such $R$) and let $\mathbb{D}$ denote the disc $D(0,R):=\{x:~{}|x\leq R|\}$, centered in the origin and having radius $R$.
+Then Problem 0 requires us to approximate all $d$ roots.
+Example 1.2.
+[Numerical approximation of real roots.] Root-finding in the real line or its segment (real root-finding) is highly important because in many applications, e.g., to optimization in algebraic geometry and geometric modeling, only real roots of a polynomial are of interest, and typically they are much less numerous than all $d$ complex roots (cf.
+Kac [66], Edelman and Kostlan [39]).
+More generally consider Problem 0 where $\mathbb{D}$ denotes the minimal rectangle in the complex plane that covers the $\epsilon/\sqrt{2}$-neighborhood of a fixed real line segment.
+To solve this problem we should approximate both real and near-real roots, that is, their perturbations due to small input and rounding errors.
+The known real root-finders (cf.
+[80, Chs.
+2 and 3]) do not address this important issue.
+Major motivation for modern study of Problem 0 comes from application to approximation of the eigenvalues of a matrix (see our Secs.
+1.3 and 1.5) and the solution of a multivariate polynomial system of equations [29], both problems being highly important for the theory and practice of modern computations.
+[82, Introduction] specifies some applications of polynomial root-finding to computer algebra, signal processing, geometric modeling, control, and financial mathematics.
+Furthermore, Problem 0 for a domain $\mathbb{D}$ containing all roots is closely linked to polynomial factorization (see Secs.
+24 and 25), which has applications to time series analysis, Wiener filtering, noise variance estimation, co-variance matrix computation, and the study of multichannel systems [150, 18, 8, 35, 36, 147].
 Hundreds of efficient root-finders have been proposed [80, 82] and keep appearing (see [122, 22, 40, 70, 76, 127, 108, 23, 24, 64, 110, 111, 75, 60, 62, 81, 63, 59, 115], and the bibliography therein).
-Smale in [129] proposed to study Problem 1 in terms of arithmetic complexity and computational precision and later (jointly with Blum, Cucker, and Shub) (cf.
-[13, 28]) linked this study to rounding errors and conditioning of the roots.
-In [130] Schönhage proposed to estimate the complexity of polynomial root-finders in terms of bit operations involved and presented an algorithm that approximated all zeros of $p$ within $\epsilon=1/2^{b}$ by using $\tilde{O}(bd^{3})$ bit operations, that is, $bd^{3}$ bit operations up to a polylogarithmic factor in $bd$.
-Intensive study in the 1980s and 1990s has culminated in [92] (cf.
-also [93, 99] and surveys in [95, 96]) with our divide and conquer root-finder that decreased Schönhage’s record to $\tilde{O}((b+d)d^{2})$.
-This near optimal bound, that is, optimal up to polylogarithmic factors, implies that the root-finder runs nearly as fast as one accesses the coefficients with the precision required for the task.
-If we only consider polynomials whose zeros and tiny clusters of zeros are pairwise sufficiently isolated (as we specify in Def.
-B.1 of Sec.
-B), then the known lower bounds on the computational precision and the bit operation complexity of root-finding decrease by a factor of $d$, and so do the upper bounds supported by the algorithms of [92, 98] as well; hence the algorithms remain near optimal.
-1.2 Reduction of root-finding problems to the case of the unit disc
-Hereafter ${}^{T}$ shows transposition of a vector or a matrix, $\mathbb{R}^{k}$ and $\mathbb{C}^{k}$ denote the spaces of the column vectors $(v_{1},\dots,v_{k})^{T}$ of length $k$ with real and complex coordinates, respectively, and $D(c,\rho):=\{x:~{}|x-c|\leq\rho\}$ and $C(c,\rho):=\{x:~{}|x-c|=\rho\}$ denote the disc of radius $\rho>0$ centered at $c\in\mathbb{C}$ and its boundary circle, respectively.
-Next we link the above Root-finding problem (i) to its case (ii) where the Region of Interest $\mathbb{S}$ is the unit disc $D(0,1)$.
-Problem 0: Root approximation in the unit disc.
-INPUT: positive $\epsilon\in\mathbb{R}$, $d\in\mathbb{N}$, and $(p_{0},p_{1},\dots,p_{d})^{T}\in\mathbb{C}^{d}$, defining the polynomial
-$p=p(x):=\sum_{i=0}^{d}p_{i}x^{i}:=p_{d}\prod_{j=1}^{d}(x-z_{j}),$ (1.1)
-APPROXIMATE within $\epsilon$ all its zeros lying in the unit disc $D(0,1)$. [^5] We will also refer to $z_{1},\dots,z_{d}$ as “roots of equation $p=0$” or just “roots”.
-(i) Two ways from the zeros of $p$ lying in the unit disc to all its zeros.
-(a) Given an algorithm $\mathcal{A}$ that solves Problem 0, we can compute approximations $\tilde{y}_{j}$ within $\epsilon$ to the reciprocals $y_{j}=1/x_{j}$ of all zeros $x_{j}$ of $p$ lying outside the disc $D(0,1)$ [130] – simply by applying this algorithm to the reverse polynomial
+In [133] Smale proposed to study Problem 0 in terms of arithmetic complexity and computational precision and later, jointly with Blum, Cucker, and Shub (cf.
+[14, 30]), linked this study to rounding errors and conditioning of the roots.
+In [134] Schönhage proposed to estimate the complexity of polynomial root-finders in terms of the number of bit operations involved and presented an algorithm that solved Problem 0 for a domain that contains all roots by using $\tilde{O}(bd^{3})$ bit operations, [^2] Here and hereafter we assume that $b\geq d$, recalling that error amplification factors $2^{cd}$ are inevitable in polynomial root-finding, e.g., in routine shift of the variable.
+that is, $bd^{3}$ bit operations up to a polylogarithmic factor in $bd$.
+His work prompted intensive study of root-finding by researchers in the Theory of Computing in the 1980s and 1990s, culminated at ACM STOC 1995 with our near-optimal divide and conquer root-finder of [96] (also see its journal versions [97, 103] and surveys [99, 100]); it supports bound $\tilde{O}(bd^{2})$ and showed the information lower bound $0.25(d+1)bd$ (see Cor.
+24.2 in Sec.
+24.2).
+If we only consider polynomials whose zeros and tiny clusters of zeros are pairwise sufficiently isolated, as we specify in Def.
+24.1 of Sec.
+24, then the lower bounds on the computational precision and the bit operation complexity of root-finding decrease by a factor of $d$ (see Sec.
+24), and so do the upper bounds supported by the algorithm of [96, 97, 103] as well; hence the algorithm remains near-optimal (see Thms.
+24.1 and 24.4 and Cor.
+24.3 in Sec.
+24).
+1.2 Reduction to the case of the unit disc
+Hereafter T shows transposition of a vector or a matrix, $\mathbb{R}^{k}$ and $\mathbb{C}^{k}$ denote the spaces of the column vectors $(v_{1},\dots,v_{k})^{T}$ of length $k$ with real and complex coordinates, respectively, and $D(c,\rho):=\{x:~{}|x-c|\leq\rho\}$ and $C(c,\rho):=\{x:~{}|x-c|=\rho\}$ denote the disc of radius $\rho>0$ centered at $c\in\mathbb{C}$ and its boundary circle, respectively.
+Furthermore, let Problem $\widehat{0}$ denote Problem 0 where the domain $\mathbb{D}$ is the unit disc $D(0,1)$; we can extend its solution to root-finding in the disc $D(c,\rho)$ for any pair of complex $c$ and positive $\rho$ by means of applying the linear map
+$x\mapsto\frac{x-c}{\rho},~{}p(x)\mapsto t(x)=p\Big{(}\frac{x-c}{\rho}\Big{)},~{}D(c,\rho)\mapsto D(0,1),$ (1.3)
+which shifts the variable $x$ by $c$ and then scales $x-c$ by $1/\rho$.
+Next we reduce Problem 0 for any domain $\mathbb{D}$ to Problem $\widehat{0}$.
+(i) Two ways from the roots in the unit disc to all roots. Given a value
+$R\geq\max_{j=1}^{d}\{|x_{j}|\}$ (1.4)
+(see [52] for simple expressions for such $R$), we can scale the variable $x\mapsto y=\frac{x}{R}$, $p(x)\mapsto p(y)=p(\frac{x}{R})$ and approximate every root $x_{y}$, for $j=1,\dots,d$, within $\epsilon R$ by solving Problem $\widehat{0}$ for $p(y)$.
+Schönhage in [134] reduced approximation of all $d$ roots to Problem $\widehat{0}$ without assuming (1.4) or using (1.3) as follows.
+Given an algorithm $\mathcal{A}$ that solves Problem $\widehat{0}$, compute also approximations $\tilde{y}_{j}$ within $\epsilon$ to the reciprocals $y_{j}=1/z_{j}$ of all roots lying outside the disc $D(0,1)$ – simply by applying this algorithm to the reverse polynomial
 $p_{\rm rev}(x):=x^{d}p\Big{(}\frac{1}{x}\Big{)}=\sum_{i=0}^{d}p_{i}x^{d-i},~{}p_{\rm rev}(x)=p_{0}\prod_{j=1}^{d}\Big{(}x-\frac{1}{z_{j}}\Big{)}~{}{\rm if}~{}p_{0}\neq 0.$ (1.5)
-Write $\Delta_{j}:=\tilde{y}_{j}-y_{j}$ and deduce that $\frac{1}{y_{j}}-\frac{1}{\tilde{y}_{j}}=\frac{\Delta_{j}}{\tilde{y}_{j}~{}y_{j}}$, and so $\frac{\tilde{y}_{j}}{y_{j}}-1=\frac{\Delta_{j}}{y_{j}}=\Delta_{j}x_{j}$.
-Recall that $|\Delta_{j}|\leq\epsilon$ by assumption and obtain that $|\frac{\tilde{y}_{j}}{y_{j}}-1|\leq\epsilon|x_{j}|$, thus arriving at the following
+Write $\Delta_{j}:=\tilde{y}_{j}-y_{j}$ and deduce that $\frac{1}{y_{j}}-\frac{1}{\tilde{y}_{j}}=\frac{\Delta_{j}}{\tilde{y}_{j}~{}y_{j}}$, and so $\frac{\tilde{y}_{j}}{y_{j}}-1=\frac{\Delta_{j}}{y_{j}}=\Delta_{j}z_{j}$.
+Recall that $|\Delta_{j}|\leq\epsilon$ by assumption and obtain that $|\frac{\tilde{y}_{j}}{y_{j}}-1|\leq\epsilon|z_{j}|$, thus arriving at the following
 Proposition 1.1.
 The reciprocal of an approximation within $\epsilon$ to a zero of $p_{\rm rev}(x)$ in the disc $D(0,1)$ approximates the associated zero of $p(x)$ lying outside that disc within a relative error at most $\epsilon$.
-(b) If in addition to an algorithm $\mathcal{A}$ we are given a value
-$R\geq\max_{j=1}^{d}\{|x_{j}|\},$ (1.3)
-then we can scale the variable $x\mapsto y=\frac{x}{R}$, $p(x)\mapsto p(y)=p(\frac{x}{R})$ and then approximate every root $x_{y}$, for $j=1,\dots,d$, within $\epsilon R$ by solving Problem 0.
-(ii) From the zeros of $p$ in the unit disc to all zeros in a Region of Interest.
-For any pair of complex $c$ and positive $\rho$ we can reduce root-finding in the disc $D(c,\rho)$ to Problem 0 by means of applying the linear map
-$x\mapsto\frac{x-c}{\rho},~{}p(x)\mapsto t(x)=p\Big{(}\frac{x-c}{\rho}\Big{)},~{}D(c,\rho)\mapsto D(0,1),$ (1.4)
-which shifts the variable $x$ by $c$ and then scales $x-c$ by $1/\rho$.
-Now, given any bounded region of interest $\mathbb{S}$ on the complex plane, we can successively (a) cover it with a minimal disc $D(c,\rho)$ (see Remark 1.4 of Sec.
-1.6.2 for closer covers), (b) map this disc into the disc $D(0,1)$, (c) solve Problem 0 for the resulting polynomial $t(x)$ of (1.4), and finally, (d) apply map (1.4) in the reverse direction to obtain approximations to all zeros of $p$ lying in the intersection of the $\epsilon$-neighborhood $\mathbb{S}_{\epsilon}$ of the region $\mathbb{S}$ with the covering disc $D(c,\rho)$.
-1.3 Black Box Polynomial Root-finding and Eigenvalue Approximation: Problem Statements
-A nontrivial variant of Newton’s iterations in the FOCS 2016 paper [73] by Louis and Vempala approximates within a fixed positive $\epsilon\leq 1/2$ an absolutely largest zero $z$ of a real-rooted polynomial $p$ of a degree $d$ at the cost of its evaluation with a precision in $O((b+\log(R))\log(R~{}d)))$ at $O((b+\log(R))\log(d))$ points for $b:=\log_{2}(R/\epsilon)$ and a fixed $R$ satisfying (1.3).
-While the previous polynomial root-finders only worked for polynomials given in the monomial basis, that is, with their coefficients, the algorithm of [73] never involves them and works for a black box polynomial $p$ – given with an oracle (black box subroutine) for its evaluation.
-Next we modify Problem 0 accordingly.
-Problem 1: Black Box Polynomial Root Approximation.
-INPUT: a positive $\epsilon\in\mathbb{R}$, $d\in\mathbb{N}$, and a $d$th degree black box polynomial $p$.
-APPROXIMATE within $\epsilon$ the zeros of $p$ lying in the unit disc $D(0,1)$.
-By applying their root-finder to the characteristic polynomial $p(x)=\det(xI-M)$ of a $d\times d$ symmetric matrix $M$, Louis and Vempala approximated its absolutely largest eigenvalue by using a record Las Vegas expected number of bit operations.
-[^6] Under Las Vegas randomization one allows output errors but detects them at a dominated computational cost.
-Accordingly, we extend Problem 0 as follows:
+(ii) From roots in the unit disc to roots in a bounded domain.
+Given a bounded domain $\mathbb{D}$ on the complex plane, we can successively (a) cover it with a minimal disc $D(c,\rho)$, (b) map this disc into the disc $D(0,1)$, (c) solve Problem $\widehat{0}$ for the resulting polynomial $t(x)$ of (1.3), and finally, (d) use the map (1.3) again – this time to approximate all roots lying in the intersection of the disc $D(c,\rho)$ with the domain $\mathbb{D}$.
+In this way we also approximate extraneous roots lying in $D(c,\rho)$ but not in $\mathbb{D}$; in Sec.
+2.2 we control the computational cost of that additional work.
+1.3 Black box root-finders; the isolation of roots, eigenvalues, and their clusters
+A nontrivial variant of Newton’s iterations by Louis and Vempala [76], at FOCS 2016, approximates within a fixed positive $\epsilon\leq 1/2$ an absolutely largest zero of a real-rooted polynomial $p$ of a degree $d$ at the cost of its evaluation with a precision in $O((b+\log(R))\log(R~{}d)))$ at $O((b+\log(R))\log(d))$ points for $b:=\log_{2}(R/\epsilon)$ and a fixed $R$ satisfying (1.4).
+While the previous polynomial root-finders have been devised for polynomials given in the monomial basis, that is, with their coefficients, the algorithm of [76] never involves them and works for a black box polynomial $p$ – given with an oracle (black box subroutine) for its evaluation.
+[76] shows benefits of extension of such algorithms to approximation of the eigenvalues of a matrix.
+Next we state the relevant tasks.
+Problem 1 (Black Box Polynomial Root Approximation) is Problem 0 for a black box polynomial $p$.
 Problem 2: Approximation of the Eigenvalues of a Matrix. [^3] As in [76], the problem covers or can be extended to the case where an input is a polynomial matrix, aka matrix polynomial.
-INPUT: positive $\epsilon\in\mathbb{R}$, $d\in\mathbb{N}$, and a $d\times d$ matrix $M$.
-APPROXIMATE within $\epsilon$ the eigenvalues of $M$ lying in the unit disc $D(0,1)$.
-The algorithms for Problems 0, 1, and 2 may output multiple approximations within $\epsilon$ to the same roots and eigenvalues, but in the next modification we avoid such a confusion.
-Problems $0^{*}$, $1^{*}$, $2^{*}$: Isolation of Clusters of the Zeros of $p$ and the Eigenvalues of $M$.
-Under the assumptions of Problems 0, 1, and 2 compute at most $d$ disjoint discs of radii at most $\epsilon$ (hereafter $\epsilon$-discs) covering all zeros or eigenvalues that lie in the disc $D(0,1)$ and also compute the number of the zeros or eigenvalues lying in each $\epsilon$-disc.
+INPUT: $\epsilon=1/2^{b}>0$, integers $m$ and $d$ such that $0<m\leq d$, and a $d\times d$ matrix $M$, and a complex domain $\mathbb{D}$ that contains precisely $m$ eigenvalues of $M$ counted with their multiplicities.
+OUTPUT: Approximations within $\epsilon$ to each of all $m$ eigenvalues of $M$ lying in the domain $\mathbb{D}$.
+The output of Problems 0, 1, and 2 may include multiple approximations within $\epsilon$ to the same root or eigenvalue, but the next modifications exclude such a confusion.
+Problems $0^{*}$, $1^{*}$, $2^{*}$: Isolation of Clusters of the Zeros of a polynomial and the Eigenvalues of a matrix.
+Under the assumptions of Problems 0, 1, and 2 compute at most $d$ disjoint discs of radii at most $\epsilon$ (hereafter $\epsilon$-discs) covering all roots or eigenvalues that lie in the domain $\mathbb{D}$; also compute the number of the roots or eigenvalues lying in each $\epsilon$-disc.
 Remark 1.1.
 If we collapse all zeros of $p$ lying in every $\epsilon$-disc into a single multiple zero placed in the center of the disc, then these zeros will lie at the distances greater than $2\epsilon$ from each other.
-If polynomial $p$ is square-free (having only simple zeros) and has integer coefficients, then we would arrive at the solution of the well-known Problem of Root-isolation (see Problem ISOL and Remark C.4 in Sec.
-C.5).
-These comments can be immediately extended to Problem $*{}^{*}$.
-1.4 Some Benefits of Black Box Polynomial Root-finders
-(i) Black box polynomial root-finders are particularly efficient for $d$th degree polynomials $p$ that can be evaluated fast, say, in $O(\log(d))$ arithmetic operations (ops), such as Mandelbrot-like polynomials $p:=M_{k}(x)$, where $M_{0}:=x,~{}M_{i+1}(x):=xM_{i}^{2}(x)+1$, $i=0,1,\dots,k$, $d=2^{k}-1$, and the sums of a bounded number of shifted monomials, such as $p:=\alpha(x-a)^{d}+\beta(x-b)^{d}+\gamma(x-c)^{d}$ for six constants $a$, $b$, $c$, $\alpha$, $\beta$, and $\gamma$.
-For comparison, Horner’s algorithm uses $2d$ ops to evaluate at a point $x$ a general polynomial $p=p(x)$ of (1.1) given with its $d+1$ coefficients, and this bound is optimal (see [85, 66] or [135, Section “Pan’s method"]).
-One can evaluate $p$ at $d_{+}>d$ points and then interpolate to it, but this would destroy sparsity, blow up the precision and bit operation cost of the computations, and dramatically slow down root-finding in a region containing a small number of roots.
-(ii) The known efficient polynomial root-finders extensively shift (translate) and scale variable $x$ and in particular routinely apply map (1.4) to reduce root-finding in any disc $D(c,\rho)$ to root-finding in the unit disc $D(0,1)$.
-All this, however, is done at the price of losing sparsity and blowing up computational precision and bit operation complexity of root-finding.
-(iii) The same problems severely limit application of the DLG classical root-squaring iterations of Dandelin 1826, Lobachevsky 1834, and Gräffe 1837 (see [49, 82, 109] for the history), which compute polynomials $p_{2^{i}}(x):=\prod_{j=1}^{d}(x-z_{j}^{2^{i}})$, for $i=0,1,\dots$, as follows:
-$p_{0}(x):=\frac{p(x)}{p_{d}},~{}p_{2^{i+1}}(x^{2}):=(-1)^{d}p_{2^{i}}(x)p_{2^{i}}(-x),~{}i=0,1,\dots$ (1.5)
-Among various helpful properties of iterations (1.5), root-squaring also squares the distances from the origin to the roots, thus moving the external roots away from the unit circle $C(0,1)$, that is, strengthens its isolation from the roots.
-By avoiding computation of the coefficients of $t(x)$ and $p_{2^{i}}(x)$ of maps (1.4) and (1.5), respectively, black box polynomial root-finders keep all benefits of using these maps but avoid the cited notorious problems.
-For yet another benefit we can apply black box root-finders to polynomials in Bernstein form [71], whose evaluation is known to be numerically stable.
-Remark 1.2.
-Polynomial root-finding in the 19-th and well into the 20-th century relied on these iterations combined with Vieta’s formulas, which link the polynomial coefficients $p_{0},p_{1},\dots,p_{d}$ to the elementary symmetric polynomials of the roots.
-The combination enabled efficient approximation of the minimal and maximal distance of the roots from the origin and due to the map (1.4) from any complex point (see [82, 109, 111] for details).
-The approach became obsolete because of severe problems of numerical stability of the DLG iterations, but they disappeared in the case of root-finders that work for black box polynomials.
-Extensive tests in [111] with standard test polynomials have consistently demonstrated efficiency of a single particular implementation of this approach, but even the straightforward implementation of iterations (1.5) promises to be highly efficient as well.
-1.5 Our complexity estimates
-We adopt the model of [73] but otherwise rely on distinct approach and techniques (with few exceptions listed below).
-Indeed, the root-finder of [73] cannot help solve even Problem 1: even in the explicit setting, where the coefficients of $p$ are given, repeatedly deflating $p$, by peeling off large roots, would accumulate errors that grow too quickly, and it is not entirely clear how to even do this in the black box setting.
-Even if this recipe worked it would only solve the well-known and much easier problem of root-finding for a real-rooted polynomial (cf.
+If a polynomial $p$ is square-free (having only simple zeros) and has integer coefficients, then Problem 0 turns into the well-known Problem of Root-isolation (see Problem ISOL and Remark 25.4 in Sec.
+25.5).
+These comments can be immediately extended to Problems 1 and 2.
+1.4 More on Black Box Polynomial Root-finders
+1.4.1 Why we should use them
+One can perform black box evaluation at $q>d$ points $x_{1},\dots,x_{q}$ and then interpolate to $p$.
+The bit operation cost of interpolation, however, is unbounded unless the $q$ points are well isolated from the roots, but it is a challenge to yield such isolation.
+Furthermore, even if the isolation is ensured, the cost of interpolation exceeds the cost of the solution of Problems 0, 0, 1 ,1, 2, and 2 in the cases where domain $\mathbb{D}$ contains only a small number of roots.
+Moreover, here are three examples of important benefits of using black box algorithms even where an input polynomial is given with its coefficients.
+(i) Application of a black box root-finder to the characteristic polynomial $p(x)=\det(xI-M)$ of a $d\times d$ matrix $M$ only involves the values of this polynomial but not its coefficients, whose overall size exceeds the size (a norm) of the matrix $M$ by a factor of $d$.
+Consequently, extension of such a root-finder to Problem 2 saves a factor of $d$ in bit precision and bit operation count versus the known solutions of Problems 2 and 2, based on [134, Sec.
+21].
+(ii) Map (1.3) is a valuable root-finding tool, but its application is limited because it destroys sparsity and blows up computational precision and bit operation cost.
+Similar comments can be applied to maps (3.6) and (3.7) of Sec.
+3.4.
+Black box polynomial root-finders keep all benefits of using those maps but avoid the latter notorious problems by involving no coefficients of $t(x)$ and $p_{h}(x)$ of these maps.
+(iii) Black box polynomial root-finders are particularly efficient where a $d$th degree polynomial $p$ can be evaluated fast, say, in $O(\log(d))$ arithmetic operations (hereafter (ops) such as a Mandelbrot-like polynomial $p:=M_{k}(x)$, where $M_{0}:=x,~{}M_{i+1}(x):=xM_{i}^{2}(x)+1$, $i=0,1,\dots,k$, $d=2^{k}-1$, or the sum of a bounded number of shifted monomials, such as $p:=\alpha(x-a)^{d}+\beta(x-b)^{d}+\gamma(x-c)^{d}$ for six constants $a$, $b$, $c$, $\alpha$, $\beta$, and $\gamma$.
+For comparison, if we compute the coefficients of such a polynomial and then use them to evaluate it at a point $x$ by applying Horner’s algorithm, then we would need $2d$ ops, and this bound is optimal (see [89, 69] or [139, Section “Pan’s method"]).
+1.4.2 What we adopt and what cannot adopt for black box polynomial root-finding
+The known polynomial root-finders, except [76], extensively use the coefficients of $p$, and we must be careful in adopting their techniques for Problems 1 and 1.
+On the other hand, the root-finder of [76] cannot help solve even Problem 1.
+Indeed, repeatedly deflating $p$, by peeling off large roots, can accumulate errors that grow too quickly, together with the coefficient length of the factors (notice, e.g., the swell of the coefficients of $p(x)=x^{d}+1$ in transition to its factors of degree $d/2$, say).
+This is true even if one performs deflation in the black box setting implicitly, based on Eqn.
+(3.3) of Sec.
+3.
+Moreover, even in that setting the coefficient swell in deflation is translated into swelling of the values of the high order derivatives involved in the algorithm of [76].
+Even if the algorithm of [76] worked for Problem 1 it would only solve the well-known and much easier problem of root-finding for a real-rooted polynomial (cf.
 [93, 25, 33]).
 Moreover, the resulting algorithm for that much easier problem would involve more than $b^{2}d^{2}$ bit operations, which would by a factor of $b$ exceed the upper bound of [96, 103] for general polynomial root-finding.
-[^8] For root-finding problems one should assume that $b$ exceeds $d$ because the precision of representation of $p$ (input length) increases by a factor of $d$ in routine shifts and scaling of the variable $x$.
-Recursive combination of [73] with the Inverse Power Iterations [46] would enable approximation of all eigenvalues of a symmetric (although not unsymmetric) matrix, together with its eigenvectors.
-One can be challenged to estimate the increased computational complexity of the solution of this extended problem of matrix computations, but we study other approaches.
-In particular we accelerate approximation of $m$ zeros of $p$ for $m\leq d$ by a factor of $m$ by means of incorporation of fast multipoint polynomial evaluation of [40, 81, 65, 78], which we could not use if we successively approximated zeros of $p$ or matrix eigenvalues by using deflation.
-As in [73] we reduce root-finding to the evaluation of $p(x)$ at sufficiently many points.
-The ops involved in this evaluation are more numerous than the other ops involved in our algorithms, even in the case where $p(x)=x^{d}$.
-Unlike [73] we use Las Vegas randomization – to control with a high probability (whp) precision and bit-operation complexity of our computation.
-In particular we readily prove that whp $p(x)$ and $p^{\prime}(x)$ are not too close to 0 for a random variable $x$, and so we perform division safely whp when we approximate Newton’s ratio NR$(x):=\frac{p(x)}{p^{\prime}(x)}$ or its reciprocal – Newtons Inverse Ratio NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ at random points $x$.
-For these basic steps of our root-finders we fix a complex $\delta$ whose absolute value is sufficiently small in context and approximate $p^{\prime}(x)$ with the ratio $\frac{p(x)-p(x-\delta)}{\delta}$.
-This ratio is the first order divided difference, equal to $p^{\prime}(y)$ for some $y\in[x-\delta,x]$ by virtue of Taylor-Lagrange’s formula, [^9] Unlike [73] we do not need to approximate higher order derivatives and thus involve no higher order divided differences; by using Las Vegas randomization, however, we still ensure whp the same bound $O(b\log(d))$ of [73] on the computational precision.
-and so $y\mapsto x$ and $p^{\prime}(y)\mapsto p^{\prime}(x)$ as $\delta\mapsto 0$.
-Substitute the ratio for $p^{\prime}(x)$ and obtain that
-${\rm NIR}(x)\approx\frac{p(x)-p(x-\delta)}{p(x)\delta}=\frac{p^{\prime}(y)}{p(x)}~{}{\rm for~{}some}~{}y\in[x-\delta,x].$ (1.6)
-Our root-finders approximate the roots lying in the unit disc $D(0,1)$ together with some roots lying nearby.
-At the end we only keep the roots lying in the disc $D(0,1+\epsilon)$ but at the intermediate steps may approximate some extraneous roots in a larger domain, but we modify subdivision in Sec.
-1.6.2 so that all extraneous roots lie in the disc $D(0,1+\beta)$ for any fixed positive $\beta$.
+Recursive combination of [76] with the Inverse Power Iterations [48] could lead to approximation of all eigenvalues of a symmetric (although not unsymmetric) matrix, together with its eigenvectors.
+One can be challenged to estimate the increased complexity of this extended computations, but we study other approaches.
+By applying a root-finder to Given a polynomials $p$ of (1.1) given with its coefficients We accelerate known algorithms for Problem 0 by a factor of $m/\log^{2}(m)$ by incorporating fast multipoint polynomial evaluation algorithms of [42, 84, 68, 81], but we could not use them if we successively approximated zeros of $p$ or matrix eigenvalues by using deflation.
+Thus, instead of applying high order Newton’s iterations of [76] and deflation, we recall and extend subdivision iterations of [149, 52, 128, 101].
+1.5 Our complexity estimates
 Definition 1.1.
-Hereafter $m=m(\beta)$ denotes the number of the zeros of $p$ lying in the disc $D(0,1+\beta)$ for a fixed positive constant $\beta$ of our choice.
-[^10] See our Cor.
-1.2 and compare more restrictive choice of $\beta=2$ in [22, 23].
-Our Las Vegas randomized root-finders support the Main Theorems 1-4 below.
+$m=m(\mathbb{D},\beta)$ for $\beta>0$ is the number of roots of polynomial $p(x)$ in the $\beta$-neighborhood of a domain $\mathbb{D}$ on the complex plane.
+Main Theorems 1-4 below hold for $m$ of Def. 1.1, any fixed constant $\beta>0$, and either $p(x)=\det(xI-M)$ in the case of Problems 2 and 2 or $p(x)$ of (1.1) otherwise.
 Theorem 1.1.
-(Main Theorem 1.) One can solve Problems 1 and $*{}^{*}$ by applying Las Vegas randomized queries of the polynomial evaluation oracle with a computational precision $O(b\log(d))$ at $\tilde{O}(dm)$ locations for $b:=\log_{2}(1/\epsilon)$ and $m$ of Def.
-1.1.
-By following [73] one can readily extend Main Theorem 1 to matrix eigenvalue approximation:
+(Main Theorem 1.) One can solve Problems 1 and 1 by querying the polynomial evaluation oracle at $\tilde{O}(m^{2})$ locations with a computational precision $O(b\log(d))$.
 Theorem 1.2.
-(Main Theorem 2.) One can solve Problems 2 and $*{}^{*}$ at Las Vegas randomized bit operation cost
-$\mathbb{B}_{\rm eigen}=\tilde{O}(\tilde{b}d^{\omega}m)$ (1.7)
-for $m$ of Def.
-1.1, $\omega$ denoting an exponent of feasible or unfeasible matrix multiplication (MM),  [^11] The current records are about 2.7734 for exponents of feasible MM [86, 103, 105, 69, 54], unbeaten since 1982, and about 2.372 for exponents of unfeasible MM [6, 35].
-The exponents below 2.77 are only reached for unfeasible MM – supported by algorithms that are slower than the classical $d\times d$ MM unless $\log_{2}(\log_{2}(\log_{2}(d))))>100$, say.
-and $\tilde{b}:=\log_{2}(||M||_{F}/\epsilon)$ for the Frobenius norm $||M||_{F}$ of a matrix $M$.
-Thm.
-1.1 implies the bound $\tilde{O}(bd^{2}m)$ on Las Vegas randomized bit operation complexity of Problem 1 because one can evaluate any polynomial $p$ of degree $d$ in $2d$ ops and perform an op with the precision of $s$ bits by using $O(s\log(s)\log(\log(s)))=\tilde{O}(s)$ bit operations, which extends the celebrated bound
-$\mu(s):=O(s\log(s)\log(\log(s)))=\tilde{O}(s)$ (1.8)
-on the number of bit operations for multiplication of two integers modulo $2^{s}$ [137] (see [66, 130, 131] for this extension).
-In the case of the Mandelbrot polynomials $p:=M_{k}(x)$ and the sums of a constant number of shifted monomials one can immediately decrease by a factor of $d/\log(d)$ both of the above upper bounds of $2d$ ops and $\tilde{O}(bd^{2}m)$ bit-operations.
-Even in the worst case of general polynomial $p$ given with its coefficients, however, we decrease the upper bound on the bit operation cost of our root-finders by roughly a factor of $d$ – to match the near optimal bound of [92, 99] in our next theorem.
-We achieve this by combining our root-finders supporting Main Theorem 1 with the known algorithms for fast multipoint polynomial evaluation [40, 81, 65, 78].
+(Main Theorem 2.) One can solve Problems 2 and 2 at Las Vegas randomized bit operation cost
+$\mathbb{B}_{\rm eigen}=\tilde{O}(\tilde{b}d^{\omega}m^{2})$ (1.6)
+for $\tilde{b}:=\log_{2}(||M||_{F}/\epsilon)$, the Frobenius norm $||M||_{F}$ of a matrix $M$, and $\omega$ denoting an exponent of feasible or unfeasible matrix multiplication (MM).
+[^4] The current records are about 2.7734 for exponents of feasible MM [90, 107, 109, 72, 57], unbeaten since 1982, and about 2.372 for exponents of unfeasible MM [6, 37].
+The exponents below 2.77 have only been reached at the expense of blowing up the size of MM by immense factors.
 Theorem 1.3.
-(Main Theorem 3.) One can solve Problems 0 and $*{}^{*}$ at Las Vegas randomized bit operation cost $\tilde{O}((b+d)dm)$ for $b:=\log_{2}(1/\epsilon)$, $m$ of Def.
-1.1, and general polynomial $p$ of (1.1) normalized such that $||p(x)||_{1}:=\sum_{i=1}^{d}|p_{i}|=1$.
-Finally, we decrease our upper bounds of Thms.
-1.1-1.3 by a factor of $d$ under an additional assumption on the disposition of the zeros of $p$ and eigenvalues of $M$, namely, under the following model: [^12] Its assumption is a little stronger than that of our Def.
-B.1, supporting such a decrease in [92, 99], but it is in good accordance with the results of extensive tests in [57, 60] for standard test polynomials (see Sec.
-13.4.1 and the beginning of Sec.
-4.1).
-Random Root Model: The zeros of $p$ (in particular for $p$ denoting the characteristic polynomial of a matrix $M$) are independent identically distributed (iid) random variables sampled under the uniform probability distribution on the disc $D(0,R)$ for a large $R\in\mathbb{R}$ such that the value $\frac{R^{2}}{d}-1$ exceeds a positive constant.
+(Main Theorem 3.) One can solve Problems 0 and 0 at Las Vegas randomized bit operation cost $\tilde{O}(bdm))$ for a polynomial $p$ of (1.1) normalized such that $||p(x)||_{1}:=\sum_{i=1}^{d}|p_{i}|=1$.
+Random Root Model: The zeros of $p$ (in particular for $p$ denoting the characteristic polynomial of a matrix $M$) are independent identically distributed (iid) random variables sampled under the uniform probability distribution on the disc $D(0,R)$ for a large $R\in\mathbb{R}$ such that the value $\frac{R^{2}}{d}-1$ exceeds a positive constant.555(i) The assumption of the model is a little stronger than that of our Def.
+24.1, supporting such a decrease in [96, 103].
+(ii) Random coefficient models may seem more natural for polynomial root-finding but are useless because no known probability distribution withstands shift of the variable $x$.
 Theorem 1.4.
 (Main Theorem 4) The upper bounds of Thms. 1.1-1.3 on computational precision and bit operation complexity can be decreased by a factor of $d$ under the Random Root Model.
-In view of Remarks 1.1 and C.4 we can readily extend our Main Theorems to Polynomial Root Isolation.
-1.6 Our algorithms: design and analysis (outline)
-For a polynomial $p$ given in monomial basis (with its coefficients) our deterministic algorithms of [92, 99] support the Main Theorems 3 and 4, but our present Las Vegas randomized algorithms does this in the case of a black box polynomial as well.
-1.6.1 Subdivision root-finders and reduction to e/i tests
-We depart from the popular subdivision root-finders, traced back to [144, 50, 124, 97], which extend root-finding in a line segment by means of bisection to polynomial root-finding in a square on the complex plane and have been extensively used in Computational Geometry under the name of Quad-tree Construction.
-Let a square contain precisely $m>0$ zeros of $p$ (with multiplicities).
-Call it suspect, recursively divide every suspect square into four congruent sub-squares, and to each of them apply exclusion/inclusion (e/i) test, to decide whether the square contains any root or does not.
+In view of Remarks 1.1 and 25.4 we can readily extend our Main Theorems to Polynomial Root Isolation.
+2 Accelerated subdivision polynomial root-finders: overview and reduction to soft exclusion/inclusion tests
+For a polynomial $p$ given in monomial basis (with its coefficients) the deterministic algorithm of [96, 103] supports our Main Theorems 3 and 4, but not our Main Theorems 1 and 2.
+This is because the algorithm of [96, 103] cannot be applied to a black box polynomial, unlike our present Las Vegas randomized algorithms, which we outline next and elaborate upon in Parts II–IV.
+Namely, in the next two sections we reduce our root-finders to recursive application of exclusion/inclusion (e/i) test, which decides whether a fixed complex disc contains roots or not.
+In Secs.
+3.3.2 and 3.4 we outline our two approaches to devising e/i tests for a black box polynomial and unit disc $\mathcal{D}=D(0,1)$.
+We cover these two approaches in some detail in Part II, made up of Secs.
+612, and Part III, made up of Secs.
+1318, respectively, and describe yet another fast e/i test in Sec.
+23.4.
+Our root-finders of Parts II and III incorporate these e/i tests and are quite fast but still only support root approximation (rather than isolation) and the complexity estimates exceeding those of our Main Theorems by a factor of $b$.
+In Part IV, made up of Secs.
+1923 and outlined in Sec.
+4, we fix both deficiencies, thus proving our Main Theorems.
+In Sec.
+5.4 we list some tentative directions to further progress.
+2.1 Subdivision root-finders and reduction to e/i tests
+We depart from the popular subdivision root-finders, traced back to [149, 52, 128, 101].
+They extend root-finding in a line segment by means of bisection to polynomial root-finding in a square on the complex plane and have been extensively used in Computational Geometry under the name of Quad-tree Construction.
+Next we outline them.
+(A) First let a square on the complex plane contain precisely $m>0$ roots (with multiplicities) and let us approximate them.
+Call the square suspect, recursively divide every suspect square into four congruent sub-squares, and to each of them apply exclusion/inclusion (e/i) test in order to decide whether the square contains a root or does not.
 If it does, call the square suspect and process it in the next iteration; otherwise discard it (see Fig.
 1a).
+In Sec.
+9 and then in our Part IV we use a natural extension of e/i tests to $m$-tests, which decide whether a disc in the complex plane covering at most $m$ roots does or does not contain $m$ roots.
 (a) Four roots marked by asterisks lie in suspect sub-squares; the other (empty) sub-squares are discarded.
@@
 Figure 1:
-Observation 1.1.
-A subdivision iteration applied to a square with $m$ zeros of $p$ (i) decreases by twice the diameter of a suspect square and (ii) processes at most $4m$ suspect squares, (iii) whose centers approximate every zero of $p$ within the distance of one half of their diameter.
+Observation 2.1.
+A subdivision iteration applied to a square with $m$ roots (i) decreases by twice the diameter of a suspect square and (ii) processes at most $4m$ suspect squares, (iii) whose centers approximate every root within the distance of one half of their diameter.
 Proof.
 Readily verify claim (i).
-Claim (ii) holds because a zero of $p$ can lie in at most four suspect squares (and even in at most two unless it is the shared vertex of the four).
+Claim (ii) holds because a root can lie in at most four suspect squares, and even in at most two unless it is the shared vertex of the four.
 Claim (iii) holds because a subdivision iteration never discards a square containing a root.
@@
 Corollary 2.1.
-Let subdivision iterations be applied to a square $S$ with side length $\Delta$ (and diameter $\Delta\sqrt{2}$) containing precisely $m$ zeros of $p$.
-Then (i) the $k$-th iteration outputs at most $4m$ suspect squares whose centers approximate within $\epsilon:=\Delta/2^{k-0.5}$ all the $m$ zeros of $p$ lying in $S$, so that $k\leq\lceil\log_{2}(\frac{\Delta}{2\epsilon}\sqrt{2})\rceil$, and (ii) the first $k$ iterations together involve at most $4mk$ e/i tests.
-1.6.2 Soft e/i tests and compression of a Region of Interest
-Instead of testing whether a suspect square $S$ contains some zeros of $p$, Henrici in [50] tested whether the minimal disc $D(c,\rho)$ covering $S$ does this.
-We adopt his approach, which implies that the zeros of $p$ lying outside the square but in the disc can affect the result of our e/i test.
-Furthermore, in the presence of rounding errors no e/i test can decide whether a zero of $p$ lies in or outside a disc if it actually lies on or very close to its boundary circle.
-To address both problems we fix $\sigma>1$ and apply soft e/i tests as follows.
+Let subdivision iterations be applied to a square $S$ with side length $\Delta$ and diameter $\Delta\sqrt{2}$ containing precisely $m$ roots.
+Then (i) the $k$-th iteration outputs at most $4m$ suspect squares whose centers approximate within $\epsilon:=\Delta/2^{k-0.5}$ all the $m$ roots lying in $S$, so that $k\leq\lceil\log_{2}(\frac{\Delta}{2\epsilon}\sqrt{2})\rceil$, and (ii) the first $k$ iterations together involve at most $4mk$ e/i tests.
+(B) Seeking roots in any fixed domain $\mathbb{D}$ in the complex plane, we apply subdivision iterations to its minimal covering square $S$ and while performing them also discard all squares that do not intersect the domain $\mathbb{D}$, even if they contain roots.
+Then the $(\Delta_{k}\sqrt{2})$-neighborhood of $\mathbb{D}$ covers the union $\mathbb{U}_{k}$ of all suspect squares defined after the $k$th subdivision iteration, while $\Delta_{k}=\Delta_{0}/2^{k}$ fast converges to 0 as $k$ increases.
+Hence the neighborhood fast converges to the domain $\mathbb{D}$: the width of the neighborhood decreases by a factor of $2^{k}$ in $k$ subdivision iterations, still at the price of performing at most $4km$ e/i tests.
+Remark 2.1.
+[Real and near-real root-finding.] In the case of Problem 0 of Example 1.2, generalize bisection as follows: (i) recursively partition (subdivide) every suspect rectangle into two congruent rectangles, to each apply e/i test, and then discards a rectangle if it contains no roots but call it suspect and subdivides it otherwise, (ii) stop where the rectangles have diameter at most $2\epsilon$, so that their centers approximate the roots within $\epsilon$.
+2.2 Subdivision iterations with soft e/i tests
+Instead of testing whether a suspect square $S$ contains some roots, Henrici in [52] tested whether the minimal disc $D(c,\rho)$ covering $S$ does this.
+Furthermore, in the presence of rounding errors no e/i test can decide whether a root lies in or outside a disc if it actually lies on or very close to its boundary circle.
+To address both problems we fix $\sigma>1$ and apply soft e/i tests.
 Definition 2.1.
-Given a real $\sigma>1$, a polynomial $p$, a square $S$, and its minimal covering disc $D(c,\rho)$, the $\sigma$-soft e/i test reports either exclusion or $\sigma$-soft inclusion.
-Exclusion certifies that the square $S$ contains no zeros of $p$ and therefore should be discarded from subdivision process.
+[See the end of Sec.
+6 for a generalization, covering in particular soft $m$-tests.] Given a real $\sigma>1$, a polynomial $p$, a square $S$, and its minimal covering disc $D(c,\rho)$, the $\sigma$-soft e/i test reports either exclusion or $\sigma$-soft inclusion.
+Exclusion certifies that the square $S$ contains no roots and therefore should be discarded from subdivision process.
 $\sigma$-soft inclusion certifies that the disc $D(c,\sigma\rho)$, said to be the $\sigma$-cover of $S$, contains a root; then the square $S$ is said to be suspect and is subdivided into four congruent sub-squares in the next subdivision step.
@@
 1b), but the test ends as soon as it verifies any of them.
-Let us estimate how transition to $\sigma$-soft e/i tests affects Cor.
-1.1.131313We counter the adverse impact of rounding errors by increasing $\sigma$ but only need tiny increase for small $\epsilon$ and will ignore it in our estimates.
-Observation 1.2.
-(i) The number $n(\sigma)$ of the $\sigma$-covers of suspect squares that contains a fixed zero of $p$ is $O(1)$, that is, bounded by a constant, if so is $\sigma$.
-(ii) In particular $n(\sigma)\leq 4$ if $\sigma<\sqrt{2}$.
-Remark 1.3.
-We only use $\sigma$-soft e/i tests for $\sigma<\sqrt{2}$ in our root-finders.
-Seeking roots in the unit disc $D(0,1)$ we apply subdivision iterations to its minimal covering square $S(0,1)=\{x:~{}|\Re(x)|\leq 1,~{}|\Im(x)|\leq 1\}$, having side length 2 (see Fig.
-1b) and $\sigma$-cover $D(0,\sigma\sqrt{2})$.
-Observation 1.3.
-Let $U_{i}=U_{i}(\sigma)$ denote the union of the $\sigma$-covers of all suspect squares processed at the $i$-th subdivision step with $\sigma$-soft e/i tests.
-Then (i) $U_{i+1}(\sigma)\subset U_{i}(\sigma)$, that is, domain $U_{i}(\sigma)$ shrinks as $i$ increases, (ii) $U_{i}(\sigma)\subset D(0,2)$ for all $i\geq 0$ if $\sigma<\sqrt{2}$, and (iii) $U_{h}(\sigma)\subset D(1+\beta)$ for any fixed positive $\beta$ and some $h=h(\sigma,\beta)$, which is a constant, $O(1)$, if so is $\sigma$.
-Under the Random Root Model the numbers of roots lying in the discs $D(0,1)$ and $D(0,2)$ differ by a constant factor $f=O(1)$ whp, and so for $\sigma<\sqrt{2}$ the estimates of Cor.
-1.1 hold up to that factor whp, due to claim (ii) of the latter observation.
-Based on its claims (i) and (iii) we next extend Cor. 1.1 deterministically, by incorporating the following two recipes of Extension and Delay for discarding squares in subdivision process.
-(a) Extend discarding: In addition to discarding squares according to $\sigma$-soft e/i tests, discard all squares that lie outside the disc $D(0,1)$.
-(b) Delay discarding: For a fixed positive $\beta$ begin applying $\sigma$-soft e/i tests and discarding squares only at the $h$th subdivision step for $h=h(\sigma,\beta)$ of claim (iii) of Observation 1.3.
-Clearly, the $h$th subdivision iteration may process up to $4^{h}=O(1)$ suspect squares, although a little fewer for larger $h$ if the extension of discarding is applied.
-Now we can extend Cor. 1.1 as follows.
+Observation 2.2.
+Let $\gamma(\sigma)$ denote the maximal number of $\sigma$-covers at a fixed subdivision step that share a complex point, e.g., a root.
+Then $\gamma(\sigma)$ is in $O(1)$ if so is $\sigma$, and $\gamma(\sigma)\leq 4$ if $\sigma<\sqrt{2}$.
 Corollary 2.2.
-Fix $\beta>0$ and $\sigma>1$, let the disc $D(0,1+\beta)$ contain precisely $m=m(\beta)$ zeros of $p$ (cf.
-Def.
-1.1), and apply subdivision iterations (with $\sigma$-soft e/i tests under the recipes of extension and delay of discarding) to a minimal square $S(0,1)$ covering the unit disc $D(0,1)$.
-Then (i) the $k$-th iteration involves at most $n(\sigma)m$ suspect squares for $n(\sigma)$ of Observation 1.2, $n(\sigma)\leq 4$ for $\sigma<\sqrt{2}$; the centers of those squares approximate within $2^{1.5-k}$ all the zeros of $p$ lying in the disc $D(0,1)$ and (ii) the first $k$ subdivision iterations involve no suspect squares and $\sigma$-soft e/i tests for $k<h=h(\sigma,\beta)$ and at most $(k-h-1)n(\sigma)m+4^{h}$ suspect squares and $\sigma$-soft e/i tests for $k\geq h$.
-Remark 1.4.
-We extend Cor. 1.1 for subdivision process applied to the unit disc $D(0,1)$ but can similarly do this for any fixed Region of Interest.
-1.6.3 Cauchy-based root-counting and e/i tests
-We have highly efficient heuristic algorithms based on the DLG iterations (1.5) that closely approximate the smallest distance to the roots from a complex point (see Remark 1.2); this can be immediately extended to soft e/i tests.
-Next we cover efficient e/i tests that also enable formal support of our Main Theorems 1– 4.
-Let #$\#(\mathcal{D})$ denote the number of zeros of $p$ lying in a complex domain $\mathcal{D}$ with the boundary $\Gamma$ and recall from [1] that
+Cor.
+2.1 still holds if e/i tests are replaced with $\sigma$-soft e/i tests for $\sigma<\sqrt{2}$, while a subdivision root-finder using $\sigma$-soft e/i tests involves $O(km)$ such tests if $\sigma$ is a constant.
+3 Fast soft e/i tests
+3.1 Background: Newton’s Inverse Ratio and its approximation
+We reduce e/i tests of subdivision root-finders to the multipoint evaluation of the ratio
+${\rm NIR}(x)=\frac{p^{\prime}(x)}{p(x)}=\sum_{j=1}^{d}\frac{1}{x-z_{j}},$ (3.1)
+which we call Newton’s inverse ratio. To prove the latter well-known and much used equation, differentiate $p(x)=p_{d}\prod_{j=1}^{d}(x-z_{j})$.
+The arithmetic operations (hereafter ops) involved in our evaluation of NIR$(x)$ are more numerous than the other ops involved in our root-finders, even in the case where $p(x)=x^{d}$.
+By following [76] we fix a complex $\delta$, with $|\delta|$ small in context, and approximate $p^{\prime}(x)$ with the first order divided difference $\frac{p(x)-p(x-\delta)}{\delta}$, equal to $p^{\prime}(y)$ for some $y\in[x-\delta,x]$, by virtue of Taylor-Lagrange’s formula.
+Hence $y\mapsto x$ and $p^{\prime}(y)\mapsto p^{\prime}(x)$ as $\delta\mapsto 0$, and
+${\rm NIR}(x)\approx\frac{p(x)-p(x-\delta)}{p(x)\delta}=\frac{p^{\prime}(y)}{p(x)}~{}{\rm for~{}some}~{}y\in[x-\delta,x].$ (3.2)
+We deduce the same bound $O(b\log(d))$ of [76] on computational precision of our root-finders, but its derivation is greatly simplified versus [76] because we only use the first order derivatives and the first order divided differences.
+Unlike [76] we estimate bit-operation cost of our root-finders and allow random choice of $x$ to control the value of the ratio $\frac{|p(x)-p(x-\delta)|}{p(x)\delta}$ with a high probability (whp).
+This is Las Vegas randomization, and so is any randomization in polynomial root-finding, as we show next.
+3.2 Background: Error detection and correction
+A randomized root-finder can lose some roots, although with a low probability, but we can detect such a loss at the end of root-finding process, simply by observing that among the $m\leq d$ roots lying in an input disc only $m-w$ tame roots, say, $z_{w+1},z_{w+1},\dots,z_{m}$ have been closely approximated, [^6] We call a complex $c$ a tame root for an error tolerance $\epsilon$ if $c$ lies in an isolated disc $D(c,\epsilon)$, whose index #$\#(D(c,\epsilon))$ we can readily compute (cf.
+Sec.
+23.4).
+while $w>0$ wild roots, say, $z_{1},z_{2},\dots,z_{w}$ remain at large.
+We can compute the coefficients of the wild factor $q(x):=p(x)/\prod_{j=w+1}^{m}(x-z_{j})$, say, by applying evaluation-interpolation techniques, then apply to $q(x)$ the same or another root-finder, and repeat this computation recursively until we approximate all $m$ roots in the disc.
+[^7] By this recipe we detect output errors at the very end of computations in any root-finder that can only lose roots but cannot wrongly generate them.
+Our subdivision root-finders belong to this class, but with them we can detect the loss of a root even earlier – whenever in a subdivision process applied to a square with $m$ roots we detect that $m$ exceeds the sum of the indices of all suspect squares, that is, the overall number of roots contained in them.
+Readily bounds the complexity and error probability of this recursive process:
+Proposition 3.1.
+Let the above recipe for error detection and correction be applied to the output of a randomized root-finder performing at the cost $\alpha$ (under a fixed model), let it be correct with a probability $1/\beta$ for $\beta>1$ and let the algorithm be applied recursively, each time with error detection at the cost $\delta$.
+Then in $v$ recursive applications, the output is certified to be correct at the cost of less than $(\alpha+\delta)v$ with a probability at least $1-1/\beta^{v}$.
+Explicit deflation can blow up the coefficient length and destroy sparsity unless the degree $w$ is sufficiently small, but we avoid these mishaps by combining a fixed black box polynomial root-finder with implicit deflation, that is, by applying that root-finder to the quotient polynomial $q(x)$ without computing its coefficients – based on the following expression, implied by (3.1):
+$\frac{q^{\prime}(x)}{q(x)}=\frac{p^{\prime}(x)}{p(x)}-\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}.$ (3.3)
+One can accelerate low precision numerical approximation of the sum $\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}$ simultaneously at many points $x$ by applying Fast Multipole Method (FMM) of [50, 28, 32].
+3.3 Cauchy-based root-counting and e/i tests
+3.3.1 Introductory comments
+Let #$\#(\mathcal{D})$ denote the number of roots lying in a complex domain $\mathcal{D}$ with the boundary $\Gamma$ and recall from [1] that
 […]
@@
 This is more than enough for our e/i tests, for which we only need to know whether $s_{0}=0$ or not.
-In Secs.
-1.6.4 and 1.6.5 we outline our two approaches to devising e/i tests for a black box polynomial based on (1.9) for $\mathcal{D}=D(0,1)$ We cover these two approaches in some detail in Part II, made up of Secs.
-29, and Part III, made up of Secs.
-10-15, respectively.
-The subdivision root-finders that incorporate these e/i tests are quite fast but only support root approximation (rather than isolation) and the complexity estimates exceeding those of our Main Theorems by a factor of $b$.
-In Part IV, made up of Secs.
-16-20 and outlined in Sec.
-1.6.7, we fix both deficiencies, thus proving our Main Theorems.
-1.6.4 Our soft e/i tests in Part II: outline
+3.3.2 Cauchy-based soft e/i tests: outline
 In Part II we approximate the integral of (3.4) with finite sums made up of the scaled values of NIR$(x)$ at the scaled $q$th roots of unity.
@@
 and devise randomized $\sigma$-soft e/i tests as follows:
-(i) choose a real $\rho$ at random in a fixed range covered by $[1,\sigma]$,
-(ii) fix a sufficiently large integer $q$ of order $m\log(d)$, and
-(iii) compute approximation $s_{0,q}$ to the integral (1.9) for the disc $D(0,\rho)$ and the polynomial $t(x)=p(\rho x)$ replacing $p(x)$.
+(i) Choose a real $\rho$ at random in a fixed range covered by $[1,\sigma]$.
+(ii) Fix a sufficiently large integer $q$ of order $m\log(d)$.
 […]
-(iv) Obtain #$\#(D(0,\rho))$ as the closest integer approximation of $s_{0,q}$, prove that the recipe is correct whp, and extend the root-counter to e/i test.
-(v) Repeat the test $v$ times to decrease the probability of its output error exponentially in $v$.
-The $v$ tests are reduced to the evaluation of $p(x)$ at $O(mv\log(d))$ points; in Part II we also devise a deterministic soft e/i test at the price of increasing the above bound on the number of points to order of $m^{2}\log(d)$.
-1.6.5 Our generic soft e/i test in Part III
-(i) In Part III we decrease our estimates of Part II for $q$ in e/i tests by a factor of $m$ by avoiding approximation of the integer #$\#(D(0,1))$ and only deciding whether this integer vanishes or not, which is sufficient for e/i tests.
-In these e/i tests we elaborate upon the following generic one, where we use $h$-lifting of roots generalizing DLG root-squaring (1.5):
+(iv) Conclude that exclusion holds whp if $|s_{0,q}|<1/2$.
+(v) Repeat the test $v$ times to decrease the probability of its output error exponentially in $v$, from $\frac{1}{\gamma}$ to $\frac{1}{\gamma^{v}}$ for $\gamma>1$ of our choice, say, for $\gamma=2$.
+The $v$ tests are reduced to the evaluation of $p(x)$ at $O(mv\log(d))$ points.
+3.4 Soft e/i test with root-lifting: generic algorithm
+In Part III we decrease the complexity estimates of Part II for $q$ in e/i tests by a factor of $m$ by avoiding approximation of the integer #$\#(D(0,1))$ and only deciding whether this integer is 0 or not, which is sufficient for e/i tests.
+In these e/i tests we elaborate upon the generic one (see Alg.
+3.1 below), where we use $h$-lifting of roots:
 $p_{0}(x):=\frac{p(x)}{p_{d}},~{}p_{h}(x):=\prod_{g=0}^{h-1}p_{0}(\zeta_{h}^{g}x)=(-1)^{d}\prod_{j=1}^{d}(x-z_{j}^{h}),~{}\zeta_{h}:=\exp\Big{(}\frac{2\pi{\bf i}}{h}\Big{)},$ (3.6)
 for $h=1,2,\dots$ and $t_{2^{k}}(x)=p_{k}(x)$ for all $k$.
+$h$-lifting maps roots to their $h$th powers; this keeps the unit disc $D(0,1)$ and its boundary circle $C(0,1)$ intact but moves the external roots away from the circle, thus strengthening its isolation from those roots.
+Remark 3.1.
+The $h$-lifting generalizes the DLG classical root-squaring iterations of Dandelin 1826, Lobachevsky 1834, and Gräffe 1837 (cf. [51]):
+$p_{0}(x):=\frac{p(x)}{p_{d}},~{}p_{2^{i}}(x^{2}):=(-1)^{d}p_{2^{i-1}}(x)p_{2^{i-1}}(-x)=\prod_{j=1}^{d}(x-z_{j}^{2^{i}}),~{}i=1,2,\dots.$ (3.7)
+Polynomial root-finding in the 19-th and well into the 20-th century relied on iterations (3.7) combined with Vieta’s formulas, which link the polynomial coefficients $p_{0},p_{1},\dots,p_{d}$ to the elementary symmetric polynomials of the roots.
+The combination enabled efficient approximation of the minimal and maximal distance of the roots from the origin and due to the map (1.3) from any complex point (see [86, 113, 115] for details).
+The approach became obsolete because of severe problems of numerical stability of the DLG iterations, but they disappeared in the case of root-finders that work for black box polynomials.
+Extensive tests with standard test polynomials have consistently demonstrated efficiency of the particular implementation of this approach in [115], but even the straightforward implementation of iterations (3.7) promises to be highly efficient as well.
+We will elaborate upon the following generic algorithm in Sec. 14.
 Algorithm 3.1.
 A generic soft e/i test.
-INPUT: a black box polynomial $p(x)$ of a degree $d$, given by an oracle for its evaluation, a real $\sigma>1$ and a positive integer $h$ (both specified in Sec. 11).
-INITIALIZATION: Fix a positive $q\in\mathbb{N}$.
+INPUT: $\sigma>1$, an integer $d>0$, and a $d$-th degree black box polynomial $p(x)$ given by an oracle for its evaluation.
+INITIALIZATION: Fix a positive $q\in\mathbb{N}$ and a positive integer $h$.
 COMPUTATIONS:
 (1) Select $q$ points $x_{1},\dots,x_{q}$ on the circle $C(0,1)$.
-(2) By extending (1.6) to $p_{h}(x)$ of (1.5) (we specify this extension in (10.5)) closely approximate
-$v_{i}:=|{\rm NIR}_{p_{h}}(x_{i})|=\Big{|}\frac{p^{\prime}_{h}(x_{i})}{p_{h}(x_{i})}\Big{|},~{}i=1,\dots,q.$ (1.12)
-(3) Certify that #$(D(0,\sigma))>0$, that is, claim $\sigma$-soft inclusion, if $1+d/v_{j}\leq\sigma^{h}$ for some $i\leq q$. Otherwise, claim that #$(D(0,1))=0$, that is, claim exclusion.
-(ii) To prove correctness of soft inclusion claim of Alg. 1.1, first differentiate $p(x)=p_{d}\prod_{j=1}^{d}(x-z_{j})$ and obtain the following well-known equation:
-${\rm NIR}(x):=\frac{p^{\prime}(x)}{p(x)}=\sum_{j=1}^{d}\frac{1}{x-z_{j}}.$ (1.13)
+(2) By extending (3.2) to $p_{h}(x)$ of (3.7) (we specify this extension in (13.3)) approximate the values
+$v_{g}:=|{\rm NIR}_{p_{h}}(x_{g})|=\Big{|}\frac{p^{\prime}_{h}(x_{g})}{p_{h}(x_{g})}\Big{|},~{}g=1,\dots,q.$ (3.8)
+(3) If $1+d/v_{g}\leq\sigma^{h}$ for some $g\leq q$, then certify that #$(D(0,\sigma))>0$, that is, claim $\sigma$-soft inclusion. Otherwise, claim that #$(D(0,1))=0$, that is, claim exclusion.
+3.5 Soft e/i test with root-lifting: correctness of inclusion claim
 Eqn. (3.1) implies two useful corollaries:
@@
 Corollary 3.2.
-#$\#(D(c,\rho))>0$ if $\rho\leq d~{}|p(c)/p^{\prime}(c)|$.
+It holds that $r_{d}(c)\leq d~{}\Big{|}\frac{p(c)}{p^{\prime}(c)}\Big{|}$, that is, #$\#\Big{(}D(c,d~{}\Big{|}\frac{p(c)}{p^{\prime}(c)}\Big{|}\Big{)}\Big{)}>0$, for any complex $c$.
 Proof.
@@
 $\Big{|}\frac{p^{\prime}(c)}{p(c)}\Big{|}=\Big{|}\sum_{j=1}^{d}\frac{1}{c-z_{j}}\Big{|}\leq\sum_{j=1}^{d}\frac{1}{|c-z_{j}|}<\frac{d}{\rho},~{}{\rm and~{}so}~{}\Big{|}\frac{p(c)}{p^{\prime}(c)}\Big{|}>\frac{\rho}{d}.$
-We use Cor. 1.3 later but next apply Cor. 1.4 to prove the following lemma, which implies correctness of the soft inclusion claim of Alg. 1.1.
+We use Cor.
+3.1 later but next apply Cor.
+3.2 to prove correctness of the inclusion claim of Alg.
+3.1.
+The bound of Cor.
+3.1 on $r_{d}(c)$ is too large for that, but a similar bound on $r_{d}(c,p_{h})=(r_{d}(c))^{h}$, implies that $r_{d}(c)$ is close to 1 already for $h$ of order $\log(d)$.
 Lemma 3.1.
@@
 Alg.
-1.1).
+3.1).
 Deduce that #$(D(0,\sigma))>0$ for $\sigma^{h}=1+d/v_{i}$, that is, Alg.
 3.1 is correct when it claims $\sigma$-soft inclusion.
-Observation 1.4.
-$\sigma$-soft inclusion test by means of Alg. 1.1 is correct where
+Observation 3.1.
+$\sigma$-soft inclusion test by means of Alg. 3.1 is correct where
 $d\geq 2,~{}x_{1}=1,~{}\frac{1}{|{\rm NIR}(1)|}=4\sqrt{d}-\frac{1}{d},~{}h=40\log_{2}(d),~{}{\rm and}~{}\sigma<1.1.$
@@
 $1+\rho=1+\frac{d}{|{\rm NIR}(1)|}=4d\sqrt{d}=2^{2+1.5\log_{2}(d)}.$
-Apply Lemma 1.1 and obtain that #$(D(0,\sigma))>0$ for $\sigma=(1+\rho)^{1/h}=2^{\nu}~{}{\rm for}~{}\nu=\frac{2+1.5\log_{2}(d)}{40\log_{2}(d)}$.
+Apply Lemma 3.1 and obtain that #$(D(0,\sigma))>0$ for $\sigma=(1+\rho)^{1/h}=2^{\nu}~{}{\rm and}~{}\nu=\frac{2+1.5\log_{2}(d)}{40\log_{2}(d)}$.
 Deduce that $\nu\leq\frac{7}{80}$ because $\frac{2}{40\log_{2}(d)}\leq 0.05$ for $d\geq 2$ and because $\frac{2+1.5\log_{2}(d)}{40\log_{2}(d)}<\frac{3}{80}$.
 Hence $\sigma\leq 2^{\frac{7}{80}}=1.062527\dots<1.1$. ∎
-(b) We prove correctness of the exclusion claim of Alg.
-1.1 separately for its two specifications, both of which strengthen isolation of the unit circle $C(0,1)$ by means of root-squaring (1.5) or root-lifting (1.11).
-In Sec.
+3.6 Soft e/i test with root-lifting: further analysis (outline)
+(i) Correctness of the exclusion claim of Alg.
+3.1.
+We prove it separately for its two specifications, both of which strengthen isolation of the unit circle $C(0,1)$ by means of root-squaring (3.7) or root-lifting (3.6).
+(a) In Sec.
 14.3 we choose any integer $q$ exceeding $d$ and then prove deterministic correctness of exclusion claim, provided that $|{{\rm NIR}_{p_{h}}(x_{i})}|<1/(2\sqrt{q})$ for all $x_{i}$ being $q$th roots of unity.
@@
 We also extend this deterministic e/i test by allowing any $q>m$ (rather than $q>d$) provided that $|{{\rm NIR}_{p_{h}}(x_{i})}|<1/(3\sqrt{q})$ for all $x_{i}$ being $q$th roots of unity.
-In Sec.
+(b) In Sec.
 14.4 we sample $q=O(\log(d))$ iid random values of $x_{i}$ under the uniform probability distribution on the circle $C(0,1)$ and then prove that the exclusion claim is correct whp under the Random Root Model if $|{{\rm NIR}_{p_{h}}(x_{i})}|\leq 1/d^{a}$ for a sufficiently large constant $a$.
-(iv) Counting query locations.
-Approximation of $p_{h}(x)$ is reduced to approximation of $p(x)$ at all the $h$th roots of $x$ for $h$ being a power of 2 based on (1.5) or for any integer $h>1$ if we apply (1.11), and so Alg.
+(ii) Counting query locations.
+Eqn.
+(3.7) reduces approximation of $p_{h}(x)$ at $x=c$ to approximation of $p(x)$ at all the $h$th roots of $c$.
+We can fix $h>1$ being a power of 2 and apply (3.7) or fix any integer $h>1$ and apply (3.6).
+In both cases Alg.
 3.1 evaluates polynomial $p(x)$ either at $2jh$ points for $j\leq q$, when it claims inclusion, or at $2qh$ points, when it claims exclusion.
-In all these estimates we choose factor $h$ in $O(\log(d))$, and so we can ignore this factor in our $\tilde{O}$ estimates and say that our soft e/i test by means of Alg.
-1.1 evaluates polynomial $p(x)$ at $\tilde{O}(m)$ points.
-Combine these estimates with Cor.
-1.2 to bound by $\tilde{O}(bm^{2})$ the overall number of points $x$ at which we evaluate $p(x)$ in subdivision root-finders incorporating our e/i tests.
+We always choose $h=O(\log(d))$ and then can ignore a factor of $h$ in our $\tilde{O}$ estimates.
+In our deterministic soft e/i test based on Alg.
+3.1 we evaluate polynomial $p(x)$ at $\tilde{O}(m)$ points, combine these estimates with Cor.
+2.2, and bound by $\tilde{O}(bm^{2})$ the overall number of points $x$ at which we evaluate $p(x)$ in subdivision root-finders incorporating our e/i tests.
 Under the Random Root Model of Sec.
-1.5 the zeros of $p$ are isolated whp from the unit circle $C(0,1)$, and we decrease the bound $\tilde{O}(bm^{2})$ to $\tilde{O}(bm)$ whp by exploit their even stronger isolation ensured by means of $O(\log(\log(d)))$ root-squaring steps (1.5) or alternatively with $O(\log(d))$ steps of root-lifting of (1.11).
-1.6.6 Computational precision and bit-operation complexity
+1.5 the roots are reasonably well isolated whp from the unit circle $C(0,1)$.
+We strengthen this isolation by applying $O(\log(\log(d)))$ root-squaring steps (3.7) or $O(\log(d))$ steps of root-lifting of (3.6) and then whp decrease the overall bit operation count from $\tilde{O}(bm^{2})$ to $\tilde{O}(bm)$.
+(iii) Estimation of computational precision and bit-operation complexity.
 Louis and Vempala in [76] bound by $O(b\log(d))$ the computational precision of their novel high order Newton’s iterations.
-We also prove this bound for our root-finders by using a small part of their analysis together with Las Vegas randomization and some rather customary techniques of numerical analysis.
-We extend this bound combined with the bound on the number of queries in our root-finders to the bit operation complexity estimates: we use the known fast algorithms for multipoint polynomial evaluation.
-We satisfy the assumptions of the latter algorithms by using map (1.4).
-We devise two groups of supporting algorithms in Parts II and III of the paper, and in both cases arrive at the estimates for the numbers of evaluation points and bit-operations involved in our algorithms that exceed those of our Main Theorems by factors of $b$ for the roots and $\tilde{b}$ for the eigenvalues.
-1.6.7 Acceleration of root-finding in Part IV: outline
-In Part IV we accelerate our algorithms to yield the Main Theorems. Here is an outline of that acceleration.
+We combine a small part of their advanced analysis with Las Vegas randomization and some customary techniques of error analysis to extend this estimate to our root-finders.
+For the solution of Problems 0, 1, and 3, where the coefficients are given, we apply fast algorithm of [81] for multipoint polynomial evaluation to decrease the estimated number of queries and the bit operation cost bound.
+We devise two groups of efficient e/i tests in Parts II and III of this paper, and in both cases arrive at the estimates for the numbers of (i) evaluation points, that is, of the oracle queries, and (ii) bit-operations involved in our algorithms.
+Both of the estimates exceed those of our Main Theorems by factors of $b$ for the roots and $\tilde{b}$ for the eigenvalues.
+4 Acceleration of subdivision root-finding by means of compression: outline
+In Part IV we accelerate the algorithms of both Parts II and III by these factors (of $b$ for roots and $\tilde{b}$ for eigenvalues) to prove the Main Theorems in two ways.
+Next we outline that acceleration.
 (i) Component tree and compact components.
 Apply subdivision process for a fixed $\sigma$, $1<\sigma<\sqrt{2}$, to Problem 1 and partition the union of all suspect squares tested at every subdivision step into connected components.
-Represent all partitions of components with a component tree as follows.
+Represent all partitions of components with a component tree, $\mathbb{T}$, as follows.
 At most $m$ components, said to be the leaves, are covered by isolated $\epsilon$-discs for a fixed error tolerance $\epsilon\geq 0$.
@@
 Recall that every component contains $O(m)$ suspect squares whose diameter is halved at every subdivision step.
-Hence in $O(\log_{2}(mb))$ steps, for $b:=\log_{2}(1/\epsilon)$, every component has diameter below $\epsilon/2$, and then its $\sigma$-cover lies in an $\epsilon$-disc.
+Hence in $O(\log(b))$ subdivision iterations, for $b:=\log_{2}(1/\epsilon)$, every component has diameter below $\epsilon/2$, and then its $\sigma$-cover lies in an $\epsilon$-disc.
 Call a component compact if it consists of at most four squares sharing a vertex (see Fig.
@@
 (ii) Acceleration of subdivision process with compression of components.
-The bound $O(m)$ does not generally hold for the overall number of suspect squares at all subdivision steps.
-This is because subdivision steps become inefficient where a component stays compact and not broken in many consecutive subdivision steps (see Fig.
-3), but we replace these steps with compression of such a component $C$ into a near minimal disc $D$ that shares with $C$ all the zeros of $p$ lying in it.
+Subdivision process becomes quite inefficient where a component stays compact and not broken into more than one component in many consecutive subdivision steps (see Fig.
+3).
+To accelerate it, we replace these steps with compression of such a component into an $\epsilon$-disc or a near-minimal disc $D$ that still keeps all roots lying in the component.
+In Sec.
+19 we comment on two customary but inferior alternatives to our compression, relying on deflation or solely on functional iterations.
 Figure 3: Five blue marked centers of suspect squares converge to a green-marked root cluster with linear rate; two red marked Schröder’s iterates of (23.1) converge with quadratic rate.
-If the disc $D$ is an $\epsilon$-disc we just output it.
-Otherwise we apply subdivision iterations to its minimal superscribing square $S$.
-Since we compress the component into a near minimal disc $D$, we obtain readily that these iterations very soon output more than one component or a sub-domain of an $\epsilon$-disc.
+If the compression outputs $\epsilon$-disc, we just output it and do not subdivide it anymore.
+Otherwise we apply subdivision iterations to the minimal superscribing square $S$ of the disc $D$.
+Since the disc is near-minimal, the iterations very soon output more than one component or a sub-domain of an $\epsilon$-disc.
 From this we deduce that our subdivision process with compression reduces the solution of Problem 1 to at most $22m-$ 2 compression steps and $O(m)$ soft e/i steps overall.
-(iii) Compression of components. We reduce a compression of a compact component to the following main stages (see more details in Sec. 20.5):
-(i) compute a point $c$ lying in or near the minimal disc that covers all zeros of $p$ lying in the component $C$ and let $c$ denote the center of the compressed disc that we seek;
-(ii) compute a close upper bound $\rho$ on the distance from $c$ to the farthest zero of $p$ among its at most $m$ zeros lying in the compact component $C$.
-We prove that Schröder’s single celebrated iteration (20.1) following $O(\log(d))$ auxiliary subdivision steps performs step (i) above, that is, computes the desired center $c$.
-Beginning stage (ii) above we have a range $[\rho_{-},\rho_{+}]$ where the bound $\rho$ lies, for $10<\rho_{-}<\rho_{+}<$ 1.
-Hence we can approximate it in $\lceil\log_{2}(1/\epsilon)\rceil$ bisection steps, but we greatly accelerate convergence to $\rho$ by applying bisection to the range of the exponents, $[\log_{2}(\rho_{-}),\log_{2}(\rho_{+})]$.
-We extend our soft e/i tests to define the pivot step of such a modified bisection.
-Summarizing we reduce the solution of Problem 1 to $\tilde{O}(m)$ soft e/i tests and $\tilde{O}(m)$ Schröder’s iterations.
-In this way we accelerate the root-finders of Parts II and III by a factor of $b$.
-Together with our recipes for detection and correction of output errors of root-finders in the next subsection, this completes the proof of our Main Theorems.
+(iii) Compression of components. We reduce compression of a compact component $C$ to the following two main stages (see more details in Sec. 23.5):
+(a) compute a point $c^{\prime}$ lying in or near the minimal disc that covers all roots lying in the component $C$ and let $c^{\prime}$ denote the center of the compressed disc that we seek;
+(b) compute a close upper bound $\rho^{\prime}$ on the maximal distance from $c^{\prime}$ to the roots in the component $C$ and output a compressed disc $D(c^{\prime},\rho^{\prime})$.
+We prove that Schröder’s single iteration (23.1) following $O(\log(d))$ auxiliary subdivision steps performs step (a) above, that is, computes the desired center $c^{\prime}$.
+At the beginning of step (b) above we have a range $[\rho_{-},\rho_{+}]$ for the bound $\rho^{\prime}$, $0<\epsilon\leq\rho_{-}<\rho_{+}\leq 1$.
+Hence we can approximate it in $\lceil b\rceil$ bisection steps, for $b=\log_{2}(1/\epsilon)$ but we greatly accelerate convergence to $\rho$ by applying bisection to the range of the logarithms, $[\log_{2}(\rho_{-}),\log_{2}(\rho_{+})]=[b,\log_{2}(\rho_{+})]$, for $\log_{2}(\rho_{+})\leq 0$.
+We extend our fast soft e/i tests of Part II to fast soft $m$-tests (see part (A) of Sec. 2.1) and use them to define the pivot step of such a modified bisection.
+Summarizing we reduce the solution of Problems 0 and 1 to $\tilde{O}(m)$ soft e/i tests and $m$ tests and $\tilde{O}(m)$ Schröder’s iterations.
+In this way we accelerate the root-finders of both Parts II and III by a factor of $b/\log(b)$ to complete the proof of our Main Theorems.
 Remark 4.1.
@@
 Remark 4.2.
-For bounding precision of computing, Louis and Vempala in [73] presented a major technical innovation, which enabled us to deduce a priori estimates for the bit operation cost of our subdivision root-finders versus routine a posteriori estimates in the previous papers on subdivision root-finding (see, e.g., [118, 23]).
+By using some of the advanced techniques of Louis and Vempala in [76], we readily bound the precision of computing in our root-finders.
+This enables us to deduce a priori estimates for the bit operation cost of our subdivision root-finders versus inferior known a posteriori estimates (see, e.g., [122, 24]).
 In particular, the bit operation cost estimates of the Main Theorem of [24] contain the term
@@
 with “Discr" standing for “Discriminant", such that
-$\log(1/{\rm Discr}(p(x/2^{b})))=(d-1)db\log(1/{\rm Discr}(p(x))).$
+$\log(1/{\rm Discr}(p(x/2^{b})))=(d+1)db\log(1/{\rm Discr}(p(x))).$
 Subdivision iterations of the known root-finders, in particular, of those of [24], scale the variable $x$ order of $m$ times by factors of $2^{b}$.
 Hence the bit-complexity specified in the Main Theorem of [24] has order of $mbd^{2}$, which exceeds our bound and that of [96, 103] by a factor of $d$.
-1.7 Error detection and correction
-A randomized root-finder can lose some roots, although with a low probability, but we can detect such a loss at the end of root-finding process, simply by observing that among the $m\leq d$ roots lying in an input disc only $m-w$ tame roots, say, $z_{w+1},z_{w+1},\dots,z_{m}$ have been closely approximated, [^14] We call a complex $c$ a tame root for an error tolerance $\epsilon$ if $c$ lies in an isolated disc $D(c,\epsilon)$, whose index #$\#(D(c,\epsilon))$ we can readily compute (cf.
-Sec.
-20.4).
-while $w>0$ wild roots, say, $z_{1},z_{2},\dots,z_{w}$ remain at large.
-We can compute the coefficients of the wild factor $q(x):=p(x)/\prod_{j=w+1}^{m}(x-z_{j})$, say, by applying evaluation-interpolation techniques, then apply to $q(x)$ the same or another root-finder, and repeat this computation recursively until we approximate all $m$ roots in the disc.
-[^15] By this recipe we detect output errors at the very end of computations in any root-finder that can only lose roots but cannot wrongly generate them.
-Our subdivision root-finders belong to this class, but with them we can detect the loss of a root even earlier – whenever in a subdivision process applied to a square with $m$ roots we detect that $m$ exceeds the sum of the indices of all suspect squares, that is, the overall number of roots contained in them.
-Readily bounds the complexity and error probability of this recursive process:
-Proposition 1.2.
-Let the above recipe for error detection and correction be applied to the output of a randomized root-finder performing at the cost $\alpha$ (under a fixed model), let it be correct with a probability $1/\beta$ for $\beta>1$ and let the algorithm be applied recursively, each time with error detection at the cost $\delta$.
-Then in $v$ recursive applications, the output is certified to be correct at the cost of less than $(\alpha+\delta)v$ with a probability at least $1-1/\beta^{v}$.
-Explicit deflation can blow up the coefficient length and destroy sparsity unless the degree $w$ is sufficiently small, but we can avoid these mishaps by combining a fixed black box polynomial root-finder with implicit deflation, that is, by applying that root-finder to the quotient polynomial $q(x)$ without computing its coefficients – based on the following expression, implied by (1.13):
-$\frac{q^{\prime}(x)}{q(x)}=\frac{p^{\prime}(x)}{p(x)}-\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}.$ (1.14)
-In many cases one can accelerate simultaneous numerical approximation of the sum $\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}$ at many points $x$ by applying Fast Multipole Method (FMM) of [48, 27, 30].
-1.8 Our novelties and related works
-1.8.1 Our novelties (a short list)
+5 Our novelties, related works, implementation, and further progress
+5.1 Our novelties (a short list)
 Root-finding for polynomial given in the monomial basis, with their coefficients, has been studied for four millennia and very intensively in the last decades, but we had to introduce a number of nontrivial novel techniques (i) to handle efficiently black box polynomials and (ii) to accelerate the known subdivision algorithms to a near optimal level.
-Some of our techniques for black box polynomial root-finding, e.g., randomization for root-finding, numerical stabilization of root-squaring, acceleration of black box subdivision root-finding by means of compression of a disc without losing any zero of $p$, acceleration of bisection by means of applying it to the exponent range, approximation of root radii and near real roots, and even our simple but novel recipes for detection of output errors of polynomial root-finders can be of independent interest.
-1.8.2 Related works
-Unlike deterministic solution in [92, 99] we use Las Vegas randomization but propose algorithms that work for a black box polynomial, are simpler to implement, and in the case of a polynomial $p$ given with its coefficients support the same near optimal complexity bound of [92, 99].
+Some of our techniques such as randomization, numerical stabilization of root-squaring, compression of a disc without losing roots, bisection of logarithms, approximation of root radii and near-real roots, and even our simple error detection recipe can be of independent interest.
+5.2 Related works
 The pioneering paper [76] proposed a novel algorithm for approximation of an absolutely largest zero of a real-rooted black box polynomial, estimated the computational precision of that algorithm, and extended it to approximation of an absolutely largest eigenvalue of a symmetric matrix.
-We adopt their model of computation, their extension to approximation of matrix eigenvalues, and some of their patterns for estimating the computational precision of root-finders, but otherwise use distinct approach and techniques.
-Namely, we accelerate the classical subdivision root-finders, based on e/i tests.
-All known e/i test operate with the coefficients of $p$: they involve higher order derivatives $p^{(i)}(x)$, $i=1,2,\dots,d$ [50, 124], Newton’s identities and fast root-squaring algorithms based on convolution [97], or Pellet’s theorem [22, 23].
-Neither of these approaches works for black box polynomials.
-Our soft e/i tests involve the classical root-squaring techniques of Dandelin 1826, Lobachevsky 1834, and Gräffe 1837, but we stabilize them numerically avoiding notorious problems that severely limit all previous applications of maps (1.4) of shift and scaling and (1.5) of DLG iterations.
-Renegar in [124] accelerated subdivision process by a factor of $b/\log(b)$ by incorporating Newton’s root-finding iterations; the papers [97, 22, 23] combined his approach with alternative functional root-finding iterations.
-We apply a distinct technique of compression of a disc instead.
-Estimating the bit operation complexity of our root-finders we use [130, Eqn.
-(9.4)] and [78, Thm.
+From that paper we adopt the model of computation, extension to approximation of matrix eigenvalues, and some patterns for estimation of the computational precision of root-finders, but otherwise use distinct approach and techniques.
+Namely, instead of Newton’s iterations in [76], we accelerate the classical subdivision root-finders, based on e/i tests.
+Its closest variations in [128, 101, 23, 24] still use inferior e/i tests, for which we proposed two fast novel approaches – based on approximation of Cauchy integrals and on root-lifting; these tests run much faster than the tests in the cited papers and, unlike the latter tests, can be applied to a black box polynomial.
+The papers [128, 101, 23, 24] apply functional iterations (such as Newton’s and Schröders) to accelerate convergence of subdivision process but we apply our novel techniques of disc compression instead.
+Estimating the bit operation complexity of our root-finders we use [134, Eqn.
+(9.4)] and [81, Thm.
 2] for multipoint polynomial evaluation.
-[^16] The ingenious algorithm of Moroz supporting his [78, Thm.
+[^8] The ingenious algorithm of Moroz supporting his [81, Thm.
 2] departs from the simple observation that a polynomial $p(x-c)\mod(x-c)^{\ell-1}$ approximates $p(x)$ in the disc $D(c,\rho)$ within $O(\rho^{\ell+1})$ for small $\rho$.
 To approximate $p(c)$ for any $c\in D(0,1)$ Moroz used $\tilde{O}((\ell+\tau)d)$ bit operations to compute a set of complex $c_{i}$ and positive $\rho_{i}$ for $i=1,\dots,K$ and $K=O(d/\ell)$ such that $D(0,1)\subseteq\cup_{i=1}^{K}D(c_{i},\rho_{i})$; then he ensured approximation to $p(c)$ for every point $c\in D(0,1)$ within $\epsilon:=3\sum_{{}_{i}=0}^{d}|p_{i}|/2^{\ell}$ by means of approximation within $\epsilon/2$ of the lower degree polynomial $p(x-c_{i})\mod(x-c)^{\ell+1}$ for an appropriate $i$.
-At this point, application of [65, Alg.
+At this point, application of [68, Alg.
 5.3 and Thm.
-3.9] yielded new record fast algorithm for multipoint polynomial approximation.
-In Sec.
-11.4 we accelerate e/i tests by a factor of $m$ under a random root model and accordingly accelerate the solution of Problems 1-3 and $*{}^{*}$$*{}^{*}$ by such a factor.
-Moroz in [78] and Imbach and Moroz in [56] accelerated their solution under an upper bound on condition numbers of all zeros of $p(x)$ instead.
-In spite of the appeal to the concept of condition number, used for the design and analysis of polynomial root-finding in [13, 28, 21], this is a little stronger restriction than bounding cluster sizes of Def.
-B.1, under which [92, 99] proved such a complexity bound.
-Without using the algorithm of [78, Thm.
-2], our current algorithms still support such a near optimal bit operation cost bound for Problem 1 with no restriction on an input polynomial $p$ in the case where $m^{2}=\tilde{O}(d)$, $1\leq m\leq d$ (see [108, Remark 20.1]) as well as under the Random Root Model of Sec.
-1.5.
+3.9] yields new record fast algorithm for multipoint approximation of a polynomial given with its coefficients.
+The root-finders of Moroz in [81] and Imbach and Moroz in [59] use $\tilde{O}(bd)$ bit operations for polynomials given with their coefficients and having only well-conditioned zeros.
+In spite of the appeal to the concept of condition number, used for the design and analysis of polynomial root-finding in [14, 30, 22], the upper bound on the condition numbers of all zeros of $p(x)$ in [81, 59] is a little stronger restriction than the bound on the cluster sizes of Def.
+24.1, under which [96, 103] proved the same complexity bound.
+Unlike our present root-finders, the root-finders of [81, 59] are not efficient for the worst case inputs and extensively use the coefficients.
 5.3 Implementation: State of the Art
-The algorithm of [130] was implemented by Xavier Gourdon in [44] for the Magma and PARI/GP computer algebra systems, while the near optimal but more involved algorithm of [92, 99] has never been implemented.
-Our current algorithms are much simpler for implementation than that of [92, 99] and, if properly implemented, can become user’s choice root-finders.
-Already incorporation in 2020 of some of our techniques into the previous best implementation of subdivision root-finder of [61] has enabled its noticeable acceleration (see [57]).
-The more advanced publicly available implementation in [60] “relies on novel ideas and techniques" from [108] (see [60, Sec.
+The near-optimal but quite involved algorithm of [96, 103] has never been implemented, while the slower algorithm of [134] has been implemented by Xavier Gourdon in [46] for the Magma and PARI/GP computer algebra systems.
+By now it is superseded by implementation of various non-optimal complex polynomial root-finders such as MATLAB “roots”, Maple “solve”, CCluster, and MPSolve; the latter one relies on Ehrlich’s iterations and is user’s current choice for approximation of all $d$ complex roots (in Sec.
+26.4 we list some competitive implementations of Newton’s iterations).
+Current user’s choice for real root-finding is the algorithm of [70], although it has already been superseded in [62] according to extensive tests for standard test polynomials.
+We present algorithms that are rather simple for implementation and promise to compete for user’s choice.
+Already incorporation in 2020 of some of our techniques into the previous best implementation of subdivision root-finder [64] has enabled its noticeable acceleration (see [60]).
+The more advanced publicly available implementation in [63] “relies on novel ideas and techniques" from [112] (see [63, Sec.
 1.2]).
-According to the results of extensive tests with standard test polynomials, the algorithm has performed dramatically faster than [61] or any other known variant of complex subdivision root-finder, has at least competed with user’s choice package MPSolve even for approximation of all $d$ zeros of a polynomial $p$ given with its coefficients, [^17] The complexity of a subdivision root-finder in a region is proportional to the number of roots in it, while MPSolve is about as fast and slow for all roots as for their fixed subset.
-MPSolve implements Ehrlich-Aberth’s root-finding iterations, which empirically converge to all $d$ roots very fast right from the start but only under empirical support and only under initialization that operates with the coefficients of $p$ [125].
-and tended to supersede MPSolve significantly as the degree $d$ of $p$ grew large.
-Such a performance could only be a dream for the designers of subdivision root-finders before 2022, but there is a large room for further improvement because [60] still inherits some earlier algorithms from [22, 61], which are slower by order of magnitude than our current ones (see some specific directions next).
-1.10 Tentative directions for further progress
-Next we comment on some specific directions for significant (and perhaps dramatic) acceleration of the root-finders of [57, 60], based partly on our present work and partly on its expected further advance.
-Since our algorithms are near optimal, we only seek relaxation of the assumptions of the Random Root Model and empirical acceleration.
-(A) Acceleration of e/i tests. They are the most used ingredient of our subdivision root-finders as well as those of [57, 60].
-(i) In Part III we by-pass approximation of the Cauchy integral of (1.9) and only evaluate the polynomial $p$ at or near the unit circle $C(0,1)$ (see Sec. 11.4).
-(ii) We achieve additional acceleration under the Random Root Model because we only need to perform root-lifting and then to evaluate $p$ at a small number of random points $x$.
-The cost of root-lifting is dominated in our formal study but can substantially slow down implementation.
-We conjecture that empirically the algorithm should work without lifting or with a small number of lifting steps.
-(iii) E/i tests by means of approximation of the distance from a complex point to the roots seems to be competitive or maybe even superior empirically (see Remark 1.2).
-(iv) We can try to exploit potential benefits of fast numerical approximation of Cauchy integrals with the Exponentially Convergent Adaptive Trapezoidal Quadrature (see [141]), efficiently implemented in Boost C++ libraries.
-This promises no significant acceleration for the worst case inputs but may be useful empirically.
-(B) Acceleration of compression of a disc.
-We accelerate the compression in [57, 60] at two stages:
-(i) The algorithms of Secs. 20.2 and 20.3 are much faster than those of [57, 60], based on Cauchy sum, for computation of a point that lie in or near the convex hull of the root set in a fixed disc.
-(ii) The algorithm of Sec.
-12.1 runs significantly faster than those in [57, 60] for approximation of the smallest root radius from a given point on the complex plane, and those based on the DLG iterations (see Remark 1.2) are potentially competitive.
-Notice that we specified and analyzed extension from approximation of the distance to roots to approximation of the distance to the $m$ closest root in a disc containing precisely $m$ roots (see Sec.
-12).
-(C) Acceleration of real and near real root-finding. Our novel algorithm of Sec. 9 can be readily implemented and seems to be the first that efficiently solves this important problem.
-(D) Acceleration of approximation of matrix eigenvalues.
-We adopt extension of root-finding to proposed in [73] because it implies record low upper estimates in the case of matrices and matrix polynomials of very large size, but this extension relies on the computation of the coefficients of the characteristic polynomial of an input matrix or matrix polynomial, which is prone to numerical stability problems and avoided in practice of numerical matrix computations.
-In Remark 7.1 of Sec.
-7 we point out an alternative extension, which avoids such a notorious stage and is dramatically faster where the input matrix or matrix polynomial is data sparse.
+According to the results of extensive tests with standard test polynomials, the algorithm has performed dramatically faster than [64], has at least competed with user’s choice package MPSolve even for approximation of all $d$ zeros of a polynomial $p$ given with its coefficients, [^9] The computational cost of subdivision root-finders in a domain $\mathbb{D}$ as well as the root-finder of [96, 103] decreases at least proportionally to the number of roots in $\mathbb{D}$, while Ehrlich’s, Weierstrass’s, and a number of other functional iterations approximate the zeros of $p$ in a domain $\mathbb{D}$ almost as slow as all its $d$ complex roots.
+Thus already implementation in [64] of subdivision iterations, slightly outperformed MPSolve for root-finding in a domain $\mathbb{D}$ containing a small number of roots although is by far inferior for the approximation of all $d$ roots.
+and tended to supersede MPSolve more and more significantly as the degree $d$ of $p$ grew large.
+There is still a large room for further improvement because the root-finder of [63] inherits some earlier sub-algorithms from [23, 64], which are slower by order of magnitude than our current ones.
+Next we comment on some promising directions to further improvement of [63], but also notice that adjustment of our near-optimal subdivision root-finder in Remark 2.1 to approximation of real and near-real roots and consequently of matrix eigenvalues should be at least competitive with the current user’s choice [70] for real root-finding as well as with its significant improvement in [62]; moreover, it more efficiently counters the impact of the input and rounding errors.
+5.4 Further progress and promise
+(A) Acceleration of e/i tests.
+E/i tests are basic ingredient of subdivision root-finders.
+[60, 63] used the Cauchy-based e/i tests of our Part II and their heuristic acceleration.
+It is still a challenge to compare their empirical performance with that of the formally supported e/i tests of our Part III and Sec.
+23.4, based on root-lifting.
+Furthermore, we can try to exploit potential benefits of fast numerical approximation of Cauchy integrals with the Exponentially Convergent Adaptive Trapezoidal Quadrature (see [146]), efficiently implemented in Boost C++ libraries.
+(B) Acceleration of the compression of a disc.
+We assume that the unit disc $D(0,1)$ contains $m$ roots and is isolated from the external $d-m$ roots, and we compress that disc towards the minimal disc containing the same $m$ roots.
+We reduce this task essentially to (i) computing a point in or near the convex hull of the root set in $D(0,1)$ and (ii) approximation of the $m$th smallest root radius from a given point in $D(0,1)$.
+For task (i) we dramatically accelerate our algorithm of [60, 63] because our algorithms of Secs.
+23.2 and 23.3 for task (i), using Schröder’s iterations, deterministic, and formally supported, run dramatically faster than those of [60, 63], based on approximation of the average of all roots in $D(0,1)$.
+For task (ii) we can apply our Las Vegas randomized algorithms of Secs.
+8 and 9 combined or the alternative of [60, 63], based on the algorithm of our Sec.
+15.1, which reduces this task to performing a constant number of e/i tests.
+Both approaches support the same asymptotic cost bound up to polylog factors, but fro $m\gg\log^{2}(d)$ the second approach is faster by a factor of $m/\log^{2}(d)$ where it uses fast heuristic e/i tests.
+Such an improvement, however, may be lost in practice because of the impact of polylog factors.
+Implementation of compression in [60, 63] is slowed down with verification based on Pellet’s theorem, but we can drop this verification because we readily detect highly unlikely output errors of our Las Vegas root-finders.
+(C) Fast approximation of the eigenvalues of a matrix.
+[20] explored reduction of approximation of the eigenvalues of a matrix to root-finding by means of Ehrlich’s iterations, reduced to recursive computations of NIR$(x)$, and showed benefits of this approach.
+We can readily extend this reduction and its benefits by using subdivision iterations instead of Ehrlich’s.
+Furthermore, win the case of subdivision we obtain additional benefits, e.g., our algorithms are much faster for approximation of a small number of eigenvalues that lie in a fixed domain and unlike Ehrlich’s iterations isolate eigenvalues, rather than just approximate them.
+(D) Acceleration of real and near real root-finding.
+In Remark 2.1 we adjust our near-optimal subdivision iterations to real and near-real root-finding, which counters the impact of input and rounding errors and can be readily extended to approximation of the eigenvalues of a matrix.
+(E) Acceleration by using fast multipoint polynomial evaluation.
+Mere incorporation of fast algorithms of [81, 59] should greatly accelerate the performance of [60, 63] in the case of root-finding for polynomials given with their coefficients.
 5.5 Organization of the rest of the paper
 In Parts II and III we devise and analyze our two alternative e/i tests.
-Deterministic e/i test of Part II is slower than that of Part III, but the root-finders of both parts use Las Vegas randomization to bound computational precision and bit-operation cost.
-These bounds are roughly the same in both parts, except thatin Part III we yield acceleration by a factor of $d$ under the Random Root Model.
-All the bit-operation complexity bounds of Parts II and III are is inferior by a factor of $b$ to that of Thm.
-1.3.
-We complete the proof of Thm. 1.3 in Part IV by means of our novel acceleration of subdivision iterations based on compression of a disc without losing its roots.
+Deterministic e/i test of Part III is as fast as the randomized one of Part II.
+We use Las Vegas randomization in both parts to bound computational precision and bit-operation cost, and in Part III we yield acceleration by a factor of $d$ under the Random Root Model.
+The bit-operation complexity bounds of Parts II and III are inferior by a factor of $b$ to that of [96, 97, 103], but we match it in Part IV by means of our acceleration of the algorithms of both Parts II and III based on compression of a disc without losing its roots.
+In Part V we cover some important root-finding subjects that involve no substitution iterations, namely, factorization, root isolation, and functional iterations.
+In Secs.
+24 and 25 we outline the fundamental results of Schönhage from [134] on polynomial factorization, its links to root-finding and root-isolation, and some relevant bit-complexity estimates, including lower bounds.
 In Sec.
-A of the Appendix we comment on applications and implementation of root-finding.
-In its Secs.
-B and C we outline the fundamental results of Schönhage from [130] on polynomial factorization, its links to root-finding, and root-isolation and some relevant bit-complexity estimates.
-In Sec.
-D we discuss root-finding by means of functional iterations, namely, Newton’s, Weierstrass’s (aka Durand-Kerner’s), and Ehrlich’s (aka Aberth’s), and strengthen them with incorporation of Fast Multipole Algorithm (FMM).
+26 we discuss root-finding by means of functional iterations, namely, Newton’s, Weierstrass’s (aka Durand-Kerner’s), and Ehrlich’s (aka Aberth’s), and their acceleration by means of incorporation of Fast Multipole Algorithm (FMM) and fast multipoint polynomial evaluation as well as their initialization both where the coefficients of $p$ are unknown or given to us.
 5.6 Independence of Part III from Part II
-In Parts II an III we solve Problems 0, 1, $*{}^{*}$, and $*{}^{*}$ in two different ways.
+In Parts II an III we solve Problems 0, 1, 0, and 1 in two different ways.
 In Part II we rely on approximation of a Cauchy integral associated with root-counting in a disc on the complex plane.
-In Part III we rely on lifting the zeros of $p$ to a fixed power $h=O(\log(d))$.
+In Part III we rely on lifting the roots to a fixed power $h=O(\log(d))$.
 In addition, in Part III, unlike Part II, we devise a faster root-finder under a random root model.
@@
 PART II: Classical subdivision root-finder with novel exclusion/inclusion (e/i) tests based on approximation of Cauchy integrals
-[…]
+Organization of the presentation in Part II: We devote the next section to some basic definitions.
+In Sec.
+7 we study root-counting in a disc on the complex plane by means of approximation of a Cauchy integral over its boundary circle by finite sums.
+[^10] We do not use the results of Sec.
+7.5, but they can be of independent interest.
+In Sec.
+8 we apply this study to devise and analyze root-counters and e/i tests in a disc on the complex plane.
+In Sec.
+9 we extend our e/i tests to estimating root radii, that is, the distances from a fixed complex point to the roots.
+In Secs.
+10 we estimate precision of computing of our root-finders.
+We devote short Sec.
+11 to their extension to approximation of the eigenvalues of a matrix or a matrix polynomial.
+In Sec.
+12 we estimate the bit operation complexity of our root-finders assuming that $p(x)$ is a general polynomial of (1.1) represented with its coefficients.
 6 Background: basic definitions
+  • “Wlog" stands for “without loss of generality".
   • $S(c,\rho):=\{x:~{}|\Re(c-x)|\leq\rho$, $|\Im(c-x)|\leq\rho\},~{}D(c,\rho):=\{x:~{}|x-c|\leq\rho\}$,
@@
     denote a square, a disc, a circle (circumference), and an annulus (ring) on the complex plane, respectively.
-  • $\Delta(\mathbb{R})$, $X(\mathbb{R})$, $CH(\mathbb{R})$, and $\#(\mathbb{R})$ denote the diameter, the root set, the convex hull, and the index (root set’s cardinality) of a region $\mathbb{R}$ on the complex plane, respectively.
-  • A disc $D(c,\rho)$, a circle $C(c,\rho)$, or a square $S(c,\rho)$ is said to be $\theta$-isolated for $\theta>1$ if $X(D(c,\rho))=X(D(c,\theta\rho))$, $X(C(c,\rho))=X(A(c,\rho/\theta,\rho/\theta))$, or $X(S(c,\rho))=X(S(c,\theta\rho))$, respectively.
+  • $\Delta(\mathbb{S})$, $X(\mathbb{S})$, $CH(\mathbb{S})$, and #$\#(\mathbb{S})$ denote the diameter, the root set, the convex hull, and the index (root set’s cardinality) of a set $\mathbb{S}$ in the complex plane, respectively.
+  • A disc $D(c,\rho)$, a circle $C(c,\rho)$, or a square $S(c,\rho)$ is said to be $\theta$-isolated (from external roots) for $\theta>1$ if $X(D(c,\rho))=X(D(c,\theta\rho))$, $X(C(c,\rho))=X(A(c,\rho/\theta,\rho/\theta))$, or $X(S(c,\rho))=X(S(c,\theta\rho))$, respectively.
   • The largest upper bound on such a value $\theta$ is said to be the isolation of the disc $D(c,\rho)$, the circle $C(c,\rho)$, or the square $S(c,\rho)$, respectively (see Fig.
-3), and is denoted $i(D(c,\rho))$, $i(C(c,\rho))$, and $i(S(c,\rho))$, respectively.
-  • $r_{1}(c,t)=|y_{1}-c|,\dots,r_{d}(c,t)=|y_{d}-c|$ in non-increasing order are the $d$ root radii, that is, the distances from a complex center $c$ to the zeros $y_{1}\dots,y_{d}$ of a $d$th degree polynomial $t(x)$.
+4), and is denoted $i(D(c,\rho))$, $i(C(c,\rho))$, and $i(S(c,\rho))$, respectively.
+  • $r_{1}(c,t)=|y_{1}-c|,\dots,r_{d}(c,t)=|y_{d}-c|$ in non-increasing order are the $d$ root radii, that is, the distances from a complex point $c$ to the zeros $y_{1}\dots,y_{d}$ of a $d$th degree polynomial $t(x)$.
 $r_{j}(c):=r_{j}(c,p)$, $r_{j}:=r_{j}(0)$ for $j=1,\dots,d$.
-[…]
-Observation 2.1.
+  • Generalize Def.
+2.1 of $\sigma$-soft e/i tests as follows: for $p(x)$ of (1.1), $\sigma>1$, and integers $\ell$ and $m$ such that $1\leq\ell\leq m\leq d$ and #$\#(D(0,1))\leq m$, a $\sigma$-soft $\ell$-test, or just $\ell$-test for short (so that e/i test turns into 1-test), either outputs 1 and stops if it detects that $r_{d-\ell+1}\leq\sigma$, that is, #$\#(D(0,\sigma))\geq\ell$, or outputs 0 and stops if it detects that $r_{d-\ell+1}>1$, that is, [^11] Bounds $r_{d-\ell+1}\leq\sigma$ and $r_{d-\ell+1}>1$ can hold simultaneously (cf.
+Fig.
+1b), but as soon as an $\ell$-test verifies any of them, it stops without checking if the other bound also holds.
+#$\#(D(0,1))<\ell$.
+$\ell$-testc,ρ, aka $\ell$-test for the disc $D(c,\rho)$, is an $\ell$-test applied to the polynomial $t(y)$ of (1.3).
+Observation 6.1.
 Eqn. (1.3) maps the zeros $z_{j}$ of $p(x)$ into the zeros $y_{j}=\frac{z_{j}-c}{\rho}$ of $t(y)$, for $j=1,\dots,d$, and preserves the index #$\#(D(c,\rho))$ and the isolation $i(D(c,\rho))$.
@@
 7.2 Cauchy sums in the unit disc as weighted power sums of the roots
-Cor. 3.1 of the following lemma expresses the $h$th Cauchy sum $s_{h,q}$ as the sum of the $h$th powers of the roots $z_{j}$ with the weights $\frac{1}{1-x^{q}}$.
+Cor. 7.1 of the following lemma expresses the $h$th Cauchy sum $s_{h,q}$ as the sum of the $h$th powers of the roots $z_{j}$ with the weights $\frac{1}{1-x^{q}}$.
 Lemma 7.1.
@@
 $\frac{1}{1-y}=\sum_{u=0}^{\infty}y^{u}~{}{\rm for}~{}y=\frac{z}{\zeta^{g}}.$
-We can apply it because $|y|=|z|$ while $|z|<1$ by assumption.
+We can apply it because $|y|=|z|$, while $|z|<1$ by assumption.
 Sum the above expressions in $g$, recall that
@@
 Hence (7.3) holds in the case where $|z|>1$ as well. ∎
-Sum equations (3.3) for $z=z_{j}$ and $j=1,\dots,d$, combine Eqns. 1.13 and (3.2), and obtain
+Sum equations (7.3) for $z=z_{j}$ and $j=1,\dots,d$, combine Eqns. 3.1 and (7.2), and obtain
 Corollary 7.1.
-For the zeros $z_{j}$ of $p(x)$ and all $h$, the Cauchy sums $s_{h,q}$ of (3.2) satisfy $s_{h,q}=\sum_{j=1}^{d}\frac{z_{j}^{h}}{1-z_{j}^{q}}~{}{\rm unless}~{}z_{j}^{q}=1~{}{\rm for~{}some}~{}j.$
+For the roots $z_{j}$ and all $h$, the Cauchy sums $s_{h,q}$ of (7.2) satisfy $s_{h,q}=\sum_{j=1}^{d}\frac{z_{j}^{h}}{1-z_{j}^{q}}$ unless $z_{j}^{q}=1~{}{\rm for~{}some}~{}j.$
 7.3 Approximation of the power sums in the unit disc
-Cor. 3.1 implies that the values $|s_{h,q-}s_{h}|$ decrease exponentially in $q-h$ if the unit disc $D(0,1)$ is isolated.
+Cor. 7.1 implies that the values $|s_{h,q-}s_{h}|$ decrease exponentially in $q-h$ if the unit disc $D(0,1)$ is isolated.
 Theorem 7.1.
-[130, Eqn.
+[134, Eqn.
 (12.10)].
@@
 Proof.
-Enumerate the roots $z_{j}$ so that $|z_{j}|<1$ if and only if $j\leq d_{\rm in}$; then Cor. 3.1 implies that
+Enumerate the roots $z_{j}$ so that $|z_{j}|<1$ if and only if $j\leq d_{\rm in}$; then Cor. 7.1 implies that
 $s_{h,q}-s_{h}=s_{h,q,{\rm in}}+s_{h,q,{\rm out}}$ (7.6)
@@
 Example 7.1.
-Let the unit circle $C(0,1)$ be 2-isolated. Then Thm. 3.1 implies that $|s_{0,q}-s_{0}|<1/2$ for $d\leq 1,000$ if $q=10$ and for $d\leq 1,000,000$ if $q=20$.
+Let the unit circle $C(0,1)$ be 2-isolated. Then Thm. 7.1 implies that $|s_{0,q}-s_{0}|<1/2$ for $d\leq 1,000$ if $q=10$ and for $d\leq 1,000,000$ if $q=20$.
 Algorithm 7.1.
@@
 COMPUTATIONS: Compute Cauchy sum $s_{0,q}$ and output an integer $\bar{s}_{0}$ closest to it.
-Observation 3.1.
+Observation 7.1.
 (i) Alg.
@@
 Proof.
-Thm. 3.1 immediately implies claim (ii) but also implies that #$\#(D(0,1))>0$ if $\bar{s}_{0}>0$ unless #$\#(A(0,1/\theta,\theta))>0$. In both cases #$\#(D(0,\theta))>0$. ∎
+Thm. 7.1 immediately implies claim (ii) but also implies that #$\#(D(0,1))>0$ if $\bar{s}_{0}>0$ unless #$\#(A(0,1/\theta,\theta))>0$. In both cases #$\#(D(0,\theta))>0$. ∎
 7.4 Extension to any disc
-Map (1.4), preserving #$\#(D(0,1))$ (cf. Observation 2.1), enables us to extend definition (3.2) of Cauchy sum to any disc $D(c,\rho)$ as follows:
+Map (1.3), preserving #$\#(D(0,1))$ (cf. Observation 6.1), enables us to extend definition (7.2) of Cauchy sum to any disc $D(c,\rho)$ as follows:
 $s_{0,q}(p,c,\rho):=s_{0,q}(t,0,1):=\frac{\rho}{q}\sum_{g=0}^{q-1}\zeta^{g}~{}\frac{p^{\prime}(c+\rho\zeta^{g})}{p(c+\rho\zeta^{g})},$ (7.8)
 for $\zeta$ of (3.5), that is, $s_{h,q}(p,c,\rho)$ is the Cauchy sums $s_{h,q}(t,0,1)$ of the zeros of the polynomial $t(y)=p(\frac{y-c}{\rho})$ in the unit disc $D(0,1)$.
-Observation 3.2.
+Observation 7.2.
 Given a polynomial $p(x)$, a complex $c$, a positive $\rho$, a positive integer $q$, and the $q$-th roots of unity, we can compute $s_{0,q}(p,c,\rho)$ by means of the evaluation of NIR$(\zeta^{g})$ for $g=0,1,\dots,q-1$, reduced to evaluation of $p$ at $2q$ points, and in addition performing $q+1$ divisions, $12q-$ 1 multiplications, and $12q-$ 1 additions.
-[…]
+Refer to Alg.
+7.1 applied to the polynomial $t(y)$ of (1.3) for an isolation parameter $\theta>1$ and $q=\lfloor\log_{\theta}(4d+2)\rfloor$ as Alg.
+7.1c,ρ,q and also as Alg.
+7.1 applied to $\theta$-isolated circle $C(c,\rho)$ for $2q$ evaluation points.
+Combine Observations 6.1 and 7.2 to extend Observation 7.1 as follows.
 Proposition 7.1.
-[…]
+(i) Alg.
+7.1c,ρ,θ evaluates polynomial $p(x)$ at $q=\lfloor\log_{\theta}(4d+2)\rfloor$ points $x$ and in addition performs $25q-$ 2 ops.
+(ii) It outputs $\bar{s}_{0}=\#(D(c,\rho))$ if the circle $C(c,\rho)$ is $\theta$-isolated.
+(iii) If the algorithm outputs $\bar{s}_{0}>0$, then #$\#(D(c,\theta\rho))>0$.
 By applying Thm.
@@
 [^13] We will not use the results of this subsection, but they can be of independent interest.
-Recall that $s_{0,q}=\sum_{j=1}^{d}\frac{1}{1-z_{j}^{q}}$ unless $z_{j}^{q}=1$ for some $q$ (cf. Cor. 3.1), write $z:=z_{j}^{q}$ and estimate the contribution of a roots $z_{j}$ to $\Re(s_{0,q})$.
+Recall that $s_{0,q}=\sum_{j=1}^{d}\frac{1}{1-z_{j}^{q}}$ unless $z_{j}^{q}=1$ for some $q$ (cf. Cor. 7.1), write $z:=z_{j}^{q}$ and estimate the contribution of a roots $z_{j}$ to $\Re(s_{0,q})$.
 Proposition 7.3.
@@
 Corollary 7.2.
-Under the assumptions of Prop. 3.3 it holds that
+Under the assumptions of Prop. 7.3 it holds that
 $\Re\Big{(}\frac{1}{1-z}\Big{)}\geq 0~{}{\rm if}~{}\rho\leq 1,$ (7.11)
@@
 ∎
-Next we outline but do not elaborate upon some potential links of Cor. 3.2 to root-counting. Let the unit disc $D(0,1)$ be $\theta$-isolated, combine bounds (3.15) and (3.16), and obtain
+Next we outline but do not elaborate upon some potential links of Cor. 7.2 to root-counting. Let the unit disc $D(0,1)$ be $\theta$-isolated, combine bounds (7.15) and (7.16), and obtain
 $|\Re(s_{0,q})-s_{0}|\leq\frac{d}{\theta^{q}-1}.$
@@
 Proof.
-Combine Cor. 3.1 for $h=0$ with equation (3.12). ∎
+Combine Cor. 7.1 for $h=0$ with equation (7.12). ∎
 8 Root-counting and $\ell$-tests
 8.1 Overview
-We seek fast root-counters and $\ell$-tests based on computing Cauchy sum $s_{0,q}$ for a smaller integer $q$, but the output can deceive us if we decrease $q$ too much: the polynomial $p(x):=(1-wx^{d-m})(x^{m}-dx+w)$ has zero near the origin if $w\approx 0$, while NIR$(x)=p^{\prime}(x)/p(x)$ vanishes for $x$ equal to any $(d-1)$st root of unity.
-Such deceptive e/i tests have been extremely rare guests in the extensive experiments in [57, 60] for $q$ of order $\log^{2}(d))$ and were completely excluded where the values $|s_{h,q}|$ were required to be small for $h=0,1,2$, but formal support for this empirical behavior is still a challenge.
+We seek fast root-counters and $\ell$-tests based on computing Cauchy sum $s_{0,q}$ for a smaller integer $q$, but the output can deceive us if we decrease $q$ too much: the polynomial $p(x):=x^{d}-dx+w$ has zero near the origin if $w\approx 0$, while NIR$(x)=p^{\prime}(x)/p(x)$ vanishes for $x$ equal to any $(d-1)$st root of unity.
+Such deceptive e/i tests have been very rare guests in the extensive experiments in [60, 63] for $q$ of order $\log^{2}(d))$ and were completely excluded where the values $|s_{h,q}|$ were required to be small for $h=0,1,2$, but formal support for this empirical behavior is still a challenge.
 In the next two subsections we formally support Las Vegas randomized root-counters and $\ell$-tests that evaluate polynomial $p$ at $O(m\log(d))$ points.
-One can probably obtain similar results by computing the Cauchy sum $s_{0,q}$ for the polynomial $p(ax)$ at $O(m\log(d))$ random values $a\in C(0,1)$, but we leave formal support of this conjecture as a research challenge.
-Recall that randomization in our root-finders is of Las Vegas type (see Sec.
-1.7), while we apply randomization in Sec.
-6 anyway – to bound computational precision in our root-finders.
-For the sake of completeness we also present a deterministic $\ell$-test in Sec.
-4.4.
+Combine this estimate with Cor. 2.2 for $k$ of order $b$ and obtain Las Vegas randomized solution of Problems 0, 1, 0, and 1 at the cost of evaluation of the polynomial $p$ at $O(m^{2}b)$ points.
 8.2 A basic randomized root-counter
@@
 INITIALIZATION: Sample a random value $w$ in the range $[0.2,0.4]$ under the uniform probability distribution in that range and write $\rho:=2^{w}$ and $q=\lfloor 10m\gamma\log_{2}(4d+2)\rfloor$.
-[…]
+COMPUTATIONS: Apply Alg. 7.10,ρ,q.
 Proposition 8.1.
-Alg. 4.1 queries evaluation oracle at $q=\lfloor 10m\gamma\log_{2}(4d+2)\rfloor$ locations and outputs the value $\bar{s}_{0}=\#(D(0,\rho))$ with a probability at least $1-1/\gamma$.
+Alg. 8.1 queries evaluation oracle at $q=\lfloor 10m\gamma\log_{2}(4d+2)\rfloor$ locations and outputs the value $\bar{s}_{0}=\#(D(0,\rho))$ with a probability at least $1-1/\gamma$.
 Proof.
@@
 8.1 outputs $\bar{s}_{0}=s_{0}=\#(D(0,\rho))$ by virtue of claim (ii) of Prop.
-3.1.
+7.1.
 ∎
-4.3 Refining the bound on error probability
-By reapplying Alg. 4.1 $v$ times we evaluate the polynomial $p$ at $v$ times more points, but we will prove that this decreases the Las Vegas error probability dramatically – below $1/\gamma^{v}$.
+8.3 Refining the bound on the error probability
+When we apply Alg.
+8.1 $v$ times, the number of query locations of the evaluation oracle grows linearly – by a factor of $v$, while the bound on Las Vegas error probability of outputting errors decreases exponentially – from $1/\gamma$ to $1/\gamma^{v}$.
 First let $\ell=1$ or $\ell=m$ and then again reduce the root radius approximation problem to the decision problem of $\ell$-test.
@@
 7.1 if $\ell=1$; similarly the output 0 is certified if $\ell=m$.
-Next extend the $\ell$-test by means of applying Alg. 4.1 for $v$ iid random variables $w$ in the range $0.2\leq w\leq 0.4$. Call this $\ell$-test Alg. $\ell$.
+Next extend the $\ell$-test by means of applying Alg. 8.1 for $v$ iid random variables $w$ in the range $0.2\leq w\leq 0.4$. Call this $\ell$-test Alg. $\ell$.
 […]
-Recall Prop. 4.1 and then readily verify the following proposition.
+Recall Prop. 8.1 and then readily verify the following proposition.
 Proposition 8.2.
@@
 $\ell$ evaluates $p(x)$ at $\lfloor 10m\gamma\log_{2}(4d+2)\rfloor v$ points; this evaluation involves more ops than the other stages of the algorithm, even in the cases of the Mandelbrot and sparse polynomials $p$ (cf.
-Observation 3.2).
+Observation 7.2).
 (ii) Output 1 of Alg.
-4.1v,1 and the output 0 of Alg.
-4.1v,$m$ are certified.
+8.1v,1 and the output 0 of Alg.
+8.1v,$m$ are certified.
 (iii) Output 0 of Alg.
-4.1v,1 and output 1 of Alg.
-4.1v,$m$ are correct with a probability at least $1-1/\gamma^{v}$.
+8.1v,1 and output 1 of Alg.
+8.1v,$m$ are correct with a probability at least $1-1/\gamma^{v}$.
 The proposition bounds the cost of $\ell$-tests for $\ell=1$ (e/i tests) and $\ell=m$.
@@
 Assign the values 0 or 1 to the output integers $g$ as follows: compute the average of the $v$ outputs in $v$ independent applications of the test and output $g$ denoting the integer closest to that average; if the average is 1/2, write $g=0$ to break ties.
-This extension of Alg. $\ell$ is only interesting for $1<\ell<m$; otherwise its cost/error estimates are inferior to those of Prop. 4.2.
+This extension of Alg. $\ell$ is only interesting for $1<\ell<m$; otherwise its cost/error estimates are inferior to those of Prop. 8.2.
 Proposition 8.3.
@@
 ∎
-4.4 Deterministic $\ell$-test
-In view of our remark at the end of Sec.
-4.1, randomization is only a nominal restriction for $\ell$-tests, but next we get rid of it, although at the price of increasing the numbers of points of evaluation of $p$ and bit operations involved in the tests by a factor of $m$.
-[…]
-Next we elaborate upon this test.
-Without loss of generality (wlog) let $c=0$, $\theta\rho=1$, $C(c,\theta\rho)=C(0,1)$.
-Algorithm 4.2.
-Deterministic $\ell$-test.
-INPUT: A black box polynomial $p$ of a degree $d$ and two positive integers $\ell$ and $m\geq\ell$ such that the unit disc $D(0,1)$ contains at most $m$ roots.
-INITIALIZATION: Fix
-$\sigma>1,~{}\theta=\sigma^{\frac{1}{4m+1}},~{}q=\lfloor\log_{\theta}(4d+2)\rfloor,~{}\rho_{i}=\frac{1}{\theta^{2i+1}},~{}i=0,1,\dots,2m.$ (4.1)
-OUTPUT: 0 if #$\#(D(0,1/\sigma))<\ell$ or 1 if #$\#(D(0,1))\geq\ell$.
-[…]
-Output 1 if at least $m+1$ integers $\bar{s}_{0}$ output by the algorithm (among all $12m+$ 1 such integers) exceed $\ell-1$.
-Otherwise output 0.
-(For $\ell=1$ we can output 1 already on a single output $\bar{s}_{0}>0$; for $\ell=m$ we can output 0 already on a single output $\bar{s}_{0}<m$.)
-Proposition 4.4.
-Alg. 4.2 is a $\sigma$-soft $\ell$-test for $\sigma>1$ of our choice (see (4.1)).
-Proof.
-[…]
-[…]
-To complete the proof, combine this property with the relationships $\rho_{2m}=1/\theta^{4m+1}$ (cf.
-(4.1)), $\rho_{0}=1/\theta<1$, and #$\#(D(0,\rho_{i+1}))\leq\#(D(0,\rho_{i}))$, which hold because $D(0,\rho_{i+1})\subseteq(D(0,\rho_{i})$ for $i=0,1,\dots,2m-1$.
-∎
-Observation 4.1.
-[Cf.
-Observation 3.2.] Alg.
-4.2 evaluates polynomial $p(x)$ at $(2m+1)q$ points for $q=\lfloor\log_{\theta}(4d+2)\rfloor=\lfloor(4m+1)\log_{\sigma}(4d+2)\rfloor$, which dominates the cost of the remaining ops involved, even in the cases of the Mandelbrot and sparse input polynomials and which is in $O(m^{2}\log(d))$ if $\sigma-1$ exceeds a positive constant.
 9 Root radius estimation based on $\ell$-test
-In this section we assume that we are given an algorithm for $\ell$-test and go back from the decision problem of $\ell$-test to approximation of $r_{d-\ell+1}(c^{\prime})$, the $\ell$th smallest root radius from a fixed complex point $c^{\prime}$; we assume that $r_{d-\ell+1}(c^{\prime})$ lies in a fixed range $[a_{-},a_{+}]$.
-[^22] For our root-finders we only need to approximate $r_{d-\ell+1}(c^{\prime})$ for $\ell=1$ and $\ell=m$ and only with a constant bound on the relative error, but we will estimate $r_{d-\ell+1}(c^{\prime})$ within $1/2^{b}$ for any real $b$ and $\ell$, $1\leq\ell\leq m$, which is of independent interest for root-finding.
-Wlog let this be the range $[\epsilon,1]$ for a fixed small positive $\epsilon=1/2^{b}$.
-The bisection algorithm is too slow for us, and we accelerate it by applying it to approximate the exponent of a target value.
-Call that algorithm bisection of exponent or BoE and apply it to bracket the $\ell$th smallest root radius $r_{d-\ell+1}$ in a range $[\rho_{-},\rho_{+}]$ such that $0<\rho_{<}-\rho_{+}\leq\epsilon$ or $0<\rho_{+}<\beta\rho_{-}$ for two fixed tolerance values $\epsilon=1/2^{b}$ and $\beta=2^{\phi}>1$.
-[…]
-Since the test is soft, so is BoE as well: for a range $[b_{-},b_{+}]$ with midpoint $\widehat{b}=0.5(b_{+}+b_{-})$ regular rigid bisection continues the search in the range $[b_{-},\widehat{b}]$ or $[\widehat{b},b_{+}]$, but our soft BoE continues it in the range $[b_{-},(1+\alpha)\widehat{b}]$ or $[(1-\alpha)\widehat{b},b_{+}]$ for a fixed small positive $\alpha$.
-The approximations still converges linearly and just a little slower than the regular rigid bisection: the approximation error decreases by a factor of $\frac{2}{1+\alpha}$ rather than by twice.
-[^23] If we applied BoE to the range $[0,1]$, we would have to bisect the infinite range $[-\infty,0]$, but we deal with the narrower range $[-b,0]$ since we have narrowed the range $[0,1]$ to $[\epsilon,1]$.
+In this section we assume that we are given an algorithm for $\ell$-test and seek approximation of $r_{d-\ell+1}(c^{\prime})$, the $\ell$th smallest root radius from a fixed complex point $c^{\prime}$, known to lie in a fixed range $[a_{-},a_{+}]$.
+[^14] For our root-finders we only need to approximate $r_{d-\ell+1}(c^{\prime})$ for $\ell=1$ and $\ell=m$, but we estimate $r_{d-\ell+1}(c^{\prime})$ for any $\ell$, $1\leq\ell\leq m$, which is of independent interest for root-finding.
+Wlog let this be the range $[\epsilon,1]$ for a fixed small positive $\epsilon=1/2^{b}$ and seek a range $[\rho_{-},\rho_{+}]$ such that $0\leq\rho_{-}\leq r_{d-\ell+1}(c^{\prime})\leq\rho_{+}$ and either $\rho_{+}\leq\epsilon$ or else $0<\rho_{+}<\beta\rho_{-}$ for two fixed tolerance values $\epsilon=1/2^{b}$ and $\beta=2^{\phi}>1$.
+We can obtain the desired approximation by applying bisection algorithm to the range $[\epsilon,1]$, but this is too slow for us, and we apply bisection of logarithms or BoL to approximate the logarithm of the target value in the range $[-b,0]=[\log_{2}(\epsilon),\log_{2}(1)]$.
+Assume that besides the above tolerance values the input of BoL includes a softness bound $\sigma=1+\alpha$, for a fixed small positive parameter $\alpha<1/2$, and a $\sigma$-soft $\ell$-testc,ρ, to be used as the pivot of BoL.
+Since the test is soft, so is BoL as well: for a range $[b_{-},b_{+}]$ with midpoint $\widehat{b}=0.5(b_{+}+b_{-})$ regular rigid bisection continues the search in the range $[b_{-},\widehat{b}]$ or $[\widehat{b},b_{+}]$, but our soft BoL continues it in the range $[b_{-},(1+\alpha)\widehat{b}]$ or $[(1-\alpha)\widehat{b},b_{+}]$ for a fixed small positive $\alpha$.
+Accordingly we slightly increase range of search for $r_{d-\ell+1}(c^{\prime})$ to $[-\bar{b},0]$ for $\bar{b}:=\log_{2}(\frac{\sigma}{\epsilon})$.
+BoL still converges linearly and just a little slower than the regular rigid bisection: the approximation error decreases by a factor of $\frac{2}{1+\alpha}$ rather than by twice.
+[^15] If we applied BoL to the range $[0,1]$, we would have to bisect the infinite range $[-\infty,0]$, but we deal with the narrower range $[-b,0]$ since we have narrowed the range $[0,1]$ to $[\epsilon,1]$.
 Algorithm 9.1.
 Bracketing the $\ell$-th smallest root radius based on $\ell$-test.
-[…]
+INPUT: Two integers $\ell$ and $m$ such that $1\leq\ell\leq m$ and #$\#(D(0,1))\leq m$, $\sigma=1+\alpha$ for $20<\alpha<1/$ 2, a $\sigma$-soft $\ell$-test0,ρ, and two tolerance values $\beta:=2^{\phi}>1$ and $\epsilon=1/2^{b}$, $b\geq 1$.
 OUTPUT: A range $[\rho_{-},\rho_{+}]$ such that $[\rho_{-},\rho_{+}]=[0,\epsilon]$ or $\rho_{-}\leq r_{d-\ell+1}\leq\rho_{+}\leq\beta\rho_{-}$.
@@
 This implies correctness of Alg.
-5.1.
+9.1.
 Let us estimate the number $\mathbb{I}$ of its invocations of an $\ell$-test.
 Proposition 9.1.
-Alg. 5.1 invokes an $\ell$-test $\mathbb{I}$ times where $\mathbb{I}\leq\lceil\log_{\nu}(\frac{\bar{b}}{\phi})\rceil+1$ for $\nu=\frac{2}{1+\alpha}$, $\mathbb{I}=O(\log(b/\phi))$.
+Alg. 9.1 invokes an $\ell$-test $\mathbb{I}$ times where $\mathbb{I}\leq\lceil\log_{\nu}(\frac{\bar{b}}{\phi})\rceil+1$ for $\nu=\frac{2}{1+\alpha}$, $\mathbb{I}=O(\log(b/\phi))$.
 Proof.
-The computations in Alg. 5.1 stop after the first application of an $\ell$-test if it outputs 1. It remains to count the subsequent $\ell$-tests if it outputs 0.
+The computations in Alg. 9.1 stop after the first application of an $\ell$-test if it outputs 1. It remains to count the subsequent $\ell$-tests if it outputs 0.
 Each test decreases the length $\bar{b}=b_{+}-b_{-}$ of its input range $[b_{-},b_{+}]$ for the exponent to $0.5(1+\alpha)\bar{b}$.
@@
 Let us specify these estimates for Alg.
-5.1.
+9.1.
 Corollary 9.1.
@@
 9.1 approximates the $\ell$-th smallest root radius $|z_{j}|$ within the relative error $2^{\phi}$ or determines that $|z_{j}|\leq\epsilon$, at the cost of evaluation of polynomial $p(x)$ at $\mathbb{I}\cdot\mathcal{NR}$ points for $\mathbb{I}=O(\log(\frac{b}{\phi}))$ estimated in Prop.
-5.1.
+9.1.
 10 Precision of computing in our root-finders
 10.1 Doubling computational precision towards its a posteriori estimation
-Given an algorithm $\mathbb{A}$ for black box polynomial computations and target tolerance TOL to output error, how should one choose computational precision?
-A simple practical recipe of [10, 14, 21] was the basis for highly successful performance root-finders of MPSolve: namely, first apply algorithm $\mathbb{A}$ with a lower precision and then recursively double it until an output error decreases below TOL, at which point it outputs precision within a factor of rwo from optimal, thus also supporting near optimal a posteriori bit operation complexity estimate for algorithm $\mathbb{A}$.
-In the rest of this section we obtain a priori estimate by proving that precision of $O(b\log(d))$ bits is sufficient for correctness of our subdivision root-finders.
+Given an algorithm $\mathbb{A}$ for black box polynomial computations and target tolerance TOL to output errors, how should one choose computational precision?
+Here is a simple practical recipe from [10, 15, 22] and MPSolve for obtaining a posteriori estimate within a factor of two from optimal: first apply algorithm $\mathbb{A}$ with a lower precision, then recursively double it and stop where an output error decreases below TOL.
+In the rest of this section we obtain a priori estimate for the precision of $O(b\log(d))$ bits.
 10.2 Reduction to approximation of $x$NIR$(x)$ within $\frac{3}{8}$
-Next we estimate precision of computations in our deterministic and randomized e/i tests that supports approximation of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ required in our root-finders; we only need to consider these tests performed for $\theta$-isolated discs $D(c,\rho)$ with $\theta-1>0$ of order $\frac{1}{m}$ where we approximate NIR$(x)$ with the first order divided difference of (1.6).
-Map (1.4) reduces such a test for $p(x)$ to the disc $D(0,1)$ for the polynomial $t(x)=p(\frac{x-c}{\rho})$ and decreases the error estimate by a factor of $\rho$, which is at most $2^{b}$ because we only apply our e/i tests to the disc of radius greater than $\epsilon=1/2^{b}$.
-By examining the impact of map (1.4) on the error and precision bounds we observe that the latter bound can increase by at most a factor of $b$.
-Thus we arrive at the claimed precision bound $O(b\log(d))$ in case of any disc $D(c,\rho)$ by proving bound $O(\log(d))$ in the case of the unit disc $D(0,1)$.
-The output of an e/i test is an integer, and we only need to certify that the output errors of Algs. 4.2 and 4.1 are less than $\frac{1}{2}$.
+In our e/i tests we approximate NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ with the first order divided difference of (3.2); we only apply the tests to $\theta$-isolated discs $D(c,\rho)$ for $\theta-1>0$ of order $\frac{1}{m}$ and $\rho\geq\epsilon=1/2^{b}$.
+Map (1.3) reduces such a test for $p(x)$ to a test in the disc $D(0,1)$ for the polynomial $t(x)=p(\frac{x-c}{\rho})$ replacing $p(x)$ and decreases the error estimate by a factor of $\rho$.
+Hence map (1.3) increases the precision bound by at most a factor of $b$.
+Thus we obtain the claimed precision bound in $O(b\log(d))$ in case of any disc $D(c,\rho)$ if we prove bound $O(\log(d))$ in the case of the unit disc $D(0,1)$.
+The output of an e/i test is an integer, and we only need to certify that the output error bound of Alg.
+9.1 is less than $\frac{1}{2}$.
+The error can be magnified by a factor of $v$ in $v$ applications of Alg.
+9.1 but this factor is immaterial in our notation $\tilde{O}$.
 Our e/i tests compute the $q$ values $\rho x\frac{p^{\prime}(c+\rho x)}{p(c+\rho x)}=x\frac{t^{\prime}(x)}{t(x)}$, for $x=\zeta^{g}$, $\zeta$ of (7.2), and $g=0,1,\dots,q-1$, sum these values, and divide the sum by $q$.
@@
 According to straightforward error analysis (cf.
-[20, Appendix A of Ch.3] and the references therein) rounding errors introduced by these additions and division are dominated, and we can ignore them; then it is sufficient to compute $\zeta^{g}{\rm NIR}(\zeta^{g})$ within, say, $\frac{3}{8}$ for every $g$.
+[21, Part A of Ch.3] and the references therein) rounding errors introduced by these additions and division are dominated, and we can ignore them; then it is sufficient to compute $\zeta^{g}{\rm NIR}(\zeta^{g})$ within, say, $\frac{3}{8}$ for every $g$.
 Indeed, in that case the sum of $q$ error bounds is at most $\frac{3q}{8}$ and decreases below $\frac{1}{2}$ in division by $q$.
-6.3 Precision of representation of $x$NIR$(x)$ within $\frac{1}{8}$
+10.3 Precision of representation of NIR$(x)$ within $\frac{1}{8}$
 Recall Eqn. (3.1) and represent $x{\rm NIR}(x)$ for $|x|=1$ within $\frac{1}{8}$ by using a precision of $O(\log(d))$ bits.
@@
 Thus, we shall increase the above error bound $\frac{1}{8}$ by adding to it the upper bounds $\frac{1}{8}$ on $\alpha:=|\frac{xt^{\prime}(x)}{t(x)}-\frac{xt^{\prime}(y)}{t(x)}|$ for $|x|=1$ and on the rounding error $\beta$ of computing $x\frac{t(x)-t(x-\delta)}{\delta t(x)}$.
-Next we estimate $\alpha$ in terms of $d$, $\theta$, and $\delta$ – by extending the proof of [73, Lemma 3.6], then readily ensure that $\alpha<\frac{1}{8}$ by choosing $\delta$ with a precision $\log_{2}(1/|\delta|)$ of order $\log(d)$, and finally estimate $\beta$.
+Next we estimate $\alpha$ in terms of $d$, $\theta$, and $\delta$ – by extending the proof of [76, Lemma 3.6], then readily ensure that $\alpha<\frac{1}{8}$ by choosing $\delta$ with a precision $\log_{2}(1/|\delta|)$ of order $\log(d)$, and finally estimate $\beta$.
 We simplify our estimates by dropping the factor $x$ where $|x|=1$.
 10.5 Precision of approximation with divided difference
-Recall the following lemma, which is [73, Fact 3.5].
+Recall the following lemma, which is [76, Fact 3.5].
 Lemma 10.1.
@@
 Proposition 10.2.
-Under the assumptions of Prop. 6.1 let $w=w(x):=\max_{j=1}\frac{1}{|x-y_{j}|}$. Then
+Under the assumptions of Prop. 10.1 let $w=w(x):=\max_{j=1}\frac{1}{|x-y_{j}|}$. Then
 $|t(x)|\alpha=|\frac{t(x)-t(x-\delta)}{\delta}-t^{\prime}(x)|\leq|t(x)\frac{(d~{}w)^{2}\delta}{1-\delta~{}d~{}w}|$.
@@
 The claimed bound on $|\xi^{\prime}|$ follows from (3.1).
-It remains to apply the first five lines of the proof of [73, Lemma 3.6] with $f$, $\alpha$, $n$, and $\xi$ replaced by $t$, $-\delta$, $d$, and $w$, s respectively.
+It remains to apply the first five lines of the proof of [76, Lemma 3.6] with $f$, $\alpha$, $n$, and $\xi$ replaced by $t$, $-\delta$, $d$, and $w$, s respectively.
 Namely, first obtain from Taylor’s expansion that $t(x)-t(x-\delta)=\sum_{j=0}^{\infty}\frac{\delta^{j}}{j!}t^{(j)}(x)-t(x)$.
@@
 $\Big{|}\frac{t(x)-t(x-\delta)}{\delta}-t^{\prime}(x)\Big{|}\leq\Big{|}\frac{f(x)}{\delta}\sum_{j=2}^{\infty}\delta^{j}w^{j}d^{j}\Big{|}\leq\Big{|}f(x)\frac{(d~{}w)^{2}~{}\delta}{1-\delta~{}d~{}w}\Big{|}.$
-Recall that in our estimates for $\alpha$ and $\delta$ we can apply Prop. 6.1 for $w\leq\frac{\theta}{\theta-1}$ and obtain
+In our estimates for $\alpha$ and $\delta$ we apply Prop. 10.1 for $w\leq\frac{\theta}{\theta-1}$ and obtain
 Corollary 10.2.
-Under the assumptions of Prop. 6.1, let $w\leq\frac{\theta}{\theta-1}$. Then
+Under the assumptions of Prop. 10.1, let $w\leq\frac{\theta}{\theta-1}$. Then
 $\alpha<1/8~{}{\rm for}~{}|\delta|\leq\frac{(\theta-1)^{2}}{10d^{2}\theta^{2}}.$
@@
 Proposition 10.3.
-Suppose that a black box oracle computes for us the values $t(y)$ within a relative error bounds $\nabla(y)$ for $y=x$, $y=x-\delta$, and $\delta$ of Cor.
-6.2.
+Suppose that a black box oracle has computed the values $t(y)$ within a relative error bounds $\nabla(y)$ for $y=x$, $y=x-\delta$, and $\delta$ of Cor.
+10.2.
 Then we can ensure that $|\beta|<\frac{1}{8}$ by choosing a proper $\nabla(y)$ of order $1/d^{O(1)}$, represented with a precision of $\log(\frac{1}{\nabla(y)})=O(\log(d))$ bits.
@@
 Hence $|\frac{t(x-\delta)}{t(x)}|\leq 1+|(1+\alpha^{\prime})\delta|$ for $\alpha^{\prime}=\frac{t^{\prime}(u)-t(x)}{t(x)}$.
-By extending Cor. 6.2 obtain $|\alpha^{\prime}|\leq\frac{1}{8}$.
+By extending Cor. 10.2 obtain $|\alpha^{\prime}|\leq\frac{1}{8}$.
 Hence $|\frac{t(x-\delta)}{t(x)}|\leq w:=1+\frac{9|\delta|}{8}$, while $|\beta\delta|\leq w\frac{|\nabla(x)|+|\nabla(x-\delta)|}{1-|\nabla(x)|}$.
@@
 ∎
-By combining Cors. 6.1 and 6.2 and Prop. 6.3 obtain
+By combining Cors. 10.1 and 10.2 and Prop. 10.3 obtain
 Corollary 10.3.
@@
 11 Approximation of matrix eigenvalues
-Given a $d\times d$ matrix $M$ that has characteristic polynomial $p(x)=\det(xI-M)$, our root-finders can approximate all $m$ its zeros (that is, the eigenvalues of $M$) lying in a fixed disc $D$ on the complex plane isolated from external eigenvalues.
-Recall Gershgorin’s bound for any eigenvalue $\lambda$ of a matrix $M=(a_{i,j})_{i,j=1}^{d}$ [134, Thm. 1.3.2]: $|\lambda-a_{i,i}|\leq\sum_{j\neq i}|a_{i,j}|$ for some $i$, $1\leq i\leq d$.
+Given a $d\times d$ matrix $M$ that has characteristic polynomial $p(x)=\det(xI-M)$, our root-finders approximate all $m$ its zeros (that is, the eigenvalues of $M$) lying in a fixed disc $D$ on the complex plane isolated from external eigenvalues.
+Recall Gershgorin’s bound for any eigenvalue $\lambda$ of a matrix $M=(a_{i,j})_{i,j=1}^{d}$ [138, Thm. 1.3.2]: $|\lambda-a_{i,i}|\leq\sum_{j\neq i}|a_{i,j}|$ for some $i$, $1\leq i\leq d$.
 Hence, clearly, $|\lambda|\leq||M||_{F}$.
@@
 Theorem 11.1.
-One can evaluate the determinant of a $d\times d$ integer matrix $A$ by using an expected number of $O(d^{\omega}\log^{2}(d)\log(|A||_{F}))$ bit operations.
+One can evaluate the determinant of a $d\times d$ integer matrix $A$ by using a Las Vegas expected number of $O(d^{\omega}\log^{2}(d)\log(|A||_{F}))$ bit operations.
 This implies a record bit operation complexity bound $\tilde{O}(m^{2}\tilde{b}^{2}d^{\omega})$ (cf.
-(1.7)) for approximation within $\epsilon$ of all $m$ eigenvalues of $M$ lying in the disc $D$ provided that $\tilde{b}=\log_{2}(||M||_{F}/\epsilon)$.
+(1.6)) for Las Vegas approximation within $\epsilon$ of all $m$ eigenvalues of $M$ lying in the disc $D$ provided that $\tilde{b}=\log_{2}(||M||_{F}/\epsilon)$.
 Remark 11.1.
-Our root-finders are essentially reduced to evaluation of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ at sufficiently many points $x$.
-Hence we can rely on the following expressions for approximation of the eigenvalues of a matrix or a matrix polynomial:
+Our root-finders are essentially reduced to the evaluation of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ at sufficiently many points $x$, and we can apply the following expressions:
 $\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}((xI-M)^{-1})}$
@@
 $\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}(M^{-1}(x)M^{\prime}(x))}$
-where $M(x)$ is a matrix polynomial and $p(x)=\det(M(x))$ [19, Eqn.
+where $M(x)$ is a matrix polynomial and $p(x)=\det(M(x))$ [20, Eqn.
 (5)].
-[^25] These expressions have been obtained based on Jacobi’s formula ${\rm d}(\ln(\det(A)))={\rm trace}(A^{-1}{\rm d}A),$ which links differentials of a nonsingular matrix $A$ and of its determinant.
+[^17] These expressions have been obtained based on Jacobi’s formula ${\rm d}(\ln(\det(A)))={\rm trace}(A^{-1}{\rm d}A),$ which links differentials of a nonsingular matrix $A$ and its determinant.
 Based on the above expressions one can reduce approximation of eigenvalues of a matrix or a matrix polynomial to its inversion.
-This computation tends to be more numerically stable than computing determinants and is significantly accelerated in the highly important case where where an input matrix or matrix polynomial can be inverted fast, e.g., is data sparse.
-8 Bit operation cost of the classical subdivision root-finding with novel e/i tests based on approximation of a Cauchy integral
+This computation tends to be more numerically stable than computing determinants and is significantly accelerated in the highly important case where an input matrix or matrix polynomial can be inverted fast, e.g., is data sparse.
+12 Bit operation cost of the classical subdivision root-finding with e/i tests based on approximation of a Cauchy integral
 12.1 Our goal and a basic theorem
-Next we estimate the bit operation complexity of our root-finders by combining them with fast multipoint evaluation of a polynomial by means of the algorithm of Fiduccia [40] and Moenck and Borodin [81] (see [18, Sec.
+Next we estimate the bit operation complexity of our root-finders by combining them with fast multipoint evaluation of a polynomial by means of the algorithm of Fiduccia [42] and Moenck and Borodin [84] (see [19, Sec.
 4.5] for historical account), non-trivially extended in [81] to support the following theorem:
@@
 Proof.
-For $q=d$ this is [78, Thm.
+For $q=d$ this is [81, Thm.
 2], based on using novel data structure for hyperbolic approximation of $p$, proposed and analyzed by Moroz in [81].
@@
 $\mathbb{B}_{\rm Kirr}=\tilde{O}((d+q)(b+q)),$ (12.2)
-due to [65, Thm.
+due to [68, Thm.
 3.9, Alg.
@@
 (12.1) and (12.2) on the overall bit operation complexity by computing with multiple precision for larger values $b$.
-Moroz obtained [78, Thm.
+Moroz obtained [81, Thm.
 2] by applying bound (12.2) to his lower degree polynomials approximating $p(x)$.
@@
 12.2 Reduction of our root-finders to multipoint polynomial approximation
-Given a complex $c$, a pair of positive $b$ and $\rho$, and $d+1$ coefficients of a polynomial $p=p(x)$ of (1.1) such that the disc $D(c,\rho)$ is isolated and contains precisely $m$ zeros of $p$, we approximate all these zeros within $R/2^{b}$ for $R=|c|+\rho$ by applying subdivision root-finding iterations with deterministic or Las Vegas randomized e/i tests of Sec.
-4.
+Given a complex $c$, a pair of positive $b$ and $\rho$, and $d+1$ coefficients of a polynomial $p=p(x)$ of (1.1) such that the disc $D(c,\rho)$ is isolated and contains precisely $m$ roots, we approximate all these roots within $R/2^{b}$ for $R=|c|+\rho$ by applying subdivision root-finding iterations with e/i tests of Sec.
+8.
 Then every subdivision step is reduced to application of e/i tests to $\bar{m}=O(m)$ discs $D(c_{\lambda},\rho_{\lambda})$ for $\rho_{\lambda}\geq R/2^{b}$ and all $\lambda$.
-Such a test amounts to approximation of $p(x)$ at $O(q\log(d))$ equally spaced points $x$ on each of $v$ circles sharing the center $c_{\lambda}$ and having radii in the range $[\rho_{\lambda}/\sigma,\rho_{\lambda}\sigma]$ for $1<\sigma<\sqrt{2}$.
-Here $v=2m+1$ and $q=m^{2}$ or $v=O(1)$ and $q=m$ in our e/i tests of Sec.
-4, while $q=1$ under the random root model.
+Such a test, based on our e/i test of Sec.
+8, amounts to approximation of $p(x)$ at $O(m\log(d))$ equally spaced points $x$ on each of $v$ circles sharing the center $c_{\lambda}$ and having radii in the range $[\rho_{\lambda}/\sigma,\rho_{\lambda}\sigma]$ for $1<\sigma<\sqrt{2}$.
 We only use Thm.
@@
 [^19] Scaling $p$ and $x$ above is done once and for all e/i tests; to compensate for this we only need to change our bound on the output precision $b$ by $\lceil|\log_{2}R|\rceil$ bits.
-Due to Cor. 6.3 it is sufficient to ensure the bound $1/(|t(x)|2^{\bar{b}})=1/d^{O(1)}$; we achieve this by choosing any $\bar{b}\geq\log(\frac{1}{|t(x)|})+O(\log(d))$.
+Due to Cor. 10.3 it is sufficient to ensure the bound $1/(|t(x)|2^{\bar{b}})=1/d^{O(1)}$; we achieve this by choosing any $\bar{b}\geq\log(\frac{1}{|t(x)|})+O(\log(d))$.
 In our e/i tests we only need to approximate $t(x)$ at the points $x$ lying on $\theta_{\lambda}$-isolated circles where $\frac{1}{\theta_{\lambda}-1}=O(m)$.
@@
 Now substitute $||t(x)||_{1}\leq 2^{b+d}$ and $\log(\frac{||t(x)||_{1}}{|t(x)|})=O(d\log(m))$ (cf.
-[130, Eqn.(9.4)]) [^28] This bound is readily deduced from the following non-trivial result [130, Thm.
+[134, Eqn.(9.4)]) [^20] This bound is readily deduced from the following non-trivial result [134, Thm.
 4.2]: Let $t(x)=f\prod_{j=1}^{m}(x-z_{j})\prod_{j=m+1}^{d}(\frac{x}{z_{j}}-1)$, $||t(x)||_{1}=1$, $|z_{j}|<1$ for $j\leq m$, and $|z_{j}|>1$ for $j>m$.
@@
 Remark 12.1.
-For $m=d$ the above bound on $\log(\frac{||t(x)||_{1}}{|t(x)|})$ of [130, Eqn.(9.4)]) is sharp up to a constant factor and is reached for the polynomial $t(x)=(x+1-\frac{1}{m})^{d}/(2-\frac{1}{m})^{d}$, $\rho_{\lambda}=1$, and $x=-1$.
-For $m<d$ [130, Thm.
+For $m=d$ the above bound on $\log(\frac{||t(x)||_{1}}{|t(x)|})$ of [134, Eqn.(9.4)]) is sharp up to a constant factor and is reached for the polynomial $t(x)=(x+1-\frac{1}{m})^{d}/(2-\frac{1}{m})^{d}$, $\rho_{\lambda}=1$, and $x=-1$.
+For $m<d$ [134, Thm.
 4.5] implies a little stronger bounds, namely, $\log(\frac{||t(x)||_{1}}{|t(x)|})=O(d+m\log(m))$ and hence $\bar{b}=O(m\log(m)+d+b)$ because all discs $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ lie in $\theta$-isolated unit disc $D(0,1)$ for $\theta-1$ exceeding a positive constant.
@@
 for the overall bit operation cost of our $O(b)$ subdivision steps, which includes the time-complexity of the evaluation oracle.
-Based on [65, Alg.
+Based on [68, Alg.
 5.3] and Eqn.
-(8.2) obtain the following (a little larger but still reasonably small) upper bound on the bit operation cost of our subdivision root-finders,
+(12.2) obtain a little larger upper bound on the bit operation cost of our subdivision root-finders:
 $\mathbb{B}_{\rm roots}^{\prime}=\tilde{O}((m^{2}+d+b)(m^{2}+d)b).$ (12.4)
@@
 Transition from $p(x)$ to $t_{c,\rho}(x)$, however, involves scaling of the variable $x$ by factors ranging from $R/2^{b}$ to $R=\max_{j-1}^{d}|z_{j}|$ (besides root-squaring and shift of the variable).
-Hence for every $i$, $i=0,1,\dots,d$, the coefficient $t_{i}$ must be computed within $1/2^{b-i}$ to approximate the zeros of $p$ within $1/2^{b}$ (see Observation B.2).
-This involves at least $bg$ bits for $g=1,\dots,d$, that is, at least $0.5(d-1)db$ bits and $0.25(d-1)db$ bit-operations overall for a single e/i test.
-[…]
+Hence for every $i$, $i=0,1,\dots,d$, the coefficient $t_{i}$ must be computed within $1/2^{b-i}$ to approximate the zeros of $p$ within $1/2^{b}$ (see Observation 24.2).
+This involves at least $bg$ bits for $g=1,\dots,d$, that is, at least $0.5(d+1)db$ bits and $0.25(d+1)db$ bit-operations overall for a single e/i test.
+A single subdivision iteration for Problem 1m can involve up to $4m$ e/i tests and hence up to at least $0.25(d+1)dbm$ bit-operations.
+This is by a factor of $m$ greater than for the root-finder of [96, 103] overall.
 The bit-count above and the bit operation cost bounds for the root-finders of [96, 103, 24] decrease by a factor of $d$ where the minimal pairwise distance between the zeros of $p$ exceeds a positive constant, but the ratio of the bounds of [24] and [96, 103] remains at least $m$, where $m$ can be as large as $d$.
 Similar argument shows that even representation of the leading coefficients $t_{d}$ of $t(x)$ alone after scaling for a singe e/i test involves odder of $bd$ bits, which is amplified to $d^{2}b$ for $t_{d}^{d}$ in every e/i test (see the end of Sec.
-B.4) and up to $d^{2}bm$ per a subdivision iteration with $m$ e/i tests.
+24.4) and up to $d^{2}bm$ per a subdivision iteration with $m$ e/i tests.
 It may be interesting that a shift by $c$ such that $|c|\leq 2$ is much less costly than scaling:
 Theorem 12.2.
-Bit operation cost of the shift of the variable, [132, Lemma 2.3] (cf.
-[65, Lemma 3.6]).
-Given a positive $b$, the coefficients of a polynomial $p(x):=\sum_{i=0}^{d}p_{i}x^{i}$ such that $||p(x)||_{1}\leq 1$, and a complex $c$ such that $|c|\leq 2$, one can approximate within $1/2^{b}$ the coefficients $t_{0},t_{1},\dots,t_{d}$ of the polynomial $t(x):=\sum_{i=0}^{d}t_{i}x^{i}:=p(x-c)$ by performing $O(\mu((b+d)d))$ bit operations for $\mu(s)$ of (1.8).
+Bit operation cost of the shift of the variable, [136, Lemma 2.3] (cf.
+[68, Lemma 3.6]).
+Given a positive $b$, the coefficients of a polynomial $p(x):=\sum_{i=0}^{d}p_{i}x^{i}$ such that $||p(x)||_{1}\leq 1$, and a complex $c$ such that $|c|\leq 2$, one can approximate within $1/2^{b}$ the coefficients $t_{0},t_{1},\dots,t_{d}$ of the polynomial $t(x):=\sum_{i=0}^{d}t_{i}x^{i}:=p(x-c)$ by performing $O(\mu((b+d)d))$ bit operations for
+$\mu(s):=O(s\log(s)=\tilde{O}(s)$ (12.5)
+bounding the number of bit operations for integer multiplication modulo $2^{s}$ [53] (cf. [142, 69, 134, 135] ).
 We can relax the restriction $|c|\leq 2$ above by means of scaling the variable $x\mapsto xc$, but as we showed above, this would greatly increase the bound on the overall number of bit operations involved.
-9 Real and near real root-finding
-9.1 Background, outline, and organization of our presentation
-The algorithm of [67] specializes subdivision of [23] to real root-finding for univariate polynomial represented in monomial basis.
-Currently this is user’s choice algorithm, although we have accelerated it in [59] by narrowing the region for the search of the real roots, based on application of Newton’s polygon – by following [130] and [88, Sec.4] (cf.
-[89, Sec.
-5]).
-Unlike the customary real root-finders, next we approximate both real and near real roots, lying on and near a fixed real line segment.
-This covers approximation of real zeros of a polynomial given by its values or coefficients that are only available numerically with rounding errors.
-Taylor’s shifts with scaling (1.4) readily reduce root-finding and root-counting for $p(x)$ in an isolated line segment to the case of the real line segment $[-1,1]$.
-Zhukovsky’s map of the next subsection further reduces the task to root-finding and root-counting for an auxiliary polynomial $s(z)$ on isolated unit circle $C(0,1)$ (see Remark 9.1).
-If the polynomial $s(z)$ has the same root set on the unit circle $C(0,1)$ and in the annulus $A(0,\theta_{-},\theta_{+})$, for $0<\theta_{-}<1<\theta_{+}$, and if its the boundary circles are isolated, then we obtain an efficient root-counter, by applying our Alg.
-3.1 to the isolated discs $D(0,\theta_{-})$ and $D(0,\theta_{+})$.
-We can readily extend our subdivision algorithms to the the case of an input annulus rather than a disc.
-For a fixed error bound $\epsilon>0$, its complexity is dominated by the cost of the evaluation of NIR$(x)$ at $O(w\log(1/\epsilon))$ points for $w$ denoting the number of real zeros of $p(x)$ in a fixed range; realistically this number tends to be much smaller than $d$ (cf.
-[63, 37]).
-We organize our presentation in this section as follows.
-In Sec.
-9.2 we specify the two-way map, based on Zhukovsky’s function, from the real segment $[-1,1]$ to the unit circle $C(0,1)$ and backward.
-The map $[-1,1]\Longleftrightarrow C(0,1)$ can be extended to the neighborhoods of the segment and the circle (cf.
-Remark 9.1).
-We can readily compute the number $w$ of roots on the isolated unit circle $C(0,1)$ and can deflate a factor $f$ of $p$ with root set made up of the $w$ zeros of $p$ in $C(0,1)$ by extending deflation algorithms of [130] from a disc to an annulus.
-In Secs.
-9.3 and 9.4 we extend these computations to the segment $[-1,1]$ and its neighborhood and consequently to approximation of both real and near real roots.
-9.2 Zhukovsky’s function
-For back and forth transition between the segment and the circle we can apply the two-to-one Zhukovsky’s function $z=J(x)$ and its one-to-two inverse, for complex variables $x$ and $z$.
-It maps the circle $C(0,1)$ to the segment $[-1,1]$ and vice versa:
-$x=J(z):=\frac{1}{2}\Big{(}z+\frac{1}{z}\Big{)};~{}z=J^{-1}(x):=x\pm\sqrt{x^{2}-1}.$ (9.1)
-9.3 A real root-finder
-Algorithm 9.1.
-Root-finding on a line segment.
-Input:
-A polynomial $p=p(x)$ of (1.1).
-Output:
-The number $w$ of its zeros on the unit segment $[-1,1]$ and approximations to all these zeros.
-Computations:
-  1. 1.
-    Compute the values $v_{h}=p(\Re(\zeta_{2d}^{h}))$ of the polynomial $p$ at the Chebyshev points $\Re(\zeta_{2d}^{h})=\cos(\frac{\pi h}{2d})$ for $h=0,1,\dots,2d-1$ and $\zeta_{2d}$ of (1.10).
-  2. 2.
-    Interpolate to the polynomial $s(z)$ of degree $2d$ such that
-    $s(z)=z^{d}p(x)~{}{\rm for}~{}x=\frac{1}{2}(z+z^{-1})$ (9.2)
-    from its values
-    $s(\zeta_{2d}^{h})=(-1)^{h}v_{h}~{}{\rm for}~{}h=0,1,\dots,2d-1$
-    by means of Inverse DFT. [Recall that $\zeta_{2d}^{dh}=(-1)^{h}$.]
-  3. 3.
-    Approximate a factor (e.g., the monic factor) $g=g(z)$ of $p$ whose root set is made up of all roots of the polynomial $s(z)$ that lie on the unit circle $C(0,1)$.
-[The zeros of $s(z)$ appear in pairs: $s(z)=0$ if $s(z)=0$ by virtue of (9.2).
-Furthermore, if $p(1)=0$ and/or $p(-1)=0$, then 1 and/or $-1$ are also the zeros of $g(z)$ with double multiplicity.] Approximate a factor (e.g., the monic factor) $g=g(z)$ of $p$ whose root set is made up of all zeros of the polynomial $s(z)$ that lie on the unit circle $C(0,1)$.
-At this stage extend deflation algorithms of [130] to the the case of an input annulus about the unit circle $C(0,1)$.
-Output $w=0.5\deg(g)$.
-  4. 4.
-    By applying MPSolve, subdivision, or another root-finder approximate all $2w$ zeros of the polynomial $g(z)$.
-Let $y_{1},\dots,y_{w}$ denote the first $w$ of them in the order of increasing their arguments from 0.
-  5. 5.
-    Compute and output the values $z_{j}=\frac{1}{2}(y_{j}+\frac{1}{y_{j}})$ for $j=1,\dots,w$; these are the $w$ zeros of $p$ lying in the segment $[-1,1]$.
-Correctness of this algorithm follows from (9.1) and (9.2).
-Its Stage 2 is DFT.
-Its Stage 1 of the evaluation at Chebyshev’s points can be performed by means of the algorithm of [43] or Discrete Cosine Transform (cf.
-[98, Sec.
-3.11 and the notes to it in Sec.
-3.12]).
-In both cases the computation is simplified if $d$ is a power of 2; we ensure this by replacing $p$ with $x^{u}p$ for the minimal non-negative integer $u$ such that $d+u$ is a power of 2.
-The complexity at stage 3 is dominated by the cost of evaluation of $p(x)$ at $O(w\log(1/\epsilon))$ points for $\epsilon$ bounding the output errors.
-The converse transition from the unit circle $C(0,1)$ to the unit segment $[-1,1]$ enables us to simplify stages 4 and 5 of Alg.
-9.1 by moving from $g(z)$ to a polynomial of degree $w$ whose all zeros lie in the segment $[-1,1]$.
-Then again we achieve this by means of evaluation and interpolation:
-Algorithm 9.2.
-Transition to the unit segment.
-  1. 1.
-    Compute the values $u_{h}=g(\zeta_{2K}^{h})$ of the polynomial $g(z)$ at the $2K$-th roots of unity $\zeta_{2K}^{h}$ for $h=0,1,\dots,K-1$ and $K>w$.
-  2. 2.
-    Interpolate to the polynomial $f(x)$ of degree at most $w$ from its values $f(\Re(\zeta_{2K}^{h}))=(-1)^{h}u_{h}$ at the Chebyshev points $\Re(\zeta_{2K}^{h})=\cos(\frac{\pi h}{2K})$, for $h=0,1,\dots,K-1$.
-[Recall that $\zeta_{2K}^{Kh}=(-1)^{h}$.]
-  3. 3.
-    Approximate the $w$ zeros of the polynomial $f=f(x)$ by applying MPSolve, real subdivision, e.g., those of [67], or another real polynomial root-finder.
-We propose to perform steps 1 and 2 above by means of forward DFT and inverse Cosine Transforms, which we would apply to the polynomial $z^{v}g(z)$, replacing $g(x)$, where $v$ is the minimal non-negative integer such that $w+v$ is a power 2.
-Remark 9.1.
-Represent complex numbers as $z:=u+v\bf i$. Then Zhukovsky’s map transforms a circle $C(0,\rho)$ for $\rho\neq 1$ into the ellipse $E(0,\rho)$ whose points $(u,v)$ satisfy the following equation,
-$\frac{u^{2}}{s^{2}}+\frac{v^{2}}{t^{2}}=1~{}{\rm for}~{}s=\frac{1}{2}\Big{(}\rho+\frac{1}{\rho}\Big{)},~{}~{}t=\frac{1}{2}\Big{(}\rho-\frac{1}{\rho}\Big{)}.$
-Consequently it transforms the annulus $A(0,\frac{1}{\theta},\theta)$ into the domain bounded by the ellipses $E(0,\frac{1}{\theta})$ and $E(0,\theta)$, and so the circle $C(0,1)$ is $\theta$-isolated if and only if no roots of $p$ lie in the latter domain.
-9.4 Real root-counting
-We can simplify Alg. 9.1 and use only evaluations if we restrict our task to counting the roots that lie on or near a line segment.
-Here is a high level description of this algorithm, where we do not specify the parameters $\theta$ and $q$ and assume that the input includes the polynomial $s(z)$.
-Algorithm 9.3.
-Root-counting on a line segment.
-A real $\theta>1$ and the polynomial $s(z)$ of equation (9.2) such that the unit circle $C(0,1)$ is t $\eta$-isolated for $\eta\geq\theta^{2}$, that is, the annulus $A(0,\frac{1}{\theta^{2}},\theta^{2})$ contains no zeros of $s(z)$ except possibly some on the unit circle $C(0,1)$.
-The number $w$ of the roots in the unit segment $[-1,1]$ or FAILURE.
-  1. 1.
-    Compute the polynomial $s_{-}(z)=s(\frac{z}{\sqrt{\theta}})$.
-  2. 2.
-    Choose a sufficiently large $q$ and compute Cauchy sums $\tilde{s}_{0,q}$ of the roots of $s_{-}(z)$ in the unit disc $D(0,1)$.
-  3. 3.
-    If the value $s_{0,q}$ is sufficiently close to an integer $j$, then output $w:=d-j$ and stop. Otherwise output FAILURE.
-Let us prove correctness of the algorithm.
-By assumption the annulus $A(0,\frac{1}{\theta},1)$ contains no roots, except possibly some on the circle $C(0,1)$, and so $j$ is the number of the zeros of $s(z)$ in the disc $D(0,\frac{1}{\theta})$ by virtue of Observation 3.1.
-Clearly, the polynomial $s(z)$ of degree $2d$ has as many zeros lying in and outside the unit disc, and so precisely $2d-2j$ zeros of $s(z)$ lie on the unit circle $C(0,1)$.
-This is twice as many zeros as $p(x)$ has on the unit segment $[-1,1]$.
 PART III: Classical subdivision root-finder with novel e/i tests based on root-lifting
@@
 In Secs.
-11 and 12 we cover our e/i tests and their extensions to $m$-tests, respectively.
+14 and 15 we cover our e/i tests and their extensions to $m$-tests, respectively.
 In Secs.
-13 and 15 we estimate computational precision of our root-finders and their bit operation cost, respectively.
+16 and 18 we estimate computational precision of our root-finders and their bit operation cost, respectively.
 In short Sec.
@@
 We recall or redefine some concepts from the Introduction and Part II.
-  • ${\rm NR}_{t}(x):=\frac{t(x)}{t^{\prime}(x)}$ is Newton’s ratio, ${\rm NIR}_{t}(x):=\frac{t^{\prime}(x)}{t(x)}$ is Newton’s inverse ratio for a polynomial $t(x)$, ${\rm NIR}_{t}(x){\rm NR}_{t}(x)=1$.
-  • We write NR$(x):={\rm NR}_{p}(x)$ and NIR$(x):={\rm NIR}_{p}(x)$.
-  • We approximate NIR$(x)$ at the cost of evaluation of $p(x)$ at two points by using the first order divided differences (cf. (1.6)).
+  • ${\rm NR}_{t}(x):=\frac{t(x)}{t^{\prime}(x)}$ is Newton’s ratio, ${\rm NIR}_{t}(x):=\frac{t^{\prime}(x)}{t(x)}$ is Newton’s inverse ratio for a polynomial $t(x)$.
+We write NR$(x):={\rm NR}_{p}(x)$ and NIR$(x):={\rm NIR}_{p}(x)$ (cf.
+(3.2)).
   • Define squares, discs, circles (circumferences), and annuli on the complex plane: $S(c,\rho):=\{x:~{}|\Re(c-x)|\leq\rho$, $|\Im(c-x)|\leq\rho\}$$D(c,\rho):=\{x:~{}|x-c|\leq\rho\},$ $C(c,\rho):=\{x:~{}|x-c|=\rho\}$,  $A(c,\rho,\rho^{\prime}):=\{x:~{}\rho\leq|x-c|\leq\rho^{\prime}\}.$
-  • $\Delta(\mathbb{S})$, $CH(\mathbb{S})$, $X(\mathbb{S})$ and #$\#(\mathbb{S})$ denote the diameter, root set, convex hull, and index (root set’s cardinality) of a set, or region $\mathbb{S}$ on the complex plane, respectively.
+  • $\Delta(\mathbb{S})$, $CH(\mathbb{S})$, $X(\mathbb{S})$ and #$\#(\mathbb{S})$ denote the diameter, root set, convex hull, and index (root set’s cardinality) of a set $\mathbb{S}$ on the complex plane, respectively.
   • A disc $D(c,\rho)$, a circle $C(c,\rho)$, or a square $S(c,\rho)$ is $\theta$-isolated for $\theta\geq 1$ if $X(D(c,\rho))=X(D(c,\theta\rho))$, $X(C(c,\rho))=X(A(c,\rho/\theta,\rho\theta)$ or $X(S(c,\rho))=X(S(c,\theta\rho)),$ respectively.
     The isolation of a disc $D(c,\rho)$, a circle $C(c,\rho)$, or a square $S(c,\rho)$ is the largest upper bound on such a value $\theta$ (see Fig.
-4) and is denoted $i(D(c,\rho))$, $i(C(c,\rho))$, or $i(S(c,\rho))$, respectively.
+5) and is denoted $i(D(c,\rho))$, $i(C(c,\rho))$, or $i(S(c,\rho))$, respectively.
     We say isolated for “$\theta$-isolated" if $\theta-1$ exceeds a positive constant.
@@
   • $|u|:=||{\bf u}||_{1}:=\sum_{i=0}^{q-1}|u_{i}|$, $||{\bf u}||:=||{\bf u}||_{2}:=(\sum_{i=0}^{q-1}|u_{i}|^{2})^{\frac{1}{2}}$, and $|u|_{\rm max}:=||{\bf u}||_{\infty}:=\max_{i=0}^{q-1}|u_{i}|$ denote the 1-, 2-, and $\infty$-norm, respectively, of a polynomial $u(x)=\sum_{i=0}^{q-1}u_{i}x^{i}$ and its coefficient vector [^22] Here and hereafter ${\bf v}$ denotes a column vector, while ${\bf v}^{T}$ denotes its transpose.
-${\bf u}=(u_{i})_{i=0}^{d-1}$, and then (see [46, Eqs.
+${\bf u}=(u_{i})_{i=0}^{d-1}$, and then (see [48, Eqs.
 (2.2.5)–(2.2.7)])
     $||{\bf u}||_{\infty}\leq||{\bf u}||_{2}\leq||{\bf u}||_{1}\leq{\sqrt{q}}~{}||{\bf u}||_{2}\leq q||{\bf u}||_{\infty}.$ (13.1)
-  • $r_{1}(c,t)=|y_{1}-c|,\dots,r_{d}(c,t)=|y_{d}-c|$ in non-increasing order are the $d$ root radii, that is, the distances from a complex center $c$ to the roots $y_{1},\dots,y_{d}$ of a $d$th degree polynomial $t(x)$.
+  • $r_{1}(c,t)=|y_{1}-c|,\dots,r_{d}(c,t)=|y_{d}-c|$ in non-increasing order are the $d$ root radii, that is, the distances from a complex point $c$ to the roots $y_{1},\dots,y_{d}$ of a $d$th degree polynomial $t(x)$.
 Write $r^{h}_{j}(c,t):=(r_{j}(c,t))^{h}$, $r_{j}(c):=r_{j}(c,p)$.
-  • Generalize e/i tests as follows:
-    Definition 10.1.
-[…]
-Observation 10.1.
+  • Generalize Def.
+2.1 of $\sigma$-soft e/i tests as follows: For $p(x)$ of (1.1), $\sigma>1$, and integers $\ell$ and $m$ such that $1\leq\ell\leq m\leq d$ and #$\#(D(0,1))\leq m$, a $\sigma$-soft $\ell$-test, or just $\ell$-test for short (in particular e/i test turns into 1-test), either outputs 1 and stops if it detects that $r_{d-\ell+1}\leq\sigma$, that is, #$\#(D(0,\sigma))\geq\ell$, or outputs 0 and stops if it detects that $r_{d-\ell+1}>1$, that is, [^23] Bounds $r_{d-\ell+1}\leq\sigma$ and $r_{d-\ell+1}>1$ can hold simultaneously (cf.
+Fig.
+1b), but as soon as an $\ell$-test verifies any of them, it stops without checking if the other bound also holds.
+#$\#(D(0,1))<\ell$.
+An $\ell$-test applied to the polynomial $t(y)$ of (1.3) is said to be an $\ell$-test for the disc $D(c,\rho)$ and/or the circle $C(c,\rho)$ and also to be an $\ell$-testc,ρ.
+[^24] We will only use $\ell$-tests for $\ell=1$ (which are e/i tests) and $\ell=m$.
+Observation 13.1.
 Taylor’s shift, or translation, of the variable $x\mapsto y:=x-c$ and its scaling combined (see (1.3)) map the zeros $z_{j}$ of $p(x)$ into the zeros $y_{j}=\frac{z_{j}-c}{\rho}$ of $t(y)$, for $j=1,\dots,d$, and preserve the index and the isolation of the disc $D(c,\rho)$.
@@
 ${\rm(iii)}~{}p^{\prime}_{\rm rev}(x)=dx^{d-1}p\Big{(}\frac{1}{x}\Big{)}-x^{d-2}p^{\prime}\Big{(}\frac{1}{x}\Big{)}~{}{\rm for}~{}x\neq 0,~{}p_{d}=p_{\rm rev}(0),~{}p_{d-1}=p^{\prime}_{\rm rev}(0).$
-${\rm(iv)}~{}~{}~{}~{}\frac{p_{\rm rev}(x)}{p^{\prime}_{\rm rev}(x)}=\frac{1}{x}-\frac{{\rm NIR}(x)}{x^{2}},~{}x\neq 0,~{}{\rm NIR}(0)=\frac{p_{d-1}}{p_{d}}=-\sum_{j=1}^{d}z_{j}.~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
 […]
-We will not use the following theorem but include it for the sake of completeness.
-Theorem 10.1.
-For $p(x)$ of (1.1) and $R>1$ it holds that
-$\Big{|}\frac{p(R)}{R^{d}}-p_{d}\Big{|}\leq\frac{1}{R-1}\max_{i<d}|p_{i}|,$ (10.3)
-$\Big{|}\frac{Rp^{\prime}(R)-dp(R)}{R^{d-1}}-p_{d-1}\Big{|}\leq\frac{1}{R-1}\max_{0<i<d}|p_{i}|.$ (10.4)
-Proof.
-To prove bound (10.3) notice that
-$\frac{p(R)}{R^{d}}-p_{d}=\sum_{i=0}^{d-1}p_{i}R^{i-d}=\frac{1}{R}\sum_{i=0}^{d-1}p_{i}R^{i-d+1},$
-and so
-$\Big{|}\frac{p(R)}{R^{d}}-p_{d}\Big{|}\leq\frac{1}{R}\sum_{i=0}^{d-1}|p_{i}|R^{i-d+1}\leq\max_{i<d}|p_{i}|\sum_{i=0}^{d-1}|p_{i}|R^{i-d}\leq\frac{1}{R-1}\max_{i<d}|p_{i}|.$
-Likewise, to prove bound (10.4) notice that
-$Rp^{\prime}(R)-dp(R)-dR^{d-1}p_{d-1}=\sum_{i=1}^{d-1}ip_{i}R^{i-1}.$
-Hence
-$\frac{Rp^{\prime}(R)-dp(R)}{R^{d-1}}-p_{d-1}=\frac{1}{R}\sum_{i=1}^{d-1}\frac{i}{d}p_{i}R^{i-d+1},$
-and so
-$\Big{|}\frac{Rp^{\prime}(R)-dp(R)}{R^{d-1}}-p_{d-1}\Big{|}\leq\max_{0<i<d}|p_{i}|\sum_{i=1}^{d-1}R^{i-d}\leq\frac{1}{R-1}\max_{0<i<d}|p_{i}|.$
 13.3 Root-squaring and root-lifting
@@
 […]
-Observation 10.2.
+Observation 13.2.
 (i) Root-lifting preserves the index #$\#(D(0,1))$ of the unit disc.
 (ii) $h$-lifting (achieved with root-squaring for $h=2^{k}$ and a positive integer $k$) maps the values $i(D(0,1))$, $\sigma(D(0,1))$, $\eta(D(0,1))$, and $r_{j}(0,p)$ for all $j$, that is, the isolation of the unit disc $D(0,1)$, its softness, its rigidity and root radii, to their $h$-th powers.
-10.4 The distance to a closest zero of $p$ and fast inclusion tests
-Next we estimate the distance from the origin to the closest zero $z_{d}$ of $p$.
-The distance to its farthest zero $z_{1}$ is given by the same estimates applied to the reverse polynomial $p_{\rm rev}(x)$.
-For boundons $r_{d}$ and $r_{1}$ in terms of the coefficients of $p$, see [77, Sec. 1.10]; in particular see Fujihara’a bounds of [39]:
+13.4 The distance to a closest root and fast inclusion tests
+Next we estimate the distance from the origin to the closest root. The distance to the farthest root $z_{1}$ is given by the same estimates applied to the reverse polynomial $p_{\rm rev}(x)$.
+For bounds $r_{d}$ and $r_{1}$ in terms of the coefficients of $p$, see [80, Sec. 1.10]; in particular according to Fujihara’a bounds of [41]:
 $r_{1}\leq 2\max_{i=1}^{d}|p_{d-j}/p_{d}|^{1/j}$ unless $p_{d}=0$ and $r_{d}\geq 2\min_{i=1}^{d}|p_{j}/p_{0}|^{1/j}$ unless $p_{0}=0$.
-For a complex $c$ and positive integers $j$ and $h$, Eqn. (1.13) and Observation 10.2 combined imply the following relationships independent of $p_{d}$:
+For a complex $c$ and positive integers $j$ and $h$, Eqn. (3.1) and Observation 13.2 combined imply the following relationships independent of $p_{d}$:
 $r_{d}(c)\leq d~{}|{\rm NR}(c)|,$ (13.4)
@@
 Map (1.3) and Cor.
-1.3 together reduce an e/i test for any disc to the test for the unit disc $D=D(0,1)$, and we devise soft e/i tests for that disc assuming that #$\#(D(0,2))\leq m$ for a fixed $m$, $1\leq m\leq d$ (cf.
-claim (iv) of Cor.
-1.2).
+3.1 together reduce an e/i test for any disc to the test for the unit disc $D=D(0,1)$, and we devise soft e/i tests for that disc assuming that #$\#(D(0,2))\leq m$ for a fixed $m$, $1\leq m\leq d$.
 Our Alg.
@@
 14.3 are deterministic specifications of Alg.
-1.1, which is always correct when it claims inclusion, by virtue of Lemma 1.1.
+3.1, which is always correct when it claims inclusion, by virtue of Lemma 3.1.
 We prove that both of these specifications are also correct when they output exclusion.
 Namely, Algs.
-11.1 and 11.2 evaluate ${\rm NIR}(x)$ at $hq$ equally spaced points of the unit circle $C(0,1)$ for $h=O(\log(d))$ and correctly claim exclusion if at all these points $|{\rm NIR}(x)|\leq\frac{1}{2\sqrt{q}}$ for $q>d$ or if $|{\rm NIR}(x)|\leq\frac{1}{3\sqrt{q}}$ for $q>m$, respectively.
+14.1 and 14.2 evaluate ${\rm NIR}(x)$ at $hq$ equally spaced points of the unit circle $C(0,1)$ for $h=O(\log(d))$ and correctly claim exclusion if at all these points $|{\rm NIR}(x)|\leq\frac{1}{2\sqrt{q}}$ for $q>d$ or if $|{\rm NIR}(x)|\leq\frac{1}{3\sqrt{q}}$ for $q>m$, respectively.
 Our Example 14.1 shows that we cannot devise such deterministic algorithms for $q$ decreased below $m$, but in Sec.
@@
 Basic deterministic e/i test.
-INPUT: a polynomial $p(x)$ of a degree $d$ and $\sigma$, $1<\sigma<\sqrt{2}$.
+INPUT: a black box polynomial $p(x)$ of a degree $d$ given by an oracle for its evaluation and $\sigma$, $1<\sigma<\sqrt{2}$.
 INITIALIZATION: Fix two integers $q>d$ and $h:=\lceil\log_{\sigma}(1+2d\sqrt{q})\rceil$, such that $\sigma^{h}\leq 1+2d\sqrt{q}$.
@@
 Stop and certify $\sigma$-soft inclusion #$(D(0,\sigma))>0$ if $1+d/v_{g}\leq\sigma^{h}$ for an integer $g<q$. Otherwise stop and certify exclusion, that is, #$(D(0,1))=0$.
-Alg. 11.1 outputs $\sigma$-soft inclusion correctly by virtue of Lemma 1.1. Next we prove that it also correctly claims exclusion.
+Alg. 14.1 outputs $\sigma$-soft inclusion correctly by virtue of Lemma 3.1. Next we prove that it also correctly claims exclusion.
 Lemma 14.1.
@@
 14.1 outputs exclusion, then $1+\frac{d}{v_{g}}>\sigma^{h}$ for $g=0,1,\dots,q-1$, where $\sigma^{h}>1+2d\sqrt{q}$ according to initialization of Alg.
-11.1.
+14.1.
 Hence $\frac{1}{v_{g}}>2\sqrt{q}$, and so the assumptions of Prop.
 14.1 hold for $p_{h}(x)$ replacing $p(x)$.
-Therefore, the theorem implies that the disc $D(0,1)$ contains no zeros of $p_{h}(x)$, but then it contains no zeros of $p(x)$ as well (see Observation 10.2 (iii)), and so exclusion is correct.
+Therefore, the theorem implies that the disc $D(0,1)$ contains no zeros of $p_{h}(x)$, but then it contains no zeros of $p(x)$ as well (see Observation 13.2 (iii)), and so exclusion is correct.
 Readily verify the following estimates.
@@
 Otherwise certify exclusion, #$(D(0,1))=0$.
-Prop. 10.2, for $p_{h}$ replacing $p$, and the assumed bound $\frac{d-m}{\theta^{h}-1}\leq\frac{1}{6\sqrt{q}}$ together imply
+Prop. 13.2, for $p_{h}$ replacing $p$, and the assumed bound $\frac{d-m}{\theta^{h}-1}\leq\frac{1}{6\sqrt{q}}$ together imply
 $\frac{m}{r_{d}(\zeta^{g},p_{h})}\geq v_{g}-\frac{d-m}{\theta^{h}-1}\geq v_{g}-\frac{1}{6\sqrt{q}}~{}{\rm for~{}all}~{}g.$
-Alg. 11.2 claims inclusion if $v_{g}\geq\frac{1}{3\sqrt{q}}$ for some $g$, but then
+Alg. 14.2 claims inclusion if $v_{g}\geq\frac{1}{3\sqrt{q}}$ for some $g$, but then
 $\frac{m}{r_{d}(\zeta^{g},p_{h})}\geq\frac{1}{6\sqrt{q}},~{}r_{d}(\zeta^{g},p_{h})\leq 6m\sqrt{q},$
@@
 $r_{d}(0,p_{h})\leq 1+6m\sqrt{q}~{}\leq\sigma^{h}~{}{\rm for}~{}|\zeta|=1.$
-Now Observation 10.2 (iii) implies that Alg. 11.2 claims inclusion correctly.
+Now Observation 13.2 (iii) implies that Alg. 14.2 claims inclusion correctly.
 Next prove that it also claims exclusion correctly.
@@
 ∎
-Readily extend Prop. 11.2 to Alg. 11.2 by decreasing its bounds on $q$ and $h$ provided that #$\#(D(0,\theta))=m$ for $\theta\geq 1+6(d-m)\sqrt{q}$.
+Readily extend Prop. 14.2 to Alg. 14.2 by decreasing its bounds on $q$ and $h$ provided that #$\#(D(0,\theta))=m$ for $\theta\geq 1+6(d-m)\sqrt{q}$.
 Proposition 14.4.
@@
 14.4 A randomized e/i test
-13.4.1.
+14.4.1.
 Introductory comments.
 Would Algs.
-11.1 and 11.2 stay correct if we apply them at fewer evaluation points?
-They have consistently stayed correct for $q$ of order $\log^{2}(d)$ in extensive numerical tests of [57, 60] applied to $p(x)$ rather than to $p_{i}(x)$ for $i>0$ (that is, used no root-lifting).
+14.1 and 14.2 stay correct if we apply them at fewer evaluation points?
+They have consistently stayed correct for $q$ of order $\log^{2}(d)$ in extensive numerical tests of [60, 63] applied to $p(x)$ rather than to $p_{i}(x)$ for $i>0$.
 The tests reported exclusion where the values $|s_{h,q}|$ were reasonably small for $s_{h,q}:=\frac{1}{q}\sum_{g=0}^{q-1}\zeta^{(h+1)g}~{}\frac{p^{\prime}(\zeta^{g})}{p(\zeta^{g})},~{}{\rm for}~{}h=0,1,2$ (cf.
@@
 Then $p(x)=(1-u)(x^{d}-dx+u)$, $p^{\prime}(x)=(1-u)d(x^{d-1}-1)$ and NIR$(x)=0$ for $x$ being any $(d-1)$-st roots of unity.
-Extension of our e/i test to $q=d-1$ would imply that #$\#(D(0,1))=0$, while $p(y)=0$ for $u\approx 0$ and some $y\approx 0$.
+Then our e/i test for $q=d-1$ would imply that #$\#(D(0,1))=0$, while $p(y)=0$ for $u\approx 0$ and some $y\approx 0$.
 (ii) For $0<m<d$ notice that $p^{\prime}(x)=uf(x)+m(1-ux^{d-m})(x^{m-1}-1)$ where $f(x)=(d-m)x^{d-m-1}(x^{m}-mx+u)$ and so $uf(x)\approx 0$ while $1-ux^{d-m}\approx 1$ for $|x|\leq 1$ and $u\approx 0$.
@@
 Next we formally support this observation by introducing probability structure in the space of polynomials $p(x)$ and $t(x)=p(\frac{x-c}{\rho})$.
-Shifts destroy all customary probability distributions in the space of the coefficients of $p(x)$, but we prove correctness whp of e/i tests that evaluate $p(x)$ at $O(\log(d))$ under the Random Root Model.
+Shifts destroy all customary probability distributions in the space of the coefficients of $p(x)$, but we prove that our e/i tests that evaluate $p(x)$ at $O(\log(d)$ points are correct whp under the Random Root Model.
 […]
 Lemma 14.2.
-Fix any constant $\beta>1$ and a sufficiently large integer $h$ of order $\log(d)$. Then under the Random Root Model, Alg. 1.1(rrm) is correct with a probability at least $1-P$, where
+Fix any constant $\beta>1$ and a sufficiently large integer $h$ of order $\log(d)$. Then under the Random Root Model, Alg. 3.1(rrm) is correct with a probability at least $1-P$, where
 $P\leq\frac{\gamma d}{\pi R^{2}}~{}{\rm for}~{}\gamma:=\pi\max\Big{\{}\frac{\beta^{2}}{\sigma^{2}},O\Big{(}\frac{1}{\beta^{2h}}\Big{)}\Big{\}}.$ (14.4)
 Proof.
-By virtue of Lemma 1.1 Alg. 1.1(rrm) outputs $\sigma$-soft inclusion correctly and therefore can be incorrect only where it outputs exclusion for a disc $D(0,1)$ containing a zero of $p$.
-Let this zero be $z_{1}$, say.
+By virtue of Lemma 3.1 Alg. 3.1(rrm) outputs $\sigma$-soft inclusion correctly and, therefore, can be incorrect only where it outputs exclusion for a disc $D(0,1)$ containing a root.
+Let this root be $z_{1}$, say.
 Then fix the values of the random variables $z_{2},\dots,z_{d}$ and $x_{1}\in C(0,1)$ and let $\gamma$ denote the area (Lebesgue’s measure) of the set $\mathbb{S}$ of the values of the variable $z_{1}$ for which
 $v_{1}=|{\rm NIR}_{t_{h}}(x_{1})|<\frac{d}{\sigma^{h}-1}.$ (14.5)
-Then Alg. 1.1(rrm) is incorrect with a probability at most
+Then Alg. 3.1(rrm) is incorrect with a probability at most
 $P=\frac{\gamma d}{\pi R^{2}}$ (14.6)
@@
 $\lambda_{g+1}:=z^{g+1}-\bar{z}^{g+1},~{}\lambda_{g}:=z^{g}-\bar{z}^{g},~{}{\rm and~{}so}~{}z^{g+1}=z\bar{z}^{g}+z\lambda_{g},~{}|\lambda_{h}|\leq\nu_{h}~{}{\rm for}~{}h=g,g+1.$
-Therefore $\lambda_{g+1}+\bar{z}^{g+1}=\lambda_{g}z+\bar{z}^{g}z$, and so
+Therefore, $\lambda_{g+1}+\bar{z}^{g+1}=\lambda_{g}z+\bar{z}^{g}z$, and so
 $z-\bar{z}=\frac{\lambda_{g+1}-z\lambda_{g}}{\bar{z}^{g}},~{}|\bar{z}-z|\leq\frac{|\lambda_{g+1}|+|z|\cdot|\lambda_{g}|}{|\bar{z}|^{g}}<\frac{\nu_{g+1}+|z|\nu_{g}}{|\bar{z}|^{g}}=O\Big{(}\frac{d}{(\sigma|\bar{z}|)^{g}}\Big{)}.$
@@
 Then under the Random Root Model the probability $P$ of the output error for Alg.
-1.1(rrm) is less than a constant $\nabla<1$.
+3.1(rrm) is less than a constant $\nabla<1$.
 We can decrease the error probability below $\nabla^{v}$ by applying the algorithm $v$ times, in particular below $d^{a}$ for any fixed positive constant $a$ and for $v=O(\log(d))$.
-Cor.
-1.2 reduces classical subdivision root-finder to performing $O(mb)$ soft e/i tests, but in Part IV we decrease that bound to $w=O(m\log(b))$ tests.
+Based on Cor.
+2.2 we reduce classical subdivision root-finder to performing $O(mb)$ soft e/i tests, but in Part IV we decrease that bound to $w=O(m\log(b))$ tests.
 These tests magnify the error bounds by at most a factor of $w$.
@@
 Corollary 14.1.
-Under Random Root Model one can perform all $O(m\log(b)))$ soft Las Vegas randomized e/i tests involved in a subdivision root-finder which ensure output error bound $\epsilon=1/2^{b}$ whp provided that $\log(b)=d^{O(1)}$ and that every test is reduced to $v=O(\log(d))$ applications of Alg.
-1.1(rrm).
+Under Random Root Model one only need to apply a subdivision root-finder that involves $O(m\log(b)))$ soft Las Vegas randomized e/i tests to ensure the output error bound $\epsilon=1/2^{b}$ whp provided that $\log(b)=d^{O(1)}$ and that every test is reduced to $v=O(\log(d))$ applications of Alg.
+3.1(rrm).
 Remark 14.1.
-It is tempting to combine root-lifting or root-squaring with algorithms of [111], which for approximate Cauchy integral (1.9), to obtain alternative randomized e/i test supporting Cor.
-11.1, but for that we would have to raise disc isolation from order of $1+1/m$ to a constant, and this would require lifting of order $m$.
+It is tempting to combine root-lifting or root-squaring with algorithms of [115], which approximate Cauchy integral (3.4), to obtain alternative randomized e/i test supporting Cor.
+14.1, but for that we would have to raise disc isolation from order of $1+1/m$ to a constant, and this would require too costly lifting of order $m$.
 14.5 Counting the evaluation queries of our root-finders
-Combine Cors. 1.2 and 11.1 and Prop. 11.4 and obtain
+Combine Cors. 2.2 and 14.1 and Prop. 14.4 and obtain
 Corollary 14.2.
-We can reduce root-finding to deterministic evaluation of $p(x)$ at $O(m^{2}b\log(d))$ points as well as evaluation at $O(mb\log^{2}(d))$ points under the Random Root Model.
+We can reduce root-finding to deterministic evaluation of $p(x)$ at $O(m^{2}b\log(d))$ points as well as to evaluation at $O(mb\log^{2}(d))$ points under the Random Root Model.
 15 Reduction of $m$-test to e/i tests
-Our algorithm in Sec.
-12.1 reduces $m$-test to e/i tests motivated by application in Part IV.
-For these application our slower alternative reduction in Sec.
-12.2 is inferior, but we include it because it relaxes isolation assumption and works for any disc, which can be of independent interest.
+Our two algorithms of Secs.
+15.1 and 15.2 reduce $m$-test for a disc $D$ to bounded number of e/i tests.
+Namely, the algorithm of Sec.
+15.2 uses just a single e/i test where $m=d$ but up to order of $\log(d)$ e/i tests otherwise, while the algorithm of Sec.
+15.1 uses $O(1)$ e/i tests for any $m$.
+15.1 $m$-test by means of e/i tests for any disc
+Our alternative reduction of $m$-test for $1\leq m\leq d$ to a bounded number of e/i tests can be applied in the case of any disc, not necessarily isolated.
+If at least one of the e/i tests fails, then definitely there is a root in a narrow annulus about the boundary circle of an input disc, and hence the internal disc of the annulus would not contain $m$ roots.
+Otherwise we would obtain a root-free concentric annulus about the boundary circle of an input disc and then reduce $m$-test to root-counting in an isolated disc, which we can perform at a low cost (see, e.g., claim (ii) of Prop.
+7.1).
+Algorithm 15.1.
+Reduction of $m$-test to e/i tests.
+INPUT: a polynomial $p(x)$ of a degree $d$, two values $\bar{\theta}>1$ and $\bar{\sigma}>1$, an integer $m$ such that $1\leq m\leq d$ and #$\#(D(0,\bar{\sigma}))\leq m$, and two algorithms $\mathbb{ALG}~{}0$ and $\mathbb{ALG}~{}1$, which for a black box polynomial perform $\bar{\sigma}$-soft e/i test for in any disc and root-counting in an isolated disc, respectively.
+INITIALIZATION (see Fig. 6): Fix a positive $\rho$ such that $\bar{\sigma}=O(\rho)$, say, $\rho=\bar{\sigma}/5$, and compute the integer
+$v=\lceil 2\pi/\phi\rceil$ (15.1)
+where $\phi$ is the angle defined by an arc of $C(0,1)$ with end points at the distance $\rho$, and so $v=O(1)$ for $\bar{\sigma}=O(\rho)$. Then compute the values
+$c_{j}=\exp(\phi_{j}{\bf i}),~{}j=0,1,\dots,v,$ (15.2)
+such that $|c_{j}-c_{j-1}|=\rho$ for $j=1,2,\dots,v$ and $\phi_{j}-\phi_{j-1}=\phi$.
+$c_{1}$$c_{2}$$c_{3}$$c_{4}$$c_{5}$$c_{6}$$c_{7}$$c_{8}$$c_{9}$$c_{0}$
+Figure 6: $v=9$, $|c_{j}-c_{j-1}|=\rho>|c_{9}-c_{0}|$ for $j=1,2,\dots,9$.
+OUTPUT: Either an upper bound $r_{d-m+1}<1-\rho\frac{\sqrt{3}}{2}$ or one of the two lower bounds $r_{d-m+1}\geq 1-\bar{\sigma}\rho$ or $r_{d-m+1}>1+(\sqrt{3}-1)\rho$.
+COMPUTATIONS: Apply algorithm $\mathbb{ALG}~{}0$ (soft e/i test) to the discs $D(c_{j},\rho)$, $j=0,1,\dots,v$.
+(a) Stop and certify that $r_{d-m+1}\geq 1-\bar{\sigma}\rho$, that is, #$\#(D(0,1-\bar{\sigma}\rho))<m$ for the open disc $(D(0,1-\bar{\sigma}\rho)$, unless algorithm $\mathbb{ALG}~{}0$ outputs exclusion in all its $v+1$ applications.
+(b) If it outputs exclusion $v+1$ times, then apply algorithm $\mathbb{ALG}~{}1$ (root-counter) to the $\theta$-isolated disc $D(0,1)$ for
+${}\theta:=\min\Big{\{}1+(\sqrt{3}-1)\rho,\frac{1}{1-0.5\rho\sqrt{3}}\Big{\}}.$ (15.3)
+Let $\bar{s}_{0}=\#(D(0,1))$ denotes its output integer. Then conclude that
+(i) $r_{d-m+1}\leq 1-\rho\frac{\sqrt{3}}{2}$, and so #$\#(D(0,1-\rho\frac{\sqrt{3}}{2}))=m$ if $\bar{s}_{0}=m$, while
+(ii) $r_{d-m+1}\geq 1+(\sqrt{3}-1)\rho$, and so #$\#(D(0,1+(\sqrt{3}-1)\rho))<m$ for the open disc $D(0,1+(\sqrt{3}-1)\rho)$ if $\bar{s}_{0}<m$.
+To prove correctness of Alg. 15.1 we need the following result.
+Theorem 15.1.
+For the above values $c_{0},\dots,c_{v}$, and $\rho$, the domain $\cup_{j=0}^{v-1}D(c_{j},\rho)$ covers the annulus $A(0,1-0.5\rho\sqrt{3},1+(\sqrt{3}-1)\rho)$.
+Proof.
+Let $c_{j-1}$ and $c_{j}$ be the centers of two neighboring circles $C(c_{j-1},\rho)$ and $C(c_{j},\rho)$, let $z_{j}$ and $z^{\prime}_{j}$ be their intersection points such that $|z_{j}|=|z_{j}^{\prime}|-|z_{j}-z^{\prime}_{j}|$ for $j=1,2,\dots,v$, observe that the values $|z_{j}-z_{j}^{\prime}|$, $|z_{j}|$, $|z^{\prime}_{j}|$, and $|c_{j-1}-c_{j}|$ are invariant in $j$, and hence so is the annulus $A(0,|z_{j}|,|z_{j}^{\prime}|)$ of width $|z_{j}-z_{j}^{\prime}|$ as well; notice that it lies in the domain $\cup_{j=0}^{v-1}D(c_{j},\rho)$.
+$c_{0}$$c_{1}$$z$$z^{\prime}$
+Figure 7: $|c_{1}-c_{0}|=|c_{0}-z|=|c_{1}-z|=|c_{0}-z^{\prime}|=|c_{1}-z^{\prime}|=\rho$.
+Let us prove that $|z_{j}-z_{j}^{\prime}|=\rho\sqrt{3}$ for $j=1$, and hence also for $j=2,\dots,v$.
+Write $z:=z_{1}$ and $z^{\prime}:=z_{1}^{\prime}$ and observe (see Fig.
+7) that $|c_{1}-c_{0}|=|c_{0}-z|=|c_{1}-z|=|c_{0}-z^{\prime}|=|c_{1}-z^{\prime}|=\rho$, and so the points $c_{0}$, $c_{1}$, and $z$ are the vertices of an equilateral triangle with a side length $\rho$.
+Hence $|z-z^{\prime}|=\rho\sqrt{3}$.
+The circle $C(c_{j},\rho)$ intersects the line segment $[z_{j},z_{j}^{\prime}]$ at a point $y_{j}$ such that $|y_{j}-z_{j}|>\rho\sqrt{3}/2$ and $|y_{j}-z_{j}^{\prime}|>(\sqrt{3}-1)\rho$ for $\rho<1$, and this implies the theorem if $|c_{v}-c_{0}|=\rho$.
+If, however, $|c_{v}-c_{0}|<\rho$ and if $C(c_{0},\rho)\cap C(c_{v},\rho)=\{\bar{z},\bar{z}\}$, then $|\bar{z}-\bar{z}^{\prime}|>|z-z^{\prime}|$, and the theorem follows in this case as well.
+∎
+Corollary 15.1.
+The disc $D(0,1)$ is $\theta$-isolated for $\theta$ of (15.3) provided that the discs $D(c_{j},\rho)$ contain no roots for $j=0,1,\dots,v$.
+Correctness of Alg.
+15.1.
+(i) If algorithm $\mathbb{ALG}~{}1$ outputs inclusion in at least one of its applications, then one of the discs $D(c_{j},\bar{\sigma}\rho)$ contains a root, and then Alg.
+15.1 correctly reports that $r_{d-m+1}\geq 1-\sigma\bar{\rho}$.
+(ii) Otherwise Cor. 15.1 holds, and then correctness of Alg. 15.1 follows from correctness of $\mathbb{ALG}~{}1$.
+Softness is at most $\frac{1}{1-\bar{\sigma}\rho}$, $\frac{1}{1-0.5\rho\sqrt{3}}$, and $1+(\sqrt{3}-1)\rho$ in cases (a), (b,i), and (b,ii), respectively.
+Computational complexity.
+Alg.
+15.1 invokes algorithm $\mathbb{ALG}~{}0$ at most $v+1$ times for $v=O(1)$ of (15.1).
+If $\mathbb{ALG}~{}0$ outputs exclusion $v+1$ times, then algorithm $\mathbb{ALG}~{}1$ is invoked and runs at a dominated cost.
+Remark 15.1.
+One can modify Alg.
+15.1 as follows (see [63]): instead of fixing a positive $\rho$ and then computing $v$ of (15.1) fix an integer $v>0$, apply algorithm $\mathbb{ALG}~{}0$ at $v$ equally spaced points $c_{j}=\exp(\frac{2j\pi{\bf i}}{v}),~{}j=0,1,\dots,v-1$, let $\phi=\frac{2\pi}{v}$, and then define $\rho$.
+We also arrive at this algorithm if we compute or guess this value $\rho$ and then apply Alg.
+15.1, but notice that $c_{v}=c_{0}$ for $c_{j}$ of (15.2) and only invoke algorithm $\mathbb{ALG}~{}0$ for $c_{0},\dots,c_{v-1}$.
 15.2 Simple reduction to a single e/i test in the case of an isolated disc
@@
 Now ensure that $V_{f}<\frac{1}{2\sqrt{q}}$, say, by requiring that $\max_{g=0}^{q-1}|{\rm NIR}_{fg}(\zeta^{g})|<\frac{1}{4\sqrt{q}}$ because bound (15.5) implies that $\frac{d-m}{\theta-1}<\frac{1}{4\sqrt{q}}$.
-12.2 $m$-test by means of e/i tests for any disc
-Our alternative reduction of $m$-test for $1\leq m\leq d$ to a bounded number of e/i tests can be applied in the case of any disc, not necessarily isolated.
-If at least one of the e/i tests fails, then definitely there is a root in a narrow annulus about the boundary circle of an input disc, and hence the internal disc of the annulus would not contain $m$ roots.
-Otherwise we would obtain a root-free concentric annulus about the boundary circle of an input disc and then reduce $m$-test to root-counting in an isolated disc, which we can perform at a low cost (see, e.g., claim (ii) of Prop.
-3.1).
-Algorithm 12.1.
-Reduction of $m$-test to e/i tests.
-INPUT: a polynomial $p(x)$ of a degree $d$, two values $\bar{\theta}>1$ and $\bar{\sigma}>1$, an integer $m$ such that $1\leq m\leq d$ and #$\#(D(0,\bar{\sigma}))\leq m$, and two algorithms $\mathbb{ALG}~{}0$ and $\mathbb{ALG}~{}1$, which a black box polynomial perform $\bar{\sigma}$-soft e/i test for in any disc and root-counting in an isolated disc, respectively.
-INITIALIZATION (see Fig. 6): Fix a positive $\rho$ and compute the integer
-$v=\lceil 2\pi/\phi\rceil$ (12.4)
-where $\phi$ is the angle defined by an arc of $C(0,1)$ with end points at the distance $\rho$. Then compute the values
-$c_{j}=\exp(\phi_{j}{\bf i}),~{}j=0,1,\dots,v,$ (12.5)
-such that $|c_{j}-c_{j-1}|=\rho$ for $j=1,2,\dots,v$ and $\phi_{j}-\phi_{j-1}=\phi$.
-$c_{1}$$c_{2}$$c_{3}$$c_{4}$$c_{5}$$c_{6}$$c_{7}$$c_{8}$$c_{9}$$c_{0}$
-Figure 6: $v=9$, $|c_{j}-c_{j-1}|=\rho>|c_{9}-c_{0}|$ for $j=1,2,\dots,9$.
-OUTPUT: Either an upper bound $r_{d-m+1}<1-\rho\frac{\sqrt{3}}{2}$ or one of the two lower bounds $r_{d-m+1}\geq 1-\bar{\sigma}\rho$ or $r_{d-m+1}>1+(\sqrt{3}-1)\rho$.
-COMPUTATIONS: Apply algorithm $\mathbb{ALG}~{}0$ (soft e/i test) to the discs $D(c_{j},\rho)$, $j=0,1,\dots,v$.
-(a) Stop and certify that $r_{d-m+1}\geq 1-\bar{\sigma}\rho$, that is, #$\#(D(0,1-\bar{\sigma}\rho))<m$ for the open disc $(D(0,1-\bar{\sigma}\rho)$, unless algorithm $\mathbb{ALG}~{}0$ outputs exclusion in all its $v+1$ applications.
-(b) If it outputs exclusion $v+1$ times, then apply algorithm $\mathbb{ALG}~{}1$ (root-counter) to the $\theta$-isolated disc $D(0,1)$ for
-${}\theta:=\min\Big{\{}1+(\sqrt{3}-1)\rho,\frac{1}{1-0.5\rho\sqrt{3}}\Big{\}}.$ (12.6)
-Let $\bar{s}_{0}=\#(D(0,1))$ denotes its output integer. Then conclude that
-(i) $r_{d-m+1}\leq 1-\rho\frac{\sqrt{3}}{2}$, and so #$\#(D(0,1-\rho\frac{\sqrt{3}}{2}))=m$ if $\bar{s}_{0}=m$, while
-(ii) $r_{d-m+1}\geq 1+(\sqrt{3}-1)\rho$, and so #$\#(D(0,1+(\sqrt{3}-1)\rho))<m$ for the open disc $D(0,1+(\sqrt{3}-1)\rho)$ if $\bar{s}_{0}<m$.
-To prove correctness of Alg. 12.1 we need the following result.
-Theorem 12.1.
-For the above values $c_{0},\dots,c_{v}$, and $\rho$, the domain $\cup_{j=0}^{v-1}D(c_{j},\rho)$ covers the annulus $A(0,1-0.5\rho\sqrt{3},1+(\sqrt{3}-1)\rho)$.
-Proof.
-Let $c_{j-1}$ and $c_{j}$ be the centers of two neighboring circles $C(c_{j-1},\rho)$ and $C(c_{j},\rho)$, let $z_{j}$ and $z^{\prime}_{j}$ be their intersection points such that $|z_{j}|=|z_{j}^{\prime}|-|z_{j}-z^{\prime}_{j}|$ for $j=1,2,\dots,v$, observe that the values $|z_{j}-z_{j}^{\prime}|$, $|z_{j}|$, $|z^{\prime}_{j}|$, and $|c_{j-1}-c_{j}|$ are invariant in $j$, and hence so is the annulus $A(0,|z_{j}|,|z_{j}^{\prime}|)$ of width $|z_{j}-z_{j}^{\prime}|$ as well; notice that it lies in the domain $\cup_{j=0}^{v-1}D(c_{j},\rho)$.
-$c_{0}$$c_{1}$$z$$z^{\prime}$
-Figure 7: $|c_{1}-c_{0}|=|c_{0}-z|=|c_{1}-z|=|c_{0}-z^{\prime}|=|c_{1}-z^{\prime}|=\rho$.
-Let us prove that $|z_{j}-z_{j}^{\prime}|=\rho\sqrt{3}$ for $j=1$, and hence also for $j=2,\dots,v$.
-Write $z:=z_{1}$ and $z^{\prime}:=z_{1}^{\prime}$ and observe (see Fig.
-7) that $|c_{1}-c_{0}|=|c_{0}-z|=|c_{1}-z|=|c_{0}-z^{\prime}|=|c_{1}-z^{\prime}|=\rho$, and so the points $c_{0}$, $c_{1}$, and $z$ are the vertices of an equilateral triangle with a side length $\rho$.
-Hence $|z-z^{\prime}|=\rho\sqrt{3}$.
-The circle $C(c_{j},\rho)$ intersects the line segment $[z_{j},z_{j}^{\prime}]$ at a point $y_{j}$ such that $|y_{j}-z_{j}|>\rho\sqrt{3}/2$ and $|y_{j}-z_{j}^{\prime}|>(\sqrt{3}-1)\rho$ for $\rho<1$, and this implies the theorem if $|c_{v}-c_{0}|=\rho$.
-If, however, $|c_{v}-c_{0}|<\rho$ and if $C(c_{0},\rho)\cap C(c_{v},\rho)=\{\bar{z},\bar{z}\}$, then $|\bar{z}-\bar{z}^{\prime}|>|z-z^{\prime}|$, and the theorem follows in this case as well.
-∎
-Corollary 12.1.
-The disc $D(0,1)$ is $\theta$-isolated for $\theta$ of (12.6) provided that the discs $D(c_{j},\rho)$ contain no roots for $j=0,1,\dots,v$.
-Correctness of Alg.
-12.1.
-(i) If algorithm $\mathbb{ALG}~{}1$ outputs inclusion in at least one of its applications, then one of the discs $D(c_{j},\bar{\sigma}\rho)$ contains a root, and then Alg.
-12.1 correctly reports that $r_{d-m+1}\geq 1-\sigma\bar{\rho}$.
-(ii) Otherwise Cor. 12.1 holds, and then correctness of Alg. 12.1 follows from correctness of $\mathbb{ALG}~{}1$.
-Softness is at most $\frac{1}{1-\bar{\sigma}\rho}$, $\frac{1}{1-0.5\rho\sqrt{3}}$, and $1+(\sqrt{3}-1)\rho$ in cases (a), (b,i), and (b,ii), respectively.
-Computational complexity.
-Alg.
-12.1 invokes algorithm $\mathbb{ALG}~{}0$ at most $v+1$ times for $v$ of (12.4).
-If $\mathbb{ALG}~{}0$ outputs exclusion $v+1$ times, then algorithm $\mathbb{ALG}~{}1$ is invoked.
-Remark 12.1.
-One can modify Alg.
-12.1 as follows (see [60]): instead of fixing a positive $\rho$ and then computing $v$ of (12.4) fix an integer $v>0$, apply algorithm $\mathbb{ALG}~{}0$ at $v$ equally spaced points $c_{j}=\exp(\frac{2j\pi{\bf i}}{v}),~{}j=0,1,\dots,v-1$, let $\phi=\frac{2\pi}{v}$, and then define $\rho$.
-We also arrive at this algorithm if we compute or guess this value $\rho$ and then apply Alg.
-12.1, but notice that $c_{v}=c_{0}$ for $c_{j}$ of (12.5) and only invoke algorithm $\mathbb{ALG}~{}0$ for $c_{0},\dots,c_{v-1}$.
 16 Precision of computing
 16.1 Doubling computational precision towards its a posteriori estimation
-Given an algorithm $\mathbb{A}$ for black box polynomial computations and target tolerance TOL to output error, how should one choose computational precision?
-A simple practical recipe of [10, 14, 21] was the basis for highly successful empirical performance of MPSolve: namely, first apply the algorithm $\mathbb{A}$ with a lower precision and then recursively double it until an output error decreases below TOL; at that point it outputs precision within a factor of two from optimal, thus also supporting near optimal a posteriori bit operation complexity estimate for algorithm $\mathbb{A}$.
+Given an algorithm $\mathbb{A}$ for black box polynomial computations and target tolerance TOL to the output errors, how should one choose computational precision?
+Here is a simple practical recipe from [10, 15, 22] and MPSolve for obtaining a posteriori estimate within a factor of two from optimal: first apply algorithm $\mathbb{A}$ with a lower precision, then recursively double it and stop where an output error decreases below TOL.
 16.2 A priori estimates for computational precision
-In the rest of this section we deduce a priori estimate for computational precision by proving that a precision of $O(b\log(d))$ bits is sufficient for supporting correctness of our subdivision root-finders.
-Such an estimate is deduced in [73] based on advanced non-trivial techniques, but we only use a very small part of them because the paper [73] works with higher order divided differences, while we only use the first order ones combined with Las Vegas randomization and some rather routine techniques for numerical error estimation.
-Estimating computational precision of our soft e/i tests we assume that they are applied to the unit disc $D(0,1)$ because we can extend the results to any disc $D(c,\rho)$ by applying map (1.4) and increasing the upper bounds on the precision by a factor of $\log_{2}(1/\rho)$, which is at most $b$ because we only need to handle discs of radii $\rho>\epsilon=1/2^{b}$ and not $\epsilon$-discs.
-Therefore, our bound $O(\log(d))$ on the computational precision of our e/i tests for the disc $D(0,1)$ will imply the target bound $O(b\log(d))$ for testing any disc.
-We also make Assumption 0 (Relative error bound) that we can request and obtain from evaluation oracle (black box subroutine) the values of $p$ with a relative error bound in $1/d^{\nu}$ for any fixed constant $\nu>0$ of our choice; we can represent these value with the bit-precision $\log_{2}(d)+\log_{2}(\nu)$.
+In the rest of this section we deduce a priori estimate of $O(b\log(d))$ bits for computational precision, proving that this precision supports correctness of our subdivision root-finders.
+Such an estimate is deduced in [76] based on advanced non-trivial techniques, of which we only use a very small part.
+Estimating computational precision of our soft e/i tests we assume that they are applied to the unit disc $D(0,1)$ because we can extend the results to any disc $D(c,\rho)$ by applying map (1.3) and increasing the upper bounds on the precision by a factor of $\log_{2}(1/\rho)$.
+This factor is at most $b$ because we only handle discs of radii $\rho>\epsilon=1/2^{b}$ until we stop reaching $\epsilon$-discs.
+Therefore, our bound $O(\log(d))$ on the computational precision of our e/i tests for the disc $D(0,1)$ will imply the bound $O(b\log(d))$ for any disc.
+We also make Assumption 0 that we can request and obtain from evaluation oracle (black box subroutine) the values of $p$ with a relative error bound in $1/d^{\nu}$ for any fixed constant $\nu>0$ of our choice; we can represent these values with the bit-precision $\log_{2}(d)+\log_{2}(\nu)$.
 In our e/i tests we only need to compare the values $|{\rm NIR}_{t_{h}}(x)|$ at some complex points $x$ with some fixed values $v_{d}\geq 1/d^{\beta}$ for a constant $\beta>0$, and next we estimate that under Assumption 0 we can indeed do this by performing our computations with a precision of $O(\log(d))$ bits.
 Multiplication with a precision $s$ contributes a relative error at most $2^{s}$ (cf.
-[20, Eqn.
+[21, Eqn.
 A.3 of Ch.
-3]) and otherwise just sums the relative error bounds of the factors because we can ignore the dominated impact of higher order errors.
-Hence we deduce that $h=O(\log(d))$ multiplications for root-lifting to their $h$th powers increase these bounds at most by a factor of $h$, which is immaterial since we allow relative errors of order $1/d^{\nu}$ for a constant $\nu>0$ of our choice.
-We only need to apply this analysis to approximation of NIR$(x)$ for $p_{h}(x)$ based on expression (1.6).
-In our analysis we also use the following
-Assumption 1 (Isolation Assumption): Whp all $q$ evaluation points $x\in C(0,1)$ of an e/i test for the unit disc $D(0,1)$ lie at least at the distance $\phi:=\frac{2\pi}{mqw}$ from all zeros of $p$ for any fixed $w>1$ of our choice.
-We justify this assumption by applying our e/i tests under random rotation $x\mapsto vx$ for the random variable $v$ specified in our next proposition.
+3]) and otherwise just sums the relative error bounds of the factors ignoring the dominated impact of higher order errors.
+Hence we deduce that $h=O(\log(d))$ multiplications for root-lifting to their $h$th powers increase these bounds by at most a factor of $h$, which is immaterial since we allow relative errors of order $1/d^{\nu}$ for a constant $\nu>0$ of our choice.
+We only apply this analysis to approximation of NIR$(x)$ for $p_{h}(x)$ based on (3.2) and also use the following
+Assumption 1: Whp all $q$ evaluation points $x\in C(0,1)$ of an e/i test for the unit disc $D(0,1)$ lie at the distance at least $\phi:=\frac{2\pi}{mqw}$ from all roots for any fixed $w>1$ of our choice.
+This assumption holds because we apply our e/i tests under random rotation $x\mapsto vx$ for the random variable $v$ specified in our next proposition.
 Proposition 16.1.
-Let $\phi_{g}:=\min_{j=1}^{d}|v\zeta^{g}-z_{j}|$ and let $\phi:=\min_{g=0}^{q-1}\phi_{g}$ for the zeros $z_{j}$ of $p$, $\zeta$ of (1.10) and random variable $v$ sampled from a fixed arc of the circle $C(0,1)$ of length $2\pi/q$ under the uniform probability distribution on that arc.
+Let $\phi_{g}:=\min_{j=1}^{d}|v\zeta^{g}-z_{j}|$ and let $\phi:=\min_{g=0}^{q-1}\phi_{g}$ for the roots $z_{j}$, $\zeta$ of (3.5) and random variable $v$ sampled from a fixed arc of the circle $C(0,1)$ of length $2\pi/q$ under the uniform probability distribution on that arc.
 Let $i(D(0,1))=m$.
@@
 ∎
-Next we prove that precision of $O(\log(d))$ bits supports our e/i tests under Assumptions 0 and 1 by applying Prop.
+Next we prove that a precision of $O(\log(d))$ bits supports our e/i tests under Assumptions 0 and 1 by applying Prop.
 16.1 for sufficiently large $w=d^{O(1)}$.
@@
 Thus we shall increase the above error bound by adding upper bounds $\alpha$ on $|\frac{t^{\prime}(x)}{t(x)}-\frac{t^{\prime}(y)}{t(x)}|$ and $\beta$ on the rounding error of computing $\frac{t(x)-t(x-\delta)}{\delta t(x)}$, for $x=\zeta^{g}$ and $g=1,2,\dots,q$.
-In the proof of the next theorem we use the following lemma, which is [73, Fact 3.5].
+In the proof of the next theorem we use the following lemma, which is [76, Fact 3.5].
 Lemma 16.1.
@@
 Proposition 16.3.
-Under the assumptions of Prop. 13.2 it holds that
+Under the assumptions of Prop. 16.2 it holds that
 $|t(x)\alpha|=|\frac{t(x)-t(x-\delta)}{\delta}-t^{\prime}(x)|\leq|t(x)\frac{(d\xi^{\prime})^{2}\delta}{1-d\xi^{\prime}\delta}|$ for $\xi^{\prime}=\frac{1}{\rho}$.
@@
 The claimed bound on $|\xi^{\prime}|$ follows from (3.1).
-It remains to apply the first five lines of the proof of [73, Lemma 3.6] with $f$, $\alpha$, $n$, and $\xi$ replaced by $t$, $-\delta$, $d$, and $w$, respectively.
+It remains to apply the first five lines of the proof of [76, Lemma 3.6] with $f$, $\alpha$, $n$, and $\xi$ replaced by $t$, $-\delta$, $d$, and $w$, respectively.
 Namely, first obtain from Taylor’s expansion that $t(x)-t(x-\delta)=\sum_{j=0}^{\infty}\frac{\delta^{j}}{j!}t^{(j)}(x)-t(x)$.
@@
 Suppose that the values $t(y)$ have been computed by a black box oracle within a relative error bounds $\nabla(y)$ for $y=x$, $y=x-\delta$, and $\delta$ of Cor.
-13.2.
+16.2.
 Then one can ensure that $\beta<\frac{1}{d^{\bar{\eta}}}$ for any fixed constant $\bar{\eta}$ by choosing a proper $\nabla(y)$ of order $1/d^{O(1)}$, represented with a precision of $O(\log(d))$ bits, which we can do under Assumption 0.
@@
 the proof of Prop.
-6.1).
+10.1).
 Hence $w:=|\frac{t(x-\delta)}{t(x)}|\leq d^{O(1)}$ and $|\beta\delta|\leq|w|\frac{|\nabla(x)|+|\nabla(x-\delta)|}{1-|\nabla(x)|}$.
@@
 ∎
-Combine Cors. 13.1 and 13.2 and Prop. 13.4 to obtain
+Combine Cors. 16.1 and 16.2 and Prop. 16.4 to obtain
 Corollary 16.3.
@@
 We derived our precision bound $O(\log(d))$ (in the case of root-finding in the unit disc $D(0,1)$) based on operating entirely with NIR$(x)$.
-For comparison, if we relied on computing $p(x)$ and on the bound $\min_{x:~{}|x|=1}(|p(x)|/\sum_{i=0}^{d}|p_{i}|)\geq(\frac{\theta-1}{2\theta})^{d}$ of [130, Eqn.
+For comparison, if we relied on computing $p(x)$ and on the bound $\min_{x:~{}|x|=1}(|p(x)|/\sum_{i=0}^{d}|p_{i}|)\geq(\frac{\theta-1}{2\theta})^{d}$ of [134, Eqn.
 (9.4)], we would have only obtained a bound of order $m^{2}d$ on the computational precision of our root-finders.
@@
 Given a $d\times d$ matrix $M$ that has characteristic polynomial $p(x)=\det(xI-M)$, our root-finders can approximate all $m$ its zeros (that is, the eigenvalues of $M$) lying in a fixed disc $D$ on the complex plane isolated from external eigenvalues.
-Recall Gershgorin’s bound for any eigenvalue $\lambda$ of a matrix $M=(a_{i,j})_{i,j=1}^{d}$ [134, Thm. 1.3.2]: $|\lambda-a_{i,i}|\leq\sum_{j\neq i}|a_{i,j}|$ for some $i$, $1\leq i\leq d$.
+Recall Gershgorin’s bound for any eigenvalue $\lambda$ of a matrix $M=(a_{i,j})_{i,j=1}^{d}$ [138, Thm. 1.3.2]: $|\lambda-a_{i,i}|\leq\sum_{j\neq i}|a_{i,j}|$ for some $i$, $1\leq i\leq d$.
 Hence, clearly, $|\lambda|\leq||M||_{F}$.
 Combine this bound with the precision estimates of the previous section for our root-finders [^27] For a posteriori estimates, one can combine our root-finders with the recipe of doubling the precision of computing of (see Sec.
-13.1).
+16.1).
 and with the following result of Storjohann [137] (cf.
@@
 Remark 17.1.
-Our root-finders are essentially reduced to evaluation of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ at sufficiently many points $x$.
-Hence we can rely on the following expressions for approximation of the eigenvalues of a matrix or a matrix polynomial:
+Our root-finders are essentially reduced to the evaluation of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ at sufficiently many points $x$, and we can rely on the following expressions for approximation of the eigenvalues of a matrix or a matrix polynomial:
 $\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}((xI-M)^{-1})}$
@@
 $\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}(M^{-1}(x)M^{\prime}(x))}$
-where $M(x)$ is a matrix polynomial and $p(x)=\det(M(x))$ [19, Eqn.
+where $M(x)$ is a matrix polynomial and $p(x)=\det(M(x))$ [20, Eqn.
 (5)].
@@
 1.5 we can whp decrease this bound by a factor of $m$.
-Next, however, we decrease this bound roughly by a factor of $d$ by using benefits of fast multipoint evaluation of a polynomial by means of the algorithm of Fiduccia [40] and Moenck and Borodin [81] (see [18, Sec.
+Next, however, we decrease this bound roughly by a factor of $d$ by using benefits of fast multipoint evaluation of a polynomial by means of the algorithm of Fiduccia [42] and Moenck and Borodin [84] (see [19, Sec.
 4.5] for historical account), non-trivially extended in [81] to support the following theorem:
@@
 Proof.
-For $q=d$ this is [78, Thm.
+For $q=d$ this is [81, Thm.
 2], based on using novel data structure for hyperbolic approximation of $p$, proposed and analyzed by Guillaume Moroz in [81].
@@
 $\mathbb{B}_{\rm Kirr}=\tilde{O}((d+q)(b+q)),$ (18.2)
-due to [65, Thm.
+due to [68, Thm.
 3.9, Alg.
@@
 (18.1) and (18.2) on the overall bit operation complexity by computing with multiple precision for larger values $b$.
-Moroz obtained [78, Thm.
+Moroz obtained [81, Thm.
 2] by applying bound (18.2) to his lower degree polynomials approximating $p(x)$.
@@
 18.1 Reduction of our root-finders to multipoint polynomial approximation
-Given a complex $c$, a pair of positive $b$ and $\rho$, and $d+1$ coefficients of a polynomial $p=p(x)$ of (1.1) such that the disc $D(c,\rho)$ is isolated and contains precisely $m$ zeros of $p$, we approximate all these zeros within $R/2^{b}$ for $R=|c|+\rho$ by applying subdivision root-finding iterations with e/i tests of Sec.
-11.
+Given a complex $c$, a pair of positive $b$ and $\rho$, and $d+1$ coefficients of a polynomial $p=p(x)$ of (1.1) such that the disc $D(c,\rho)$ is isolated and contains precisely $m$ roots, we approximate all these roots within $R/2^{b}$ for $R=|c|+\rho$ by applying subdivision root-finding iterations with e/i tests of Sec.
+14.
 Then every subdivision step is reduced to application of our e/i tests to $\bar{m}=O(m)$ discs $D(c_{\lambda},\rho_{\lambda})$, where $\rho_{\lambda}\geq R/2^{b}$ for all $\lambda$.
@@
 Assumption 1 of Sec.
-13.2).
-Then $\log(\frac{1}{|t(x)|})=O(d\log(m))$ by virtue of the bound of [130, Eqn.
+16.2).
+Then $\log(\frac{1}{|t(x)|})=O(d\log(m))$ by virtue of the bound of [134, Eqn.
 (9.4)]).
-[^40] That bound follows from the estimate $|f|\geq 1/2^{d}$ of [130, Thm.
+[^32] That bound follows from the estimate $|f|\geq 1/2^{d}$ of [134, Thm.
 4.2], which holds provided that $t(x)=f\prod_{j=1}^{m}(x-x_{j})\prod_{j=m+1}^{d}(\frac{x}{x_{j}}-1)$, $||t(x)||_{1}=1$, $|x_{j}|<1$ for $j\leq m$, and $|x_{j}|>1$ for $j>m$.
@@
 For $m=d$, the above bound on $\log(\frac{1}{|t(x)|})$ is sharp up to a constant factor and is reached for the polynomial $t(x)=(x+1-\frac{1}{m})^{d}/(2-\frac{1}{m})^{d}$, $\rho_{\lambda}=1$, and $x=-1$.
-For $m<d$, [130, Thm.
+For $m<d$, [134, Thm.
 4.5] implies a little stronger bounds, namely, $\log(\frac{1}{|t(x)|})=O(d+m\log(m))$ and hence $\bar{b}=O(m\log(m)+d+b)$ because all discs $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ lie in $\theta$-isolated unit disc $D(0,1)$ for $\theta-1$ exceeding a positive constant.
@@
 1.5 we can decrease this bound by a factor of $m$ whp and can extend this decrease to the related subsequent estimates.
-Based on [65, Alg. 5.3] and Eqn. (15.2) we obtain a little larger but still reasonably small upper bound
+Based on [68, Alg. 5.3] and Eqn. (18.2) we obtain a little larger but still reasonably small upper bound
 $\mathbb{B}_{\rm roots}^{\prime}=\tilde{O}(((m^{2}+d)(m^{2}+d+b)b).$ (18.4)
-In Part IV we decrease the number of subdivision steps of our root-finders from order of $b$ to $O(\log(b))$, thus decreasing bounds (15.3) and (15.4) by a factor of $b$.
-As by-product we ensure $\epsilon$-isolation and arrive at our Main Theorems 2, 3, and 4.
+In Part IV we decrease the number of subdivision steps of our root-finders from order of $b$ to $O(\log(b))$, thus decreasing bounds (18.3) and (18.4) by a factor of $b$, and arrive at our Main Theorems 2, 3, and 4.
 PART IV: Compression of a Disc and Acceleration of Subdivision Root-Finders
-16 Overview
-Classical subdivision iterations converge to all zeros of a polynomial $p$ with linear convergences rate (see Sec.
-1.6), but clearly this is very slow when they process a component $\mathbb{C}$ made up of a single suspect square or up to four of them sharing a vertex.
-In such cases the algorithms of [124, 97] dramatically accelerate convergence to superlinear rate by means of Newton’s or Schröder’s iterations (see Fig.
-7), but the authors of [124, 97] have not controlled the bit operation cost of their algorithms.
+19 Overview and two alternative approaches
+Classical subdivision iterations converge to the roots with linear convergence rate (see Sec.
+2), which is slow when many consecutive iterations process a component $C$ made up of a single suspect square or up to four of them sharing a vertex.
+In this case the algorithms of [128, 101] dramatically accelerate convergence by means of applying Newton’s or Schröder’s iterations (see Fig.
+8), which converge with superlinear rate to the convex hull of all roots lying in $C$.
+The diameter of the convex hull, however, can be too large, and then [128, 101] apply subdivision iterations again.
+The resulting root-finder recursively intertwines subdivision iterations with Newton’s or Schröder’s, and this complicates the complexity estimates for a root-finder.
+[128, 101] have only estimated the number of ops but not of the bit operations involved.
+[23, 24] estimated the bit operation complexity, but only under a very strong assumption about the output (see Remark 4.2 and Sec.
+12.4).
 Figure 8: Five blue marked centers of suspect squares converge to a green-marked root cluster with linear rate; two red marked Schröder’s iterates converge with quadratic rate.
-Alternatively, one can apply algorithms of [130] to split $p(x)$ into the product of two polynomials $f(x)$ and $g(x)$ whose zero sets lie in and outside a disc covering the isolated component $\mathbb{C}$, respectively, and then apply a root-finder to the polynomials $f=f(x)$ and $g=g(x)$.
-Splitting of $p(x)$ into factors, however, does not preserve sparsity of $p(x)$, and the factors may have significantly larger coefficient length than $p(x)$.
-(E.g., splitting $p(x)=x^{1000}+1$ into two factors of degrees 300 and 700.) [^42] The near optimal root-finder of [92, 99] relies on by splitting $p(x)$ into factors of comparable degrees, but under the assumption of computing with a rather high precision of order $d\log(d)$.
-These problems disappear if one of the factors, say, $g$ has a small degree: one can deflate it, approximate its zeros, and then approximate the zeros of $f(x)=p(x)/g(x)$ by applying implicit deflation (cf.
-(1.14)).
-We stay within subdivision framework and like [124, 97, 22, 23] partition the union of all suspect squares at each subdivision step into pairwise isolated connected components but propose a novel algorithm that compresses such a component $\mathbb{C}$ into a near minimal disc containing all roots that lie in the component.
-This algorithm enables us to decrease the estimates of Secs. 8.3 and 15.2 for the overall complexity of root-finding by this factor, thus providing formal support for our Main Theorems 2-4.
+Instead of Newton’s or Schröder’s iterations, one can apply the algorithm of [134], which splits $p(x)$ into the product of two polynomials $f(x)$ and $g(x)$ whose zero sets lie in and outside the minimal disc covering the component $C$, respectively.
+Then one can apply a root-finder to the polynomials $f=f(x)$ and $g=g(x)$.
+Splitting of $p(x)$ into factors does not preserve sparsity of $p(x)$, and the factors may have significantly larger coefficient length than $p(x)$.
+(E.g., split $p(x)=x^{1000}+1$ into two factors of degrees 300 and 700.)
+Overall asymptotic bit-complexity for high precision root-finding can still be near-optimal if a root-finder recursively splits $p(x)$ into factors of comparable degrees.
+This is achieved in [96, 103], but the resulting root-finder is much involved and extensively uses the coefficients of $p$.
+The problems caused by splitting disappear if one of the factors, say, $g$ has a small degree; then we can deflate it, approximate its zeros readily, and approximate the zeros of $f(x)=p(x)/g(x)$ by applying to it a black box root-finder and using implicit deflation based on (3.3) or the equation NIR$(p)=$NIR(f)+limit-$(f)+$NIR$(g)$ for $p=fg$.
+One would lose optimality, however, by always deflating low-degree factors: to deflate $k$ times one would need at least order of $bkmd$ bit operations overall, that is, would exceed the lower bound $0.25(d+1)db$ of Cor.
+24.2 by at least a factor of $k$.
+We stay with subdivision framework but instead of Newton’s or Schröder’s iterations, used in [128, 101], apply our novel algorithm that compresses a component isolated from its external roots into a near-minimal disc that keeps all internal roots.
+This enables us to decrease the estimates of Secs.
+12.3 and 18.2 for the overall complexity of root-finding by a factor of $b$, thus completing proof of our Main Theorems 2-4.
 We organize our presentation in this part as follows.
@@
 In particular we observe in Sec.
-18.4 that the whole subdivision process involves at most $12m-$ 1 compression steps.
+21.4 that the whole subdivision process involves at most $12m-$ 1 compression steps, reduced to $O(m)$ e/i tests.
 In Sec.
-19 we prove that in addition it only involves $O(m)$ e/i steps.
-Together with our estimates for the complexity of an e/i step this enables us to bound the overall complexity of our root-finders and reach near optimal bound on the bit operation complexity of our root-finders.
-We reach that bound in Sec.
-20 by devising and analyzing our alternative compression algorithm.
-In that section we also present an alternative fast root-counter in an isolated disc.
-We use some results of Part I, Alg.
-5.1 of Part II, and Sec.
-10 (Background) of Part III.
-Otherwise our presentation in this part of the paper is self-contained.
-Our fast root-counter of Sec.
-20.4 can serve as alternative to root-counters of Part II.
+22 we prove that in addition it also involves $O(m)$ e/i steps.
+We reach a near-optimal bound on the bit operation complexity of our root-finders in Sec.
+23 by devising and analyzing our compression algorithm and recalling our estimates for the complexity of an e/i step.
+In that section we also present a fast root-counter in an isolated disc.
+We use some results of Part I, Alg. 9.1 of Part II, and Sec. 13 (Background) of Part III. Otherwise our presentation in this part of the paper is self-contained.
 20 Component tree and compact components
-17.1 Component tree: definition
-Partition the union of all suspect squares at a subdivision step into components; a step can break some of them into more than one component.
-Represent all such partitions with a component tree $\mathbb{T}$.
-At most $m$ components, said to be the leaves of the tree, can be covered by isolated discs of radius at most $\epsilon$ for a fixed tolerance $\epsilon\geq 0$.
-The other vertices are components-parents, each broken by a subdivision step into at least two components-children.
+20.1 Component tree, edge length, non-adjacent suspect squares, and compact components
+Recall the definition of the component tree $\mathbb{T}$ of Sec.
+4, which represent partition of the union of suspect squares at every subdivision step and representation of such partition with the component tree $\mathbb{T}$ of Sec.
+4.
+The tree has at most $m$ components-leaves that represent the output $\epsilon$-discs and at most $m-1$ other vertices, which represent components-parents, each broken by a subdivision step into at least two components-children.
 The initial suspect square is the root of the tree $\mathbb{T}$.
-The edges of the tree connect parents and children.
-The tree does not shows modifications of components between their partitions represented by the edges.
-Observation 17.1.
-Having at most $m$ leaves, the tree $\mathbb{T}$ has at most $m-1$ other nodes.
-These upper bounds are sharp – attained where a subdivision process defines the complete binary tree $\mathbb{T}$ with $m$ leaves.
-17.2 Edge length, non-adjacent suspect squares, and compact components
-The edges of the tree $\mathbb{T}$ represent subdivision steps between partitions of components. Cor. 17.1 will bound the number of such steps; one can view this as the edge length.
+The edges of the tree connect parents and children, representing subdivision steps between partitions of components.
+Cor.
+20.1 below bounds the number of such steps; one can consider this number the edge length.
 Call a pair of suspect squares non-adjacent if they share no vertices.
 Call a component compact if it contains just a single non-adjacent suspect square, that is, consists of at most four suspect squares sharing a vertex (see Fig. 10).
-Observation 17.2.
+Figure 9: Two discs superscribing two non-adjacent suspect squares are not minimal but do not overlap.
+Observation 20.1.
 (See Fig. 9.) The $\sigma$-covers of two non-adjacent suspect squares are separated by at least the distance $a\Delta$ for the side length $\Delta$ and separation coefficient
@@
 20.2 Isolation of compact components
-Observation 17.3.
+Observation 20.2.
 (i) A compact component is surrounded by a frame of discarded squares (see Fig.
-8) and hence is covered by a 2-isolated square.
+10) and hence is covered by a 2-isolated square.
 (ii) If it stays compact in $h$ successive subdivision steps, then its minimal covering square and disc become $2^{h+1}$-isolated and $2^{h}\sqrt{2}$-isolated, respectively.
@@
 Fig.
-7), but then we accelerate this process by interrupting it with compression steps in at most three iterations.
-Next we specify compression for a 3.5-isolated unit disc $D=D(0,1)$ covering a component; map (1.4) extends compression to any 3.5-isolated disc $D(c,\rho)$; the map $\rho\epsilon\mapsto\epsilon$ adjusts the error tolerance $\epsilon$ to scaling by $\rho$.
-Observation 17.3 implies that we can readily ensure $\theta$-isolation at the cost of performing $O(\log(\theta))$ additional subdivision steps.
+8), but then we accelerate this process by interrupting it with compression steps in at most three iterations.
+Next we specify compression for a 3.5-isolated unit disc $D=D(0,1)$ covering a component; map (1.3) extends compression to any 3.5-isolated disc $D(c,\rho)$; the map $\rho\epsilon\mapsto\epsilon$ adjusts the error tolerance $\epsilon$ by scaling by $\rho$.
+Observation 20.2 implies that we can readily ensure $\theta$-isolation at the cost of performing $O(\log(\theta))$ additional subdivision steps.
 Algorithm 21.1.
 $\epsilon$-compression (high level description).
-INPUT: a positive $\epsilon\leq 1$ and a black box polynomial $p(x)$ of a degree $d$ such that [^43] Newton’s and Schröder’s iterations of [124, 97] accelerate convergence of subdivision iterations to a root cluster in a disc if the disc is $5d^{2}$-isolated, while (18.1) only requires 3.5-isolation.
+INPUT: a positive $\epsilon\leq 1$ and a black box polynomial $p(x)$ of a degree $d$ such that [^34] Newton’s and Schröder’s iterations of [128, 101] accelerate convergence of subdivision iterations to a root cluster in a disc if the disc is $5d^{2}$-isolated, while Alg.
+21.1 only requires 3.5-isolation (see (21.1)).
 $i(D(0,1))\geq 3.5~{}{\rm and}~{}\#(D(0,1))\geq 1.$ (21.1)
 OUTPUT: an integer $m$, a complex $c^{\prime}$, and a positive $\rho^{\prime}$ such that $m=\#(D(0,1))$, (i) $X(D(c^{\prime},\rho^{\prime}))=X(D(0,1))$, and (ii) $\rho^{\prime}\leq\epsilon$ or $8|z_{g}-z_{h}|>5\rho^{\prime}$ for a pair of roots $z_{g},z_{h}\in D(c^{\prime},\rho^{\prime})$.
+[Hence the disc $D(c^{\prime},\rho^{\prime})$ is 5/8-rigid.]
 INITIALIZATION: Fix $\delta\geq 0$ such that $8\delta\leq\epsilon\leq 1$.
-COMPUTATIONS: Compute $m=\#(D(0,1))$, $c^{\prime}$, $\rho^{\prime}$ such that
+COMPUTATIONS: Compute $m=\#(D(0,1))$, $c^{\prime}$, and $\rho^{\prime}$ such that
 $|c^{\prime}-c|\leq\delta~{}{\rm and}~{}\rho\leq\rho^{\prime}\leq\max\{\epsilon,\rho+2\delta\}$ (21.2)
@@
 Furthermore, #$\#(D(c^{\prime},\rho^{\prime}))=m$ since $\rho^{\prime}=\rho:=r_{d+1-m}(c^{\prime})$, while #$\#(D(0,1))=m$ by assumption.
-Now (18.1) implies that #$\#(D(0,3))=m$, which supports claim (i) for the output of Alg. 18.1.
-Next let $|c^{\prime}-z_{g}|=\rho:=r_{d+1-m}(c^{\prime})$ for a zero $z_{g}$ of $p(x)$.
+Now (21.1) implies that #$\#(D(0,3))=m$, which supports claim (i) for the output of Alg. 21.1.
+Next let $|c^{\prime}-z_{g}|=\rho:=r_{d+1-m}(c^{\prime})$ for a root $z_{g}$.
 Then all $m$ roots in $D(0,1)$ also lie in the disc $D(c^{\prime},\rho)=D(c^{\prime},\rho^{\prime})$ by virtue of claim (i), just proved.
@@
 (21.2)).
-Hence $|z_{g}-z_{h}|>3\rho^{\prime}/4$ (implying claim (ii) for the output of Alg. 18.1) because $42\delta\leq\epsilon/4<\rho^{\prime}/$ 4.
+Hence $|z_{g}-z_{h}|>3\rho^{\prime}/4$ (implying claim (ii) for the output of Alg. 21.1) because $42\delta\leq\epsilon/4<\rho^{\prime}/$ 4.
 Next extend this correctness proof to the general case based on bounds (21.2).
@@
 Therefore $D(c^{\prime},\rho^{\prime})\subseteq D(0,3.5)$ because $8\delta\leq\epsilon\leq 1$ by assumption.
-Now recall that #$\#(D(0,3.5))=\#(D(0,1))=m$, while #$\#(D(c^{\prime},\rho^{\prime}))\geq\#(D(c^{\prime},\rho))=m$ for $\rho^{\prime}\geq\rho$ and deduce claim (i) for the output of Alg. 18.1.
-Next wlog assume that $c=0$, $c^{\prime}=\delta$, and $z_{g}=c^{\prime}+\rho$ for a root $z_{g}$.
+Now recall that #$\#(D(0,3.5))=\#(D(0,1))=m$, while #$\#(D(c^{\prime},\rho^{\prime}))\geq\#(D(c^{\prime},\rho))=m$ for $\rho^{\prime}\geq\rho$ and deduce claim (i) for the output of Alg. 21.1.
+Next assume wlog that $c=0$, $c^{\prime}=\delta$, and $z_{g}=c^{\prime}+\rho$ for a root $z_{g}$.
 Then $\Re(z_{g})\geq\rho-\delta\geq\rho^{\prime}-3\delta$ because $|c^{\prime}|=|c^{\prime}-c|\leq\delta$ and $\rho^{\prime}\leq\rho+2\delta$ by assumptions.
@@
 Therefore, $|z_{g}-z_{h}|\geq\Re(z_{g})-\Re(z_{h})>\rho^{\prime}-3\delta\geq\rho^{\prime}-\frac{3}{8}\epsilon$.
-This implies claim (ii) for the output of Alg. 18.1 since $\rho^{\prime}>\epsilon$.
+This implies claim (ii) for the output of Alg. 21.1 since $\rho^{\prime}>\epsilon$.
 21.3 Elaboration upon the compression algorithm
-We can readily compute $m=\#(D(0,1))$ for isolated input disc $D(0,1)$ of Alg.
-18.1, e.g., by applying Alg.
-3.1, and can extend this recipe to any disc $D(c,\rho)$ by applying map (1.4) and Prop.
-3.1.
-Moreover, in Sec.
-20.4 we describe a very fast alternative root-counter.
-Furthermore, based on Prop.
-3.2 we can compute close approximations $\tilde{s}_{1}$ to the sum $s_{1}$ of the $m$ zeros of $p$ lying in that disc and then closely approximate their average $c^{\prime}:=s_{1}/m$, lying in the convex hull $CH(D(c,\rho))$.
-Finally, we can approximate the $m$th smallest root radius $r_{d-m+1}(c^{\prime},p)$ by applying our root radius estimation of Sec.
-5, based on $\ell$-tests for $\ell=m$.
-There we can apply $\ell$-tests of Part II or combine e/i tests of Sec.
-11 with reduction of $m$-test to e/i tests in Sec.
-12.
-Computation of a sufficiently close approximation to the sum $s_{1}$ of the zeros of $p$ lying in the disc $D(0,1)$ and of the average $s_{1}/m$ is too costly to support Thms.
-1.1 and 1.3 (we would need to evaluate $p(x)$ at $O(b\log(d)$ points and similarly for bit operation cost, but we fix this deficiency in Sec.
-20 by presenting another variant of compression algorithm.
+We readily (i) count roots in isolated input disc $D(0,1)$ of Alg.
+21.1 by applying the fast root-counter of Sec.
+23.4 or alternatively Alg.
+7.1 and Prop.
+7.1 and (ii) compute a complex point $c^{\prime}$ lying in or near the convex hull of the roots in $C$ by applying fast algorithm of Sec.
+23.
+Finally, we estimate the root radius $\rho^{\prime}:=r_{d-m+1}(c^{\prime})$ by applying Alg.
+9.1, which incorporates the $m$-test of of Sec.
+23.4, or alternatively Alg.
+7.1 and Prop.
+7.1.
 21.4 Counting $\epsilon$-compression steps
-$c^{\prime}$ approximates within $\epsilon$ all $m$ roots lying in the disc $D(0,1)$ if $\rho^{\prime}\leq\epsilon$ for the output value $\rho^{\prime}$ of Alg.
-18.1.
-Then call the compression final and stop computations.
+Recall from Sec.
+1.3 that the output disc $D(c^{\prime},\rho^{\prime})$ of Alg.
+21.1 is said to be an $\epsilon$-disc if $\rho^{\prime}\leq\epsilon$; in this case call the compression by Alg.
+21.1 final and output that disc.
 Otherwise call the square $S=S(c^{\prime},\rho^{\prime})$ suspect and apply to it subdivision iterations.
 Extend Lemma 20.1 to the case where $|z_{g}-z_{h}|\geq\frac{5}{8}\rho^{\prime}$ for a pair of roots $z_{g},z_{h}\in D(c^{\prime},\rho^{\prime})$, combine it with claim (ii) for the output of Alg.
-18.1, and obtain that at most $2.5+\log_{2}(8\sigma m/5)$ subdivision steps break the square $S$ into more than one component.
-Until that moment apply no new compression steps and hence apply less than $2m$ such steps overall – no more than there are nodes in the partition tree $\mathbb{T}$.
-Now elaborate upon Observation 17.1 as follows.
+21.1, and obtain that at most $2.5+\log_{2}(8\sigma m/5)$ subdivision steps break the square $S$ into more than one component.
+From part (ii) of Sec. 4 recall the bound $22m-$ 2 on the number of compression steps and obtain
 Theorem 21.1.
 For a fixed $\epsilon\geq 0$ let subdivision and compression steps be applied to a component $C$ of suspect squares containing at most $m$ roots.
-Then at most $m$ final compression steps among at most $12m-$ 1 compression steps overall output isolated discs of radii at most $\epsilon$ that cover all roots lying in $C$.
-Corollary 18.1.
-Under the assumptions of Thm. 18.1 all subdivision steps process at most $36m-$ 3 compact components, made up of at most $1224m-$ 12 suspect squares; this involves at most $1224m-$ 12 e/i tests.
+Then the computations involve at most $66m-$ 6 compact components, made up of at most $2424m-$ 24 suspect squares; that is, these computations involve at most $2424m-$ 24 e/i tests.
 Proof.
-A compact component stays unbroken and uncompressed in at most $3$ consecutive subdivision steps, which end with a compression step or follow a non-final compression step, that is, such components are processed in at most $36m-$ 3 subdivision , which involve at most $1224m-$ 12 e/i tests overall.
+A compact component stays unbroken and uncompressed in at most $3$ consecutive subdivision steps, which end with a final compression step or follow a non-final compression step and end with breaking a component.
+Hence all such components are processed in at most $66m-$ 6 subdivision step, each involving at most 4 e/i tests.
 ∎
@@
 22.1 An upper estimate
-The upper estimate $12m-$ 1 of Thm.
-18.1 is sharp – reached in the case of the complete binary tree $\mathbb{T}$ and implying that our accelerated subdivision process can involve $12m-$ 1 suspect squares (e/i steps).
-This reaches lower estimate up to a constant factor:
+The upper estimate $22m-$ 2 of part (ii) of Sec.
+4 on the number of compression steps is sharp – reached in the case of the complete binary tree $\mathbb{T}$.
+Each compression step is preceded by an e/i test, and so we perform at least $22m-$ 2 e/i steps.
+According to the following result, which is the main theorem of this section, this lower bound on the number of e/i tests involved is sharp up to a constant factor:
 Theorem 22.1.
@@
 22.2 The first lemma
-Our proof of Thm. 19.1 begins with
+We begin our proof of Thm. 22.1 with
 Lemma 22.1.
@@
 Lemma 22.2.
-Fix two integers $i\geq 0$ and $g\geq 3$, a value $\sigma$ of softness, $1<\sigma<\sqrt{2}$, define the separation fraction $a:=2-\sigma\sqrt{2}$ of Observation 17.2.
+Fix two integers $i\geq 0$ and $g\geq 3$, fix a value $\sigma$ of softness, $1<\sigma<\sqrt{2}$, and define the separation fraction $a:=2-\sigma\sqrt{2}$ of Observation 20.1.
 For all $j$ let $j$-th subdivision step process $v_{j}$ components made up of $w_{j}$ suspect squares, $n_{j}$ of which are non-adjacent, let $\Delta_{j}$ denote their side length, and let the $j$-th steps perform no compression for $j=i,i+1,\dots,i+g$.
@@
 ∎
-By virtue of Cor.
-18.1 all compact components together involve $O(m)$ suspect squares and e/i tests; it remains to prove that we only involve $O(m)$ other e/i tests, for which $n_{j}\geq 2$ for all $j$.
 Next we extend Lemma 22.2 by allowing compression steps.
-We only need to extend the equation $2\Delta_{j+1}=\Delta_{j}$ to the case where the $j$th subdivision step for $i\leq j\leq i+g$ may perform compression, after which we apply subdivision process to a single suspect square with side length at most $4\Delta_{j}$.
-The next two subdivision steps perform only five e/i tests, that is, at most $510m-$ 5 e/i tests are performed at all pairs of subdivision steps following all (at most $12m-$ 1) compression steps.
-Then again, we can exclude all these e/i tests and bound by $O(m)$ the number of e/i tests in the remaining subdivision process, which is not interrupted with compression steps anymore.
-Furthermore, we resume subdivision process after a compression step with a suspect square whose side length decreased (in two discounted subdivision steps) from at most $4\Delta_{j}$ to at most $\Delta_{j}$.
+We only need to extend the equation $2\Delta_{j+1}=\Delta_{j}$ to the case where the $j$th subdivision step for $i\leq j\leq i+g$ performs non-final compression, each followed with subdivision steps applied to a single suspect square with side length at most $4\Delta_{j}$.
+The next two subdivision steps perform only five e/i tests, making up at most $105m-$ 10 e/i tests performed at all pairs of subdivision steps following all (at most $m-2$) non-final compression steps.
+It remains to bound by $O(m)$ the number of e/i tests in the remaining subdivision steps, not interrupted with compression steps.
+Furthermore, after a compression step we resume subdivision process with a suspect square whose side length bound decreases in two next subdivision steps from at most $4\Delta_{j}$ to at most $\Delta_{j}$.
 Hence we can apply to it Lemma 22.2, because now it accesses the same number $n_{j}$ of non-adjacent suspect squares and an increased number $\sigma_{j}$ of all suspect squares (cf. (22.1)).
 22.4 Counting e/i tests at partition steps
-Let $h$ be the height of our partition tree $\mathbb{T}$ and count suspect squares processed in subdivision steps that partition components, temporarily ignoring the other steps.
+Let $h$ denote the height of our partition tree $\mathbb{T}$ and count suspect squares processed in subdivision steps that partition components, temporarily ignoring the other steps.
 Clearly, bounds (22.1) still hold and can only be strengthened. Notice that $v_{i}\geq 1$ for all $i$, sum $v_{i+g}$ as well as bounds (22.1) for a fixed $g>0$ and $i=0,1,\dots,h-g$, and obtain
@@
 22.5 Extension to counting all e/i steps
-Next consider a non-compact component, count all suspect squares processed in all subdivision steps applied until it is partitioned, and then prove in Lemma 19.3 that up to a constant factor the overall number of these suspect squares is given by their number at the partition step.
-Such a result will extend to all subdivision steps our count $n^{\prime}=O(m)$ of (19.3), so far restricted to partition steps.
+Next consider a non-compact component, count all suspect squares processed in all subdivision steps applied until it is partitioned, and then prove in Lemma 22.3 that up to a constant factor the overall number of these suspect squares is bounded by their overall number at all partition steps.
+Such a result extends to all subdivision steps the bound $n^{\prime}=O(m)$ of (22.3), so far restricted to partition steps.
 Together with Lemma 22.1 this will complete our proof of Thm.
-19.1.
+22.1.
 Lemma 22.3.
@@
 Then $\sum_{j=0}^{k-1}n_{j}\leq n_{k}\frac{10\sqrt{2}}{a}$ for the separation fraction $a:=2-\sigma\sqrt{2}$ (cf.
-Observation 17.2.)
+Observation 20.1.)
 Proof.
@@
 (ii) Our proof of that lemma covers the worst case where a long chains between the roots $x$ and $y$ can consist of up to $4m$ suspect squares, but for typical inputs such chains are much shorter and can consist of just a few suspect squares.
-(iii) In the proof of Lemma 19.3 we assume that $n_{i}-1\geq n_{i}/2$; this holds for any $n_{i}\geq 2$ but the bounds become stronger as $n_{i}$ increases.
+(iii) In the proof of Lemma 22.3 we assume that $n_{i}-1\geq n_{i}/2$; this holds for any $n_{i}\geq 2$, but the bounds are strengthened as $n_{i}$ increases.
 (iv) In a component made up of five suspect squares at least two are non-adjacent, and so the bound of Lemma 22.1 can be strengthened at least in this case.
@@
 21.1 compresses an input disc $D(0,1)$ into $\eta$-rigid disc for $\eta=5/8$, but our proof of Thm.
-19.1 only involves $\eta$ the argument in Sec.
-18.4, and this argument and therefore the whole proof of the theorem can be immediately extended if it relies on compression into any rigid disc.
-20 Computation of a point near root set and root-counting in a disc
+22.1 only involves $\eta$ in the argument in Sec.
+21.4, and this argument and therefore the whole proof of the theorem can be immediately extended if it relies on compression into any rigid disc.
+23 A point near root set and root-count in a disc
+We prove that the algorithms of this section work where an input disc is $d^{\nu}$-isolated for a fixed positive constant $\nu$.
+In application to compression we begin with an isolated disc, say, 3.5-isolated, but can turn it into $d^{\nu}$-isolated for any fixed positive constant $\nu$ by applying $h$ additional subdivision steps or $h$-lifting for $h=O(\log(d))$.
+This comment can be extended to application of e/i test of Sec.
+23.4 instead of Cauchy-based e/i tests of Part II.
 23.1 Computation of a point near root set: introductory comments
-Recall Schröder’s iterations [128], celebrated for fast convergence to a root of multiplicity $m$ [77, Sec. 5.4]:
-$z=y-m~{}{\rm NR}(y).$ (20.1)
-Due to map (1.4) we only need to consider the case where the input disc $D$ is the unit disc $D(0,1)$. We first assume that it contains all $d$ zeros of $p$ and readily verify
-Observation 20.1.
-Schröder’s single iteration (20.1) for $m=d$ and $p(x)=p_{d}(x-z_{1})^{d}$ outputs the zero $z_{1}$ of $p(x)$ unless $y=z_{1}$.
-Next we extend this observation by proving that the output $z$ of Schröder’s iteration (20.1) lies near the convex hull $CH(X(D))$ provided that $m=d$, a disc $D$ contain all $d$ zeros of $p$, and point $y$ lies far enough from the disc $D$.
+Recall Schröder’s iterations [132],
+$z=y-m~{}{\rm NR}(y),$ (23.1)
+which fast converge to a root of multiplicity $m$ [80, Sec.
+5.4].
+Due to map (1.3) we only need to consider the case where the input disc $D$ is the unit disc $D(0,1)$.
+We first assume that it contains all $d$ roots and readily verify
+Observation 23.1.
+Schröder’s single iteration (23.1) for $m=d$ and $p(x)=p_{d}(x-z_{1})^{d}$ outputs the root $z_{1}$ unless $y=z_{1}$.
+Next we extend this observation by proving that the output $z$ of Schröder’s iteration (23.1) lies near the convex hull $CH(X(D))$ provided that $m=d$, a disc $D$ contains all $d$ roots, and point $y$ lies far enough from the disc $D$.
 Here are our specific estimates.
@@
 Let $\bar{x}\in CH$.
-Then $|z-\bar{x}|\leq\alpha~{}\Delta$ for $\alpha:=\frac{(|y|-1)^{2}}{(y|-2)^{2}}$.
+Then $|z-\bar{x}|\leq\alpha~{}\Delta$ for $\alpha:=\frac{(|y|-1)^{2}}{(y|-2)^{2}}=(1+\frac{1}{(y|-2})^{2}$, and so $\alpha=2.25$ for $u|=4$ and $\alpha=25/16$ for $u|=6$, $\alpha\mapsto 1$ $|y|\mapsto\infty$.
 Proof.
@@
 Substitute the bounds $\Delta\leq 2$ and $|\bar{x}|\leq 1$ and obtain the theorem. ∎
-Since $|z-\bar{x}|\geq r_{d}(z,p)$, Thm. 20.2 implies
 Corollary 23.1.
-Under the assumptions of Thm. 20.2 it holds that $r_{d}(z)\leq 2.25\Delta$ for $|y|\geq 4$, while $r_{d}(z)\leq 25\Delta/16$ for $|y|\geq 6$.
-Readily verify
-Observation 20.2.
-Let #$\#(D)=m$ for a disc $D$ and let $\Delta:=\Delta(CH(X(D)))$ denote the diameter of the convex hull of the root set $X(D)$ in that disc.
-Write $r_{d-m}(z,p):=0$, $r:=r_{d-m+1}(z,p)$, and $\rho:=r_{d}(z,p)$ for a complex $z$ and let $r+\Delta\leq r_{d-m}(z,p)$, which implies that $r$ and $\rho$ do not depend on the $d-m$ zeros of $p$ lying outside the disc $D$.
-Then $r-\rho\leq\Delta\leq 2r$.
-Corollary 20.2.
-If $\rho\leq\nu\Delta$ under the assumptions of Observation 20.2, then the disc $D(z,r)$ is $1/(2+2\nu)$-rigid.
-Now we devise our fast compression algorithm.
+Under the assumptions of Thm. 23.2 the disc $D(z,r_{1}(z))$ is $\frac{1}{2+2\alpha}$-rigid.
+Proof.
+Verify the claim in the case where $x_{d}=0$, $x_{1}=\Delta=1$, e.g., $0\leq x_{j}\leq 1$ for all $j$, and $z=1+\alpha$ and notice that this is a worst case input. ∎
+Now we devise our fast compression algorithm inthe case where $m=d$.
 Algorithm 23.1.
@@
 #$\#(D(0,1))=d.$ (23.4)
-OUTPUT: a complex $z$, and a positive $r$ such that (i) $X(D(z,r))=X(D(0,1))$ r equivalently #$\#(D(z,r))=d$, and (ii) $r\leq\epsilon$ or the disc $D(z,r)$ is rigid.
+OUTPUT: a complex $z$, and a positive $r$ such that #$\#(D(z,r))=d$ and either $r\leq\epsilon$ or otherwise the disc $D(z,r)$ is $\frac{2}{13}$-rigid.
 INITIALIZATION: Fix $y\geq 4$.
 COMPUTATIONS:
-(i) Compute and output $z=y-d{\rm NR}~{}(y)$ (cf. (20.1)).
-(ii) Compute an approximation $\tilde{r}\geq 0$ by applying Alg. 5.1 for $m=d$, a constant $\beta>1$, and $\epsilon$ replaced by $\epsilon/4$).
+(i) Compute and output $z=y-d~{}{\rm NR}~{}(y)$.
+(ii) Compute an approximation $\tilde{r}\geq 0$ to $r_{1}(z)$ by applying Alg. 9.1 for $m=d$, a constant $\beta>1$, and $\epsilon$ replaced by $\epsilon/4$.
 If $\tilde{r}\leq 3\epsilon/4$, output $z$ and $r=\epsilon$, defining $\epsilon$-compression into the disc $D(z,r)$ of radius $\epsilon$.
-Otherwise output $r=\tilde{r}+0.25\epsilon$, defining compression into the rigid disc $D(z,r)$.
+Otherwise output $r=\tilde{r}+0.25\epsilon$, defining compression into rigid disc $D(z,r)$.
 Correctness of the algorithm is readily verified if Alg.
-5.1 outputs a range $[\rho_{-}\rho_{+}]$ of length at most $\epsilon/4$, for in this case $r_{1}\leq r=\tilde{r}+\epsilon/4\leq\epsilon$, and #$\#(D(z,\epsilon))=d$.
-Otherwise $1/\beta\leq\tilde{r}/r_{1}\leq\beta$ for a constant $\beta>1$ of Alg.
-5.1, while the disc $D(x,r_{1})$ is $\frac{1}{2\nu+2}$-rigid by virtue of Cor.
-20.2 where $\nu\leq 2.25$ by virtue of Cor.
-20.1 since $y>4$ in our case.
-It follows that the disc $\bar{D}:=D(z,r_{1}(z))$ is rigid.
-This compression queries the evaluation oracle twice while performing Schröder’s single iteration for computing $z$ and otherwise runs at the cost of performing Alg.
-5.1, estimated in Cor.
-5.1.
+9.1 outputs a range $[\rho_{-},\rho_{+}]$ of length at most $\epsilon/4$, for in this case $r_{1}\leq r=\tilde{r}+\epsilon/4\leq\epsilon$, and #$\#(D(z,\epsilon))=d$.
+Otherwise $1/\beta\leq\tilde{r}/r_{1}\leq\beta$ for a constant $\beta\geq 1$ of Alg.
+9.1, while the disc $D(x,r_{1})$ is $\frac{1}{2\alpha+2}$-rigid by virtue of Cor.
+23.1 where $\alpha\leq 2.25$ by virtue of THm.
+23.1 because $y\geq 4$ in our case.
+It follows that the disc $\bar{D}:=D(z,r_{1}(z))$ is $\frac{2}{13}$-rigid.
+Remark 23.1.
+By the same argument we can ensure that unless $\tilde{r}\leq 3\epsilon/4$ the disc $D(z,r_{1}(z))$ is $\eta$-isolated for any $\eta>4$ if we initialize Alg. 23.1 at a sufficiently large $y$.
+This compression queries evaluation oracle twice while performing Schröder’s single iteration for computing $z$ and otherwise runs at the cost of performing Alg.
+9.1, estimated in Cor.
+9.1.
 We can decrease the latter cost bound for a large class of polynomials $p$ by choosing larger tolerance $\epsilon$.
-Then we yield desired compression if the algorithm outputs $r$ exceeding 1/4 of that tolerance value.
+Then we will yield desired compression if the algorithm outputs $r$ exceeding $\epsilon/4$.
 Otherwise we can recursively repeat the computations for tolerance recursively decreased to $\epsilon/4$.
@@
 Now suppose that #$\#(D(0,1))=m<d$ and $i(D(0,1))\geq\theta$ for a sufficiently large $\theta$ of order $d^{O(1)}$.
-Then verify that the $d-m$ external zeros of $p$, lying outside the disc $D(0,1)$, little affect the proof of Thm.
+Then verify that the $d-m$ external roots, lying outside the disc $D(0,1)$, little affect the proof of Thm.
 23.1 and our whole study in the previous subsection.
-Thus we first perform $O(\log(d))$ subdivision steps (see Observation 17.3).
+Thus we first perform $O(\log(d))$ subdivision steps (see Observation 20.2).
 We do not need compression if these steps break the component into more than one component or compress it into an $\epsilon$-disc.
 Otherwise the component remains compact and hence in $O(\log(d))$ subdivision steps shrinks enough to become $d^{\nu}$-isolated for any fixed constant $\nu$, in which case we apply our compression algorithm.
-20.4 A fast root-counter based on Schröder’s iterations
+23.4 A fast root-counter and a fast e/i test based on Schröder’s iterations
 Let the unit disc be sufficiently strongly isolated and let a single Schröder’s iteration (23.1) be applied at a point $y$ lying sufficiently far from the unit disc $D(0,1)$.
-Then our study in this subsection implies that the iteration outputs a point $z=z(m)$ lying in or near the disc $D(0,1)$, say, in the disc $D(0,2)$, if and only if #$\#(D(0,1))=m$, that is, we arrive at a root-counter in $D(0,1)$ that evaluates $p(x)$ at two points.
+Then our study in this subsection implies that the iteration outputs a point $z=z(m)$ lying in or near the disc $D(0,1)$, say, in the disc $D(0,2)$, if and only if #$\#(D(0,1))=m$, that is, we count roots in $D(0,1)$ at the cost of evaluation of $p(x)$ at two points.
 Indeed, if #$\#(D(0,1))=m$, then, as we have showed in this section, the iteration outputs a point $z=z(m)$ lying near the convex hull of the root set in $D(0,1)$ and hence lying in or near the disc $D(0,1)$, say, in the disc $D(0,2)$ if the value $|y|$ is large enough.
-In this case $|m{\rm NR}(y)-y|\leq 2$, and so $|{\rm NR}(y)-y|/m\leq 2/m$.
-If #$\#(D(0,1))=\bar{m}\neq m$, however, then the iteration outputs a point $z=z(m)$ that lies at the distance at least $|\bar{m}-m|\cdot|{\rm NR}(y)|-2$ from the disc $D(0,2)$, which must be positive if $|y|$ is sufficiently (although reasonably) large.
+In this case $|m~{}{\rm NR}(y)-y|\leq 2$, and so $|{\rm NR}(y)-y|/m\leq 2/m$.
+If #$\#(D(0,1))=\bar{m}\neq m$, then the iteration outputs a point $z=z(m)$ that lies at the distance at least $|\bar{m}-m|\cdot|{\rm NR}(y)|-2$ from the disc $D(0,2)$, which must be large if $|y|$ is sufficiently (although reasonably) large.
+Extension to a fast e/i test.
+Now suppose that the disc $D(0,1)$ is sufficiently well isolated and apply this root-counter to the polynomial $xp(x)$ for a large $y$ and $m=1$, in which case Schröder’s iterations turn into Newton’s.
+In this case root-counter will show that the disc $D(0,1)$ contains no roots of $p(x)$ if and only if its output lies in that disc.
 23.5 Overall complexity of our root-finders
 We have reduced our root-finding in an isolated disc to performing at most $12m-$ 1 compression steps and in addition $O(m)$ e/i tests.
-We have also reduced a compression step to (i) performing $O(\log(d))$ additional subdivision steps (each included to yield isolation of an input disc for compression and reduced to $O(m)$ e/i tests), (ii) application of Schröder’s single iteration, which evaluates polynomial $p(x)$ at two points, and (iii) performing $O(\log(d))$ e/i tests for approximation of the root radius $r_{d-m+1}(z,p)$.
-Summarize these estimates, combine them with the estimates of the previous subsection, and arrive at the complexity bounds of our Main Theorems.
-Remark 20.1.
-If we incorporate Kirrinnis’s algorithm for multipoint polynomial evaluation, [65, Alg.
-5.3], rather than Moroz’s algorithm of [78], then our root-finders would slow down but still would stay very fast and still would use near optimal number of bit operations in the cases where $m^{2}=\tilde{O}(d)$ and whp under the Random Root Model.
-Indeed, by using our compression steps we decrease the complexity estimate of (8.4) and (15.4) by a factor of $b/\log(b)$ to arrive at the bound $\mathbb{B}_{\rm roots}^{\prime}=\tilde{O}(((m^{2}+d)(m^{2}+d+b))$ on the number of bit operations.
-This bound turns into $\tilde{O}((d+b)d)$ if $m^{2}=\tilde{O}(d)$ and whp under the Random Root Model.
-APPENDIX
-Appendix A Applications and implementation of root-finders: brief overview
-Numerous areas of mathematics and computational mathematics studied in the last 100 years had no direct link to polynomial root-finding.
-Furthermore, many computational problems arising in the sciences, engineering, business management, and statistics have been linearized and then solved by using tools from linear algebra, linear programming, fast Fourier transform (FFT), and Fast Multipole Method (FMM).
-[^44] Both FFT and FMM are among the Top Ten Algorithms of 20-th century (cf.
-[26]).
-For the history of FFT see D.E.
-Knuth [66]; on FMM see L.F.
-Greengard and V.
-Rokhlin Jr.
-[48] and [147].
-Such tools sometimes involve root-finding for polynomials but usually of smaller degree $d$, where the available subroutines are sufficiently effective in most cases.
-Root-finding for large degree polynomials, however, still has important applications to various areas of Scientific Computing, in particular to Computer Algebra, signal processing, geometric modeling, control, financial mathematics, and complex dynamics (see McNamee and Pan [79, Introduction]).
-Unlike the near optimal polynomial root-finders of [92, 99], various non-optimal complex polynomial root-finders have been implemented (see e.g., MATLAB “roots”, Maple “solve”, MPSolve, and CCluster).
-In particular, Imbach, P., and Yap have implemented the root-finder of [22] in [61].
-Since its creation in 2000 the package MPSolve by Bini and Fiorentino [14], revised in 2014 by Bini and Robol [21], has been remaining user’s choice for approximation of all $d$ complex roots of a polynomial $p$.
-It implements Ehrlich’s iterations of 1967 [36], aka Aberth’s of 1973 [3], which empirically converge very fast globally – right from the start, under proper initialization, involving the coefficients of $p$.
-Such convergence has been consistently observed in decades-long worldwide applications of these iterations as well as in the case of some other root-finders by means of functional iterations, notably Weierstrass’s iterations [143], aka Durand and Kerner’s [29, 64]).
-Unsuccessful attempts of supporting this empirical behavior formally (e.g., see [4, 51]), however, have recently ended with surprising proofs that Weierstrass’s and Ehrlich’s iterations diverge in the case of inputs that lie on some complex manifolds, being an open domain in the case of Weierstrass’s iterations (see Reinke, Schleicher, and M.
-Stoll [127, 125]).
-Serious competition to MPSolve came in 2001 from the package EigenSolve by Fortune [41], but the later version of MPSolve [21] combines the benefits of both packages.
-An implementations of Newton’s iterations for approximation of all roots by Randig, Schleicher, and R.
-Stoll in [126] and by Schleicher and R.
-Stoll in [138] competed with MPSolve in 2017, but then MPSolve has regained upper hand by means of incorporation of Fast Multipole Method (FMM), even though such implementation is simpler and can be more efficient for Newton’s than Ehrlich’s iterations (see our Sec.
-D.4).
-The root-finder of [22], efficiently implemented in [61], slightly outperforms MPSolve for root-finding in a region of the complex plane containing a small number of roots, although it is by far inferior for the approximation of all $d$ roots of a polynomial $p(x)$.
-[^45] The computational cost of root-finders of [92, 99, 22, 23] decreases at least proportionally to the number of roots in a region of interest such as a disc or a rectangle, while Ehrlich’s, Weierstrass’s, and a number of other functional iterations approximate the roots in such a region almost as slow as all complex roots.
-[57, 60] have already superseded MPSolve according to extensive tests for standard test polynomials, and our present root-finders are significantly and probably even dramatically faster (see Sec.
-1.9).
-Real polynomial root-finding is highly important because in many applications, e.g., to optimization in algebraic geometry and geometric modeling, only real roots of a polynomial are of interest and typically are much less numerous than all $d$ complex roots (cf.
-Kac [63], Edelman and Kostlan [37]).
-Current user’s choice for this task is the root-finder in [67], although it has been already superseded in [59] according to extensive tests for standard test polynomials.
-Our novel algorithm of Sec.
-9 can be readily implemented and seems to be the first that efficiently solves the important numerical problem of approximation of both real and near real zeros of a polynomial.
-Appendix B Polynomial Factorization, its links to Root-finding, and some bit-complexity estimates
-B.1 Polynomial Factorization and its links to Root-finding (the results)
+We have also reduced a compression step to (i) performing $O(\log(d))$ additional subdivision steps (included to strengthen isolation of input discs for compression), (ii) application of Schröder’s single iteration, which evaluates polynomial $p(x)$ at two points, and (iii) performing $O(\log(b))$ times $m$-tests for approximation of the root radius $r_{d-m+1}(z,p)$.
+Overall our root-finders apply e/i tests $O(m\log(d))$ times, $m$-tests $O(m\log(b))$ times, and at most $22m-$ 2 Schröder’s iterations.
+Clearly, the computational cost of Schröder’s is dominated, while in Sec.
+15 we reduced an $m$-test is to $O(1)$ e/i tests, specified in Sec.
+14.
+Based on the latter results we can strengthen our estimates of Sec.
+18 and arrive at the complexity bounds of our Main Theorems.
+Alternatively, instead of estimating the cost of e/i tests and reduction to them of $m$- tests, we can use our estimates for the complexity of $\ell$- tests of Part II for $\ell=1$ and $\ell=m$ and would still arrive at our Main Theorems 2 – 4.
+Part V: Factorization, Root Isolation, and Functional Iterations
+24 Polynomial Factorization, its links to Root-finding, and some bit-complexity estimates
+24.1 Polynomial Factorization and its links to Root-finding (the results)
 Hereafter write $|u|=\sum_{i=0}^{d}|u_{i}|$ for a polynomial $u(x)=\sum_{i=0}^{d}u_{i}x^{i}$.
-The root-finders of [130, 92, 99] solve Problems 0 and $*{}^{*}$ by extending the the solution of the following problem of independent importance:
+The root-finders of [134, 96, 103] solve Problems 0 and 0 by extending the the solution of the following problem of independent importance:
 Problem F.
-Approximate Polynomial Factorization: [^46] The solution of Problem F can be also extended to isolation of the zeros of a square-free polynomial with integer coefficients and approximation of matrix eigenvalues (see our Remark C.4 and Sec.
-7) and has various other applications to modern computations, e.g., to time series analysis, Wiener filtering, noise variance estimation, co-variance matrix computation, and the study of multichannel systems [145, 17, 8, 33, 34, 142].
+Approximate Polynomial Factorization: [^35] The solution of Problem F can be also extended to isolation of the zeros of a square-free polynomial with integer coefficients and approximation of matrix eigenvalues (see our Remark 25.4 and Sec.
+11) and has various other applications to modern computations, e.g., to time series analysis, Wiener filtering, noise variance estimation, co-variance matrix computation, and the study of multichannel systems [150, 18, 8, 35, 36, 147].
 For a polynomial of degree $d$ the known sharp upper and lower bounds on the input precision and number of bit operations for root-finding exceed those for numerical factorization by a factor of $d$.
 Given a real $b$ and $d+1$ coefficients $p_{0},p_{1},\dots,p_{d}$ of (1.1), compute $2d$ complex numbers $u_{j},v_{j}$ for $j=1,\dots,d$ such that
-$|p-\prod_{j=1}^{d}(u_{j}x-v_{j})|\leq\epsilon|p|~{}{\rm for}~{}\epsilon=1/2^{b}.$ (B.1)
+$|p-\prod_{j=1}^{d}(u_{j}x-v_{j})|\leq\epsilon|p|~{}{\rm for}~{}\epsilon=1/2^{b}.$ (24.1)
 Hereafter $\mathbb{B}_{0}(b,d)$ and $\mathbb{B}_{F}(b,d)$ denote the optimal number of bit operations required for the solution of Problems 0 and F, respectively, for an error bound $\epsilon=1/2^{b}$, and next we specify a class of polynomials whose zeros make up no large clusters (cf.
 [14, 30]).
-Definition B.1.
-Polynomials with bounded cluster sizes.
-The zero set of $p$ consists of clusters, made up of a constant number of zeros, each having a diameter at most $1/2^{3-b/d}$; furthermore, these clusters are pairwise separated by the distances at least $1/d^{\tilde{O}(1)}$.
+Definition 24.1.
+Polynomials with bounded root cluster sizes.
+The zero set of $p$ consists of clusters, each made up of a constant number of zeros and having a diameter at most $1/2^{3-b/d}$; furthermore, these clusters are pairwise separated by the distances at least $1/d^{\tilde{O}(1)}$.
 [96, 97, 99, 100, 103] proved the following near optimal bound for Problem F and readily extended it to near optimal solution of Problem 0.
-Theorem B.1.
+Theorem 24.1.
 $\mathbb{B}_{F}(b,d)=\tilde{O}((b+d)d)$.
-In Sec. B.3 we reduce Problem 0 to Problem F and vice versa by applying Schönhage’s results of [130, Cor. 4.3 and Sec. 19]:
-Theorem B.2.
+In Sec. 24.3 we reduce Problem 0 to Problem F and vice versa by applying Schönhage’s results of [134, Cor. 4.3 and Sec. 19]:
+Theorem 24.2.
 (i) $\mathbb{B}_{0}(b,d)=\tilde{O}(\mathbb{B}_{F}(bd,d))$ for any polynomial $p(x)$ of degree $d$, (ii) $\mathbb{B}_{F}(b,d)=\tilde{O}(\mathbb{B}_{0}(b+d+\log(d),d))$ for any polynomial $p(x)$ of degree $d$, and (iii) $\mathbb{B}_{0}(b,d)=\tilde{O}(\mathbb{B}_{F}2(b,d))$ for polynomials of Def.
-B.1, with bounded cluster sizes.
-By combining Thms. B.1 and B.2 obtain the following estimates.
-Theorem B.3.
+24.1, with bounded cluster sizes.
+By combining Thms. 24.1 and 24.2 obtain the following estimates.
+Theorem 24.3.
 One can solve Problem 0 by using (i) $\tilde{O}((b+d)d^{2})$ bit operations for any polynomial $p$ of (1.1) and (ii) $\tilde{O}((b+d)d)$ bit operations for a polynomial $p$ with bounded cluster sizes.
-Corollary B.1.
+Corollary 24.1.
 One can solve Problem 0 by using $\tilde{O}((b+d)d+(b+m)m^{2})$ bit operation for $m$ of Def. 1.1 of Sec. 1.5. The bound turns into $\tilde{O}((b+d)d$ if $m^{2}=\tilde{O}(d)$.
@@
 (ii) apply the algorithm that supports Thm.
-B.3 to the $m$th degree factor of the polynomial $\prod_{j=1}^{d}(u_{j}x-v_{j})$ of (B.1) that shares with this polynomial all its zeros lying in the disc $D(1+\beta)$ for positive $\beta$ of Def.
-1.1.474747We can slightly speed up stage (i) because at that stage we actually only need to approximate factorization of the factor $f$ of degree $m$ of $p$ that shares with $p$ its zeros in $D$.
-B.2 Lower bounds on the bit operation complexity
+24.3 to the $m$th degree factor of the polynomial $\prod_{j=1}^{d}(u_{j}x-v_{j})$ of (24.1) that shares with this polynomial all its zeros lying in the disc $D(1+\beta)$ for positive $\beta$ of Def.
+1.1.363636We can slightly speed up stage (i) because at that stage we actually only need to approximate factorization of the factor $f$ of degree $m$ of $p$ that shares with $p$ its zeros in $D$.
+24.2 Lower bounds on the bit operation complexity
 Readily verify the following information lower bounds.
-Observation B.1.
+Observation 24.1.
 To solve Problem 0 or F one must access $(d+1)b$ bits of input coefficients and perform at least $0.5(d+1)b$ bit operations – at most one operation for two input bits.
 Next strengthen the lower bound for Problem 0.
-Observation B.2.
+Observation 24.2.
 Let $p(x)=(x-v_{1})^{j}f(x)$ for an integer $j>0$ and a monic polynomial $f(x)$ of degree $d-j$.
@@
 Then the polynomial $p_{j}(x)=p(x)-2^{-jb}f(x)$ has the $j$ zeros $v_{1}+\zeta^{g}2^{-b}$ for $g=0,1,\dots,j-1$.
-Observation B.2 implies that one must access the coefficient $p_{d-j}$ within $2^{-jb}$ to approximate within $2^{-b}$ a $j$-multiple zero of $p$ or its zero in a cluster of $j$ zeros, and so a root-finder must process at least $jb$ bits of the coefficient $p_{d-j}$.
+Observation 24.2 implies that one must access the coefficient $p_{d-j}$ within $2^{-jb}$ to approximate within $2^{-b}$ a $j$-multiple zero of $p$ or its zero in a cluster of $j$ zeros, and so a root-finder must process at least $jb$ bits of the coefficient $p_{d-j}$.
 Sum these bounds for $j=1,\dots,d$ and obtain
-Corollary B.2.
-One must access at least $0.5(d-1)db$ bits of the coefficients of $p$ and perform at least $0.25(d-1)db$ bit operations to approximate even a single zero of $p$ within $2^{-b}$.
-Likewise, deduce lower bounds on the number of bit operations required for Problem 0.
-Corollary B.3.
-To solve Problem 0 and even to approximate within $\epsilon$ a single zero of $p$ in the disc $D$, one must access at least $0.5b\max\{d+1,(m-1)m\}$ bits of input coefficients and perform at least $0.25b\max\{d+1,(m-1)m\}$ bit operations for $b=\log_{2}(R/\epsilon)$ and $R=|c|+\rho$.
+Corollary 24.2.
+One must access at least $0.5(d+1)db$ bits of the coefficients of a monic polynomial $p$ and perform at least $0.25(d+1)db$ bit operations to approximate even its single within $2^{-b}$.
 The lower bounds of this subsection imply that the algorithm of [96, 103] and its extension supporting Thm.
-B.1 solve Problems 0 and F of large size by using optimal number of bit operations up to polylog factors or, in simple terms.
-B.3 Linking Factorization and Root-finding (with more details)
-Next we prove Thm. B.2, by applying the results of [130, Cor. 4.3 and Sec. 19].
+24.1 solve Problems 0 and F of large size by using optimal number of bit operations up to polylogarithmic factors.
+24.3 Linking Factorization and Root-finding (with more details)
+Next we prove Thm. 24.2, by applying the results of [134, Cor. 4.3 and Sec. 19].
 Recall that $\mathbb{B}_{0}(b,d)$ and $\mathbb{B}_{F}(b,d)$ denote the optimal number of bit operations for the solution of Problems 0 and F, respectively, for an error bound $\epsilon=1/2^{b}$, and furthermore, recall the lower bounds
-$\mathbb{B}_{0}(b,d)\geq 0.5(d+1)b~{}{\rm and}~{}\mathbb{B}_{F}(b,d)\geq 0.25(d-1)db$ (B.2)
-of [92, 99] and the following result (cf. [130, Thm. 19.1]).
-Theorem B.4.
+$\mathbb{B}_{0}(b,d)\geq 0.5(d+1)b~{}{\rm and}~{}\mathbb{B}_{F}(b,d)\geq 0.25(d+1)db$ (24.2)
+of [96, 103] and the following result (cf. [134, Thm. 19.1]).
+Theorem 24.4.
 Let $r_{1}=\max_{j=1}^{d}|z_{j}|\leq 1$.
 Then a solution of Problem 0 within an error bound $1/2^{b}$ can be recovered at a dominated bit operation cost from the solution of Problem F within an error bound $1/2^{bd+2}$, and so $\mathbb{B}_{0}(b,d)\leq\mathbb{B}_{F}(bd+2,d)$.
-Based on [130, Cor. 4.3] we can also readily relax the assumption that $p(x)$ is monic.
-Furthermore, in view of (B.2), relaxing the bound $r_{1}\leq 1$ would little affect the latter estimate because we can apply map (1.4), which changes the error bound of Problem 0 by at most $\log_{2}(\max\{\rho,1/\rho\})$, for $\rho\geq 1/2^{b}$.
-[132, Thm. 2.7], which we reproduce in Sec. C.4 as Thm. C.3, shows an alternative way to relaxing both assumptions – that $r_{1}\leq 1$ and $p(x)$ is monic.
-Schönhage in [130, Sec. 19] has greatly strengthened the estimate of Thm. B.4 unless “the zeros of $p$ are clustered too much". Namely, he proved (see [130, Eqn. (19.8)]) the following result.
-Lemma B.1.
-Assume that the polynomial $p(x)$ is monic, the values $u_{j}=1$ for all $j$ and complex $v_{1},\dots,v_{d}$ satisfy bound (B.1), and $|z_{j}-v_{j}|\leq r:=2^{2-b/d}$ for all $j$.
+Based on [134, Cor. 4.3] we can also readily relax the assumption that $p(x)$ is monic.
+Furthermore, in view of (24.2), relaxing the bound $r_{1}\leq 1$ would little affect the latter estimate because we can apply map (1.3), which changes the error bound of Problem 0 by at most $\log_{2}(\max\{\rho,1/\rho\})$, for $\rho\geq 1/2^{b}$.
+[136, Thm. 2.7], which we reproduce in Sec. 25.4 as Thm. 25.3, shows an alternative way to relaxing both assumptions – that $r_{1}\leq 1$ and $p(x)$ is monic.
+Schönhage in [134, Sec. 19] has greatly strengthened the estimate of Thm. 24.4 unless “the zeros of $p$ are clustered too much". Namely, he proved (see [134, Eqn. (19.8)]) the following result.
+Lemma 24.1.
+Assume that the polynomial $p(x)$ is monic, the values $u_{j}=1$ for all $j$ and complex $v_{1},\dots,v_{d}$ satisfy bound (24.1), and $|z_{j}-v_{j}|\leq r:=2^{2-b/d}$ for all $j$.
 Furthermore, write $\Delta_{j,1}:=|v_{j}-v_{1}|$ for $j=2,3,\dots,d$ and suppose that $1.25\Delta_{j,1}\leq r$ for $j=2,3,\dots,d$ and $\Delta_{1}:=\min_{j}\Delta_{j,1}>r$.
 Then $|z_{1}-v_{1}|\leq 2^{2-b/m}/|\Delta_{1}-r|^{\frac{d-m}{m}}$.
-If $1/|\Delta-r|^{\frac{d-m}{m}}=2^{\tilde{O}(d)}$, then Lemma B.1 implies that $|z_{1}-v_{1}|\leq 1/2^{b/m-\tilde{O}(d)}$, and we can similarly estimate $|v_{j}-z_{j}|$ for all $j=1,2,\dots,d$.
-By extending these estimates we strengthen Thm. B.4 as follows.
-Corollary B.4.
+If $1/|\Delta-r|^{\frac{d-m}{m}}=2^{\tilde{O}(d)}$, then Lemma 24.1 implies that $|z_{1}-v_{1}|\leq 1/2^{b/m-\tilde{O}(d)}$, and we can similarly estimate $|v_{j}-z_{j}|$ for all $j=1,2,\dots,d$.
+By extending these estimates we strengthen Thm. 24.4 as follows.
+Corollary 24.3.
 Under bounding cluster sizes of Def.
-B.1, it holds that $|z_{j}-v_{j}|\leq 1/2^{b/m-\tilde{O}(d)},~{}j=1,2,\dots,d$.
+24.1, it holds that $|z_{j}-v_{j}|\leq 1/2^{b/m-\tilde{O}(d)},~{}j=1,2,\dots,d$.
 Hence if $m=O(1)$ is a constant, then a solution of Problem 0 within an error bound $1/2^{b}$ can be recovered at a dominated bit operation cost from the solution of Problem F with a precision $1/2^{\bar{b}}$ for a sufficiently large $\bar{b}=\tilde{O}(b+d)$, and so $\mathbb{B}_{1}(b,d)\leq\mathbb{B}_{2}(\tilde{O}(b+d),d)$.
-For converse reduction of Problem F to Problem 0 recall [130, Cor. 4.3] in the case of $k=2$, covering the product of two polynomials:
-Theorem B.5.
+For converse reduction of Problem F to Problem 0 recall [134, Cor. 4.3] in the case of $k=2$, covering the product of two polynomials:
+Theorem 24.5.
 Let $p=gh$ for three polynomials $p$ of (1.1), $g$, and $h$. Then $||g||_{1}||h||_{1}\leq 2^{d-1}||p||_{1}$.
 Now consider the products $p=gh$ for $g=x-z_{j}$, for $j=1,\dots,d$.
-For a fixed $j$, an error of at most $\epsilon/2^{d}$ for the zero $z_{j}$ contributes at most $\epsilon||h||_{1}$ to overall factorization error of Problem F.
+For a fixed $j$, an error of at most $\epsilon/2^{d}$ for the root $z_{j}$ contributes at most $\epsilon||h||_{1}$ to overall factorization error of Problem F.
 Therefore, $||h||_{1}\leq 2^{d-1}||p||_{1}/||x-z_{j}||_{1}\leq 2^{d-1}||p||_{1}$ by virtue of Thm.
-B.5.
-Hence the approximation errors of all the $d$ zeros of $p$ together contribute at most $2^{d-1}d\epsilon$ to the relative error of the solution of Problem F (ignoring dominated impact of higher order terms), and we arrive at
-Corollary B.5.
+24.5.
+Hence the approximation errors of all the $d$ roots together contribute at most $2^{d-1}d\epsilon$ to the relative error of the solution of Problem F (ignoring dominated impact of higher order terms), and we arrive at
+Corollary 24.4.
 If $b=O(b-d-\log(d))$, then Problem F for an error bound $1/2^{b}$ can be reduced at a dominated bit operation cost to Problem 0 for an error bound $1/2^{\widehat{b}}$ and a sufficiently large $\widehat{b}=O(b+d+\log(d))$, and so $\mathbb{B}_{F}(b,d)=O(\mathbb{B}_{0}(b,d))$.
@@
 Then again, the bit operation complexity of this solution of Problem F is optimal (up to polylogarithmic factors) under the random root model provided that $d=\tilde{O}(b)$ (cf.
-Observation B.1).
-B.4 Two root-finders of 2016 and 2018
+Observation 24.1).
+24.4 Two root-finders of 2016 and 2018
 Becker et al in [23, 24] claimed that their extension of subdivision root-finders of [128, 101] (i) runs within the same near optimal bit operation bounds of [96, 103] but (ii) unlike [96, 103] is easy to implement.
 This seemed to diminish incentive for further research in this area, but were the claims valid?
-Implementation of the root-finder of [22] by Imbach et al in [61] provided no empirical support to optimality claimed by Becker et al because the root-finder of [60], using some novel techniques from [108] (see [60, Sec.
+Implementation of the root-finder of [23] by Imbach et al in [64] provided no empirical support to optimality claimed by Becker et al because the root-finder of [63], using some novel techniques from [112] (see [63, Sec.
 1.2]), runs much faster.
@@
 Furthermore, as we pointed out in the introduction, the the bit-complexity specified in the Main Theorem of [24] has order of $mbd^{2}$, that is, exceeds the near optimal bound of [96, 103] by a factor of $d$.
-[…]
-Theorem B.6.
+Can one salvage the claim by Becker et al in another way?
+Next we argue that this can be problematic because the algorithms of [23, 24] for Problem 1m are some variants of subdivision iterations.
+The basic tool and the main algorithmic novelty of the variant of subdivision in [96, 103] are e/i tests based on the following theorem by Pellet [88], which extended the previous results by Cauchy and Rouché:
+Theorem 24.6.
 Given a polynomial $(x)$ of (1.1) with complex coefficients and $p_{g}=1$.
@@
 To apply Pellet’s theorem to $t(x)$ one must compute the values $|\frac{t_{i}}{t_{d}}|$ for $i=0,1,\dots,d-1$.
-Hence, by virtue of Observation B.2, one must process at least $0.5(d-1)db$ bits of the coefficients of $t(x)$ to support approximation of its zeros within $1/2^{b}$.
+Hence, by virtue of Observation 24.2, one must process at least $0.5(d+1)db$ bits of the coefficients of $t(x)$ to support approximation of its zeros within $1/2^{b}$.
 The algorithms of [24] cannot avoid this because they apply e/i tests to the polynomials obtained in $\log(\log(d))$ steps of root-squaring (cf.
 Sec.
-10.3) from polynomials $t_{c,\rho}(x)=p(\frac{x-c}{\rho})$ where $\rho>0$ can be as small as $1/2^{b}$.
-[…]
-Appendix C Recursive factorization of a polynomial and extensions to root-finding and root isolation
+13.3) from polynomials $t_{c,\rho}(x)=p(\frac{x-c}{\rho})$ where $\rho>0$ can be as small as $1/2^{b}$.
+Already a single subdivision iteration can perform more than $4m$ soft e/i tests; this means that one must process more than $2(d+1)dmb$ bits of memory and $(d+1)dmb$ bit-operations to solve Problem 1m by applying such a Pellet-based root-finder.
+25 Recursive Factorization; extensions to Root-finding and Root Isolation
 In this section we briefly recall some algorithms and complexity estimates from [134, 136].
 We first recursively decompose a polynomial into the product of linear factors and then extend the factorization to approximation and isolation of the roots.
-C.1 Auxiliary norm bounds
+25.1 Auxiliary norm bounds
 Recall two simple lemmas.
-Lemma C.1.
+Lemma 25.1.
 It holds that $|u+v|\leq|u|+|v|$ and $|uv|\leq|u|~{}|v|$ for any pair of polynomials $u=u(x)$ and $v=v(x)$.
-Lemma C.2.
+Lemma 25.2.
 Let $u=\sum_{i=0}^{d}u_{i}x^{i}$ and $|u|_{2}=(\sum_{i=0}^{d}|u_{i}|^{2})^{\frac{1}{2}}$. Then $|u|_{2}\leq|u|$.
 The following lemma relates the norms of a polynomial and its factors.
-Lemma C.3.
+Lemma 25.3.
 If $p=p(x)=\prod_{j=1}^{k}f_{j}$ for polynomials $f_{1},\dots,f_{k}$ and $\deg p\leq d$, then $\prod_{j=1}^{k}|f_{j}|\leq 2^{d}|p|_{2}\leq 2^{d}|p|.$
 Proof.
-The leftmost bound was proved by Mignotte in [75]. The rightmost bound follows from Lemma C.2. ∎
-Remark C.1.
-[132, Lemma 2.6] states with no proof a little stronger bound $\prod_{j=1}^{k}|f_{j}|\leq 2^{d-1}|p|$ under the assumptions of Lemma C.3.
+The leftmost bound was proved by Mignotte in [78]. The rightmost bound follows from Lemma 25.2. ∎
+Remark 25.1.
+[136, Lemma 2.6] states with no proof a little stronger bound $\prod_{j=1}^{k}|f_{j}|\leq 2^{d-1}|p|$ under the assumptions of Lemma 25.3.
 From various factors of the polynomial $p(x)=x^{d}-1$ such as $\prod_{j=1}^{\frac{d}{2}}(x-\zeta_{d}^{j})$ for even $d$ and $\zeta:=\zeta_{q}:=\exp(\frac{2\pi{\bf i}}{q})$ denoting a primitive $q$-th root of unity, one can see some limitations on strengthening this bound further.
-C.2 Errors and complexity
+25.2 Errors and complexity
 Suppose that we split a polynomial $p$ into a pair of factors over an isolated circle and recursively apply this algorithm to the factors until they become linear of the form $ux+v$; some or all of them can be repeated.
 Finally, we arrive at complete approximate factorization
-$p\approx p^{\ast}=p^{\ast}(x)=\prod_{j=1}^{d}(u_{j}x+v_{j}).$ (C.1)
-Next, by following [130, Sec. 5], we estimate the norm of the residual polynomial
-$\Delta^{\ast}=p^{\ast}-p.$ (C.2)
+$p\approx p^{\ast}=p^{\ast}(x)=\prod_{j=1}^{d}(u_{j}x+v_{j}).$ (25.1)
+Next, by following [134, Sec. 5], we estimate the norm of the residual polynomial
+$\Delta^{\ast}=p^{\ast}-p.$ (25.2)
 We begin with an auxiliary result.
-Theorem C.1.
+Theorem 25.1.
 Let
-$\Delta_{k}=|p-f_{1}\cdots f_{k}|\leq\frac{k\epsilon|p|}{d},$ (C.3)
-$\Delta=|f_{1}-fg|\leq\epsilon_{k}|f_{1}|,$ (C.4)
+$\Delta_{k}=|p-f_{1}\cdots f_{k}|\leq\frac{k\epsilon|p|}{d},$ (25.3)
+$\Delta=|f_{1}-fg|\leq\epsilon_{k}|f_{1}|,$ (25.4)
 for some non-constant polynomials $f_{1},\ldots,f_{k},f$ and $g$ such that
-$\epsilon_{k}d2^{d}\prod_{j=1}^{k}|f_{j}|\leq\epsilon.$ (C.5)
+$\epsilon_{k}d2^{d}\prod_{j=1}^{k}|f_{j}|\leq\epsilon.$ (25.5)
 Then
-$|\Delta_{k+1}|=|p-fgf_{2}\cdots f_{k}|\leq\frac{(k+1)\epsilon|p|}{d}.$ (C.6)
+$|\Delta_{k+1}|=|p-fgf_{2}\cdots f_{k}|\leq\frac{(k+1)\epsilon|p|}{d}.$ (25.6)
 Proof.
-$\Delta_{k+1}=|p-f_{1}\cdots f_{k}+(f_{1}-fg)f_{2}\cdots f_{k}|.$ Apply Lemma C.1 and deduce that $\Delta_{k+1}\leq\Delta_{k}+|(f_{1}-fg)f_{2}\cdots f_{k}|$ and, furthermore, that
+$\Delta_{k+1}=|p-f_{1}\cdots f_{k}+(f_{1}-fg)f_{2}\cdots f_{k}|.$ Apply Lemma 25.1 and deduce that $\Delta_{k+1}\leq\Delta_{k}+|(f_{1}-fg)f_{2}\cdots f_{k}|$ and, furthermore, that
 $|(f_{1}-fg)f_{2}\cdots f_{k}|\leq|f_{1}-fg|~{}|f_{2}\cdots f_{k}|=\Delta|f_{2}\cdots f_{k}|.$
-Combine the latter inequalities and obtain $\Delta_{k+1}\leq\Delta_{k}+\Delta|f_{2}\cdots f_{k}|$. Combine this bound with (C.3)-(C.5) and Lemmas C.1 and C.3 and obtain (C.6). ∎
-Write $f_{1}:=f$ and $f_{k+1}=g.$ Then (C.6) turns into (C.3) for $k$ replaced by $k+1.$ Now compute one of the factors $f_{j}$ as in (C.4), apply Thm.
-C.1, then recursively continue splitting the polynomial $p$ into factors of smaller degrees, and finally arrive at factorization (C.1) with
+Combine the latter inequalities and obtain $\Delta_{k+1}\leq\Delta_{k}+\Delta|f_{2}\cdots f_{k}|$. Combine this bound with (25.3)-(25.5) and Lemmas 25.1 and 25.3 and obtain (25.6). ∎
+Write $f_{1}:=f$ and $f_{k+1}=g.$ Then (25.6) turns into (25.3) for $k$ replaced by $k+1.$ Now compute one of the factors $f_{j}$ as in (25.4), apply Thm.
+25.1, then recursively continue splitting the polynomial $p$ into factors of smaller degrees, and finally arrive at factorization (25.1) with
 $|\Delta^{\ast}|\leq\epsilon|p|$
-for $\Delta^{*}$ of (C.2). Let us call this computation Recursive Splitting Process provided that it begins with $k=1$ and $f_{1}=p$ and ends with $k=d$.
-Theorem C.2.
-In order to support (C.3) for $0<\epsilon\leq 1$ and $j=1,2,\ldots,d$ in the Recursive Splitting Process it is sufficient to choose $\epsilon_{k}$ in (C.4) satisfying
-$\epsilon_{k}~{}d~{}2^{2d+1}\leq\epsilon~{}\textrm{for~{}all}~{}k.$ (C.7)
+for $\Delta^{*}$ of (25.2). Let us call this computation Recursive Splitting Process provided that it begins with $k=1$ and $f_{1}=p$ and ends with $k=d$.
+Theorem 25.2.
+In order to support (25.3) for $0<\epsilon\leq 1$ and $j=1,2,\ldots,d$ in the Recursive Splitting Process it is sufficient to choose $\epsilon_{k}$ in (25.4) satisfying
+$\epsilon_{k}~{}d~{}2^{2d+1}\leq\epsilon~{}\textrm{for~{}all}~{}k.$ (25.7)
 Proof.
-Prove bound (C.3) by induction on $j$. Clearly, the bound holds for $k=1$. It remains to deduce (C.6) from (C.3) and (C.7) for any $k$. By first applying Lemma C.3 and then bound (C.3), obtain
+Prove bound (25.3) by induction on $j$. Clearly, the bound holds for $k=1$. It remains to deduce (25.6) from (25.3) and (25.7) for any $k$. By first applying Lemma 25.3 and then bound (25.3), obtain
 $\prod_{i=1}^{k}|f_{i}|\leq 2^{d}|\prod_{i=1}^{k}f_{i}|\leq 2^{d}~{}\Big{(}1+\frac{k\epsilon}{d}\Big{)}~{}|p|.$
-The latter bound cannot exceed $2^{d+1}|{p}|$ for $k\leq d$, $\epsilon\leq 1$. Consequently (C.7) ensures (C.5), and then (C.6) follows by virtue of Thm. C.1.
-Remark C.2.
+The latter bound cannot exceed $2^{d+1}|{p}|$ for $k\leq d$, $\epsilon\leq 1$. Consequently (25.7) ensures (25.5), and then (25.6) follows by virtue of Thm. 25.1.
+Remark 25.2.
 The theorem implies that we can ensure the output precision bound $b^{\prime}$ by working with the precision of
@@
 bits throughout Recursive Splitting Process.
-C.3 Overall complexity of recursive factorization and extension to Root Isolation
+25.3 Overall complexity of Recursive Factorization and extension to Root Isolation
 The overall complexity of recursive factorization is dominated by the sum of the bounds on the complexity of all splittings into pairs of factors.
@@
 Consequently, $O(\log(d)$ recursive splitting steps are sufficient in [103], and so the overall complexity of root-finding is proportional to that of the first splitting.
-In contrast, the degrees of a factor in any splittin in [130] can have low, e.g., 1, and so there can be allows up to $d-1$ splittings overall, with each of the first $0.5d$ of them almost as expensive as the very first one.
-C.4 From factors to roots
-Theorem C.3.
-[132, Thm. 2.7]. Suppose that
+In contrast, the degrees of a factor in any splitting
+in [134] can have low, e.g., 1, and so there can be allows up to $d-1$ splittings overall, with each of the first $0.5d$ of them almost as expensive as the very first one.
+25.4 From factors to roots
+Theorem 25.3.
+[136, Thm. 2.7]. Suppose that
 $p=p_{d}\prod_{j=1}^{d}(x-z_{j}),~{}\tilde{p}=\tilde{p}_{d}\prod_{j=1}^{d}(x-y_{j}),~{}|\tilde{p}-p|\leq\epsilon|p|,~{}\epsilon\leq 2^{-7d}$
@@
 By virtue of Thm.
-C.3 for $b^{\prime}=O(bd)$ we can bound the bit operation complexity of the solution of Problem 0 by increasing the estimate for the complexity of factorization in Sec.
-C.3 by a factor of $d$.
-Corollary C.1.
+25.3 for $b^{\prime}=O(bd)$ we can bound the bit operation complexity of the solution of Problem 0 by increasing the estimate for the complexity of factorization in Sec.
+25.3 by a factor of $d$.
+Corollary 25.1.
 The bit operation complexity of the solution of Problem 0.
-Given a polynomial $p$ of degree $d$ and a real $b^{\prime}\geq 7d$, one can approximate all zeros of that polynomial within the error bound $2^{-b^{\prime}}$ by using $O(\mu(b^{\prime}+d)d^{2}\log(d)(\log^{2}d+\log(b^{\prime})))=\tilde{O}((b^{\prime}+d)d^{2})$ bit operations.
-Remark C.3.
+Given a polynomial $p$ of degree $d$ and a real $b^{\prime}\geq 7d$, one can approximate all its $d$ zeros within the error bound $2^{-b^{\prime}}$ by using $O(\mu(b^{\prime}+d)d^{2}\log(d)(\log^{2}d+\log(b^{\prime})))=\tilde{O}((b^{\prime}+d)d^{2})$ bit operations.
+Remark 25.3.
 The study in this section covers the solution of Problem 1 for the worst case input polynomial $p$.
 If its roots are isolated from each other, then the upper estimates of Thm.
-C.3 and Cor.
-C.2 can be decreased by a factor of $d$ (cf.
+25.3 and Cor.
+25.2 can be decreased by a factor of $d$ (cf.
 [103]).
-C.5 Polynomial Root Isolation
-Based on his results recalled in this appendix, Schönhage in [130, Sec. 20] estimates the bit operation complexity of the following well-known problem by reducing it to Problem F.
+25.5 Polynomial Root Isolation
+Based on his results recalled in this section, Schönhage in [134, Sec. 20] estimates the bit operation complexity of the following well-known problem by reducing it to Problem F.
 Problem ISOL.
@@
 Given a polynomial $p$ of (1.1) that has integer coefficients and is square-free, that is, has only simple roots, compute $d$ disjoint discs on the complex plane, each containing exactly one root.
-Remark C.4.
-Our Problem $*{}^{*}$ can be viewed as the Problem of Polynomial Root Cluster Isolation.
-It turns into Problem ISOL if all the coefficients of $p$ are integers and the roots are simple, and so our algorithms for Problem $*{}^{*}$ as well as the Algorithms of [92, 99] solve Problem ISOL within a record complexity bounds.
-In [130, Sec.
+Remark 25.4.
+Our Problem 0 can be viewed as the Problem of Polynomial Root Cluster Isolation.
+It turns into Problem ISOL if all the coefficients of $p$ are integers and the roots are simple, and so our algorithms for Problem 0 as well as the Algorithms of [96, 103] solve Problem ISOL within a record complexity bounds.
+In [134, Sec.
 20] Schönhage reduces to factorization Problem F the root-isolation problem of computing $d$ disjoint discs each covering a single simple root of a square-free polynomial with integer coefficients.
 Problem ISOL is readily solved if the zeros are approximated within $\epsilon=1/2^{b}<0.5\delta$ for $\delta$ denoting the minimal pairwise distance between the zeros of $p$ (see the known lower bounds on $\delta$ in Kurt Mahler and Maurice Mignotte [77, 78, 79]), and Schönhage has estimated that solution of Problem F (e.g., by means of the algorithms of [134, 94, 103]) implies isolation provided that
-$b=O((\tau+\log(d))+d)~{}{\rm and}~{}\tau=\log_{2}\max_{i=0}^{d}|p_{i}|.$ (C.8)
+$b=O((\tau+\log(d))+d)~{}{\rm and}~{}\tau=\log_{2}\max_{i=0}^{d}|p_{i}|.$ (25.8)
 This implied the following result.
-Corollary C.2.
+Corollary 25.2.
 The bit operation complexity of polynomial root isolation.
@@
 Then one can solve Problem ISOL by using $\tilde{O}(bdm)$ bit operations for $b=\frac{b^{\prime}}{m}$.
-Appendix D Newton’s, Weierstrass’s, and Ehrlich’s iterations
-D.1 Contents
-In the next subsection we comment on convergence of the three listed iterative root-finders and on their stopping criterion.
+26 Newton’s, Weierstrass’s, and Ehrlich’s iterations
+26.1 Contents
+In the next subsection we recall Newton’s, Weierstrass’s, and Ehrlich’s iterations, comment on their convergence, and acceleration by means of applying fast multipoint evaluation, which is dramatic but involves the coefficients of $p$.
 In Sec.
-D.3 we discuss the policy of doubling precision, the resulting dynamic partition of the roots into tame and wild, and the deflation of a wild factor.
-In Sec.
-D.4 we accelerate these iterations by incorporating FMM.
-In Sec.
-D.5 we compute a disc that contains all zeros of a black box polynomial $p$, which one needs in customary initialization of root-finding for a black box polynomial by means of Newton’s iterations (cf.
-[138]).
-Customary initialization is more demanding for Weierstrass’s and Ehrlich’s iterations (see Sec.
-D.6).
-D.2 Definition
-1. Definition. Consider Newton’s, Weierstrass’s (aka Durand-Kerner’s), and Ehrlich’s (aka Aberth’s) iterations [68, 55, 126, 138, 143, 29, 64, 36, 3, 21]):
-$z\leftarrow z-N(z),~{}z_{j}\leftarrow z_{j}-W(z_{j}),~{}z_{j}\leftarrow z_{j}-E_{j}(z_{j}),~{}j=1,\dots,d,~{}{\rm where}$ (D.1)
-$N(x)=W(x)=E_{j}(x)=0~{}{\rm for~{}all}~{}j~{}{\rm if}~{}p(x)=0$ (D.2)
+26.3 we cover mre limited acceleration of these iterations by means of incorporating FMM, which can be applied to a black box polynomial.
+In Secs.
+26.5 and 27) we cover initialization of Newton’s, Weierstrass’s, and Ehrlich’s iterations, required for their fast global empirical convergence.
+For Newton’s iterations we only need to compute a disc that contains all zeros of a polynomial $p$ (cf.
+[71, 58, 143, 130, 141]).
+This is given by the well-known estimates via the coefficients, but we present simple but novel recipes in the case of black box polynomial $p$.
+The customary initialization is more demanding for Weierstrass’s and Ehrlich’s iterations but is still fast for polynomials represented with their coefficients, as we recall in Sec.
+27), where we also point out how this initialization can be extended to fast real and nearly real root-finding.
+It is still a challenge to extend such an initialization to the case of a black box polynomial.
+26.2 Definitions
+1. Definitions. Consider Newton’s, Weierstrass’s (aka Durand-Kerner’s), and Ehrlich’s (aka Aberth’s) iterations [71, 58, 130, 143, 148, 31, 67, 38, 3, 22]):
+$z\leftarrow z-N(z),~{}z_{j}\leftarrow z_{j}-W(z_{j}),~{}z_{j}\leftarrow z_{j}-E_{j}(z_{j}),~{}j=1,\dots,d,~{}{\rm where}$ (26.1)
+$N(x)=W(x)=E_{j}(x)=0~{}{\rm for~{}all}~{}j~{}{\rm if}~{}p(x)=0$ (26.2)
 and otherwise
-$N(z):=\frac{p(z)}{p^{\prime}(z)},~{}W(x):=\frac{p(x)}{l^{\prime}(x)},~{}\frac{1}{E_{j}(x)}:=\frac{1}{N(x)}-\sum_{i=1,i\neq j}^{d}\frac{1}{x-z_{i}}$ (D.3)
+$N(z):=\frac{p(z)}{p^{\prime}(z)},~{}W(x):=\frac{p(x)}{l^{\prime}(x)},~{}\frac{1}{E_{j}(x)}:=\frac{1}{N(x)}-\sum_{i=1,i\neq j}^{d}\frac{1}{x-z_{i}}$ (26.3)
 for $l(x):=\prod_{j=1}^{d}(x-z_{j})$.
-Newton’s iterations can be initiated at a single point $z$ or at many points and can approximate a single root, many or all $d$ roots at the cost decreasing as the number of roots decreases [68, 55, 126, 138].
+Newton’s iterations can be initiated at a single point $z$ or at many points and can approximate a single root, many or all $d$ roots at the cost decreasing as the number of roots decreases [71, 58, 143, 141, 85, 130].
 In contrast, Weierstrass’s and Ehrlich’s iterations begin at $d$ points for a $d$th degree polynomial and approximate a subset of its $d$ roots about as fast and almost as slow as they approximate all its $d$ roots.
@@
 [^37] Formal support for such global convergence is much weaker for Newton’s iterations (see [71, 58, 143], and the references therein) and is completely missing for Weierstrass’s and Ehrlich’s iterations.
-D.3 Doubling precision and the deflation of a wild factor
-Computational (working) precision sufficient in order to control the relative error of an output approximation to a root can be defined by root’s condition number (see the relevant discussion in [14, 21]).
-This number is not readily available, however, and instead one can recursively double precision of computing, until a fixed output accuracy is verified: first apply a fixed root-finding iterations with the IEEE double precision of 53 bits and then recursively double it as long as this is feasible and enables more accurate approximation of the roots, stopping where the required output accuracy is ensured.
-This policy for root-finding can be traced back to [14], where it was proposed for MPSolve.
-It turned out to be highly efficient and was adopted in [41] for Eigensolve, in [21] for the next version of MPSolve, and in [126, 138] for approximation of all roots by means of Newton’s iterations.
-According to recipe (D.2) the computed approximation $z$ is only updated until $p(z)=0$ or, in numerical implementation, until the roots become tame, for a fixed stopping criterion.
-In the latter case, one should restrict updating to the wild roots (cf.
-Sec.
-1.7) and should deflate the wild factor, if it has small degree.
-When a root is tamed the minimal working precision is exceeded by less than a factor of 2.
-Recall from [130] and [132, Thm.
-2.7] that working precision does not need to exceed the output precision $b^{\prime}$ by more than a factor of $d$, and so at most $O(\log(db^{\prime}))$ steps of doubling precision are sufficient.
-For every working precision the $d$ roots of $p(x)$ are partitioned into $d-w$ tame and $w$ wild roots as above; then one can apply explicit or implicit deflation (cf. (1.14)).
-[…]
-D.4 Acceleration with Fast Multipole Method (FMM)
+All expressions in (26.3) for the functions $N(x)$, $W(x)$, and $E_{j}(x)$ are already presented or can be rewritten as ratios $\frac{u(x)}{v(x)}$ for pairs of polynomials $u(x)$ and $v(x)$, whose coefficients are given or readily available at low cost if the polynomial $p(x)$ is gien with its coeffcients (see [102, Problems 3.6.1 and 3.6.3]).
+Then multipoint evaluation of such a function amounts to multipoint evaluation of $u(x)$ and $v(x)$ and division of the computed values, and one can accelerate Newton’s, Weierstrass’s, and Ehrlich’s iterations dramatically by incorporating the algorithms of [81, 59].
+26.3 Acceleration with Fast Multipole Method (FMM)
+The algorithms of [81, 59] deeply involve the coefficients of $p(x)$.
+Next we cover acceleration of Newton’s, Weierstrass’s, and Ehrlich’s iterations applied to black box polynomial root-finding.
 Given $3d$ complex values $w_{j}$, $z_{i}$, $v_{j}$ such that $z_{i}\neq v_{j}$ for all $i,j=1,\dots,d$, Fast Multipole Method (FMM) by Greengard and Rokhlin [50] and by Carrier, Greengard, and Rokhlin [28] computes the sums $\sum_{j=1}^{d}\frac{w_{j}}{x_{i}-z_{j}}$ as well as $\sum_{j=1,j\neq i}^{d}\frac{w_{j}}{z_{i}-z_{j}}$ for all $i$ by using $O(db)$ arithmetic operations, with the precision of order $b$, versus $2d^{2}$ arithmetic operations in the straightforward algorithm.
@@
 $A=D+{\bf v}{\bf e}^{T}~{}{\rm for}~{}D=\operatorname{diag}(z_{j})_{j=1}^{d},~{}{\bf e}=(1)_{j=1}^{d},~{}~{}{\bf w}=(w_{j})_{j=1}^{d},~{}w_{j}=W(z_{j}),$
-$W(x)$ of (D.3) and all $j$ (cf. [15]). Let $I_{d}$ denote the $d\times d$ identity matrix, apply the Sherman-Morrison-Woodbury formula, and obtain
+$W(x)$ of (26.3) and all $j$ (cf. [16]). Let $I_{d}$ denote the $d\times d$ identity matrix, apply the Sherman-Morrison-Woodbury formula, and obtain
 $xI_{d}-A=(xI_{d}-D)(I_{d}-(xI_{d}-D)^{-1}{\bf w}{\bf e}^{T}).$
 Hence
-$\frac{1}{p_{d}}p(x)=\det(xI_{d}-A)=-\sum_{i=1}^{d}w_{i}\prod_{j\neq i}(x-z_{j})+\prod_{j=1}^{d}(x-z_{j}).$ (D.4)
+$\frac{1}{p_{d}}p(x)=\det(xI_{d}-A)=-\sum_{i=1}^{d}w_{i}\prod_{j\neq i}(x-z_{j})+\prod_{j=1}^{d}(x-z_{j}).$ (26.4)
 Indeed, the monic polynomials of degree $d$ on both sides have the same values for $x=z_{j}$, $j=1,\dots,d$ (cf.
@@
 [22]).
-Equations (D.4) imply that
+Equations (26.4) imply that
 $p(z_{i})=-w_{i}\prod_{j\neq i}(z_{i}-z_{j}),~{}p^{\prime}(z_{i})=\prod_{j\neq i}(z_{i}-z_{j})\Big{(}1-\sum_{j\neq i}\frac{w_{i}+w_{j}}{z_{i}-z_{j}}\Big{)},$
 and therefore
-$\frac{w_{i}}{N(z_{i})}=\sum_{j\neq i}\frac{w_{i}+w_{j}}{z_{i}-z_{j}}-1.$ (D.5)
-Hence, given the $2d$ values of $z_{j}$ and $w_{j}=W(z_{j})$ for all $j$, one can fast approximate $N(z_{j})$ and consequently $E_{j}(z_{j})$ for all $j$ (cf. (D.3)) by applying FMM.
-Every iteration but the first one begins with some values $z_{j}$, $w_{j}$, and updated values $z_{j}^{\rm(new)}$ of (D.1), for $j=1,\dots,d$, and outputs updated values $w_{j}^{\rm(new)}=W(z_{j}^{\rm(new)})$ for all $j$.
-Then equations (D.3) and (D.5) define $N(z_{j}^{\rm(new)})$ and $E_{j}(z_{j}^{\rm(new)})$ for all $j$.
-To update the values $w_{j}$, apply [21, equation (17)]:
+$\frac{w_{i}}{N(z_{i})}=\sum_{j\neq i}\frac{w_{i}+w_{j}}{z_{i}-z_{j}}-1.$ (26.5)
+Hence, given the $2d$ values of $z_{j}$ and $w_{j}=W(z_{j})$ for all $j$, one can fast approximate $N(z_{j})$ and consequently $E_{j}(z_{j})$ for all $j$ (cf. (26.3)) by applying FMM.
+Every iteration but the first one begins with some values $z_{j}$, $w_{j}$, and updated values $z_{j}^{\rm(new)}$ of (26.1), for $j=1,\dots,d$, and outputs updated values $w_{j}^{\rm(new)}=W(z_{j}^{\rm(new)})$ for all $j$.
+Then equations (26.3) and (26.5) define $N(z_{j}^{\rm(new)})$ and $E_{j}(z_{j}^{\rm(new)})$ for all $j$.
+To update the values $w_{j}$, apply [22, equation (17)]:
 $w_{i}^{\rm new}=(z_{i}^{\rm new}-z_{i})\prod_{j\neq i}\frac{z_{i}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}\Big{(}\sum_{k=1}^{d}\frac{w_{k}}{z_{i}^{\rm new}-z_{k}}-1\Big{)},~{}i=1,\dots,d.$
@@
 fast in two ways.
-(i) Use the following customary trick (see [30, Remark 3]): approximate the values $\ln(R_{i})$ for $i=1,\dots,d$ fast by applying FMM and then readily approximate $R_{i}=\exp(\ln(R_{i}))$ for all $i$.
+(i) Use the following customary trick (see [32, Remark 3]): approximate the values $\ln(R_{i})$ for $i=1,\dots,d$ fast by applying FMM and then readily approximate $R_{i}=\exp(\ln(R_{i}))$ for all $i$.
 [32] proves numerical stability of the reduction to FMM of a similar albeit a little more involved rational expressions.
@@
 The same expression (3.3) defines Ehrlich’s iterations for both $f(x)$ and $p(x)$ by virtue of Thm.
-10.1.
+13.1.
 Newton’s iterations to a wild factor one additionally subtract the sum $\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}$ (cf.
@@
 To apply Weierstrass’s iterations to a wild factor one performs additional division by $t(x)$ and decreases the degree of the divisor $l^{\prime}(x)$ from $d-1$ to $w-1$.
-D.5 Computation of a disc covering all zeros of a black box polynomial
-D.5.1 Introductory comments
+26.4 MPSolve and its competition with Newton’s iterations
+Since its creation in 2000 the package MPSolve by Bini and Fiorentino [15], revised in 2014 by Bini and Robol [22], has been remaining user’s choice for approximation of all $d$ complex roots of a polynomial $p$.
+It implements Ehrlich’s iterations of 1967 [38], aka Aberth’s of 1973 [3], which empirically converge very fast globally – right from the start, under proper initialization, involving the coefficients of $p$.
+Such convergence has been consistently observed in decades-long worldwide applications of these iterations as well as in the case of some other root-finders by means of functional iterations, notably Weierstrass’s iterations [148], aka Durand and Kerner’s [31, 67]).
+Unsuccessful attempts of supporting this empirical behavior formally (e.g., see [4, 54]), however, have recently ended with surprising proofs that Weierstrass’s and Ehrlich’s iterations diverge in the case of inputs that lie on some complex manifolds, being an open domain in the case of Weierstrass’s iterations (see Reinke, Schleicher, and M.
+Stoll [131, 129]).
+Serious competition to MPSolve came in 2001 from the package EigenSolve by Fortune [43], but the later version of MPSolve [22] combines the benefits of both packages.
+An implementations of Newton’s iterations for approximation of all roots by Schleicher and R.
+Stoll in [143] competed with MPSolve in 2017, but then MPSolve has regained upper hand by means of incorporation of Fast Multipole Method (FMM), even though such implementation is simpler and can be more efficient for Newton’s than Ehrlich’s iterations (see the previous subsection).
+The competition seems to resume in 2024 (see [12, 130, 85]).
+26.5 Computation of a disc covering all zeros of a black box polynomial
+26.5.1 Introductory comments
 Given coefficients of a polynomial $p$ of a degree $d$, we can readily compute its largest and smallest root radii $r_{1}$ and $r_{d}$, respectively (cf.
@@
 Hereafter let $p(x)$ denote a black box polynomial.
-D.5.2 Probabilistic solution
+26.5.2 Probabilistic solution
 Seek a disc $D(c,\rho)$ containing no zeros of $p(x)$ in its interior, where we can choose any positive $\rho$ Sample a random $c$ in the unit disc $D(0,1)$ under the uniform probability distribution in $D(0,1)$.
@@
 By applying our e/i tests of Secs. 14.414.3 to this disc, we can verify correctness of its choice.
-D.5.3 Deterministic solution
+26.5.3 Deterministic solution
 We can deterministically compute a desired disc $D(0,\rho_{-})$ at a higher cost, based on a soft e/i test.
 […]
-D.6 $d$ root radii and initialization
-The algorithms of [68, 55, 138] approximate all $d$ roots of a polynomial $p$ by applying Newton’s iterations at $f(d)$ equally spaced points on a large circle containing inside all $d$ roots for $f(d)$ of order at least $d\log(d)$.
-Bounds (10.6) immediately define such a circle for a polynomial given by its coefficients.
-MPSolve further elaborates upon initialization.
-It first applies Bini’s heuristic recipe of [10], reused in [14, 21], for the approximation of the $d$ root radii by involving all coefficients of $p$.
+27 Fast approximation of the $d$ root radii with applications
+The customary initialization recipes of Weierstrass"s and Ehrlich’s iterations involve the $d$ root radii at the origin and consequently require to use the coefficients of $p$.
+To initialize Ehrlich’s iterations MPSolve first applies Bini’s heuristic recipe of [10], reused in [15, 22], for the approximation of the $d$ root radii.
 Given such approximations MPSolve defines $k$ concentric annuli $A(c,\rho_{i},\rho_{i}^{\prime})$ and $k$ integers $m_{i}$ such that #$\#(A(c,\rho_{i},\rho_{i}^{\prime}))=m_{i}$ for $i=1,\dots,k$, $k\leq d$, and $\sum_{i=1}^{k}m_{i}=d$.
-MPSolve initializes Ehrlich’s iterations at $d$ points $z_{j}$, $j=1,2,\dots,d$, exactly $m_{i}$ of them being equally spaced on the circle $C(c,(\rho_{i}+\rho_{i}^{\prime})/2)$.
-This heuristic initialization recipe turns out to be highly efficient, and one is challenged to test whether it can also help accelerate convergence of Newton’s iterations.
+Then MPSolve initializes Ehrlich’s iterations at $d$ points $z_{j}$, $j=1,2,\dots,d$, exactly $m_{i}$ of them being equally spaced on the circle $C(c,(\rho_{i}+\rho_{i}^{\prime})/2)$.
+This heuristic initialization rec;ipe turns out to be highly efficient.
 The algorithm supporting the next theorem also approximates the $d$ root radii but is certified, deterministic, and faster than Bini’s.
-It was first outlined and briefly analyzed in [130] for fast approximation of a single root radius and then extended in [88, Sec.
+It was first outlined and briefly analyzed in [134] for fast approximation of a single root radius and then extended in [92, Sec.
 4] (cf.
-also [89, Sec.
+also [93, Sec.
 5]) to approximation of all root radii of a polynomial.
@@
 A variation of this algorithm in [62] supports the following theorem.
-Theorem D.1.
-[59, Prop.
+Theorem 27.1.
+[62, Prop.
 3].
@@
 [62] extends the supporting algorithm to approximation of all root radii within a positive relative error bound $\Delta$ by performing $\lceil\log_{2}\log_{1+\Delta}(2d)\rceil$ DLG root-squaring iterations (3.7).
-Combine Thm. D.1 with estimates for the Boolean cost of performing these iterations implicit in the proof of [130, Cor. 14.3] and obtain
-Corollary D.1.
+Combine Thm. 27.1 with estimates for the Boolean cost of performing these iterations implicit in the proof of [134, Cor. 14.3] and obtain
+Corollary 27.1.
 Given the coefficients of a polynomial $p=p(x)$ and a positive $\Delta$, one can approximate all the $d$ root radii $|x_{1}|,\dots,|x_{d}|$ within a relative error bound $\Delta$ in $\frac{1}{d^{O(1)}}$ at a Boolean cost in $O(d\log(||p||_{\infty})+d^{2}\log^{2}(d))$.
 The algorithm supporting the corollary and said to be Alg.
-D.1a computes at a low cost $d$ suspect annuli $A(0,r_{j,-},r_{j,+})$ for $j=1,\dots,d$ (some of them may pairwise overlap), whose union $\mathbb{U}$ contains all $d$ roots.
+27.1a computes at a low cost $d$ suspect annuli $A(0,r_{j,-},r_{j,+})$ for $j=1,\dots,d$ (some of them may pairwise overlap), whose union $\mathbb{U}$ contains all $d$ roots.
 The algorithm was the point of departure for amazing progress in [59], which extends the novel ingenious algorithm of [81] for multipoint polynomial evaluation.
 We conclude this subsection with a list of some other benefits of fast approximation of the $d$ root radii.
-(i) In a subdivision step one can skip testing exclusion and discard a square unless it intersects $\mathbb{U}$.
-(ii) The search for real roots can be narrowed to the intersection $\mathbb{I}$ of $\mathbb{U}$ with the real axis (see Fig.
-10).
-[122] has tested a simple real root-finder that approximates the roots in $\mathbb{I}$.
+(i) In a subdivision step one can skip testing exclusion and discard a square unless it intersects the set $\mathbb{U}$.
+(ii) The search for real roots can be narrowed to the intersection $\mathbb{I}$ of $\mathbb{U}$ with the real axis, which consists of $2m$ segments of small length (see Fig.
+11).
+[126] has tested a simple real root-finder that approximates the roots in $\mathbb{I}$ by initiating Newton’s iterations at the centers of the segments.
 According to these tests the algorithm is highly efficient for the approximation of all isolated real roots.
-Clearly, the well-developed root-finding techniques, e.g., ones using derivatives and higher order derivatives (see [99, 62], and the references therein) can further strengthen these algorithms.
-By complementing the root radii approach with some advanced novel techniques, the paper [59] has noticeably accelerated the real root-finder [67], currently of user’s choice.
-One can extend the progress to fast approximation of near real roots by combining root radii computation with the techniques of Sec.
+One can extend Newton’s iterations at the centers of the segments by applying also our root-radii algorithm of Sec.
 9.
+By complementing the root radii approach with some advanced novel techniques, the paper [62] has noticeably accelerated the real root-finder [70], which is currently user’s choice.
 (iii) Initialization based on Alg.
-D.1a should improve global convergence of Ehrlich’s, Weierstrass’s, and other functional iterations for root-finding as well as of Kirrinnis’s factorization algorithm of [65].
-In particular Ehrlich’s iterations of MPSolve are celebrated for their very fast global empirical convergence (right from the start), but it has only been observed with such initialization; moreover, they diverge without it in an example of [125].
+27.1a should improve global convergence of Ehrlich’s, Weierstrass’s, and other functional iterations for root-finding as well as of Kirrinnis’s factorization algorithm of [68].
 […]

### v37

@@
-Fast Polynomial Root-Finding and Approximation of the Eigenvalues of a Matrix
+Polynomial Root-Finding and Algebraic Eigenvalue Problem
 Abstract
 Univariate polynomial root-finding has been studied for four millennia and very intensively in the last decades.
-Our new near-optimal root-finders approximate all zeros of a polynomial $p$ almost as fast as one accesses its coefficients with the precision required for the solution within a prescribed error bound.
-Furthermore, our root-finders can be applied to a black box polynomial, defined by an oracle (black box subroutine) for its evaluation rather than by its coefficients.
-Due to this feature our root-finders (i) support approximation of the eigenvalues of a matrix in a record Las Vegas expected bit operation time and (ii) are particularly fast for a polynomial that can be evaluated fast such as the sum of a few shifted monomials of the form $(x-c)^{d}$ or a Mandelbrot-like polynomial defined by a recurrence.
-Our divide and conquer algorithm of ACM STOC 1995 is the only other known near-optimal polynomial root-finder, but it extensively uses the coefficients, is quite involved, and has never been implemented, while according to extensive numerical experiments with standard test polynomials, already a slower initial implementation of our new root-finders competes with user’s choice package of root-finding subroutines MPSolve and supersedes it more and more significantly as the degree of a polynomial grows large.
-Our auxiliary algorithms and techniques for computations with black box polynomials can be of independent interest.
+Our new near-optimal root-finders approximate all complex zeros of a polynomial almost as fast as one accesses its coefficients with the precision required for the solution within a prescribed error bound.
+The root-finders work even for a black box polynomial, defined by an oracle (black box subroutine) for its evaluation rather than by its coefficients.
+Hence they (i) support approximation of the eigenvalues of a matrix in a record Las Vegas expected bit operation time and (ii) are particularly fast for polynomials that can be evaluated fast such as the sum of a few shifted monomials of the form $(x-c)^{d}$ or a Mandelbrot-like polynomial defined by a recurrence.
+Our divide and conquer algorithm of ACM STOC 1995 is the only other known near-optimal polynomial root-finder, but it extensively uses the coefficients, is quite involved, and has never been implemented, while according to extensive numerical experiments with standard test polynomials already a slower initial implementation of our new root-finders competes with user’s choice package of root-finding subroutines MPSolve and supersedes it more and more significantly as the degree of a polynomial grows large.
+We elaborate upon design and analysis of our algorithms, comment on their potential heuristic acceleration, and briefly cover polynomial root-finding by means of functional iterations.
+Our techniques can be of independent interest.
 Keywords:
@@
 65H05, 26C10, 30C15
-PART I: BACKGROUND AND OUR PROGRESS
+PART I: State of the Art and Our Progress (Outline)
 1 State of the Art and our Main Results
-In this part of the paper we state our main results and outline our algorithms. We present them in some detail in Parts II, III, and IV, which can be read independently of each other.
 1.1 Polynomial Root-finding: Problems, Impact, and Earlier Studies
@@
 [105]), and the first rigorous proof is due to an amateurish mathematician Argand (1806, revised in 1813).
-Already about 500 BC it was proved in Pythagorean school that the roots of the equation $x^{2}=2$ were irrational radicals, while even with the help of radicals one cannot express the roots rationally if the degree $d$ of $p(x)$ exceeds 4, according to Ruffini (1799,1813) and Abel (1824); moreover, one cannot do this already for specific equations such as $x^{5}-x-1=0$ or $x^{5}-4x-2$, according to Galois theory of 1830, 1832.
-Nevertheless, in 1924 in [149], Herman Weyl published his constructive proof of the Fundamental Theorem of Algebra where he rationally approximated within any fixed positive error tolerance all roots of $p(x)=0$ lying in a fixed square in the complex plane, that is, all roots if the square is large enough.
-This was the solution of the following Problem 0 of modern polynomial root-finding:
+Already about 500 BC it was proved in Pythagorean school that the roots of the equation $x^{2}=2$ were irrational radicals, while even with the help of radicals one cannot express the roots rationally if the degree $d$ of $p(x)$ exceeds 4, according to Ruffini (1799,1813) and Abel (1824); moreover, one cannot do this already for specific equations such as $x^{5}-x-1=0$ or $x^{5}-4x-2=0$, according to Galois theory of 1830, 1832.
+Nevertheless, in his constructive proof of the Fundamental Theorem of Algebra in [120], Herman Weyl has rationally approximated within any fixed positive error tolerance all roots of $p(x)=0$ lying in a fixed square in the complex plane, that is, all roots if the square is large enough.
+This solved the following Problem 0 of modern polynomial root-finding:
 INPUT: $d+1$ complex coefficients $p_{0},p_{1},\dots,p_{d}$ of a polynomial
 $p=p(x):=\sum_{i=0}^{d}p_{i}x^{i}:=p_{d}\prod_{j=1}^{d}(x-z_{j}),$ (1.1)
-a positive error tolerance $\epsilon=1/2^{b}>0$, an integer $m$, $1\leq m\leq d$, and a bounded domain $\mathbb{D}$ in the complex plane, e.g., a disc or a rectangle, such that $z_{j}\in\mathbb{D}$ if and only if $j\leq m$.
-OUTPUT: approximations $\tilde{z}_{j}\approx z_{j}$, $j=1,\dots,m$, to the $m$ zeros of $p(x)$ (aka roots of $p(x)=0$) [^1] Hereafter we freely say just “root” or “zero” for short, always count them with their multiplicities, and do not distinguish multiple roots from pairwise isolated root clusters of diameters less than $\epsilon$.
+a positive error tolerance $\epsilon=1/2^{b}>0$, and a bounded domain $\mathbb{D}$ in the complex plane, e.g., a disc or a rectangle.
+OUTPUT: approximations within $\epsilon$ to every zero of $p(x)$ (aka every root of $p(x)=0$) [^1] Hereafter we freely say just “root” or “zero” for short, always count them with their multiplicities, and do not distinguish multiple roots from pairwise isolated root clusters of diameters less than $\epsilon$.
 lying in the domain $\mathbb{D}$ such that
-$|\tilde{z}_{j}-z_{j}|\leq\epsilon~{}{\rm for}~{}j=1,\dots,m.$ (1.2)
+$|\tilde{z}_{j}-z_{j}|\leq\epsilon~{}{\rm if}~{}z_{j}\in\mathbb{D}.$ (1.2)
 Example 1.1.
-[Approximation of all roots.] Fix $R\geq\max_{j}|z_{j}|$ (see [52] for simple expressions for such $R$) and let $\mathbb{D}$ denote the disc $D(0,R):=\{x:~{}|x\leq R|\}$, centered in the origin and having radius $R$.
-Then Problem 0 requires us to approximate all $d$ roots.
+[Approximation of all $d$ roots.] For a real
+$R\geq\max_{j=1}^{d}\{|x_{j}|\}$ (1.3)
+(see [39] for simple expressions for such bound $R$) and $\mathbb{D}$ denoting the disc $D(0,R):=\{x:~{}|x|\leq R\}$, Problem 0 amounts to approximation of all $d$ roots $z_{j}$ within $\epsilon$.
 Example 1.2.
-[Numerical approximation of real roots.] Root-finding in the real line or its segment (real root-finding) is highly important because in many applications, e.g., to optimization in algebraic geometry and geometric modeling, only real roots of a polynomial are of interest, and typically they are much less numerous than all $d$ complex roots (cf.
-Kac [66], Edelman and Kostlan [39]).
-More generally consider Problem 0 where $\mathbb{D}$ denotes the minimal rectangle in the complex plane that covers the $\epsilon/\sqrt{2}$-neighborhood of a fixed real line segment.
-To solve this problem we should approximate both real and near-real roots, that is, their perturbations due to small input and rounding errors.
-The known real root-finders (cf.
-[80, Chs.
-2 and 3]) do not address this important issue.
-Major motivation for modern study of Problem 0 comes from application to approximation of the eigenvalues of a matrix (see our Secs.
-1.3 and 1.5) and the solution of a multivariate polynomial system of equations [29], both problems being highly important for the theory and practice of modern computations.
-[82, Introduction] specifies some applications of polynomial root-finding to computer algebra, signal processing, geometric modeling, control, and financial mathematics.
+[Numerical real root-finding.] Root-finding in the real line or its segment (real root-finding) is highly important because in many applications, e.g., to optimization in algebraic geometry and geometric modeling, only real roots of a polynomial are of interest, and typically they are much less numerous than all $d$ complex roots [50, 29].
+Now consider Problem 0 where $\mathbb{D}$ denotes the minimal rectangle in the complex plane that covers the $\epsilon$-neighborhood of a fixed real line segment.
+This task addresses approximation of both real and near-real roots, that is, their perturbations due to small input and rounding errors.
+The known real polynomial root-finders (cf.
+[63, Chs.
+2 and 3]) ignore this important issue.
+Solution of Problem 0 can be extended to the eigenvalue problem for matrices and matrix polynomials (see our Secs.
+1.3 and 1.5 and Part (C) of Sec.
+5.4) and is involved into the solution of a multivariate polynomial system of equations [21]; both of them are much studied areas of modern computations.
+[65, Introduction] specifies some applications of polynomial root-finding to computer algebra, signal processing, geometric modeling, control, and financial mathematics.
 Furthermore, Problem 0 for a domain $\mathbb{D}$ containing all roots is closely linked to polynomial factorization (see Secs.
-24 and 25), which has applications to time series analysis, Wiener filtering, noise variance estimation, co-variance matrix computation, and the study of multichannel systems [150, 18, 8, 35, 36, 147].
-Hundreds of efficient root-finders have been proposed [80, 82] and keep appearing (see [122, 22, 40, 70, 76, 127, 108, 23, 24, 64, 110, 111, 75, 60, 62, 81, 63, 59, 115], and the bibliography therein).
-In [133] Smale proposed to study Problem 0 in terms of arithmetic complexity and computational precision and later, jointly with Blum, Cucker, and Shub (cf.
-[14, 30]), linked this study to rounding errors and conditioning of the roots.
-In [134] Schönhage proposed to estimate the complexity of polynomial root-finders in terms of the number of bit operations involved and presented an algorithm that solved Problem 0 for a domain that contains all roots by using $\tilde{O}(bd^{3})$ bit operations, [^2] Here and hereafter we assume that $b\geq d$, recalling that error amplification factors $2^{cd}$ are inevitable in polynomial root-finding, e.g., in routine shift of the variable.
-that is, $bd^{3}$ bit operations up to a polylogarithmic factor in $bd$.
-His work prompted intensive study of root-finding by researchers in the Theory of Computing in the 1980s and 1990s, culminated at ACM STOC 1995 with our near-optimal divide and conquer root-finder of [96] (also see its journal versions [97, 103] and surveys [99, 100]); it supports bound $\tilde{O}(bd^{2})$ and showed the information lower bound $0.25(d+1)bd$ (see Cor.
+22 and 23), which has applications to time series analysis, Wiener filtering, noise variance estimation, co-variance matrix computation, and the study of multichannel systems [121, 11, 5, 26, 27, 118].
+Hundreds of efficient polynomial root-finders have been proposed [63, 65] and keep appearing (see [97, 15, 30, 54, 59, 99, 89, 16, 17, 49, 91, 92, 58, 45, 47, 64, 48, 44, 96], and the bibliography therein).
+In [105] Smale proposed to study Problem 0 in terms of its arithmetic complexity and computational precision.
+Together they would define its bit operation complexity, but Smale deduced no complexity estimates.
+[106] Schönhage did this in [106].
+His much involved root-finder, presented with terse complexity analysis on 46 pages, approximates all $d$ roots by using $\tilde{O}(bd^{3})$ bit operations, that is, $bd^{3}$ bit operations up to a factor in $bd$ [^2] Error amplification factors $2^{cd}$ are inevitable in polynomial root-finding, e.g., in routine shift of the variable, and assume that $b\geq d$ throughout our paper; alternatively one can replace $b$ with $b+d$ in our upper bounds on the computational cost.
+His work prompted intensive study of root-finding by researchers in the Theory of Computing in the 1980s and 1990s, culminated at ACM STOC 1995 with our near-optimal divide and conquer root-finder of [78] (also see its journal versions [79, 84] and surveys [80, 81]).
+Both upper bound in $\tilde{O}(bd^{2})$ and lower bound of $0.25(d+1)bd$ bit operations of [78, 84] decrease by a factor of $d$ for polynomials of Def.
+22.1 of Sec.
+22, whose all zeros and tiny clusters of zeros are pairwise sufficiently well isolated (see Thms.
+22.1 and 22.4 and Cor.
 22.4 in Sec.
-24.2).
-If we only consider polynomials whose zeros and tiny clusters of zeros are pairwise sufficiently isolated, as we specify in Def.
-24.1 of Sec.
-24, then the lower bounds on the computational precision and the bit operation complexity of root-finding decrease by a factor of $d$ (see Sec.
-24), and so do the upper bounds supported by the algorithm of [96, 97, 103] as well; hence the algorithm remains near-optimal (see Thms.
-24.1 and 24.4 and Cor.
-24.3 in Sec.
-24).
+22).
 1.2 Reduction to the case of the unit disc
-Hereafter T shows transposition of a vector or a matrix, $\mathbb{R}^{k}$ and $\mathbb{C}^{k}$ denote the spaces of the column vectors $(v_{1},\dots,v_{k})^{T}$ of length $k$ with real and complex coordinates, respectively, and $D(c,\rho):=\{x:~{}|x-c|\leq\rho\}$ and $C(c,\rho):=\{x:~{}|x-c|=\rho\}$ denote the disc of radius $\rho>0$ centered at $c\in\mathbb{C}$ and its boundary circle, respectively.
-Furthermore, let Problem $\widehat{0}$ denote Problem 0 where the domain $\mathbb{D}$ is the unit disc $D(0,1)$; we can extend its solution to root-finding in the disc $D(c,\rho)$ for any pair of complex $c$ and positive $\rho$ by means of applying the linear map
+Hereafter $D(c,\rho):=\{x:~{}|x-c|\leq\rho\}$ and $C(c,\rho):=\{x:~{}|x-c|=\rho\}$ denote the disc of radius $\rho>0$ centered at a complex point $c$ and its boundary circle, respectively, and Problem $\widehat{0}$ denotes Problem 0 where the domain $\mathbb{D}$ is the unit disc $D(0,1)$; we can extend its solution to root-finding in the disc $D(c,\rho)$ for any pair of complex $c$ and positive $\rho$ by means of applying the linear map
 $x\mapsto\frac{x-c}{\rho},~{}p(x)\mapsto t(x)=p\Big{(}\frac{x-c}{\rho}\Big{)},~{}D(c,\rho)\mapsto D(0,1),$ (1.4)
@@
 Next we reduce Problem 0 for any domain $\mathbb{D}$ to Problem $\widehat{0}$.
-(i) Two ways from the roots in the unit disc to all roots. Given a value
-$R\geq\max_{j=1}^{d}\{|x_{j}|\}$ (1.4)
-(see [52] for simple expressions for such $R$), we can scale the variable $x\mapsto y=\frac{x}{R}$, $p(x)\mapsto p(y)=p(\frac{x}{R})$ and approximate every root $x_{y}$, for $j=1,\dots,d$, within $\epsilon R$ by solving Problem $\widehat{0}$ for $p(y)$.
-Schönhage in [134] reduced approximation of all $d$ roots to Problem $\widehat{0}$ without assuming (1.4) or using (1.3) as follows.
-Given an algorithm $\mathcal{A}$ that solves Problem $\widehat{0}$, compute also approximations $\tilde{y}_{j}$ within $\epsilon$ to the reciprocals $y_{j}=1/z_{j}$ of all roots lying outside the disc $D(0,1)$ – simply by applying this algorithm to the reverse polynomial
+(i) Two ways from the roots in the unit disc to all roots.
+Given a value $R$ of (1.3), we can scale the variable, $x\mapsto y=\frac{x}{R}$, $p(x)\mapsto p(y)=p(\frac{x}{R})$, and then approximate every root $x_{j}$, for $j=1,\dots,d$, within $\epsilon R$ by solving Problem $\widehat{0}$ for $p(y)$.
+Here is Schönhage’s alternative reduction in [106] of Problem 0 for all $d$ roots to Problem $\widehat{0}$: Given an algorithm $\mathcal{A}$ that solves Problem $\widehat{0}$, compute also approximations $\tilde{y}_{j}$ within $\epsilon$ to the reciprocals $y_{j}=1/z_{j}$ of all roots lying outside the disc $D(0,1)$ – simply by applying this algorithm to the reverse polynomial
 $p_{\rm rev}(x):=x^{d}p\Big{(}\frac{1}{x}\Big{)}=\sum_{i=0}^{d}p_{i}x^{d-i},~{}p_{\rm rev}(x)=p_{0}\prod_{j=1}^{d}\Big{(}x-\frac{1}{z_{j}}\Big{)}~{}{\rm if}~{}p_{0}\neq 0.$ (1.5)
@@
 (ii) From roots in the unit disc to roots in a bounded domain.
-Given a bounded domain $\mathbb{D}$ on the complex plane, we can successively (a) cover it with a minimal disc $D(c,\rho)$, (b) map this disc into the disc $D(0,1)$, (c) solve Problem $\widehat{0}$ for the resulting polynomial $t(x)$ of (1.3), and finally, (d) use the map (1.3) again – this time to approximate all roots lying in the intersection of the disc $D(c,\rho)$ with the domain $\mathbb{D}$.
-In this way we also approximate extraneous roots lying in $D(c,\rho)$ but not in $\mathbb{D}$; in Sec.
-2.2 we control the computational cost of that additional work.
-1.3 Black box root-finders; the isolation of roots, eigenvalues, and their clusters
-A nontrivial variant of Newton’s iterations by Louis and Vempala [76], at FOCS 2016, approximates within a fixed positive $\epsilon\leq 1/2$ an absolutely largest zero of a real-rooted polynomial $p$ of a degree $d$ at the cost of its evaluation with a precision in $O((b+\log(R))\log(R~{}d)))$ at $O((b+\log(R))\log(d))$ points for $b:=\log_{2}(R/\epsilon)$ and a fixed $R$ satisfying (1.4).
-While the previous polynomial root-finders have been devised for polynomials given in the monomial basis, that is, with their coefficients, the algorithm of [76] never involves them and works for a black box polynomial $p$ – given with an oracle (black box subroutine) for its evaluation.
-[76] shows benefits of extension of such algorithms to approximation of the eigenvalues of a matrix.
+Given a bounded domain $\mathbb{D}$ on the complex plane, we can successively (a) cover it with a minimal disc $D(c,\rho)$, (b) map this disc into the disc $D(0,1)$, (c) solve Problem $\widehat{0}$ for the resulting polynomial $t(x)$ of (1.4), and finally (d) use the map (1.4) again – this time to approximate all roots lying in the intersection of the disc $D(c,\rho)$ with the domain $\mathbb{D}$.
+In this way we also approximate extraneous roots lying in $D(c,\rho)$ but not in $\mathbb{D}$.
+In Sec.
+2.1 we curtail such an extraneous work (see Prop.
+2.1 and Remark 2.1).
+1.3 Black box root-finders and approximation of matrix eigenvalues
+A nontrivial variant of Newton’s iterations by Louis and Vempala at FOCS 2016 [59] approximates within a fixed positive $\epsilon\leq 1/2$ an absolutely largest zero of a real-rooted polynomial $p$ of a degree $d$ at the cost of its evaluation with a precision in $O((b+\log(R))\log(R~{}d)))$ at $O((b+\log(R))\log(d))$ points for $b:=\log_{2}(R/\epsilon)$ and a fixed $R$ satisfying (1.3).
+While the previous polynomial root-finders have been devised for polynomials given in the monomial basis, that is, with their coefficients, Louis and Vempala in [59] never involve them, working for a black box polynomial $p$ – given with an oracle (black box subroutine) for its evaluation, and then extend their root-finder to efficient approximation of the eigenvalues of a matrix.
 Next we state the relevant tasks.
 Problem 1 (Black Box Polynomial Root Approximation) is Problem 0 for a black box polynomial $p$.
-Problem 2: Approximation of the Eigenvalues of a Matrix. [^3] As in [76], the problem covers or can be extended to the case where an input is a polynomial matrix, aka matrix polynomial.
-INPUT: $\epsilon=1/2^{b}>0$, integers $m$ and $d$ such that $0<m\leq d$, and a $d\times d$ matrix $M$, and a complex domain $\mathbb{D}$ that contains precisely $m$ eigenvalues of $M$ counted with their multiplicities.
-OUTPUT: Approximations within $\epsilon$ to each of all $m$ eigenvalues of $M$ lying in the domain $\mathbb{D}$.
+Problem 2: Approximation of the Eigenvalues of a Matrix. [^3] As in [59], the problem covers or can be extended from a matrix to a polynomial matrix, aka matrix polynomial.
+INPUT: $\epsilon=1/2^{\tilde{b}}>0$, an integer $d>0$, a $d\times d$ matrix $M$, and a domain $\mathbb{D}$ on the complex plane.
+OUTPUT: Approximations within $\epsilon$ to every eigenvalue of $M$ lying in the domain $\mathbb{D}$.
 The output of Problems 0, 1, and 2 may include multiple approximations within $\epsilon$ to the same root or eigenvalue, but the next modifications exclude such a confusion.
-Problems $0^{*}$, $1^{*}$, $2^{*}$: Isolation of Clusters of the Zeros of a polynomial and the Eigenvalues of a matrix.
-Under the assumptions of Problems 0, 1, and 2 compute at most $d$ disjoint discs of radii at most $\epsilon$ (hereafter $\epsilon$-discs) covering all roots or eigenvalues that lie in the domain $\mathbb{D}$; also compute the number of the roots or eigenvalues lying in each $\epsilon$-disc.
+Problems $0^{*}$, $1^{*}$, $2^{*}$: Isolation of Clusters of Roots and Eigenvalues.
+Under the assumptions of Problems 0, 1, and 2 compute at most $d$ disjoint discs of radii at most $\epsilon$ (hereafter $\epsilon$-discs) that cover all roots or eigenvalues lying in the domain $\mathbb{D}$; also compute the number of the roots or eigenvalues lying in each $\epsilon$-disc.
 Remark 1.1.
@@
 If a polynomial $p$ is square-free (having only simple zeros) and has integer coefficients, then Problem 0 turns into the well-known Problem of Root-isolation (see Problem ISOL and Remark 23.4 in Sec.
-25.5).
+23.5).
 These comments can be immediately extended to Problems 1 and 2.
@@
 1.4.1 Why we should use them
+Namely, the algorithm of Fiduccia [32] and Moenck and Borodin [67] (see [12, Sec.
+4.5] for historical account) evaluates a polynomial of degree $m$ at $m$ points in $O(m\log^{2}(m))$ ops if its coefficients are given, versus $2m^{2}$ ops involved in Horner’s algorithms.
+Numerically stable algorithms of [64, 44] yield such acceleration also in terms of bit operation cost.
 One can perform black box evaluation at $q>d$ points $x_{1},\dots,x_{q}$ and then interpolate to $p$.
 The bit operation cost of interpolation, however, is unbounded unless the $q$ points are well isolated from the roots, but it is a challenge to yield such isolation.
-Furthermore, even if the isolation is ensured, the cost of interpolation exceeds the cost of the solution of Problems 0, 0, 1 ,1, 2, and 2 in the cases where domain $\mathbb{D}$ contains only a small number of roots.
-Moreover, here are three examples of important benefits of using black box algorithms even where an input polynomial is given with its coefficients.
+Furthermore, even if the isolation is ensured, the cost of interpolation exceeds the cost of the solution of Problems 0, 0, 1, 1, 2, and 2 in the cases where domain $\mathbb{D}$ contains only a small number of roots as well as where a polynomial $p$ can be evaluated fast because root-finding for a polynomial is routinely reduced to its evaluation at sufficiently many points [63, 65].
+For example, one only needs $O(\log(d))$ ops for evaluation of a Mandelbrot-like polynomial $p:=M_{k}(x)$, where $M_{0}:=x,~{}M_{i+1}(x):=xM_{i}^{2}(x)+1$, $i=0,1,\dots,k$, $d=2^{k}-1$, or the sum of a bounded number of shifted monomials, such as $p:=\alpha(x-a)^{d}+\beta(x-b)^{d}+\gamma(x-c)^{d}$ for six constants $a$, $b$, $c$, $\alpha$, $\beta$, and $\gamma$.
+For comparison, if we compute the coefficients of such a polynomial and then use them to evaluate it at a point $x$ by applying Horner’s algorithm, then we would need $2d$ ops, and this bound is optimal (see [72, 53] or [112, Section “Pan’s method"]).
+In computational practice numerical stability of polynomial evaluation is highly important, and it is ensured when a polynomial is represented in the Bernstein basis [57].
+Transition to it from the monomial basis is costly but is not needed where a root-finder works for a black box polynomial.
+One can benefit from black box root-finders even where an input polynomial is given with its coefficients.
 (i) Application of a black box root-finder to the characteristic polynomial $p(x)=\det(xI-M)$ of a $d\times d$ matrix $M$ only involves the values of this polynomial but not its coefficients, whose overall size exceeds the size (a norm) of the matrix $M$ by a factor of $d$.
-Consequently, extension of such a root-finder to Problem 2 saves a factor of $d$ in bit precision and bit operation count versus the known solutions of Problems 2 and 2, based on [134, Sec.
+Hence extension of such a root-finder to Problem 2 saves a factor of $d$ in bit precision and bit operation count versus the known solutions of Problems 2 and 2, based on [106, Sec.
 21].
@@
 Black box polynomial root-finders keep all benefits of using those maps but avoid the latter notorious problems by involving no coefficients of $t(x)$ and $p_{h}(x)$ of these maps.
-(iii) Black box polynomial root-finders are particularly efficient where a $d$th degree polynomial $p$ can be evaluated fast, say, in $O(\log(d))$ arithmetic operations (hereafter (ops) such as a Mandelbrot-like polynomial $p:=M_{k}(x)$, where $M_{0}:=x,~{}M_{i+1}(x):=xM_{i}^{2}(x)+1$, $i=0,1,\dots,k$, $d=2^{k}-1$, or the sum of a bounded number of shifted monomials, such as $p:=\alpha(x-a)^{d}+\beta(x-b)^{d}+\gamma(x-c)^{d}$ for six constants $a$, $b$, $c$, $\alpha$, $\beta$, and $\gamma$.
-For comparison, if we compute the coefficients of such a polynomial and then use them to evaluate it at a point $x$ by applying Horner’s algorithm, then we would need $2d$ ops, and this bound is optimal (see [89, 69] or [139, Section “Pan’s method"]).
-1.4.2 What we adopt and what cannot adopt for black box polynomial root-finding
+(iii) Deflation of a factor of a polynomial is a well-known cause of the coefficient swell, e.g., for degree 300 factors of $x^{1000}+1$, which blows up precision and bit complexity of computations.
+For an exception, one can readily deflate a small degree factor $f(x)$ and approximate its zeros, but the coefficients of the high degree factor $g(x)=p(x)/f(x)$ still tend to blow up.
+We, however, avoid their computation when we apply a black box root-finder to the polynomial $g(x)$ represented by the pair of $p(x)$ and $f(x)$.
+1.4.2 Progress at FOCS 2016 and its limitations
 The known polynomial root-finders, except [59], extensively use the coefficients of $p$, and we must be careful in adopting their techniques for Problems 1 and 1.
-On the other hand, the root-finder of [76] cannot help solve even Problem 1.
-Indeed, repeatedly deflating $p$, by peeling off large roots, can accumulate errors that grow too quickly, together with the coefficient length of the factors (notice, e.g., the swell of the coefficients of $p(x)=x^{d}+1$ in transition to its factors of degree $d/2$, say).
+Besides proposing novel root-finder for a black box polynomial with extension to approximation of an eigenvalue of a matrix, Louis and Vempala in [59] presented non-trivial techniques for bounding precision of root-finding computations.
+Extension of these techniques helped us simplify dramatically the estimation of the bit operation complexity of our root-finders compared to those of [106, 78, 84, 52], which bypass estimation of computational precision.
+On the other hand, direct extension of the root-finder of [59] cannot help solve even Problem 1.
+Indeed, repeatedly deflating $p$, by peeling off large roots, can accumulate errors that grow too quickly, together with the coefficient length of the factors: notice, e.g., the swell of the coefficients of $p(x)=x^{d}+1$ in transition to its factors of degree $d/2$, say.
 This is true even if one performs deflation in the black box setting implicitly, based on Eqn.
@@
 3.
-Moreover, even in that setting the coefficient swell in deflation is translated into swelling of the values of the high order derivatives involved in the algorithm of [76].
-Even if the algorithm of [76] worked for Problem 1 it would only solve the well-known and much easier problem of root-finding for a real-rooted polynomial (cf.
+Indeed, the high order derivatives involved in the algorithm of [59] are proportional to the coefficients and together with them swell in deflation.
+Even if deflation succeeded, the algorithm of [59] would have only solved the well-known and much easier problem of root-finding for a real-rooted polynomial (cf.
 [75, 18, 25]).
-Moreover, the resulting algorithm for that much easier problem would involve more than $b^{2}d^{2}$ bit operations, which would by a factor of $b$ exceed the upper bound of [96, 103] for general polynomial root-finding.
-Recursive combination of [76] with the Inverse Power Iterations [48] could lead to approximation of all eigenvalues of a symmetric (although not unsymmetric) matrix, together with its eigenvectors.
-One can be challenged to estimate the increased complexity of this extended computations, but we study other approaches.
-By applying a root-finder to Given a polynomials $p$ of (1.1) given with its coefficients We accelerate known algorithms for Problem 0 by a factor of $m/\log^{2}(m)$ by incorporating fast multipoint polynomial evaluation algorithms of [42, 84, 68, 81], but we could not use them if we successively approximated zeros of $p$ or matrix eigenvalues by using deflation.
-Thus, instead of applying high order Newton’s iterations of [76] and deflation, we recall and extend subdivision iterations of [149, 52, 128, 101].
+Moreover, the resulting algorithm for that much easier problem would involve more than $b^{2}d^{2}$ bit operations, exceeding by a factor of $b$ the upper bound of [78, 84] for general polynomial root-finding.
+Furthermore, we accelerate the solution of Problem 0 by a factor of $m/\log^{2}(m)$ by incorporating fast multipoint polynomial evaluation of [32, 67, 52, 64], but we would lose this chance if we successively approximated the zeros of $p$ or matrix eigenvalues by using deflation.
+Thus, instead of applying high order Newton’s iterations of [59] and deflation, we recall and advance subdivision iterations of [120, 39, 100, 82].
 1.5 Our complexity estimates
+Unlike the root-finders of [106, 78, 84, 52], seeking all $d$ roots, of [59], seeking a single absolutely largest root and of [78, 84], seeking all $d$ roots or all roots in a disc well isolated from the external roots, we approximate the roots in any domain $\mathbb{D}$ in the complex plane.
+Root approximation within $\epsilon$ in a domain $\mathbb{D}$ involves the roots in its $\epsilon$-neighborhood; our complexity estimates grows with the number of roots in a little larger but very small neighborhood of $\mathbb{D}$.
 Definition 1.1.
-$m=m(\mathbb{D},\beta)$ for $\beta>0$ is the number of roots of polynomial $p(x)$ in the $\beta$-neighborhood of a domain $\mathbb{D}$ on the complex plane.
-Main Theorems 1-4 below hold for $m$ of Def. 1.1, any fixed constant $\beta>0$, and either $p(x)=\det(xI-M)$ in the case of Problems 2 and 2 or $p(x)$ of (1.1) otherwise.
+For a polynomial $p=p(x)$, a domain $\mathbb{D}$ in the complex plane having a diameter $\Delta$, and a positive $\beta$ such that $\beta\Delta>2\epsilon$, write $m=m(p,\mathbb{D},\beta)$ if precisely $m$ roots of $p(x)$ lie in the $\beta\Delta$-neighborhood of $\mathbb{D}$.
+We will prove our Main Theorems 1-4 below for $m$ of Def.
+1.1, any fixed constant $\beta>2\epsilon/\Delta$, and either $p(x)=\det(xI-M)$ in the case of Problems 2 and 2 or $p(x)$ of (1.1) otherwise.
+[^4] Hereafter, whenever we say that subdivision iterations are applied to a input domain $\mathbb{D}$ of diameter $\Delta$ that contains $m$ roots, we mean that so does its $\beta$-neighborhood.
 Theorem 1.1.
@@
 Theorem 1.2.
-(Main Theorem 2.) One can solve Problems 2 and 2 at Las Vegas randomized bit operation cost
+(Main Theorem 2.) One can solve Problems 2 and 2 at a Las Vegas randomized bit operation cost
 $\mathbb{B}_{\rm eigen}=\tilde{O}(\tilde{b}d^{\omega}m^{2})$ (1.6)
 for $\tilde{b}:=\log_{2}(||M||_{F}/\epsilon)$, the Frobenius norm $||M||_{F}$ of a matrix $M$, and $\omega$ denoting an exponent of feasible or unfeasible matrix multiplication (MM).
-[^4] The current records are about 2.7734 for exponents of feasible MM [90, 107, 109, 72, 57], unbeaten since 1982, and about 2.372 for exponents of unfeasible MM [6, 37].
+[^5] The current records are about 2.7734 for exponents of feasible MM [73, 88, 90, 56, 42], unbeaten since 1982, and about 2.372 for exponents of unfeasible MM [122].
 The exponents below 2.77 have only been reached at the expense of blowing up the size of MM by immense factors.
 Theorem 1.3.
-(Main Theorem 3.) One can solve Problems 0 and 0 at Las Vegas randomized bit operation cost $\tilde{O}(bdm))$ for a polynomial $p$ of (1.1) normalized such that $||p(x)||_{1}:=\sum_{i=1}^{d}|p_{i}|=1$.
-Random Root Model: The zeros of $p$ (in particular for $p$ denoting the characteristic polynomial of a matrix $M$) are independent identically distributed (iid) random variables sampled under the uniform probability distribution on the disc $D(0,R)$ for a large $R\in\mathbb{R}$ such that the value $\frac{R^{2}}{d}-1$ exceeds a positive constant.555(i) The assumption of the model is a little stronger than that of our Def.
-24.1, supporting such a decrease in [96, 103].
-(ii) Random coefficient models may seem more natural for polynomial root-finding but are useless because no known probability distribution withstands shift of the variable $x$.
+(Main Theorem 3.) One can solve Problems 0 and 0 at a Las Vegas randomized bit operation cost $\tilde{O}(bdm))$ for a polynomial $p$ of (1.1) normalized such that $\sum_{i=1}^{d}|p_{i}|=1$.
+Random Root Model: The zeros of $p(x)$, for $p(x)=\det(xI-M)$ in Problems 2 and 2 and for $p(x)$ of (1.1) otherwise, are independent identically distributed (iid) random variables sampled under the uniform probability distribution on the disc $D(0,R)$ for a large real $R$ such that the value $\frac{R^{2}}{d}-1$ exceeds a positive constant.666(i) The assumption of the model is a little stronger than that of our Def.
+22.1, supporting such a decrease in [78, 84].
+(ii) Random coefficient models fails at the basic step of the shift of the variable $x$, which no known probability distribution withstands.
 Theorem 1.4.
-(Main Theorem 4) The upper bounds of Thms. 1.1-1.3 on computational precision and bit operation complexity can be decreased by a factor of $d$ under the Random Root Model.
-In view of Remarks 1.1 and 25.4 we can readily extend our Main Theorems to Polynomial Root Isolation.
-2 Accelerated subdivision polynomial root-finders: overview and reduction to soft exclusion/inclusion tests
-For a polynomial $p$ given in monomial basis (with its coefficients) the deterministic algorithm of [96, 103] supports our Main Theorems 3 and 4, but not our Main Theorems 1 and 2.
-This is because the algorithm of [96, 103] cannot be applied to a black box polynomial, unlike our present Las Vegas randomized algorithms, which we outline next and elaborate upon in Parts II–IV.
-Namely, in the next two sections we reduce our root-finders to recursive application of exclusion/inclusion (e/i) test, which decides whether a fixed complex disc contains roots or not.
-In Secs.
-3.3.2 and 3.4 we outline our two approaches to devising e/i tests for a black box polynomial and unit disc $\mathcal{D}=D(0,1)$.
-We cover these two approaches in some detail in Part II, made up of Secs.
-612, and Part III, made up of Secs.
-1318, respectively, and describe yet another fast e/i test in Sec.
-23.4.
-Our root-finders of Parts II and III incorporate these e/i tests and are quite fast but still only support root approximation (rather than isolation) and the complexity estimates exceeding those of our Main Theorems by a factor of $b$.
+(Main Theorem 4.) Under the Random Root Model one can decrease by a factor of $m$ the upper bounds of Thms. 1.1-1.3 on the computational precision and bit operation complexity.
+In view of Remarks 1.1 and 23.4 we can readily extend our Main Theorems to Polynomial Root Isolation.
+2 Fast subdivision root-finding: overview and soft e/i tests
+We support Main Theorems 1 – 4 with our accelerated subdivision root-finders.
+They are reduced to recursive application of soft exclusion/inclusion (e/i) tests, which decide whether a fixed complex disc or its small neighborhood contains roots.
+In Sec.
+3 we outline our two approaches to devising e/i tests for a black box polynomial and the unit disc $\mathcal{D}=D(0,1)$.
+We expand these two outlines in Part II, made up of Secs.
+610, and Part III, made up of Secs.
+1115, respectively.
+Our root-finders of Parts II and III only approximate (rather than isolate) the roots and support complexity estimates that are inferior to those of our Main Theorems by a factor of $b$.
 In Part IV, made up of Secs.
 1621 and outlined in Sec.
-4, we fix both deficiencies, thus proving our Main Theorems.
+4, we fix both deficiencies, thus proving the Main Theorems.
 In Sec.
-5.4 we list some tentative directions to further progress.
+5.5 we further comment about the contents of our paper, including its Part V, made up of Secs.
+2225.
 2.1 Subdivision root-finders and reduction to e/i tests
-We depart from the popular subdivision root-finders, traced back to [149, 52, 128, 101].
+We depart from popular subdivision root-finders, traced back to [120, 39, 100, 82].
 They extend root-finding in a line segment by means of bisection to polynomial root-finding in a square on the complex plane and have been extensively used in Computational Geometry under the name of Quad-tree Construction.
 Next we outline them.
-(A) First let a square on the complex plane contain precisely $m>0$ roots (with multiplicities) and let us approximate them.
-Call the square suspect, recursively divide every suspect square into four congruent sub-squares, and to each of them apply exclusion/inclusion (e/i) test in order to decide whether the square contains a root or does not.
+(A) First consider root-finding in a square on the complex plane.
+Call that square suspect, recursively divide it and every other suspect square into four congruent sub-squares, and to each of them apply exclusion/inclusion (e/i) test to decide whether the square contains a root or does not.
 If it does, call the square suspect and process it in the next iteration; otherwise discard it (see Fig.
 1a).
-In Sec.
-9 and then in our Part IV we use a natural extension of e/i tests to $m$-tests, which decide whether a disc in the complex plane covering at most $m$ roots does or does not contain $m$ roots.
+In Def.
+2.1 we generalize e/i tests to $\ell$-tests, which decide whether a disc in the complex plane contains $\ell$ or fewer roots provided that it contains at most $m$ roots.
+E/i test is a 1-tests, and whenever we need $\ell$-tests for $\ell>1$ (e.g., in Sec.
+16 and Part IV), we reduce them to e/i tests (see Sec.
+18 and 13.4).
 (a) Four roots marked by asterisks lie in suspect sub-squares; the other (empty) sub-squares are discarded.
@@
 Observation 2.1.
-A subdivision iteration applied to a square with $m$ roots (i) decreases by twice the diameter of a suspect square and (ii) processes at most $4m$ suspect squares, (iii) whose centers approximate every root within the distance of one half of their diameter.
+A subdivision iteration applied to a square with $m$ roots in its $\beta$-neighborhood (cf.
+Def.
+1.1 and Remark 2.1) (i) decreases by twice the diameter of a suspect square and (ii) processes at most $4m$ suspect squares, (iii) whose centers approximate every root within the distance of one half of their diameter.
 Proof.
@@
 Corollary 2.1.
-Let subdivision iterations be applied to a square $S$ with side length $\Delta$ and diameter $\Delta\sqrt{2}$ containing precisely $m$ roots.
-Then (i) the $k$-th iteration outputs at most $4m$ suspect squares whose centers approximate within $\epsilon:=\Delta/2^{k-0.5}$ all the $m$ roots lying in $S$, so that $k\leq\lceil\log_{2}(\frac{\Delta}{2\epsilon}\sqrt{2})\rceil$, and (ii) the first $k$ iterations together involve at most $4mk$ e/i tests.
-(B) Seeking roots in any fixed domain $\mathbb{D}$ in the complex plane, we apply subdivision iterations to its minimal covering square $S$ and while performing them also discard all squares that do not intersect the domain $\mathbb{D}$, even if they contain roots.
-Then the $(\Delta_{k}\sqrt{2})$-neighborhood of $\mathbb{D}$ covers the union $\mathbb{U}_{k}$ of all suspect squares defined after the $k$th subdivision iteration, while $\Delta_{k}=\Delta_{0}/2^{k}$ fast converges to 0 as $k$ increases.
-Hence the neighborhood fast converges to the domain $\mathbb{D}$: the width of the neighborhood decreases by a factor of $2^{k}$ in $k$ subdivision iterations, still at the price of performing at most $4km$ e/i tests.
+Let subdivision iterations be applied to a square $S$ with side length $\Delta$ and diameter $\Delta^{\prime}:=\Delta\sqrt{2}$ containing precisely $m$ roots.
+Then (i) the $k$-th iteration outputs at most $4m$ suspect squares whose centers approximate within $\epsilon:=\Delta^{\prime}/2^{k+1}$ all the $m$ roots lying in $S$, so that $k+1\leq\lceil\log_{2}(\frac{\Delta^{\prime}}{\epsilon})\rceil$, and (ii) the first $k$ iterations together involve at most $4mk$ e/i tests.
+(B) Seeking roots in any fixed complex domain $\mathbb{D}$, we apply subdivision iterations to its minimal covering square $S$ with diameter $\Delta^{\prime}$ and while performing an iteration also discard all squares not intersecting the domain $\mathbb{D}$, even if they contain roots.
+The following proposition bounds the overall number of e/i tests performed in $k$ iterations.
+Proposition 2.1.
+Let $k$ subdivision iterations be applied to a complex domain $\mathbb{D}$, let precisely $m_{i}$ roots lie in the $\Delta_{i}^{\prime}$-neighborhood of $\mathbb{D}$ for $\Delta^{\prime}_{i}$ denoting the diameter of a suspect square at the $i$-th subdivision iteration, and let precisely $m_{i}$ e/i tests be performed at the $i$-thiteration, so that
+$m_{i}\leq 4^{i+1},~{}i=0,1,\dots,k-1.$ (2.1)
+Then
+(i) the $(\Delta_{h}^{\prime})$-neighborhood of $\mathbb{D}$ covers the union of all suspect squares processed in the $h$-th subdivision iteration, while $\Delta_{h}^{\prime}=\Delta_{0}^{\prime}/2^{h}$ fast converges to 0 as $h$ increases; in particular, for any fixed $\beta$ of Def.
+1.1 we have $\Delta^{\prime}_{h}\leq 0.5\beta\Delta^{\prime}_{0}$ already for $h=O(1)$.
+(ii) The first $h$ iterations involve at most $\frac{4^{h+1}}{3}$ e/i tests, that is, $O(1)$ tests if $h=O(1)$.
+Proof.
+We only prove claim (ii). Eqn. (2.1) implies that $\sum_{i=0}^{h-1}m_{i}\leq\sum_{i=0}^{h-1}4^{i+1}=\frac{4^{h+1}}{3}$. ∎
 Remark 2.1.
-[Real and near-real root-finding.] In the case of Problem 0 of Example 1.2, generalize bisection as follows: (i) recursively partition (subdivide) every suspect rectangle into two congruent rectangles, to each apply e/i test, and then discards a rectangle if it contains no roots but call it suspect and subdivides it otherwise, (ii) stop where the rectangles have diameter at most $2\epsilon$, so that their centers approximate the roots within $\epsilon$.
+Every root is defined up to within $\epsilon$, and so at the $h$-th subdivision iteration we decide whether we should discard a square or call it suspect and keep subdividing it depending on the roots that lie in the $(\Delta_{h}^{\prime}+\epsilon)$-neighborhood of $\mathbb{D}$.
+Def.
+1.1 for $\Delta:=\Delta_{0}^{\prime}$ implies that $\Delta^{\prime}_{h}\epsilon\leq\beta\Delta$ already for some $h=O(1)$, and so, the outputs of all e/i tests, except for $h=O(1)$ initial ones, depend only on the roots that lie in the $\beta\Delta$-neighborhood of $\mathbb{D}$.
+Remark 2.2.
+[Real and near-real root-finding.] In the case of Problem 0 of Example 1.2, generalize bisection root-finding as follows: (i) recursively partition (subdivide) the real line segment defining every suspect rectangle into two halves and accordingly subdivide the rectangle into two congruent rectangles; to each of them apply e/i test, and then discard a rectangle if it contains no roots but call it suspect and subdivide it otherwise, (ii) stop where a rectangle has diameter at most $2\epsilon$, so that its center approximates within $\epsilon$ all roots lying in it.
 2.2 Subdivision iterations with soft e/i tests
@@
 Furthermore, in the presence of rounding errors no e/i test can decide whether a root lies in or outside a disc if it actually lies on or very close to its boundary circle.
-To address both problems we fix $\sigma>1$ and apply soft e/i tests.
+We address both problems by fixing $\sigma>1$ and applying soft e/i tests, which are a special case for $\ell=1$ of the following more general soft $\ell$-tests.
+[^7] Besides 1-tests (that is, e/i tests) we only use $\ell$-tests for sub-discs of discs containing at most $\ell$ roots and well-isolated from the external roots.
 Definition 2.1.
-[See the end of Sec.
-6 for a generalization, covering in particular soft $m$-tests.] Given a real $\sigma>1$, a polynomial $p$, a square $S$, and its minimal covering disc $D(c,\rho)$, the $\sigma$-soft e/i test reports either exclusion or $\sigma$-soft inclusion.
-Exclusion certifies that the square $S$ contains no roots and therefore should be discarded from subdivision process.
-$\sigma$-soft inclusion certifies that the disc $D(c,\sigma\rho)$, said to be the $\sigma$-cover of $S$, contains a root; then the square $S$ is said to be suspect and is subdivided into four congruent sub-squares in the next subdivision step.
-Criteria for both exclusion and $\sigma$-soft inclusion can hold simultaneously (see Fig.
-1b), but the test ends as soon as it verifies any of them.
+For $p(x)$ of (1.1), $\sigma>1$, and an integer $\ell$, $1\leq\ell\leq d$, a $\sigma$-soft $\ell$-test, or just $\ell$-test for short (in particular an e/i test turns into a 1-test), either outputs 1 and stops if it detects that $r_{d-\ell+1}\leq\sigma$, that is, #$\#(D(0,\sigma))\geq\ell$, or outputs 0 and stops if it detects that $r_{d-\ell+1}>1$, that is, [^8] Bounds $r_{d-\ell+1}\leq\sigma$ and $r_{d-\ell+1}>1$ can hold simultaneously (cf.
+Fig.
+1b), but as soon as an $\ell$-test verifies any of them, it stops without checking if the other bound also holds.
+#$\#(D(0,1))<\ell$.
+An $\ell$-test applied to the polynomial $t(y)$ of (1.4) is said to be an $\ell$-test for the disc $D(c,\rho)$ and/or the circle $C(c,\rho)$ and also to be an $\ell$-testc,ρ.
 Observation 2.2.
 Let $\gamma(\sigma)$ denote the maximal number of $\sigma$-covers at a fixed subdivision step that share a complex point, e.g., a root.
-Then $\gamma(\sigma)$ is in $O(1)$ if so is $\sigma$, and $\gamma(\sigma)\leq 4$ if $\sigma<\sqrt{2}$.
+Then $\gamma(\sigma)\leq 4$ if $\sigma<\sqrt{2}$ and $\gamma(\sigma)$ is in $O(1)$ if so is $\sigma$.
 Corollary 2.2.
 Cor.
-2.1 still holds if e/i tests are replaced with $\sigma$-soft e/i tests for $\sigma<\sqrt{2}$, while a subdivision root-finder using $\sigma$-soft e/i tests involves $O(km)$ such tests if $\sigma$ is a constant.
+2.1 and the observation of Remark 2.1 still hold if e/i tests are replaced with $\sigma$-soft e/i tests for $\sigma<\sqrt{2}$, while a subdivision root-finder using $\sigma$-soft e/i tests involves $O(km)$ such tests if $\sigma$ is a constant.
 3 Fast soft e/i tests
-3.1 Background: Newton’s Inverse Ratio and its approximation
-We reduce e/i tests of subdivision root-finders to the multipoint evaluation of the ratio
+3.1 Background: Newton’s Inverse Ratio
+We reduce e/i tests of subdivision root-finders to multipoint evaluation of the ratio
 ${\rm NIR}(x)=\frac{p^{\prime}(x)}{p(x)}=\sum_{j=1}^{d}\frac{1}{x-z_{j}},$ (3.1)
-which we call Newton’s inverse ratio. To prove the latter well-known and much used equation, differentiate $p(x)=p_{d}\prod_{j=1}^{d}(x-z_{j})$.
-The arithmetic operations (hereafter ops) involved in our evaluation of NIR$(x)$ are more numerous than the other ops involved in our root-finders, even in the case where $p(x)=x^{d}$.
-By following [76] we fix a complex $\delta$, with $|\delta|$ small in context, and approximate $p^{\prime}(x)$ with the first order divided difference $\frac{p(x)-p(x-\delta)}{\delta}$, equal to $p^{\prime}(y)$ for some $y\in[x-\delta,x]$, by virtue of Taylor-Lagrange’s formula.
+said to be Newton’s inverse ratio. To prove the latter well-known equation, differentiate $p(x)=p_{d}\prod_{j=1}^{d}(x-z_{j})$.
+The ops involved in our evaluation of NIR$(x)$ are more numerous than the other ops involved in our root-finders, even in the case where $p(x)=x^{d}$.
+By following [59] we fix a complex $\delta$, for $|\delta|$ small in context, and approximate $p^{\prime}(x)$ with the first order divided difference $\frac{p(x)-p(x-\delta)}{\delta}$, equal to $p^{\prime}(y)$ for some $y\in[x-\delta,x]$, by virtue of Taylor-Lagrange’s formula.
 Hence $y\mapsto x$ and $p^{\prime}(y)\mapsto p^{\prime}(x)$ as $\delta\mapsto 0$, and
 ${\rm NIR}(x)\approx\frac{p(x)-p(x-\delta)}{p(x)\delta}=\frac{p^{\prime}(y)}{p(x)}~{}{\rm for~{}some}~{}y\in[x-\delta,x].$ (3.2)
-We deduce the same bound $O(b\log(d))$ of [76] on computational precision of our root-finders, but its derivation is greatly simplified versus [76] because we only use the first order derivatives and the first order divided differences.
-Unlike [76] we estimate bit-operation cost of our root-finders and allow random choice of $x$ to control the value of the ratio $\frac{|p(x)-p(x-\delta)|}{p(x)\delta}$ with a high probability (whp).
-This is Las Vegas randomization, and so is any randomization in polynomial root-finding, as we show next.
 3.2 Background: error detection and correction
-A randomized root-finder can lose some roots, although with a low probability, but we can detect such a loss at the end of root-finding process, simply by observing that among the $m\leq d$ roots lying in an input disc only $m-w$ tame roots, say, $z_{w+1},z_{w+1},\dots,z_{m}$ have been closely approximated, [^6] We call a complex $c$ a tame root for an error tolerance $\epsilon$ if $c$ lies in an isolated disc $D(c,\epsilon)$, whose index #$\#(D(c,\epsilon))$ we can readily compute (cf.
-Sec.
-23.4).
+A randomized root-finder can lose some roots, although with a low probability, but we can detect such a loss at the end of root-finding process, simply by observing that among the $m\leq d$ roots lying in an input disc $D$ only $m-w$ tame roots, say, $z_{w+1},z_{w+1},\dots,z_{m}$ have been closely approximated, [^9] We call a complex $c$ a tame root for an error tolerance $\epsilon$ if $c$ lies in an isolated disc $D(c,\epsilon)$, whose index #$\#(D(c,\epsilon))$ we can compute fast by applying Thm.
+12.1 or the algorithm in Sec.
+17.5.
 while $w>0$ wild roots, say, $z_{1},z_{2},\dots,z_{w}$ remain at large.
-We can compute the coefficients of the wild factor $q(x):=p(x)/\prod_{j=w+1}^{m}(x-z_{j})$, say, by applying evaluation-interpolation techniques, then apply to $q(x)$ the same or another root-finder, and repeat this computation recursively until we approximate all $m$ roots in the disc.
-[^7] By this recipe we detect output errors at the very end of computations in any root-finder that can only lose roots but cannot wrongly generate them.
+We can compute the coefficients of the factor $g(x):=\frac{p(x)}{f(x)}$ for $f(x):=\prod_{j=w+1}^{m}(x-z_{j})$, say, by applying evaluation-interpolation techniques, then apply to $g(x)$ a black box root-finder, and repeat this computation recursively to approximate the $w$ wild roots.
+[^10] By this recipe we detect any output error at the very end of computations in any root-finder that can only lose roots but cannot wrongly generate them.
 Our subdivision root-finders belong to this class, but with them we can detect the loss of a root even earlier – whenever in a subdivision process applied to a square with $m$ roots we detect that $m$ exceeds the sum of the indices of all suspect squares, that is, the overall number of roots contained in them.
-Readily bounds the complexity and error probability of this recursive process:
+We readily bound the complexity and error probability of this recursive process (although we will not use these estimates in the following):
 Proposition 3.1.
@@
 Then in $v$ recursive applications, the output is certified to be correct at the cost of less than $(\alpha+\delta)v$ with a probability at least $1-1/\beta^{v}$.
-Explicit deflation can blow up the coefficient length and destroy sparsity unless the degree $w$ is sufficiently small, but we avoid these mishaps by combining a fixed black box polynomial root-finder with implicit deflation, that is, by applying that root-finder to the quotient polynomial $q(x)$ without computing its coefficients – based on the following expression, implied by (3.1):
-$\frac{q^{\prime}(x)}{q(x)}=\frac{p^{\prime}(x)}{p(x)}-\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}.$ (3.3)
-One can accelerate low precision numerical approximation of the sum $\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}$ simultaneously at many points $x$ by applying Fast Multipole Method (FMM) of [50, 28, 32].
-3.3 Cauchy-based root-counting and e/i tests
-3.3.1 Introductory comments
-Let #$\#(\mathcal{D})$ denote the number of roots lying in a complex domain $\mathcal{D}$ with the boundary $\Gamma$ and recall from [1] that
-[…]
-that is, #$\#(\mathcal{D})$ is the average of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ on the boundary $\Gamma$.
-Having an approximation of the integral of (3.4) within less than 1/2, we can round it to the closest integer and obtain $s_{0}=\#(D(0,1))$.
-This is more than enough for our e/i tests, for which we only need to know whether $s_{0}=0$ or not.
-3.3.2 Cauchy-based soft e/i tests: outline
-In Part II we approximate the integral of (3.4) with finite sums made up of the scaled values of NIR$(x)$ at the scaled $q$th roots of unity.
-Namely, for a fixed $q>1$ let $\zeta:=\zeta_{q}$ denote a primitive $q$th root of unity:
-$\zeta:=\zeta_{q}:=\exp\Big{(}\frac{2\pi{\bf i}}{q}\Big{)},~{}{\bf i}:=\sqrt{-1},$ (3.5)
-and devise randomized $\sigma$-soft e/i tests as follows:
-(i) Choose a real $\rho$ at random in a fixed range covered by $[1,\sigma]$.
-(ii) Fix a sufficiently large integer $q$ of order $m\log(d)$.
-[…]
-(iv) Conclude that exclusion holds whp if $|s_{0,q}|<1/2$.
-(v) Repeat the test $v$ times to decrease the probability of its output error exponentially in $v$, from $\frac{1}{\gamma}$ to $\frac{1}{\gamma^{v}}$ for $\gamma>1$ of our choice, say, for $\gamma=2$.
-The $v$ tests are reduced to the evaluation of $p(x)$ at $O(mv\log(d))$ points.
-3.4 Soft e/i test with root-lifting: generic algorithm
-In Part III we decrease the complexity estimates of Part II for $q$ in e/i tests by a factor of $m$ by avoiding approximation of the integer #$\#(D(0,1))$ and only deciding whether this integer is 0 or not, which is sufficient for e/i tests.
-In these e/i tests we elaborate upon the generic one (see Alg.
-3.1 below), where we use $h$-lifting of roots:
-$p_{0}(x):=\frac{p(x)}{p_{d}},~{}p_{h}(x):=\prod_{g=0}^{h-1}p_{0}(\zeta_{h}^{g}x)=(-1)^{d}\prod_{j=1}^{d}(x-z_{j}^{h}),~{}\zeta_{h}:=\exp\Big{(}\frac{2\pi{\bf i}}{h}\Big{)},$ (3.6)
+Explicit deflation can blow up the coefficient length and destroy sparsity unless the degree $w$ is sufficiently small, but we avoid these mishaps by combining a fixed black box polynomial root-finder with implicit deflation (cf.
+example (iii) in Sec.
+1.4.1), that is, by applying that root-finder to the quotient polynomial $q(x)$ without computing its coefficients – based on the following expression, implied by (3.1):
+$\frac{q^{\prime}(x)}{q(x)}=\frac{p^{\prime}(x)}{p(x)}-\frac{t^{\prime}(x)}{t(x)}=\frac{p^{\prime}(x)}{p(x)}-\sum_{j=w+1}^{m}\frac{1}{x-z_{j}}.$ (3.3)
+Remark 3.1.
+One can accelerate low precision numerical approximation of the sum $\sum_{j=w+1}^{m}\frac{1}{x-z_{j}}$ simultaneously at many points $x$ by applying Fast Multipole Method (FMM) by Greengard and Rokhlin [37, 20, 24] (see more in Sec.
+24.3).
+3.3 Background: root-lifting
+Our e/i tests of Part II involve $h$-lifting of roots:
+$p_{0}(x):=\frac{p(x)}{p_{d}},~{}p_{h}(x):=\prod_{g=0}^{h-1}p_{0}(\zeta_{h}^{g}x)=(-1)^{d}\prod_{j=1}^{d}(x-z_{j}^{h}),~{}\zeta_{h}:=\exp\Big{(}\frac{2\pi\sqrt{-1}}{h}\Big{)},$ (3.4)
 for $h=1,2,\dots$ and $t_{2^{k}}(x)=p_{k}(x)$ for all $k$.
@@
 Remark 3.2.
-The $h$-lifting generalizes the DLG classical root-squaring iterations of Dandelin 1826, Lobachevsky 1834, and Gräffe 1837 (cf. [51]):
+For $h$ being a power of 2, the $h$-lifting generalizes the DLG classical root-squaring iterations of Dandelin 1826, Lobachevsky 1834, and Gräffe 1837 (cf. [38]):
 $p_{0}(x):=\frac{p(x)}{p_{d}},~{}p_{2^{i}}(x^{2}):=(-1)^{d}p_{2^{i-1}}(x)p_{2^{i-1}}(-x)=\prod_{j=1}^{d}(x-z_{j}^{2^{i}}),~{}i=1,2,\dots.$ (3.5)
 Polynomial root-finding in the 19-th and well into the 20-th century relied on iterations (3.5) combined with Vieta’s formulas, which link the polynomial coefficients $p_{0},p_{1},\dots,p_{d}$ to the elementary symmetric polynomials of the roots.
-The combination enabled efficient approximation of the minimal and maximal distance of the roots from the origin and due to the map (1.3) from any complex point (see [86, 113, 115] for details).
+The combination enabled efficient approximation of the minimal and maximal distance of the roots from the origin and, due to the map (1.4), from any complex point (see [69, 94, 96] for details).
 The approach became obsolete because of severe problems of numerical stability of the DLG iterations, but they disappeared in the case of root-finders that work for black box polynomials.
 Extensive tests with standard test polynomials have consistently demonstrated efficiency of the particular implementation of this approach in [96], but even the straightforward implementation of iterations (3.5) promises to be highly efficient as well.
-We will elaborate upon the following generic algorithm in Sec. 14.
+The paper [36] explores direct root-squaring for NIR of a black box polynomial:
+${\rm NIR}_{p_{h+1}}(x)=\frac{1}{2\sqrt{x}}\left({\rm NIR}_{p_{h}}(\sqrt{x}){\rm NIR}_{p_{h}}(-\sqrt{x})\right),~{}x\neq 0,~{}~{}h=0,1,\dots.$
+3.4 Soft e/i test with root-lifting: generic algorithm
+In Sec. 7 we elaborate upon the following generic soft e/i test applied to the unit disc $D(0,1)$.
 Algorithm 3.1.
@@
 INPUT: $\sigma>1$, an integer $d>0$, and a $d$-th degree black box polynomial $p(x)$ given by an oracle for its evaluation.
-INITIALIZATION: Fix a positive $q\in\mathbb{N}$ and a positive integer $h$.
-COMPUTATIONS:
-(1) Select $q$ points $x_{1},\dots,x_{q}$ on the circle $C(0,1)$.
-(2) By extending (3.2) to $p_{h}(x)$ of (3.7) (we specify this extension in (13.3)) approximate the values
+INITIALIZATION: Fix two positive integers $h$ and $q$ and $q$ points $x_{1},\dots,x_{q}$ on the circle $C(0,1)$.
+COMPUTATIONS: By extending (3.2) to $p_{h}(x)$ of (3.4) (see (6.3)) approximate the values
 $v_{g}:=|{\rm NIR}_{p_{h}}(x_{g})|=\Big{|}\frac{p^{\prime}_{h}(x_{g})}{p_{h}(x_{g})}\Big{|},~{}g=1,\dots,q.$ (3.6)
-(3) If $1+d/v_{g}\leq\sigma^{h}$ for some $g\leq q$, then certify that #$(D(0,\sigma))>0$, that is, claim $\sigma$-soft inclusion. Otherwise, claim that #$(D(0,1))=0$, that is, claim exclusion.
+If $1+d/v_{g}\leq\sigma^{h}$ for some $g\leq q$, then certify that #$(D(0,\sigma))>0$ ($\sigma$-soft inclusion). Otherwise, claim that #$(D(0,1))=0$ (exclusion).
+Remark 3.3.
+(6.3) and (3.4) together express $v_{g}$ for all $g$ through $2hq$ values of $p(x)$, by-palink to the values of $p_{h}(s)$ helps us o prove correctness of the algorithm.
 3.5 Soft e/i test with root-lifting: correctness of inclusion claim
@@
 Corollary 3.2.
-It holds that $r_{d}(c)\leq d~{}\Big{|}\frac{p(c)}{p^{\prime}(c)}\Big{|}$, that is, #$\#\Big{(}D(c,d~{}\Big{|}\frac{p(c)}{p^{\prime}(c)}\Big{|}\Big{)}\Big{)}>0$, for any complex $c$.
+It holds that $r_{d}(c)\leq d~{}\Big{|}\frac{p(c)}{p^{\prime}(c)}\Big{|}$, that is, #$\#\Big{(}D\Big{(}c,d~{}\Big{|}\frac{p(c)}{p^{\prime}(c)}\Big{|}\Big{)}\Big{)}>0$, for any complex $c$.
 Proof.
@@
 $\Big{|}\frac{p^{\prime}(c)}{p(c)}\Big{|}=\Big{|}\sum_{j=1}^{d}\frac{1}{c-z_{j}}\Big{|}\leq\sum_{j=1}^{d}\frac{1}{|c-z_{j}|}<\frac{d}{\rho},~{}{\rm and~{}so}~{}\Big{|}\frac{p(c)}{p^{\prime}(c)}\Big{|}>\frac{\rho}{d}.$
-We use Cor.
-3.1 later but next apply Cor.
-3.2 to prove correctness of the inclusion claim of Alg.
-3.1.
-The bound of Cor.
-3.1 on $r_{d}(c)$ is too large for that, but a similar bound on $r_{d}(c,p_{h})=(r_{d}(c))^{h}$, implies that $r_{d}(c)$ is close to 1 already for $h$ of order $\log(d)$.
+We use Cor. 3.1 later but next apply Cor. 3.2 to $p_{h}$ for a sufficiently large $h$ of order $\log(d)$ to prove correctness of the inclusion claim of Alg. 3.1.
 Lemma 3.1.
@@
 3.6 Soft e/i test with root-lifting: further analysis (outline)
-(i) Correctness of the exclusion claim of Alg.
-3.1.
-We prove it separately for its two specifications, both of which strengthen isolation of the unit circle $C(0,1)$ by means of root-squaring (3.7) or root-lifting (3.6).
+We prove (i) correctness of the exclusion claim of Alg. 3.1 separately for two specifications, both strengthening isolation of the unit circle $C(0,1)$ from the roots by means of root-lifting (3.4).
 (a) In Sec.
-14.3 we choose any integer $q$ exceeding $d$ and then prove deterministic correctness of exclusion claim, provided that $|{{\rm NIR}_{p_{h}}(x_{i})}|<1/(2\sqrt{q})$ for all $x_{i}$ being $q$th roots of unity.
-This bound is compatible with the bound of Example 3.1 because $4\sqrt{d}-\frac{1}{d}>3\sqrt{d}$ for $d\geq 2$.
-We also extend this deterministic e/i test by allowing any $q>m$ (rather than $q>d$) provided that $|{{\rm NIR}_{p_{h}}(x_{i})}|<1/(3\sqrt{q})$ for all $x_{i}$ being $q$th roots of unity.
+7.3 we choose any integer $q$ exceeding $d$ and then prove correctness of exclusion claim, provided that $|{{\rm NIR}_{p_{h}}(x_{i})}|<1/(2\sqrt{q})$ for all $x_{i}$ being $q$th roots of unity.
+This bound is compatible with the bound of Observation 3.1 because $4\sqrt{d}-\frac{1}{d}>3\sqrt{d}$ for $d\geq 2$.
+We also extend this e/i test by allowing any $q>m$ (rather than $q>d$) provided that the unit disc contains $m$ roots and $|{{\rm NIR}_{p_{h}}(x_{i})}|<1/(3\sqrt{q})$ for all $x_{i}$ being $q$th roots of unity.
 (b) In Sec.
@@
 Eqn.
-(3.7) reduces approximation of $p_{h}(x)$ at $x=c$ to approximation of $p(x)$ at all the $h$th roots of $c$.
-We can fix $h>1$ being a power of 2 and apply (3.7) or fix any integer $h>1$ and apply (3.6).
+(3.4) reduces approximation of $p_{h}(x)$ at $x=c$ to approximation of $p(x)$ at all $h$th roots of $c$.
+We can fix an integer $h>1$ and apply (3.4) or fix $h>1$ being a power of two and apply (3.4).
 In both cases Alg.
-3.1 evaluates polynomial $p(x)$ either at $2jh$ points for $j\leq q$, when it claims inclusion, or at $2qh$ points, when it claims exclusion.
-We always choose $h=O(\log(d))$ and then can ignore a factor of $h$ in our $\tilde{O}$ estimates.
-In our deterministic soft e/i test based on Alg.
-3.1 we evaluate polynomial $p(x)$ at $\tilde{O}(m)$ points, combine these estimates with Cor.
-2.2, and bound by $\tilde{O}(bm^{2})$ the overall number of points $x$ at which we evaluate $p(x)$ in subdivision root-finders incorporating our e/i tests.
+3.1 evaluates polynomial $p(x)$ either at $2jh$ points for $j\leq q$ when it claims inclusion or at $2qh$ points when it claims exclusion.
+We always choose $h=O(\log(d))$ and then ignore a factor of $h$ in our $\tilde{O}$ estimates.
+In our soft e/i test based on Alg.
+3.1 we evaluate polynomial $p(x)$ at $\tilde{O}(m)$ points and then combine these estimates with Cor.
+2.2, and bound by $\tilde{O}(bm^{2})$ the overall number of evaluation queries in the resulting root-finders.
 Under the Random Root Model of Sec.
 1.5 the roots are reasonably well isolated whp from the unit circle $C(0,1)$.
-We strengthen this isolation by applying $O(\log(\log(d)))$ root-squaring steps (3.7) or $O(\log(d))$ steps of root-lifting of (3.6) and then whp decrease the overall bit operation count from $\tilde{O}(bm^{2})$ to $\tilde{O}(bm)$.
-(iii) Estimation of computational precision and bit-operation complexity.
-Louis and Vempala in [76] bound by $O(b\log(d))$ the computational precision of their novel high order Newton’s iterations.
-We combine a small part of their advanced analysis with Las Vegas randomization and some customary techniques of error analysis to extend this estimate to our root-finders.
-For the solution of Problems 0, 1, and 3, where the coefficients are given, we apply fast algorithm of [81] for multipoint polynomial evaluation to decrease the estimated number of queries and the bit operation cost bound.
-We devise two groups of efficient e/i tests in Parts II and III of this paper, and in both cases arrive at the estimates for the numbers of (i) evaluation points, that is, of the oracle queries, and (ii) bit-operations involved in our algorithms.
-Both of the estimates exceed those of our Main Theorems by factors of $b$ for the roots and $\tilde{b}$ for the eigenvalues.
-4 Acceleration of subdivision root-finding by means of compression: outline
-In Part IV we accelerate the algorithms of both Parts II and III by these factors (of $b$ for roots and $\tilde{b}$ for eigenvalues) to prove the Main Theorems in two ways.
+We strengthen this isolation by applying $O(\log(d))$ root-lifting steps of (3.4) and then whp bound the overall number of queries by $\tilde{O}(bm)$.
+(iii) Computational precision and bit-operation complexity.
+We deduce the same bound $O(b\log(d))$ of [59] on computational precision of our root-finders; our derivation is much simpler because unlike [59] we only use the first order derivatives and divided differences.
+We combine this estimate and our count of evaluation queries in our root-finders with (i) Storjohann’s algorithm for Las Vegas computation of the determinant [109] to obtain record fast solution of Problems 2 and $2^{*}$ of matrix eigenvalue approximation (see alternative extension to eigenproblem in part (C) of Sec.
+5.4) and (ii) the complexity estimates of Guillaume Moroz [64] for his fast multipoint evaluation of a polynomial represented in monomial basis to arrive at the near-optimal complexity bounds of our Main Theorems 1 – 4 within a factor of $b$, to be removed in Part IV.
+3.7 Cauchy-based soft e/i tests: outline
+Next we outline our alternative soft e/i test of Part III. Let #$\#(\mathcal{D})$ denote the number of roots lying in a complex domain $\mathcal{D}$ with the boundary $\Gamma$ and recall from [1] that
+[…]
+that is, #$\#(\mathcal{D})$ is the average of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ on the boundary $\Gamma$.
+Having an approximation of integral (3.7) within less than 1/2, we can round it to a closest integer and obtain $s_{0}=\#(D(0,1))$.
+This root-counter defines $\ell$-tests for all $\ell\leq d$ and in particular an e/i test, for $\ell=1$.
+In Part III we approximate the integral (3.7) with finite sums of the values of $\frac{x}{q}$$(x)$ at the scaled $q$th roots of unity.
+Namely, let $\zeta:=\zeta_{q}$ denote a primitive $q$th root of unity (cf.
+(3.4)) and devise randomized $\sigma$-soft e/i tests as follows.
+Algorithm 3.2.
+[Cauchy e/i test].
+(i) Fix a constant $\gamma>1$, say, $\gamma=2$, and choose a real $\rho$ at random in a fixed range in $(1,\sigma)$.
+(ii) Fix a sufficiently large integer $q$ of order $m\log(d)$.
+[…]
+(iv) Unless $\bar{s}_{0}=0$ certify inclusion.
+Otherwise, repeat the test $v$ times for a fixed integer $v>0$ (at the cost of the evaluation of $p(x)$ at $O(mv\log(d))$ points).
+Unless inclusion is certified in these tests, claim exclusion (correctly with a probability at least $1-1/\gamma^{v}$).
+4 Faster subdivision with compression: outline
+In Part IV we accelerate the algorithms of both Parts II and III by a factor of $b/\log(b)$ to ensure pairwise isolation of clusters of roots and eigenvalues and to prove Main Theorems 1 – 4; we prove the first three of them in two ways – by devising two distinct supporting root-finders based on Parts II and III, respectively.
 Next we outline that acceleration.
 (i) Component tree and compact components.
-Apply subdivision process for a fixed $\sigma$, $1<\sigma<\sqrt{2}$, to Problem 1 and partition the union of all suspect squares tested at every subdivision step into connected components.
-Represent all partitions of components with a component tree, $\mathbb{T}$, as follows.
-At most $m$ components, said to be the leaves, are covered by isolated $\epsilon$-discs for a fixed error tolerance $\epsilon\geq 0$.
-The other (at most $m-1$) vertices of the tree are components-parents, each broken by a subdivision step into at least two components-children.
+For fixed integer $m$, positive $\epsilon$, and $\sigma$, $1<\sigma<\sqrt{2}$, apply subdivision process to Problem 1 and at every subdivision step partition the union of the suspect squares into connected components.
+Figure 2: Roots (asterisks) define compact components. A subdivision step halves their diameters and accordingly strengthens isolation of their minimal covering squares and discs.
+Represent all partitions of components with a component tree whose at most $m$ leaves represent components covered by isolated $\epsilon$-discs.
+The other (at most $m-1$) vertices of the tree represent components-parents, each broken by a subdivision step into at least two components-children.
 The initial suspect square is the root of the tree.
-The tree has at most $22m-$ 2 edges, connecting parents with their children and representing partition of the components but not showing modification of components between their partitions.
-Recall that every component contains $O(m)$ suspect squares whose diameter is halved at every subdivision step.
-Hence in $O(\log(b))$ subdivision iterations, for $b:=\log_{2}(1/\epsilon)$, every component has diameter below $\epsilon/2$, and then its $\sigma$-cover lies in an $\epsilon$-disc.
-Call a component compact if it consists of at most four squares sharing a vertex (see Fig.
+The tree has at most $22m-$ 2 edges, connecting parents with their children and representing partition of the components but not showing modification of components between partitions.
+Every component contains $O(m)$ suspect squares whose diameter is halved at every subdivision step.
+Hence in $O(b)$ subdivision iterations, for $b:=\log_{2}(2\Delta^{\prime}/\epsilon)$ and the diameter $\Delta^{\prime}$ of the initial suspect square, every component has diameter below $\epsilon/2$ and hence lies in an $\epsilon$-disc.
+(ii) Faster subdivision with compression.
+Call a component $\mathcal{C}$ compact if all (at most four) its squares share a vertex; subdivision steps fast decrease the diameter of $\mathcal{C}$ but never decrease its distance from external roots (see Fig.
 2).
-Recall that a subdivision step never includes new roots into a component and deduce that unless a component stays compact it is soon broken into more than one component or shrinks enough to be covered by an $\epsilon$-disc.
-[…]
-Figure 2: Roots (asterisks) define compact components. A subdivision step halves their diameters and accordingly strengthens isolation of their minimal covering squares and discs.
-(ii) Acceleration of subdivision process with compression of components.
-Subdivision process becomes quite inefficient where a component stays compact and not broken into more than one component in many consecutive subdivision steps (see Fig.
-3).
-To accelerate it, we replace these steps with compression of such a component into an $\epsilon$-disc or a near-minimal disc $D$ that still keeps all roots lying in the component.
-In Sec.
-19 we comment on two customary but inferior alternatives to our compression, relying on deflation or solely on functional iterations.
+Subdivision becomes inefficient where a component stays unbroken and compact in many subdivision steps (see Fig.
+3), and we replace them with its compression with no loss of roots, of which the component very soon becomes strongly isolated.
+The compression outputs either an $\epsilon$-disc, and then we stop, or a rigid disc $D$, that is, minimal up to a fixed constant factor, and then we subdivide a minimal superscribing square $S$ of $D$.
+Since the disc $D$ is rigid, the iterations very soon output more than one component or a sub-domain of an $\epsilon$-disc.
 Figure 3: Five blue marked centers of suspect squares converge to a green-marked root cluster with linear rate; two red marked Schröder’s iterates of (17.1) converge with quadratic rate.
-If the compression outputs $\epsilon$-disc, we just output it and do not subdivide it anymore.
-Otherwise we apply subdivision iterations to the minimal superscribing square $S$ of the disc $D$.
-Since the disc is near-minimal, the iterations very soon output more than one component or a sub-domain of an $\epsilon$-disc.
-From this we deduce that our subdivision process with compression reduces the solution of Problem 1 to at most $22m-$ 2 compression steps and $O(m)$ soft e/i steps overall.
-(iii) Compression of components. We reduce compression of a compact component $C$ to the following two main stages (see more details in Sec. 23.5):
-(a) compute a point $c^{\prime}$ lying in or near the minimal disc that covers all roots lying in the component $C$ and let $c^{\prime}$ denote the center of the compressed disc that we seek;
-(b) compute a close upper bound $\rho^{\prime}$ on the maximal distance from $c^{\prime}$ to the roots in the component $C$ and output a compressed disc $D(c^{\prime},\rho^{\prime})$.
-We prove that Schröder’s single iteration (23.1) following $O(\log(d))$ auxiliary subdivision steps performs step (a) above, that is, computes the desired center $c^{\prime}$.
-At the beginning of step (b) above we have a range $[\rho_{-},\rho_{+}]$ for the bound $\rho^{\prime}$, $0<\epsilon\leq\rho_{-}<\rho_{+}\leq 1$.
-Hence we can approximate it in $\lceil b\rceil$ bisection steps, for $b=\log_{2}(1/\epsilon)$ but we greatly accelerate convergence to $\rho$ by applying bisection to the range of the logarithms, $[\log_{2}(\rho_{-}),\log_{2}(\rho_{+})]=[b,\log_{2}(\rho_{+})]$, for $\log_{2}(\rho_{+})\leq 0$.
-We extend our fast soft e/i tests of Part II to fast soft $m$-tests (see part (A) of Sec. 2.1) and use them to define the pivot step of such a modified bisection.
-Summarizing we reduce the solution of Problems 0 and 1 to $\tilde{O}(m)$ soft e/i tests and $m$ tests and $\tilde{O}(m)$ Schröder’s iterations.
-In this way we accelerate the root-finders of both Parts II and III by a factor of $b/\log(b)$ to complete the proof of our Main Theorems.
+This way we solve Problem 1 by applying at most $22m-$ 2 compression steps and in addition $O(m)$ soft e/i tests, as we deduced in Sec. 19.
+(iii) Compression of components. We compress a compact component $\mathcal{C}$ into a disc $D=D(c,\rho)$ by computing:
+(a) the number of roots in a component,
+(b) a point $c$ in or near the minimal disc covering the root set of $\mathcal{C}$ (hereafter the MCD of $\mathcal{C}$) and
+(c) a reasonably close upper bound $\rho$ on the maximal distance from $c$ to the roots in $\mathcal{C}$.
+In Secs. 17.4 and 17.5 we readily perform steps (a) and (b) at a low computational cost because subdivision iterations fast strengthen the isolation of the component from external roots.
+We begin step (c) with a range $[\rho_{-},\rho_{+}]$ for the bound $\rho$, $0<1/2^{b}=\rho_{-}<\rho_{+}$.
+We can approximate $\rho$ by applying bisection to this range but proceed faster by applying bisection of the logarithms, said to be BoL, to the range $[\log_{2}(\rho_{-}),\log_{2}(\rho_{+})]=[b,\log_{2}(\rho_{+})]$, for $\log_{2}(\rho_{+})\leq 0$.
+We apply fast soft $m$-tests (see part (A) of Sec.
+2.1) to define a pivot step of our accelerated bisection, with which we ultimately reduce compression to $O(m\log(b))$ e/i tests.
+Their computational cost dominates the overall cost of the solution of Problems 0,0, 1, and 1, and this enables us to complete the proof of our Main Theorems.
 Remark 4.1.
 Many “natural" directions to proving our Main Theorems have led us into dead end.
-E.g., we defined a tree whose vertices represented non-adjacent suspect squares; then we readily bounded by $12m-$ 1 their overall number at all partition stages but could not count them so readily between these stages until we decided to count all suspect squares.
+E.g., we tried to use a tree whose vertices represented components made up of non-adjacent suspect squares; then we readily bounded by $12m-$ 1 their overall number at all partition stages but could not count them so readily between these stages until we decided to count all suspect squares.
 See more examples of this kind in the next remark and Remark 7.1.
@@
 Subdivision iterations of the known root-finders, in particular, of those of [17], scale the variable $x$ order of $m$ times by factors of $2^{b}$.
-Hence the bit-complexity specified in the Main Theorem of [24] has order of $mbd^{2}$, which exceeds our bound and that of [96, 103] by a factor of $d$.
+Hence the bit-complexity specified in the Main Theorem of [17] has order of $mbd^{2}$, which exceeds our bound and that of [78, 84] by a factor of $d$ (see more comments in Sec.
+10.4).
 5 Our novelties, related works, implementation, and further progress
 5.1 Our novelties (a short list)
-Root-finding for polynomial given in the monomial basis, with their coefficients, has been studied for four millennia and very intensively in the last decades, but we had to introduce a number of nontrivial novel techniques (i) to handle efficiently black box polynomials and (ii) to accelerate the known subdivision algorithms to a near optimal level.
+Root-finding for polynomial given in the monomial basis, with their coefficients, has been studied for four millennia and very intensively in the last decades, but we had to introduce a number of nontrivial novel techniques (i) to handle efficiently black box polynomials and (ii) to accelerate the known subdivision algorithms to a near-optimal level.
 Some of our techniques such as randomization, numerical stabilization of root-squaring, compression of a disc without losing roots, bisection of logarithms, approximation of root radii and near-real roots, and even our simple error detection recipe can be of independent interest.
 5.2 Related works
-The pioneering paper [76] proposed a novel algorithm for approximation of an absolutely largest zero of a real-rooted black box polynomial, estimated the computational precision of that algorithm, and extended it to approximation of an absolutely largest eigenvalue of a symmetric matrix.
-From that paper we adopt the model of computation, extension to approximation of matrix eigenvalues, and some patterns for estimation of the computational precision of root-finders, but otherwise use distinct approach and techniques.
-Namely, instead of Newton’s iterations in [76], we accelerate the classical subdivision root-finders, based on e/i tests.
-Its closest variations in [128, 101, 23, 24] still use inferior e/i tests, for which we proposed two fast novel approaches – based on approximation of Cauchy integrals and on root-lifting; these tests run much faster than the tests in the cited papers and, unlike the latter tests, can be applied to a black box polynomial.
-The papers [128, 101, 23, 24] apply functional iterations (such as Newton’s and Schröders) to accelerate convergence of subdivision process but we apply our novel techniques of disc compression instead.
-Estimating the bit operation complexity of our root-finders we use [134, Eqn.
-(9.4)] and [81, Thm.
+From the pioneering paper [59] we adopt the model of computation, extension of a black box polynomial root-finder to approximation of matrix eigenvalues, and the approach to estimation of the computational precision of root-finders, but otherwise we use distinct approach and techniques.
+Estimating the bit operation complexity of our root-finders we use [106, Eqn.
+(9.4)] and [64, Thm.
 2] for multipoint polynomial evaluation.
-[^8] The ingenious algorithm of Moroz supporting his [81, Thm.
+[^11] The ingenious algorithm of Moroz supporting his [64, Thm.
 2] departs from the simple observation that a polynomial $p(x-c)\mod(x-c)^{\ell-1}$ approximates $p(x)$ in the disc $D(c,\rho)$ within $O(\rho^{\ell+1})$ for small $\rho$.
 To approximate $p(c)$ for any $c\in D(0,1)$ Moroz used $\tilde{O}((\ell+\tau)d)$ bit operations to compute a set of complex $c_{i}$ and positive $\rho_{i}$ for $i=1,\dots,K$ and $K=O(d/\ell)$ such that $D(0,1)\subseteq\cup_{i=1}^{K}D(c_{i},\rho_{i})$; then he ensured approximation to $p(c)$ for every point $c\in D(0,1)$ within $\epsilon:=3\sum_{{}_{i}=0}^{d}|p_{i}|/2^{\ell}$ by means of approximation within $\epsilon/2$ of the lower degree polynomial $p(x-c_{i})\mod(x-c)^{\ell+1}$ for an appropriate $i$.
-At this point, application of [68, Alg.
+At this point, application of [52, Alg.
 5.3 and Thm.
 3.9] yields new record fast algorithm for multipoint approximation of a polynomial given with its coefficients.
-The root-finders of Moroz in [81] and Imbach and Moroz in [59] use $\tilde{O}(bd)$ bit operations for polynomials given with their coefficients and having only well-conditioned zeros.
-In spite of the appeal to the concept of condition number, used for the design and analysis of polynomial root-finding in [14, 30, 22], the upper bound on the condition numbers of all zeros of $p(x)$ in [81, 59] is a little stronger restriction than the bound on the cluster sizes of Def.
-24.1, under which [96, 103] proved the same complexity bound.
-Unlike our present root-finders, the root-finders of [81, 59] are not efficient for the worst case inputs and extensively use the coefficients.
-5.3 Implementation: State of the Art
+Ignoring Kirinnis’s paper [52], Moroz cited a rediscovery of his result by Kobel and Sagraloff, made two decades later.
+Instead of applying high order Newton’s iterations of [59], we accelerate classical subdivision root-finders.
+The closest earlier variations in [100, 82, 16, 17] use mach slower e/i tests, involving coefficients of $p$.
+Two our fast novel e/i tests – based on root-lifting and approximation of Cauchy integrals, respectively, work for a black box polynomial.
+The algorithms of [100, 82] dramatically accelerate convergence of subdivision process by means of applying Newton’s or Schröder’s iterations (see Fig.
+3).
+They converge with superlinear rate to the convex hull of all roots lying in $C$.
+The diameter of the minimal covering disc, however, can be too large, and then the papers [100, 82] apply subdivision iterations again.
+The resulting root-finders recursively intertwine subdivision iterations with Newton’s or Schröder’s, and this complicates the complexity estimates for a root-finder.
+[100, 82] have only estimated the number of ops but not of the bit operations involved.
+[16, 17] estimated the bit operation complexity in terms of the unknown minimal distance between the unknown roots.
+Specified in terms of $b$ and $d$, the estimates turned out to be inferior to ours by a factor of $d$ (see Remark 4.2 and Sec.
+10.4).
+Instead of Newton’s or Schröder’s iterations, one can apply the algorithm of [106].
+It first splits $p(x)$ into the product of two polynomials $f(x)$ and $g(x)$ whose zero sets lie in and outside the minimal disc covering the component $C$, respectively.
+Then it recursively splits the polynomial $p(x)$ into factors $f=f(x)$ and $g=g(x)$.
+The resulting algorithm of [106] is slower than ours by a factor of $d$.
+It has been accelerated to a near-optimal level in [78, 84], although that root-finder is much involved and extensively uses the coefficients of $p$, like the root-finders of [106, 100, 82, 16, 17].
+[64, 44] presented efficient root-finders for polynomials having only well-conditioned zeros, but not for worst case polynomials $p(x)$.
+Furthermore, the algorithms of [64, 44] extensively use the coefficients of $p(x)$.
+5.3 Implementation: State of the Art (briefly)
 The near-optimal but quite involved algorithm of [78, 84] has never been implemented, while the slower algorithm of [106] has been implemented by Xavier Gourdon in [34] for the Magma and PARI/GP computer algebra systems.
-By now it is superseded by implementation of various non-optimal complex polynomial root-finders such as MATLAB “roots”, Maple “solve”, CCluster, and MPSolve; the latter one relies on Ehrlich’s iterations and is user’s current choice for approximation of all $d$ complex roots (in Sec.
-26.4 we list some competitive implementations of Newton’s iterations).
+By now it is superseded by implementations of various non-optimal complex polynomial root-finders such as MATLAB “roots”, Maple “solve”, CCluster, and MPSolve; the latter one relies on Ehrlich’s iterations and is user’s current choice for approximation of all $d$ complex roots (see some competitive implementations of Newton’s iterations in [115, 102, 68]).
 Current user’s choice for real root-finding is the algorithm of [54], although it has already been superseded in [47] according to extensive tests for standard test polynomials.
-We present algorithms that are rather simple for implementation and promise to compete for user’s choice.
+We present algorithms that promise to compete for user’s choice.
 Already incorporation in 2020 of some of our techniques into the previous best implementation of subdivision root-finder [49] has enabled its noticeable acceleration (see [45]).
-The more advanced publicly available implementation in [63] “relies on novel ideas and techniques" from [112] (see [63, Sec.
+The more advanced publicly available implementation in [48] “relies on novel ideas and techniques" from [93] (see [48, Sec.
 1.2]).
-According to the results of extensive tests with standard test polynomials, the algorithm has performed dramatically faster than [64], has at least competed with user’s choice package MPSolve even for approximation of all $d$ zeros of a polynomial $p$ given with its coefficients, [^9] The computational cost of subdivision root-finders in a domain $\mathbb{D}$ as well as the root-finder of [96, 103] decreases at least proportionally to the number of roots in $\mathbb{D}$, while Ehrlich’s, Weierstrass’s, and a number of other functional iterations approximate the zeros of $p$ in a domain $\mathbb{D}$ almost as slow as all its $d$ complex roots.
+According to the extensive tests with standard test polynomials, the algorithm has performed dramatically faster than [49], has at least competed with user’s choice package MPSolve even for approximation of all $d$ zeros of a polynomial $p$ given with its coefficients, [^12] The computational cost of subdivision root-finders in a domain $\mathbb{D}$ as well as the root-finder of [78, 84] decreases at least proportionally to the number of roots in $\mathbb{D}$, while Ehrlich’s, Weierstrass’s, and a number of other functional iterations approximate the zeros of $p$ in a domain $\mathbb{D}$ almost as slow as all its $d$ complex roots.
 Thus already implementation in [49] of subdivision iterations, slightly outperformed MPSolve for root-finding in a domain $\mathbb{D}$ containing a small number of roots although is by far inferior for the approximation of all $d$ roots.
-and tended to supersede MPSolve more and more significantly as the degree $d$ of $p$ grew large.
-There is still a large room for further improvement because the root-finder of [63] inherits some earlier sub-algorithms from [23, 64], which are slower by order of magnitude than our current ones.
-Next we comment on some promising directions to further improvement of [63], but also notice that adjustment of our near-optimal subdivision root-finder in Remark 2.1 to approximation of real and near-real roots and consequently of matrix eigenvalues should be at least competitive with the current user’s choice [70] for real root-finding as well as with its significant improvement in [62]; moreover, it more efficiently counters the impact of the input and rounding errors.
-5.4 Further progress and promise
+and superseded MPSolve more and more significantly as the degree $d$ of $p$ grew large.
+There is still a large room for further improvement because the root-finder of [48] inherits some earlier sub-algorithms from [16, 49], which are slower by order of magnitude than our current ones (see the next subsection).
+Application of our near-optimal subdivision root-finder in Remark 2.2 to approximation of real and near-real roots promises to be at least competitive with user’s current choice [54] and its improvement in [47].
+It can be applied to a black box polynomial, counters the impact of input and rounding errors, and can be extended to approximation of real and near-real matrix eigenvalues.
+5.4 Further progress and promises
 (A) Acceleration of e/i tests.
 E/i tests are basic ingredient of subdivision root-finders.
-[60, 63] used the Cauchy-based e/i tests of our Part II and their heuristic acceleration.
-It is still a challenge to compare their empirical performance with that of the formally supported e/i tests of our Part III and Sec.
-23.4, based on root-lifting.
-Furthermore, we can try to exploit potential benefits of fast numerical approximation of Cauchy integrals with the Exponentially Convergent Adaptive Trapezoidal Quadrature (see [146]), efficiently implemented in Boost C++ libraries.
+[45, 48] used the Cauchy-based e/i tests of our Part III and their heuristic acceleration.
+We can devise and test similar heuristic acceleration of the formally supported e/i tests of our Part II, based on root-lifting.
+Namely, we can merely apply Alg.
+7.2 with an integer $q$ decreased dynamically from $m+1$, as long as the exclusion claim remains empirically correct.
 (B) Acceleration of the compression of a disc.
-We assume that the unit disc $D(0,1)$ contains $m$ roots and is isolated from the external $d-m$ roots, and we compress that disc towards the minimal disc containing the same $m$ roots.
-We reduce this task essentially to (i) computing a point in or near the convex hull of the root set in $D(0,1)$ and (ii) approximation of the $m$th smallest root radius from a given point in $D(0,1)$.
-For task (i) we dramatically accelerate our algorithm of [60, 63] because our algorithms of Secs.
-23.2 and 23.3 for task (i), using Schröder’s iterations, deterministic, and formally supported, run dramatically faster than those of [60, 63], based on approximation of the average of all roots in $D(0,1)$.
-For task (ii) we can apply our Las Vegas randomized algorithms of Secs.
-8 and 9 combined or the alternative of [60, 63], based on the algorithm of our Sec.
-15.1, which reduces this task to performing a constant number of e/i tests.
-Both approaches support the same asymptotic cost bound up to polylog factors, but fro $m\gg\log^{2}(d)$ the second approach is faster by a factor of $m/\log^{2}(d)$ where it uses fast heuristic e/i tests.
-Such an improvement, however, may be lost in practice because of the impact of polylog factors.
-Implementation of compression in [60, 63] is slowed down with verification based on Pellet’s theorem, but we can drop this verification because we readily detect highly unlikely output errors of our Las Vegas root-finders.
+Assuming that the unit disc $D(0,1)$ contains $m$ roots and is isolated from the $d-m$ external roots, we compress that disc towards an $\epsilon$-disc or a rigid disc, in both cases losing no roots.
+Algorithmic extras 1 (for compression of a disc with a small number of roots) and 2 of Sec.
+21 can help strengthen this approach.
+Like us, [48] reduces compression to (a) non-costly root-counting in an isolated disc, (b) computing a point $c$ lying in or near the MCD, that is, the minimal covering disc  [^13] One may use the convex hull instead of the minimal covering disc.
+of the root set in $D(0,1)$, and (c) approximation of the $m$th smallest root radius from $c$.
+At stage (b) [48] computes a close approximation of the sum $s_{1}$ of the $m$ roots lying in the disc $D(0,1)$ and then outputs $c=\frac{s_{1}}{m}$.
+In Sec.
+17 we greatly accelerate this algorithms in the case where the radius of the output disc is small.
+At stage (c) [48] applies the algorithm of our Sec.
+18, which performs a constant but fairly large number of e/i tests in a disc.
+One may test this approach versus our alternative of Sec.
+13.4 using a single e/i test for an annulus and versus the combination of the Las Vegas randomized algorithm of Sec.
+13 with BoL of Sec.
+16.
+Finally, implementation of compression in [45, 48] is slowed down at the stage of verification based on Pellet’s theorem, but we can drop this verification because we readily detect highly unlikely output errors of our root-finders at a dominated cost (see Sec.
+3.2).
 (C) Fast approximation of the eigenvalues of a matrix.
-[20] explored reduction of approximation of the eigenvalues of a matrix to root-finding by means of Ehrlich’s iterations, reduced to recursive computations of NIR$(x)$, and showed benefits of this approach.
-We can readily extend this reduction and its benefits by using subdivision iterations instead of Ehrlich’s.
-Furthermore, win the case of subdivision we obtain additional benefits, e.g., our algorithms are much faster for approximation of a small number of eigenvalues that lie in a fixed domain and unlike Ehrlich’s iterations isolate eigenvalues, rather than just approximate them.
-(D) Acceleration of real and near real root-finding.
-In Remark 2.1 we adjust our near-optimal subdivision iterations to real and near-real root-finding, which counters the impact of input and rounding errors and can be readily extended to approximation of the eigenvalues of a matrix.
+Our root-finders are essentially reduced to the evaluation of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ at sufficiently many points $x$, and we can extend them to approximation of the eigenvalues of a matrix or a matrix polynomial by following [59] or based on the following alternative expressions:
+$\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}((xI-M)^{-1})}$
+where $M$ is a matrix, while $p(x)=\det(xI-M)$ (cf. (3.1)) and
+$\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}(M^{-1}(x)M^{\prime}(x))}$
+where $M(x)$ is a matrix polynomial and $p(x)=\det(M(x))$ [13, Eqn.
+(5)].
+[^14] These expressions have been obtained based on Jacobi’s formula ${\rm d}(\ln(\det(A)))={\rm trace}(A^{-1}{\rm d}A),$ which links differentials of a nonsingular matrix $A$ and of its determinant.
+These expressions reduce approximation of the eigenvalues of a matrix or a matrix polynomial to its inversion.
+One can use the matrix inversion algorithm and its bit operation cost estimates of [110] towards alternative derivation of our Main Theorem 2 for any input matrix or matrix polynomial, but the inversion can be greatly accelerated in the highly important case where an input matrix or matrix polynomial can be inverted fast, e.g., is data sparse.
+Bini and Noferini in [13] proposed and explored such a reduction of approximation of all $d$ eigenvalues to root-finding by means of Ehrlich’s iterations and showed various benefits of this approach, but we can readily combine this reduction with subdivision iterations instead of Ehrlich’s and then obtain additional benefits: our algorithms (a) are much faster for approximation of a small number of eigenvalues that lie in any fixed domain and (b) isolate the eigenvalues instead of just approximating them.
+(D) Acceleration of real and near-real root-finding.
+In Remark 2.2 we adjust our near-optimal subdivision iterations to real and near-real root-finding, which counters the impact of input and rounding errors.
 (E) Acceleration by using fast multipoint polynomial evaluation.
@@
 5.5 Organization of the rest of the paper
-In Parts II and III we devise and analyze our two alternative e/i tests.
-Deterministic e/i test of Part III is as fast as the randomized one of Part II.
-We use Las Vegas randomization in both parts to bound computational precision and bit-operation cost, and in Part III we yield acceleration by a factor of $d$ under the Random Root Model.
-The bit-operation complexity bounds of Parts II and III are inferior by a factor of $b$ to that of [96, 97, 103], but we match it in Part IV by means of our acceleration of the algorithms of both Parts II and III based on compression of a disc without losing its roots.
-In Part V we cover some important root-finding subjects that involve no substitution iterations, namely, factorization, root isolation, and functional iterations.
+In Parts II and III we devise and analyze our two e/i tests.
+We use Las Vegas randomization in both parts to bound computational precision and bit-operation cost.
+In Part II the same randomization helps us devise fast e/i tests.
+In Part III we support that step with additional randomization.
+In Part II we further accelerate by a factor of $d$ e/i tests and overall root-finding under the Random Root Model.
+We make our presentation in Part III largely independent of Part II by reusing some definitions, results, and arguments.
+In Part IV we devise and analyze our compression algorithm and incorporate it into the algorithms of Parts II and III – to accelerate them by a factor of $b/\log(b)$ and to ensure pairwise isolation of the output roots and their clusters.
+In Part V we cover (i) low bounds on the bit-operation complexity of polynomial factorization, root-finding, and root isolation, (ii) acceleration with FMM (cf.
+Remark 3.1) and initialization of root-finding by means of functional iterations, and (iii) some additional algorithms for root radii approximation.
+PART II: Classical subdivision root-finder with novel e/i tests based on root-lifting
+Organization of Part II: We devote the next section to background.
+In Sec.
+7 we cover our e/i tests.
 In Secs.
-24 and 25 we outline the fundamental results of Schönhage from [134] on polynomial factorization, its links to root-finding and root-isolation, and some relevant bit-complexity estimates, including lower bounds.
+8 and 10 we estimate computational precision of our root-finders and their bit operation cost, respectively.
 In Sec.
-26 we discuss root-finding by means of functional iterations, namely, Newton’s, Weierstrass’s (aka Durand-Kerner’s), and Ehrlich’s (aka Aberth’s), and their acceleration by means of incorporation of Fast Multipole Algorithm (FMM) and fast multipoint polynomial evaluation as well as their initialization both where the coefficients of $p$ are unknown or given to us.
-5.6 Independence of Part III from Part II
-In Parts II an III we solve Problems 0, 1, 0, and 1 in two different ways.
-In Part II we rely on approximation of a Cauchy integral associated with root-counting in a disc on the complex plane.
-In Part III we rely on lifting the roots to a fixed power $h=O(\log(d))$.
-In addition, in Part III, unlike Part II, we devise a faster root-finder under a random root model.
-We make our presentation in Part III independent of Part II; in particular we include again some definitions, figures, and theorems.
-PART II: Classical subdivision root-finder with novel exclusion/inclusion (e/i) tests based on approximation of Cauchy integrals
-Organization of the presentation in Part II: We devote the next section to some basic definitions.
-In Sec.
-7 we study root-counting in a disc on the complex plane by means of approximation of a Cauchy integral over its boundary circle by finite sums.
-[^10] We do not use the results of Sec.
-7.5, but they can be of independent interest.
-In Sec.
-8 we apply this study to devise and analyze root-counters and e/i tests in a disc on the complex plane.
-In Sec.
-9 we extend our e/i tests to estimating root radii, that is, the distances from a fixed complex point to the roots.
-In Secs.
-10 we estimate precision of computing of our root-finders.
-We devote short Sec.
-11 to their extension to approximation of the eigenvalues of a matrix or a matrix polynomial.
-In Sec.
-12 we estimate the bit operation complexity of our root-finders assuming that $p(x)$ is a general polynomial of (1.1) represented with its coefficients.
-6 Background: basic definitions
-  • “Wlog" stands for “without loss of generality".
-  • $S(c,\rho):=\{x:~{}|\Re(c-x)|\leq\rho$, $|\Im(c-x)|\leq\rho\},~{}D(c,\rho):=\{x:~{}|x-c|\leq\rho\}$,
-    $C(c,\rho):=\{x:~{}|x-c|=\rho\}$, and $A(c,\rho,\rho^{\prime}):=\{x:~{}\rho\leq|x-c|\leq\rho^{\prime}\}$
-    denote a square, a disc, a circle (circumference), and an annulus (ring) on the complex plane, respectively.
-  • $\Delta(\mathbb{S})$, $X(\mathbb{S})$, $CH(\mathbb{S})$, and #$\#(\mathbb{S})$ denote the diameter, the root set, the convex hull, and the index (root set’s cardinality) of a set $\mathbb{S}$ in the complex plane, respectively.
-  • A disc $D(c,\rho)$, a circle $C(c,\rho)$, or a square $S(c,\rho)$ is said to be $\theta$-isolated (from external roots) for $\theta>1$ if $X(D(c,\rho))=X(D(c,\theta\rho))$, $X(C(c,\rho))=X(A(c,\rho/\theta,\rho/\theta))$, or $X(S(c,\rho))=X(S(c,\theta\rho))$, respectively.
-  • The largest upper bound on such a value $\theta$ is said to be the isolation of the disc $D(c,\rho)$, the circle $C(c,\rho)$, or the square $S(c,\rho)$, respectively (see Fig.
-4), and is denoted $i(D(c,\rho))$, $i(C(c,\rho))$, and $i(S(c,\rho))$, respectively.
-  • $r_{1}(c,t)=|y_{1}-c|,\dots,r_{d}(c,t)=|y_{d}-c|$ in non-increasing order are the $d$ root radii, that is, the distances from a complex point $c$ to the zeros $y_{1}\dots,y_{d}$ of a $d$th degree polynomial $t(x)$.
-$r_{j}(c):=r_{j}(c,p)$, $r_{j}:=r_{j}(0)$ for $j=1,\dots,d$.
-  • Generalize Def.
-2.1 of $\sigma$-soft e/i tests as follows: for $p(x)$ of (1.1), $\sigma>1$, and integers $\ell$ and $m$ such that $1\leq\ell\leq m\leq d$ and #$\#(D(0,1))\leq m$, a $\sigma$-soft $\ell$-test, or just $\ell$-test for short (so that e/i test turns into 1-test), either outputs 1 and stops if it detects that $r_{d-\ell+1}\leq\sigma$, that is, #$\#(D(0,\sigma))\geq\ell$, or outputs 0 and stops if it detects that $r_{d-\ell+1}>1$, that is, [^11] Bounds $r_{d-\ell+1}\leq\sigma$ and $r_{d-\ell+1}>1$ can hold simultaneously (cf.
-Fig.
-1b), but as soon as an $\ell$-test verifies any of them, it stops without checking if the other bound also holds.
-#$\#(D(0,1))<\ell$.
-$\ell$-testc,ρ, aka $\ell$-test for the disc $D(c,\rho)$, is an $\ell$-test applied to the polynomial $t(y)$ of (1.3).
-Observation 6.1.
-Eqn. (1.3) maps the zeros $z_{j}$ of $p(x)$ into the zeros $y_{j}=\frac{z_{j}-c}{\rho}$ of $t(y)$, for $j=1,\dots,d$, and preserves the index #$\#(D(c,\rho))$ and the isolation $i(D(c,\rho))$.
-$\rho$$\rho\theta$$\rho\theta^{2}$c********
-Figure 4: The roots are marked by asterisks. The red circle $C(c,\rho\theta)$ and the disc $D(c,\rho\theta)$ have isolation $\theta$. The disc $D(c,\rho)$ has isolation $\theta^{2}$.
-7 Cauchy root-counting in a disc
-7.1 The power sums of the roots in a disc
-Recall the following equations generalizing (3.4) (cf. [1]):
-[…]
-For $\mathcal{D}=D(0,1)$ and $\zeta$ of (3.5) we will approximate $s_{h}$ with finite sums
-$s_{h,q}:=\frac{1}{q}\sum_{g=0}^{q-1}\zeta^{(h+1)g}~{}\frac{p^{\prime}(\zeta^{g})}{p(\zeta^{g})}~{}{\rm for}~{}h=0,1,\dots,q-1$ (7.2)
-and recall that $s_{0}(D(0,1))=\#(D(0,1))$ (cf. (3.4)), while an e/i test is just the decision whether #$\#(D(0,1))>0$.
-7.2 Cauchy sums in the unit disc as weighted power sums of the roots
-Cor. 7.1 of the following lemma expresses the $h$th Cauchy sum $s_{h,q}$ as the sum of the $h$th powers of the roots $z_{j}$ with the weights $\frac{1}{1-x^{q}}$.
-Lemma 7.1.
-For a complex $z$, two integers $h\geq 0$ and $q>1$, and $\zeta$ of (3.5), it holds that
-$\frac{1}{q}\sum_{g=0}^{q-1}\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=\frac{z^{h}}{1-z^{q}}.$ (7.3)
-Proof.
-First let $|z|<1$ and obtain
-$\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=\frac{\zeta^{hg}}{1-\frac{z}{\zeta^{g}}}=\zeta^{hg}\sum_{u=0}^{\infty}\Big{(}\frac{z}{\zeta^{g}}\Big{)}^{u}=\sum_{u=0}^{\infty}\frac{z^{u}}{\zeta^{(u-h)g}}.$
-The equation in the middle follows from Newman’s expansion
-$\frac{1}{1-y}=\sum_{u=0}^{\infty}y^{u}~{}{\rm for}~{}y=\frac{z}{\zeta^{g}}.$
-We can apply it because $|y|=|z|$, while $|z|<1$ by assumption.
-Sum the above expressions in $g$, recall that
-$\zeta^{u}\neq 1~{}{\rm and}~{}\sum_{s=0}^{q-1}\zeta^{u}=0~{}{\rm for}~{}0<u<q,~{}\zeta^{q}=1,$ (7.4)
-and deduce that
-$\frac{1}{q}\sum_{g=0}^{q-1}\frac{z^{u}}{\zeta^{(u-h)g}}=z^{u}$
-for $u=h+ql$ and an integer $l$, while $\sum_{g=0}^{q-1}\frac{z^{u}}{\zeta^{(u-h)g}}=0$ for all other values $u$. Therefore
-$\frac{1}{q}\sum_{g=0}^{q-1}\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=z^{h}\sum_{l=0}^{\infty}z^{ql}.$
-Apply Newman’s expansion for $y=z^{q}$ and deduce (7.3) for $|z|<1$. By continuity extend (7.3) to the case of $|z|=1.$
-Now let $|z|>1$. Then
-$\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=-\frac{\zeta^{(h+1)g}}{z}~{}~{}\frac{1}{1-\frac{\zeta^{g}}{z}}=-\frac{\zeta^{(h+1)g}}{z}\sum_{u=0}^{\infty}\Big{(}\frac{\zeta^{g}}{z}\Big{)}^{u}=-\sum_{u=0}^{\infty}\frac{\zeta^{(u+h+1)g}}{z^{u+1}}.$
-Sum these expressions in $g$, write $u:=ql-h-1$, apply (7.4), and obtain
-$\frac{1}{q}\sum_{g=0}^{q-1}\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=-\sum_{l=1}^{\infty}\frac{1}{z^{ql-h}}=-z^{h-q}\sum_{l=0}^{\infty}\frac{1}{z^{ql}}.$
-Apply Newman’s expansion for $y=z^{q}$ again and obtain
-$\frac{1}{q}\sum_{g=0}^{q-1}\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=-\frac{z^{h-q}}{1-\frac{1}{z^{q}}}=\frac{z^{h}}{1-z^{q}}.$
-Hence (7.3) holds in the case where $|z|>1$ as well. ∎
-Sum equations (7.3) for $z=z_{j}$ and $j=1,\dots,d$, combine Eqns. 3.1 and (7.2), and obtain
-Corollary 7.1.
-For the roots $z_{j}$ and all $h$, the Cauchy sums $s_{h,q}$ of (7.2) satisfy $s_{h,q}=\sum_{j=1}^{d}\frac{z_{j}^{h}}{1-z_{j}^{q}}$ unless $z_{j}^{q}=1~{}{\rm for~{}some}~{}j.$
-7.3 Approximation of the power sums in the unit disc
-Cor. 7.1 implies that the values $|s_{h,q-}s_{h}|$ decrease exponentially in $q-h$ if the unit disc $D(0,1)$ is isolated.
-Theorem 7.1.
-[134, Eqn.
-(12.10)].
-[^12] Unlike the proof in [134], we rely on Cor.
-7.1 of independent interest.
-Let the unit disc $D(0,1)$ be $\theta$-isolated and let $d_{\rm in}$ and $d_{\rm out}=d-d_{\rm in}$ denote the numbers of the roots lying in and outside the disc $D(0,1)$, respectively.
-Then
-$|s_{h,q}-s_{h}|\leq\frac{d_{\rm in}\theta^{-h}+d_{\rm out}\theta^{h}}{\theta^{q}-1}\leq\frac{d\theta^{h}}{\theta^{q}-1}~{}{\rm for}~{}h=0,1,\dots,q-1.$ (7.5)
-Proof.
-Enumerate the roots $z_{j}$ so that $|z_{j}|<1$ if and only if $j\leq d_{\rm in}$; then Cor. 7.1 implies that
-$s_{h,q}-s_{h}=s_{h,q,{\rm in}}+s_{h,q,{\rm out}}$ (7.6)
-where
-$s_{h,q,{\rm in}}=\sum_{j=1}^{d_{\rm in}}z_{j}^{h}\Big{(}\frac{1}{1-z_{j}^{q}}-1\Big{)}~{}{\rm and}~{}s_{h,q,{\rm out}}=\sum_{j=d_{\rm in}+1}^{d}\frac{z_{j}^{h}}{1-z_{j}^{q}}=\sum_{j=d-d_{\rm out}+1}^{d}\frac{z_{j}^{h}}{1-z_{j}^{q}}.$ (7.7)
-Notice that $\frac{1}{1-z_{j}^{q}}-1=\frac{z^{q}}{1-z_{j}^{q}}$, and so $|z_{j}^{h}(\frac{1}{1-z_{j}^{q}}-1)|\leq\frac{\theta^{-h}}{\theta^{q}-1}$ for $z_{j}\in D(0,1)$, while $\Big{|}\frac{z_{j}^{h}}{1-z_{j}^{q}}\Big{|}\leq\frac{\theta^{h}}{\theta^{q}-1}$ for $z_{j}\notin D(0,1)$.
-Combine these bounds with Eqns.
-(7.6) and (7.7) and obtain (7.5).
-∎
-By choosing $q$ that exceeds $\log_{\theta}(2d+1)$ we ensure that $|s_{0}-s_{0,q}|<1/2$.
-Example 7.1.
-Let the unit circle $C(0,1)$ be 2-isolated. Then Thm. 7.1 implies that $|s_{0,q}-s_{0}|<1/2$ for $d\leq 1,000$ if $q=10$ and for $d\leq 1,000,000$ if $q=20$.
-Algorithm 7.1.
-INPUT: a black box polynomial $p$ of a degree $d$ and $\theta>1$.
-INITIALIZATION: Compute the integer $q=\lfloor\log_{\theta}(4d+2)\rfloor>\log_{\theta}(2d+1)$.
-COMPUTATIONS: Compute Cauchy sum $s_{0,q}$ and output an integer $\bar{s}_{0}$ closest to it.
-Observation 7.1.
-(i) Alg.
-7.1 evaluates $p(x)$ at $q=\lfloor\log_{\theta}(4d+2)\rfloor$ points.
-(ii) It outputs $\bar{s}_{0}=\#(D(0,1))$ if the circle $C(0,1)$ is $\theta$-isolated for $\theta>1$.
-(iii) If the algorithm outputs $\bar{s}_{0}>0$, then #$\#(D(0,\theta))>0$.
-Proof.
-Thm. 7.1 immediately implies claim (ii) but also implies that #$\#(D(0,1))>0$ if $\bar{s}_{0}>0$ unless #$\#(A(0,1/\theta,\theta))>0$. In both cases #$\#(D(0,\theta))>0$. ∎
-7.4 Extension to any disc
-Map (1.3), preserving #$\#(D(0,1))$ (cf. Observation 6.1), enables us to extend definition (7.2) of Cauchy sum to any disc $D(c,\rho)$ as follows:
-$s_{0,q}(p,c,\rho):=s_{0,q}(t,0,1):=\frac{\rho}{q}\sum_{g=0}^{q-1}\zeta^{g}~{}\frac{p^{\prime}(c+\rho\zeta^{g})}{p(c+\rho\zeta^{g})},$ (7.8)
-for $\zeta$ of (3.5), that is, $s_{h,q}(p,c,\rho)$ is the Cauchy sums $s_{h,q}(t,0,1)$ of the zeros of the polynomial $t(y)=p(\frac{y-c}{\rho})$ in the unit disc $D(0,1)$.
-Observation 7.2.
-Given a polynomial $p(x)$, a complex $c$, a positive $\rho$, a positive integer $q$, and the $q$-th roots of unity, we can compute $s_{0,q}(p,c,\rho)$ by means of the evaluation of NIR$(\zeta^{g})$ for $g=0,1,\dots,q-1$, reduced to evaluation of $p$ at $2q$ points, and in addition performing $q+1$ divisions, $12q-$ 1 multiplications, and $12q-$ 1 additions.
-Refer to Alg.
-7.1 applied to the polynomial $t(y)$ of (1.3) for an isolation parameter $\theta>1$ and $q=\lfloor\log_{\theta}(4d+2)\rfloor$ as Alg.
-7.1c,ρ,q and also as Alg.
-7.1 applied to $\theta$-isolated circle $C(c,\rho)$ for $2q$ evaluation points.
-Combine Observations 6.1 and 7.2 to extend Observation 7.1 as follows.
-Proposition 7.1.
-(i) Alg.
-7.1c,ρ,θ evaluates polynomial $p(x)$ at $q=\lfloor\log_{\theta}(4d+2)\rfloor$ points $x$ and in addition performs $25q-$ 2 ops.
-(ii) It outputs $\bar{s}_{0}=\#(D(c,\rho))$ if the circle $C(c,\rho)$ is $\theta$-isolated.
-(iii) If the algorithm outputs $\bar{s}_{0}>0$, then #$\#(D(c,\theta\rho))>0$.
-By applying Thm.
-7.1 we can estimate the errors $|s_{h,q}(t,0,1)-s_{h}(t,0,1)|$ of the approximation of the $h$th power sums $s_{h}(t,0,1)$ of the roots of $t(x)$ in the unit disc $D(0,1)$.
-Next we specify this for $h=1$.
-Proposition 7.2.
-(Cf. [63].) Let for $\theta>1$ a $\theta$-isolated disc $D(c,\rho)$ contain exactly $m$ roots. Fix a positive integer $q$ and define the Cauchy sum $s_{1,q}(p,c,\rho)$ of (7.8). Then
-$|mc+\rho s_{1,q}(p,c,\rho)-s_{1}(p,c,\rho)|\leq\frac{\rho d\theta}{\theta^{q}-1}.$
-7.5 The real part of the 0-th Cauchy sum
-Clearly, $|\Re(s_{0,q})-s_{0}|\leq|s_{0,q}-s_{0}|$ because $s_{0}$ is an integer.
-Next we estimate the bound $|\Re(s_{0,q})-s_{0}|$ for the unit disc $D(0,1)$, point out some tentative directions to algorithmic gain for root-finding, but leave further elaboration as a challenge for formal and empirical investigation.
-[^13] We will not use the results of this subsection, but they can be of independent interest.
-Recall that $s_{0,q}=\sum_{j=1}^{d}\frac{1}{1-z_{j}^{q}}$ unless $z_{j}^{q}=1$ for some $q$ (cf. Cor. 7.1), write $z:=z_{j}^{q}$ and estimate the contribution of a roots $z_{j}$ to $\Re(s_{0,q})$.
-Proposition 7.3.
-Let $z=\rho\cdot(\cos(\phi)+\sin(\phi)\bf i)$ for a non-negative $\rho$ and a real $\phi\in[0,2\pi)$. Then
-$\Re\Big{(}\frac{1}{1-z}\Big{)}=\frac{1-\rho\cos(\phi)}{1+\rho^{2}-2\rho\cos(\phi)},$ (7.9)
-$1-\Re\Big{(}\frac{1}{1-z}\Big{)}=\rho\frac{\rho-\cos(\phi)}{1+\rho^{2}-2\rho\cos(\phi)}.$ (7.10)
-Proof.
-Multiply both numerator and denominator of the fraction
-$\frac{1}{1-z}=\frac{1}{1-\rho\cdot(\cos(\phi)+\sin(\phi)\bf i)}$
-by $1-\rho\cdot(\cos(\phi)-\sin(\phi)\bf i)$, substitute $\cos^{2}(\phi)+\sin^{2}(\phi)=1$, and obtain (7.9). Then readily verify (7.10). ∎
-Corollary 7.2.
-Under the assumptions of Prop. 7.3 it holds that
-$\Re\Big{(}\frac{1}{1-z}\Big{)}\geq 0~{}{\rm if}~{}\rho\leq 1,$ (7.11)
-$\Re\Big{(}\frac{1}{1-z}\Big{)}=\frac{1}{2}~{}{\rm if}~{}\rho=1~{}{\rm and~{}if}~{}\phi\neq 0,$ (7.12)
-$\Re\Big{(}\frac{1}{1-z}\Big{)}\geq\frac{1-0.8\rho}{1+\rho^{2}-1.6\rho}~{}{\rm if}~{}\rho\leq 5/4~{}{\rm and~{}if}~{}\cos(\phi)\leq 0.8,$ (7.13)
-$\Re\Big{(}\frac{1}{1-z}\Big{)}=\frac{1}{1-\rho}~{}{\rm if}~{}\phi=0~{}{\rm and~{}if}~{}\rho\neq 1,$ (7.14)
-$\frac{1}{1+\rho}\leq\Big{|}\Re\Big{(}\frac{1}{1-z}\Big{)}\Big{|}\leq\frac{1}{|1-\rho|},$ (7.15)
-$\frac{\rho}{1+\rho}\leq\Big{|}1-\Re\Big{(}\frac{1}{1-z}\Big{)}\Big{|}\leq\frac{\rho}{|1-\rho|}.$ (7.16)
-Proof.
-Readily deduce (7.11), (7.12), and (7.14) from (7.9).
-Write
-$y:=\cos(\phi),~{}f(y):=\Re\Big{(}\frac{1}{1-z}\Big{)},~{}{\rm and}~{}g(y):=\frac{1}{f(y)}.$
-Then deduce from (7.9) that $g(y)=2+\frac{\rho^{2}-1}{1-\rho y}$ Hence $g^{\prime}(y)=\frac{(\rho^{2}-1)\rho}{(1-\rho y)^{2}}$ is a monotone increasing or monotone decreasing function, singular if and only if $y=\frac{1}{\rho}$.
-Extend these properties to $f(y)$ and $1-f(y)$ if $\rho<1$.
-This implies (7.13), (7.15), and (7.16) provided that $\rho<1$ because
-$f(1)=\frac{1}{1-\rho},~{}1-f(1)=\frac{\rho}{1-\rho},~{}f(-1)=\frac{1}{|1+\rho|},~{}1-f(-1)=\frac{\rho}{|1+\rho|},$
-and $f(0.8)=\frac{1-0.8\rho}{1+\rho^{2}-1.6\rho}$.
-The proof can be extended to the case where $\rho\geq 1$ except that the zero $y=\frac{1}{\rho}$ of $f(y)$ turns into a pole of $\frac{1}{f(y)}$.
-We complete the proof of the bounds (7.15) for all positive $\rho$ by observing that $f(\frac{1}{\rho})$ is undefined for $\rho=1$ and that otherwise $f(\frac{1}{\rho})=0$ and $11-f(\frac{1}{\rho})=$ 1.
-∎
-Next we outline but do not elaborate upon some potential links of Cor. 7.2 to root-counting. Let the unit disc $D(0,1)$ be $\theta$-isolated, combine bounds (7.15) and (7.16), and obtain
-$|\Re(s_{0,q})-s_{0}|\leq\frac{d}{\theta^{q}-1}.$
-The bound is similar to (7.5) for $h=0$ but is overly pessimistic: it is only reached where $\cos(\phi)=1$ for $z=z_{j}$ and all $j$.
-Equation (7.14) implies numerically unstable dependence of the fraction $\Re(\frac{1}{1-z})$ on $\rho$ where $\cos(\phi)\approx 1\approx\rho$, but bound (7.13) shows that outside a rather small neighborhood of the extreme values of $\cos(\phi)$ and $\rho$ the fraction is smooth and approaches 1/2 as $\rho\rightarrow 1$ and/or $\cos(\phi)\rightarrow-1$.
-Remark 7.1.
-In view of bound (7.13) random rotation of the circle $C(0,1)$ is a heuristic recipe for countering instability in $z$ of the fraction $\Re(\frac{1}{1-z})$ because $\cos(\phi)>0.8$ for only one of four equally spaced points $\phi$ on the unit circle $C(0,1)$ and hence occurs with a probability less than 1/4 for a random choice of $\phi$ under the uniform probability distribution in the range $[0,2\pi)$.
-Moreover, random rotation of the circle can limit the adverse and numerically unstable impact of boundary roots, lying on or near the boundary circle $C(0,1)$, on an e/i test.
-We conclude with a simple but surprising result on root-counting on an isolated circle. We do not use it in this paper.
-Corollary 7.3.
-Suppose that $p(1)\neq 0$ and that $p(x)$ has precisely $m$ roots in an annulus $A(0,\frac{1}{\theta},\theta)$ for a fixed $\theta>1$ and that all of them lie on the unit circle $C(0,1)$.
-Let $s^{\prime}$ and $s^{\prime\prime}$ denote the 00th Cauchy sums for the discs $D(0,\frac{1}{\theta})$ and $D(0,\theta)$, respectively, and write $s_{0,q}=s^{\prime\prime}-s^{\prime}$.
-Then #$(C(0,1))=2\Re(s_{0,q})$.
-Proof.
-Combine Cor. 7.1 for $h=0$ with equation (7.12). ∎
-8 Root-counting and $\ell$-tests
-8.1 Overview
-We seek fast root-counters and $\ell$-tests based on computing Cauchy sum $s_{0,q}$ for a smaller integer $q$, but the output can deceive us if we decrease $q$ too much: the polynomial $p(x):=x^{d}-dx+w$ has zero near the origin if $w\approx 0$, while NIR$(x)=p^{\prime}(x)/p(x)$ vanishes for $x$ equal to any $(d-1)$st root of unity.
-Such deceptive e/i tests have been very rare guests in the extensive experiments in [60, 63] for $q$ of order $\log^{2}(d))$ and were completely excluded where the values $|s_{h,q}|$ were required to be small for $h=0,1,2$, but formal support for this empirical behavior is still a challenge.
-In the next two subsections we formally support Las Vegas randomized root-counters and $\ell$-tests that evaluate polynomial $p$ at $O(m\log(d))$ points.
-Combine this estimate with Cor. 2.2 for $k$ of order $b$ and obtain Las Vegas randomized solution of Problems 0, 1, 0, and 1 at the cost of evaluation of the polynomial $p$ at $O(m^{2}b)$ points.
-8.2 A basic randomized root-counter
-Next we fix a parameter $\gamma>1$ and devise a root-counter for a disc $D(c,\rho)$ with a fixed center $c$ (wlog we let $c=0$) and a radius $\rho$ sampled at random in a fixed small range.
-Then we prove that under such a Las Vegas randomization the test fails with a probability at most $1/\gamma$.
-Algorithm 8.1.
-Basic randomized root-counter.
-INPUT: $\gamma\geq 1$ and a $d$-th degree black box polynomial $p$ having at most $m$ zeros in the disc $D(0,\sqrt{2})$.
-INITIALIZATION: Sample a random value $w$ in the range $[0.2,0.4]$ under the uniform probability distribution in that range and write $\rho:=2^{w}$ and $q=\lfloor 10m\gamma\log_{2}(4d+2)\rfloor$.
-COMPUTATIONS: Apply Alg. 7.10,ρ,q.
-Proposition 8.1.
-Alg. 8.1 queries evaluation oracle at $q=\lfloor 10m\gamma\log_{2}(4d+2)\rfloor$ locations and outputs the value $\bar{s}_{0}=\#(D(0,\rho))$ with a probability at least $1-1/\gamma$.
-Proof.
-Readily verify the bound on the number of query locations claimed in the proposition.
-By assumption, the annulus $A(0,1,\sqrt{2})$ contains at most $m$ roots.
-Hence at most $m$ root radii $r_{j}=2^{e_{j}}$, for $j=1,\dots,m^{\prime}\leq m$, lie in the range $[1,\sqrt{2}]$, that is, $0\leq e_{j}\leq 0.5$ for at most $m^{\prime}\leq m$ integers $j$.
-Fix $m^{\prime}$ intervals, centered at $e_{j}$, for $j=1,\dots,m^{\prime}$, each of a length at most $1/(5m^{\prime}\gamma)$.
-Their overall length is at most $1/(5\gamma)$.
-Let $\mathbb{U}$ denote their union.
-Sample a random $u$ in the range $[0.2,0.4]$ under the uniform probability distribution in that range and notice that Probability$(u\in\mathbb{U})\leq 1/\gamma$.
-Hence with a probability at least $1-1/\gamma$ the circle $C(0,\rho)$ is $\theta$-isolated for $\theta=2^{\frac{1}{10m\gamma}}$, in which case Alg.
-8.1 outputs $\bar{s}_{0}=s_{0}=\#(D(0,\rho))$ by virtue of claim (ii) of Prop.
-7.1.
-∎
-8.3 Refining the bound on the error probability
-When we apply Alg.
-8.1 $v$ times, the number of query locations of the evaluation oracle grows linearly – by a factor of $v$, while the bound on Las Vegas error probability of outputting errors decreases exponentially – from $1/\gamma$ to $1/\gamma^{v}$.
-First let $\ell=1$ or $\ell=m$ and then again reduce the root radius approximation problem to the decision problem of $\ell$-test.
-Namely, narrow Alg.
-8.1 to an $\ell$-test for a fixed $\ell$ in the range $1\leq\ell\leq m$, so that with a probability at least $1-1/\gamma$ this $\ell$-test outputs 0 if #$\#(D(0,1))<\ell$ and outputs 1 if #$\#(D(0,2^{0.5}))\geq\ell$.
-Moreover, output 1 is certified by virtue of claim (iii) of Prop.
-7.1 if $\ell=1$; similarly the output 0 is certified if $\ell=m$.
-Next extend the $\ell$-test by means of applying Alg. 8.1 for $v$ iid random variables $w$ in the range $0.2\leq w\leq 0.4$. Call this $\ell$-test Alg. $\ell$.
-[…]
-Recall Prop. 8.1 and then readily verify the following proposition.
-Proposition 8.2.
-(i) For two integers $1\leq\ell\leq m$ and $v\geq 1$ and a real $\gamma\geq 1$, Alg.
-$\ell$ evaluates $p(x)$ at $\lfloor 10m\gamma\log_{2}(4d+2)\rfloor v$ points; this evaluation involves more ops than the other stages of the algorithm, even in the cases of the Mandelbrot and sparse polynomials $p$ (cf.
-Observation 7.2).
-(ii) Output 1 of Alg.
-8.1v,1 and the output 0 of Alg.
-8.1v,$m$ are certified.
-(iii) Output 0 of Alg.
-8.1v,1 and output 1 of Alg.
-8.1v,$m$ are correct with a probability at least $1-1/\gamma^{v}$.
-The proposition bounds the cost of $\ell$-tests for $\ell=1$ (e/i tests) and $\ell=m$.
-This is sufficient for our root-finding, but our next natural extension covers $\ell$-tests for any integer $\ell$ in the range $1<\ell<m$.
-Assign the values 0 or 1 to the output integers $g$ as follows: compute the average of the $v$ outputs in $v$ independent applications of the test and output $g$ denoting the integer closest to that average; if the average is 1/2, write $g=0$ to break ties.
-This extension of Alg. $\ell$ is only interesting for $1<\ell<m$; otherwise its cost/error estimates are inferior to those of Prop. 8.2.
-Proposition 8.3.
-For an integer $\ell$ such that $1<\ell<m\leq d$ and real $v\geq 1$ and $\gamma>1$, Algs.
-$\ell$ queries evaluation oracle at $\lfloor 10m\gamma\log_{2}(4d+2)\rfloor v$ locations, and its output is correct with a probability at least $1-(4/\gamma)^{v/2}$.
-Proof.
-Since $D(0,1)\subset D(0,\sqrt{2})$, we only consider the following three cases:
-(i) #$\#(D(0,1))<\ell$ and #$\#(D(0,2^{0.5}))\geq\ell$,
-(ii) #$\#(D(0,2^{0.5}))<\ell$, and then also #$\#(D(0,1))<\ell$,
-(iii) #$\#(D(0,1))\geq\ell$, and then also #$\#(D(0,2^{0.5}))\geq\ell$.
-In case (i) the output $g$ is always correct.
-In case (ii) the output 0 is correct, while the output 1 is wrong and occurs if and only if more than $v/2$ tests go wrong, by outputting 1.
-Likewise, in case (iii) the output 1 is correct, while the output 0 occurs if and only if more than $v/2$ tests go wrong, by outputting 0.
-In both cases (ii) and (iii) we deal with the binomial distribution having failure probability $1/\gamma$, and so (cf.
-[26]) the probability $P$ of at least $v/2$ failures among $v$ tests is at most $\sum_{j=\bar{v}}^{v}\begin{pmatrix}v\\ j\end{pmatrix}T_{j}$ for $T_{j}=\gamma^{-j}(1-1/\gamma)^{v-j}\leq\gamma^{-j}$.
-Therefore, $P\leq\sum_{j=0}^{v}\begin{pmatrix}v\\ j\end{pmatrix}=T_{\bar{v}}2^{v}\max_{j}T_{j}$.
-Substitute $\sum_{j=0}^{v}\begin{pmatrix}v\\ j\end{pmatrix}=2^{v}$ and $\bar{v}=\lceil v/2\rceil$ and obtain $P<2^{v}T_{\bar{v}}\leq 2^{v}/\gamma^{v/2}=(4/\gamma)^{v/2}$.
-∎
-9 Root radius estimation based on $\ell$-test
-In this section we assume that we are given an algorithm for $\ell$-test and seek approximation of $r_{d-\ell+1}(c^{\prime})$, the $\ell$th smallest root radius from a fixed complex point $c^{\prime}$, known to lie in a fixed range $[a_{-},a_{+}]$.
-[^14] For our root-finders we only need to approximate $r_{d-\ell+1}(c^{\prime})$ for $\ell=1$ and $\ell=m$, but we estimate $r_{d-\ell+1}(c^{\prime})$ for any $\ell$, $1\leq\ell\leq m$, which is of independent interest for root-finding.
-Wlog let this be the range $[\epsilon,1]$ for a fixed small positive $\epsilon=1/2^{b}$ and seek a range $[\rho_{-},\rho_{+}]$ such that $0\leq\rho_{-}\leq r_{d-\ell+1}(c^{\prime})\leq\rho_{+}$ and either $\rho_{+}\leq\epsilon$ or else $0<\rho_{+}<\beta\rho_{-}$ for two fixed tolerance values $\epsilon=1/2^{b}$ and $\beta=2^{\phi}>1$.
-We can obtain the desired approximation by applying bisection algorithm to the range $[\epsilon,1]$, but this is too slow for us, and we apply bisection of logarithms or BoL to approximate the logarithm of the target value in the range $[-b,0]=[\log_{2}(\epsilon),\log_{2}(1)]$.
-Assume that besides the above tolerance values the input of BoL includes a softness bound $\sigma=1+\alpha$, for a fixed small positive parameter $\alpha<1/2$, and a $\sigma$-soft $\ell$-testc,ρ, to be used as the pivot of BoL.
-Since the test is soft, so is BoL as well: for a range $[b_{-},b_{+}]$ with midpoint $\widehat{b}=0.5(b_{+}+b_{-})$ regular rigid bisection continues the search in the range $[b_{-},\widehat{b}]$ or $[\widehat{b},b_{+}]$, but our soft BoL continues it in the range $[b_{-},(1+\alpha)\widehat{b}]$ or $[(1-\alpha)\widehat{b},b_{+}]$ for a fixed small positive $\alpha$.
-Accordingly we slightly increase range of search for $r_{d-\ell+1}(c^{\prime})$ to $[-\bar{b},0]$ for $\bar{b}:=\log_{2}(\frac{\sigma}{\epsilon})$.
-BoL still converges linearly and just a little slower than the regular rigid bisection: the approximation error decreases by a factor of $\frac{2}{1+\alpha}$ rather than by twice.
-[^15] If we applied BoL to the range $[0,1]$, we would have to bisect the infinite range $[-\infty,0]$, but we deal with the narrower range $[-b,0]$ since we have narrowed the range $[0,1]$ to $[\epsilon,1]$.
-Algorithm 9.1.
-Bracketing the $\ell$-th smallest root radius based on $\ell$-test.
-INPUT: Two integers $\ell$ and $m$ such that $1\leq\ell\leq m$ and #$\#(D(0,1))\leq m$, $\sigma=1+\alpha$ for $20<\alpha<1/$ 2, a $\sigma$-soft $\ell$-test0,ρ, and two tolerance values $\beta:=2^{\phi}>1$ and $\epsilon=1/2^{b}$, $b\geq 1$.
-OUTPUT: A range $[\rho_{-},\rho_{+}]$ such that $[\rho_{-},\rho_{+}]=[0,\epsilon]$ or $\rho_{-}\leq r_{d-\ell+1}\leq\rho_{+}\leq\beta\rho_{-}$.
-INITIALIZATION: Write $\bar{b}:=\log_{2}(\frac{\sigma}{\epsilon})$, $b_{-}:=-\bar{b}$, and $b_{+}:=0$.
-[…]
-$\rho_{-}=2^{b_{-}},~{}\rho_{+}=2^{b_{+}},~{}\rho:=2^{\widehat{b}},~{}{\rm and}~{}\widehat{b}:=\lceil(b_{-}+b_{+})/2\rceil.$ (9.1)
-Write $b_{-}:=(1-\alpha)\widehat{b}$ if the $\ell$-test outputs 0; write $b_{+}:=(1+\alpha)\widehat{b}$ if the test outputs 1. Then update $\widehat{b}$ and $\rho$ according to (9.1).
-Continue recursively: stop and output the current range $[\rho_{-},\rho_{+}]=[2^{b_{-}},2^{b_{+}}]$ if $b_{+}-b_{-}\leq\phi$, that is, if $\rho_{+}\leq\beta\rho_{-}$; otherwise update $\widehat{b}$ and $\rho$ and reapply the computations.
-The disc $D(0,\epsilon)$ contains at least $\ell$ roots, that is, the range $[0,\epsilon]$ brackets the $\ell$th smallest root radius $r_{d-\ell+1}$ if the $\sigma$-soft $\ell$-test applied to the disc $D(0,\frac{\epsilon}{\sigma})$ outputs 1.
-Otherwise, the disc $D(0,\frac{\epsilon}{\sigma})$ contains less than $\ell$ roots, and so the root radius $r_{d-\ell+1}$ lies in the range $[\frac{\epsilon}{\sigma},1]$, in which we shall continue our search.
-Readily verify that our policy of recursively updating the range $[b_{-},b_{+}]$ maintains this property throughout and that updating stops where $b_{+}-b_{-}\leq\phi$, that is, where $\rho_{+}\leq\beta\rho_{-}$.
-This implies correctness of Alg.
-9.1.
-Let us estimate the number $\mathbb{I}$ of its invocations of an $\ell$-test.
-Proposition 9.1.
-Alg. 9.1 invokes an $\ell$-test $\mathbb{I}$ times where $\mathbb{I}\leq\lceil\log_{\nu}(\frac{\bar{b}}{\phi})\rceil+1$ for $\nu=\frac{2}{1+\alpha}$, $\mathbb{I}=O(\log(b/\phi))$.
-Proof.
-The computations in Alg. 9.1 stop after the first application of an $\ell$-test if it outputs 1. It remains to count the subsequent $\ell$-tests if it outputs 0.
-Each test decreases the length $\bar{b}=b_{+}-b_{-}$ of its input range $[b_{-},b_{+}]$ for the exponent to $0.5(1+\alpha)\bar{b}$.
-Hence $i$ applications change it into $\frac{(1+\alpha)^{i}}{2^{i}}\bar{b}$ for $i=1,2,\dots$.
-These values linearly converge to 0 as $i$ increases to the infinity, and the proposition follows.
-∎
-Remark 9.1.
-In the case where we know that #$\#(D(0,\rho))\geq\ell$ for a fixed $\rho$ in the range $[\epsilon,1]$, we only need to refine the initial range $[\rho,1]$ for the root radius $r_{d-\ell+1}$.
-Clearly, $\mathbb{I}$ applications of $\ell$-test increase the number of evaluation points for the ratio $\frac{p^{\prime}(x)}{p(x)}$ by at most a factor of $\mathbb{I}$, and if the $\ell$-test is randomized, then the error probability also grows by at most a factor of $\mathbb{I}$.
-Let us specify these estimates for Alg.
-9.1.
-Corollary 9.1.
-Let precisely $m$ zeros of a black box polynomial $p(x)$ of a degree $d$ lie in the unit disc, that is, let #$\#(D(0,1))=m$, and let $i(D(0,1))\geq\theta$, for $\theta-1$ exceeding a positive constant.
-Given two positive tolerance values $\epsilon=1/2^{b}$ and $\phi$, an integer $\ell$, $1\leq\ell\leq m$, and a black box $\ell$-test, which evaluates polynomial $p(x)$ at $\mathcal{NR}$ points, Alg.
-9.1 approximates the $\ell$-th smallest root radius $|z_{j}|$ within the relative error $2^{\phi}$ or determines that $|z_{j}|\leq\epsilon$, at the cost of evaluation of polynomial $p(x)$ at $\mathbb{I}\cdot\mathcal{NR}$ points for $\mathbb{I}=O(\log(\frac{b}{\phi}))$ estimated in Prop.
-9.1.
-10 Precision of computing in our root-finders
-10.1 Doubling computational precision towards its a posteriori estimation
-Given an algorithm $\mathbb{A}$ for black box polynomial computations and target tolerance TOL to output errors, how should one choose computational precision?
-Here is a simple practical recipe from [10, 15, 22] and MPSolve for obtaining a posteriori estimate within a factor of two from optimal: first apply algorithm $\mathbb{A}$ with a lower precision, then recursively double it and stop where an output error decreases below TOL.
-In the rest of this section we obtain a priori estimate for the precision of $O(b\log(d))$ bits.
-10.2 Reduction to approximation of $x$NIR$(x)$ within $\frac{3}{8}$
-In our e/i tests we approximate NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ with the first order divided difference of (3.2); we only apply the tests to $\theta$-isolated discs $D(c,\rho)$ for $\theta-1>0$ of order $\frac{1}{m}$ and $\rho\geq\epsilon=1/2^{b}$.
-Map (1.3) reduces such a test for $p(x)$ to a test in the disc $D(0,1)$ for the polynomial $t(x)=p(\frac{x-c}{\rho})$ replacing $p(x)$ and decreases the error estimate by a factor of $\rho$.
-Hence map (1.3) increases the precision bound by at most a factor of $b$.
-Thus we obtain the claimed precision bound in $O(b\log(d))$ in case of any disc $D(c,\rho)$ if we prove bound $O(\log(d))$ in the case of the unit disc $D(0,1)$.
-The output of an e/i test is an integer, and we only need to certify that the output error bound of Alg.
-9.1 is less than $\frac{1}{2}$.
-The error can be magnified by a factor of $v$ in $v$ applications of Alg.
-9.1 but this factor is immaterial in our notation $\tilde{O}$.
-Our e/i tests compute the $q$ values $\rho x\frac{p^{\prime}(c+\rho x)}{p(c+\rho x)}=x\frac{t^{\prime}(x)}{t(x)}$, for $x=\zeta^{g}$, $\zeta$ of (7.2), and $g=0,1,\dots,q-1$, sum these values, and divide the sum by $q$.
-By slightly abusing notation write ${\rm NIR}(x):=\frac{t^{\prime}(x)}{t(x)}$ rather than ${\rm NIR}(x)=\frac{p^{\prime}(x)}{p(x)}$.
-According to straightforward error analysis (cf.
-[21, Part A of Ch.3] and the references therein) rounding errors introduced by these additions and division are dominated, and we can ignore them; then it is sufficient to compute $\zeta^{g}{\rm NIR}(\zeta^{g})$ within, say, $\frac{3}{8}$ for every $g$.
-Indeed, in that case the sum of $q$ error bounds is at most $\frac{3q}{8}$ and decreases below $\frac{1}{2}$ in division by $q$.
-10.3 Precision of representation of NIR$(x)$ within $\frac{1}{8}$
-Recall Eqn. (3.1) and represent $x{\rm NIR}(x)$ for $|x|=1$ within $\frac{1}{8}$ by using a precision of $O(\log(d))$ bits.
-Proposition 10.1.
-Write $C:=C(c,\rho)$ and assume that $|x|=1$ and the circle $C$ is $\theta$-isolated for $\theta>1$. Then $|x{\rm NIR}(x)|\leq\frac{d}{1-1/\theta}=\frac{\theta d}{\theta-1}$.
-Proof.
-Eqn.
-(3.1) implies that $|{\rm NIR}(x)|=|\sum_{j=1}^{d}\frac{1}{x-z_{j}}|\leq\sum_{j=1}^{d}\frac{1}{|x-z_{j}|}$ where $|x-z_{j}|\geq(1-1/\theta)\rho$ for $|z_{j}|<1$, while $|x-z_{j}|\geq(\theta-1)\rho$ for $|z_{j}|>1$ since $i(C)\geq\theta$.
-Combine these bounds, write $m:=\#(D(c,\rho))$, and obtain $|x{\rm NIR}(x)|\leq\frac{m}{1-1/\theta}+\frac{d-m}{\theta-1}\leq\frac{d}{1-1/\theta}$ for $|x|=1$.
-∎
-Corollary 10.1.
-One can represent $x{\rm NIR}(x)$ within $\frac{1}{8}$ for $|x|=1$ by using a precision of $3+\lceil\log_{2}(\frac{d}{1-1/\theta})\rceil$ bits, which is in $O(\log(d))$ provided that $\frac{1}{\theta-1}=d^{O(1)}$.
-10.4 What is left to estimate?
-Instead of $x{\rm NIR}(x)=\frac{xt^{\prime}(x)}{t(x)}$ we actually approximate $\frac{xt^{\prime}(y)}{t(x)}$ where $|y-x|\leq\delta$ and $t^{\prime}(y)$ is equal to the divided difference of (3.2).
-Thus, we shall increase the above error bound $\frac{1}{8}$ by adding to it the upper bounds $\frac{1}{8}$ on $\alpha:=|\frac{xt^{\prime}(x)}{t(x)}-\frac{xt^{\prime}(y)}{t(x)}|$ for $|x|=1$ and on the rounding error $\beta$ of computing $x\frac{t(x)-t(x-\delta)}{\delta t(x)}$.
-Next we estimate $\alpha$ in terms of $d$, $\theta$, and $\delta$ – by extending the proof of [76, Lemma 3.6], then readily ensure that $\alpha<\frac{1}{8}$ by choosing $\delta$ with a precision $\log_{2}(1/|\delta|)$ of order $\log(d)$, and finally estimate $\beta$.
-We simplify our estimates by dropping the factor $x$ where $|x|=1$.
-10.5 Precision of approximation with divided difference
-Recall the following lemma, which is [76, Fact 3.5].
-Lemma 10.1.
-For $t(x)=\prod_{j=1}^{d}(x-y_{j})$ and a non-negative integer $j\leq d$ it holds that
-$t^{(j)}(x)=j!~{}t(x)\sum_{S_{j,d}}~{}\prod_{j\in S_{j,d}}\frac{1}{x-y_{j}}$
-where the summation $\sum_{S_{j,d}}$ is over all subsets $S_{j,d}$ of the set $\{1,\dots,d\}$ having cardinality $j$.
-Proposition 10.2.
-Under the assumptions of Prop. 10.1 let $w=w(x):=\max_{j=1}\frac{1}{|x-y_{j}|}$. Then
-$|t(x)|\alpha=|\frac{t(x)-t(x-\delta)}{\delta}-t^{\prime}(x)|\leq|t(x)\frac{(d~{}w)^{2}\delta}{1-\delta~{}d~{}w}|$.
-Proof.
-The claimed bound on $|\xi^{\prime}|$ follows from (3.1).
-It remains to apply the first five lines of the proof of [76, Lemma 3.6] with $f$, $\alpha$, $n$, and $\xi$ replaced by $t$, $-\delta$, $d$, and $w$, s respectively.
-Namely, first obtain from Taylor’s expansion that $t(x)-t(x-\delta)=\sum_{j=0}^{\infty}\frac{\delta^{j}}{j!}t^{(j)}(x)-t(x)$.
-Substitute the expressions of Lemma 10.1 and obtain
-$t(x)-t(x-\delta)=\delta t^{\prime}(x)+\sum_{j=2}^{\infty}\delta^{j}t(x)\sum_{S_{j,d}}~{}\prod_{j\in S_{j,d}}\frac{1}{x-y_{j}}.$
-Combine this equation with the assumed bounds on $\frac{1}{|x-y_{j}|}$ and deduce that
-$\Big{|}\frac{t(x)-t(x-\delta)}{\delta}-t^{\prime}(x)\Big{|}\leq\Big{|}\frac{f(x)}{\delta}\sum_{j=2}^{\infty}\delta^{j}w^{j}d^{j}\Big{|}\leq\Big{|}f(x)\frac{(d~{}w)^{2}~{}\delta}{1-\delta~{}d~{}w}\Big{|}.$
-In our estimates for $\alpha$ and $\delta$ we apply Prop. 10.1 for $w\leq\frac{\theta}{\theta-1}$ and obtain
-Corollary 10.2.
-Under the assumptions of Prop. 10.1, let $w\leq\frac{\theta}{\theta-1}$. Then
-$\alpha<1/8~{}{\rm for}~{}|\delta|\leq\frac{(\theta-1)^{2}}{10d^{2}\theta^{2}}.$
-10.6 The impact of rounding errors
-To approximate $|x{\rm NIR}_{c,\rho}(x)|$ within $\frac{3}{8}$ it remains to approximate $\beta$ within $\frac{1}{8}$.
-Proposition 10.3.
-Suppose that a black box oracle has computed the values $t(y)$ within a relative error bounds $\nabla(y)$ for $y=x$, $y=x-\delta$, and $\delta$ of Cor.
-10.2.
-Then we can ensure that $|\beta|<\frac{1}{8}$ by choosing a proper $\nabla(y)$ of order $1/d^{O(1)}$, represented with a precision of $\log(\frac{1}{\nabla(y)})=O(\log(d))$ bits.
-Proof.
-Write $\beta\delta=\frac{t(x-\delta)(1+\nabla(x-\delta))}{t(x)(1+\nabla(x))}-\frac{t(x-\delta)}{t(x)}=\frac{t(x-\delta)}{t(x)}~{}\frac{\nabla(x-\delta)-\nabla(x)}{1+\nabla(x)}$.
-Taylor-Lagrange’s formula implies that $\frac{t(x-\delta)}{t(x)}=1+\frac{t(x-\delta)-t(x)}{t(x)}=1+\delta\frac{t^{\prime}(u)}{t(x)}$, for $u\in[x-\delta,x]$.
-Hence $|\frac{t(x-\delta)}{t(x)}|\leq 1+|(1+\alpha^{\prime})\delta|$ for $\alpha^{\prime}=\frac{t^{\prime}(u)-t(x)}{t(x)}$.
-By extending Cor. 10.2 obtain $|\alpha^{\prime}|\leq\frac{1}{8}$.
-Hence $|\frac{t(x-\delta)}{t(x)}|\leq w:=1+\frac{9|\delta|}{8}$, while $|\beta\delta|\leq w\frac{|\nabla(x)|+|\nabla(x-\delta)|}{1-|\nabla(x)|}$.
-Now choose $\nabla(y)$ such that $|\nabla(y)|\leq\frac{|\delta|}{17w}$ for $y=x$ and $y=x-\delta$, and so $w(|\nabla(x)|+|\nabla(x-\delta)|)\leq\frac{2}{17\delta}$.
-Then verify that $\log(\frac{1}{|\nabla(y)|})=O(\log(d))$, $1-|\nabla(x)|\geq 1-\frac{|\delta|}{17w}>\frac{16}{17}$, and $|\delta\beta|<\frac{|\delta|}{8}$.
-∎
-By combining Cors. 10.1 and 10.2 and Prop. 10.3 obtain
-Corollary 10.3.
-Suppose that a black box oracle can output approximations of a $d$-th degree polynomial $t(x)$ with any relative error in $1/d^{O(1)}$ for $|x|\leq 1$ and let the unit disc be $\theta$-isolated for $t(x)$ and for $\frac{1}{\theta-1}=d^{O(1)}$.
-Then we can perform e/i test for $t(x)$ on the unit disc by using a precision of $O(b\log(d))$ bits.
-11 Approximation of matrix eigenvalues
-Given a $d\times d$ matrix $M$ that has characteristic polynomial $p(x)=\det(xI-M)$, our root-finders approximate all $m$ its zeros (that is, the eigenvalues of $M$) lying in a fixed disc $D$ on the complex plane isolated from external eigenvalues.
-Recall Gershgorin’s bound for any eigenvalue $\lambda$ of a matrix $M=(a_{i,j})_{i,j=1}^{d}$ [138, Thm. 1.3.2]: $|\lambda-a_{i,i}|\leq\sum_{j\neq i}|a_{i,j}|$ for some $i$, $1\leq i\leq d$.
-Hence, clearly, $|\lambda|\leq||M||_{F}$.
-Combine this bound with the precision estimates of the previous section for our root-finders [^16] Practically, one can combine our root-finders with the recipe of doubling the precision of computing of [10, 15, 22].
-and with the following result of Storjohann [137] (cf.
-Remark 11.1):
-Theorem 11.1.
-One can evaluate the determinant of a $d\times d$ integer matrix $A$ by using a Las Vegas expected number of $O(d^{\omega}\log^{2}(d)\log(|A||_{F}))$ bit operations.
-This implies a record bit operation complexity bound $\tilde{O}(m^{2}\tilde{b}^{2}d^{\omega})$ (cf.
-(1.6)) for Las Vegas approximation within $\epsilon$ of all $m$ eigenvalues of $M$ lying in the disc $D$ provided that $\tilde{b}=\log_{2}(||M||_{F}/\epsilon)$.
-Remark 11.1.
-Our root-finders are essentially reduced to the evaluation of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ at sufficiently many points $x$, and we can apply the following expressions:
-$\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}((xI-M)^{-1})}$
-where $M$ is a matrix and $p(x)=\det(xI-M)$ (cf. (3.1)) and
-$\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}(M^{-1}(x)M^{\prime}(x))}$
-where $M(x)$ is a matrix polynomial and $p(x)=\det(M(x))$ [20, Eqn.
-(5)].
-[^17] These expressions have been obtained based on Jacobi’s formula ${\rm d}(\ln(\det(A)))={\rm trace}(A^{-1}{\rm d}A),$ which links differentials of a nonsingular matrix $A$ and its determinant.
-Based on the above expressions one can reduce approximation of eigenvalues of a matrix or a matrix polynomial to its inversion.
-This computation tends to be more numerically stable than computing determinants and is significantly accelerated in the highly important case where an input matrix or matrix polynomial can be inverted fast, e.g., is data sparse.
-12 Bit operation cost of the classical subdivision root-finding with e/i tests based on approximation of a Cauchy integral
-12.1 Our goal and a basic theorem
-Next we estimate the bit operation complexity of our root-finders by combining them with fast multipoint evaluation of a polynomial by means of the algorithm of Fiduccia [42] and Moenck and Borodin [84] (see [19, Sec.
-4.5] for historical account), non-trivially extended in [81] to support the following theorem:
-Theorem 12.1.
-Given a positive $b$, the coefficients of a polynomial $t(x):=\sum_{i=0}^{d}t_{i}x^{i}$ such that $||t(x)||_{1}=\sum_{i=0}^{d}|t_{i}|\leq 2^{\tau}$, and $q$ complex points $x_{1},\dots,x_{q}$ in the unit disc $D(0,1)$, one can approximate the values $t(x_{1}),\dots,t(x_{q})$ within $1/2^{b}$ by using $\tilde{O}((d+q)(b+\tau))$ bit operations.
-Proof.
-For $q=d$ this is [81, Thm.
-2], based on using novel data structure for hyperbolic approximation of $p$, proposed and analyzed by Moroz in [81].
-Trivially extend this theorem to the cases where $q>d$ (by considering $t(x)$ a polynomial of degree $q$ with $q-d$ leading coefficients 0) and where $q<d$ (by removing $d-q$ points $x_{i}$, with no increase of the complexity bound).
-∎
-We only apply this theorem for $\tau=0$, in which case its bit operation bound turns into
-$\mathbb{B}_{\rm Moroz}=\tilde{O}((d+q)b)$ (12.1)
-For comparison, the previous State of the Art was represented by the bound
-$\mathbb{B}_{\rm Kirr}=\tilde{O}((d+q)(b+q)),$ (12.2)
-due to [68, Thm.
-3.9, Alg.
-5.3, and Appendix A.3] by Peter Kirrinnis, which relied on combining his and Schönhage’s study of the bit operation cost of basic polynomial computations (cf.
-[134]) with the algorithm of [42, 84].
-[^18] Precision of computing required in the latter algorithm grows fast as $d$ increases; e.g., the IEEE standard double precision of the customary numerical computations is not enough to support that algorithm already for $d>50$; for $b=O(\log(d))$ one can fix this deficiency by applying the Fast Multipole Method of Greengard and Rokhlin [50, 28, 105, 106] to the equivalent task of multiplication of Vandermonde matrix by a vector.
-One can, however, support bounds of Eqns.
-(12.1) and (12.2) on the overall bit operation complexity by computing with multiple precision for larger values $b$.
-Moroz obtained [81, Thm.
-2] by applying bound (12.2) to his lower degree polynomials approximating $p(x)$.
-The algorithms of [42, 84, 68, 81] extensively involve the coefficients of $p(x)$.
-12.2 Reduction of our root-finders to multipoint polynomial approximation
-Given a complex $c$, a pair of positive $b$ and $\rho$, and $d+1$ coefficients of a polynomial $p=p(x)$ of (1.1) such that the disc $D(c,\rho)$ is isolated and contains precisely $m$ roots, we approximate all these roots within $R/2^{b}$ for $R=|c|+\rho$ by applying subdivision root-finding iterations with e/i tests of Sec.
-8.
-Then every subdivision step is reduced to application of e/i tests to $\bar{m}=O(m)$ discs $D(c_{\lambda},\rho_{\lambda})$ for $\rho_{\lambda}\geq R/2^{b}$ and all $\lambda$.
-Such a test, based on our e/i test of Sec.
-8, amounts to approximation of $p(x)$ at $O(m\log(d))$ equally spaced points $x$ on each of $v$ circles sharing the center $c_{\lambda}$ and having radii in the range $[\rho_{\lambda}/\sigma,\rho_{\lambda}\sigma]$ for $1<\sigma<\sqrt{2}$.
-We only use Thm.
-12.1 for $\tau=0$ to cover the case where a polynomial $t(x)$ with $||t(x)||_{1}\leq 1$ is evaluated at points $x_{j}\in D(0,1)$.
-To ensure these assumptions we first scale the variable $x\mapsto Rx$ to map the discs $D(c_{\lambda},\rho_{\lambda}\sigma)\mapsto D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})\subseteq D(0,1)$, for $\bar{c}_{\lambda}=c_{\lambda}/R$, $\bar{\rho}_{\lambda}=\rho_{\lambda}/R\geq 1/2^{b}$, and all $\lambda$.
-Then we write $t(x):=p(Rx)/\psi$ for $\psi=O(R^{d}||p(x)||_{1})$ such that $||t(x)||_{1}=1$ and approximate $t(x)$ within $1/2^{\bar{b}}$ or equivalently within the relative error bound $1/(|t(x)|2^{\bar{b}})$ at $\bar{q}:=qv$ points $x$ in $\bar{m}$ e/i tests.
-[^19] Scaling $p$ and $x$ above is done once and for all e/i tests; to compensate for this we only need to change our bound on the output precision $b$ by $\lceil|\log_{2}R|\rceil$ bits.
-Due to Cor. 10.3 it is sufficient to ensure the bound $1/(|t(x)|2^{\bar{b}})=1/d^{O(1)}$; we achieve this by choosing any $\bar{b}\geq\log(\frac{1}{|t(x)|})+O(\log(d))$.
-In our e/i tests we only need to approximate $t(x)$ at the points $x$ lying on $\theta_{\lambda}$-isolated circles where $\frac{1}{\theta_{\lambda}-1}=O(m)$.
-Map the disc $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ into the unit disc $D(0,1)$.
-This does not change isolation of the boundary circle and may increase the norm $||t(x)||_{1}$ by at most factor of $2^{b+d}$ because $|\bar{c}_{\lambda}|\leq 1$ and $\bar{\rho}_{\lambda}\geq 1/2^{b}$.
-Now substitute $||t(x)||_{1}\leq 2^{b+d}$ and $\log(\frac{||t(x)||_{1}}{|t(x)|})=O(d\log(m))$ (cf.
-[134, Eqn.(9.4)]) [^20] This bound is readily deduced from the following non-trivial result [134, Thm.
-4.2]: Let $t(x)=f\prod_{j=1}^{m}(x-z_{j})\prod_{j=m+1}^{d}(\frac{x}{z_{j}}-1)$, $||t(x)||_{1}=1$, $|z_{j}|<1$ for $j\leq m$, and $|z_{j}|>1$ for $j>m$.
-Then $|f|\geq 1/2^{d}$.
-and obtain that $\bar{b}=\log(\frac{1}{|t(x)|})+O(\log(d))=O(d\log(m)+b)$.
-Remark 12.1.
-For $m=d$ the above bound on $\log(\frac{||t(x)||_{1}}{|t(x)|})$ of [134, Eqn.(9.4)]) is sharp up to a constant factor and is reached for the polynomial $t(x)=(x+1-\frac{1}{m})^{d}/(2-\frac{1}{m})^{d}$, $\rho_{\lambda}=1$, and $x=-1$.
-For $m<d$ [134, Thm.
-4.5] implies a little stronger bounds, namely, $\log(\frac{||t(x)||_{1}}{|t(x)|})=O(d+m\log(m))$ and hence $\bar{b}=O(m\log(m)+d+b)$ because all discs $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ lie in $\theta$-isolated unit disc $D(0,1)$ for $\theta-1$ exceeding a positive constant.
-12.3 The bit operation complexity of our e/i tests and root-finders
-Apply Thm.
-12.1 with $q$ and $b$ replaced by $\bar{q}=O(m^{2})$ and $\bar{b}=O(b+d+m\log(m))$, respectively, [^21] Under the Random Root Model of Sec.
-1.5 we can whp decrease this bound by a factor of $m$ and then extend this decrease to the related subsequent estimates.
-and obtain that the bit operation cost of our $O(m)$ e/i tests of Sec.
-8 at any fixed subdivision step is in $\tilde{O}((b+d+m)m^{2})$ for $1\leq m\leq d$.
-This implies the upper estimate
-$\mathbb{B}_{\rm roots}=\tilde{O}((b+d)m^{2}b)$ (12.3)
-for the overall bit operation cost of our $O(b)$ subdivision steps, which includes the time-complexity of the evaluation oracle.
-Based on [68, Alg.
-5.3] and Eqn.
-(12.2) obtain a little larger upper bound on the bit operation cost of our subdivision root-finders:
-$\mathbb{B}_{\rm roots}^{\prime}=\tilde{O}((m^{2}+d+b)(m^{2}+d)b).$ (12.4)
-12.4 Higher bit operation complexity of general polynomial root-finding based on map (1.3) and root-squaring
-Our bit operation complexity estimates for root-finding rely on restriction of computations to approximation of NIR without involving coefficients of $p$.
-Knowing the coefficients of the polynomial $t(x)=p(\frac{x-c}{\rho})$ one can devise very fast e/i test for the disc $D(c,\rho)$ by applying Pellet’s theorem to $t_{h}(x)$ for $h$ of order $\log(d)$, whose zero are the $h$th powers of the zeros of $t(x)$ [23, 24].
-The paper [24] applies subdivision iterations of [149, 52, 128, 101] with e/i tests performed by means of Pellet’s theorem, which involves all coefficients of an input polynomial, thus losing the benefits where, say, $p(x)$ is the Mandelbrot polynomial.
-The tests must be applied to the discs covering all suspect squares, that is, up to $4m$ squares per subdivision iteration.
-A subdivision algorithm of [24] first maps the covering disc into the unit disc $D(0,1)$ (cf.
-(1.3)), thus losing the benefits where $p(x)$ is the sum of a small number of shifted monomials.
-Then the algorithm performs order of $\log(\log(d))$ root-squaring steps (one can apply order of $\log(d)$ root-lifting steps instead but within the same bit operation complexity estimates), and finally applies Pellet’s theorem to the resulting polynomial $t_{c,\rho}(x)=\sum_{i=0}^{d}t_{i}x^{i}$.
-Transition from $p(x)$ to $t_{c,\rho}(x)$, however, involves scaling of the variable $x$ by factors ranging from $R/2^{b}$ to $R=\max_{j-1}^{d}|z_{j}|$ (besides root-squaring and shift of the variable).
-Hence for every $i$, $i=0,1,\dots,d$, the coefficient $t_{i}$ must be computed within $1/2^{b-i}$ to approximate the zeros of $p$ within $1/2^{b}$ (see Observation 24.2).
-This involves at least $bg$ bits for $g=1,\dots,d$, that is, at least $0.5(d+1)db$ bits and $0.25(d+1)db$ bit-operations overall for a single e/i test.
-A single subdivision iteration for Problem 1m can involve up to $4m$ e/i tests and hence up to at least $0.25(d+1)dbm$ bit-operations.
-This is by a factor of $m$ greater than for the root-finder of [96, 103] overall.
-The bit-count above and the bit operation cost bounds for the root-finders of [96, 103, 24] decrease by a factor of $d$ where the minimal pairwise distance between the zeros of $p$ exceeds a positive constant, but the ratio of the bounds of [24] and [96, 103] remains at least $m$, where $m$ can be as large as $d$.
-Similar argument shows that even representation of the leading coefficients $t_{d}$ of $t(x)$ alone after scaling for a singe e/i test involves odder of $bd$ bits, which is amplified to $d^{2}b$ for $t_{d}^{d}$ in every e/i test (see the end of Sec.
-24.4) and up to $d^{2}bm$ per a subdivision iteration with $m$ e/i tests.
-It may be interesting that a shift by $c$ such that $|c|\leq 2$ is much less costly than scaling:
-Theorem 12.2.
-Bit operation cost of the shift of the variable, [136, Lemma 2.3] (cf.
-[68, Lemma 3.6]).
-Given a positive $b$, the coefficients of a polynomial $p(x):=\sum_{i=0}^{d}p_{i}x^{i}$ such that $||p(x)||_{1}\leq 1$, and a complex $c$ such that $|c|\leq 2$, one can approximate within $1/2^{b}$ the coefficients $t_{0},t_{1},\dots,t_{d}$ of the polynomial $t(x):=\sum_{i=0}^{d}t_{i}x^{i}:=p(x-c)$ by performing $O(\mu((b+d)d))$ bit operations for
-$\mu(s):=O(s\log(s)=\tilde{O}(s)$ (12.5)
-bounding the number of bit operations for integer multiplication modulo $2^{s}$ [53] (cf. [142, 69, 134, 135] ).
-We can relax the restriction $|c|\leq 2$ above by means of scaling the variable $x\mapsto xc$, but as we showed above, this would greatly increase the bound on the overall number of bit operations involved.
-PART III: Classical subdivision root-finder with novel e/i tests based on root-lifting
-Organization of Part III: We devote the next section to background.
-In Secs.
-14 and 15 we cover our e/i tests and their extensions to $m$-tests, respectively.
-In Secs.
-16 and 18 we estimate computational precision of our root-finders and their bit operation cost, respectively.
-In short Sec.
 9 we extend our study of root-finding to approximation of matrix eigenvalues.
 6 Background
-13.1 Definitions
-We recall or redefine some concepts from the Introduction and Part II.
-  • ${\rm NR}_{t}(x):=\frac{t(x)}{t^{\prime}(x)}$ is Newton’s ratio, ${\rm NIR}_{t}(x):=\frac{t^{\prime}(x)}{t(x)}$ is Newton’s inverse ratio for a polynomial $t(x)$.
+6.1 Basic definitions
+We extend the list of definitions used in the Introduction.
+  • ${\rm NR}_{t}(x):=\frac{t(x)}{t^{\prime}(x)}$ and ${\rm NIR}_{t}(x):=\frac{t^{\prime}(x)}{t(x)}$ are Newton’s ratio and Newton’s inverse ratio, respectively, for a polynomial $t(x)$.
 We write NR$(x):={\rm NR}_{p}(x)$ and NIR$(x):={\rm NIR}_{p}(x)$ (cf.
@@
   • Define squares, discs, circles (circumferences), and annuli on the complex plane: $S(c,\rho):=\{x:~{}|\Re(c-x)|\leq\rho$, $|\Im(c-x)|\leq\rho\}$$D(c,\rho):=\{x:~{}|x-c|\leq\rho\},$ $C(c,\rho):=\{x:~{}|x-c|=\rho\}$,  $A(c,\rho,\rho^{\prime}):=\{x:~{}\rho\leq|x-c|\leq\rho^{\prime}\}.$
-  • $\Delta(\mathbb{S})$, $CH(\mathbb{S})$, $X(\mathbb{S})$ and #$\#(\mathbb{S})$ denote the diameter, root set, convex hull, and index (root set’s cardinality) of a set $\mathbb{S}$ on the complex plane, respectively.
+  • $\Delta(\mathbb{S})$, $MCD(\mathbb{S})$, $X(\mathbb{S})$, and #$\#(\mathbb{S})$ denote the diameter, root set, minimal covering disc, and index (root set’s cardinality) of a set $\mathbb{S}$ on the complex plane, respectively.
   • A disc $D(c,\rho)$, a circle $C(c,\rho)$, or a square $S(c,\rho)$ is $\theta$-isolated for $\theta\geq 1$ if $X(D(c,\rho))=X(D(c,\theta\rho))$, $X(C(c,\rho))=X(A(c,\rho/\theta,\rho\theta)$ or $X(S(c,\rho))=X(S(c,\theta\rho)),$ respectively.
     The isolation of a disc $D(c,\rho)$, a circle $C(c,\rho)$, or a square $S(c,\rho)$ is the largest upper bound on such a value $\theta$ (see Fig.
-5) and is denoted $i(D(c,\rho))$, $i(C(c,\rho))$, or $i(S(c,\rho))$, respectively.
-    We say isolated for “$\theta$-isolated" if $\theta-1$ exceeds a positive constant.
+4) and is denoted $i(D(c,\rho))$, $i(C(c,\rho))$, or $i(S(c,\rho))$, respectively.
+    We say “isolated" for “$\theta$-isolated" if $\theta-1$ exceeds a positive constant.
     $\rho$$\rho\theta$$\rho\theta^{2}$c********
@@
   • A disc $D$ has rigidity $\eta(\mathbb{S})=\Delta(X(D))/\Delta(D)\leq 1$. A disc is $\eta$-rigid for any $\eta\leq\eta(D)$.
-    We say rigid for “$\eta$-rigid" if $\eta$ exceeds a positive constant.
-  • $|u|:=||{\bf u}||_{1}:=\sum_{i=0}^{q-1}|u_{i}|$, $||{\bf u}||:=||{\bf u}||_{2}:=(\sum_{i=0}^{q-1}|u_{i}|^{2})^{\frac{1}{2}}$, and $|u|_{\rm max}:=||{\bf u}||_{\infty}:=\max_{i=0}^{q-1}|u_{i}|$ denote the 1-, 2-, and $\infty$-norm, respectively, of a polynomial $u(x)=\sum_{i=0}^{q-1}u_{i}x^{i}$ and its coefficient vector [^22] Here and hereafter ${\bf v}$ denotes a column vector, while ${\bf v}^{T}$ denotes its transpose.
-${\bf u}=(u_{i})_{i=0}^{d-1}$, and then (see [48, Eqs.
+    We say “rigid" for “$\eta$-rigid" if $\eta$ exceeds a positive constant.
+  • $|u|:=||{\bf u}||_{1}:=\sum_{i=0}^{q-1}|u_{i}|$, $||{\bf u}||:=||{\bf u}||_{2}:=(\sum_{i=0}^{q-1}|u_{i}|^{2})^{\frac{1}{2}}$, and $|u|_{\rm max}:=||{\bf u}||_{\infty}:=\max_{i=0}^{q-1}|u_{i}|$ denote the 1-, 2-, and $\infty$-norm, respectively, of a polynomial $u(x)=\sum_{i=0}^{q-1}u_{i}x^{i}$ and its coefficient vector ${\bf u}=(u_{i})_{i=0}^{d-1}$, and then (see [35, Eqs.
 (2.2.5)–(2.2.7)])
@@
 Write $r^{h}_{j}(c,t):=(r_{j}(c,t))^{h}$, $r_{j}(c):=r_{j}(c,p)$.
-  • Generalize Def.
-2.1 of $\sigma$-soft e/i tests as follows: For $p(x)$ of (1.1), $\sigma>1$, and integers $\ell$ and $m$ such that $1\leq\ell\leq m\leq d$ and #$\#(D(0,1))\leq m$, a $\sigma$-soft $\ell$-test, or just $\ell$-test for short (in particular e/i test turns into 1-test), either outputs 1 and stops if it detects that $r_{d-\ell+1}\leq\sigma$, that is, #$\#(D(0,\sigma))\geq\ell$, or outputs 0 and stops if it detects that $r_{d-\ell+1}>1$, that is, [^23] Bounds $r_{d-\ell+1}\leq\sigma$ and $r_{d-\ell+1}>1$ can hold simultaneously (cf.
-Fig.
-1b), but as soon as an $\ell$-test verifies any of them, it stops without checking if the other bound also holds.
-#$\#(D(0,1))<\ell$.
-An $\ell$-test applied to the polynomial $t(y)$ of (1.3) is said to be an $\ell$-test for the disc $D(c,\rho)$ and/or the circle $C(c,\rho)$ and also to be an $\ell$-testc,ρ.
-[^24] We will only use $\ell$-tests for $\ell=1$ (which are e/i tests) and $\ell=m$.
-Observation 13.1.
+Observation 6.1.
 Taylor’s shift, or translation, of the variable $x\mapsto y:=x-c$ and its scaling combined (see (1.4)) map the zeros $z_{j}$ of $p(x)$ into the zeros $y_{j}=\frac{z_{j}-c}{\rho}$ of $t(y)$, for $j=1,\dots,d$, and preserve the index and the isolation of the disc $D(c,\rho)$.
 6.2 Reverse polynomial
-Immediately observe the following properties of the reverse polynomial of (1.5).
+Immediately observe the following properties of the reverse polynomial (1.5).
 Proposition 6.1.
@@
 ${\rm(iii)}~{}p^{\prime}_{\rm rev}(x)=dx^{d-1}p\Big{(}\frac{1}{x}\Big{)}-x^{d-2}p^{\prime}\Big{(}\frac{1}{x}\Big{)}~{}{\rm for}~{}x\neq 0,~{}p_{d}=p_{\rm rev}(0),~{}p_{d-1}=p^{\prime}_{\rm rev}(0).$
+${\rm(iv)}~{}~{}~{}~{}\frac{p_{\rm rev}(x)}{p^{\prime}_{\rm rev}(x)}=\frac{1}{x}-\frac{{\rm NIR}(x)}{x^{2}},~{}x\neq 0,~{}{\rm NIR}(0)=\frac{p_{d-1}}{p_{d}}=-\sum_{j=1}^{d}z_{j}.~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
 […]
-13.3 Root-squaring and root-lifting
-The DLG root-squaring iterations (3.7) by Dandelin 1826, Lobachevsky 1834, and Gräffe 1837 compute the polynomials $p_{2^{i}}(x):=\prod_{j=1}^{d}(x-z_{j}^{2^{i}})$, for $i=0,1,\dots$.
-By applying $h$-lifting of (3.6) we extend the definition of $p_{h}(x)$ to all $h\in\mathbb{N}$.
-As we noticed in the introduction, the known studies of DLG iterations always focus and stumble on numerically unstable computation of the coefficients of the polynomials $p_{i}(x)$ [86, 113], but we only compute their values bypassing those problems.
+6.3 Root-lifting with divided differences
 […]
-${\rm NIR}_{p_{h}}(x)=\frac{1}{{\rm NR}_{p_{h}}(x)}\approx\frac{p_{h}(x)-p_{h}(x-\delta)}{p_{h}(x)\delta}=\frac{p_{h}^{\prime}(y)}{p_{h}(x)}.$ (13.3)
+${\rm NIR}_{p_{h}}(x)\approx\frac{p_{h}(x)-p_{h}(x-\delta)}{p_{h}(x)\delta}=\frac{p_{h}^{\prime}(y)}{p_{h}(x)}.$ (6.3)
 Remark 6.1.
 […]
-Observation 13.2.
+Observation 6.2.
 (i) Root-lifting preserves the index #$\#(D(0,1))$ of the unit disc.
 (ii) $h$-lifting (achieved with root-squaring for $h=2^{k}$ and a positive integer $k$) maps the values $i(D(0,1))$, $\sigma(D(0,1))$, $\eta(D(0,1))$, and $r_{j}(0,p)$ for all $j$, that is, the isolation of the unit disc $D(0,1)$, its softness, its rigidity and root radii, to their $h$-th powers.
-13.4 The distance to a closest root and fast inclusion tests
-Next we estimate the distance from the origin to the closest root. The distance to the farthest root $z_{1}$ is given by the same estimates applied to the reverse polynomial $p_{\rm rev}(x)$.
-For bounds $r_{d}$ and $r_{1}$ in terms of the coefficients of $p$, see [80, Sec. 1.10]; in particular according to Fujihara’a bounds of [41]:
+6.4 The distance to a closest root and certified inclusion
+[63, Sec. 1.10] lists various bounds on $r_{d}$ and $r_{1}$ in terms of the coefficients of $p$, in particular [31]:
 $r_{1}\leq 2\max_{i=1}^{d}|p_{d-j}/p_{d}|^{1/j}$ unless $p_{d}=0$ and $r_{d}\geq 2\min_{i=1}^{d}|p_{j}/p_{0}|^{1/j}$ unless $p_{0}=0$.
-For a complex $c$ and positive integers $j$ and $h$, Eqn. (3.1) and Observation 13.2 combined imply the following relationships independent of $p_{d}$:
-$r_{d}(c)\leq d~{}|{\rm NR}(c)|,$ (13.4)
-$r^{h}_{j}(0,p)=r_{j}(0,p_{h}).$ (13.5)
+For a complex $c$ and positive integers $j$ and $h$, Eqn. (3.1) and Observation 6.2 combined imply that
+$r_{d}(c)\leq d~{}|{\rm NR}(c)|,~{}{\rm and}~{}r^{h}_{j}(0,p)=r_{j}(0,p_{h}).$ (6.4)
 The roots lying far from a point $c$ little affect NIR$(c)$, and we extend bounds (6.4) as follows.
@@
 Proof.
-We can assume that $|z_{j}|>\theta$ if and only if $j>m$ and then deduce from Eqn.
+Assume that $|z_{j}|>\theta$ if and only if $j>m$ and then deduce from Eqn.
 (3.1) that $|{\rm NIR}(c)|\leq\sum_{j=1}^{m}\frac{1}{|c-z_{j}|}+\sum_{j=m+1}^{d}\frac{1}{|c-z_{j}|}<\frac{m}{r_{d}(c)}+\frac{d-m}{\theta-1}$.
 ∎
-Certify inclusion into the disc $D(0,\sigma)$ if $\rho_{d}(0,p_{h})\leq\sigma^{h}$, $\sigma>1$ and $h>0$, although this bound is very poor for worst case polynomials $p$:
+Certify inclusion into the disc $D(0,\sigma)$ if $\rho_{d}(0,p_{h})\leq\sigma^{h}$, $\sigma>1$ and $h>0$, although this bound is extremely poor for worst case polynomials $p$:
 Proposition 6.3.
@@
 Notice that $p^{\prime}(0)=0\neq p(0)=-u^{d}$, $p(x)=\prod_{j=0}^{d-1}(x-u\zeta_{d}^{j})$ for $\zeta_{d}$ of (3.4). ∎
-Next we argue that such deficiency only occurs for a narrow class of polynomials $p$ and comment on some heuristic recipes to counter it.
 Notice that $\frac{1}{r_{d}(c)}\leq|{\rm NIR}(c)|=\frac{1}{d}\Big{|}\sum_{j=1}^{d}\frac{1}{c-z_{j}}\Big{|}$ by virtue of (3.1), and so approximation to the root radius $r_{d}(c,p)$ is poor if and only if severe cancellation occurs in the summation of the $d$ roots.
-Formal probabilistic support for this observation is hard, but intuitively such a cancellation only occurs for a very narrow class of polynomials $p$.
-If so, the above non-costly test certifies inclusion for a very large class of polynomials $p$ [^26] One can try some empirical recipes.
-Clearly, the problem persists for the root radius $r_{d}(w)$ where $p^{\prime}(w)$ vanishes; rotation of the variable $p(x)\leftarrow t(x)=p(ax)$ for $|a|=1$ does not fix it, but the bound $r_{d}\leq|p_{0}/p_{d}|^{1/d}$ and shifts $p(x)\leftarrow t(x)=p(x-c)$ for $c\neq 0$ combined should fix it.
-We must also study the case where e/i tests output exclusion, however, because this occurs in more than 25% of all e/i tests in a subdivision process:
+It seems that such a cancellation only occurs for a very narrow class of polynomials $p$, and if so, then the above non-costly test certifies inclusion for a very large class of polynomials $p$ [^15] Shifts of the variable $p(x)\leftarrow t(x)=p(x-c)$ for random $c$ should further enlarge this class.
+6.5 Subdivision root-finding involves exclusion
+It is not enough for us to have non-costly certification of inclusion because exclusion occurs in more than 25% of all e/i tests in a subdivision process:
 Proposition 6.4.
@@
 Notice that $\sigma_{i+1}=4(\sigma_{i}-d_{i}-f_{i})~{}{\rm for}~{}i=1,2,\dots,t+1$, sum these equations for all $i$, substitute $\sigma_{1}=4$ and $\sigma_{t+1}=0$, and obtain $3\Sigma+4\geq 4D+4F$.
-Represent subdivision process by a tree with $F\leq m$ leaves whose other nodes are components made up of suspect squares, to obtain pessimistic lower bounds $\Sigma+1\geq 2m\geq 2F$ and hence $D\geq\Sigma/4+1/2$.
+Represent subdivision process by a tree with at most $F\leq m$ leaves whose other nodes are components made up of suspect squares, to obtain pessimistic lower bounds $\Sigma+1\geq 2m\geq 2F$ and hence $D\geq\Sigma/4+1/2$.
 ∎
@@
 7.1 Overview
-Map (1.3) and Cor.
-3.1 together reduce an e/i test for any disc to the test for the unit disc $D=D(0,1)$, and we devise soft e/i tests for that disc assuming that #$\#(D(0,2))\leq m$ for a fixed $m$, $1\leq m\leq d$.
 Our Alg.
@@
 7.3 are deterministic specifications of Alg.
-3.1, which is always correct when it claims inclusion, by virtue of Lemma 3.1.
-We prove that both of these specifications are also correct when they output exclusion.
+3.1.
+By virtue of Lemma 3.1 all of them are correct when they claim inclusion, but we will prove that they are also correct when output exclusion.
 Namely, Algs.
-14.1 and 14.2 evaluate ${\rm NIR}(x)$ at $hq$ equally spaced points of the unit circle $C(0,1)$ for $h=O(\log(d))$ and correctly claim exclusion if at all these points $|{\rm NIR}(x)|\leq\frac{1}{2\sqrt{q}}$ for $q>d$ or if $|{\rm NIR}(x)|\leq\frac{1}{3\sqrt{q}}$ for $q>m$, respectively.
-Our Example 14.1 shows that we cannot devise such deterministic algorithms for $q$ decreased below $m$, but in Sec.
-14.4 we devise Las Vegas randomized e/i tests that evaluate polynomial $p(x)$ at $O(\log(d))$ points $x$ overall under the Random Root Model.
+7.1 and 7.2 evaluate ${\rm NIR}(x)$ at $hq$ equally spaced points of the unit circle $C(0,1)$ for $h=O(\log(d))$ and certify exclusion if at all these points $|{\rm NIR}_{p_{h}}(x)|\leq\frac{1}{2\sqrt{q}}$ for $q>d$ or if $|{\rm NIR}_{p_{h}}(x)|\leq\frac{1}{3\sqrt{q}}$ for $q>m$, respectively, and otherwise certify inclusion.
+Our Example 7.1 shows that we cannot devise such deterministic algorithms for $q<m$, but our Las Vegas e/i test evaluates polynomial $p(x)$ at $O(\log(d))$ points $x$ overall under the Random Root Model.
 7.2 Basic deterministic e/i test
@@
 INITIALIZATION: Fix two integers $q>d$ and $h:=\lceil\log_{\sigma}(1+2d\sqrt{q})\rceil$, such that $\sigma^{h}\leq 1+2d\sqrt{q}$.
-COMPUTATIONS: For $g=0,1,\dots,q-1$ recursively compute
+COMPUTATIONS: For $g=0,1,\dots,q-1$ recursively compute the values
 $v_{g}:=\Big{|}\frac{p^{\prime}_{h}(\zeta^{g})}{p_{h}(\zeta^{g})}\Big{|}.$ (7.1)
 Stop and certify $\sigma$-soft inclusion #$(D(0,\sigma))>0$ if $1+d/v_{g}\leq\sigma^{h}$ for an integer $g<q$. Otherwise stop and certify exclusion, that is, #$(D(0,1))=0$.
-Alg. 14.1 outputs $\sigma$-soft inclusion correctly by virtue of Lemma 3.1. Next we prove that it also correctly claims exclusion.
+Alg. 7.1 outputs $\sigma$-soft inclusion correctly by virtue of Lemma 3.1.
+Next we will prove that it also correctly claims exclusion.
 Lemma 7.1.
@@
 ∎
-Now, if Alg.
-14.1 outputs exclusion, then $1+\frac{d}{v_{g}}>\sigma^{h}$ for $g=0,1,\dots,q-1$, where $\sigma^{h}>1+2d\sqrt{q}$ according to initialization of Alg.
-14.1.
-Hence $\frac{1}{v_{g}}>2\sqrt{q}$, and so the assumptions of Prop.
-14.1 hold for $p_{h}(x)$ replacing $p(x)$.
-Therefore, the theorem implies that the disc $D(0,1)$ contains no zeros of $p_{h}(x)$, but then it contains no zeros of $p(x)$ as well (see Observation 13.2 (iii)), and so exclusion is correct.
+If Alg.
+7.1 outputs exclusion, then $1+\frac{d}{v_{g}}>\sigma^{h}$ for $g=0,1,\dots,q-1$.
+Hence $\frac{1}{v_{g}}>2\sqrt{q}$ because $\sigma^{h}>1+2d\sqrt{q}$ according to initialization of Alg.
+7.1.
+Therefore, the assumptions of Prop.
+7.1 hold for $p_{h}(x)$ replacing $p(x)$, and hence the disc $D(0,1)$ contains no zeros of $p_{h}(x)$, but then it contains no zeros of $p(x)$ as well (see Observation 6.2 (iii)), and so exclusion is correct.
 Readily verify the following estimates.
@@
 […]
-Next we supply some relevant estimates.
+Next we supply relevant estimates.
 Algorithm 7.2.
@@
 Otherwise certify exclusion, #$(D(0,1))=0$.
-Prop. 13.2, for $p_{h}$ replacing $p$, and the assumed bound $\frac{d-m}{\theta^{h}-1}\leq\frac{1}{6\sqrt{q}}$ together imply
+Prop. 6.2, for $p_{h}$ replacing $p$, and the assumed bound $\frac{d-m}{\theta^{h}-1}\leq\frac{1}{6\sqrt{q}}$ together imply
 $\frac{m}{r_{d}(\zeta^{g},p_{h})}\geq v_{g}-\frac{d-m}{\theta^{h}-1}\geq v_{g}-\frac{1}{6\sqrt{q}}~{}{\rm for~{}all}~{}g.$
-Alg. 14.2 claims inclusion if $v_{g}\geq\frac{1}{3\sqrt{q}}$ for some $g$, but then
+Alg. 7.2 claims inclusion if $v_{g}\geq\frac{1}{3\sqrt{q}}$ for some $g$, but then
 $\frac{m}{r_{d}(\zeta^{g},p_{h})}\geq\frac{1}{6\sqrt{q}},~{}r_{d}(\zeta^{g},p_{h})\leq 6m\sqrt{q},$
@@
 $r_{d}(0,p_{h})\leq 1+6m\sqrt{q}~{}\leq\sigma^{h}~{}{\rm for}~{}|\zeta|=1.$
-Now Observation 13.2 (iii) implies that Alg. 14.2 claims inclusion correctly.
-Next prove that it also claims exclusion correctly.
+Now Observation 6.2 (iii) implies that Alg. 7.2 claims inclusion correctly.
+Next we prove that it also claims exclusion correctly.
 Proposition 7.3.
@@
 ∎
-Readily extend Prop. 14.2 to Alg. 14.2 by decreasing its bounds on $q$ and $h$ provided that #$\#(D(0,\theta))=m$ for $\theta\geq 1+6(d-m)\sqrt{q}$.
+Readily extend Prop. 7.2 to Alg. 7.2 by decreasing its bounds on $q$ and $h$ provided that #$\#(D(0,\theta))=m$ for $\theta\geq 1+6(d-m)\sqrt{q}$.
 Proposition 7.4.
 Alg.
-14.2 is a $\sigma$-soft e/i test: it certifies either inclusion or exclusion at the cost of evaluation of $p(x)$ at $2hj$ or $2hq$ points $x$, respectively, for three integers $h$, $j$, and $q$ such that $1\leq j\leq q$, $q>m:=\#(D(0,\theta)),~{}\sigma^{h}>1+6m\sqrt{q}$, $\theta\geq 1+6(d-m)\sqrt{q}$, and $h:=\lceil\log_{\sigma}(1+6m\sqrt{q})\rceil$, so that $h=O(\log(m))$ if $\sigma-1$ exceeds a positive constant.
+7.2 is a $\sigma$-soft e/i test: it certifies either inclusion or exclusion at the cost of evaluation of $p(x)$ at $2hj$ or $2hq$ points $x$, respectively, for three integers $h$, $j$, and $q$ such that $1\leq j\leq q$, $q>m:=\#(D(0,\theta)),~{}\sigma^{h}>1+6m\sqrt{q}$, $\theta\geq 1+6(d-m)\sqrt{q}$, and $h:=\lceil\log_{\sigma}(1+6m\sqrt{q})\rceil$, so that $h=O(\log(m))$ if $\sigma-1$ exceeds a positive constant, and in this case we only need $O(m\log(m)$ queries of the oracle to perform an e/i test.
 7.4 A randomized e/i test
-14.4.1.
+7.4.1.
 Introductory comments.
 Would Algs.
-14.1 and 14.2 stay correct if we apply them at fewer evaluation points?
+7.1 and 7.2 stay correct if we apply them at fewer evaluation points?
 They have consistently stayed correct for $q$ of order $\log^{2}(d)$ in extensive numerical tests of [45, 48] applied to $p(x)$ rather than to $p_{i}(x)$ for $i>0$.
@@
 Eqn.
-7.2 in Part II), and reported inclusion otherwise.
+12.2 in Part III), and reported inclusion otherwise.
 Clearly, $|s_{h,q}|\leq\delta$ for all $h$ if $|p(\zeta^{g})|\leq\delta/q$ for $g=0,1,\dots,q-1$, while inclusion is certified unless $|p(\zeta^{g})|$ is small for all $g$.
@@
 Clearly, [45, 48] have never encountered worst case inputs in their extensive tests.
-Next we formally support this observation by introducing probability structure in the space of polynomials $p(x)$ and $t(x)=p(\frac{x-c}{\rho})$.
-Shifts destroy all customary probability distributions in the space of the coefficients of $p(x)$, but we prove that our e/i tests that evaluate $p(x)$ at $O(\log(d)$ points are correct whp under the Random Root Model.
+Next, in line with this observation, we prove that our e/i tests, which evaluate $p(x)$ at $O(\log(d)$ points, are correct whp under the Random Root Model.
 […]
@@
 $z-\bar{z}=\frac{\lambda_{g+1}-z\lambda_{g}}{\bar{z}^{g}},~{}|\bar{z}-z|\leq\frac{|\lambda_{g+1}|+|z|\cdot|\lambda_{g}|}{|\bar{z}|^{g}}<\frac{\nu_{g+1}+|z|\nu_{g}}{|\bar{z}|^{g}}=O\Big{(}\frac{d}{(\sigma|\bar{z}|)^{g}}\Big{)}.$
-Hence $|z-\bar{z}|=O(1/\beta^{g})$ for a fixed $\beta>1$. Combine this bound with (14.6) and (14.7), choosing sufficiently large $g$ of order $\log d$ unless $\sigma|\bar{z}|<\beta$.
+Hence $|z-\bar{z}|=O(1/\beta^{g})$ for a fixed $\beta>1$. Combine this bound with (7.6) and (7.7), choosing sufficiently large $g$ of order $\log(d)$ unless $\sigma|\bar{z}|<\beta$.
 Combine this bound with (7.6) and (7.7), recall that $h=g$ or $h=g+1$, and obtain the lemma. ∎
-For a fixed $\sigma>1$, e.g., for $\sigma=1.2$, choose a constant $\beta>1$ and sufficiently large $g=O(\log(d))$ such that $\gamma<\pi$ in (14.4).
+For a fixed $\sigma>1$, e.g., for $\sigma=1.2$, choose a constant $\beta>1$ and a sufficiently large $g=O(\log(d))$ such that $\gamma<\pi$ in (7.4).
 Then under the Random Root Model the probability $P$ of the output error for Alg.
@@
 We can decrease the error probability below $\nabla^{v}$ by applying the algorithm $v$ times, in particular below $d^{a}$ for any fixed positive constant $a$ and for $v=O(\log(d))$.
-Based on Cor.
-2.2 we reduce classical subdivision root-finder to performing $O(mb)$ soft e/i tests, but in Part IV we decrease that bound to $w=O(m\log(b))$ tests.
-These tests magnify the error bounds by at most a factor of $w$.
-Since $m\leq d$, we can ensure correctness of the resulting e/i test whp provided that $\log(b)=o(d^{a-1})$.
 We arrive at the following
 Corollary 7.1.
-Under Random Root Model one only need to apply a subdivision root-finder that involves $O(m\log(b)))$ soft Las Vegas randomized e/i tests to ensure the output error bound $\epsilon=1/2^{b}$ whp provided that $\log(b)=d^{O(1)}$ and that every test is reduced to $v=O(\log(d))$ applications of Alg.
-3.1(rrm).
+Under Random Root Model one only needs to query an evaluation oracle to perform an e/i test correct;y whp.
 Remark 7.1.
-It is tempting to combine root-lifting or root-squaring with algorithms of [115], which approximate Cauchy integral (3.4), to obtain alternative randomized e/i test supporting Cor.
-14.1, but for that we would have to raise disc isolation from order of $1+1/m$ to a constant, and this would require too costly lifting of order $m$.
+It is tempting to combine root-lifting or root-squaring with the algorithms of Part III to obtain an alternative randomized e/i test supporting Cor.
+7.1, but for that we would need to raise disc isolation from order of $1+1/m$ to a constant; this would require too costly lifting of order $m$.
 7.5 Counting the evaluation queries of our root-finders
-Combine Cors. 2.2 and 14.1 and Prop. 14.4 and obtain
+Combine Cors. 2.2 and 7.1 and Prop. 7.4 and obtain
 Corollary 7.2.
-We can reduce root-finding to deterministic evaluation of $p(x)$ at $O(m^{2}b\log(d))$ points as well as to evaluation at $O(mb\log^{2}(d))$ points under the Random Root Model.
-15 Reduction of $m$-test to e/i tests
-Our two algorithms of Secs.
-15.1 and 15.2 reduce $m$-test for a disc $D$ to bounded number of e/i tests.
-Namely, the algorithm of Sec.
-15.2 uses just a single e/i test where $m=d$ but up to order of $\log(d)$ e/i tests otherwise, while the algorithm of Sec.
-15.1 uses $O(1)$ e/i tests for any $m$.
-15.1 $m$-test by means of e/i tests for any disc
-Our alternative reduction of $m$-test for $1\leq m\leq d$ to a bounded number of e/i tests can be applied in the case of any disc, not necessarily isolated.
-If at least one of the e/i tests fails, then definitely there is a root in a narrow annulus about the boundary circle of an input disc, and hence the internal disc of the annulus would not contain $m$ roots.
-Otherwise we would obtain a root-free concentric annulus about the boundary circle of an input disc and then reduce $m$-test to root-counting in an isolated disc, which we can perform at a low cost (see, e.g., claim (ii) of Prop.
-7.1).
-Algorithm 15.1.
-Reduction of $m$-test to e/i tests.
-INPUT: a polynomial $p(x)$ of a degree $d$, two values $\bar{\theta}>1$ and $\bar{\sigma}>1$, an integer $m$ such that $1\leq m\leq d$ and #$\#(D(0,\bar{\sigma}))\leq m$, and two algorithms $\mathbb{ALG}~{}0$ and $\mathbb{ALG}~{}1$, which for a black box polynomial perform $\bar{\sigma}$-soft e/i test for in any disc and root-counting in an isolated disc, respectively.
-INITIALIZATION (see Fig. 6): Fix a positive $\rho$ such that $\bar{\sigma}=O(\rho)$, say, $\rho=\bar{\sigma}/5$, and compute the integer
-$v=\lceil 2\pi/\phi\rceil$ (15.1)
-where $\phi$ is the angle defined by an arc of $C(0,1)$ with end points at the distance $\rho$, and so $v=O(1)$ for $\bar{\sigma}=O(\rho)$. Then compute the values
-$c_{j}=\exp(\phi_{j}{\bf i}),~{}j=0,1,\dots,v,$ (15.2)
-such that $|c_{j}-c_{j-1}|=\rho$ for $j=1,2,\dots,v$ and $\phi_{j}-\phi_{j-1}=\phi$.
-$c_{1}$$c_{2}$$c_{3}$$c_{4}$$c_{5}$$c_{6}$$c_{7}$$c_{8}$$c_{9}$$c_{0}$
-Figure 6: $v=9$, $|c_{j}-c_{j-1}|=\rho>|c_{9}-c_{0}|$ for $j=1,2,\dots,9$.
-OUTPUT: Either an upper bound $r_{d-m+1}<1-\rho\frac{\sqrt{3}}{2}$ or one of the two lower bounds $r_{d-m+1}\geq 1-\bar{\sigma}\rho$ or $r_{d-m+1}>1+(\sqrt{3}-1)\rho$.
-COMPUTATIONS: Apply algorithm $\mathbb{ALG}~{}0$ (soft e/i test) to the discs $D(c_{j},\rho)$, $j=0,1,\dots,v$.
-(a) Stop and certify that $r_{d-m+1}\geq 1-\bar{\sigma}\rho$, that is, #$\#(D(0,1-\bar{\sigma}\rho))<m$ for the open disc $(D(0,1-\bar{\sigma}\rho)$, unless algorithm $\mathbb{ALG}~{}0$ outputs exclusion in all its $v+1$ applications.
-(b) If it outputs exclusion $v+1$ times, then apply algorithm $\mathbb{ALG}~{}1$ (root-counter) to the $\theta$-isolated disc $D(0,1)$ for
-${}\theta:=\min\Big{\{}1+(\sqrt{3}-1)\rho,\frac{1}{1-0.5\rho\sqrt{3}}\Big{\}}.$ (15.3)
-Let $\bar{s}_{0}=\#(D(0,1))$ denotes its output integer. Then conclude that
-(i) $r_{d-m+1}\leq 1-\rho\frac{\sqrt{3}}{2}$, and so #$\#(D(0,1-\rho\frac{\sqrt{3}}{2}))=m$ if $\bar{s}_{0}=m$, while
-(ii) $r_{d-m+1}\geq 1+(\sqrt{3}-1)\rho$, and so #$\#(D(0,1+(\sqrt{3}-1)\rho))<m$ for the open disc $D(0,1+(\sqrt{3}-1)\rho)$ if $\bar{s}_{0}<m$.
-To prove correctness of Alg. 15.1 we need the following result.
-Theorem 15.1.
-For the above values $c_{0},\dots,c_{v}$, and $\rho$, the domain $\cup_{j=0}^{v-1}D(c_{j},\rho)$ covers the annulus $A(0,1-0.5\rho\sqrt{3},1+(\sqrt{3}-1)\rho)$.
-Proof.
-Let $c_{j-1}$ and $c_{j}$ be the centers of two neighboring circles $C(c_{j-1},\rho)$ and $C(c_{j},\rho)$, let $z_{j}$ and $z^{\prime}_{j}$ be their intersection points such that $|z_{j}|=|z_{j}^{\prime}|-|z_{j}-z^{\prime}_{j}|$ for $j=1,2,\dots,v$, observe that the values $|z_{j}-z_{j}^{\prime}|$, $|z_{j}|$, $|z^{\prime}_{j}|$, and $|c_{j-1}-c_{j}|$ are invariant in $j$, and hence so is the annulus $A(0,|z_{j}|,|z_{j}^{\prime}|)$ of width $|z_{j}-z_{j}^{\prime}|$ as well; notice that it lies in the domain $\cup_{j=0}^{v-1}D(c_{j},\rho)$.
-$c_{0}$$c_{1}$$z$$z^{\prime}$
-Figure 7: $|c_{1}-c_{0}|=|c_{0}-z|=|c_{1}-z|=|c_{0}-z^{\prime}|=|c_{1}-z^{\prime}|=\rho$.
-Let us prove that $|z_{j}-z_{j}^{\prime}|=\rho\sqrt{3}$ for $j=1$, and hence also for $j=2,\dots,v$.
-Write $z:=z_{1}$ and $z^{\prime}:=z_{1}^{\prime}$ and observe (see Fig.
-7) that $|c_{1}-c_{0}|=|c_{0}-z|=|c_{1}-z|=|c_{0}-z^{\prime}|=|c_{1}-z^{\prime}|=\rho$, and so the points $c_{0}$, $c_{1}$, and $z$ are the vertices of an equilateral triangle with a side length $\rho$.
-Hence $|z-z^{\prime}|=\rho\sqrt{3}$.
-The circle $C(c_{j},\rho)$ intersects the line segment $[z_{j},z_{j}^{\prime}]$ at a point $y_{j}$ such that $|y_{j}-z_{j}|>\rho\sqrt{3}/2$ and $|y_{j}-z_{j}^{\prime}|>(\sqrt{3}-1)\rho$ for $\rho<1$, and this implies the theorem if $|c_{v}-c_{0}|=\rho$.
-If, however, $|c_{v}-c_{0}|<\rho$ and if $C(c_{0},\rho)\cap C(c_{v},\rho)=\{\bar{z},\bar{z}\}$, then $|\bar{z}-\bar{z}^{\prime}|>|z-z^{\prime}|$, and the theorem follows in this case as well.
-∎
-Corollary 15.1.
-The disc $D(0,1)$ is $\theta$-isolated for $\theta$ of (15.3) provided that the discs $D(c_{j},\rho)$ contain no roots for $j=0,1,\dots,v$.
-Correctness of Alg.
-15.1.
-(i) If algorithm $\mathbb{ALG}~{}1$ outputs inclusion in at least one of its applications, then one of the discs $D(c_{j},\bar{\sigma}\rho)$ contains a root, and then Alg.
-15.1 correctly reports that $r_{d-m+1}\geq 1-\sigma\bar{\rho}$.
-(ii) Otherwise Cor. 15.1 holds, and then correctness of Alg. 15.1 follows from correctness of $\mathbb{ALG}~{}1$.
-Softness is at most $\frac{1}{1-\bar{\sigma}\rho}$, $\frac{1}{1-0.5\rho\sqrt{3}}$, and $1+(\sqrt{3}-1)\rho$ in cases (a), (b,i), and (b,ii), respectively.
-Computational complexity.
-Alg.
-15.1 invokes algorithm $\mathbb{ALG}~{}0$ at most $v+1$ times for $v=O(1)$ of (15.1).
-If $\mathbb{ALG}~{}0$ outputs exclusion $v+1$ times, then algorithm $\mathbb{ALG}~{}1$ is invoked and runs at a dominated cost.
-Remark 15.1.
-One can modify Alg.
-15.1 as follows (see [63]): instead of fixing a positive $\rho$ and then computing $v$ of (15.1) fix an integer $v>0$, apply algorithm $\mathbb{ALG}~{}0$ at $v$ equally spaced points $c_{j}=\exp(\frac{2j\pi{\bf i}}{v}),~{}j=0,1,\dots,v-1$, let $\phi=\frac{2\pi}{v}$, and then define $\rho$.
-We also arrive at this algorithm if we compute or guess this value $\rho$ and then apply Alg.
-15.1, but notice that $c_{v}=c_{0}$ for $c_{j}$ of (15.2) and only invoke algorithm $\mathbb{ALG}~{}0$ for $c_{0},\dots,c_{v-1}$.
-15.2 Simple reduction to a single e/i test in the case of an isolated disc
-Clearly, we can immediately reduce the $d$-test in the unit disc $D(0,1)$ for $p(x)$ to performing 1-test (that is, e/i test) in that disc for the reverse polynomial $p_{\rm rev}(x)$ of (1.5).
-Let us next reduce an $m$-test, for $0<m<d$, to a single application of 1-test provided that an input disc $D$ with $m$ roots is $\theta$-isolated for $\theta$ of order $\log(d)$; we can ensure that provision by applying root-lifting or root-squaring, in both cases at the cost of evaluation of $p(x)$ at $O(\log(d))$ points if an initial disc is just isolated.
-Proposition 15.1.
-Let the root sets of two polynomials $f(y)$ and $g(y)$ of degrees $m$ and $d-m$, respectively, lie outside the disc $D(0,1)$ and in the disc $D(0,1/\theta)$, respectively, for a sufficiently large real $\theta$.
-Then
-${\rm NIR}_{f}(z)={\rm NIR}_{fg}(z)-\frac{d-m}{z}+\Delta~{}{\rm for}~{}|z|=1~{}{\rm and}~{}|\Delta|\leq\frac{d-m}{\theta-1}.$
-Proof.
-Notice that $\frac{1}{z-y_{j}}-\frac{1}{z}=\frac{y_{j}}{(z-y_{j})z}$, recall that $|y_{j}|\leq\frac{1}{\theta}$ for $\theta>1$ and all $d-m$ roots $y_{1},\dots,y_{d-m}$ of $g(y)$ by assumption, and deduce that
-$\Big{|}\frac{1}{z-y_{j}}-\frac{1}{z}\Big{|}\leq\frac{1}{\theta-1}~{}{\rm for}~{}|z|=1~{}{\rm and}~{}j=1,\dots,d-m.$ (15.4)
-[…]
-Next fix an integer $q$ such that $m<q\leq 2m$, e.g., $q:=m+1$ or $q:=2^{h}$ for $h=\lceil\log_{2}(m+1)$. Assume that $\theta-1$ exceeds a positive constant and ensure that
-$\theta>1+4(d-m)\sqrt{q}$ (15.5)
-by applying root-lifting or root-squaring at the cost of evaluation of $p(x)$ at $O(\log(d))$ points.
-[…]
-[…]
-$V_{f}:=\max_{g=0}^{q-1}|{\rm NIR}_{fg}(\zeta^{g})-(d-m)\zeta^{-g}|.$ (15.6)
-Now ensure that $V_{f}<\frac{1}{2\sqrt{q}}$, say, by requiring that $\max_{g=0}^{q-1}|{\rm NIR}_{fg}(\zeta^{g})|<\frac{1}{4\sqrt{q}}$ because bound (15.5) implies that $\frac{d-m}{\theta-1}<\frac{1}{4\sqrt{q}}$.
+We can reduce root-finding to deterministic evaluation of $p(x)$ at $O(m^{2}b\log(d))$ points as well as to the evaluation at $O(mb\log^{2}(d))$ points under the Random Root Model.
 8 Precision of computing
@@
 8.2 A priori estimates for computational precision
-In the rest of this section we deduce a priori estimate of $O(b\log(d))$ bits for computational precision, proving that this precision supports correctness of our subdivision root-finders.
-Such an estimate is deduced in [76] based on advanced non-trivial techniques, of which we only use a very small part.
+In the rest of this section we deduce a priori estimate of $O(b\log(d))$ bits for computational precision.
+Namely, we prove that this precision supports correctness of our subdivision root-finders; our analysis includes some advanced techniques of [59].
 Estimating computational precision of our soft e/i tests we assume that they are applied to the unit disc $D(0,1)$ because we can extend the results to any disc $D(c,\rho)$ by applying map (1.4) and increasing the upper bounds on the precision by a factor of $\log_{2}(1/\rho)$.
-This factor is at most $b$ because we only handle discs of radii $\rho>\epsilon=1/2^{b}$ until we stop reaching $\epsilon$-discs.
+This factor is at most $b$ because we only handle discs of radii $\rho>\epsilon=1/2^{b}$ until we reach $\epsilon$-discs and stop.
 Therefore, our bound $O(\log(d))$ on the computational precision of our e/i tests for the disc $D(0,1)$ will imply the bound $O(b\log(d))$ for any disc.
-We also make Assumption 0 that we can request and obtain from evaluation oracle (black box subroutine) the values of $p$ with a relative error bound in $1/d^{\nu}$ for any fixed constant $\nu>0$ of our choice; we can represent these values with the bit-precision $\log_{2}(d)+\log_{2}(\nu)$.
-In our e/i tests we only need to compare the values $|{\rm NIR}_{t_{h}}(x)|$ at some complex points $x$ with some fixed values $v_{d}\geq 1/d^{\beta}$ for a constant $\beta>0$, and next we estimate that under Assumption 0 we can indeed do this by performing our computations with a precision of $O(\log(d))$ bits.
+We also make Assumption 0 that we can request and obtain from evaluation oracle (black box subroutine) the values of $p$ with a relative error bound within $1/d^{\nu}$ for any fixed constant $\nu>0$ of our choice; we can represent these values with the bit-precision $\log_{2}(d)+\log_{2}(\nu)$.
+In our e/i tests we only need to compare the values $|{\rm NIR}_{t_{h}}(x)|$ at some complex points $x=v_{d}\geq 1/d^{\beta}$ for a constant $\beta>0$, and next we estimate that under Assumption 0 we can indeed do this by performing our computations with a precision of $O(\log(d))$ bits.
 Multiplication with a precision $s$ contributes a relative error at most $2^{s}$ (cf.
-[21, Eqn.
+[14, Eqn.
 A.3 of Ch.
-3]) and otherwise just sums the relative error bounds of the factors ignoring the dominated impact of higher order errors.
+3]) and otherwise just sums the relative error bounds of the factors for we ignore the dominated impact of higher order errors.
 Hence we deduce that $h=O(\log(d))$ multiplications for root-lifting to their $h$th powers increase these bounds by at most a factor of $h$, which is immaterial since we allow relative errors of order $1/d^{\nu}$ for a constant $\nu>0$ of our choice.
-We only apply this analysis to approximation of NIR$(x)$ for $p_{h}(x)$ based on (3.2) and also use the following
+We apply this analysis to approximation of NIR$(x)$ for $p_{h}(x)$ based on (3.2) and under the following
 Assumption 1: Whp all $q$ evaluation points $x\in C(0,1)$ of an e/i test for the unit disc $D(0,1)$ lie at the distance at least $\phi:=\frac{2\pi}{mqw}$ from all roots for any fixed $w>1$ of our choice.
-This assumption holds because we apply our e/i tests under random rotation $x\mapsto vx$ for the random variable $v$ specified in our next proposition.
+This assumption holds whp because we apply our e/i tests under random rotation $x\mapsto vx$ for random variable $v$ specified in our next proposition.
 Proposition 8.1.
@@
 Then $\psi_{j}=0$ for at least $d-m$ integers $j$, and $\max_{j=1}^{d}\psi_{j}\leq\frac{\pi^{2}}{mqw}$ for sure.
-Hence $\sum_{j=1}^{d}\psi_{j}<\frac{\pi^{2}}{qw}$, and so $\phi_{g}\leq\frac{2\pi}{mqw}$ for a fixed $g$ with a probability at most $\frac{\pi}{qw}$.
-Therefore, $\phi\leq\frac{2\pi}{mqw}$ with a probability at most $\frac{\pi}{w}$.
+Hence $\sum_{j=1}^{d}\psi_{j}<\frac{\pi^{2}}{qw}$, and so $\phi_{g}\leq\frac{2\pi}{mqw}$ for a fixed $g$ with a probability at most $\frac{\pi}{qw}$, and the lemma follows.
 ∎
-Next we prove that a precision of $O(\log(d))$ bits supports our e/i tests under Assumptions 0 and 1 by applying Prop.
-16.1 for sufficiently large $w=d^{O(1)}$.
-We will (i) prove that $|{\rm NIR}(z)|=d^{O(1)}$, (ii) approximate this value within $d^{\nu}\delta$ for any fixed constant $\nu$ and for $\delta$ of (3.2) chosen such that $\log(\frac{1}{\delta})=O(\log(d))$, and (iii) estimate that such a bound withstands the impact of rounding errors of computations for (3.2).
+Next, by applying Prop.
+8.1 for a sufficiently large $w=d^{O(1)}$, we prove that a precision of $O(\log(d))$ bits supports our e/i tests under Assumptions 0 and 1.
+We will (i) prove that $|{\rm NIR}(z)|=d^{O(1)}$, (ii) will approximate this value within $d^{\nu}\delta$ for any fixed constant $\nu$ and for $\delta$ of (3.2) chosen such that $\log(\frac{1}{\delta})=O(\log(d))$, and (iii) will estimate that such a bound withstands the impact of rounding errors of computations for (3.2).
 Eqn. (3.1) immediately implies
@@
 Thus we shall increase the above error bound by adding upper bounds $\alpha$ on $|\frac{t^{\prime}(x)}{t(x)}-\frac{t^{\prime}(y)}{t(x)}|$ and $\beta$ on the rounding error of computing $\frac{t(x)-t(x-\delta)}{\delta t(x)}$, for $x=\zeta^{g}$ and $g=1,2,\dots,q$.
-In the proof of the next theorem we use the following lemma, which is [76, Fact 3.5].
+In the proof of the next theorem we use the following lemma, which is [59, Fact 3.5].
 Lemma 8.1.
@@
 Proposition 8.3.
-Under the assumptions of Prop. 16.2 it holds that
-$|t(x)\alpha|=|\frac{t(x)-t(x-\delta)}{\delta}-t^{\prime}(x)|\leq|t(x)\frac{(d\xi^{\prime})^{2}\delta}{1-d\xi^{\prime}\delta}|$ for $\xi^{\prime}=\frac{1}{\rho}$.
+Under the assumptions of Prop. 8.2 it holds that
+$|t(x)\alpha|=\Big{|}\frac{t(x)-t(x-\delta)}{\delta}-t^{\prime}(x)\Big{|}\leq\Big{|}t(x)\frac{(d\xi^{\prime})^{2}\delta}{1-d\xi^{\prime}\delta}\Big{|}~{}{\rm for}~{}\xi^{\prime}=\frac{1}{\rho}.$
 Proof.
 The claimed bound on $|\xi^{\prime}|$ follows from (3.1).
-It remains to apply the first five lines of the proof of [76, Lemma 3.6] with $f$, $\alpha$, $n$, and $\xi$ replaced by $t$, $-\delta$, $d$, and $w$, respectively.
+It remains to apply the first five lines of the proof of [59, Lemma 3.6] with $f$, $\alpha$, $n$, and $\xi$ replaced by $t$, $-\delta$, $d$, and $w$, respectively.
 Namely, first obtain from Taylor’s expansion that $t(x)-t(x-\delta)=\sum_{j=0}^{\infty}\frac{\delta^{j}}{j!}t^{(j)}(x)-t(x)$.
@@
 Proposition 8.4.
-Suppose that the values $t(y)$ have been computed by a black box oracle within a relative error bounds $\nabla(y)$ for $y=x$, $y=x-\delta$, and $\delta$ of Cor.
-16.2.
+Suppose that the values $t(y)$ have been computed by an oracle within a relative error bounds $\nabla(y)$ for $y=x$, $y=x-\delta$, and $\delta$ of Cor.
+8.2.
 Then one can ensure that $\beta<\frac{1}{d^{\bar{\eta}}}$ for any fixed constant $\bar{\eta}$ by choosing a proper $\nabla(y)$ of order $1/d^{O(1)}$, represented with a precision of $O(\log(d))$ bits, which we can do under Assumption 0.
@@
 Taylor-Lagrange’s formula implies that
-$\frac{t(x-\delta)}{t(x)}=1+\frac{t(x-\delta)-t(x)}{t(x)}=1+\delta\frac{t^{\prime}(u)}{t(x)}$, for $u\in[x-\delta,x]$.
+$\frac{t(x-\delta)}{t(x)}=1+\frac{t(x-\delta)-t(x)}{t(x)}=1+\delta\frac{t^{\prime}(u)}{t(x)},~{}{\rm for}~{}u\in[x-\delta,x].$
 Hence $|\frac{t(x-\delta)}{t(x)}|\leq 1+|\delta\alpha^{\prime}|+|\frac{t^{\prime}(x)}{t(x)}|$ for $\alpha^{\prime}=\frac{t^{\prime}(u)-t(x)}{t(x)}$.
@@
 the proof of Prop.
-10.1).
-Hence $w:=|\frac{t(x-\delta)}{t(x)}|\leq d^{O(1)}$ and $|\beta\delta|\leq|w|\frac{|\nabla(x)|+|\nabla(x-\delta)|}{1-|\nabla(x)|}$.
-Now choose $\nabla(y)$ such that $|\nabla(y)|\leq\frac{|\delta|}{20|w|}$ for $y=x$ and $y=x-\delta$.
-Then verify that $\log(\frac{1}{|\nabla(y)|})=O(\log(d))$, $1-|\nabla(x)|\geq 1-\frac{|\delta|}{20|w|}>\frac{19}{20}$, and $|\delta\beta|<\frac{|\delta|}{8}$.
-∎
-Combine Cors. 16.1 and 16.2 and Prop. 16.4 to obtain
+14.1).
+Hence
+$w:=\Big{|}\frac{t(x-\delta)}{t(x)}\Big{|}\leq d^{O(1)}~{}{\rm and}~{}|\beta\delta|\leq|w|\frac{|\nabla(x)|+|\nabla(x-\delta)|}{1-|\nabla(x)|}.$
+Now choose $\nabla(y)$ such that $|\nabla(y)|\leq\frac{|\delta|}{20|w|}$ for $y=x$ and $y=x-\delta$. Then verify that
+$\log\Big{(}\frac{1}{|\nabla(y)|}\Big{)}=O(\log(d)),~{}1-|\nabla(x)|\geq 1-\frac{|\delta|}{20|w|}>\frac{19}{20},~{}{\rm and}~{}|\delta\beta|<\frac{|\delta|}{8}.$
+Combine Cors. 8.1 and 8.2 and Prop. 8.4 to obtain
 Corollary 8.3.
-Given a black box (oracle) for the evaluation of a polynomial $p(x)$ of a degree $d$ with a relative error of any order in $1/d^{O(1)}$ (this holds under Assumption 0), we can perform our black box subdivision root-finders with a precision of $O(\log(d))$ bits.
+Given an oracle for the evaluation of a polynomial $p(x)$ of a degree $d$ with a relative error of any order in $1/d^{O(1)}$ (this holds under Assumption 0), we can perform our black box subdivision root-finders with a precision of $O(\log(d))$ bits.
 Remark 8.1.
 We derived our precision bound $O(\log(d))$ (in the case of root-finding in the unit disc $D(0,1)$) based on operating entirely with NIR$(x)$.
-For comparison, if we relied on computing $p(x)$ and on the bound $\min_{x:~{}|x|=1}(|p(x)|/\sum_{i=0}^{d}|p_{i}|)\geq(\frac{\theta-1}{2\theta})^{d}$ of [134, Eqn.
-(9.4)], we would have only obtained a bound of order $m^{2}d$ on the computational precision of our root-finders.
+For comparison, if we relied on computing $p(x)$ and on the bound $\min_{x:~{}|x|=1}(|p(x)|/\sum_{i=0}^{d}|p_{i}|)\geq(\frac{\theta-1}{2\theta})^{d}$ of [106, Eqn.
+(9.4)], we would have only needed a bound of order $m^{2}d$ on the computational precision of our root-finders.
 9 Approximation of matrix eigenvalues
-Given a $d\times d$ matrix $M$ that has characteristic polynomial $p(x)=\det(xI-M)$, our root-finders can approximate all $m$ its zeros (that is, the eigenvalues of $M$) lying in a fixed disc $D$ on the complex plane isolated from external eigenvalues.
-Recall Gershgorin’s bound for any eigenvalue $\lambda$ of a matrix $M=(a_{i,j})_{i,j=1}^{d}$ [138, Thm. 1.3.2]: $|\lambda-a_{i,i}|\leq\sum_{j\neq i}|a_{i,j}|$ for some $i$, $1\leq i\leq d$.
+Given a $d\times d$ matrix $M$ that has characteristic polynomial $p(x)=\det(xI-M)$, our root-finders can approximate all its $m$ zeros (that is, the eigenvalues of $M$) lying in a fixed disc $D$ on the complex plane isolated from its external zeros (or eigenvalues of $M$).
+Recall Gershgorin’s bound for any eigenvalue $\lambda$ of a matrix $M=(a_{i,j})_{i,j=1}^{d}$ [111, Thm. 1.3.2]:
+$|\lambda-a_{i,i}|\leq\sum_{j\neq i}|a_{i,j}|~{}{\rm for~{}some}~{}i,~{}1\leq i\leq d.$
 Hence, clearly, $|\lambda|\leq||M||_{F}$.
-Combine this bound with the precision estimates of the previous section for our root-finders [^27] For a posteriori estimates, one can combine our root-finders with the recipe of doubling the precision of computing of (see Sec.
-16.1).
-and with the following result of Storjohann [137] (cf.
-Remark 17.1):
+Combine this bound with the a priori precision estimates of the previous section for our root-finders [^16] For a posteriori estimates for the bit operation complexity of Problems 0, $0^{*}$, 2, and 2, one can combine our root-finders with the recipe of doubling the precision of computing of Sec.
+8.1.
+and with the following result of Storjohann [109]:
 Theorem 9.1.
 One can evaluate the determinant of a $d\times d$ integer matrix $A$ by using an expected number of $O(d^{\omega}\log^{2}(d)\log(|A||_{F}))$ bit operations.
-This implies a record bit operation complexity bound $\tilde{O}(m^{2}\tilde{b}^{2}d^{\omega})$ for any matrix $M$ which decreases whp by a factor of $m$ for matrices $M$ whose eigenvalues are iid random variables sampled under the uniform probability distribution from a large disc on the complex plane (cf.
-(1.6)) for approximation within $\epsilon$ of all $m$ eigenvalues of $M$ lying in the disc $D$ provided that $\tilde{b}=\log_{2}(||M||_{F}/\epsilon)$.
-This supports a bound exceeding (1.6) by a factor of $b$, to be removed with acceleration of our root-finding in Part III.
-Remark 17.1.
-Our root-finders are essentially reduced to the evaluation of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ at sufficiently many points $x$, and we can rely on the following expressions for approximation of the eigenvalues of a matrix or a matrix polynomial:
-$\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}((xI-M)^{-1})}$
-where $M$ is a matrix and $p(x)=\det(xI-M)$ (cf. (3.1)) and
-$\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}(M^{-1}(x)M^{\prime}(x))}$
-where $M(x)$ is a matrix polynomial and $p(x)=\det(M(x))$ [20, Eqn.
-(5)].
-[^28] These expressions have been obtained based on Jacobi’s formula ${\rm d}(\ln(\det(A)))={\rm trace}(A^{-1}{\rm d}A),$ which links differentials of a nonsingular matrix $A$ and of its determinant.
-Based on the above expressions one can reduce approximation of eigenvalues of a matrix or a matrix polynomial to its inversion.
-This computation tends to be more numerically stable than computing determinants and is significantly accelerated the highly important case where an input matrix or matrix polynomial can be inverted fast, e.g., is data sparse.
-18 Bit operation complexity of general polynomial root-finding
-Next we estimate the bit operation complexity of our root-finders, which accelerate classical subdivision root-finding by incorporating our e/i tests and are applied to a general polynomial $p$ of (1.1) whose coefficients are given with a precision sufficient to support our root-finding.
-We could simply combine the precision bound $O(b\log(d))$ of Sec.
-16 for our e/i tests with the estimate $O(bm)$ of Cor.
-14.2 for the number of evaluation points for our root-finders, the bound of $2d$ ops for the evaluation of $p(x)$, and (12.5) for bit-operation cost of an op and then would arrive at the bit-operation complexity bound $\tilde{O}(b^{2}dm)$ for our root-finders.
-[^29] Recall that under the Random Root Model of Sec.
-1.5 we can whp decrease this bound by a factor of $m$.
-Next, however, we decrease this bound roughly by a factor of $d$ by using benefits of fast multipoint evaluation of a polynomial by means of the algorithm of Fiduccia [42] and Moenck and Borodin [84] (see [19, Sec.
-4.5] for historical account), non-trivially extended in [81] to support the following theorem:
+For any $d\times d$ matrix $M$ we arrive at a record bit operation complexity bound $\tilde{O}(m^{2}\tilde{b}bd^{\omega})$ for approximation within $\epsilon$ of all $m$ eigenvalues of any matrix $M$ lying in the disc $D$ provided that $\tilde{b}=\log_{2}(||M||_{F}/\epsilon)$.
+This bound exceeds (1.6) by a factor of $b$ and can be decreased by a factor of $m$ under the Random Root Model for the roots of the characteristic polynomial of $M$.
+10 Bit operation complexity of polynomial root-finding
+10.1 Straightforward extension of the precision bound and a direction to saving evaluation queries
+Recall the bound
+$\mu(s):=O(s\log(s)=\tilde{O}(s)$ (10.1)
+on the number of bit operations for integer multiplication modulo $2^{s}$ [40] (cf.
+[114, 53, 106, 107]), combine it with the estimate $O(bm)$ of Cor.
+7.2 for the number of evaluation queries in our root-finders, the upper bound $2d$ ops for the evaluation of a polynomial $p(x)$ given with its coefficients, and the precision bound $O(b\log(d))$ of Sec.
+8, and obtain the bit-operation complexity bound $\tilde{O}(b^{2}dm)$ for our root-finders.
+[^17] By extending precision bound of [59] we greatly simplified derivation of our bit-complexity estimates versus the papers [106, 52], each of 46 pages, which bypass precision bound and had to use long tedious step-by-step analysis of their much involved root-finders.
+Since $p(x)$ is given with its coefficients, we apply fast multipoint evaluation of a polynomial in [32, 67] and use fewer ops, roughly by a factor of $d/\log^{2}(d)$.
+Peter Kirrinnis in [52] and Moroz in [64] have non-trivially extended this progress to saving bit-operations:
 Theorem 10.1.
@@
 Proof.
-For $q=d$ this is [81, Thm.
-2], based on using novel data structure for hyperbolic approximation of $p$, proposed and analyzed by Guillaume Moroz in [81].
+For $q=d$ this is [64, Thm.
+2], based on using novel data structure for hyperbolic approximation of $p$, proposed and analyzed by Moroz in [64].
 Trivially extend this theorem to the cases where $q>d$ (by considering $t(x)$ a polynomial of degree $q$ with $q-d$ leading coefficients 0) and where $q<d$ (by removing $d-q$ points $x_{i}$, with no increase of the complexity bound).
 ∎
-We only apply this theorem for $\tau=0$, in which case its bit operation cost bound turns into
-$\mathbb{B}_{\rm Moroz}=\tilde{O}((d+q)b)$ (18.1)
-For comparison, the previous State of the Art was represented by the bound
-$\mathbb{B}_{\rm Kirr}=\tilde{O}((d+q)(b+q)),$ (18.2)
-due to [68, Thm.
+We only apply Thm. 10.1 for $\tau=0$, in which case its bit operation cost bound turns into
+$\mathbb{B}_{\rm Moroz}=\tilde{O}((d+q)b).$ (10.2)
+For comparison, the State of the Art before [64] was the bound
+$\mathbb{B}_{\rm Kirr}=\tilde{O}((d+q)(b+q))$ (10.3)
+of [52, Thm.
 3.9, Alg.
-5.3, and Appendix A.3] by Peter Kirrinnis, which relied on combining his and Schönhage’s study of bit operation complexity of basic polynomial computations (cf.
+5.3, and Appendix A.3] by Kirrinnis, who combined his and Schönhage’s study of the bit operation complexity of basic polynomial computations (cf.
 [106]) with the algorithm of [32, 67].
-[^30] Precision of computing required in the latter algorithm grows fast as $d$ increases; e.g., the IEEE standard double precision of the customary numerical computations is not enough to support that algorithm already for $d>50$; for $b=O(\log(d))$ one can fix this deficiency by applying the Fast Multipole Method of Greengard and Rokhlin [50, 28, 105, 106] to the equivalent task of multiplication of Vandermonde matrix by a vector.
+[^18] Precision of computing required in the latter algorithm grows fast as $d$ increases; e.g., the IEEE standard double precision of the customary numerical computations is not enough to support that algorithm already for $d>50$; for $b=O(\log(d))$ one can fix this deficiency by reducing the task to multiplication of Vandermonde matrix by a vector applying FMM (cf.
+[86, 87]).
 One can, however, support bounds of Eqns.
 (10.2) and (10.3) on the overall bit operation complexity by computing with multiple precision for larger values $b$.
-Moroz obtained [81, Thm.
+Moroz obtained [64, Thm.
 2] by applying bound (10.3) to his lower degree polynomials approximating $p(x)$.
-The algorithms of [42, 84, 68, 81] extensively involve the coefficients of $p(x)$.
 Bound (10.2) combined with the precision bound $O(\log(d))$ of Sec.
-16 implies the bit operation cost bound $\tilde{O}((d+q)b)$ dominated by an estimated bit operation cost $\mathbb{B}$ for the evaluation of $p(x)$ with relative error $1/d^{O(1)}$ at $\tilde{O}(db)$ points.
-It remains to estimate $\mathbb{B}$, which is our next goal.
+8 implies the bit operation cost bound $\tilde{O}((d+q)b)$, dominated by an estimated bit operation cost $\mathbb{B}$ for the evaluation of $p(x)$ with relative error $1/d^{O(1)}$ at $\tilde{O}(db)$ points.
+It remains to estimate $\mathbb{B}$, which we do next.
 10.2 Reduction of our root-finders to multipoint polynomial approximation
 Given a complex $c$, a pair of positive $b$ and $\rho$, and $d+1$ coefficients of a polynomial $p=p(x)$ of (1.1) such that the disc $D(c,\rho)$ is isolated and contains precisely $m$ roots, we approximate all these roots within $R/2^{b}$ for $R=|c|+\rho$ by applying subdivision root-finding iterations with e/i tests of Sec.
-14.
+7.
 Then every subdivision step is reduced to application of our e/i tests to $\bar{m}=O(m)$ discs $D(c_{\lambda},\rho_{\lambda})$, where $\rho_{\lambda}\geq R/2^{b}$ for all $\lambda$.
@@
 Due to Cor.
-16.3 it is sufficient to ensure the bound $1/(|t(x)|2^{\bar{b}})=1/d^{O(1)}$; we achieve this by choosing $\bar{b}\geq\log(\frac{1}{|t(x)|})+O(\log(d))$.
+8.3 it is sufficient to ensure the bound $1/(|t(x)|2^{\bar{b}})=1/d^{O(1)}$; we deduce it by choosing $\bar{b}\geq\log(\frac{1}{|t(x)|})+O(\log(d))$.
 In our e/i tests we only need to approximate $t(x)$ at the points $x$ lying on $\theta_{\lambda}$-isolated circles where $\frac{1}{\theta_{\lambda}-1}=O(m)$ (cf.
 Assumption 1 of Sec.
-16.2).
-Then $\log(\frac{1}{|t(x)|})=O(d\log(m))$ by virtue of the bound of [134, Eqn.
+8.2).
+Then $\log(\frac{1}{|t(x)|})=O(d\log(m))$ by virtue of [106, Eqn.
 (9.4)]).
-[^32] That bound follows from the estimate $|f|\geq 1/2^{d}$ of [134, Thm.
+[^20] That bound follows from the estimate $|f|\geq 1/2^{d}$ of [106, Thm.
 4.2], which holds provided that $t(x)=f\prod_{j=1}^{m}(x-x_{j})\prod_{j=m+1}^{d}(\frac{x}{x_{j}}-1)$, $||t(x)||_{1}=1$, $|x_{j}|<1$ for $j\leq m$, and $|x_{j}|>1$ for $j>m$.
@@
 For $m=d$, the above bound on $\log(\frac{1}{|t(x)|})$ is sharp up to a constant factor and is reached for the polynomial $t(x)=(x+1-\frac{1}{m})^{d}/(2-\frac{1}{m})^{d}$, $\rho_{\lambda}=1$, and $x=-1$.
-For $m<d$, [134, Thm.
+For $m<d$, [106, Thm.
 4.5] implies a little stronger bounds, namely, $\log(\frac{1}{|t(x)|})=O(d+m\log(m))$ and hence $\bar{b}=O(m\log(m)+d+b)$ because all discs $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ lie in $\theta$-isolated unit disc $D(0,1)$ for $\theta-1$ exceeding a positive constant.
@@
 $\mathbb{B}_{\rm roots}=\tilde{O}((b+d)(m^{2}+d)b)$ (10.4)
-for the overall bit operation cost of our $O(b)$ subdivision steps.
-[^33] Under the Random Root Model of Sec.
-1.5 we can decrease this bound by a factor of $m$ whp and can extend this decrease to the related subsequent estimates.
-Based on [68, Alg. 5.3] and Eqn. (18.2) we obtain a little larger but still reasonably small upper bound
+for the overall bit operation cost of our $O(b)$ subdivision steps. Based on [52, Alg. 5.3] and Eqn. (10.3) we obtain a little larger but still reasonably small upper bound
 $\mathbb{B}_{\rm roots}^{\prime}=\tilde{O}(((m^{2}+d)(m^{2}+d+b)b).$ (10.5)
-In Part IV we decrease the number of subdivision steps of our root-finders from order of $b$ to $O(\log(b))$, thus decreasing bounds (18.3) and (18.4) by a factor of $b$, and arrive at our Main Theorems 2, 3, and 4.
-PART IV: Compression of a Disc and Acceleration of Subdivision Root-Finders
-19 Overview and two alternative approaches
-Classical subdivision iterations converge to the roots with linear convergence rate (see Sec.
-2), which is slow when many consecutive iterations process a component $C$ made up of a single suspect square or up to four of them sharing a vertex.
-In this case the algorithms of [128, 101] dramatically accelerate convergence by means of applying Newton’s or Schröder’s iterations (see Fig.
-8), which converge with superlinear rate to the convex hull of all roots lying in $C$.
-The diameter of the convex hull, however, can be too large, and then [128, 101] apply subdivision iterations again.
-The resulting root-finder recursively intertwines subdivision iterations with Newton’s or Schröder’s, and this complicates the complexity estimates for a root-finder.
-[128, 101] have only estimated the number of ops but not of the bit operations involved.
-[23, 24] estimated the bit operation complexity, but only under a very strong assumption about the output (see Remark 4.2 and Sec.
-12.4).
-Figure 8: Five blue marked centers of suspect squares converge to a green-marked root cluster with linear rate; two red marked Schröder’s iterates converge with quadratic rate.
-Instead of Newton’s or Schröder’s iterations, one can apply the algorithm of [134], which splits $p(x)$ into the product of two polynomials $f(x)$ and $g(x)$ whose zero sets lie in and outside the minimal disc covering the component $C$, respectively.
-Then one can apply a root-finder to the polynomials $f=f(x)$ and $g=g(x)$.
-Splitting of $p(x)$ into factors does not preserve sparsity of $p(x)$, and the factors may have significantly larger coefficient length than $p(x)$.
-(E.g., split $p(x)=x^{1000}+1$ into two factors of degrees 300 and 700.)
-Overall asymptotic bit-complexity for high precision root-finding can still be near-optimal if a root-finder recursively splits $p(x)$ into factors of comparable degrees.
-This is achieved in [96, 103], but the resulting root-finder is much involved and extensively uses the coefficients of $p$.
-The problems caused by splitting disappear if one of the factors, say, $g$ has a small degree; then we can deflate it, approximate its zeros readily, and approximate the zeros of $f(x)=p(x)/g(x)$ by applying to it a black box root-finder and using implicit deflation based on (3.3) or the equation NIR$(p)=$NIR(f)+limit-$(f)+$NIR$(g)$ for $p=fg$.
-One would lose optimality, however, by always deflating low-degree factors: to deflate $k$ times one would need at least order of $bkmd$ bit operations overall, that is, would exceed the lower bound $0.25(d+1)db$ of Cor.
-24.2 by at least a factor of $k$.
-We stay with subdivision framework but instead of Newton’s or Schröder’s iterations, used in [128, 101], apply our novel algorithm that compresses a component isolated from its external roots into a near-minimal disc that keeps all internal roots.
-This enables us to decrease the estimates of Secs.
-12.3 and 18.2 for the overall complexity of root-finding by a factor of $b$, thus completing proof of our Main Theorems 2-4.
-We organize our presentation in this part as follows.
-In the next section we study a tree representing recursive partition of the union of suspect squares in a subdivision process into connected components.
+10.4 Higher bit operation complexity of general polynomial root-finding based on Pellet’s theorem, map (1.4), and root-lifting
+Recall the followings theorem of Pellet [71], which extends the previous results by Cauchy and Rouché:
+Theorem 10.2.
+Given a monic polynomial $t(x)=\sum_{i=0}^{d}t_{i}x^{i}$ with complex coefficients and $t_{d}=1$, let the polynomial $s(x)=x^{d}+p_{d-1}x^{d-1}+\cdots+|p_{i+1}|x^{i+1}-|p_{i}|x^{i}|+p_{i-1}|x^{i-1}+\cdots+|p_{0}|$ have two distinct positive roots $r$ and $R$, $r<R$.
+Then the polynomial $t(x)$ has exactly $i$ zeros in or on the circle $C(0,r)=\{z:~{}|z|=r\}$ and no zeros in the annulus $A(0,r,R)=\{z:~{}r<|z|<R\}$.
+One must know the coefficients of a polynomial $t(x)$ to apply this theorem and then can perform very fast an e/i test for a disc centered at the origin.
+Such soft e/i tests were the basis of the subdivision root-finders of [16, 17], which apply them to polynomials $t(x)$ obtained from $p(x)$ of (1.1) by means of shifts of the variable $x$, scaling it by factors ranging from $R/2^{b}$ to $R=\max_{j-1}^{d}|z_{j}|$, and order of $\log(\log(d))$ root-squaring steps.
+To apply Pellet’s theorem to such a polynomial $t(x)$ one must compute the values $|\frac{t_{i}}{t_{d}}|$ for $i=0,1,\dots,d-1$.
+Because of scaling alone, the coefficient $t_{i}$ must be computed within $1/2^{b-i}$ to approximate the zeros of $p$ within $1/2^{b}$ for every $i$, $i=0,1,\dots,d$.
+As in the proof of Observation 22.2, it follows that one must process at least $0.5(d+1)db$ bits of the coefficients of $t(x)$ and must perform at least $0.25(d+1)db$ bit operations to support even a single e/i tests towards approximation of the zeros of $p(x)$ within $1/2^{b}$.
+Already a single subdivision iteration for Problem 1m can involve up to $4m$ soft e/i tests and hence up to at least $0.25(d+1)dbm$ bit-operations.
+This is by a factor of $d$ exceeds the bound of [78, 84] for approximation ofall $m$ roots in an isolated disc.
+Such a bit-count and the bit operation cost bounds for the root-finders of [78, 84, 17] decrease by a factor of $d$ where the minimal pairwise distance between the zeros of $p$ exceeds a positive constant; the ratio of the bounds of [17] and [78, 84] remains at least $d$.
+Similar argument shows that already for a singe e/i test representation of the leading coefficients $t_{d}$ of $t(x)$ alone under scaling involves order of $bd$ bits.
+Moreover, this is amplified to $d^{2}b$ for $t_{d}^{d}$ in every e/i test and up to $d^{2}bm$ per a subdivision iteration with $m$ e/i tests.
+It may be interesting that a shift by $c$ such that $|c|\leq 2$ is much less costly than scaling: [^21] We can relax the restriction $|c|\leq 2$ above by means of scaling the variable $x$, but as we showed above, scaling can greatly increase the bound on the overall number of bit operations involved.
+Theorem 10.3.
+Bit operation cost of the shift of the variable, [108, Lemma 2.3] (cf.
+[52, Lemma 3.6]).
+Given a positive $b$, the coefficients of a polynomial $p(x):=\sum_{i=0}^{d}p_{i}x^{i}$ such that $||p(x)||_{1}\leq 1$, and a complex $c$ such that $|c|\leq 2$, one can approximate within $1/2^{b}$ the coefficients $t_{0},t_{1},\dots,t_{d}$ of the polynomial $t(x):=\sum_{i=0}^{d}t_{i}x^{i}:=p(x-c)$ by performing $O(\mu((b+d)d))$ bit operations for $\mu(s)$ of (10.1).
+PART III: Classical subdivision root-finder with Cauchy-based novel e/i tests
+Next we elaborate upon Alg. 3.2 and analyze it.
+Organization of the presentation in Part III: We devote the next section to some basic definitions.
 In Sec.
-21 we devise and analyze our first compression algorithm.
-In particular we observe in Sec.
-21.4 that the whole subdivision process involves at most $12m-$ 1 compression steps, reduced to $O(m)$ e/i tests.
+12 we count roots in a disc on the complex plane by means of approximation of a Cauchy integral over its boundary circle.
 In Sec.
-22 we prove that in addition it also involves $O(m)$ e/i steps.
-We reach a near-optimal bound on the bit operation complexity of our root-finders in Sec.
-23 by devising and analyzing our compression algorithm and recalling our estimates for the complexity of an e/i step.
-In that section we also present a fast root-counter in an isolated disc.
-We use some results of Part I, Alg. 9.1 of Part II, and Sec. 13 (Background) of Part III. Otherwise our presentation in this part of the paper is self-contained.
-20 Component tree and compact components
-20.1 Component tree, edge length, non-adjacent suspect squares, and compact components
-Recall the definition of the component tree $\mathbb{T}$ of Sec.
-4, which represent partition of the union of suspect squares at every subdivision step and representation of such partition with the component tree $\mathbb{T}$ of Sec.
+13 we analyze the resulting root-counters and e/i tests in a disc.
+In Secs.
+14 we estimate the computational precision supporting our root-finders.
+In Sec.
+15 we estimate their bit operation complexity assuming that a polynomial $p(x)$ is represented with its coefficients.
+11 Background: basic definitions
+  • “Wlog" stands for “without loss of generality".
+  • $S(c,\rho):=\{x:~{}|\Re(c-x)|\leq\rho$, $|\Im(c-x)|\leq\rho\},~{}D(c,\rho):=\{x:~{}|x-c|\leq\rho\}$,
+    $C(c,\rho):=\{x:~{}|x-c|=\rho\}$, and $A(c,\rho,\rho^{\prime}):=\{x:~{}\rho\leq|x-c|\leq\rho^{\prime}\}$
+    denote a square, a disc, a circle (circumference), and an annulus (ring) on the complex plane, respectively.
+  • The largest upper bound on such a value $\theta$ is said to be the isolation of the disc $D(c,\rho)$, the circle $C(c,\rho)$, or the square $S(c,\rho)$, respectively (see Fig.
+4), and is denoted $i(D(c,\rho))$, $i(C(c,\rho))$, and $i(S(c,\rho))$, respectively.
+  • $r_{1}(c,t)=|y_{1}-c|,\dots,r_{d}(c,t)=|y_{d}-c|$ in non-increasing order are the $d$ root radii, that is, the distances from a complex point $c$ to the zeros $y_{1}\dots,y_{d}$ of a $d$th degree polynomial $t(x)$.
+$r_{j}(c):=r_{j}(c,p)$, $r_{j}:=r_{j}(0)$ for $j=1,\dots,d$.
+Observation 11.1.
+Eqn. (1.4) maps the zeros $z_{j}$ of $p(x)$ into the zeros $y_{j}=\frac{z_{j}-c}{\rho}$ of $t(y)$, for $j=1,\dots,d$, and preserves the index #$\#(D(c,\rho))$ and the isolation $i(D(c,\rho))$.
+12 Cauchy root-counting in a disc
+12.1 The power sums of the roots in a disc
+Generalize (3.7) as follows (cf. [1]):
+[…]
+For $\mathcal{D}=D(0,1)$ and $\zeta$ of (3.4) approximate $s_{h}$ with finite sums (cf. [106, 117])
+$s_{h,q}:=\frac{1}{q}\sum_{g=0}^{q-1}\zeta^{(h+1)g}~{}\frac{p^{\prime}(\zeta^{g})}{p(\zeta^{g})}~{}{\rm for}~{}h=0,1,\dots,q-1$ (12.2)
+and recall that $s_{0}(D(0,1))=\#(D(0,1))$ (cf. (3.7)), while an e/i test is just the decision whether #$\#(D(0,1))>0$.
+12.2 Cauchy sums in the unit disc as weighted power sums of the roots
+Cor. 12.1 of the following lemma expresses the $h$th Cauchy sum $s_{h,q}$ as the sum of the $h$th powers of the roots $z_{j}$ with the weights $\frac{1}{1-x^{q}}$.
+Lemma 12.1.
+For a complex $z$, two integers $h\geq 0$ and $q>1$, and $\zeta$ of (3.4), it holds that
+$\frac{1}{q}\sum_{g=0}^{q-1}\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=\frac{z^{h}}{1-z^{q}}.$ (12.3)
+Proof.
+First let $|z|<1$ and obtain
+$\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=\frac{\zeta^{hg}}{1-\frac{z}{\zeta^{g}}}=\zeta^{hg}\sum_{u=0}^{\infty}\Big{(}\frac{z}{\zeta^{g}}\Big{)}^{u}=\sum_{u=0}^{\infty}\frac{z^{u}}{\zeta^{(u-h)g}}.$
+The equation in the middle follows from Newman’s expansion
+$\frac{1}{1-y}=\sum_{u=0}^{\infty}y^{u}~{}{\rm for}~{}y=\frac{z}{\zeta^{g}}.$
+We can apply it because $|y|=|z|$, while $|z|<1$ by assumption.
+Sum the above expressions in $g$, recall that
+$\zeta^{u}\neq 1~{}{\rm and}~{}\sum_{s=0}^{q-1}\zeta^{u}=0~{}{\rm for}~{}0<u<q,~{}\zeta^{q}=1,$ (12.4)
+and deduce that
+$\frac{1}{q}\sum_{g=0}^{q-1}\frac{z^{u}}{\zeta^{(u-h)g}}=z^{u}$
+for $u=h+ql$ and an integer $l$, while $\sum_{g=0}^{q-1}\frac{z^{u}}{\zeta^{(u-h)g}}=0$ for all other values $u$. Therefore
+$\frac{1}{q}\sum_{g=0}^{q-1}\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=z^{h}\sum_{l=0}^{\infty}z^{ql}.$
+Apply Newman’s expansion for $y=z^{q}$ and deduce (12.3) for $|z|<1$. By continuity extend (12.3) to the case of $|z|=1.$
+Now let $|z|>1$. Then
+$\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=-\frac{\zeta^{(h+1)g}}{z}~{}~{}\frac{1}{1-\frac{\zeta^{g}}{z}}=-\frac{\zeta^{(h+1)g}}{z}\sum_{u=0}^{\infty}\Big{(}\frac{\zeta^{g}}{z}\Big{)}^{u}=-\sum_{u=0}^{\infty}\frac{\zeta^{(u+h+1)g}}{z^{u+1}}.$
+Sum these expressions in $g$, write $u:=ql-h-1$, apply (12.4), and obtain
+$\frac{1}{q}\sum_{g=0}^{q-1}\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=-\sum_{l=1}^{\infty}\frac{1}{z^{ql-h}}=-z^{h-q}\sum_{l=0}^{\infty}\frac{1}{z^{ql}}.$
+Apply Newman’s expansion for $y=z^{q}$ again and obtain
+$\frac{1}{q}\sum_{g=0}^{q-1}\frac{\zeta^{(h+1)g}}{\zeta^{g}-z}=-\frac{z^{h-q}}{1-\frac{1}{z^{q}}}=\frac{z^{h}}{1-z^{q}}.$
+Hence (12.3) holds in the case where $|z|>1$ as well. ∎
+Sum equations (12.3) for $z=z_{j}$ and $j=1,\dots,d$, combine Eqns. 3.1 and (12.2), and obtain
+Corollary 12.1.
+For the roots $z_{j}$ and all $h$, the Cauchy sums $s_{h,q}$ of (12.2) satisfy $s_{h,q}=\sum_{j=1}^{d}\frac{z_{j}^{h}}{1-z_{j}^{q}}$ unless $z_{j}^{q}=1~{}{\rm for~{}some}~{}j.$
+12.3 Approximation of the power sums in the unit disc
+Cor. 12.1 implies that the values $|s_{h,q-}s_{h}|$ decrease exponentially in $q-h$ if the unit disc $D(0,1)$ is isolated.
+Theorem 12.1.
+[106, Eqn.
+(12.10)].
+[^22] Unlike the proof in [106], we rely on Cor.
+12.1 of independent interest.
+Thm.
+12.1 does not follow from [117, Thms.
+2.1 and 2.2] but can possibly be proved with some techniques of [117].
+Let the unit circle $C(0,1)$ be $\theta$-isolated and let $d_{\rm in}$ and $d_{\rm out}=d-d_{\rm in}$ denote the numbers of the roots lying in and outside the disc $D(0,1)$, respectively.
+Then
+$|s_{h,q}-s_{h}|\leq\frac{d_{\rm in}\theta^{-h}+d_{\rm out}\theta^{h}}{\theta^{q}-1}\leq\frac{d\theta^{h}}{\theta^{q}-1}~{}{\rm for}~{}h=0,1,\dots,q-1.$ (12.5)
+Proof.
+Enumerate the roots $z_{j}$ so that $|z_{j}|<1$ if and only if $j\leq d_{\rm in}$; then Cor. 12.1 implies that
+$s_{h,q}-s_{h}=s_{h,q,{\rm in}}+s_{h,q,{\rm out}}$ (12.6)
+where
+$s_{h,q,{\rm in}}=\sum_{j=1}^{d_{\rm in}}z_{j}^{h}\Big{(}\frac{1}{1-z_{j}^{q}}-1\Big{)}~{}{\rm and}~{}s_{h,q,{\rm out}}=\sum_{j=d_{\rm in}+1}^{d}\frac{z_{j}^{h}}{1-z_{j}^{q}}=\sum_{j=d-d_{\rm out}+1}^{d}\frac{z_{j}^{h}}{1-z_{j}^{q}}.$ (12.7)
+Notice that $\frac{1}{1-z_{j}^{q}}-1=\frac{z^{q}}{1-z_{j}^{q}}$, and so $|z_{j}^{h}(\frac{1}{1-z_{j}^{q}}-1)|\leq\frac{\theta^{-h}}{\theta^{q}-1}$ for $z_{j}\in D(0,1)$, while $\Big{|}\frac{z_{j}^{h}}{1-z_{j}^{q}}\Big{|}\leq\frac{\theta^{h}}{\theta^{q}-1}$ for $z_{j}\notin D(0,1)$.
+Combine these bounds with Eqns.
+(12.6) and (12.7) and obtain (12.5).
+∎
+By choosing $q$ that exceeds $\log_{\theta}(2d+1)$ we ensure that $|s_{0}-s_{0,q}|<1/2$.
+Example 12.1.
+Let the unit circle $C(0,1)$ be 2-isolated. Then Thm. 12.1 implies that $|s_{0,q}-s_{0}|<1/2$ for $d\leq 1,000$ if $q=10$ and for $d\leq 1,000,000$ if $q=20$.
+Algorithm 12.1.
+INPUT: a black box polynomial $p$ of a degree $d$ and $\theta>1$.
+INITIALIZATION: Compute the integer $q=\lfloor\log_{\theta}(4d+2)\rfloor>\log_{\theta}(2d+1)$.
+COMPUTATIONS: Compute Cauchy sum $s_{0,q}$ and output an integer $\bar{s}_{0}$ closest to it.
+Observation 12.1.
+(i) Alg.
+12.1 evaluates $p(x)$ at $q=\lfloor\log_{\theta}(4d+2)\rfloor$ points.
+(ii) It outputs $\bar{s}_{0}=\#(D(0,1))$ if the circle $C(0,1)$ is $\theta$-isolated for $\theta>1$.
+(iii) If the algorithm outputs $\bar{s}_{0}>0$, then #$\#(D(0,\theta))>0$.
+Proof.
+Thm. 12.1 immediately implies claim (ii) but also implies that #$\#(D(0,1))>0$ if $\bar{s}_{0}>0$ unless #$\#(A(0,1/\theta,\theta))>0$. In both cases #$\#(D(0,\theta))>0$. ∎
+12.4 Extension to any disc
+Map (1.4), preserving #$\#(D(0,1))$ (cf. Observation 11.1), enables us to extend definition (12.2) of Cauchy sum to any disc $D(c,\rho)$ as follows:
+$s_{0,q}(p,c,\rho):=s_{0,q}(t,0,1):=\frac{\rho}{q}\sum_{g=0}^{q-1}\zeta^{g}~{}\frac{p^{\prime}(c+\rho\zeta^{g})}{p(c+\rho\zeta^{g})},$ (12.8)
+for $\zeta=\zeta_{q}$ of (3.4), that is, $s_{h,q}(p,c,\rho)$ is the Cauchy sums $s_{h,q}(t,0,1)$ of the zeros of the polynomial $t(y)=p(\frac{y-c}{\rho})$ in the unit disc $D(0,1)$.
+Observation 12.2.
+Given a polynomial $p(x)$, a complex $c$, a positive $\rho$, a positive integer $q$, and the $q$-th roots of unity, we can compute $s_{0,q}(p,c,\rho)$ by means of the evaluation of NIR$(\zeta^{g})$ for $g=0,1,\dots,q-1$, reduced to evaluation of $p$ at $2q$ points, and in addition performing $q+1$ divisions, $12q-$ 1 multiplications, and $12q-$ 1 additions.
+Refer to Alg.
+12.1 applied to the polynomial $t(y)$ of (1.4) for an isolation parameter $\theta>1$ and $q=\lfloor\log_{\theta}(4d+2)\rfloor$ as Alg.
+12.1c,ρ,q and also as Alg.
+12.1 applied to $\theta$-isolated circle $C(c,\rho)$ for $2q$ evaluation points.
+Combine Observations 11.1 and 12.2 to extend Observation 12.1 as follows.
+Proposition 12.1.
+(i) Alg.
+12.1c,ρ,q evaluates polynomial $p(x)$ at $q=\lfloor\log_{\theta}(4d+2)\rfloor$ points $x$ and in addition performs $25q-$ 2 ops.
+(ii) It outputs $\bar{s}_{0}=\#(D(c,\rho))$ if the circle $C(c,\rho)$ is $\theta$-isolated.
+(iii) If the algorithm outputs $\bar{s}_{0}>0$, then #$\#(D(c,\theta\rho))>0$.
+By applying Thm.
+12.1 we can estimate the errors $|s_{h,q}(t,0,1)-s_{h}(t,0,1)|$ of the approximation of the $h$th power sums $s_{h}(t,0,1)$ of the roots of $t(x)$ in the unit disc $D(0,1)$.
+Next we specify this for $h=1$.
+Proposition 12.2.
+(Cf. [48].) Let for $\theta>1$ a $\theta$-isolated disc $D(c,\rho)$ contain exactly $m$ roots. Fix a positive integer $q$ and define the Cauchy sum $s_{1,q}(p,c,\rho)$ of (12.8). Then
+$|mc+\rho s_{1,q}(p,c,\rho)-s_{1}(p,c,\rho)|\leq\frac{\rho d\theta}{\theta^{q}-1}.$
+13 Root-counting and $\ell$-tests
+13.1 Overview
+We seek fast root-counters and $\ell$-tests based on computing Cauchy sum $s_{0,q}$ for a smaller integer $q$, but the output can deceive us if we decrease $q$ below $m$ (see Example 7.1).
+Extensive tests for $q$ of order $\log^{2}(d))$ in [45, 48] have never encountered such unlucky inputs, and we only provide some indirect formal support for this empirical behavior in Sec. 7.4.
+In the next two subsections we support Las Vegas randomized root-counters and $\ell$-tests that evaluate polynomial $p$ at $O(m\log(d))$ points.
+Combine this estimate with Cor. 2.2 for $k$ of order $b$ and obtain Las Vegas randomized solution of Problems 0, 1, 0, and 1 at the cost of evaluation of the polynomial $p$ at $O(m^{2}b)$ points.
+13.2 A basic randomized root-counter
+Next we fix a parameter $\gamma>1$ and devise a root-counter for a disc $D(c,\rho)$ with a fixed center $c$ (wlog we let $c=0$) and a radius $\rho$ sampled at random in a fixed small range.
+Then we prove that under such a Las Vegas randomization the test fails with a probability at most $1/\gamma$.
+Algorithm 13.1.
+Basic randomized root-counter.
+INPUT: $\gamma\geq 1$ and a $d$-th degree black box polynomial $p$ having at most $m$ zeros in the disc $D(0,\sqrt{2})$.
+INITIALIZATION: Sample a random value $w$ in the range $[0.2,0.4]$ under the uniform probability distribution in that range and write $\rho:=2^{w}$ and $q=\lfloor 10m\gamma\log_{2}(4d+2)\rfloor$.
+COMPUTATIONS: Apply Alg. 12.10,ρ,q.
+Proposition 13.1.
+Alg. 13.1 queries evaluation oracle at $q=\lfloor 10m\gamma\log_{2}(4d+2)\rfloor$ locations and outputs the value $\bar{s}_{0}=\#(D(0,\rho))$ with a probability at least $1-1/\gamma$.
+Proof.
+Readily verify the bound on the number of query locations claimed in the proposition.
+By assumption, the annulus $A(0,1,\sqrt{2})$ contains at most $m$ roots.
+Hence at most $m$ root radii $r_{j}=2^{e_{j}}$, for $j=1,\dots,m^{\prime}\leq m$, lie in the range $[1,\sqrt{2}]$, that is, $0\leq e_{j}\leq 0.5$ for at most $m^{\prime}\leq m$ integers $j$.
+Fix $m^{\prime}$ intervals, centered at $e_{j}$, for $j=1,\dots,m^{\prime}$, each of a length at most $1/(5m^{\prime}\gamma)$.
+Their overall length is at most $1/(5\gamma)$.
+Let $\mathbb{U}$ denote their union.
+Sample a random $u$ in the range $[0.2,0.4]$ under the uniform probability distribution in that range and notice that Probability$(u\in\mathbb{U})\leq 1/\gamma$.
+Hence with a probability at least $1-1/\gamma$ the circle $C(0,\rho)$ is $\theta$-isolated for $\theta=2^{\frac{1}{10m\gamma}}$, in which case Alg.
+13.1 outputs $\bar{s}_{0}=s_{0}=\#(D(0,\rho))$ by virtue of claim (ii) of Prop.
+12.1.
+∎
+13.3 Refining the bound on the error probability
+When we apply Alg.
+13.1 $v$ times, the number of query locations of the evaluation oracle grows linearly – by a factor of $v$, while the bound on Las Vegas error probability of outputting errors decreases exponentially – from $1/\gamma$ to $1/\gamma^{v}$.
+First let $\ell=1$ or $\ell=m$ and then again reduce the root radius approximation problem to the decision problem of $\ell$-test.
+Namely, narrow Alg.
+13.1 to an $\ell$-test for a fixed $\ell$ in the range $1\leq\ell\leq m$, so that with a probability at least $1-1/\gamma$ this $\ell$-test outputs 0 if #$\#(D(0,1))<\ell$ and outputs 1 if #$\#(D(0,2^{0.5}))\geq\ell$.
+Moreover, output 1 is certified by virtue of claim (iii) of Prop.
+12.1 if $\ell=1$; similarly the output 0 is certified if $\ell=m$.
+Next extend the $\ell$-test by means of applying Alg. 13.1 for $v$ iid random variables $w$ in the range $0.2\leq w\leq 0.4$. Call this $\ell$-test Alg. $\ell$.
+[…]
+Recall Prop. 13.1 and then readily verify the following proposition.
+Proposition 13.2.
+(i) For two integers $1\leq\ell\leq m$ and $v\geq 1$ and a real $\gamma\geq 1$, Alg.
+$\ell$ evaluates $p(x)$ at $\lfloor 10m\gamma\log_{2}(4d+2)\rfloor v$ points; this evaluation involves more ops than the other stages of the algorithm, even in the cases of the Mandelbrot and sparse polynomials $p$ (cf.
+Observation 12.2).
+(ii) Output 1 of Alg.
+13.1v,1 and the output 0 of Alg.
+13.1v,$m$ are certified.
+(iii) Output 0 of Alg.
+13.1v,1 and output 1 of Alg.
+13.1v,$m$ are correct with a probability at least $1-1/\gamma^{v}$.
+The proposition bounds the cost of $\ell$-tests for $\ell=1$ (e/i tests) and $\ell=m$.
+This is sufficient for our root-finding, but our next natural extension covers $\ell$-tests for any integer $\ell$ in the range $1<\ell<m$.
+Assign the values 0 or 1 to the output integers $g$ as follows: compute the average of the $v$ outputs in $v$ independent applications of the test and output $g$ denoting the integer closest to that average; if the average is 1/2, write $g=0$ to break ties.
+This extension of Alg. $\ell$ is only interesting for $1<\ell<m$; otherwise its cost/error estimates are inferior to those of Prop. 13.2.
+Proposition 13.3.
+For an integer $\ell$ such that $1<\ell<m\leq d$ and real $v\geq 1$ and $\gamma>1$, Algs.
+$\ell$ queries evaluation oracle at $\lfloor 10m\gamma\log_{2}(4d+2)\rfloor v$ locations, and its output is correct with a probability at least $1-(4/\gamma)^{v/2}$.
+Proof.
+Since $D(0,1)\subset D(0,\sqrt{2})$, we only consider the following three cases:
+(i) #$\#(D(0,1))<\ell$ and #$\#(D(0,2^{0.5}))\geq\ell$,
+(ii) #$\#(D(0,2^{0.5}))<\ell$, and then also #$\#(D(0,1))<\ell$,
+(iii) #$\#(D(0,1))\geq\ell$, and then also #$\#(D(0,2^{0.5}))\geq\ell$.
+In case (i) the output $g$ is always correct.
+In case (ii) the output 0 is correct, while the output 1 is wrong and occurs if and only if more than $v/2$ tests go wrong, by outputting 1.
+Likewise, in case (iii) the output 1 is correct, while the output 0 occurs if and only if more than $v/2$ tests go wrong, by outputting 0.
+In both cases (ii) and (iii) we deal with the binomial distribution having failure probability $1/\gamma$, and so (cf.
+[19]) the probability $P$ of at least $v/2$ failures among $v$ tests is at most $\sum_{j=\bar{v}}^{v}\begin{pmatrix}v\\ j\end{pmatrix}T_{j}$ for $T_{j}=\gamma^{-j}(1-1/\gamma)^{v-j}\leq\gamma^{-j}$.
+Therefore, $P\leq\sum_{j=0}^{v}\begin{pmatrix}v\\ j\end{pmatrix}=T_{\bar{v}}2^{v}\max_{j}T_{j}$.
+Substitute $\sum_{j=0}^{v}\begin{pmatrix}v\\ j\end{pmatrix}=2^{v}$ and $\bar{v}=\lceil v/2\rceil$ and obtain $P<2^{v}T_{\bar{v}}\leq 2^{v}/\gamma^{v/2}=(4/\gamma)^{v/2}$.
+∎
+13.4 E/i test for an annulus
+The power sums of the roots lying in an annulus $A(c,\rho,\rho^{\prime})$ can be expressed as the difference of the power sums of the roots lying in the two discs $D(c,\rho^{\prime})$ and $D(c,\rho)$, and our algorithms and their analysis can be readily extended, based on Thm.
+12.1 supporting e/i tests for an annulus.
+We can immediately replace the $m$-test of Sec.
+18 with such e/i tests.
+14 Estimation of the precision of computing
+14.1 Doubling computational precision towards its a posteriori estimation
+As in Part II we can obtain near-optimal posteriori estimates for computational precision by recursively doubling it.
+Next we obtain its a priori estimate of $O(b\log(d))$ bits. The derivation is a little simpler than in Part II due to observations in the next subsection.
+14.2 Reduction of a priori estimates to approximation of $x$NIR$(x)$ within $\frac{3}{8}$
+In our e/i tests we approximate NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ with the first order divided difference of (3.2); whp we only need to apply the tests to $\theta$-isolated discs $D(c,\rho)$ for $\theta-1>0$ of order $\frac{1}{m}$ and $\rho\geq\epsilon=1/2^{b}$.
+Map (1.4) reduces such a test for $p(x)$ to a test in the disc $D(0,1)$ for the polynomial $t(x)=p(\frac{x-c}{\rho})$ replacing $p(x)$ and decreases the error estimate by a factor of $\rho$.
+Hence map (1.4) increases the precision bound by at most a factor of $b$, and we obtain the claimed precision bound in $O(b\log(d))$ in case of any disc $D(c,\rho)$ from the bound $O(\log(d))$ for the unit disc $D(0,1)$.
+The output of an e/i test is an integer, and we only need to certify that the output error bound of Alg.
+16.1 is less than $\frac{1}{2}$.
+The error can be magnified by a factor of $v$ in $v$ applications of Alg.
+16.1 but this factor is immaterial in our notation $\tilde{O}$.
+Our e/i tests compute the $q$ values $\rho x\frac{p^{\prime}(c+\rho x)}{p(c+\rho x)}=x\frac{t^{\prime}(x)}{t(x)}$, for $x=\zeta^{g}$, $\zeta$ of (12.2), and $g=0,1,\dots,q-1$, sum these values, and divide the sum by $q$.
+By slightly abusing notation write ${\rm NIR}(x):=\frac{t^{\prime}(x)}{t(x)}$ rather than ${\rm NIR}(x)=\frac{p^{\prime}(x)}{p(x)}$.
+According to straightforward error analysis (cf.
+[14, Part A of Ch.3] and the references therein) rounding errors introduced by these additions and division are dominated, and we can ignore them; then it is sufficient to compute $\zeta^{g}{\rm NIR}(\zeta^{g})$ within, say, $\frac{3}{8}$ for every $g$.
+Indeed, in that case the sum of $q$ error bounds is at most $\frac{3q}{8}$ and decreases below $\frac{1}{2}$ in division by $q$.
+14.3 Precision of representation of NIR$(x)$ within $\frac{1}{8}$
+Recall Eqn. (3.1) and represent $x{\rm NIR}(x)$ for $|x|=1$ within $\frac{1}{8}$ by using a precision of $O(\log(d))$ bits.
+Proposition 14.1.
+Write $C:=C(c,\rho)$ and assume that $|x|=1$ and the circle $C$ is $\theta$-isolated for $\theta>1$. Then $|x{\rm NIR}(x)|\leq\frac{d}{1-1/\theta}=\frac{\theta d}{\theta-1}$.
+Proof.
+Eqn.
+(3.1) implies that $|{\rm NIR}(x)|=|\sum_{j=1}^{d}\frac{1}{x-z_{j}}|\leq\sum_{j=1}^{d}\frac{1}{|x-z_{j}|}$ where $|x-z_{j}|\geq(1-1/\theta)\rho$ for $|z_{j}|<1$, while $|x-z_{j}|\geq(\theta-1)\rho$ for $|z_{j}|>1$ since $i(C)\geq\theta$.
+Combine these bounds, write $m:=\#(D(c,\rho))$, and obtain $|x{\rm NIR}(x)|\leq\frac{m}{1-1/\theta}+\frac{d-m}{\theta-1}\leq\frac{d}{1-1/\theta}$ for $|x|=1$.
+∎
+Corollary 14.1.
+One can represent $x{\rm NIR}(x)$ within $\frac{1}{8}$ for $|x|=1$ by using a precision of $3+\lceil\log_{2}(\frac{d}{1-1/\theta})\rceil$ bits, which is in $O(\log(d))$ provided that $\frac{1}{\theta-1}=d^{O(1)}$.
+14.4 What is left to estimate?
+Instead of $x{\rm NIR}(x)=\frac{xt^{\prime}(x)}{t(x)}$ we actually approximate $\frac{xt^{\prime}(y)}{t(x)}$ where $|y-x|\leq\delta$ and $t^{\prime}(y)$ is equal to the divided difference of (3.2).
+Thus, we shall increase the above error bound $\frac{1}{8}$ by adding to it the upper bounds $\frac{1}{8}$ on $\alpha:=|\frac{xt^{\prime}(x)}{t(x)}-\frac{xt^{\prime}(y)}{t(x)}|$ for $|x|=1$ and on the rounding error $\beta$ of computing $x\frac{t(x)-t(x-\delta)}{\delta t(x)}$.
+Next we estimate $\alpha$ in terms of $d$, $\theta$, and $\delta$ – by extending the proof of [59, Lemma 3.6], then readily ensure that $\alpha<\frac{1}{8}$ by choosing $\delta$ with a precision $\log_{2}(1/|\delta|)$ of order $\log(d)$, and finally estimate $\beta$.
+We simplify our estimates by dropping the factor $x$ where $|x|=1$.
+14.5 Precision of approximation with divided difference
+Recall the following lemma, which is [59, Fact 3.5].
+Lemma 14.1.
+For $t(x)=\prod_{j=1}^{d}(x-y_{j})$ and a non-negative integer $j\leq d$ it holds that
+$t^{(j)}(x)=j!~{}t(x)\sum_{S_{j,d}}~{}\prod_{j\in S_{j,d}}\frac{1}{x-y_{j}}$
+where the summation $\sum_{S_{j,d}}$ is over all subsets $S_{j,d}$ of the set $\{1,\dots,d\}$ having cardinality $j$.
+Proposition 14.2.
+Under the assumptions of Prop. 14.1 let $w=w(x):=\max_{j=1}\frac{1}{|x-y_{j}|}$. Then
+$|t(x)|\alpha=|\frac{t(x)-t(x-\delta)}{\delta}-t^{\prime}(x)|\leq|t(x)\frac{(d~{}w)^{2}\delta}{1-\delta~{}d~{}w}|$.
+Proof.
+The claimed bound on $|\xi^{\prime}|$ follows from (3.1).
+It remains to apply the first five lines of the proof of [59, Lemma 3.6] with $f$, $\alpha$, $n$, and $\xi$ replaced by $t$, $-\delta$, $d$, and $w$, s respectively.
+Namely, first obtain from Taylor’s expansion that $t(x)-t(x-\delta)=\sum_{j=0}^{\infty}\frac{\delta^{j}}{j!}t^{(j)}(x)-t(x)$.
+Substitute the expressions of Lemma 14.1 and obtain
+$t(x)-t(x-\delta)=\delta t^{\prime}(x)+\sum_{j=2}^{\infty}\delta^{j}t(x)\sum_{S_{j,d}}~{}\prod_{j\in S_{j,d}}\frac{1}{x-y_{j}}.$
+Combine this equation with the assumed bounds on $\frac{1}{|x-y_{j}|}$ and deduce that
+$\Big{|}\frac{t(x)-t(x-\delta)}{\delta}-t^{\prime}(x)\Big{|}\leq\Big{|}\frac{f(x)}{\delta}\sum_{j=2}^{\infty}\delta^{j}w^{j}d^{j}\Big{|}\leq\Big{|}f(x)\frac{(d~{}w)^{2}~{}\delta}{1-\delta~{}d~{}w}\Big{|}.$
+In our estimates for $\alpha$ and $\delta$ we apply Prop. 14.1 for $w\leq\frac{\theta}{\theta-1}$ and obtain
+Corollary 14.2.
+Under the assumptions of Prop. 14.1, let $w\leq\frac{\theta}{\theta-1}$. Then
+$\alpha<1/8~{}{\rm for}~{}|\delta|\leq\frac{(\theta-1)^{2}}{10d^{2}\theta^{2}}.$
+14.6 The impact of rounding errors
+To approximate $|x{\rm NIR}_{c,\rho}(x)|$ within $\frac{3}{8}$ it remains to approximate $\beta$ within $\frac{1}{8}$.
+Proposition 14.3.
+Suppose that a black box oracle has computed the values $t(y)$ within a relative error bounds $\nabla(y)$ for $y=x$, $y=x-\delta$, and $\delta$ of Cor.
+14.2.
+Then we can ensure that $|\beta|<\frac{1}{8}$ by choosing a proper $\nabla(y)$ of order $1/d^{O(1)}$, represented with a precision of $\log(\frac{1}{\nabla(y)})=O(\log(d))$ bits.
+Proof.
+Write $\beta\delta=\frac{t(x-\delta)(1+\nabla(x-\delta))}{t(x)(1+\nabla(x))}-\frac{t(x-\delta)}{t(x)}=\frac{t(x-\delta)}{t(x)}~{}\frac{\nabla(x-\delta)-\nabla(x)}{1+\nabla(x)}$.
+Taylor-Lagrange’s formula implies that $\frac{t(x-\delta)}{t(x)}=1+\frac{t(x-\delta)-t(x)}{t(x)}=1+\delta\frac{t^{\prime}(u)}{t(x)}$, for $u\in[x-\delta,x]$.
+Hence $|\frac{t(x-\delta)}{t(x)}|\leq 1+|(1+\alpha^{\prime})\delta|$ for $\alpha^{\prime}=\frac{t^{\prime}(u)-t(x)}{t(x)}$.
+By extending Cor. 14.2 obtain $|\alpha^{\prime}|\leq\frac{1}{8}$.
+Hence $|\frac{t(x-\delta)}{t(x)}|\leq w:=1+\frac{9|\delta|}{8}$, while $|\beta\delta|\leq w\frac{|\nabla(x)|+|\nabla(x-\delta)|}{1-|\nabla(x)|}$.
+Now choose $\nabla(y)$ such that $|\nabla(y)|\leq\frac{|\delta|}{17w}$ for $y=x$ and $y=x-\delta$, and so $w(|\nabla(x)|+|\nabla(x-\delta)|)\leq\frac{2}{17\delta}$.
+Then verify that $\log(\frac{1}{|\nabla(y)|})=O(\log(d))$, $1-|\nabla(x)|\geq 1-\frac{|\delta|}{17w}>\frac{16}{17}$, and $|\delta\beta|<\frac{|\delta|}{8}$.
+∎
+By combining Cors. 14.1 and 14.2 and Prop. 14.3 obtain
+Corollary 14.3.
+Suppose that a black box oracle can output approximations of a $d$-th degree polynomial $t(x)$ with any relative error in $1/d^{O(1)}$ for $|x|\leq 1$ and let the unit disc be $\theta$-isolated for $t(x)$ and for $\frac{1}{\theta-1}=d^{O(1)}$.
+Then we can perform e/i test for $t(x)$ on the unit disc by using a precision of $O(b\log(d))$ bits.
+15 Bit operation cost of Cauchy e/i tests and root-finding
+We omit extension of the precision bound to the bit complexity estimate for the eigenvalue problem because it can be verbatim the same as in Part II.
+15.1 Our goal and a basic theorem
+Next we estimate the bit operation complexity of our root-finders by combining them with fast multipoint evaluation of a polynomial by means of the algorithm of [32, 67] non-trivially extended in [52, 64] to support the following theorem:
+Theorem 15.1.
+Given a positive $b$, the coefficients of a polynomial $t(x):=\sum_{i=0}^{d}t_{i}x^{i}$ such that $||t(x)||_{1}=\sum_{i=0}^{d}|t_{i}|\leq 2^{\tau}$, and $q$ complex points $x_{1},\dots,x_{q}$ in the unit disc $D(0,1)$, one can approximate the values $t(x_{1}),\dots,t(x_{q})$ within $1/2^{b}$ by using $\tilde{O}((d+q)(b+\tau))$ bit operations.
+Proof.
+For $q=d$ this is [64, Thm.
+2], based on using novel data structure for hyperbolic approximation of $p$, proposed and analyzed in [64].
+Trivially extend this theorem to the cases where $q>d$ (by considering $t(x)$ a polynomial of degree $q$ with $q-d$ leading coefficients 0) and where $q<d$ (by removing $d-q$ points $x_{i}$, with no increase of the complexity bound).
+∎
+We only apply this theorem for $\tau=0$, in which case its bit operation bound turns into
+$\mathbb{B}_{\rm Moroz}=\tilde{O}((d+q)b)$ (15.1)
+For comparison, the previous State of the Art can be represented by the bound
+$\mathbb{B}_{\rm Kirr}=\tilde{O}((d+q)(b+q)),$ (15.2)
+due to [52, Thm.
+3.9, Alg.
+5.3, and Appendix A.3] based on combining Kirrinnis’s and Schönhage’s study of the bit operation cost of basic polynomial computations (cf.
+[106]) with the algorithm of [32, 67].
+[^23] Precision of computing required in the latter algorithm grows fast as $d$ increases; e.g., the IEEE standard double precision of the customary numerical computations is not enough to support that algorithm already for $d>50$; for $b=O(\log(d))$ one can fix this deficiency by applying FMM (see Remark 3.1) to the equivalent task of multiplication of Vandermonde matrix by a vector.
+One can, however, support bounds of Eqns.
+(15.1) and (15.2) on the overall bit operation complexity by computing with multiple precision for larger values $b$.
+Moroz obtained [64, Thm.
+2] by applying bound (15.2) to his lower degree polynomials approximating $p(x)$.
+15.2 Acceleration with fast multipoint polynomial approximation
+Given a complex $c$, a pair of positive $b$ and $\rho$, and $d+1$ coefficients of a polynomial $p=p(x)$ of (1.1) such that the disc $D(c,\rho)$ is isolated and contains precisely $m$ roots, we approximate all these roots within $R/2^{b}$ for $R=|c|+\rho$ by applying subdivision root-finding iterations with e/i tests of Sec.
+13.
+Then every subdivision step is reduced to application of e/i tests to $\bar{m}=O(m)$ discs $D(c_{\lambda},\rho_{\lambda})$ for $\rho_{\lambda}\geq R/2^{b}$ and all $\lambda$.
+Such a test, based on our e/i test of Sec.
+13, amounts to approximation of $p(x)$ at $O(m\log(d))$ equally spaced points $x$ on each of $v$ circles sharing the center $c_{\lambda}$ and having radii in the range $[\rho_{\lambda}/\sigma,\rho_{\lambda}\sigma]$ for $1<\sigma<\sqrt{2}$.
+We only use Thm.
+15.1 for $\tau=0$ to cover the case where a polynomial $t(x)$ with $||t(x)||_{1}\leq 1$ is evaluated at points $x_{j}\in D(0,1)$.
+To ensure these assumptions we first scale the variable $x\mapsto Rx$ to map the discs $D(c_{\lambda},\rho_{\lambda}\sigma)\mapsto D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})\subseteq D(0,1)$, for $\bar{c}_{\lambda}=c_{\lambda}/R$, $\bar{\rho}_{\lambda}=\rho_{\lambda}/R\geq 1/2^{b}$, and all $\lambda$.
+Then we write $t(x):=p(Rx)/\psi$ for $\psi=O(R^{d}||p(x)||_{1})$ such that $||t(x)||_{1}=1$ and approximate $t(x)$ within $1/2^{\bar{b}}$ or equivalently within the relative error bound $1/(|t(x)|2^{\bar{b}})$ at $\bar{q}:=qv$ points $x$ in $\bar{m}$ e/i tests.
+[^24] Scaling $p$ and $x$ above is done once and for all e/i tests; to compensate for this we only need to change our bound on the output precision $b$ by $\lceil|\log_{2}R|\rceil$ bits.
+Due to Cor. 14.3 it is sufficient to ensure the bound $1/(|t(x)|2^{\bar{b}})=1/d^{O(1)}$; we deduce it by choosing any $\bar{b}\geq\log(\frac{1}{|t(x)|})+O(\log(d))$.
+In our e/i tests we only need to approximate $t(x)$ at the points $x$ lying on $\theta_{\lambda}$-isolated circles where $\frac{1}{\theta_{\lambda}-1}=O(m)$.
+Map the disc $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ into the unit disc $D(0,1)$.
+This does not change isolation of the boundary circle and may increase the norm $||t(x)||_{1}$ by at most factor of $2^{b+d}$ because $|\bar{c}_{\lambda}|\leq 1$ and $\bar{\rho}_{\lambda}\geq 1/2^{b}$.
+Now substitute $||t(x)||_{1}\leq 2^{b+d}$ and $\log(\frac{||t(x)||_{1}}{|t(x)|})=O(d\log(m))$ (cf.
+[106, Eqn.(9.4)]) [^25] This bound is readily deduced from the following non-trivial result [106, Thm.
+4.2]: Let $t(x)=f\prod_{j=1}^{m}(x-z_{j})\prod_{j=m+1}^{d}(\frac{x}{z_{j}}-1)$, $||t(x)||_{1}=1$, $|z_{j}|<1$ for $j\leq m$, and $|z_{j}|>1$ for $j>m$.
+Then $|f|\geq 1/2^{d}$.
+and obtain that $\bar{b}=\log(\frac{1}{|t(x)|})+O(\log(d))=O(d\log(m)+b)$.
+Remark 15.1.
+For $m=d$ the above bound on $\log(\frac{||t(x)||_{1}}{|t(x)|})$ of [106, Eqn.(9.4)] is sharp up to a constant factor and is reached for the polynomial $t(x)=(x+1-\frac{1}{m})^{d}/(2-\frac{1}{m})^{d}$, $\rho_{\lambda}=1$, and $x=-1$.
+For $m<d$ [106, Thm.
+4.5] implies a little stronger bounds, namely, $\log(\frac{||t(x)||_{1}}{|t(x)|})=O(d+m\log(m))$ and hence $\bar{b}=O(m\log(m)+d+b)$ because all discs $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ lie in $\theta$-isolated unit disc $D(0,1)$ for $\theta-1$ exceeding a positive constant.
+15.3 The bit operation complexity estimates
+Apply Thm.
+15.1 with $q$ and $b$ replaced by $\bar{q}=O(m^{2})$ and $\bar{b}=O(b+d+m\log(m))$, respectively, [^26] Under the Random Root Model of Sec.
+1.5 we can whp decrease this bound by a factor of $m$ and then extend this decrease to the related subsequent estimates.
+and obtain that the bit operation cost of our $O(m)$ e/i tests of Sec.
+13 at any fixed subdivision step is in $\tilde{O}((b+d+m)m^{2})$ for $1\leq m\leq d$.
+This implies the upper estimate
+$\mathbb{B}_{\rm roots}=\tilde{O}((b+d)m^{2}b)$ (15.3)
+for the overall bit operation cost of our $O(b)$ subdivision steps, which includes the time-complexity of the evaluation oracle.
+Based on [52, Alg.
+5.3] and Eqn.
+(15.2) obtain a little larger upper bound on the bit operation cost of our subdivision root-finders:
+$\mathbb{B}_{\rm roots}^{\prime}=\tilde{O}((m^{2}+d+b)(m^{2}+d)b).$ (15.4)
+PART IV: Acceleration of Root-Finding With Compression
+In this part of the paper we decrease the number of subdivision steps of our root-finders of Parts II and III by a factor of $b/\log(b)$, thus arriving at our Main Theorems 1 – 4.
+In Sec.
+17 we extend the outline of this acceleration from Sec.
 4.
-The tree has at most $m$ components-leaves that represent the output $\epsilon$-discs and at most $m-1$ other vertices, which represent components-parents, each broken by a subdivision step into at least two components-children.
-The initial suspect square is the root of the tree $\mathbb{T}$.
-The edges of the tree connect parents and children, representing subdivision steps between partitions of components.
-Cor.
-20.1 below bounds the number of such steps; one can consider this number the edge length.
-Call a pair of suspect squares non-adjacent if they share no vertices.
-Call a component compact if it contains just a single non-adjacent suspect square, that is, consists of at most four suspect squares sharing a vertex (see Fig. 10).
+It relies on an algorithm for compression of an isolated disc, which we devise and analyze also in Sec.
+17.
+The algorithm involves estimation of the $\ell$th smallest root radius, which is of independent interest and is covered in Sec.
+16.
+It is reduced to recursive application of an $\ell$-test; we present it in Sec.
+13 for any positive $\ell\leq d$ and any input disc, but in Sec.
+18 we specify simpler tests that are sufficient in application to compression.
+In Sec.
+20 we estimate the overall complexity of the resulting root-finder, which intertwines at most $22m-$ 2 compression steps with e/i tests, whose number we estimate in Sec.
+19.
+In Sec.
+21 we further accelerate polynomial root-finding for a large class of inputs.
+16 Estimation of the $\ell$-th smallest root radius
+In this section we narrow a fixed range $[0,R]$ for the $\ell$-th smallest root radius $r_{d-\ell+1}(c)$ from a fixed complex point $c$.
+To simplify our exposition we let $R=1$ and $c=0$, wlog due to (1.4).
+Problem RR.
+Compute a range $[\rho_{-},\rho_{+}]$ such that $0\leq\rho_{-}\leq r_{d-\ell+1}\leq\rho_{+}$ and $\rho_{+}\leq\epsilon=1/2^{b}$ or $\rho_{+}\leq\beta\rho_{-}$ for fixed $b>0$ and $\beta=2^{\phi}>1$, provided that we are given a positive integer $\ell\leq d$, initial range $[0,1]$ for $r_{d-\ell+1}$, and a $\sigma$-soft $\ell$-test for a fixed $\sigma>1$.
+We first use this test to check whether $r_{d-\ell+1}\leq 1/2^{b}$.
+If so, we are done.
+Otherwise we seek $r_{d-\ell+1}$ in the range $[\frac{1}{2^{b}\sigma},1]$.
+We apply binary search, said to be bisection of logarithms or BoL, to the range $[-\bar{b},0]=[-\log_{2}(\frac{\sigma}{\epsilon}),\log_{2}(1)]$, for $\bar{b}=\log_{2}(\frac{\sigma}{\epsilon})$, to approximate $\log_{2}(r_{d-\ell+1})$.
+Since the test is soft, so is BoL as well: for a range $[b_{-},b_{+}]$ with midpoint $\widehat{b}=0.5(b_{+}+b_{-})$ regular rigid bisection continues the search in the range $[b_{-},\widehat{b}]$ or $[\widehat{b},b_{+}]$, but our soft BoL continues it in the range $[b_{-},(1+\alpha)\widehat{b}]$ or $[(1-\alpha)\widehat{b},b_{+}]$ for a fixed small positive $\alpha$.
+Accordingly we slightly increase the range of search for $r_{d-\ell+1}$ to $[-\bar{b},0]$ for $\bar{b}:=\log_{2}(\frac{\sigma}{\epsilon})$.
+BoL still converges linearly and just a little slower than the regular rigid bisection: the approximation error decreases by a factor of $\frac{2}{1+\alpha}$ rather than by twice.
+[^27] If we applied BoL to the range $[0,1]$, we would have to bisect the infinite range $[-\infty,0]$, but we deal with the narrower range $[-b,0]$ since we have narrowed the range $[0,1]$ to $[\epsilon,1]$.
+Algorithm 16.1.
+Bracketing the $\ell$-th smallest root radius based on $\ell$-test.
+INPUT: An integer $\ell$ such that $1\leq\ell\leq d$ and #$\#(D(0,1))\leq\ell$, $\sigma=1+\alpha$ for $20<\alpha<1/$ 2, a $\sigma$-soft $\ell$-test0,ρ for a positive $\rho$ (see Sec.
+13.4 for such tests), and two tolerance values $\beta:=2^{\phi}>1$ and $\epsilon=1/2^{b}$, $b\geq 1$.
+OUTPUT: A solution $[\rho_{-},\rho_{+}]$ of Problem RR.
+INITIALIZATION: Write $\bar{b}:=\log_{2}(\frac{\sigma}{\epsilon})$, $b_{-}:=-\bar{b}$, and $b_{+}:=0$.
+[…]
+$\rho_{-}=2^{b_{-}},~{}\rho_{+}=2^{b_{+}},~{}\rho:=2^{\widehat{b}},~{}{\rm and}~{}\widehat{b}:=\lceil(b_{-}+b_{+})/2\rceil.$ (16.1)
+Write $b_{-}:=(1-\alpha)\widehat{b}$ if the $\ell$-test outputs 0; write $b_{+}:=(1+\alpha)\widehat{b}$ if the test outputs 1. Then update $\widehat{b}$ and $\rho$ according to (16.1).
+Continue recursively: stop and output the current range $[\rho_{-},\rho_{+}]=[2^{b_{-}},2^{b_{+}}]$ if $b_{+}-b_{-}\leq\phi$, that is, if $\rho_{+}\leq\beta\rho_{-}$; otherwise update $\widehat{b}$ and $\rho$ and reapply the computations.
+The disc $D(0,\epsilon)$ contains at least $\ell$ roots, that is, the range $[0,\epsilon]$ brackets the $\ell$th smallest root radius $r_{d-\ell+1}$ if the $\sigma$-soft $\ell$-test applied to the disc $D(0,\frac{\epsilon}{\sigma})$ outputs 1.
+Otherwise, the disc $D(0,\frac{\epsilon}{\sigma})$ contains less than $\ell$ roots, and so the root radius $r_{d-\ell+1}$ lies in the range $[\frac{\epsilon}{\sigma},1]$, in which case we continue our search.
+Readily verify that our policy of recursively updating the range $[b_{-},b_{+}]$ maintains this property throughout and that updating stops where $b_{+}-b_{-}\leq\phi$, that is, where $\rho_{+}\leq\beta\rho_{-}$.
+This implies correctness of Alg.
+16.1.
+Let us estimate the number $\mathbb{I}$ of the invocations of an $\ell$-test.
+Proposition 16.1.
+Alg.
+16.1 invokes an $\ell$-test $\mathbb{I}$ times where $\mathbb{I}\leq\lceil\log_{\nu}(\frac{\bar{b}}{\phi})\rceil+1$ for $\nu=\frac{2}{1+\alpha}$, and so $\mathbb{I}=O(\log(b))$ provided that $\nu-1$, $\sigma-1$, and $\phi$ are positive constants.
+Proof.
+Alg.
+16.1 stops if the $\ell$-test outputs 1.
+Otherwise the test decreases the length $\bar{b}=b_{+}-b_{-}$ of its input range $[b_{-},b_{+}]$ for the exponent to $0.5(1+\alpha)\bar{b}$.
+$i$ applications change it into $\frac{(1+\alpha)^{i}}{2^{i}}\bar{b}$ for $i=1,2,\dots$, and the proposition follows.
+∎
+Remark 16.1.
+In the case where we know that #$\#(D(0,\rho))\geq\ell$ for a fixed $\rho$ in the range $[\epsilon,1]$, we only need to refine the initial range $[\rho,1]$ for the root radius $r_{d-\ell+1}$.
+Clearly, $\mathbb{I}$ applications of $\ell$-test increase the number of evaluation points for the ratio $\frac{p^{\prime}(x)}{p(x)}$ by at most a factor of $\mathbb{I}$, and if the $\ell$-test is randomized, then the error probability also grows by at most a factor of $\mathbb{I}$.
+Let us specify these estimates for Alg.
+16.1.
+Corollary 16.1.
+Let precisely $\ell$ zeros of a black box polynomial $p(x)$ of a degree $d$ lie in the unit disc, that is, let #$\#(D(0,1))=\ell$, and let $i(D(0,1))\geq\sigma$, for $\sigma-1$ exceeding a positive constant.
+Given two positive tolerance values $\epsilon=1/2^{b}$ and $\phi$ and a black box $\sigma$-soft $\ell$-test, which evaluates the polynomial $p(x)$ at $\mathcal{NR}$ points, Alg.
+16.1 approximates the $\ell$-th smallest root radius $|z_{j}|$ within the relative error $2^{\phi}$ or determines that $|z_{j}|\leq\epsilon$, at the cost of evaluation of polynomial $p(x)$ at $\mathbb{I}\cdot\mathcal{NR}$ points for $\mathbb{I}$ estimated in Prop.
+16.1, $\mathbb{I}=O(\log(b))$ provided that $\nu-1$ and $\phi$ are positive constants.
+17 Compression of an isolated disc
+17.1 A motivation: compact components in the union of suspect squares
+As in Sec.
+4, partition the union of suspect squares at every subdivision iteration into components and associate breaking (partition of) components into sub-components with component tree, which has at most $22m-$ 2 edges.
+Next we will bound the numbers of subdivision steps (and hence of e/i tests) at and between partitions of components.
+We call a pair of suspect squares non-adjacent if they share no vertices.
+Call a component compact if it contains just a single non-adjacent suspect square, that is, consists of at most four suspect squares sharing a vertex (see Fig.
+5).
+Figure 5: Roots (asterisks) define compact components. A subdivision step halves their diameters and at least doubles isolation of their minimal covering squares and discs.
 Figure 6: Two discs superscribing two non-adjacent suspect squares are not minimal but do not overlap.
-Observation 20.1.
+Observation 17.1.
 (See Fig. 6.) The $\sigma$-covers of two non-adjacent suspect squares are separated by at least the distance $a\Delta$ for the side length $\Delta$ and separation coefficient
@@
 Corollary 17.1.
-A component is broken in at most $1+\lfloor 0.5+\log_{2}(m)\rfloor$ successive subdivision steps unless it is compact.
-20.2 Isolation of compact components
-Observation 20.2.
+A component of suspect squares is broken in at most $1+\lfloor 0.5+\log_{2}(m)\rfloor$ successive subdivision steps unless it is compact.
+Observation 17.2.
 (i) A compact component is surrounded by a frame of discarded squares (see Fig.
-10) and hence is covered by a 2-isolated square.
+5) and hence is covered by a 2-isolated square.
 (ii) If it stays compact in $h$ successive subdivision steps, then its minimal covering square and disc become $2^{h+1}$-isolated and $2^{h}\sqrt{2}$-isolated, respectively.
@@
 ∎
-21 $\epsilon$-compression: algorithm and analysis
-21.1 Compression algorithm: high level description
-Recall that the classical subdivision iterations converge slowly where they successively process only compact components (cf.
-Fig.
-8), but then we accelerate this process by interrupting it with compression steps in at most three iterations.
-Next we specify compression for a 3.5-isolated unit disc $D=D(0,1)$ covering a component; map (1.3) extends compression to any 3.5-isolated disc $D(c,\rho)$; the map $\rho\epsilon\mapsto\epsilon$ adjusts the error tolerance $\epsilon$ by scaling by $\rho$.
-Observation 20.2 implies that we can readily ensure $\theta$-isolation at the cost of performing $O(\log(\theta))$ additional subdivision steps.
-Algorithm 21.1.
-$\epsilon$-compression (high level description).
-INPUT: a positive $\epsilon\leq 1$ and a black box polynomial $p(x)$ of a degree $d$ such that [^34] Newton’s and Schröder’s iterations of [128, 101] accelerate convergence of subdivision iterations to a root cluster in a disc if the disc is $5d^{2}$-isolated, while Alg.
-21.1 only requires 3.5-isolation (see (21.1)).
-$i(D(0,1))\geq 3.5~{}{\rm and}~{}\#(D(0,1))\geq 1.$ (21.1)
-OUTPUT: an integer $m$, a complex $c^{\prime}$, and a positive $\rho^{\prime}$ such that $m=\#(D(0,1))$, (i) $X(D(c^{\prime},\rho^{\prime}))=X(D(0,1))$, and (ii) $\rho^{\prime}\leq\epsilon$ or $8|z_{g}-z_{h}|>5\rho^{\prime}$ for a pair of roots $z_{g},z_{h}\in D(c^{\prime},\rho^{\prime})$.
-[Hence the disc $D(c^{\prime},\rho^{\prime})$ is 5/8-rigid.]
-INITIALIZATION: Fix $\delta\geq 0$ such that $8\delta\leq\epsilon\leq 1$.
-COMPUTATIONS: Compute $m=\#(D(0,1))$, $c^{\prime}$, and $\rho^{\prime}$ such that
-$|c^{\prime}-c|\leq\delta~{}{\rm and}~{}\rho\leq\rho^{\prime}\leq\max\{\epsilon,\rho+2\delta\}$ (21.2)
-for $\rho:=r_{d-m+1}(c^{\prime})$ and some $c\in CH(X(D(0,1)))$.
-21.2 Correctness of compression algorithm
-First prove correctness in a simple case where $\epsilon=\delta=0$, $c^{\prime}=c$, $\rho^{\prime}=\rho$.
-In this case $c^{\prime}=c\in CH(X(D(0,1)))\subset D(0,1)$, and so $\rho^{\prime}=\rho=r_{d+1-m}(c^{\prime})\leq 2$. Hence $D(c^{\prime},\rho^{\prime})\subseteq D(0,3)$.
-Furthermore, #$\#(D(c^{\prime},\rho^{\prime}))=m$ since $\rho^{\prime}=\rho:=r_{d+1-m}(c^{\prime})$, while #$\#(D(0,1))=m$ by assumption.
-Now (21.1) implies that #$\#(D(0,3))=m$, which supports claim (i) for the output of Alg. 21.1.
-Next let $|c^{\prime}-z_{g}|=\rho:=r_{d+1-m}(c^{\prime})$ for a root $z_{g}$.
-Then all $m$ roots in $D(0,1)$ also lie in the disc $D(c^{\prime},\rho)=D(c^{\prime},\rho^{\prime})$ by virtue of claim (i), just proved.
-Since $c^{\prime}=c\in CH(X(D(0,1))$, however, there is a root, $z_{h}$, lying in $D(c^{\prime},\rho^{\prime})$ but not in $D(z_{g},\rho)$, such that $|z_{g}-z_{h}|>\rho\geq\rho^{\prime}-2\delta$ (cf.
-(21.2)).
-Hence $|z_{g}-z_{h}|>3\rho^{\prime}/4$ (implying claim (ii) for the output of Alg. 21.1) because $42\delta\leq\epsilon/4<\rho^{\prime}/$ 4.
-Next extend this correctness proof to the general case based on bounds (21.2).
-Figure 10: Roots (asterisks) define compact components. A subdivision step halves their diameters and at least doubles isolation of their minimal covering squares and discs.
-Notice that $c^{\prime}\in D(0,1+\delta)$ because $c\in CH(X(D(0,1)))\subset D(0,1)$ and $|c^{\prime}-c|\leq\delta$.
-Hence $\rho=r_{d-m+1}(c^{\prime})\leq 2+\delta$, and so $\rho^{\prime}\leq\rho+2\delta\leq 2+3\delta$ (cf. (21.2)) and $D(c^{\prime},\rho^{\prime})\subseteq D(0,3+4\delta)$.
-Therefore $D(c^{\prime},\rho^{\prime})\subseteq D(0,3.5)$ because $8\delta\leq\epsilon\leq 1$ by assumption.
-Now recall that #$\#(D(0,3.5))=\#(D(0,1))=m$, while #$\#(D(c^{\prime},\rho^{\prime}))\geq\#(D(c^{\prime},\rho))=m$ for $\rho^{\prime}\geq\rho$ and deduce claim (i) for the output of Alg. 21.1.
-Next assume wlog that $c=0$, $c^{\prime}=\delta$, and $z_{g}=c^{\prime}+\rho$ for a root $z_{g}$.
-Then $\Re(z_{g})\geq\rho-\delta\geq\rho^{\prime}-3\delta$ because $|c^{\prime}|=|c^{\prime}-c|\leq\delta$ and $\rho^{\prime}\leq\rho+2\delta$ by assumptions.
-Furthermore, $\Re(z_{h})<0$ for some root $z_{h}$ because $c=0\in CH(X(D(0,1)))$.
-(Actually, $\Re(z_{h})$ reaches its maximum $-\frac{z_{h}}{m-1}$ where $z_{h}$ is a simple root and $-\frac{z_{h}}{m-1}$ is a root of multiplicity $m-1$.)
-Therefore, $|z_{g}-z_{h}|\geq\Re(z_{g})-\Re(z_{h})>\rho^{\prime}-3\delta\geq\rho^{\prime}-\frac{3}{8}\epsilon$.
-This implies claim (ii) for the output of Alg. 21.1 since $\rho^{\prime}>\epsilon$.
-21.3 Elaboration upon the compression algorithm
-We readily (i) count roots in isolated input disc $D(0,1)$ of Alg.
-21.1 by applying the fast root-counter of Sec.
-23.4 or alternatively Alg.
-7.1 and Prop.
-7.1 and (ii) compute a complex point $c^{\prime}$ lying in or near the convex hull of the roots in $C$ by applying fast algorithm of Sec.
-23.
-Finally, we estimate the root radius $\rho^{\prime}:=r_{d-m+1}(c^{\prime})$ by applying Alg.
-9.1, which incorporates the $m$-test of of Sec.
-23.4, or alternatively Alg.
-7.1 and Prop.
-7.1.
-21.4 Counting $\epsilon$-compression steps
-Recall from Sec.
-1.3 that the output disc $D(c^{\prime},\rho^{\prime})$ of Alg.
-21.1 is said to be an $\epsilon$-disc if $\rho^{\prime}\leq\epsilon$; in this case call the compression by Alg.
-21.1 final and output that disc.
-Otherwise call the square $S=S(c^{\prime},\rho^{\prime})$ suspect and apply to it subdivision iterations.
-Extend Lemma 20.1 to the case where $|z_{g}-z_{h}|\geq\frac{5}{8}\rho^{\prime}$ for a pair of roots $z_{g},z_{h}\in D(c^{\prime},\rho^{\prime})$, combine it with claim (ii) for the output of Alg.
-21.1, and obtain that at most $2.5+\log_{2}(8\sigma m/5)$ subdivision steps break the square $S$ into more than one component.
-From part (ii) of Sec. 4 recall the bound $22m-$ 2 on the number of compression steps and obtain
+Cor.
+17.1 bounds the number of subdivision steps between partitions of non-compact components, but compact components can be processed very slowly (see Fig.
+3).
+For a remedy, we intertwine subdivision process with compression of MCD of a component that stays compact in three subdivision steps, in which case the MCD is $4\sqrt{2}$-isolated by virtue of Observation 17.2.
+Next we elaborate upon such compression of MCDs.
+17.2 The compression problem and flowchart of its solution
+Problem COMPR. INPUT: a positive $\epsilon\leq 1$, a constant $\eta>0$, a black box polynomial $p(x)$ of a degree $d$, and an isolated disc $D$ with a non-empty root set.
+OUTPUT: an integer $m$, a complex $c\in D$, and positive $\rho$ and $\eta$ such that #$\#(D(c,\rho))=\#(D)$ and the disc $D(c,\rho)$ is $\eta$-rigid unless $\rho\leq\epsilon$.
+Flowchart of a compression algorithm.
+To solve Problem COMPR we successively compute (a) $m=\#(D)$, (b) complex point $c$ lying in or near the minimal covering disc of the root set $X(D)$, and (c) positive $\rho$ and and $\eta$ such that #$\#(D(c,\rho))=m$ and the disc $D(c,\rho)$ is $\eta$-rigid unless $\rho\leq\epsilon$.
+Wlog assume that $D$ is a unit disc $D(0,1)$ and (a) compute $m=\#(D)$ for isolated unit disc $D:=D(0,1)$ by using $O(\log(d))$ queries of an evaluation oracle (see Observations 12.1 and 12.2).
+Next we elaborate upon the algorithms for stages (b) and (c).
+17.3 A point in or near the minimal covering disc for the $d$ roots
+Definition 17.1.
+A disc is said to be strongly isolated if it is $d^{\nu}$-isolated for a sufficiently large constant $\nu$ of our choice. In particular every disc containing all $d$ roots is strongly isolated.
+Next we successively compress the unit disc $D=D(0,1)$ in the s where it contains $d$ roots, is strongly isolated, and is just isolated.
+Schröder’s iterations [104],
+$z=y-m~{}{\rm NR}(y),$ (17.1)
+fast converge to a root of multiplicity $m$ [63, Sec. 5.4] and for $m=d$ outputs a point in or near the MSD of the the set of the $d$ roots if point $y$ lies far enough from that set:
 Theorem 17.1.
-For a fixed $\epsilon\geq 0$ let subdivision and compression steps be applied to a component $C$ of suspect squares containing at most $m$ roots.
-Then the computations involve at most $66m-$ 6 compact components, made up of at most $2424m-$ 24 suspect squares; that is, these computations involve at most $2424m-$ 24 e/i tests.
+Write $D:=D(0,1)$ and $\Delta:=\Delta(X(D))$, let #$\#(D)=d$, fix $y$ such that $|y|>2$, let $z:=y-d~{}{\rm NR(y)}$ and $\bar{x}\in MCD(D)$.
+Then $|z-\bar{x}|\leq\alpha~{}\Delta$ for $\alpha:=\frac{(|y|-1)^{2}}{(y|-2)^{2}}=(1+\frac{1}{(y|-2})^{2}$, and so $\alpha=2.25$ for $|y|=4$, $\alpha=25/16$ for $|y|=6$, and $\alpha\mapsto 1$ as $|y|\mapsto\infty$.
 Proof.
-A compact component stays unbroken and uncompressed in at most $3$ consecutive subdivision steps, which end with a final compression step or follow a non-final compression step and end with breaking a component.
-Hence all such components are processed in at most $66m-$ 6 subdivision step, each involving at most 4 e/i tests.
-∎
-22 Counting e/i tests
-22.1 An upper estimate
-The upper estimate $22m-$ 2 of part (ii) of Sec.
-4 on the number of compression steps is sharp – reached in the case of the complete binary tree $\mathbb{T}$.
-Each compression step is preceded by an e/i test, and so we perform at least $22m-$ 2 e/i steps.
-According to the following result, which is the main theorem of this section, this lower bound on the number of e/i tests involved is sharp up to a constant factor:
-Theorem 22.1.
-Let subdivision iterations with compression be applied to a suspect square containing at most $m$ roots. Then they involve $O(m)$ e/i tests overall.
-22.2 The first lemma
-We begin our proof of Thm. 22.1 with
-Lemma 22.1.
-In every component $C$ made up of $\sigma(C)$ suspect squares we can choose $n(C)\geq\sigma(C)/5$ non-adjacent suspect squares.
-Proof.
-[…]
-22.3 The second lemma
-Lemma 22.2.
-Fix two integers $i\geq 0$ and $g\geq 3$, fix a value $\sigma$ of softness, $1<\sigma<\sqrt{2}$, and define the separation fraction $a:=2-\sigma\sqrt{2}$ of Observation 20.1.
-For all $j$ let $j$-th subdivision step process $v_{j}$ components made up of $w_{j}$ suspect squares, $n_{j}$ of which are non-adjacent, let $\Delta_{j}$ denote their side length, and let the $j$-th steps perform no compression for $j=i,i+1,\dots,i+g$.
-Then
-$w_{i+g}\geq(n_{i}-v_{i})a2^{g-0.5}~{}{\rm if}~{}\Delta_{i}\geq 2^{g}\Delta_{i+g}.$ (22.1)
-Proof.
-At the $i$th step a pair of non-adjacent suspect squares of the same component is separated by at least the gap $a\Delta_{i}$, and there are at least $n_{i}-v_{i}$ such gaps with overall length $(n_{i}-v_{i})a\Delta_{i}$.
-Neither of these gaps disappears at the $(i+g)$th step, and so they must be covered by $w_{i+g}$ suspect squares, each of diameter $\sqrt{2}~{}\Delta_{i+g}=\Delta_{i}/2^{g-0.5}$ under (22.1).
-Hence $(n_{i}-v_{i})a\Delta_{i}\leq w_{i+g}\Delta_{i}/2^{g-0.5}$.
-∎
-Next we extend Lemma 22.2 by allowing compression steps.
-We only need to extend the equation $2\Delta_{j+1}=\Delta_{j}$ to the case where the $j$th subdivision step for $i\leq j\leq i+g$ performs non-final compression, each followed with subdivision steps applied to a single suspect square with side length at most $4\Delta_{j}$.
-The next two subdivision steps perform only five e/i tests, making up at most $105m-$ 10 e/i tests performed at all pairs of subdivision steps following all (at most $m-2$) non-final compression steps.
-It remains to bound by $O(m)$ the number of e/i tests in the remaining subdivision steps, not interrupted with compression steps.
-Furthermore, after a compression step we resume subdivision process with a suspect square whose side length bound decreases in two next subdivision steps from at most $4\Delta_{j}$ to at most $\Delta_{j}$.
-Hence we can apply to it Lemma 22.2, because now it accesses the same number $n_{j}$ of non-adjacent suspect squares and an increased number $\sigma_{j}$ of all suspect squares (cf. (22.1)).
-22.4 Counting e/i tests at partition steps
-Let $h$ denote the height of our partition tree $\mathbb{T}$ and count suspect squares processed in subdivision steps that partition components, temporarily ignoring the other steps.
-Clearly, bounds (22.1) still hold and can only be strengthened. Notice that $v_{i}\geq 1$ for all $i$, sum $v_{i+g}$ as well as bounds (22.1) for a fixed $g>0$ and $i=0,1,\dots,h-g$, and obtain
-$\sum_{i=0}^{h-g}n_{i}-\lambda\sum_{i=0}^{h-g}\sigma_{i+g}\leq\sum_{i=0}^{h-g}v_{i+g}~{}{\rm for}~{}\lambda:=\frac{\sqrt{2}}{a2^{g}}.$ (22.2)
-Notice that
-$\sum_{i=0}^{h-g}v_{i+g}\leq\sum_{i=0}^{h}v_{i}-\sum_{i=0}^{g-1}v_{i}<2m-g,$
-Combine the above estimate and the bounds $n_{i}\leq m$ and (cf. Lemma 22.1) $w_{i}\leq 5n_{i},$ for all $i$, write $n^{\prime}:=\sum_{i=0}^{h}n_{i}$, obtain
-$\sum_{i=0}^{h-g}n_{i}\leq n^{\prime}-\sum_{i=h-g+1}^{h}n_{i}\leq n^{\prime}-mg,~{}\sum_{i=0}^{h-g}w_{i+g}\leq\sum_{i=0}^{h}w_{i}\leq 5n^{\prime},$
-and so
-[…]
-Obtain that
-$0.5n^{\prime}\leq 2m-1-g+mg~{}{\rm for}~{}g=1+\Big{\lceil}\log_{2}\Big{(}\frac{5\sqrt{2}}{a}\Big{)}\Big{\rceil}.$
-Hence
-$n^{\prime}\leq 4m-2-2g+2mg,~{}{\rm and~{}so}~{}n^{\prime}=O(m)~{}{\rm for}~{}g=O(1).$ (22.3)
-22.5 Extension to counting all e/i steps
-Next consider a non-compact component, count all suspect squares processed in all subdivision steps applied until it is partitioned, and then prove in Lemma 22.3 that up to a constant factor the overall number of these suspect squares is bounded by their overall number at all partition steps.
-Such a result extends to all subdivision steps the bound $n^{\prime}=O(m)$ of (22.3), so far restricted to partition steps.
-Together with Lemma 22.1 this will complete our proof of Thm.
-22.1.
-Lemma 22.3.
-Let $k+1$ successive subdivision steps with $\sigma$-soft e/i tests be applied to a non-compact component for $\sigma$ of our choice in the range $1<\sigma<\sqrt{2}$.
-Let the $j$-th step output $n_{j}$ non-adjacent suspect squares for $j=0,1,\dots,k$.
-Then $\sum_{j=0}^{k-1}n_{j}\leq n_{k}\frac{10\sqrt{2}}{a}$ for the separation fraction $a:=2-\sigma\sqrt{2}$ (cf.
-Observation 20.1.)
-Proof.
-Substitute $g=k-i$ into (22.3) and obtain $w_{k}\geq(n_{i}-v_{i})a2^{k-i-0.5}$, for $i=0,1,\dots,k-1$.
-Recall that $w_{j}\leq 5n_{j}$ (cf. Lemma 22.1) and that in our case $v_{i}=1$ and $n_{i}\geq 2$ and obtain that $n_{i}-v_{i}\geq n_{i}/2$ for all $i$.
-Hence $5n_{k}\geq n_{i}a2^{k-i-1.5}$ for all $i$, and so
-$\sum_{i=0}^{k-1}n_{i}\leq n_{k}\frac{10\sqrt{2}}{a}\sum_{j=1}^{k}2^{-j}<n_{k}\frac{10\sqrt{2}}{a}.$
-Remark 22.1.
-The overhead constants hidden in the "O" notation of our estimates are fairly large but overly pessimistic.
-E.g., (i) the roots $x$ and $y$ involved in the proof of Lemma 22.2 are linked by a path of suspect squares, but we deduce the estimates of the lemma by using only a part of this path, namely, the part that covers the gaps between non-adjacent suspect squares.
-(ii) Our proof of that lemma covers the worst case where a long chains between the roots $x$ and $y$ can consist of up to $4m$ suspect squares, but for typical inputs such chains are much shorter and can consist of just a few suspect squares.
-(iii) In the proof of Lemma 22.3 we assume that $n_{i}-1\geq n_{i}/2$; this holds for any $n_{i}\geq 2$, but the bounds are strengthened as $n_{i}$ increases.
-(iv) In a component made up of five suspect squares at least two are non-adjacent, and so the bound of Lemma 22.1 can be strengthened at least in this case.
-Remark 22.2.
-Alg.
-21.1 compresses an input disc $D(0,1)$ into $\eta$-rigid disc for $\eta=5/8$, but our proof of Thm.
-22.1 only involves $\eta$ in the argument in Sec.
-21.4, and this argument and therefore the whole proof of the theorem can be immediately extended if it relies on compression into any rigid disc.
-23 A point near root set and root-count in a disc
-We prove that the algorithms of this section work where an input disc is $d^{\nu}$-isolated for a fixed positive constant $\nu$.
-In application to compression we begin with an isolated disc, say, 3.5-isolated, but can turn it into $d^{\nu}$-isolated for any fixed positive constant $\nu$ by applying $h$ additional subdivision steps or $h$-lifting for $h=O(\log(d))$.
-This comment can be extended to application of e/i test of Sec.
-23.4 instead of Cauchy-based e/i tests of Part II.
-23.1 Computation of a point near root set: introductory comments
-Recall Schröder’s iterations [132],
-$z=y-m~{}{\rm NR}(y),$ (23.1)
-which fast converge to a root of multiplicity $m$ [80, Sec.
-5.4].
-Due to map (1.3) we only need to consider the case where the input disc $D$ is the unit disc $D(0,1)$.
-We first assume that it contains all $d$ roots and readily verify
-Observation 23.1.
-Schröder’s single iteration (23.1) for $m=d$ and $p(x)=p_{d}(x-z_{1})^{d}$ outputs the root $z_{1}$ unless $y=z_{1}$.
-Next we extend this observation by proving that the output $z$ of Schröder’s iteration (23.1) lies near the convex hull $CH(X(D))$ provided that $m=d$, a disc $D$ contains all $d$ roots, and point $y$ lies far enough from the disc $D$.
-Here are our specific estimates.
-23.2 The case where a disc contains all $d$ roots
-Theorem 23.1.
-Write $D:=D(0,1)$, $CH:=CH(D)$, and $\Delta:=\Delta(D)$, fix $y$ such that $|y|>2$, let $z:=y-d~{}{\rm NR(y)}$, which is (23.1) for $m=d$.
-Let $\bar{x}\in CH$.
-Then $|z-\bar{x}|\leq\alpha~{}\Delta$ for $\alpha:=\frac{(|y|-1)^{2}}{(y|-2)^{2}}=(1+\frac{1}{(y|-2})^{2}$, and so $\alpha=2.25$ for $u|=4$ and $\alpha=25/16$ for $u|=6$, $\alpha\mapsto 1$ $|y|\mapsto\infty$.
-Proof.
-Recall that $|\bar{x},z_{1},\dots,z_{d}\in D(0,1)$, $|z_{j}|\leq 1$ and $|\bar{x}-z_{j}|\leq\Delta$ for $j=1,\dots,d$. Hence (cf. (3.1))
+Recall that $\bar{x},z_{1},\dots,z_{d}\in X(D)$, $|z_{j}|\leq 1$ and $|\bar{x}-z_{j}|\leq\Delta$ for $j=1,\dots,d$. Hence (cf. (3.1))
 $\Big{|}{\rm NIR}(y)-\frac{d}{y-\bar{x}}\Big{|}\leq\sum_{j=1}^{d}\Big{|}\frac{1}{y-z_{j}}-\frac{1}{y-\bar{x}}\Big{|}\leq\sum_{j=1}^{d}\frac{|\bar{x}-z_{j}|}{|y-\bar{x}|\cdot|y-z_{j}|}\leq\frac{\Delta d}{(|y|-1)^{2}}.$
@@
 Corollary 17.2.
-Under the assumptions of Thm. 23.2 the disc $D(z,r_{1}(z))$ is $\frac{1}{2+2\alpha}$-rigid.
+Under the assumptions of Thm. 17.2 the disc $D(z,r_{1}(z))$ is $\frac{1}{2+2\alpha}$-rigid.
 Proof.
 Verify the claim in the case where $x_{d}=0$, $x_{1}=\Delta=1$, e.g., $0\leq x_{j}\leq 1$ for all $j$, and $z=1+\alpha$ and notice that this is a worst case input. ∎
-Now we devise our fast compression algorithm inthe case where $m=d$.
+17.4 Compression algorithm for a disc containing $d$ roots
+Next we compress a disc containing $d$ roots towards their minimal covering disc.
 Algorithm 17.1.
 $\epsilon$-compression 1.
-INPUT: a positive tolerance $\epsilon=1/2^{b}\leq 1$ and a black box polynomial $p(x)$ of a degree $d$ such that
+INPUT: a positive $\epsilon=1/2^{b}\leq 1$ and a black box polynomial $p(x)$ of a degree $d$ such that
 #$\#(D(0,1))=d.$ (17.4)
-OUTPUT: a complex $z$, and a positive $r$ such that #$\#(D(z,r))=d$ and either $r\leq\epsilon$ or otherwise the disc $D(z,r)$ is $\frac{2}{13}$-rigid.
+OUTPUT: a complex $z$, and a positive $r$ such that #$\#(D(z,r))=d$ and the disc $D(z,r)$ is $\frac{2}{13}$-rigid unless $r\leq\epsilon$.
 INITIALIZATION: Fix $y\geq 4$.
@@
 (i) Compute and output $z=y-d~{}{\rm NR}~{}(y)$.
-(ii) Compute an approximation $\tilde{r}\geq 0$ to $r_{1}(z)$ by applying Alg. 9.1 for $m=d$, a constant $\beta>1$, and $\epsilon$ replaced by $\epsilon/4$.
-If $\tilde{r}\leq 3\epsilon/4$, output $z$ and $r=\epsilon$, defining $\epsilon$-compression into the disc $D(z,r)$ of radius $\epsilon$.
-Otherwise output $r=\tilde{r}+0.25\epsilon$, defining compression into rigid disc $D(z,r)$.
+(ii) Compute an approximation $\tilde{r}\geq 0$ to $r_{1}(z)$ by applying Alg. 16.1 for $\epsilon$ replaced by $\epsilon/4$.
+Output $z$ and choose $r$ to define $\epsilon$-compression into the disc $D(z,r)$.
+Namely, output $r=\epsilon$ if $\tilde{r}\leq 3\epsilon/4$.
+Otherwise output $r=\tilde{r}+0.25\epsilon$, then the disc $D(z,r)$ is rigid.
 Correctness of the algorithm is readily verified if Alg.
@@
 Otherwise $1/\beta\leq\tilde{r}/r_{1}\leq\beta$ for a constant $\beta\geq 1$ of Alg.
-9.1, while the disc $D(x,r_{1})$ is $\frac{1}{2\alpha+2}$-rigid by virtue of Cor.
+16.1, while the disc $D(x,r_{1})$ is $\frac{1}{2\alpha+2}$-rigid by virtue of Cor.
 17.2 where $\alpha\leq 2.25$ by virtue of Thm.
@@
 Remark 17.1.
-By the same argument we can ensure that unless $\tilde{r}\leq 3\epsilon/4$ the disc $D(z,r_{1}(z))$ is $\eta$-isolated for any $\eta>4$ if we initialize Alg. 23.1 at a sufficiently large $y$.
-This compression queries evaluation oracle twice while performing Schröder’s single iteration for computing $z$ and otherwise runs at the cost of performing Alg.
-9.1, estimated in Cor.
-9.1.
-We can decrease the latter cost bound for a large class of polynomials $p$ by choosing larger tolerance $\epsilon$.
-Then we will yield desired compression if the algorithm outputs $r$ exceeding $\epsilon/4$.
-Otherwise we can recursively repeat the computations for tolerance recursively decreased to $\epsilon/4$.
-23.3 Extension to any well-isolated disc
-Now suppose that #$\#(D(0,1))=m<d$ and $i(D(0,1))\geq\theta$ for a sufficiently large $\theta$ of order $d^{O(1)}$.
-Then verify that the $d-m$ external roots, lying outside the disc $D(0,1)$, little affect the proof of Thm.
-23.1 and our whole study in the previous subsection.
-Thus we first perform $O(\log(d))$ subdivision steps (see Observation 20.2).
-We do not need compression if these steps break the component into more than one component or compress it into an $\epsilon$-disc.
-Otherwise the component remains compact and hence in $O(\log(d))$ subdivision steps shrinks enough to become $d^{\nu}$-isolated for any fixed constant $\nu$, in which case we apply our compression algorithm.
-23.4 A fast root-counter and a fast e/i test based on Schröder’s iterations
-Let the unit disc be sufficiently strongly isolated and let a single Schröder’s iteration (23.1) be applied at a point $y$ lying sufficiently far from the unit disc $D(0,1)$.
-Then our study in this subsection implies that the iteration outputs a point $z=z(m)$ lying in or near the disc $D(0,1)$, say, in the disc $D(0,2)$, if and only if #$\#(D(0,1))=m$, that is, we count roots in $D(0,1)$ at the cost of evaluation of $p(x)$ at two points.
-Indeed, if #$\#(D(0,1))=m$, then, as we have showed in this section, the iteration outputs a point $z=z(m)$ lying near the convex hull of the root set in $D(0,1)$ and hence lying in or near the disc $D(0,1)$, say, in the disc $D(0,2)$ if the value $|y|$ is large enough.
-In this case $|m~{}{\rm NR}(y)-y|\leq 2$, and so $|{\rm NR}(y)-y|/m\leq 2/m$.
-If #$\#(D(0,1))=\bar{m}\neq m$, then the iteration outputs a point $z=z(m)$ that lies at the distance at least $|\bar{m}-m|\cdot|{\rm NR}(y)|-2$ from the disc $D(0,2)$, which must be large if $|y|$ is sufficiently (although reasonably) large.
-Extension to a fast e/i test.
-Now suppose that the disc $D(0,1)$ is sufficiently well isolated and apply this root-counter to the polynomial $xp(x)$ for a large $y$ and $m=1$, in which case Schröder’s iterations turn into Newton’s.
-In this case root-counter will show that the disc $D(0,1)$ contains no roots of $p(x)$ if and only if its output lies in that disc.
-23.5 Overall complexity of our root-finders
-We have reduced our root-finding in an isolated disc to performing at most $12m-$ 1 compression steps and in addition $O(m)$ e/i tests.
-We have also reduced a compression step to (i) performing $O(\log(d))$ additional subdivision steps (included to strengthen isolation of input discs for compression), (ii) application of Schröder’s single iteration, which evaluates polynomial $p(x)$ at two points, and (iii) performing $O(\log(b))$ times $m$-tests for approximation of the root radius $r_{d-m+1}(z,p)$.
-Overall our root-finders apply e/i tests $O(m\log(d))$ times, $m$-tests $O(m\log(b))$ times, and at most $22m-$ 2 Schröder’s iterations.
-Clearly, the computational cost of Schröder’s is dominated, while in Sec.
-15 we reduced an $m$-test is to $O(1)$ e/i tests, specified in Sec.
-14.
-Based on the latter results we can strengthen our estimates of Sec.
-18 and arrive at the complexity bounds of our Main Theorems.
-Alternatively, instead of estimating the cost of e/i tests and reduction to them of $m$- tests, we can use our estimates for the complexity of $\ell$- tests of Part II for $\ell=1$ and $\ell=m$ and would still arrive at our Main Theorems 2 – 4.
-Part V: Factorization, Root Isolation, and Functional Iterations
-24 Polynomial Factorization, its links to Root-finding, and some bit-complexity estimates
-24.1 Polynomial Factorization and its links to Root-finding (the results)
+If we initialize Alg. 17.1 at a sufficiently large $y$, then the same argument ensures that the disc $D(z,r_{1}(z))$ is $\eta$-isolated for any $\eta>4$ unless $\tilde{r}\leq 3\epsilon/4$.
+This compression queries evaluation oracle twice while performing Schröder’s single iteration for computing $z$ and otherwise runs at the cost of performing Alg. 16.1, estimated in Cor. 16.1.
+We decrease the latter cost for a large class of polynomials $p$ by choosing a larger tolerance $\epsilon$.
+Then we yield desired compression if the algorithm outputs $r$ exceeding $\epsilon/4$.
+Otherwise we recursively repeat the computations for the tolerance recursively decreasing to $\epsilon/4$.
+17.5 Compression of an isolated disc
+If $m<d$ and $i(D)\geq d^{\nu}$ for a sufficiently large constant $\nu$, then Eqn.
+(3.1) implies that the $d-m$ external roots, lying outside the disc $D$, little affect the proof of Thm.
+17.1 and our whole study in the previous two subsections as long as we apply Schröder’s iteration at a complex point $y$ lying far enough from both disc $D$ and the circle $C(0,d^{\nu})$.
+The MCD of a compact component can be isolated but not strongly.
+Then, however, we apply subdivision steps to that component and in order of $\log(d)$ steps have it broken or strongly isolated by virtue of Observation 17.2.
+Hence we can extend our study in the previous subsections.
+$O(\log(d))$ subdivision steps for compact components involve $O(m\log(d))$ e/i tests.
+The factor $\log(d)$ is immaterial for the $\tilde{O}$ estimates of our Main Theorems, but we remove it next.
+17.6 From isolation to strong isolation
+Let $s_{1}$ denote the sum of all $m$ roots in $D$.
+Clearly, the point $s_{1}/m$ lies in MCD and even in the convex hull of the root set $X(D)$.
+We compute approximation $c=s_{1}(b^{\prime})/m$ of $s_{1}/m$ within $1/2^{b^{\prime}}$ at the cost of $O(b^{\prime})$ queries of an evaluation oracle (see Thm.
+12.1).
+For $b^{\prime}$ of order $\log(d)$ we yield error bound $1/d^{\nu}$ for a fixed positive constant $\nu$ of our choice at the cost of $O(\log(d))$ queries of an evaluation oracle 12.1).
+By applying Alg. 16.1 we compute approximation $\rho$ of $r_{d-m+1}(c)$ also within $1/d^{\nu}$, and so
+$|c-s_{1}/m|\leq 1/d^{\nu}~{}{\rm and}~{}|\rho-r_{d-m+1}(c)|\leq 1/d^{\nu}.$ (17.5)
+The rigidity $\eta=\eta(D(c,\rho))$ of the disc $D(c,\rho)$ is given by the value $\frac{1}{\rho}\max^{*}|x_{i}-x_{j}|$ where maximum is over all pairs of roots lying in the disc $D(c,\rho)$.
+Now notice that
+$\max^{*}|x_{i}-x_{j}|>0.5~{}r_{d-m+1}(s_{1}/m).$ (17.6)
+Indeed, because of invariance of the claim under map (1.4), we can assume that $s_{1}=0$ and $x_{i}<0$. Then $\Re(x_{j})>0$ for some $j$, and (17.6) follows.
+Furthermore, $|r_{d-m+1}(s_{1}/m)-r_{d-m+1}(c)|\leq|s_{1}/m-c|$; combine this bound with Eqn.
+(17.5) and obtain that $|r_{d-m+1}(s_{1}/m)-\rho|\leq 2/d^{\nu}$.
+Together with (17.6) this estimate implies that
+$\max^{*}|x_{i}-x_{j}|>0.5(\rho-2/d^{\nu}).$
+If $\rho\geq 4/d^{\nu}$, then $\rho-2/d^{\nu}\geq\rho/2$, the disc $D(c,\rho)$ is 0.25-rigid, and compression is completed.
+If $\rho\leq 4/d^{\nu}$, then the disc $D(c,\rho)$, sharing with the unit disc $D(0,1))$ all its $m$ roots, is strongly isolated, namely, is $d^{\nu}/4$-isolated; then we can apply to it the compression algorithms of the previous sections.
+18 Reduction of $m$-test to e/i tests
+To complete our compression algorithms we should supply a $\sigma$-soft $\ell$-test for the input of Alg. 16.1, for which we can apply the low cost Las Vegas $\ell$-test in Sec. 13.
+In application to compression, however, we can apply two alternative tests because we only need either (i) a $\sigma$-soft $m$-test for a sub-disc $D(0,\rho)$ of the unit disc $D(0,1)$ containing $m$ roots where $\rho<1$ or alternatively (ii) e/i test for the annulus $A(0,\rho,1)$ of Sec.
+13.4.
+(i) A $\sigma$-soft $m$-test for a sub-disc.
+Clearly, we can immediately reduce $d$-test in the unit disc $D(0,1)$ for $p(x)$ to performing 1-test (that is, e/i test) for that disc and for the reverse polynomial $p_{\rm rev}(x)$ of (1.5), but next we reduce $m$-test for any $m$, $1\leq m\leq d$, and for any disc $D$ to bounded number of e/i tests for sub-discs.
+If at least one of the e/i tests fails, then definitely there is a root in a narrow annulus about the boundary circle of an input disc, and then the internal disc of the annulus would not contain $m$ roots.
+Otherwise we obtain a root-free concentric annulus about the boundary circle of an input disc and then reduce $m$-test to root-counting in an isolated disc, which we can perform at a dominated cost (see claim (ii) of Prop.
+12.1).
+Algorithm 18.1.
+Reduction of $m$-test to e/i tests.
+INPUT: a polynomial $p(x)$ of a degree $d$, two values $\bar{\theta}>1$ and $\bar{\sigma}>1$, an integer $m$ such that $1\leq m\leq d$ and #$\#(D(0,\bar{\sigma}))\leq m$, and two algorithms $\mathbb{ALG}~{}0$ and $\mathbb{ALG}~{}1$, which for a black box polynomial perform $\bar{\sigma}$-soft e/i test for any disc and root-counting in an isolated disc, respectively.
+INITIALIZATION (see Fig. 7): Fix a positive $\rho$ such that $\bar{\sigma}=O(\rho)$, say, $\rho=\bar{\sigma}/5$, and compute the integer
+$v=\lceil 2\pi/\phi\rceil$ (18.1)
+where $\phi$ is the angle defined by an arc of $C(0,1)$ with end points at the distance $\rho$, and so $v=O(1)$ for $\bar{\sigma}=O(\rho)$. Then compute the values
+$c_{j}=\exp(\phi_{j}\sqrt{-1}),~{}j=0,1,\dots,v,$ (18.2)
+such that $|c_{j}-c_{j-1}|=\rho$ for $j=1,2,\dots,v$ and $\phi_{j}-\phi_{j-1}=\phi$.
+$c_{1}$$c_{2}$$c_{3}$$c_{4}$$c_{5}$$c_{6}$$c_{7}$$c_{8}$$c_{9}$$c_{0}$
+Figure 7: $v=9$, $|c_{j}-c_{j-1}|=\rho>|c_{9}-c_{0}|$ for $j=1,2,\dots,9$.
+OUTPUT: Either an upper bound $r_{d-m+1}<1-\rho\frac{\sqrt{3}}{2}$ or one of the two lower bounds $r_{d-m+1}\geq 1-\bar{\sigma}\rho$ or $r_{d-m+1}>1+(\sqrt{3}-1)\rho$.
+COMPUTATIONS: Apply algorithm $\mathbb{ALG}~{}0$ (soft e/i test) to the discs $D(c_{j},\rho)$, $j=0,1,\dots,v$.
+(a) Stop and certify that $r_{d-m+1}\geq 1-\bar{\sigma}\rho$, that is, #$\#(D(0,1-\bar{\sigma}\rho))<m$ for the open disc $(D(0,1-\bar{\sigma}\rho)$, unless algorithm $\mathbb{ALG}~{}0$ outputs exclusion in all its $v+1$ applications.
+(b) If it outputs exclusion $v+1$ times, then apply algorithm $\mathbb{ALG}~{}1$ (root-counter) to the $\theta$-isolated disc $D(0,1)$ for
+${}\theta:=\min\Big{\{}1+(\sqrt{3}-1)\rho,\frac{1}{1-0.5\rho\sqrt{3}}\Big{\}}.$ (18.3)
+Let $\bar{s}_{0}=\#(D(0,1))$ denotes its output integer. Then conclude that
+(i) $r_{d-m+1}\leq 1-\rho\frac{\sqrt{3}}{2}$, and so #$\#(D(0,1-\rho\frac{\sqrt{3}}{2}))=m$ if $\bar{s}_{0}=m$, while
+(ii) $r_{d-m+1}\geq 1+(\sqrt{3}-1)\rho$, and so #$\#(D(0,1+(\sqrt{3}-1)\rho))<m$ for the open disc $D(0,1+(\sqrt{3}-1)\rho)$ if $\bar{s}_{0}<m$.
+To prove correctness of Alg. 18.1 we need the following result.
+Theorem 18.1.
+For the above values $c_{0},\dots,c_{v}$, and $\rho$, the domain $\cup_{j=0}^{v-1}D(c_{j},\rho)$ covers the annulus $A(0,1-0.5\rho\sqrt{3},1+(\sqrt{3}-1)\rho)$.
+Proof.
+Let $c_{j-1}$ and $c_{j}$ be the centers of two neighboring circles $C(c_{j-1},\rho)$ and $C(c_{j},\rho)$, let $z_{j}$ and $z^{\prime}_{j}$ be their intersection points such that $|z_{j}|=|z_{j}^{\prime}|-|z_{j}-z^{\prime}_{j}|$ for $j=1,2,\dots,v$, observe that the values $|z_{j}-z_{j}^{\prime}|$, $|z_{j}|$, $|z^{\prime}_{j}|$, and $|c_{j-1}-c_{j}|$ are invariant in $j$, and hence so is the annulus $A(0,|z_{j}|,|z_{j}^{\prime}|)$ of width $|z_{j}-z_{j}^{\prime}|$ as well; notice that it lies in the domain $\cup_{j=0}^{v-1}D(c_{j},\rho)$.
+$c_{0}$$c_{1}$$z$$z^{\prime}$
+Figure 8: $|c_{1}-c_{0}|=|c_{0}-z|=|c_{1}-z|=|c_{0}-z^{\prime}|=|c_{1}-z^{\prime}|=\rho$.
+Let us prove that $|z_{j}-z_{j}^{\prime}|=\rho\sqrt{3}$ for $j=1$ and hence also for $j=2,\dots,v$.
+Write $z:=z_{1}$ and $z^{\prime}:=z_{1}^{\prime}$ and observe (see Fig.
+8) that $|c_{1}-c_{0}|=|c_{0}-z|=|c_{1}-z|=|c_{0}-z^{\prime}|=|c_{1}-z^{\prime}|=\rho$, and so the points $c_{0}$, $c_{1}$, and $z$ are the vertices of an equilateral triangle with a side length $\rho$.
+Hence $|z-z^{\prime}|=\rho\sqrt{3}$.
+The circle $C(c_{j},\rho)$ intersects the line segment $[z_{j},z_{j}^{\prime}]$ at a point $y_{j}$ such that $|y_{j}-z_{j}|>\rho\sqrt{3}/2$ and $|y_{j}-z_{j}^{\prime}|>(\sqrt{3}-1)\rho$ for $\rho<1$, and this implies the theorem if $|c_{v}-c_{0}|=\rho$.
+If, however, $|c_{v}-c_{0}|<\rho$ and if $C(c_{0},\rho)\cap C(c_{v},\rho)=\{\bar{z},\bar{z}\}$, then $|\bar{z}-\bar{z}^{\prime}|>|z-z^{\prime}|$, and the theorem follows in this case as well.
+∎
+Corollary 18.1.
+The disc $D(0,1)$ is $\theta$-isolated for $\theta$ of (18.3) provided that the discs $D(c_{j},\rho)$ contain no roots for $j=0,1,\dots,v$.
+Correctness of Alg.
+18.1.
+(i) If algorithm $\mathbb{ALG}~{}1$ outputs inclusion in at least one of its applications, then one of the discs $D(c_{j},\bar{\sigma}\rho)$ contains a root, and then Alg.
+18.1 correctly reports that $r_{d-m+1}\geq 1-\sigma\bar{\rho}$.
+(ii) Otherwise Cor. 18.1 holds, and then correctness of Alg. 18.1 follows from correctness of $\mathbb{ALG}~{}1$.
+Softness is at most $\frac{1}{1-\bar{\sigma}\rho}$, $\frac{1}{1-0.5\rho\sqrt{3}}$, and $1+(\sqrt{3}-1)\rho$ in cases (a), (b,i), and (b,ii), respectively.
+Computational complexity.
+Alg.
+18.1 invokes algorithm $\mathbb{ALG}~{}0$ at most $v+1$ times for $v=O(1)$ of (18.1).
+If $\mathbb{ALG}~{}0$ outputs exclusion $v+1$ times, then algorithm $\mathbb{ALG}~{}1$ is invoked and runs at a dominated cost.
+Remark 18.1.
+One can modify Alg.
+18.1 as follows (see [48]): instead of fixing a positive $\rho$ and then computing $v$ of (18.1) fix an integer $v>0$, apply algorithm $\mathbb{ALG}~{}0$ at $v$ equally spaced points $c_{j}=\exp(\frac{2j\pi\sqrt{-1}}{v}),~{}j=0,1,\dots,v-1$, let $\phi=\frac{2\pi}{v}$, and then define $\rho$.
+We also arrive at this algorithm if we compute or guess this value $\rho$ and then apply Alg.
+18.1 but notice that $c_{v}=c_{0}$ for $c_{j}$ of (18.2) and only invoke algorithm $\mathbb{ALG}~{}0$ for $c_{0},\dots,c_{v-1}$.
+19 Counting e/i tests
+19.1 An upper estimate
+The upper estimate $22m-$ 2 of part (ii) of Sec. 4 on the number of compression steps is sharp – reached in the case of the complete binary tree $\mathbb{T}$.
+Besides compression steps our subdivision root-finders involve e/i tests.
+Let us count them.
+Each compression step is preceded by an e/i test, and so we may need to perform at least $22m-$ 2 e/i tests.
+According to the main theorem of this section below, this lower bound is sharp up to a constant factor:
+Theorem 19.1.
+Let subdivision iterations with compression be applied to a suspect square containing at most $m$ roots. Then they involve $O(m)$ e/i tests overall.
+19.2 The first lemma
+We begin our proof of Thm. 19.1 with
+Lemma 19.1.
+In every component $C$ made up of $\sigma(C)$ suspect squares we can choose $n(C)\geq\sigma(C)/5$ non-adjacent suspect squares.
+Proof.
+[…]
+19.3 The second lemma
+Lemma 19.2.
+Fix two integers $i\geq 0$ and $g\geq 3$, fix a value $\sigma$ of softness, $1<\sigma<\sqrt{2}$, and define the separation fraction $a:=2-\sigma\sqrt{2}$ of Observation 17.1.
+For all $j$ let $j$-th subdivision step process $v_{j}$ components made up of $w_{j}$ suspect squares, $n_{j}$ of which are non-adjacent, let $\Delta_{j}$ denote their side length, and let the $j$-th steps perform no compression for $j=i,i+1,\dots,i+g$.
+Then
+$w_{i+g}\geq(n_{i}-v_{i})a2^{g-0.5}~{}{\rm if}~{}\Delta_{i}\geq 2^{g}\Delta_{i+g}.$ (19.1)
+Proof.
+At the $i$th step a pair of non-adjacent suspect squares of the same component is separated by at least the gap $a\Delta_{i}$, and there are at least $n_{i}-v_{i}$ such gaps with overall length $(n_{i}-v_{i})a\Delta_{i}$.
+Neither of these gaps disappears at the $(i+g)$th step, and so they must be covered by $w_{i+g}$ suspect squares, each of diameter $\sqrt{2}~{}\Delta_{i+g}=\Delta_{i}/2^{g-0.5}$ under (19.1).
+Hence $(n_{i}-v_{i})a\Delta_{i}\leq w_{i+g}\Delta_{i}/2^{g-0.5}$.
+∎
+Next we extend Lemma 19.2 by allowing compression steps.
+We only need to extend the equation $2\Delta_{j+1}=\Delta_{j}$ to the case where the $j$th subdivision step for $i\leq j\leq i+g$ performs non-final compression, each followed with subdivision steps applied to a single suspect square with side length at most $4\Delta_{j}$.
+The next two subdivision steps perform only five e/i tests, making up at most $105m-$ 10 e/i tests performed at all pairs of subdivision steps following all (at most $m-2$) non-final compression steps.
+Now we only need to bound by $O(m)$ the number of e/i tests in the remaining subdivision steps, not interrupted with compression steps.
+After a compression step we resume subdivision process with a suspect square whose side length bound decreases in the next two subdivision steps from at most $4\Delta_{j}$ to at most $\Delta_{j}$.
+Hence we can apply to it Lemma 19.2, because now it accesses the same number $n_{j}$ of non-adjacent suspect squares and an increased number $\sigma_{j}$ of all suspect squares (cf. (19.1)).
+19.4 Counting e/i tests at partition steps
+Let $h$ denote the height of our component tree and count suspect squares processed in subdivision steps that partition components, temporarily ignoring the other steps.
+Clearly, bounds (19.1) still hold and can only be strengthened. Notice that $v_{i}\geq 1$ for all $i$, sum $v_{i+g}$ as well as bounds (19.1) for a fixed $g>0$ and $i=0,1,\dots,h-g$, and obtain
+$\sum_{i=0}^{h-g}n_{i}-\lambda\sum_{i=0}^{h-g}\sigma_{i+g}\leq\sum_{i=0}^{h-g}v_{i+g}~{}{\rm for}~{}\lambda:=\frac{\sqrt{2}}{a2^{g}}.$ (19.2)
+Notice that
+$\sum_{i=0}^{h-g}v_{i+g}\leq\sum_{i=0}^{h}v_{i}-\sum_{i=0}^{g-1}v_{i}<2m-g,$
+Combine the above estimate and the bounds $n_{i}\leq m$ and (cf. Lemma 19.1) $w_{i}\leq 5n_{i},$ for all $i$, write $n^{\prime}:=\sum_{i=0}^{h}n_{i}$, obtain
+$\sum_{i=0}^{h-g}n_{i}\leq n^{\prime}-\sum_{i=h-g+1}^{h}n_{i}\leq n^{\prime}-mg,~{}\sum_{i=0}^{h-g}w_{i+g}\leq\sum_{i=0}^{h}w_{i}\leq 5n^{\prime},$
+and so
+[…]
+Obtain that
+$0.5n^{\prime}\leq 2m-1-g+mg~{}{\rm for}~{}g=1+\Big{\lceil}\log_{2}\Big{(}\frac{5\sqrt{2}}{a}\Big{)}\Big{\rceil}.$
+Hence
+$n^{\prime}\leq 4m-2-2g+2mg,~{}{\rm and~{}so}~{}n^{\prime}=O(m)~{}{\rm for}~{}g=O(1).$ (19.3)
+19.5 Counting all e/i steps
+Next consider a non-compact component, count all suspect squares processed in all subdivision steps applied until it is partitioned, and then prove in Lemma 19.3 that up to a constant factor the overall number of these suspect squares is bounded by their overall number at all partition steps.
+Such a result will extend to all subdivision steps the bound $n^{\prime}=O(m)$ of (19.3), so far restricted to partition steps.
+Together with Lemma 19.1 this will complete our proof of Thm.
+19.1.
+Lemma 19.3.
+Let $k+1$ successive subdivision steps with $\sigma$-soft e/i tests be applied to a non-compact component for $\sigma$ of our choice in the range $1<\sigma<\sqrt{2}$.
+Let the $j$-th step output $n_{j}$ non-adjacent suspect squares for $j=0,1,\dots,k$.
+Then $\sum_{j=0}^{k-1}n_{j}\leq n_{k}\frac{10\sqrt{2}}{a}$ for the separation fraction $a:=2-\sigma\sqrt{2}$ (cf.
+Observation 17.1.)
+Proof.
+Substitute $g=k-i$ into (19.3) and obtain $w_{k}\geq(n_{i}-v_{i})a2^{k-i-0.5}$, for $i=0,1,\dots,k-1$.
+Recall that $w_{j}\leq 5n_{j}$ (cf. Lemma 19.1) and that in our case $v_{i}=1$ and $n_{i}\geq 2$ and obtain that $n_{i}-v_{i}\geq n_{i}/2$ for all $i$.
+Hence $5n_{k}\geq n_{i}a2^{k-i-1.5}$ for all $i$, and so
+$\sum_{i=0}^{k-1}n_{i}\leq n_{k}\frac{10\sqrt{2}}{a}\sum_{j=1}^{k}2^{-j}<n_{k}\frac{10\sqrt{2}}{a}.$
+Remark 19.1.
+The overhead constants hidden in the "O" notation of our estimates are fairly large but overly pessimistic.
+E.g., (i) the roots $x$ and $y$ involved in the proof of Lemma 19.2 are linked by a path of suspect squares, but we deduce the estimates of the lemma by using only a part of this path, namely, the part that covers the gaps between non-adjacent suspect squares.
+(ii) Our proof of that lemma covers the worst case where a long chains between the roots $x$ and $y$ can consist of up to $4m$ suspect squares, but for typical inputs such chains are much shorter and can consist of just a few suspect squares.
+(iii) In the proof of Lemma 19.3 we assume that $n_{i}-1\geq n_{i}/2$; this holds for any $n_{i}\geq 2$, but the bounds are strengthened as $n_{i}$ increases.
+(iv) In a component made up of five suspect squares at least two are non-adjacent, and so the bound of Lemma 19.1 can be strengthened at least in this case.
+20 Overall complexity of our root-finders
+We reduced our root-finding in a square with $m$ roots in its close neighborhood to $O(m)$ e/i tests in addition to at most $22m-$ 2 compression steps.
+We reduced compression first of an isolated and then of strongly isolated disc to (i) counting roots in an isolated disc ($O(\log(d))$ evaluation queries) (ii) application of Schröder’s single iteration (two evaluation queries), and (iii) approximation of the root radius $r_{d-m+1}(z,p)$, that is, performing $O(\log(b))$ $m$-tests.
+In Sec.
+18 we reduced an $m$-test to $O(1)$ e/i tests, and so the overall complexity of subdivision iterations is dominated by the complexity of performing $O(m\log(b))$ e/i tests and $O(m\log(d))$ queries of an evaluation oracle.
+By combining the latter estimates with those of Sec.
+7 for an e/i test, we complete the proof of Main Theorems 1 – 4.
+21 Algorithmic extras
+Next we accelerate our root-finders for a large class of inputs.
+1.
+To compress the isolated unit disc $D:=D(0,1)$ containing a small number $m$ of roots $z_{1},\dots,z_{m}$, compute $s_{1},\dots,s_{m}$ the first $m$ powers sums of the roots lying in $D$, then recover the coefficients of the factor $f(x)=\prod_{j=1}^{m}(x-z_{j})$ from Newton’s identities (cf.
+[83, Eqn.
+(2.5.4)]), approximate its zeros, and apply a black box root-finder to the polynomial $p(x)/f(x)$ represented by the pair of polynomials $p(x)$ and $f(x)$.
+2.
+In the case of a large integer $m$ the latter algorithm suffers from swelling the coefficients of the factor $f(x)$.
+For $m=d$ the algorithm [102] empirically approximates all $m$ roots at a rather low cost, which provides more than just compression of $D$.
+Can we extend this approach to compression in the case where $m<d$ and the disc $D$ is sufficiently strongly isolated?
+This is a research challenge, but we propose a similar low cost heuristic recipe which, as we conjecture, accelerates our root-finding for a large class of inputs.
+Let the disc $D=D(0,1)$ be strongly isolated and let $i(D)\geq d^{2\nu}$ for a reasonably large $\nu$, and then instead of Newton’s apply Schröder’s single iteration (17.1) at sufficiently many, say 100, equally spaced points $y$ of the circle $C:=C(0,d^{\nu})$.
+Let $Z$ denote the set of images $z=y-m{\rm NR}(y)$.
+By extending our study in Sec.
+17.5, deduce that all these images lie in the $\alpha\Delta$-neighborhood of the root set $X(D)$ for $\Delta:=\Delta(D)$ and $\alpha$ fast converging to 1 as $\nu$ grows large and hence that $\Delta(Z)\leq(2\alpha+1)\Delta$.
+We may run into worst case inputs where $Z$ is a singleton and $\Delta(Z)=0$, but we conjecture that under a random rotation of the set of points $y$, the ratio $\Delta(Z)/\Delta$ exceeds a positive constant whp, and then we can narrow the initial range of BoL in Sec.
+16 from $[-b,0]$ to $[\log_{2}(\Delta(Z)),0]$ at a low cost of performing Schröder’s iteration (17.1) at reasonably many, say 100, points.
+Part V: Factorization, Root Isolation, Functional Iterations, and Root Radii Estimation
+We organize this part of the paper as follows.
+In Secs.
+22 and 23 we cover some results by Schönhage [106] on polynomial factorization and its links to root-finding and root-isolation, and our lower and upper bit-complexity estimates for these problems.
+In Sec.
+24 we recall root-finders by means of Newton’s, Weierstrass’s [119] (aka Durand-Kerner’s [23, 51]), and Ehrlich’s of 1967 [28], aka Aberth’s of 1973 [2] iterations, accelerate them by means of incorporation of FMM, and initialize Newton’s iterations for a black box input polynomial.
+In Sec.
+25 we recall fast approximation of all $d$ root radii where the coefficients of $p$ are given.
+22 Polynomial Factorization, Root-finding and Root Isolation:
+Linking three Problems and their bit operation complexity
+22.1 Linking Factorization and Root-finding: upper bounds on the bit operation complexity
 Hereafter write $|u|=\sum_{i=0}^{d}|u_{i}|$ for a polynomial $u(x)=\sum_{i=0}^{d}u_{i}x^{i}$.
-The root-finders of [134, 96, 103] solve Problems 0 and 0 by extending the the solution of the following problem of independent importance:
+The root-finders of [106, 78, 84] solve Problems 0 and 0 by extending the solution of the following problem of independent importance:
 Problem F.
-Approximate Polynomial Factorization: [^35] The solution of Problem F can be also extended to isolation of the zeros of a square-free polynomial with integer coefficients and approximation of matrix eigenvalues (see our Remark 25.4 and Sec.
-11) and has various other applications to modern computations, e.g., to time series analysis, Wiener filtering, noise variance estimation, co-variance matrix computation, and the study of multichannel systems [150, 18, 8, 35, 36, 147].
+Approximate Polynomial Factorization: [^28] The solution of Problem F can be also extended to isolation of the zeros of a square-free polynomial with integer coefficients (see our Remark 23.4) and has various other applications to modern computations, listed in Sec.
+1.1.
 For a polynomial of degree $d$ the known sharp upper and lower bounds on the input precision and number of bit operations for root-finding exceed those for numerical factorization by a factor of $d$.
@@
 The zero set of $p$ consists of clusters, each made up of a constant number of zeros and having a diameter at most $1/2^{3-b/d}$; furthermore, these clusters are pairwise separated by the distances at least $1/d^{\tilde{O}(1)}$.
-[96, 97, 99, 100, 103] proved the following near optimal bound for Problem F and readily extended it to near optimal solution of Problem 0.
+In [78, 84] we proved the following near-optimal bound for Problem F.
 Theorem 22.1.
 $\mathbb{B}_{F}(b,d)=\tilde{O}((b+d)d)$.
-In Sec. 24.3 we reduce Problem 0 to Problem F and vice versa by applying Schönhage’s results of [134, Cor. 4.3 and Sec. 19]:
+In Sec. 22.3 we reduce Problem 0 to Problem F and vice versa by applying Schönhage’s results of [106, Cor. 4.3 and Sec. 19]:
 Theorem 22.2.
 (i) $\mathbb{B}_{0}(b,d)=\tilde{O}(\mathbb{B}_{F}(bd,d))$ for any polynomial $p(x)$ of degree $d$, (ii) $\mathbb{B}_{F}(b,d)=\tilde{O}(\mathbb{B}_{0}(b+d+\log(d),d))$ for any polynomial $p(x)$ of degree $d$, and (iii) $\mathbb{B}_{0}(b,d)=\tilde{O}(\mathbb{B}_{F}2(b,d))$ for polynomials of Def.
-24.1, with bounded cluster sizes.
-By combining Thms. 24.1 and 24.2 obtain the following estimates.
+22.1, with bounded cluster sizes.
+By combining Thms. 22.1 and 22.2 obtain the following estimates.
 Theorem 22.3.
@@
 Corollary 22.1.
-One can solve Problem 0 by using $\tilde{O}((b+d)d+(b+m)m^{2})$ bit operation for $m$ of Def. 1.1 of Sec. 1.5. The bound turns into $\tilde{O}((b+d)d$ if $m^{2}=\tilde{O}(d)$.
+One can solve Problem 0 by using $\tilde{O}((b+d)d+(b+m)m^{2})$ bit operation for $m$ of Def. 1.1 of Sec. 1.5. The bound turns into $\tilde{O}((b+d)d)$ if $m^{2}=\tilde{O}(d)$.
 Proof.
@@
 22.3 to the $m$th degree factor of the polynomial $\prod_{j=1}^{d}(u_{j}x-v_{j})$ of (22.1) that shares with this polynomial all its zeros lying in the disc $D(1+\beta)$ for positive $\beta$ of Def.
-1.1.363636We can slightly speed up stage (i) because at that stage we actually only need to approximate factorization of the factor $f$ of degree $m$ of $p$ that shares with $p$ its zeros in $D$.
-24.2 Lower bounds on the bit operation complexity
+1.1.292929We can slightly speed up stage (i) because at that stage we actually only need to approximate factorization of the factor $f$ of degree $m$ of $p$ that shares with $p$ its zeros in $D$.
+22.2 Linking Factorization and Root-finding: lower bounds on the bit operation complexity
 Readily verify the following information lower bounds.
-Observation 24.1.
+Observation 22.1.
 To solve Problem 0 or F one must access $(d+1)b$ bits of input coefficients and perform at least $0.5(d+1)b$ bit operations – at most one operation for two input bits.
-Next strengthen the lower bound for Problem 0.
-Observation 24.2.
-Let $p(x)=(x-v_{1})^{j}f(x)$ for an integer $j>0$ and a monic polynomial $f(x)$ of degree $d-j$.
-Fix a real $b$.
-Then the polynomial $p_{j}(x)=p(x)-2^{-jb}f(x)$ has the $j$ zeros $v_{1}+\zeta^{g}2^{-b}$ for $g=0,1,\dots,j-1$.
-Observation 24.2 implies that one must access the coefficient $p_{d-j}$ within $2^{-jb}$ to approximate within $2^{-b}$ a $j$-multiple zero of $p$ or its zero in a cluster of $j$ zeros, and so a root-finder must process at least $jb$ bits of the coefficient $p_{d-j}$.
+Next we deduce a stronger lower bound for Problem 0.
+Observation 22.2.
+Let $p(x)=x^{d}+p_{d-j}x^{d-j}$ for an integer $j$, $1<j<d$, and complex $p_{d-j}$ such that $|p_{d-j}|=1/2^{bj}$.
+Then the polynomial $p(x)$ has the $j$ zeros $\zeta^{g}2^{-b}$ for $g=0,1,\dots,j-1$, which vanish in the transition to $x^{d}$.
+Observation 22.2 implies that one must access the coefficient $p_{d-j}$ within $2^{-jb}$ to distinguish the above $j$ zeros of $p(x)$ from the value 0 or from the zeros of the cluster of $j$ zeros of the polynomials obtained from $x^{d}$ by means of infinitesimal perturbation of the coefficient $p_{d-j}$.
 Sum these bounds for $j=1,\dots,d$ and obtain
 Corollary 22.2.
-One must access at least $0.5(d+1)db$ bits of the coefficients of a monic polynomial $p$ and perform at least $0.25(d+1)db$ bit operations to approximate even its single within $2^{-b}$.
+One must access at least $0.5(d+1)db$ bits of the coefficients of a monic polynomial $p$ and perform at least $0.25(d+1)db$ bit operations to approximate even its single zero within $2^{-b}$ if no restriction is imposed on the number of roots in an input complex domain $\mathbb{D}$ and its $\beta$-neighborhood.
 The lower bounds of this subsection imply that the algorithm of [78, 84] and its extension supporting Thm.
-24.1 solve Problems 0 and F of large size by using optimal number of bit operations up to polylogarithmic factors.
-24.3 Linking Factorization and Root-finding (with more details)
-Next we prove Thm. 24.2, by applying the results of [134, Cor. 4.3 and Sec. 19].
+22.1 solve Problems 0 for $m=d$ and F by using optimal number of bit operations up to polylogarithmic factors.
+Let us extend this study to Problem 0 for $m<d$. In this case isolate from one another the two sets made up of $m$ and $d-m$ roots.
+Observation 22.3.
+Let $p(x)=(x^{m}+p_{d-j}x^{m-j})(x-z_{1})^{d-m}$ for an integer $j$, $1<j<m$, complex $p_{d-j}$ such that $|p_{d-j}|=1/2^{bj}$, and sufficiently large $|z_{1}|$.
+Then the polynomial $p(x)$ has the $j$ zeros $\zeta^{g}2^{-b}$ for $g=0,1,\dots,j-1$, which vanish in the transition to $p(x)=x^{m}(x-z_{1})^{d-m}$.
+Now we extend Cor. 22.2 as follows.
+Corollary 22.3.
+For an integer $m$, $1\leq m\leq d$, one must access at least $0.5(m+1)mb+db$ bits of the coefficients of a monic polynomial $p$ and perform at least $0.25(m+1)mb+0.5db$ bit operations to approximate within $2^{-b}$ even its single zero lying in the $\beta$-neighborhood of a complex domain $\mathbb{D}$ for a fixed constant $\beta>0$ if this neighborhood contains at most $m$ roots.
+Proof.
+Obtain the lower bounds $0.5(m+1)mb$ and $0.25(m+1)mb$ as for Cor.
+22.2.
+Then notice that a $\delta$-perturbation of any trailing coefficient $p_{i}$ of $p(x)$ for $i\leq d-m$ can move the roots lying in the $\beta$-neighborhood of $\mathbb{D}$ by order of $\delta$.
+∎
+Our Main Theorem 3 implies that these lower bounds are near-optimal for $m$ of order $d$.
+22.3 Linking Factorization and Root-finding: proof of Thm. 22.2
+Next we prove Thm. 22.2, by applying the results of [106, Cor. 4.3 and Sec. 19].
 Recall that $\mathbb{B}_{0}(b,d)$ and $\mathbb{B}_{F}(b,d)$ denote the optimal number of bit operations for the solution of Problems 0 and F, respectively, for an error bound $\epsilon=1/2^{b}$, and furthermore, recall the lower bounds
 $\mathbb{B}_{0}(b,d)\geq 0.5(d+1)b~{}{\rm and}~{}\mathbb{B}_{F}(b,d)\geq 0.25(d+1)db$ (22.2)
-of [96, 103] and the following result (cf. [134, Thm. 19.1]).
+of [78, 84] and the following result (cf. [106, Thm. 19.1]).
 Theorem 22.4.
@@
 Then a solution of Problem 0 within an error bound $1/2^{b}$ can be recovered at a dominated bit operation cost from the solution of Problem F within an error bound $1/2^{bd+2}$, and so $\mathbb{B}_{0}(b,d)\leq\mathbb{B}_{F}(bd+2,d)$.
-Based on [134, Cor. 4.3] we can also readily relax the assumption that $p(x)$ is monic.
+Based on [106, Cor. 4.3] we can also readily relax the assumption that $p(x)$ is monic.
 Furthermore, in view of (22.2), relaxing the bound $r_{1}\leq 1$ would little affect the latter estimate because we can apply map (1.4), which changes the error bound of Problem 0 by at most $\log_{2}(\max\{\rho,1/\rho\})$, for $\rho\geq 1/2^{b}$.
-[136, Thm. 2.7], which we reproduce in Sec. 25.4 as Thm. 25.3, shows an alternative way to relaxing both assumptions – that $r_{1}\leq 1$ and $p(x)$ is monic.
-Schönhage in [134, Sec. 19] has greatly strengthened the estimate of Thm. 24.4 unless “the zeros of $p$ are clustered too much". Namely, he proved (see [134, Eqn. (19.8)]) the following result.
+[108, Thm. 2.7], which we reproduce in Sec. 23.4 as Thm. 23.3, shows an alternative way to relaxing both assumptions that $r_{1}\leq 1$ and $p(x)$ is monic.
+Schönhage in [106, Sec. 19] has greatly strengthened the estimate of Thm. 22.4 unless “the zeros of $p$ are clustered too much". Namely, he proved (see [106, Eqn. (19.8)]) the following result.
 Lemma 22.1.
@@
 If $1/|\Delta-r|^{\frac{d-m}{m}}=2^{\tilde{O}(d)}$, then Lemma 22.1 implies that $|z_{1}-v_{1}|\leq 1/2^{b/m-\tilde{O}(d)}$, and we can similarly estimate $|v_{j}-z_{j}|$ for all $j=1,2,\dots,d$.
-By extending these estimates we strengthen Thm. 24.4 as follows.
+By extending these estimates we strengthen Thm. 22.4 as follows.
 Corollary 22.4.
 Under bounding cluster sizes of Def.
-24.1, it holds that $|z_{j}-v_{j}|\leq 1/2^{b/m-\tilde{O}(d)},~{}j=1,2,\dots,d$.
-Hence if $m=O(1)$ is a constant, then a solution of Problem 0 within an error bound $1/2^{b}$ can be recovered at a dominated bit operation cost from the solution of Problem F with a precision $1/2^{\bar{b}}$ for a sufficiently large $\bar{b}=\tilde{O}(b+d)$, and so $\mathbb{B}_{1}(b,d)\leq\mathbb{B}_{2}(\tilde{O}(b+d),d)$.
-For converse reduction of Problem F to Problem 0 recall [134, Cor. 4.3] in the case of $k=2$, covering the product of two polynomials:
+22.1, it holds that $|z_{j}-v_{j}|\leq 1/2^{b/m-\tilde{O}(d)},~{}j=1,2,\dots,d$.
+Hence if $m=O(1)$ is a constant, then a solution of Problem 0 within an error bound $1/2^{b}$ can be recovered at a dominated bit operation cost from the solution of Problem F with a precision $1/2^{\bar{b}}$ for a sufficiently large $\bar{b}=\tilde{O}(b+d)$, and so $\mathbb{B}_{0}(b,d)\leq\mathbb{B}_{F}(\tilde{O}(b+d),d)$.
+For converse reduction of Problem F to Problem 0 recall [106, Cor. 4.3] in the case of $k=2$, covering the product of two polynomials:
 Theorem 22.5.
@@
 Therefore, $||h||_{1}\leq 2^{d-1}||p||_{1}/||x-z_{j}||_{1}\leq 2^{d-1}||p||_{1}$ by virtue of Thm.
-24.5.
+22.5.
 Hence the approximation errors of all the $d$ roots together contribute at most $2^{d-1}d\epsilon$ to the relative error of the solution of Problem F (ignoring dominated impact of higher order terms), and we arrive at
 Corollary 22.5.
-If $b=O(b-d-\log(d))$, then Problem F for an error bound $1/2^{b}$ can be reduced at a dominated bit operation cost to Problem 0 for an error bound $1/2^{\widehat{b}}$ and a sufficiently large $\widehat{b}=O(b+d+\log(d))$, and so $\mathbb{B}_{F}(b,d)=O(\mathbb{B}_{0}(b,d))$.
+If $b=O(b-d-\log(d))$, then Problem F for an error bound $1/2^{b}$ can be reduced at a dominated bit operation cost to Problem 0 for an error bound $1/2^{\widehat{b}}$ and a sufficiently large $\widehat{b}=O(b+d+\log(d))$; hence and so $\mathbb{B}_{F}(b,d)=O(\mathbb{B}_{0}(b,d))$.
 It follows that the complexity estimate of Thm.
@@
 Then again, the bit operation complexity of this solution of Problem F is optimal (up to polylogarithmic factors) under the random root model provided that $d=\tilde{O}(b)$ (cf.
-Observation 24.1).
-24.4 Two root-finders of 2016 and 2018
-Becker et al in [23, 24] claimed that their extension of subdivision root-finders of [128, 101] (i) runs within the same near optimal bit operation bounds of [96, 103] but (ii) unlike [96, 103] is easy to implement.
-This seemed to diminish incentive for further research in this area, but were the claims valid?
-Implementation of the root-finder of [23] by Imbach et al in [64] provided no empirical support to optimality claimed by Becker et al because the root-finder of [63], using some novel techniques from [112] (see [63, Sec.
-1.2]), runs much faster.
-Unfortunately, formal support for optimality claimed by Becker et al has been missing as well.
-Indeed, for all proofs the paper [23] refers to a preprint of [24], with a distinct list of the authors and with the main algorithms and the Main Theorem significantly different from those of [23].
-Furthermore, as we pointed out in the introduction, the the bit-complexity specified in the Main Theorem of [24] has order of $mbd^{2}$, that is, exceeds the near optimal bound of [96, 103] by a factor of $d$.
-Can one salvage the claim by Becker et al in another way?
-Next we argue that this can be problematic because the algorithms of [23, 24] for Problem 1m are some variants of subdivision iterations.
-The basic tool and the main algorithmic novelty of the variant of subdivision in [96, 103] are e/i tests based on the following theorem by Pellet [88], which extended the previous results by Cauchy and Rouché:
-Theorem 24.6.
-Given a polynomial $(x)$ of (1.1) with complex coefficients and $p_{g}=1$.
-Let the polynomial $s(x)=x^{d}+p_{d-1}x^{d-1}+\cdots+|p_{i+1}|x^{i+1}-|p_{i}|x^{i}|+p_{i-1}|x^{i-1}+\cdots+|p_{0}|$ have two distinct positive roots $r$ and $R$, $r<R$.
-Then the polynomial $p$ has exactly $i$ zeros in or on the circle $C(0,r)=\{z:~{}|z|=r\}$ and no zeros in the annulus $A(0,r,R)=\{z:~{}r<|z|<R\}$.
-[23, 24] apply Pellet’s theorem on each subdivision step to more than $4m$ polynomials $t(x)=\sum_{i=0}^{d}t_{i}x^{i}$ obtained from the polynomial $p$ by means of shifting the variable $x$, its scaling by up to $2^{b}$, and root-squaring.
-To apply Pellet’s theorem to $t(x)$ one must compute the values $|\frac{t_{i}}{t_{d}}|$ for $i=0,1,\dots,d-1$.
-Hence, by virtue of Observation 24.2, one must process at least $0.5(d+1)db$ bits of the coefficients of $t(x)$ to support approximation of its zeros within $1/2^{b}$.
-The algorithms of [24] cannot avoid this because they apply e/i tests to the polynomials obtained in $\log(\log(d))$ steps of root-squaring (cf.
-Sec.
-13.3) from polynomials $t_{c,\rho}(x)=p(\frac{x-c}{\rho})$ where $\rho>0$ can be as small as $1/2^{b}$.
-Already a single subdivision iteration can perform more than $4m$ soft e/i tests; this means that one must process more than $2(d+1)dmb$ bits of memory and $(d+1)dmb$ bit-operations to solve Problem 1m by applying such a Pellet-based root-finder.
-25 Recursive Factorization; extensions to Root-finding and Root Isolation
+Observation 22.1).
+23 Recursive Factorization, its bit complexity, and Root Isolation
 In this section we briefly recall some algorithms and complexity estimates from [106, 108].
@@
 Remark 23.1.
-[136, Lemma 2.6] states with no proof a little stronger bound $\prod_{j=1}^{k}|f_{j}|\leq 2^{d-1}|p|$ under the assumptions of Lemma 25.3.
-From various factors of the polynomial $p(x)=x^{d}-1$ such as $\prod_{j=1}^{\frac{d}{2}}(x-\zeta_{d}^{j})$ for even $d$ and $\zeta:=\zeta_{q}:=\exp(\frac{2\pi{\bf i}}{q})$ denoting a primitive $q$-th root of unity, one can see some limitations on strengthening this bound further.
+[108, Lemma 2.6] states with no proof a little stronger bound $\prod_{j=1}^{k}|f_{j}|\leq 2^{d-1}|p|$ under the assumptions of Lemma 23.3.
+From various factors of the polynomial $p(x)=x^{d}-1$ such as $\prod_{j=1}^{\frac{d}{2}}(x-\zeta_{d}^{j})$ for even $d$ and $\zeta_{q}$ of (3.4), one can see some limitations on strengthening this bound.
 23.2 Errors and complexity
@@
 $p\approx p^{\ast}=p^{\ast}(x)=\prod_{j=1}^{d}(u_{j}x+v_{j}).$ (23.1)
-Next, by following [134, Sec. 5], we estimate the norm of the residual polynomial
+Next, by following [106, Sec. 5], we estimate the norm of the residual polynomial
 $\Delta^{\ast}=p^{\ast}-p.$ (23.2)
@@
 $|(f_{1}-fg)f_{2}\cdots f_{k}|\leq|f_{1}-fg|~{}|f_{2}\cdots f_{k}|=\Delta|f_{2}\cdots f_{k}|.$
-Combine the latter inequalities and obtain $\Delta_{k+1}\leq\Delta_{k}+\Delta|f_{2}\cdots f_{k}|$. Combine this bound with (25.3)-(25.5) and Lemmas 25.1 and 25.3 and obtain (25.6). ∎
+Combine the latter inequalities and obtain $\Delta_{k+1}\leq\Delta_{k}+\Delta|f_{2}\cdots f_{k}|$. Combine this bound with (23.3)-(23.5) and Lemmas 23.1 and 23.3 and obtain (23.6). ∎
 Write $f_{1}:=f$ and $f_{k+1}=g.$ Then (23.6) turns into (23.3) for $k$ replaced by $k+1.$ Now compute one of the factors $f_{j}$ as in (23.4), apply Thm.
-25.1, then recursively continue splitting the polynomial $p$ into factors of smaller degrees, and finally arrive at factorization (25.1) with
+23.1, then recursively continue splitting the polynomial $p$ into factors of smaller degrees, and finally arrive at factorization (23.1) with
 $|\Delta^{\ast}|\leq\epsilon|p|$
@@
 $\prod_{i=1}^{k}|f_{i}|\leq 2^{d}|\prod_{i=1}^{k}f_{i}|\leq 2^{d}~{}\Big{(}1+\frac{k\epsilon}{d}\Big{)}~{}|p|.$
-The latter bound cannot exceed $2^{d+1}|{p}|$ for $k\leq d$, $\epsilon\leq 1$. Consequently (25.7) ensures (25.5), and then (25.6) follows by virtue of Thm. 25.1.
+The latter bound cannot exceed $2^{d+1}|{p}|$ for $k\leq d$, $\epsilon\leq 1$. Consequently (23.7) ensures (23.5), and then (23.6) follows by virtue of Thm. 23.1.
 Remark 23.2.
@@
 bits throughout Recursive Splitting Process.
-25.3 Overall complexity of Recursive Factorization and extension to Root Isolation
+23.3 Overall complexity of Recursive Factorization
 The overall complexity of recursive factorization is dominated by the sum of the bounds on the complexity of all splittings into pairs of factors.
@@
 This is due to additional techniques, which ensure that in every recursive splitting of a polynomial into the product of two factors the ratio of their degrees is at most 11.
-Consequently, $O(\log(d)$ recursive splitting steps are sufficient in [103], and so the overall complexity of root-finding is proportional to that of the first splitting.
-In contrast, the degrees of a factor in any splitting
-in [134] can have low, e.g., 1, and so there can be allows up to $d-1$ splittings overall, with each of the first $0.5d$ of them almost as expensive as the very first one.
+Consequently, $O(\log(d)$ recursive splitting steps are sufficient in [84], and so the overall complexity of root-finding is in $\tilde{O}(db)$ – proportional to that of the first splitting.
+In contrast, the degrees of a factor in any splitting in [106] can be low, e.g., 1, and so there can be up to $d-1$ splittings overall, each of the first $0.5d$ of them being almost as expensive as the very first one.
 23.4 From factors to roots
 Theorem 23.3.
-[136, Thm. 2.7]. Suppose that
+[108, Thm. 2.7]. Suppose that
 $p=p_{d}\prod_{j=1}^{d}(x-z_{j}),~{}\tilde{p}=\tilde{p}_{d}\prod_{j=1}^{d}(x-y_{j}),~{}|\tilde{p}-p|\leq\epsilon|p|,~{}\epsilon\leq 2^{-7d}$
@@
 23.5 Polynomial Root Isolation
-Based on his results recalled in this section, Schönhage in [134, Sec. 20] estimates the bit operation complexity of the following well-known problem by reducing it to Problem F.
-Problem ISOL.
-Polynomial Root Isolation.
+Based on his results recalled in this section, Schönhage in [106, Sec. 20] estimates the bit operation complexity of the following well-known problem by reducing it to Problem F.
+Problem ISOL: Polynomial Root Isolation.
 Given a polynomial $p$ of (1.1) that has integer coefficients and is square-free, that is, has only simple roots, compute $d$ disjoint discs on the complex plane, each containing exactly one root.
 Remark 23.4.
-Our Problem 0 can be viewed as the Problem of Polynomial Root Cluster Isolation.
+Problem 0 can be viewed as the Problem of Polynomial Root Cluster Isolation.
 It turns into Problem ISOL if all the coefficients of $p$ are integers and the roots are simple, and so our algorithms for Problem 0 as well as the Algorithms of [78, 84] solve Problem ISOL within a record complexity bounds.
-In [134, Sec.
-20] Schönhage reduces to factorization Problem F the root-isolation problem of computing $d$ disjoint discs each covering a single simple root of a square-free polynomial with integer coefficients.
-Problem ISOL is readily solved if the zeros are approximated within $\epsilon=1/2^{b}<0.5\delta$ for $\delta$ denoting the minimal pairwise distance between the zeros of $p$ (see the known lower bounds on $\delta$ in Kurt Mahler and Maurice Mignotte [77, 78, 79]), and Schönhage has estimated that solution of Problem F (e.g., by means of the algorithms of [134, 94, 103]) implies isolation provided that
+In [106, Sec.
+20] Schönhage reduces Problem ISOL to factorization Problem F.
+Problem ISOL is readily solved if the zeros are approximated within $\epsilon=1/2^{b}<0.5\delta$ for $\delta$ denoting the minimal pairwise distance between the zeros of $p$ (see the known lower bounds on $\delta$ in Kurt Mahler and Maurice Mignotte [60, 61, 62]).
+Schönhage has estimated that the solution of Problem F (e.g., presented in [106, 52, 76, 77, 84]) implies isolation provided that
 $b=O((\tau+\log(d))+d)~{}{\rm and}~{}\tau=\log_{2}\max_{i=0}^{d}|p_{i}|.$ (23.8)
-This implied the following result.
 Corollary 23.2.
@@
 24.1 Contents
-In the next subsection we recall Newton’s, Weierstrass’s, and Ehrlich’s iterations, comment on their convergence, and acceleration by means of applying fast multipoint evaluation, which is dramatic but involves the coefficients of $p$.
+In the next subsection we recall Newton’s, Weierstrass’s, and Ehrlich’s iterations and comment on their convergence and acceleration by means of fast multipoint evaluation, which is dramatic but involves the coefficients of $p$.
 In Sec.
-26.3 we cover mre limited acceleration of these iterations by means of incorporating FMM, which can be applied to a black box polynomial.
+24.3 we cover more limited acceleration by means of incorporation of FMM, which can be applied to a black box polynomial.
 In Secs.
-26.5 and 27) we cover initialization of Newton’s, Weierstrass’s, and Ehrlich’s iterations, required for their fast global empirical convergence.
+24.5 and 25 we cover initialization of Newton’s, Weierstrass’s, and Ehrlich’s iterations, required for their fast global empirical convergence.
 For Newton’s iterations we only need to compute a disc that contains all zeros of a polynomial $p$ (cf.
 [55, 43, 115, 102, 113]).
-This is given by the well-known estimates via the coefficients, but we present simple but novel recipes in the case of black box polynomial $p$.
+This is given by the well-known estimates via the coefficients, but we present simple novel initialization recipes in the case of a black box polynomial $p$.
 The customary initialization is more demanding for Weierstrass’s and Ehrlich’s iterations but is still fast for polynomials represented with their coefficients, as we recall in Sec.
-27), where we also point out how this initialization can be extended to fast real and nearly real root-finding.
-It is still a challenge to extend such an initialization to the case of a black box polynomial.
+25, where we also discuss applications of this initialization in particular to fast real and nearly real root-finding.
+It is still a challenge to extend such initialization to the case of a black box polynomial that would be faster than our radii algorithm of Sec.
+16.
 24.2 Definitions
-1. Definitions. Consider Newton’s, Weierstrass’s (aka Durand-Kerner’s), and Ehrlich’s (aka Aberth’s) iterations [71, 58, 130, 143, 148, 31, 67, 38, 3, 22]):
+Consider Newton’s, Weierstrass’s, and Ehrlich’s iterations [55, 43, 102, 115, 119, 23, 51, 28, 2, 15]):
 $z\leftarrow z-N(z),~{}z_{j}\leftarrow z_{j}-W(z_{j}),~{}z_{j}\leftarrow z_{j}-E_{j}(z_{j}),~{}j=1,\dots,d,~{}{\rm where}$ (24.1)
@@
 Newton’s iterations can be initiated at a single point $z$ or at many points and can approximate a single root, many or all $d$ roots at the cost decreasing as the number of roots decreases [55, 43, 115, 113, 68, 102].
-In contrast, Weierstrass’s and Ehrlich’s iterations begin at $d$ points for a $d$th degree polynomial and approximate a subset of its $d$ roots about as fast and almost as slow as they approximate all its $d$ roots.
-Near the roots Newton’s and Weierstrass’s iterations converge with quadratic rate, Ehrlich’s iterations with cubic rate [145], but empirically all of them consistently converge very fast globally, right from the start, assuming a proper initialization, which involves the coefficients of $p$ (see [15, 22]).
-[^37] Formal support for such global convergence is much weaker for Newton’s iterations (see [71, 58, 143], and the references therein) and is completely missing for Weierstrass’s and Ehrlich’s iterations.
-All expressions in (26.3) for the functions $N(x)$, $W(x)$, and $E_{j}(x)$ are already presented or can be rewritten as ratios $\frac{u(x)}{v(x)}$ for pairs of polynomials $u(x)$ and $v(x)$, whose coefficients are given or readily available at low cost if the polynomial $p(x)$ is gien with its coeffcients (see [102, Problems 3.6.1 and 3.6.3]).
-Then multipoint evaluation of such a function amounts to multipoint evaluation of $u(x)$ and $v(x)$ and division of the computed values, and one can accelerate Newton’s, Weierstrass’s, and Ehrlich’s iterations dramatically by incorporating the algorithms of [81, 59].
+Weierstrass’s and Ehrlich’s iterations begin at $d$ points for a $d$th degree polynomial and approximate a subset of its $d$ roots about as fast and almost as slow as all its $d$ roots.
+Near the roots Newton’s and Weierstrass’s iterations converge with quadratic rate, Ehrlich’s iterations with cubic rate [116].
+Empirically, however, all of them consistently converge very fast globally, right from the start, if they are initialized properly.
+[^30] Formal support for fast global convergence is only known for for Newton’s iterations empirically they still run faster than according to the known estimates (see [102]).
+The known fast initialization algorithms involve the coefficients of $p$ (see [9, 15]).
+All expressions in (24.3) for the functions $N(x)$, $W(x)$, and $E_{j}(x)$ can be rewritten as ratios $\frac{u(x)}{v(x)}$ for pairs of polynomials $u(x)$ and $v(x)$, whose coefficients are given or readily available at a low computational cost if the polynomial $p(x)$ is given with its coefficients (see [83, Problems 3.6.1 and 3.6.3]).
+Then multipoint evaluation of such a ratio amounts to multipoint evaluation of $u(x)$ and $v(x)$ and division of the computed values, and one can accelerate Newton’s, Weierstrass’s, and Ehrlich’s iterations by a factor $d/\log^{2}(d)$ by means of incorporation of the algorithms of [64, 44].
 24.3 Acceleration with Fast Multipole Method (FMM)
@@
 Next we cover acceleration of Newton’s, Weierstrass’s, and Ehrlich’s iterations applied to black box polynomial root-finding.
-Given $3d$ complex values $w_{j}$, $z_{i}$, $v_{j}$ such that $z_{i}\neq v_{j}$ for all $i,j=1,\dots,d$, Fast Multipole Method (FMM) by Greengard and Rokhlin [50] and by Carrier, Greengard, and Rokhlin [28] computes the sums $\sum_{j=1}^{d}\frac{w_{j}}{x_{i}-z_{j}}$ as well as $\sum_{j=1,j\neq i}^{d}\frac{w_{j}}{z_{i}-z_{j}}$ for all $i$ by using $O(db)$ arithmetic operations, with the precision of order $b$, versus $2d^{2}$ arithmetic operations in the straightforward algorithm.
-The paper [121] proposes and extensively tests its acceleration to Superfast Multipole Method.
+Given $3d$ complex values $w_{j}$, $z_{i}$, $v_{j}$ such that $z_{i}\neq v_{j}$ for all $i,j=1,\dots,d$, FMM computes the sums $\sum_{j=1}^{d}\frac{w_{j}}{x_{i}-z_{j}}$ as well as $\sum_{j=1,j\neq i}^{d}\frac{w_{j}}{z_{i}-z_{j}}$ for all $i$ by using $O(db)$ arithmetic operations, with the precision of order $b$, versus $2d^{2}$ arithmetic operations in the straightforward algorithm.
 The papers [86, 88] list the recent bibliography on FMM and related computations with Hierarchical Semiseparable (HSS) matrices.
@@
 The acceleration based on FMM is practically valuable in spite of the limitation of the working and output precision to the order of $\log(d)$ and the resulting limitation on the decrease of the bit operation complexity.
-Next we incorporate FMM into Newton’s, Weierstrass’s, and Ehrlich’s iterations. First relate the roots of a polynomial $p(x)$ to the eigenvalues of a diagonal-plus-rank-one matrix
+Next we incorporate FMM into Newton’s, Weierstrass’s, and Ehrlich’s iterations.
+First relate the roots of a polynomial $p(x)$ to the eigenvalues of a diagonal-plus-rank-one matrix [^31] Hereafter T denotes transposition of a vector.
 $A=D+{\bf v}{\bf e}^{T}~{}{\rm for}~{}D=\operatorname{diag}(z_{j})_{j=1}^{d},~{}{\bf e}=(1)_{j=1}^{d},~{}~{}{\bf w}=(w_{j})_{j=1}^{d},~{}w_{j}=W(z_{j}),$
@@
 Hence, given the $2d$ values of $z_{j}$ and $w_{j}=W(z_{j})$ for all $j$, one can fast approximate $N(z_{j})$ and consequently $E_{j}(z_{j})$ for all $j$ (cf. (24.3)) by applying FMM.
-Every iteration but the first one begins with some values $z_{j}$, $w_{j}$, and updated values $z_{j}^{\rm(new)}$ of (26.1), for $j=1,\dots,d$, and outputs updated values $w_{j}^{\rm(new)}=W(z_{j}^{\rm(new)})$ for all $j$.
+Every iteration but the first one begins with some values $z_{j}$, $w_{j}$ and the updated values $z_{j}^{\rm(new)}$ of (24.1), for $j=1,\dots,d$, and outputs the updated values $w_{j}^{\rm(new)}=W(z_{j}^{\rm(new)})$ for all $j$.
 Then equations (24.3) and (24.5) define $N(z_{j}^{\rm(new)})$ and $E_{j}(z_{j}^{\rm(new)})$ for all $j$.
-To update the values $w_{j}$, apply [22, equation (17)]:
+To update the values $w_{j}$, apply [15, equation (17)]:
 $w_{i}^{\rm new}=(z_{i}^{\rm new}-z_{i})\prod_{j\neq i}\frac{z_{i}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}\Big{(}\sum_{k=1}^{d}\frac{w_{k}}{z_{i}^{\rm new}-z_{k}}-1\Big{)},~{}i=1,\dots,d.$
@@
 fast in two ways.
-(i) Use the following customary trick (see [32, Remark 3]): approximate the values $\ln(R_{i})$ for $i=1,\dots,d$ fast by applying FMM and then readily approximate $R_{i}=\exp(\ln(R_{i}))$ for all $i$.
-[32] proves numerical stability of the reduction to FMM of a similar albeit a little more involved rational expressions.
+(i) Use the following customary trick (see [24, Remark 3]): approximate the values $\ln(R_{i})$ for $i=1,\dots,d$ fast by applying FMM and then readily approximate $R_{i}=\exp(\ln(R_{i}))$ for all $i$.
+[24] proves numerical stability of the reduction to FMM of a similar although a little more involved rational expressions.
 (ii) Observe that
@@
 $R_{i}\approx\exp(\sigma_{i})\prod_{j\in\mathbb{J}^{\prime}_{i}}\Big{(}1-\frac{z_{j}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}\Big{)},~{}\sigma_{i}=\sum_{j\in\mathbb{J}_{i}}\frac{z_{j}-z_{j}^{\rm new}}{z_{i}^{\rm new}-z_{j}^{\rm new}},~{}i=1,\dots,d.$
-By applying FMM fast approximate $\sigma_{i}$ for all $i$ and compute the values $\prod_{j\in\mathbb{J}^{\prime}}\frac{z_{i}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}$ for all $i$ in at most $4\sum_{i=1}^{d}~{}|\mathbb{J}_{i}^{\prime}|-d$ arithmetic operations for $|\mathbb{J}^{\prime}_{i}|$ denoting the cardinality of the set $\mathbb{J}^{\prime}_{i}$.
+By applying FMM approximate $\sigma_{i}$ for all $i$ and compute the values $\prod_{j\in\mathbb{J}^{\prime}}\frac{z_{i}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}$ for all $i$ in at most $4\sum_{i=1}^{d}~{}|\mathbb{J}_{i}^{\prime}|-d$ arithmetic operations for $|\mathbb{J}^{\prime}_{i}|$ denoting the cardinality of the set $\mathbb{J}^{\prime}_{i}$.
 Compare the latter cost bound with order of $d^{2}$ arithmetic operations in the straightforward algorithm.
@@
 The same expression (3.3) defines Ehrlich’s iterations for both $f(x)$ and $p(x)$ by virtue of Thm.
-13.1.
-Newton’s iterations to a wild factor one additionally subtract the sum $\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}$ (cf.
+6.1.
+Applying Newton’s iterations to a wild factor one additionally subtract the sum $\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}$ (cf.
 (3.3)) and can apply FMM to accelerate numerical computation of this sum.
-To apply Weierstrass’s iterations to a wild factor one performs additional division by $t(x)$ and decreases the degree of the divisor $l^{\prime}(x)$ from $d-1$ to $w-1$.
+Applying Weierstrass’s iterations to a wild factor one performs additional division by $t(x)$ and decreases the degree of the divisor $l^{\prime}(x)$ from $d-1$ to $w-1$.
 24.4 MPSolve and its competition with Newton’s iterations
 Since its creation in 2000 the package MPSolve by Bini and Fiorentino [9], revised in 2014 by Bini and Robol [15], has been remaining user’s choice for approximation of all $d$ complex roots of a polynomial $p$.
-It implements Ehrlich’s iterations of 1967 [38], aka Aberth’s of 1973 [3], which empirically converge very fast globally – right from the start, under proper initialization, involving the coefficients of $p$.
-Such convergence has been consistently observed in decades-long worldwide applications of these iterations as well as in the case of some other root-finders by means of functional iterations, notably Weierstrass’s iterations [148], aka Durand and Kerner’s [31, 67]).
+It implements Ehrlich’s iterations, which empirically converge very fast globally, right from the start, under proper initialization involving the coefficients of $p$.
+Such convergence has been consistently observed in decades-long worldwide applications of these iterations as well as in the case of some other root-finders by means of functional iterations, notably Weierstrass’s iterations.
 Unsuccessful attempts of supporting this empirical behavior formally (e.g., see [3, 41]), however, have recently ended with surprising proofs that Weierstrass’s and Ehrlich’s iterations diverge in the case of inputs that lie on some complex manifolds, being an open domain in the case of Weierstrass’s iterations (see Reinke, Schleicher, and M.
@@
 An implementations of Newton’s iterations for approximation of all roots by Schleicher and R.
-Stoll in [143] competed with MPSolve in 2017, but then MPSolve has regained upper hand by means of incorporation of Fast Multipole Method (FMM), even though such implementation is simpler and can be more efficient for Newton’s than Ehrlich’s iterations (see the previous subsection).
-The competition seems to resume in 2024 (see [12, 130, 85]).
+Stoll in [115] competed with MPSolve in 2017, but then MPSolve has regained upper hand by means of incorporation of FMM.
+As we showed in the previous subsection, however, such implementation is simpler and can be more efficient for Newton’s than Ehrlich’s iterations.
+The Ehrlich-Newton competition seems to resume in 2024 (see [7, 102, 68]).
 24.5 Computation of a disc covering all zeros of a black box polynomial
+Customary initialization of Ehrlich’s and Weierstrass’s iterations involves approximation of all root radii, for which we can $d$ times apply our algorithm of Sec.
+16, greatly accelerated in the next section if $p(x)$ is given with its coefficients.
+In that case we immediately estimate the largest root radius $r_{1}$ to initialize Newton’s iterations of [115, 113, 102].
+In this subsection we obtain such estimates for a black box polynomial.
 24.5.1 Introductory comments
-Given coefficients of a polynomial $p$ of a degree $d$, we can readily compute its largest and smallest root radii $r_{1}$ and $r_{d}$, respectively (cf.
+Given the coefficients of a polynomial $p$ of a degree $d$, we can readily compute its largest and smallest root radii $r_{1}$ and $r_{d}$, respectively (cf.
 [31, 39]).
-Next we compute a range $[\rho_{-},\rho_{+}]$ that brackets all root radii or, equivalently, compute an upper estimate for the largest root radius $r_{1}$ and a low estimate for the smallest root radius $r_{d}$ of $p(x)$.
-We can apply the estimates of Sec.
-13.4 to the reverse polynomial $p_{\rm rev}(x)$ if its coefficients are available, but next we study this problem for a black box polynomial.
+Next we compute a range $[\rho_{-},\rho_{+}]$ that brackets all root radii or, equivalently, compute an upper estimate for the largest root radius $r_{1}$ and a low estimate for the smallest root radius $r_{d}$ of a black box polynomial $p(x)$ of a degree $d$.
 Its solution is required, e.g., in order to initialize algorithms for all $d$ roots by means of functional iterations.
 Suppose that we have an algorithm that computes a disc $D(0,\rho_{-})$ containing no zeros of a black box polynomial $p(x)$; clearly, $\rho_{-}\leq r_{d}$.
-By applying this algorithm to the reverse polynomial $p_{\rm rev}(x)$ we will obtain the upper bound $1/\rho_{-,{\rm rev}}$ on $r_{1}$.
+By applying this algorithm to the reverse polynomial $p_{\rm rev}(x)$ we obtain the upper bound $1/\rho_{-,{\rm rev}}$ on $r_{1}$.
 It remains to describe computation of the disc $D(0,\rho_{-})$. Empirically we can do this by applying the algorithms of [94, 36], but next we study formally some alternative algorithms.
@@
 We obtain a low bound on $r_{d}$ by computing a disc that contains no zeros of a black box polynomial $p(x)$.
-Hereafter let $p(x)$ denote a black box polynomial.
+In the rest of this section $p(x)$ denotes a black box polynomial.
 24.5.2 Probabilistic solution
@@
 24.5.3 Deterministic solution
-We can deterministically compute a desired disc $D(0,\rho_{-})$ at a higher cost, based on a soft e/i test.
-[…]
+We can deterministically compute a desired disc $D(0,\rho_{-})$ at a higher computational cost, based on a soft e/i test.
+Namely, apply 1-test to $d+1$ triples of parameters $c_{i}$, $\theta_{i}$, and $\rho_{i}$, for $i=0,1,\dots,d$, such that the discs $D(c_{i},\theta_{i}\rho_{i})$ have no overlap pairwise.
+Then, clearly, at most $d$ discs contain roots because $p(x)$ has only $d$ zeros, and so at least one of the $d+1$ discs contains no roots.
+Deterministic e/i tests of Secs.
+7.2 and 7.3 involve evaluation of $p(x)$ at $22d+$ 2 points.
+We obtain a desired disc $D$ by applying these tests to at most $d+1$ discs and thus computing $p(x)$ at at most $2(d+1)^{2}$ points.
 25 Fast approximation of the $d$ root radii with applications
-The customary initialization recipes of Weierstrass"s and Ehrlich’s iterations involve the $d$ root radii at the origin and consequently require to use the coefficients of $p$.
+The customary initialization recipes of Weierstrass’s and Ehrlich’s iterations involve the $d$ root radii at the origin and consequently require to use the coefficients of $p$.
 To initialize Ehrlich’s iterations MPSolve first applies Bini’s heuristic recipe of [6], reused in [9, 15], for the approximation of the $d$ root radii.
@@
 Then MPSolve initializes Ehrlich’s iterations at $d$ points $z_{j}$, $j=1,2,\dots,d$, exactly $m_{i}$ of them being equally spaced on the circle $C(c,(\rho_{i}+\rho_{i}^{\prime})/2)$.
-This heuristic initialization rec;ipe turns out to be highly efficient.
-The algorithm supporting the next theorem also approximates the $d$ root radii but is certified, deterministic, and faster than Bini’s.
-It was first outlined and briefly analyzed in [134] for fast approximation of a single root radius and then extended in [92, Sec.
+This heuristic initialization recipe turns out to be highly efficient.
+The algorithm supporting the next theorem also approximates the $d$ root radii.
+It appeared much earlier than Bini’s, is deterministic, formally supported, and faster than Bini’s.
+The algorithm has been devised and briefly analyzed in [106] for fast approximation of a single root radius and then readily extended in [74, Sec.
 4] (cf.
-also [93, Sec.
+also [75, Sec.
 5]) to approximation of all root radii of a polynomial.
@@
 Theorem 25.1.
-[62, Prop.
+[47, Prop.
 3].
@@
 [47] extends the supporting algorithm to approximation of all root radii within a positive relative error bound $\Delta$ by performing $\lceil\log_{2}\log_{1+\Delta}(2d)\rceil$ DLG root-squaring iterations (3.5).
-Combine Thm. 27.1 with estimates for the Boolean cost of performing these iterations implicit in the proof of [134, Cor. 14.3] and obtain
+Combine Thm. 25.1 with estimates for the Boolean cost of performing these iterations implicit in the proof of [106, Cor. 14.3] and obtain
 Corollary 25.1.
@@
 The algorithm supporting the corollary and said to be Alg.
-27.1a computes at a low cost $d$ suspect annuli $A(0,r_{j,-},r_{j,+})$ for $j=1,\dots,d$ (some of them may pairwise overlap), whose union $\mathbb{U}$ contains all $d$ roots.
-The algorithm was the point of departure for amazing progress in [59], which extends the novel ingenious algorithm of [81] for multipoint polynomial evaluation.
+25.1a computes at a low cost $d$ suspect annuli $A(0,r_{j,-},r_{j,+})$ for $j=1,\dots,d$ (some may overlap), whose union $\mathbb{U}$ contains all $d$ roots.
+The algorithm was the point of departure for amazing progress in [44], which extends the ingenious algorithm of [64] for multipoint polynomial evaluation.
 We conclude this subsection with a list of some other benefits of fast approximation of the $d$ root radii.
@@
 (ii) The search for real roots can be narrowed to the intersection $\mathbb{I}$ of $\mathbb{U}$ with the real axis, which consists of $2m$ segments of small length (see Fig.
-11).
-[126] has tested a simple real root-finder that approximates the roots in $\mathbb{I}$ by initiating Newton’s iterations at the centers of the segments.
-According to these tests the algorithm is highly efficient for the approximation of all isolated real roots.
+9).
+[98] has tested a real root-finder that approximates the roots in $\mathbb{I}$ by initiating Newton’s iterations at the centers of the segments.
+According to these tests the algorithm is highly efficient for the approximation of all well-isolated real roots.
 One can extend Newton’s iterations at the centers of the segments by applying also our root-radii algorithm of Sec.
-9.
+16.
 By complementing the root radii approach with some advanced novel techniques, the paper [47] has noticeably accelerated the real root-finder [54], which is currently user’s choice.
 (iii) Initialization based on Alg.
-27.1a should improve global convergence of Ehrlich’s, Weierstrass’s, and other functional iterations for root-finding as well as of Kirrinnis’s factorization algorithm of [68].
-[…]
+25.1a should improve global convergence of Ehrlich’s, Weierstrass’s, and other functional iterations for root-finding as well as of Kirrinnis’s factorization algorithm of [52].
+(iv) Suppose that Alg.
+25.1a has computed an annulus $A(0,\rho_{-},\rho_{+})$ containing a small number $u$ of the zeros of $p$ and isolated from its external zeros of $p$.
+Then we can extend the root-finding recipes of [106] from a disc to an annulus as follows.
+(a) Readily compute the first $u$ power sums of the zeros of $p$ lying in the discs $D(0,\rho_{+})$ and $D(0,\rho_{-})$ and then (b) their differences equal to the power sums of the zeros of $p$ lying in the annuls $A(0,\rho_{-},\rho_{+})$.
+(c) By solving the system of Newton’s identities, compute the coefficients of the factor $f$ of the $u$th degree that shares with $p$ its root set in the annulus (cf.
+[106, Sec.
+12]).
+(d) Approximate the zeros of $f$, at a lower cost since $u$ is small.
 We can apply this recipe to every annulus containing a small number of zeros of $p$ and isolated from its external zeros and finally approximate the remaining wild zeros of $p$ by applying implicit deflation (cf.